### v3

# Debiased Inference for Dynamic Nonlinear Panels with Multi-dimensional Heterogeneities

Abstract
We introduce a generic class of dynamic nonlinear heterogeneous parameter models that incorporate individual and time effects in both the intercept and slope. To address the incidental parameter problem inherent in this class of models, we develop an analytical bias correction procedure to construct a bias-corrected likelihood. The resulting maximum likelihood estimators are automatically bias-corrected. Moreover, likelihood-based tests statistics—including likelihood-ratio, Lagrange-multiplier, and Wald tests—follow the limiting chi-square distribution under the null hypothesis. Simulations demonstrate the effectiveness of the proposed correction method, and an empirical application on the labor force participation of single mothers underscores its practical importance.
Keywords:
incidental parameter problem; bias correction; fixed effects; maximum likelihood
JEL Classification:
C23
1 Introduction
Panel data models with two-way fixed effects are widely used in empirical research to control for unobserved heterogeneity that may correlate with covariates. Individual effects capture time-invariant, individual-specific characteristics, while time effects account for shifts that are common across individuals but vary over time. Traditionally, these models address level heterogeneity by incorporating fixed effects only in the intercept, adjusting for baseline outcome differences across individuals and time periods. However, many empirical applications reveal substantial variation in how outcomes respond to covariates (Browning and Carro, 2007). When such response heterogeneity are present and correlated with covariates, failing to account for them can lead to biased estimates and invalid inference (Su and Chen, 2013).
To illustrate, consider the classic problem of estimating the impact of children on a mother’s labor force participation. Research indicates that mothers with more children may differ systematically from those with fewer or none (Blau and Robins 1988; Angrist and Evans 1998; Kleven et al. 2019). First, mothers who choose to have more children might be inherently less inclined to work. Second, mothers with more children may be those whose labor supply is less sensitive to family size, perhaps due to better financial resources or access to childcare. In a panel dataset tracking households over time, we can control for the first type of level heterogeneity by including fixed effects in the intercept. However, to address the second type of response heterogeneity, we must include fixed effects in the slope. Given that labor force participation is a binary outcome and typically exhibits serial correlation, this example underscores the need for a dynamic nonlinear panel data model with fixed effects in both the intercept and slope.
In this paper, we introduce a general class of dynamic nonlinear heterogeneous parameter models (DN-HP) that incorporate individual and time fixed effects in both the intercept and slope. This flexible framework accommodates multi-dimensional heterogeneities and encompasses many commonly used panel data models. To address the incidental parameter problem inherent in DN-HP models, we propose a likelihood-based analytical bias correction approach. This method ensures the validity of the maximum likelihood estimator and likelihood-based inference for the structural parameters of DN-HP models. As a preview, Figure 1 presents simulated boxplots of the estimation bias in a DN-HP logit model, comparing maximum likelihood estimators from our bias-corrected likelihoods and the original likelihoods. The results reveal that the maximum likelihood estimators obtained from the bias-corrected likelihoods significantly reduce estimation bias while maintaining comparable mean square errors.
Figure 1: Comparison of Uncorrected and Corrected Maximum Likelihood Estimators
Notes: Model. $Y_{it}=\mathds{1}\{\rho Y_{it-1}+(\beta+\alpha_{1,i}+\gamma_{1,t})X_{it}+\alpha_{2,i}+\gamma_{2,t}+\varepsilon_{it}>0\}$ where $\mathds{1}\{\cdot\}$ denotes the indicator function, $\varepsilon_{it}$ is standard logistic independent of $X_{it},$ and the true values $\rho_{0}=\beta_{0}=0.5$. ($\widehat{\beta},\widehat{\rho})$ are the uncorrected estimators and $(\widehat{\beta}_{L},\widehat{\rho}_{L})$ are the corrected estimators. $1000$ replications. For details on the data-generating process, please refer to Section 4.
1.1 Literature Review
The estimation and inference of fixed effects models in the presence of the incidental parameter problem have been extensively studied. Early work, such as Chamberlain (1980), Arellano and Bond (1991), and Lancaster (2002), focused on frameworks with only individual effects in the intercept, addressing fixed-$T$ consistency (short panels) for structural parameter estimators in specific models. In recent decades, the availability of long panel datasets has motivated a large-$N,T$ framework, where the cross-sectional size $N$ and the time dimension $T$ grow proportionally to infinity. Under this rectangular asymptotic, estimators remain consistent but typically exhibit a non-negligible bias of order $O(1/T)$ due to the presence of individual effects in the intercept. To enable valid inference, researchers have developed various bias correction methods for a broad range of models in a maximum likelihood framework. For instance, Hahn and Newey (2004) and Hahn and Kuersteiner (2011) propose “parameter-based” bias correction methods to eliminate the bias from the maximum likelihood estimator. Two alternative approaches have also been explored: “score-based” and “likelihood-based” methods. Woutersen (2002) and Li et al. (2003) modify the profiled score to remove the bias, while Bester and Hansen (2009) and Arellano and Hahn (2016) identify and correct the bias directly for the likelihood. Most of these methods are analytical, estimating the bias using explicit formulas. Alternatively, numerical procedures are available, estimating the bias through methods such as the Jackknife (Dhaene and Jochmans, 2015), the bootstrap (Bellio et al., 2023), and integrated likelihoods (Arellano and Bonhomme, 2009).
For two-way fixed effects models, incorporating time effects in the intercept generates an additional bias of $O(1/N)$. Moon and Weidner (2017) propose a parameter-based analytical correction for dynamic linear models with two-way interactive effects. For nonlinear models, Fernández-Val and Weidner (2016) develop parameter-based bias correction techniques with additive fixed effects, while Chen et al. (2021) focus on interactive fixed effects in static models. Alternatively, Jochmans and Otsu (2019) propose a likelihood-based analytical bias correction for static models, allowing fixed effects to enter the model in an arbitrary but known form. Other model-specific solutions include Bai (2009) and Charbonneau (2017), among others. For an extensive review of the incidental parameter problem for models with two-way fixed effects in the intercept, see Fernández-Val and Weidner (2018).
To our knowledge, our paper represents one of the first efforts to extend intercept-only two-way fixed effects models into multi-dimensional two-way fixed effects models, accounting for both level and response heterogeneities. Two prior studies in this line of research are Keane and Neal (2020) and Lu and Su (2023). Both examine linear heterogeneous parameter models with additive effects in both the intercept and slope. Keane and Neal (2020) propose an iterative procedure, which they refer to as the “mean observation OLS,” to eliminate the bias, while Lu and Su (2023) rely on a parameter-based Jackknife approach and provide a uniform inference theory for structural parameters. Our study significantly expands the model framework and introduces a likelihood-based analytical bias correction technique to ensure valid estimation and inference. Neither our DN-HP models nor our bias correction procedure have been discussed in the current literature.
1.2 This Paper
We assume that the multi-dimensional fixed effects enter the log-likelihood additively. Identification restrictions are imposed on the fixed effects, whose dimension grows with the sample size. Our corrected likelihood is derived using two Taylor expansions: the expansion of the score function and the expansion of the original likelihood, both with respect to the fixed effects. Under appropriate dominance conditions, we explicitly derive the two additive bias terms in the original likelihoods of DN-HP models. These bias terms are expressed as quadratic forms involving the score and Hessian of the fixed effects. We estimate the derived bias terms under weak serial dependence (conditional strong mixing) and conditional cross-sectional independence.
We rigorously show that the maximum likelihood estimator (MLE) based on the bias-corrected likelihood of the DN-HP model retains a limiting normal distribution with zero mean (automatically bias-corrected), consistent with the limiting distribution of the classical MLE in the absence of incidental parameters. Beyond the bias-corrected MLE, this likelihood-based bias correction approach also facilitates hypothesis testing using the likelihood ratio (LR), Lagrange multiplier (LM), and Wald statistics. We demonstrate that these three test statistics, derived from the bias-corrected likelihood, are asymptotically equivalent, sharing the same asymptotic $\chi^{2}$ distribution under the null hypothesis.
We demonstrate the finite-sample performance of our method through Monte Carlo simulations, comparing it to both the original likelihood and the Jackknife. In the empirical application, we apply the DN-HP logit model to study the aforementioned question of how the number of children affects a mother’s labor force participation, with a focus on single-mother households. Our analysis reveals that this impact varies significantly across individuals and over time. Neglecting such variation can lead to biased estimates and markedly different conclusions, highlighting the need to address both level and response heterogeneities, as well as the value of our framework in empirical research.
The remainder of the paper is organized as follows. Section 2 describes our settings, the identification restriction, the incidental parameter problem under DN-HP, and gives an overview of our bias correction procedures. Section 3 provides the bias expansion of the likelihood and the formula for our corrected likelihood. The statistical properties of the corrected likelihood. We rigorously show that the estimator of the structural parameter, and the LR, LM, and Wald statistics all follow their standard asymptotic distribution. Section 4 presents simulation studies to demonstrate our method’s performance. Here we show that our correction procedure reduces the bias significantly, without increasing the root mean squared errors. In addition, we also show that our procedure restores the test size of the LR test to its nominal level. We also compare our procedure with the Jackknife and find that our procedure outperforms the Jackknife. Section 5 applies our method to study single mothers’ labor force participation. Proofs, technical details, and additional results are provided in the appendix.
Notations. In what follows, we denote $\iota_{n}$ to be an $n\times 1$ vector of ones and for $\mathbb{I}_{n}$ to be an $n\times n$ identity matrix. Let $\otimes$ be the kronecker product and $\mathds{1}\left\{\cdot\right\}$ the indicator function. In addition, for each square matrix $A_{i}$, $i=1,\ldots,n$, $\operatorname*{diag}\left\{A_{1},\ldots,A_{n}\right\}$ represent a block diagonal matrix, where each matrix $A_{i}$ is the $i$-th diagonal element.
2 Bias in Heterogeneous Parameter Model
2.1 DN-HP Model
Consider the responses $Y_{it},$ which follow the conditional density $f\left(Y_{it};\theta_{0}+\alpha_{i}^{0}+\gamma_{t}^{0}\right),$ given covariates $X_{it}$ (omitted for simplicity). Here, $i=1,\ldots,N$ represents individuals, and $t=1,\ldots,T$ represents time periods. The density is known up to $\theta_{0}$, $\alpha_{i}^{0}$, and $\gamma_{t}^{0},$ and the covariates $X_{it}$ may include lagged $Y_{it}$. Specifically, $\theta_{0}$ is the $K$-dimensional structural parameter of interest, where $K$ is fixed and known. The nuisance heterogeneous effects $\alpha_{i}^{0}=\left(\alpha_{1,i}^{0},\ldots,\alpha_{K,i}^{0}\right)^{\prime}$ and $\gamma_{t}^{0}=\left(\gamma_{1,t}^{0},\ldots,\gamma_{K,t}^{0}\right)^{\prime}$ may be correlated with $X_{it}$. Throughout the paper, we treat $\alpha_{i}^{0}$ and $\gamma_{t}^{0}$ as fixed effects (unknown parameters) by conditioning on them in the density.
Denote $\phi:=\left(\alpha_{1}^{\prime},\ldots,\alpha_{N}^{\prime},\gamma_{1}^{\prime},\ldots,\gamma_{T}^{\prime}\right)^{\prime}$ and define the quasi-likelihood function as
$l(\theta,\phi):=\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}l_{i,t}\left(\theta,\phi\right),\qquad l_{i,t}(\theta,\phi):=\log f\left(Y_{it};\theta+\alpha_{i}+\gamma_{t}\right).$
Note that if the data are independent across both $i$ and $t,$ $l(\theta,\phi)$ represents the exact likelihood. However, when observations for each unit $i$ exhibit serial correlation, $l(\theta,\phi)$ is misspecified even though the likelihood for each observation $Y_{it}\mid X_{it},$ $l_{i,t}(\theta,\phi),$ is correctly specified. Our asymptotic theory assumes conditional cross-sectional independence and weak time-serial dependence (Assumption 1). For simplicity, we refer to $l(\theta,\phi)$ as the likelihood function hereafter. Here is an example that illustrates how our model captures greater heterogeneity than conventional two-way effects panel data models.
Example 1 (heterogeneous parameter model)
Consider a probit model
$Y_{it}=\mathds{1}\left\{\pi_{it}\left(\theta,\alpha_{i},\gamma_{t}\right)+\varepsilon_{it}>0\right\}$
where the linear index $\pi_{it}(\cdot)$ varies across both individuals and time periods. The error term $\varepsilon_{it}$ follows a standard normal distribution denoted by $\Phi$. Without of loss of generality, let $X_{it}\in\mathbb{R},$ and the parameter vectors $\theta=\left(\theta_{1},\theta_{2}\right)^{\prime},$ $\alpha_{i}=\left(\alpha_{1,i},\alpha_{2,i}\right)^{\prime},$ and $\gamma_{t}=\left(\gamma_{1,t},\gamma_{2,t}\right)^{\prime}.$ We accommodate the linear index of the form:
$\pi_{it}\left(\theta,\alpha_{i},\gamma_{t}\right)=\left(\theta_{1}+\alpha_{1,i}+\gamma_{1,t}\right)+X_{it}\left(\theta_{2}+\alpha_{2,i}+\gamma_{2,t}\right).$ (2.1)
In the special case where $\alpha_{i}=\alpha_{1,i}$ and $\gamma_{t}=\gamma_{1,t},$ (2.1) simplifies to
$\pi_{it}\left(\theta,\alpha_{i},\gamma_{t}\right)=\left(\theta_{1}+\alpha_{1,i}+\gamma_{1,t}\right)+X_{it}\theta_{2},$
which corresponds to the traditional two-way fixed effects model. Assume that $\varepsilon_{it}$ is independent of $X_{it}.$ The likelihood for individual $i$ at time $t$ is given by
$l_{i,t}\left(\theta,\phi\right):=Y_{it}\log\Phi\left(\pi_{it}\left(\theta,\alpha_{i},\gamma_{t}\right)\right)+\left(1-Y_{it}\right)\log\Phi\left(-\pi_{it}\left(\theta,\alpha_{i},\gamma_{t}\right)\right).$
Due to the additive form of the parameters, the structural parameters, $\theta,$ is not identified in the likelihood. To resolve this, we reparameterize the likelihood by setting $\alpha_{N}=-\sum_{i=1}^{N-1}\alpha_{i}$ and $\gamma_{T}=-\sum_{t=1}^{T-1}\gamma_{t},$ thereby imposing the constraints $\sum_{i=1}^{N}\alpha_{i}=0$ and $\sum_{t=1}^{T}\gamma_{t}=0.$ Let $\psi:=\left(\alpha_{1}^{\prime},\ldots,\alpha_{N-1}^{\prime},\gamma_{1}^{\prime},\ldots,\gamma_{T-1}^{\prime}\right)^{\prime}.$ The reparameterized likelihood is then $l\left(\theta,D^{\prime}\psi\right)=:l^{*}\left(\theta,\psi\right)$ for $D:=\operatorname*{diag}\{D_{1},D_{2}\}$ where $D_{1}:=(\mathbb{I}_{N-1},-\iota_{N-1})\otimes\mathbb{I}_{K}$ and $D_{2}:=(\mathbb{I}_{T-1},-\iota_{T-1})\otimes\mathbb{I}_{K}.$
2.2 Incidental Parameter Problem
In this section, we illustrate the incidental parameter problem (IPP) that arises in our heterogeneous parameter model. For a given $\theta,$ define
$\psi(\theta):=\arg\max_{\psi\in\Lambda_{\theta}}\mathbb{E}[l^{*}(\theta,\psi)],\quad\widehat{\psi}(\theta):=\arg\max_{\psi\in\Lambda_{\theta}}l^{*}(\theta,\psi)$ (2.2)
where $\Lambda_{\theta}$ is a compact subset of $\mathbb{R}^{(N+T-2)K}$ for fixed $N$ and $T$ (Assumption 1) and $\mathbb{E}$ denotes expectations with respect to the true conditional density. Specifically, $\mathbb{E}$ represents the conditional expectation given the true values of fixed effects and the covariates. For notational simplicity, we use $\mathbb{E}$ when no confusion arises. Let the score and Hessian be
$s\left(\theta,\psi\right):=\partial_{\psi}l^{*}\left(\theta,\psi\right),\quad H\left(\theta,\psi\right):=\partial_{\psi\psi^{\prime}}l^{*}\left(\theta,\psi\right),$
where $\partial_{\psi}$ and $\partial_{\psi\psi^{\prime}}$ represent the first and second partial derivatives with respect to $\psi$. By (2.2), we know that $s(\theta,\widehat{\psi}\left(\theta\right))=0.$ Given $\theta,$ consider a Taylor expansion of $s(\theta,\widehat{\psi}\left(\theta\right))$ around $\psi(\theta)$:
$0=s\left(\theta,\psi\left(\theta\right)\right)+H\left(\theta,\psi(\theta)\right)\big{(}\widehat{\psi}(\theta)-\psi(\theta)\big{)}+r^{s}(\theta,a),$ (2.3)
where
$r^{s}\left(\theta,a\right)$ $:=\big{(}r_{\psi_{1}}^{s}\left(\theta,a_{1}\right),\ldots,r_{\psi_{(N+T-2)K}}^{s}(\theta,a_{(N+T-2)K})\big{)}^{\prime},$
$r_{\psi_{j}}^{s}\left(\theta,a_{j}\right)$ $:=\frac{1}{2}\big{(}\widehat{\psi}\left(\theta\right)-\psi(\theta)\big{)}^{\prime}G_{\psi_{j}}(\theta,a_{j})\big{(}\widehat{\psi}(\theta)-\psi(\theta)\big{)},$
$G_{\psi_{j}}(\theta,a_{j})$ $:=\partial_{\psi_{j}}H(\theta,a_{j}\widehat{\psi}(\theta)+(1-a_{j})\psi\left(\theta\right))$
with $a=(a_{1},\ldots,a_{(N+T-2)K})^{\prime}$ for each $a_{j}\in(0,1),$ and $\psi_{j}$ denotes the $j$th component of $\psi.$ Here $G_{\psi_{j}}(\theta,a_{j})$ is the derivative of each element of $H(\theta,\psi)$ with respect to $\psi_{j},$ evaluated at the point $a_{j}\widehat{\psi}\left(\theta\right)+\left(1-a_{j}\right)\psi\left(\theta\right).$ For notational simplicity, we write
$s(\theta):=s\left(\theta,\psi\left(\theta\right)\right)\quad\text{and}\quad H\left(\theta\right):=H\left(\theta,\psi(\theta)\right).$ (2.4)
From (2.3), by Assumption 1, we obtain that
$\widehat{\psi}(\theta)-\psi(\theta)$ $=[-{H}(\theta)]^{-1}\big{(}s(\theta)+r^{s}\left(\theta,a\right)\big{)}.$ (2.5)
Next, define the profile likelihoods as $\widehat{l}(\theta):=l^{*}(\theta,\psi)\mid_{\psi=\widehat{\psi}(\theta)}$ and $l(\theta):=l^{*}(\theta,\psi)\mid_{\psi={\psi}(\theta)}$. A Taylor expansion of $\widehat{l}(\theta)$ around $\psi(\theta)$ gives
$\widehat{l}(\theta)=l(\theta)+s^{\prime}\left(\theta\right)\big{(}\widehat{\psi}(\theta)-\psi(\theta)\big{)}+\frac{1}{2}\big{(}\widehat{\psi}(\theta)-\psi(\theta)\big{)}^{\prime}{H}(\theta)\big{(}\widehat{\psi}(\theta)-\psi(\theta)\big{)}+r^{l}\left(\theta,a_{0}\right)$ (2.6)
where $r^{l}\left(\theta,a_{0}\right):=1/3\big{(}\widehat{\psi}(\theta)-\psi(\theta)\big{)}^{\prime}r^{s}(\theta,a_{0}\,\iota_{(N+T-2)K})$ for some $a_{0}\in(0,1).$ Combining (2.5) and (2.6), we have
$\widehat{l}(\theta)=l(\theta)-\frac{1}{2}s^{\prime}(\theta)[{H}(\theta)]^{-1}s(\theta)+r(\theta)$ (2.7)
where $r(\theta):=1/2(r^{s}\left(\theta,a\right))^{\prime}[{H}(\theta)]^{-1}r^{s}\left(\theta,a\right)+r^{l}\left(\theta,a_{0}\right).$ Here we suppress $a$ and $a_{0}$ in $r(\theta)$ as they are constants. Under mild regularity conditions, we show that
$\sqrt{NT}\mathbb{E}\big{\{}s^{\prime}(\theta)[{H}(\theta)]^{-1}s(\theta)\big{\}}\nrightarrow 0$
and that $r(\theta)$ is a remainder term that is negligible for inference (see Section 3). This implies that the presence of $s^{\prime}(\theta)[{H}(\theta)]^{-1}s(\theta)$ in (2.7), due to the estimation of $\psi(\theta),$ contributes the MLE of $\theta_{0}$ based on $\widehat{l}(\theta).$ Consequently, the structural estimators and $\widehat{l}(\theta)$ based test statistics no longer follow the standard asymptotic distributions as in classical maximum likelihood theory. This is referred to as the IPP in this context. We use the following example to illustrate IPP.
Example 2 (heterogeneous means)
Assume that $Y_{it}$ follows a normal distribution with mean $\theta_{0}+\alpha_{i}^{0}+\gamma_{t}^{0}$ and unit variance, independent across both $i$ and $t.$ $\theta_{0}$ is the structural parameter of interest, while $\alpha_{i}^{0}$ and $\gamma_{t}^{0}$ are fixed effects. The likelihood function is
$l\left(\theta,\phi\right)=\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}-\frac{1}{2}\left(Y_{it}-\theta-\alpha_{i}-\gamma_{t}\right)^{2}-\frac{1}{2}\log(2\pi).$
Define $\widetilde{Y}_{it}:=Y_{it}-\mathbb{E}Y_{it}.$ Under the identification condition, the reparameterized likelihood $l^{\ast}(\theta,\psi)=l(\theta,D^{\prime}\psi).$ Given $\theta,$ maximizing $l^{\ast}(\theta,\psi)$ with respect to $\psi$ yields
$\widehat{\alpha}_{i}(\theta)-\alpha_{i}^{0}$ $=$ $\frac{1}{T}\sum_{t=1}^{T}\widetilde{Y}_{it}-\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}\widetilde{Y}_{it},\quad i=1,\ldots,N-1,$ (2.8)
$\widehat{\gamma}_{t}(\theta)-\gamma_{t}^{0}$ $=$ $\frac{1}{N}\sum_{i=1}^{N}\widetilde{Y}_{it}-\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}\widetilde{Y}_{it},\quad t=1,\ldots,T-1.$ (2.9)
Let $\big{(}\widehat{\alpha}_{1}(\theta),\ldots,\widehat{\alpha}_{N-1}(\theta),\widehat{\gamma}_{1}(\theta),\ldots,\widehat{\gamma}_{T-1}(\theta)\big{)}^{\prime}=:\widehat{\psi}(\theta)$ and $\big{(}{\alpha}_{1}^{0},\ldots,{\alpha}_{N-1}^{0},{\gamma}_{1}^{0},\ldots,{\gamma}_{T-1}^{0}\big{)}^{\prime}=:\psi_{0}.$ Define $\widehat{\alpha}_{N}(\theta):=-\sum_{i=1}^{N-1}\widehat{\alpha}_{i}(\theta)$ and $\widehat{\gamma}_{T}(\theta):=-\sum_{t=1}^{T-1}\widehat{\gamma}_{t}(\theta).$ Note that (2.8) and (2.9) also hold for $i=N$ and $t=T$ respectively. Denote $\phi_{0}:=({\alpha}_{1}^{0},\ldots,{\alpha}_{N-1}^{0},{\alpha}_{N}^{0},{\gamma}_{1}^{0},\ldots,{\gamma}_{T-1}^{0},{\gamma}_{T}^{0})^{\prime}$ and $\widehat{\phi}(\theta):=\big{(}\widehat{\alpha}_{1}(\theta),\ldots,\widehat{\alpha}_{N-1}(\theta),\widehat{\alpha}_{N}(\theta),\widehat{\gamma}_{1}(\theta),\ldots,\widehat{\gamma}_{T-1}(\theta),\widehat{\gamma}_{T}(\theta)\big{)}^{\prime}.$ Note that
$\widehat{l}(\theta)-l(\theta)$ $=$ $l(\theta,\widehat{\phi}(\theta))-l(\theta,\phi_{0})$
$=$ $\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}\big{[}Y_{it}(\theta)-\frac{1}{2}(\widehat{\alpha}_{i}(\theta)-\alpha_{i}^{0})\big{]}(\widehat{\alpha}_{i}(\theta)-\alpha_{i}^{0})$
$+\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}\big{[}Y_{it}(\theta)-\frac{1}{2}(\widehat{\gamma}_{t}(\theta)-\gamma_{t}^{0})\big{]}(\widehat{\gamma}_{t}(\theta)-\gamma_{t}^{0})$
where $Y_{it}(\theta):=Y_{it}-\theta-\alpha_{i}^{0}-\gamma_{t}^{0}.$ Since $Y_{it}$ is independent across $i$ and $t,$ by substituting (2.8) and (2.9), we obtain
$\mathbb{E}\big{[}\widehat{l}(\theta)-l(\theta)\big{]}$ $=\frac{1}{T}\cdot\frac{1}{2}\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}\mathbb{E}\big{[}\widetilde{Y}_{it}^{2}\big{]}+\frac{1}{N}\cdot\frac{1}{2}\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}\mathbb{E}\big{[}\widetilde{Y}_{it}^{2}\big{]}+o\big{(}1/\sqrt{NT}\big{)}$ (2.10)
$=\frac{1}{2}\frac{1}{T}+\frac{1}{2}\frac{1}{N}+o\big{(}1/\sqrt{NT}\big{)}.$
This leads to the conclusion that $\sqrt{NT}\mathbb{E}\big{[}\widehat{l}(\theta)-l(\theta)\big{]}\nrightarrow 0,$ revealing the IPP.
2.3 Bias Correction
The expansion in (2.10) from Example 2 motivates the correction of the likelihood $\widehat{l}(\theta)$ by subtracting the first two terms on the right-hand side of (2.10), ensuring that the corrected likelihood converges to $l(\theta)$ at a rate faster than $1/\sqrt{NT}.$ This paper proposes an analytical bias correction procedure by estimating the bias term $1/2\mathbb{E}\{s^{\prime}(\theta)[{H}(\theta)]^{-1}s(\theta)\}$ from (2.7) under $N/T\rightarrow\kappa\in(0,\infty)$. Specifically, let $s_{\breve{\alpha}}(\theta)$ represent the first $(N-1)K$ elements of $s(\theta),$ and $s_{\breve{\gamma}}(\theta)$ represent the last $(T-1)K$ elements of $s(\theta),$ where $\breve{\alpha}:=(\alpha_{1}^{\prime},\ldots,\alpha_{N-1}^{\prime})^{\prime}$ and $\breve{\gamma}:=(\gamma_{1}^{\prime},\ldots,\gamma_{T-1}^{\prime})^{\prime}.$ Accordingly, we partition $[\mathbb{E}H(\theta)]^{-1}$ as
$[\mathbb{E}H(\theta)]^{-1}=:\begin{pmatrix}H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta),&H_{\breve{\alpha}\breve{\gamma}^{\prime}}^{\ast}(\theta)\\ H_{\breve{\gamma}\breve{\alpha}^{\prime}}^{\ast}(\theta),&H_{\breve{\gamma}\breve{\gamma}^{\prime}}^{\ast}(\theta)\end{pmatrix},$
where $H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)$ and $H_{\breve{\gamma}\breve{\gamma}^{\prime}}^{\ast}\left(\theta\right)$ are $(N-1)K\times(N-1)K$ and $(T-1)K\times(T-1)K$ diagonal blocks, respectively, and $H_{\breve{\alpha}\breve{\gamma}^{\prime}}^{\ast}(\theta)=[H_{\breve{\gamma}\breve{\alpha}^{\prime}}^{\ast}(\theta)]^{\prime}.$ Lemma 1 (see Section 3) shows that $1/2\mathbb{E}\{s^{\prime}(\theta)[{H}(\theta)]^{-1}s(\theta)\}$ can be asymptotically decomposed into two bias components: $1/2\mathbb{E}\{s_{\breve{\alpha}}^{\prime}(\theta)[H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)]s_{\breve{\alpha}}(\theta)\}=:{B}_{\breve{\alpha}}(\theta)$ and $1/2\mathbb{E}\{s_{\breve{\gamma}}^{\prime}(\theta)[H_{\breve{\gamma}\breve{\gamma}^{\prime}}^{\ast}(\theta)]s_{\breve{\gamma}}(\theta)\}=:{B}_{\breve{\gamma}}(\theta),$ resulting from the estimation of the individual and time effects respectively. By estimating the two bias components, we construct the corrected likelihood as follows:
$L(\theta):=\widehat{l}(\theta)+\widehat{B}_{\breve{\alpha}}(\theta)+\widehat{B}_{\breve{\gamma}}(\theta).$ (2.11)
The explicit forms of $\widehat{B}_{\breve{\alpha}}(\theta)$ and $\widehat{B}_{\breve{\gamma}}(\theta)$ are defined in Section 3. Subsequently, Theorem 2 demonstrates that
$L(\theta)-l(\theta)=o_{\mathbb{P}}\big{(}1/\sqrt{NT}\big{)},$
implying that the corrected likelihood no longer suffers from the IPP. Define the MLE of $\theta_{0}$ as the maximizer of $L(\theta):$
$\widehat{\theta}_{L}=\arg\max_{\theta\in\Theta}L(\theta),$ (2.12)
where $\Theta$ is a compact subset of $\mathbb{R}^{K}.$ Under mild regularity conditions, this MLE is asymptotically unbiased, and its asymptotic properties align with the standard results for classical MLE in the independent and identically distributed (i.i.d.) case. Moreover, the test statistics constructed from the corrected likelihood such as the Wald, LR, and LM also follow the limiting $\chi^{2}$ distribution under the null hypothesis.
Since the corrected likelihood depends implicitly on $\widehat{\psi}(\theta),$ which maximizes the reparameterized likelihood function $l^{\ast}(\theta,\psi),$ obtaining $\widehat{\theta}_{L}$ from $L(\theta)$ involves a two-layer optimization. In the inner layer, $\widehat{\psi}(\theta)$ is found by maximizing $l^{\ast}(\theta,\psi)$ using the Newton-Raphson algorithm with analytical derivatives. In the outer layer, we maximize $L(\theta)$ with respect to $\theta,$ substituting $\widehat{\psi}(\theta)$ and using numerical derivatives.
3 Asymptotic Theory
In this section, we develop the asymptotic theory for our bias correction procedure, as well as for the estimation and testing based on the corrected likelihood function. Let $\phi(\theta):=D^{\prime}\psi(\theta)$ and $\widehat{\phi}(\theta):=D^{\prime}\widehat{\psi}(\theta).$ The following assumptions are introduced.
Assumption 1
  • (i) Suppose $N/T\rightarrow\kappa$ for some $0<\kappa<\infty$ as $N$ and $T\to\infty.$
  • (ii) Let $\Theta$ be a compact subset of $\mathbb{R}^{K}$ with $\theta_{0}$ in its interior. For each $\theta\in\Theta,$ there exists a compact subset $\Lambda_{\theta}$ that contains both $\widehat{\phi}(\theta)$ and $\phi(\theta)$ in its interior, such that $l(\theta,\phi)$ is three-times continuously differentiable with respect to $\phi\in\Lambda_{\theta}$ almost surely. Let $w_{it}=(Y_{it},X_{it}^{\prime})^{\prime},$ and assume there exists a function $g(w_{it})>0$ independent of $\theta$ and $\phi,$ such that
    $\sup_{\theta\in\Theta}\sup_{\phi\in\Lambda_{\theta}}\left|\partial_{\phi_{1}\ldots\phi_{S}}l_{i,t}(\theta,\phi)\right|<g(w_{it}),$
    where $\phi_{s}$ represents any element of $\phi$ and $S\in\{0,1,2,3\}$ ($S=0$ understood as no differentiation), and $\sup_{N,T}\max_{1\leq i\leq N,1\leq t\leq T}\mathbb{E}\left[g^{\eta}(w_{it})\right]<\infty,$ for some $\eta>2.$
  • (iii) For each $k=1,\ldots,K$, $\sum_{i=1}^{N}\alpha_{ik}=\sum_{t=1}^{T}\gamma_{tk}=0$.
  • (iv) Given any fixed $\theta\in\Theta,$ for all $\phi\in\Lambda_{\theta}$ if $\phi\neq\phi(\theta),$ $\mathbb{P}\big{(}l(\theta,\phi)\neq l(\theta,\phi(\theta))\big{)}>0.$
  • (v) Conditional on $\phi(\theta),$$\left\{w_{it}\right\}$ are independent across $i$. Denote $\mathcal{A}_{it}$ and $\mathcal{B}_{it}$ to be the $\sigma$-algebras generated, respectively, by $\left\{w_{is}:1\leq s\leq t\right\}$ and $\left\{w_{is}:t\leq s\leq T\right\}$. $\left\{w_{it}\right\}$ is conditional strong mixing with mixing coefficient
    $a_{i}\left(m\right):=\sup_{\theta\in\Theta}\sup_{t\geq 1}\sup_{A\in\mathcal{A}_{it},B\in\mathcal{B}_{it+m}}\left|\mathbb{P}\left(A\cap B\mid\phi(\theta)\right)-\mathbb{P}\left(A\mid\phi(\theta)\right)\mathbb{P}\left(B\mid\phi(\theta)\right)\right|,$
    and, $\sup_{N}\max_{1\leq i\leq N}\sum_{m=0}^{\infty}\left[a_{i}\left(m\right)\right]^{1-2/\eta}<\infty,$ for some $\eta>2.$
  • (vi) For $\theta\in\Theta,$ $\sqrt{NT}\,\mathbb{E}\left[{H}(\theta)\right]$ is nonsingular when $N$ and $T$ are large.
Assumption 1 requires that both $N$ and $T$ are large and proportional, which is standardis standard in the literature. Assumption 1 imposes smoothness and boundedness conditions on the likelihood. Specifically, it requires that the likelihood and its derivatives for unit $i$ at time $t$ are unformly bounded by a measurable function whose moments of order greater than 2 exist. Assumption 1 is the identification condition for the incidental parameters (fixed effects) as they enter the likelihood additively. Given Assumption 1, Assumption 1 ensures that for each fixed $\theta,$ $\psi(\theta)$ defined in (2.2) is the unique maximizer. Assumption 1 assumes that the data are independent across units $i$ and conditional strong mixing across time $t.$ Under Assumptions 1 and 1, we can apply Davydov’s inequality to control the convergence rate of certain expectations of summations. Assumption 1 requires that the identified incidental parameter Hessian is invertible.
The following lemma states the bias induced by IPP, as shown in (2.7).
Lemma 1
Under Assumption 1, we have
$\mathbb{E}\left\{s^{\prime}(\theta)[{H}\left(\theta\right)]^{-1}s(\theta)\right\}=\mathbb{E}\left\{s_{\breve{\alpha}}^{\prime}(\theta)H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)s_{\breve{\alpha}}(\theta)\right\}+\mathbb{E}\left\{s_{\breve{\gamma}}^{\prime}(\theta)H_{\breve{\gamma}\breve{\gamma}^{\prime}}^{\ast}(\theta)s_{\breve{\gamma}}(\theta)\right\}+o\big{(}1/\sqrt{NT}\big{)},$
where $o(\cdot)$ is uniform in $\theta,$ $s_{\breve{\alpha}}(\theta)$ represent the first $(N-1)K$ elements of $s(\theta),$ and $s_{\breve{\gamma}}(\theta)$ represent the last $(T-1)K$ elements of $s(\theta).$
The following theorem presents the bias terms of $\mathbb{E}[\widehat{l}(\theta)]$ to approximate $\mathbb{E}[l(\theta)].$
Theorem 1
Under Assumption 1, as $N$ and $T\rightarrow\infty,$ we have
$\mathbb{E}[l(\theta)]=\mathbb{E}[\widehat{l}(\theta)]+B_{\breve{\alpha}}(\theta)+B_{\breve{\gamma}}(\theta)+o\big{(}1/\sqrt{NT}\big{)},$
where $o(\cdot)$ is uniform over $\theta\in\Theta.$ The bias terms $B_{\breve{\alpha}}(\theta)$ and $B_{\breve{\gamma}}(\theta)$ are defined as follows:
$B_{\breve{\alpha}}(\theta):=\frac{1}{2}\mathbb{E}\big{[}{s}_{\breve{\alpha}}^{\prime}(\theta)H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta){s}_{\breve{\alpha}}(\theta)\big{]},\quad B_{\breve{\gamma}}(\theta):=\frac{1}{2}\mathbb{E}\big{[}{s}_{\breve{\gamma}}^{\prime}(\theta){H}_{\breve{\gamma}\breve{\gamma}^{\prime}}^{\ast}(\theta){s}_{\breve{\gamma}}(\theta)\big{]}.$
Since $B_{\breve{\alpha}}(\theta)=O\big{(}1/\sqrt{NT}\big{)}$ and $B_{\breve{\gamma}}(\theta)=O\big{(}1/\sqrt{NT}\big{)}$ as $N$ and $T\to\infty,$ these bias terms introduce a non-zero limiting center to the MLE based on $\widehat{l}(\theta).$ Therefore, it is necessary to eliminate the effect of this bias. In the following, we estimate these bias terms and add them back to $\widehat{l}(\theta).$ Let $\psi(\theta)=:(\alpha^{\prime}_{1}(\theta),\ldots,\alpha^{\prime}_{N-1}(\theta),\gamma^{\prime}_{1}(\theta),\ldots,\gamma^{\prime}_{T-1}(\theta))^{\prime},$ $-\sum_{i=1}^{N-1}\alpha_{i}(\theta)=:\alpha_{N}(\theta),$ and $-\sum_{t=1}^{T-1}\gamma_{t}(\theta)=:\gamma_{T}(\theta).$ Then, $\phi(\theta)=(\alpha^{\prime}_{1}(\theta),\ldots,\alpha^{\prime}_{N}(\theta),\gamma^{\prime}_{1}(\theta),\ldots,\gamma^{\prime}_{T}(\theta))^{\prime}.$ Noting that $\mathbb{E}[s_{\breve{\alpha}}(\theta)]=0$ and $\mathbb{E}[s_{\breve{\gamma}}(\theta)]=0,$ we have
$B_{\breve{\alpha}}(\theta)$ $=\frac{1}{2}\mathbb{E}\big{\{}[s_{\breve{\alpha}}(\theta)-\mathbb{E}s_{\breve{\alpha}}(\theta)]^{\prime}H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}\left(\theta\right)[s_{\breve{\alpha}}(\theta)-\mathbb{E}s_{\breve{\alpha}}(\theta)]\big{\}}$
$=\frac{1}{2}\text{{trace}}\big{\{}H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)\Sigma^{\breve{\alpha}}(\theta)\big{\}},$ (3.1)
where
$\Sigma^{\breve{\alpha}}(\theta)$ $:=\Sigma_{1}^{\breve{\alpha}}(\theta)+\Sigma_{2}^{\breve{\alpha}}(\theta),$
$\Sigma_{1}^{\breve{\alpha}}(\theta)$ $:=\text{{diag}}\big{\{}\mathbb{E}\big{[}\widetilde{s}_{1}^{\alpha}\left(\theta\right)\left(\widetilde{s}_{1}^{\alpha}\left(\theta\right)\right)^{\prime}\big{]},\ldots,\mathbb{E}\big{[}\widetilde{s}_{N-1}^{\alpha}\left(\theta\right)\left(\widetilde{s}_{N-1}^{\alpha}\left(\theta\right)\right)^{\prime}\big{]}\big{\}},$
$\Sigma_{2}^{\breve{\alpha}}(\theta)$ $:=(\iota_{N-1}\iota_{N-1}^{\prime})\otimes\mathbb{E}\big{[}\widetilde{s}_{N}^{\alpha}\left(\theta\right)\left(\widetilde{s}_{N}^{\alpha}\left(\theta\right)\right)^{\prime}\big{]},$
$\widetilde{s}_{i}^{\alpha}\left(\theta\right)$ :=$:=\big{\{}\partial_{\alpha_{i}}l\left(\theta,\phi\right)-\mathbb{E}[\partial_{\alpha_{i}}l\left(\theta,\phi\right)]\big{\}}\big{|}_{\phi=\phi\left(\theta\right)},\quad\text{$i=1,\ldots,N.$}$
Similarly, we have
$B_{\breve{\gamma}}(\theta)$ $=\frac{1}{2}\mathbb{E}\big{\{}[s_{\breve{\gamma}}(\theta)-\mathbb{E}s_{\breve{\gamma}}(\theta)]^{\prime}H_{\breve{\gamma}\breve{\gamma}^{\prime}}^{\ast}\left(\theta\right)[s_{\breve{\gamma}}(\theta)-\mathbb{E}s_{\breve{\gamma}}(\theta)]\big{\}}$
$=\frac{1}{2}\text{{trace}}\big{\{}H_{\breve{\gamma}\breve{\gamma}^{\prime}}^{\ast}(\theta)\Sigma^{\breve{\gamma}}(\theta)\big{\}},$ (3.2)
where
$\Sigma^{\breve{\gamma}}(\theta)$ $:=\mathbb{E}\big{\{}[s_{\breve{\gamma}}(\theta)-\mathbb{E}s_{\breve{\gamma}}(\theta)][s_{\breve{\gamma}}(\theta)-\mathbb{E}s_{\breve{\gamma}}(\theta)]^{\prime}\big{\}}=\mathbb{E}\big{[}\widetilde{S}_{\breve{\gamma}}(\theta)(\widetilde{S}_{\breve{\gamma}}(\theta))^{\prime}\big{]}$
$\widetilde{S}_{\breve{\gamma}}(\theta)$ $:=\big{(}(\widetilde{s}_{1}^{\gamma}(\theta))^{\prime},\ldots,(\widetilde{s}_{T-1}^{\gamma}(\theta))^{\prime}\big{)}^{\prime}-\iota_{T-1}\otimes\widetilde{s}_{T}^{\gamma}(\theta)$
$\widetilde{s}_{t}^{\gamma}(\theta)$ :=$:=\big{\{}\partial_{\gamma_{t}}l\left(\theta,\phi\right)-\mathbb{E}[\partial_{\gamma_{t}}l\left(\theta,\phi\right)]\big{\}}\big{|}_{\phi=\phi(\theta)}.$
Next, we estimate the two bias terms. We first estimate the Hessian matrix $H(\theta)=H(\theta,\psi(\theta))$ by substituting $\widehat{\psi}(\theta).$ Define $\widehat{H}(\theta):=H(\theta,\widehat{\psi}(\theta)).$ We have $\widehat{H}(\theta)=D\big{[}\partial_{\phi\phi^{\prime}}l(\theta,\phi)\mid_{\phi=\widehat{\phi}(\theta)}\big{]}D^{\prime}$ with $\widehat{\phi}(\theta)=D^{\prime}\widehat{\psi}(\theta).$ Following (3) and (3), we estimate ${B}_{\breve{\alpha}}(\theta)$ and ${B}_{\breve{\gamma}}(\theta)$ by
$\widehat{B}_{\breve{\alpha}}(\theta):=\frac{1}{2}\operatorname*{trace}\left(\widehat{H}_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}\left(\theta\right)\widehat{\Sigma}_{\tau}^{\breve{\alpha}}(\theta)\right)\quad\text{and}\quad\widehat{B}_{\breve{\gamma}}(\theta):=\frac{1}{2}\operatorname*{trace}\left(\widehat{H}_{\breve{\gamma}\breve{\gamma}^{\prime}}^{\ast}(\theta)\widehat{\Sigma}_{\nu}^{\breve{\gamma}}(\theta)\right),$
where $\widehat{H}_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)$ and $\widehat{H}_{\breve{\gamma}\breve{\gamma}^{\prime}}^{\ast}(\theta)$ are the first $(N-1)K\times(N-1)K$ and last $(T-1)K\times(T-1)K$ blocks of
$[\widehat{H}(\theta)]^{-1}=:\begin{pmatrix}\widehat{H}_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}\left(\theta\right),&\widehat{H}_{\breve{\alpha}\breve{\gamma}^{\prime}}^{\ast}(\theta)\\ \widehat{H}_{\breve{\gamma}\breve{\alpha}^{\prime}}^{\ast}\left(\theta\right),&\widehat{H}_{\breve{\gamma}\breve{\gamma}^{\prime}}^{\ast}(\theta)\end{pmatrix},$
and $\widehat{\Sigma}_{\tau}^{\breve{\alpha}}(\theta)$ and $\widehat{\Sigma}_{\nu}^{\breve{\gamma}}(\theta)$ are the estimators of $\Sigma^{\breve{\alpha}}(\theta)$ and $\Sigma^{\breve{\gamma}}(\theta),$ defined as follows. For some $\tau\rightarrow\infty$ such that $\tau/T\rightarrow 0$,
$\widehat{\Sigma}_{\tau}^{\breve{\alpha}}(\theta)$ $:=\widehat{\Sigma}_{1,\tau}^{\breve{\alpha}}(\theta)+\widehat{\Sigma}_{2,\tau}^{\breve{\alpha}}(\theta),$
$\widehat{\Sigma}_{1,\tau}^{\breve{\alpha}}(\theta)$ $:=\frac{1}{N^{2}T^{2}}\sum_{t=1}^{T}\sum_{s=1}^{T}\mathds{1}\left\{\left|t-s\right|\leq\tau\right\}\operatorname*{diag}\big{\{}\widehat{s}_{1,t}^{\alpha}(\theta)\big{(}\widehat{s}_{1,s}^{\alpha}(\theta)\big{)}^{\prime},\ldots,\widehat{s}_{N-1,t}^{\alpha}(\theta)\big{(}\widehat{s}_{N-1,s}^{\alpha}(\theta)\big{)}^{\prime}\big{\}},$
$\widehat{\Sigma}_{2,\tau}^{\breve{\alpha}}(\theta)$ $:=\left(\iota_{N-1}\iota_{N-1}^{\prime}\right)\otimes\frac{1}{N^{2}T^{2}}\sum_{t=1}^{T}\sum_{s=1}^{T}\mathds{1}\left\{\left|t-s\right|\leq\tau\right\}\widehat{s}_{N,t}^{\alpha}(\theta)\big{(}\widehat{s}_{N,s}^{\alpha}(\theta)\big{)}^{\prime},$
$\widehat{s}_{i,t}^{\alpha}(\theta)$ $:=\bigg{[}\partial_{\alpha_{i}}l_{i,t}(\theta,\phi)-\frac{1}{T}\sum_{t=1}^{T}\partial_{\alpha_{i}}l_{i,t}(\theta,\phi)\bigg{]}\bigg{|}_{\phi=\widehat{\phi}(\theta)}.$
Here, $\tau$ is a truncation parameter commonly used in the literature (see, e.g., Hahn and Kuersteiner (2011)). When the data exhibit serial dependence, we set $\tau=1$ or $\tau=2$ as in Fernández-Val and Weidner (2016). Our simulations show that the results for $\tau=1$ and $\tau=2$ are similar. For data without serial dependence, we set $\tau=0.$ Additionally, we define
$\widehat{\Sigma}^{\breve{\gamma}}_{\nu}(\theta)$ $:=\frac{1}{N^{2}T^{2}}\sum_{i=1}^{N}\sum_{j=1}^{N}\mathds{1}\left\{\left|i-j\right|\leq\upsilon\right\}\widehat{S}^{\breve{\gamma}}_{i}(\theta)\big{(}\widehat{S}^{\breve{\gamma}}_{j}(\theta)\big{)}^{\prime}$
$\widehat{S}^{\breve{\gamma}}_{i}(\theta)$ $:=\big{(}(\widehat{s}_{i,1}^{\gamma}(\theta))^{\prime},\ldots,(\widehat{s}_{i,T-1}^{\gamma}(\theta))^{\prime}\big{)}^{\prime}-\iota_{T-1}\otimes\widehat{s}_{i,T}^{\gamma}(\theta)$
$\widehat{s}_{i,t}^{\gamma}(\theta)$ $:=\bigg{[}\partial_{\gamma_{t}}l_{i,t}\left(\theta,\phi\right)-\frac{1}{N}\sum_{i=1}^{N}\partial_{\gamma_{t}}l_{i,t}\left(\theta,\phi\right)\bigg{]}\bigg{|}_{\phi=\widehat{\phi}(\theta)},\quad i=1,\ldots,N,\ t=1,\ldots,T.$
Specifically, when we set $\upsilon=0$ given conditional cross-sectional independence, $\widehat{\Sigma}^{\breve{\gamma}}_{\nu}(\theta)$ simplifies to
$\widehat{\Sigma}^{\breve{\gamma}}_{0}(\theta)=\frac{1}{N^{2}T^{2}}\sum_{i=1}^{N}\widehat{S}^{\breve{\gamma}}_{i}(\theta)\big{(}\widehat{S}^{\breve{\gamma}}_{i}(\theta)\big{)}^{\prime}.$
Thus, using the estimated bias, we obtain the corrected likelihood $L(\theta):$
$L(\theta)=\widehat{l}(\theta)+\widehat{B}_{\breve{\alpha}}(\theta)+\widehat{B}_{\breve{\gamma}}(\theta)$
as defined in (2.11).
The following theorem demonstrates that the difference between the corrected likelihood $L(\theta)$ and and the infeasible likelihood $l(\theta)$ is of order $o_{\mathbb{P}}\big{(}1/\sqrt{NT}\big{)},$ implying that $L(\theta)$ no longer suffers from IPP.
Theorem 2
Under Assumption 1, the corrected likelihood satisfies
$L(\theta)=l(\theta)+o_{\mathbb{P}}\big{(}1/\sqrt{NT}\big{)},$
as $N$ and $T\to\infty.$
Compared to $\widehat{l}(\theta),$ the difference between the corrected likelihood and the infeasible likelihood is negligible. The corrected likelihood is derived from the expansion (2.7). Specifically, (2.7) is based on two Taylor expansions: the expansion of the score function for incidental parameters in (2.3) and the expansion of the original likelihood in (2.6). It is seen that the bias terms originate from the quadratic form on the right-hand side of (2.7), expressed in terms of the score and Hessian of the incidental parameters. Under Assumption 1, the quadratic form can be further decomposed into two additive sub-quadratic forms (Lemma 1), capturing the bias contributions from the individual and time dimensions, respectively. Consequently, our bias-corrected likelihood represents the original likelihood augmented by the estimation of these two additive bias terms.
Last, we demonstrate that the MLE obtained from the corrected likelihood, $\widehat{\theta}_{L}=\arg\max_{\theta}L(\theta),$ follows an limiting normal distribution with zero mean. Additional regularity conditions are required.
Assumption 2
  • (i) The profile likelihood function $l(\theta)$ is twice differentiable with respect to $\theta\in\Theta,$ where $\Theta$ is a compact subset of $\mathbb{R}^{K}$ with $\theta_{0}$ in its interior.
  • (ii) For all $\theta\in\Theta$ and $\theta\neq\theta_{0},$ $\mathbb{P}\big{(}l(\theta)\neq l(\theta_{0})\big{)}>0.$
  • (iii) For large $N$ and $T,$
    $\mathbb{E}\big{[}\sup_{\theta\in\Theta}|l(\theta)|\big{]}<\infty.$
  • (iv) For large $N$ and $T,$$\{\|\triangledown_{\theta}l(\theta)\|,\theta\in\Theta\}$ is uniformly integrable, and
    $\mathbb{E}\big{[}\sup_{\theta\in\Theta}\|\triangledown_{\theta\theta^{\prime}}l(\theta)\|\big{]}<\infty$
    where $\triangledown_{\theta}$ and $\triangledown_{\theta\theta^{\prime}}$ denote the first and second total derivatives with respect to $\theta.$ Here, $\|\triangledown_{\theta\theta^{\prime}}l(\theta)\|$ defines the Euclidean norm of $\triangledown_{\theta\theta^{\prime}}l(\theta).$
  • (v) The Hessian matrix $\mathbb{E}[\triangledown_{\theta\theta^{\prime}}l(\theta_{0})]$ is nonsingular for large $N$ and $T$.
Assumption 2 requires the differentiability likelihood on a compact parameter space of $\theta.$ Assumption 2 serves as an identification condition for $\theta$ ensuring that $\mathbb{E}[l(\theta)]$ is uniquely maximized at $\theta_{0}.$ Assumption 2 shows a dominance condition implying that $1/(NT)\sum_{i=1}^{N}\sum_{t=1}^{T}l_{i,t}(\theta,\phi(\theta))$ converges uniformly in probability to $\mathbb{E}[l(\theta)]$ over $\theta.$ Consequently, Assumptions 2-2 guarantee that $\widehat{\theta}_{L}$ is an asymptotically consistent estimator of $\theta_{0}.$ Assumption 2 establishes a dominance condition for the Hessian, so that for any consistent estimator $\widetilde{\theta}$ of $\theta_{0},$
$1/(NT)\sum_{i=1}^{N}\sum_{t=1}^{T}l_{i,t}(\widetilde{\theta},\phi(\widetilde{\theta}))\stackrel{{\scriptstyle\mathbb{P}}}{{\longrightarrow}}\mathbb{E}[\triangledown_{\theta\theta^{\prime}}l(\theta_{0})].$
Assumption 2 ensures that the Hessian for $\theta$ is invertible.
Corollary 1
Under Assumptions 1 and 2, we have
$\sqrt{NT}(\widehat{\theta}_{L}-\theta_{0})\stackrel{{\scriptstyle\mathbb{D}}}{{\longrightarrow}}\mathcal{N}(0,\Sigma)$
as $N$ and $T\to\infty,$ where $\Sigma=-\lim_{N,T\to\infty}\big{(}\mathbb{E}[\triangledown_{\theta\theta^{\prime}}l(\theta_{0})]\big{)}^{-1},$ and $\mathcal{N}(0,\Sigma)$ denotes the normal distribution with mean 0 and variance $\Sigma.$
We consider a generic nonlinear null hypothesis $H_{0}:R(\theta_{0})=0,$ where $R(\theta)$ is an $r\times 1$ vector-valued function independent of the data for some $r\leq K.$ Assume that $R(\theta)$ is continuously differentiable, with its Jacobian denoted as $J(\theta):=\triangledown_{\theta^{\prime}}R(\theta),$ and that $\operatorname*{rank}[J(\theta)]=r.$ Let $\widehat{\theta}_{R}:=\arg\max_{\theta}L(\theta)$ subject to $R(\theta)=0.$ Define the test statistics:
$\widehat{\xi}_{LR}:=-2NT\big{\{}L(\widehat{\theta}_{R})-L(\widehat{\theta}_{L})\big{\}},$
$\widehat{\xi}_{LM}:=-NT\big{\{}\triangledown_{\theta^{\prime}}L(\widehat{\theta}_{R})[\triangledown_{\theta\theta^{\prime}}L(\widehat{\theta}_{R})]^{-1}\triangledown_{\theta}L(\widehat{\theta}_{R})\big{\}},$
$\widehat{\xi}_{Wald}:=-NT\big{\{}R(\widehat{\theta}_{L})^{\prime}\big{(}J(\widehat{\theta}_{L})[\triangledown_{\theta\theta^{\prime}}L(\widehat{\theta}_{L})]^{-1}J(\widehat{\theta}_{L})^{\prime}\big{)}^{-1}R(\widehat{\theta}_{L})\big{\}}.$
The following corollary states that the hypothesis testing based on the corrected likelihood aligns with standard results in the context of MLE.
Corollary 2
Suppose Assumptions 1 and 2. Under the null hypothesis $H_{0}:R(\theta_{0})=0,$ the statistics $\widehat{\xi}_{LR},$ $\widehat{\xi}_{LM},$ and $\widehat{\xi}_{Wald}$ are asymptotically equivalent, following a common limiting $\chi^{2}(r)$ distribution with degrees of freedom $r$.
4 Simulations
In the simulation study, we consider data generating process (DGP) from a dynamic logit model as follows:
$Y_{it}=\mathds{1}\{\rho Y_{it-1}+(\beta+\alpha_{1,i}+\gamma_{1,t})X_{it}+\alpha_{2,i}+\gamma_{2,t}+\varepsilon_{it}>0\}$ (4.1)
where the disturbance term $\varepsilon_{it}$ follows a standard logistic distribution, independent across both $i$ and $t,$ and also independent of the regressor $X_{it}.$ According to the identification restriction, we generate true values for the incidental parameters $\{\alpha_{1,i}^{0},\gamma_{1,t}^{0}\}$ from $\mathcal{N}(0,0.01)$, after demeaning. $\{\alpha_{2,i}^{0},\gamma_{2,t}^{0}\}$ are also generated from $\mathcal{N}(0,0.01)$ where we only demean $\gamma_{2,t}^{0}$. Given these parameters, $X_{it}$ is independently generated from $\mathcal{N}\big{(}\sum_{k=1}^{2}(\alpha_{k,i}^{0}+\gamma_{k,t}^{0}),0.25\big{)}$ across both $i$ and $t.$ The true structural parameter is set as $\theta_{0}=(\rho_{0},\beta_{0})^{\prime}=(0.5,0.5)^{\prime}.$ The initial $Y_{i0}=\mathds{1}\{(\beta_{0}+\alpha_{1,i}^{0})X_{i0}+\alpha_{2,i}^{0}+\varepsilon_{i0}>0\}$ with $X_{i0}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0},0.25)$ and $\varepsilon_{i0}$ following a standard logistic distribution. Model (4.1) also encompasses the static logit model as a specific case where $\rho\equiv 0.$ We simulate this static case with the same DGP as the dynamic model, except for $\rho_{0}=0$ ($\theta_{0}=\beta_{0}$). We demonstrate the finite sample performance of the MLE of $\theta_{0}$ by maximizing our bias-corrected likelihood $L(\theta),$ and compare it with the uncorrected (original) MLE obtained from $\widehat{l}(\theta)$ without bias correction. In addition, we compare hypothesis testing in terms of size and power based on both the corrected and uncorrected likelihoods. We consider the sample size $(N,T)\in\{(30,30),(60,60),(90,90)\},$ with 1000 replications for each DGP.
Table 1 reports the estimation bias and RMSE of our bias-corrected and the uncorrected MLEs for both the static ($\theta_{0}=\beta_{0}$) and dynamic ($\theta_{0}=(\rho_{0},\beta_{0})^{\prime}$) logit models. In the static model, $\widehat{\beta}$ represents the uncorrected MLE from $\widehat{l}(\theta)$, and $\widehat{\beta}_{L}$ is our bias-corrected MLE. For the dynamic model, $(\widehat{\beta},\widehat{\rho})$ denotes the uncorrected MLEs, while $(\widehat{\beta}_{L}^{(\tau)},\widehat{\rho}_{L}^{(\tau)})^{\prime}$ represents our bias-corrected estimates, with the truncation parameter $\tau$ set to $\tau=1$ or $2$. In the static model, ${Y_{it}}$ are generated independently across time $t$ where we set $\tau=0$ for the estimation, while the data in the dynamic model exhibit serial dependence.
Table 1: Comparison between Bias-corrected and Uncorrected MLEs
$(N,T)$ $(30,30)$ $(60,60)$ $(90,90)$
Bias RMSE Bias RMSE Bias RMSE
Static Model
$\widehat{\beta}$ $0.158$ $0.253$ $0.075$ $0.109$ $0.051$ $0.072$
$\widehat{\beta}_{L}$ $0.050$ $0.175$ $0.018$ $0.073$ $0.010$ $0.049$
Dynamic Model
$\widehat{\beta}$ $0.165$ $0.261$ $0.074$ $0.111$ $0.046$ $0.071$
$\widehat{\beta}_{L}^{(1)}$ $0.059$ $0.181$ $0.017$ $0.077$ $0.006$ $0.051$
$\widehat{\beta}_{L}^{(2)}$ $0.061$ $0.183$ $0.018$ $0.078$ $0.007$ $0.051$
$\widehat{\rho}$ $-0.107$ $0.201$ $-0.048$ $0.093$ $-0.033$ $0.058$
$\widehat{\rho}_{L}^{(1)}$ $-0.044$ $0.159$ $-0.014$ $0.076$ $-0.010$ $0.047$
$\widehat{\rho}_{L}^{(2)}$ $-0.038$ $0.161$ $-0.009$ $0.077$ $-0.006$ $0.046$
[…]
    Notes: The true structure parameters are $\theta_{0}=(\beta_{0},\rho_{0})^{\prime}=(0.5,0.5)^{\prime}$ for the dynamic model, and $\theta_{0}=\beta_{0}=0.5$ for the static model. In the static model, $\widehat{\beta}$ and $\widehat{\beta}_{L}$ represent the uncorrected MLE and our bias-corrected MLE, respectively. In the dynamic model, $(\widehat{\beta},\widehat{\rho})$ is the uncorrected MLE and $(\widehat{\beta}_{L}^{(\tau)},\widehat{\rho}_{L}^{(\tau)})^{\prime}$ represents our bias-corrected MLE with the truncation parameter $\tau$ set to $\tau=1$ or $2$. The number of replications is $1000$.
The results in Table 1 reveal that the uncorrected MLEs are severely biased in both models, while our bias-corrected estimators substantially reduce the bias. For instance, in the static model with $(N,T)=(30,30)$, the bias for $\widehat{\beta}_{L}$ is only $0.05$, which is less than $1/3$ of the significant bias of $0.158$ for $\widehat{\beta}$. As the sample size increases to $(N,T)=(90,90)$, $\widehat{\beta}_{L}$ is nearly unbiased, while the bias for $\widehat{\beta}$ remaining five times larger than that for $\widehat{\beta}_{L}$. Furthermore, the RMSEs form our corrected MLEs are consistently lower than those of the uncorrected MLEs, especially with smaller sample sizes, underscoring the advantage of our bias correction method. Additionally, in the dynamic model, the choice of $\tau$ has little impact on the performance of our bias-corrected estimators.
We also conduct LR tests using both bias-corrected and uncorrected likelihoods. Consider the null hypothesis $H_{0}:\theta_{0}=\theta_{0}+\delta$ with $\delta\in\{0,\pm 0.2,\pm 0.15,\pm 0.1\}.$ We report the empirical size ($\delta=0$) and power ($\delta\neq 0$) at a $5\%$ significance level. Table 2 presents the empirical size based on corrected and uncorrected likelihoods, where the corrected likelihoods use truncation $\tau=0$ for the static model and $\tau=1$ for the dynamic model. The results show that the LR tests using uncorrected likelihoods suffer from severe size distortion, while the empirical size from our bias-corrected likelihoods maintains the significance level as the sample size increases. Additionally, our bias-corrected tests significantly reduce Type-I error even for small sample size.
Table 2: Empirical Size of LR Tests Based on Bias-corrected and Uncorrected Likelihoods
Static Model Dynamic Model
$(N,T)$ Uncorrected Corrected Uncorrected Corrected
$(30,30)$ $16\%$ $9\%$ $18\%$ $10\%$
$(60,60)$ $15\%$ $5\%$ $17\%$ $7\%$
$(90,90)$ $18\%$ $6\%$ $16\%$ $5\%$
[…]
    Notes: The null hypothesis $H_{0}:(\beta_{0},\rho_{0})^{\prime}=(0.5,0.5)^{\prime}$ for the dynamic model, and $H_{0}:\beta_{0}=0.5$ for the static model, with a significance level $5\%$. The number of replications is $1000$.
Table 3 presents the empirical power of LR tests based on our bias-corrected likelihoods. The empirical power increases as both the value of $\delta$ and the sample size grow. For $(N,T)=(90,90)$, when $\delta=\pm 0.2$, the rejection rates reach nearly 100% for both the static and dynamic models. We also perform LM and Wald tests, reporting empirical size and power under the same DGPs in the appendix. The results align with those of the LR tests. Additionally, we perform simulations using probit models with results provided in the appendix.
Table 3: Empirical Power of LR Tests Based on Bias-corrected Likelihoods
$\delta$
$(N,T)$ $-0.20$ $-0.15$ $-0.10$ $0.10$ $0.15$ $0.20$
Static Model
$(30,30)$ $45\%$ $32\%$ $23\%$ $10\%$ $14\%$ $21\%$
$(60,60)$ $86\%$ $63\%$ $37\%$ $25\%$ $49\%$ $74\%$
$(90,90)$ $100\%$ $93\%$ $65\%$ $50\%$ $84\%$ $98\%$
Dynamic Model
$(30,30)$ $48\%$ $33\%$ $21\%$ $18\%$ $28\%$ $43\%$
$(60,60)$ $95\%$ $74\%$ $40\%$ $45\%$ $75\%$ $94\%$
$(90,90)$ $100\%$ $98\%$ $76\%$ $78\%$ $98\%$ $100\%$
[…]
    Notes: The null hypothesis $H_{0}:(\beta_{0},\rho_{0})^{\prime}=(0.5+\delta,0.5+\delta)^{\prime}$ for the dynamic model, and $H_{0}:\beta_{0}=0.5+\delta$ for the static model, with a significance level $5\%,$ where $\delta\in\{\pm 0.2,\pm 0.15,\pm 0.1\}$. The number of replications is $1000$.
Finally, we compare our bias correction method to the Jackknife approach by Fernández-Val and Weidner (2016), who consider a probit model in their simulation. We consider a similar setup, using the same settings as in (4.1), except that the disturbance term $\varepsilon_{it}$ follows a standard normal distribution. Table 4 displays the estimation bias and RMSE of the uncorrected MLEs, our bias-corrected MLEs, and the Jackknife bias-corrected estimators, both for the static ($\theta_{0}=\beta_{0}$) and dynamic ($\theta_{0}=(\rho_{0},\beta_{0})^{\prime}$) probit models. In the static model, $\widehat{\beta}_{L}$ is our bias-corrected MLE with $\tau=0$, and $\widehat{\beta}_{J}$ represents the Jackknife-corrected MLE. In the dynamic model, $(\widehat{\beta}_{L},\widehat{\rho}_{L})^{\prime}$ is our bias-corrected MLE with $\tau=1$, and $(\widehat{\beta}_{J},\widehat{\rho}_{J})^{\prime}$ is the Jackknife-corrected MLE. Table 4 shows that our bias-corrected MLEs outperform both the uncorrected and Jackknife bias-corrected estimators, with consistently smaller RMSEs across all DGPs. The bias correction by our method is significantly greater than that of the Jackknife approach, especially when the sample size is small. For example, in the static model with $(N,T)=(30,30)$, the bias of our corrected MLE is only $0.048$, while the bias of the Jackknife estimator is $-1.228,$ over twenty-five times larger.
Table 4: Comparison with the Jackknife Bias Correction Method
$(N,T)$ $(30,30)$ $(60,60)$ $(90,90)$
Bias RMSE Bias RMSE Bias RMSE
Static Model
$\widehat{\beta}$ $0.144$ $0.197$ $0.061$ $0.081$ $0.041$ $0.052$
$\widehat{\beta}_{L}$ $0.048$ $0.123$ $0.011$ $0.051$ $0.007$ $0.031$
$\widehat{\beta}_{J}$ $-1.228$ $5.930$ $-0.022$ $0.058$ $-0.007$ $0.032$
Dynamic Model
$\widehat{\beta}$ $0.172$ $0.262$ $0.069$ $0.088$ $0.042$ $0.054$
$\widehat{\beta}_{L}$ $0.072$ $0.207$ $0.017$ $0.053$ $0.007$ $0.032$
$\widehat{\beta}_{J}$ $-1.339$ $3.873$ $-0.029$ $0.101$ $-0.009$ $0.034$
$\widehat{\rho}$ $-0.050$ $0.122$ $-0.028$ $0.056$ $-0.017$ $0.037$
$\widehat{\rho}_{L}$ $-0.025$ $0.104$ $-0.014$ $0.049$ $-0.007$ $0.032$
$\widehat{\rho}_{J}$ $0.003$ $0.115$ $-0.001$ $0.050$ $0.000$ $0.033$
[…]
    Notes: $(\beta_{0},\rho_{0})^{\prime}=(0.5,0.5)^{\prime}$ for the dynamic model, and $\beta_{0}=0$ for the static model. Accordingly, $(\widehat{\beta},\widehat{\rho})^{\prime}$ (or $\widehat{\beta}$) denotes the uncorrected MLE; $(\widehat{\beta}_{L},\widehat{\rho}_{L})^{\prime}$ (or $\widehat{\beta}_{L}$) is our bias-corrected MLE with truncation parameter $\tau=1$ for the dynamic model or $\tau=0$ for the static model; and $(\widehat{\beta}_{J},\widehat{\rho}_{J})^{\prime}$ (or $\widehat{\beta}_{J}$) represents the Jackknife bias-corrected estimators. The number of replications is $1000$.
5 Empirical Analysis
In this section we apply our likelihood-based bias correction method to examining the determinants of the labor force participation (LFP) decision of single mothers. In particular, we look at the impact of the number of children on the decision of the mother to engage in paid employment.
An extensive literature in labor economics has studied the labor supply decisions of married women (Killingsworth and Heckman, 1986; Angrist and Evans, 1998; Blau and Kahn, 2007). These studies have uncovered the impacts of a variety of economic variables, including female market wage and husband income (Mincer, 1962), education (Heath and Jayachandran, 2018), childcare costs (Connelly, 1992), the cost of home technology (Greenwood et al., 2016), and culture norms (Fernández, 2013). Among these variables, the number of children consistently emerges as one of the most important determinants of female labor supply (Nakamura and Nakamura, 1992). This is unsurprising since women continue to bear a disproportionate share of child-rearing responsibilities (Aguero and Marks, 2008).
In contrast to the substantial body of research on the labor supply behavior of married women, the labor supply decisions of single mothers have received relatively limited attention, with only a few exceptions (Kimmel, 1998; Blundell et al., 2016). Single mothers, however, face distinctive challenges when it comes to balancing work and child-rearing responsibilities due to the absence of a second earner in the household. Moreover, their employment decisions may have a more pronounced impact on their children’s well-being compared to the decisions of married women and should thus be of great importance to economists and policymakers.
To study the labor supply decisions of single mothers, we compile a data set from waves $20-30$ of the Panel Study of Income Dynamics (PSID), which span the period of 1987 to 1997. Our sample includes only single mothers, defined as unmarried female household heads with at least one child. The dependent variable is labor force participation status, defined as whether the individual worked any hours during the interview year. Our main explanatory variable is the number of children under 18 in the household. Following Dhaene and Jochmans (2015), we restrict our sample to individuals aged $18-60$ whose LFP status changed at least once during the period. The final sample comprises $N=90$ single mothers whose labor supply status is continuously observed over $T=10$ years from 1987 to 1996. Further details on data construction are provided in Appendix C. On average, 27 percent of these individuals switched in and out of the labor force in each year (Figure C.1--C.2). Table C.1 presents the summary statistics.
Let $Y_{it}\in\left\{0,1\right\}$ denote the labor force participation status of single mother $i$ in year $t$. Let $X_{it}$ denote her number of children. To assess the impact of $X_{it}$ on $Y_{it}$, we estimate the following two logit regressions:
$Y_{it}$ $=\mathds{1}\left\{\rho Y_{it-1}+\beta X_{it}+\mu_{i}+\varphi_{t}+\varepsilon_{it}>0\right\}$ (5.1)
$Y_{it}$ $=\mathds{1}\left\{\rho Y_{it-1}+\left(\beta+\alpha_{i}+\gamma_{t}\right)X_{it}+\mu_{i}+\varphi_{t}+\varepsilon_{it}>0\right\},$ (5.2)
where $\varepsilon_{it}$ follows a standard logistic distribution, and $\sum_{i}\mu_{i}=\sum_{i}\alpha_{i}=\sum_{i}\gamma_{t}=0$. Model (5.1) is a dynamic homogeneous slope model where $\beta$ represents the average effect of an additional child on the log odds of a single mother’s LFP. Model (5.2) is a dynamic heterogeneous slope model, where $\beta_{it}=\beta+\alpha_{i}+\gamma_{t}$ captures the heterogeneous impact $X_{it}$ may have on the log odds of $Y_{it}$. Model (5.2) nests model (5.1) as a special case when $\beta_{it}$ does not depend on $i$ and $t$.
As discussed in section 1 , when the true impact of $X_{it}$ varies among the population, model (5.1) is appropriate only if $\beta_{it}$ and $X_{it}$ are uncorrelated. In practice, they may correlate due to self-selection. For instance, single mothers who choose to have more children might be those with sufficient financial resources or access to childcare services, such that their labor force participation is less affected by additional children. Alternatively, a correlation between $\beta_{it}$ and $X_{it}$ could arise from common time trends. As we track single mother households from 1987 to 1996, more children were born while factors such as rising childcare and schooling costs, improvements in home technology, and evolving welfare policies changed the impact of each additional child on their mother’s LFP. In both scenarios, model (5.2) would be the appropriate model to use as it controls for unobserved heterogeneity in $\beta_{it}$ by incorporating individual and time fixed effects into the slope and allowing them to correlate with $X_{it}$.
Given our small sample size, direct estimation of both models could suffer from severe incidental parameter problems, which we address using our bias correction procedure. Table 5 presents the estimation results. Panel A reports results for model (5.1) under the homogeneous slope assumption. For each parameter, we provide the MLE and bias-corrected estimates, along with their standard errors. Examining the bias-corrected estimates, we observe that an increase in the number of children appears to reduce a single mother’s likelihood of labor force participation, while past LFP status strongly predicts current LFP. However, this relationship shifts dramatically once we account for slope heterogeneity. Panel B presents the results for model (5.2), allowing for heterogeneous slopes. Here, the bias-corrected estimates reveal a positive average impact of additional children on a mother’s labor supply: each additional child on average increases the log odds of labor force participation by 0.126, which corresponds to an average increase of 5.5 percent in the likelihood of employment. This indicates that, on average, single mothers with more children are indeed more likely to work, likely due to the financial demands of supporting multiple children independently. Additionally, unlike the homogeneous slope model, the serial correlation in labor force participation remains statistically significant but diminishes in magnitude, suggesting that what appears to be strong state dependence in LFP under the homogeneous slope model may in part be driven by unobserved heterogeneity in how children affect labor supply decisions across individuals. Together, these findings underscore the importance of controlling for unobserved slope heterogeneity, as failing to do so can lead to substantially different conclusions.
In addition to highlighting the difference between homogeneous and heterogeneous slope models, Table 5 demonstrates the importance of bias-correction in estimating dynamic nonlinear panel data models. Examining the results in Panel B, we observe that the MLE overestimates the impact of the number of children while underestimating the degree of serial correlation in LFP. Bias correction results in an 18 percent lower estimate of the former and a 57 percent higher estimate of the latter. Moreover, the corrected estimates yield smaller standard errors, improving statistical significance across all parameters.
Finally, in Figure 2, we plot the distribution of the estimated $\beta_{it}$ from model (5.2), with summary statistics in Table C.2. Each $\beta_{it}$ estimate is constructed as the sum of the bias-corrected common effect $\beta$ and the estimated individual and time effects $\alpha_{i}$ and $\gamma_{t}$. These estimates reveal substantial heterogeneity in how the number of children influences single mothers’ labor force participation. While the mean effect is positive, over half of the individual-time observations show a decrease in LFP with each additional child, suggesting that the average effect alone may inadequately capture the diverse impacts across individuals and over time. This reinforces the importance of accounting for both individual and temporal variations in slope coefficients, as the impact of children on labor force participation can vary significantly based on unobserved factors unique to each mother and time period.
In conclusion, our analysis demonstrates the efficacy of our likelihood-based bias correction method for estimating dynamic nonlinear models of labor supply with multidimensional heterogeneities. We show the importance of controlling for unobserved slope heterogeneity when it may correlate with regressors and the value of panel data models that accommodate two-way fixed effects in both the intercept and the slope. Our findings reveal that the number of children has, on average, a positive influence on a single mother’s labor force participation, but such effect is highly heterogeneous across individuals and time periods. This heterogeneity underscores the complex nature of labor supply decisions and the unique challenges each single mother faces in balancing work and child-rearing responsibilities.
Table 5: Single Mother Labor Force Participation: Estimation Results
         Number of Children          Lagged Participation
Uncorrected Corrected Uncorrected Corrected
                Panel A. Homogeneous Slope
Model parameter -0.954 -0.880 0.584 0.535
(0.240) (0.218) (0.113) (0.103)
                Panel B. Heterogeneous Slope
Model parameter 0.153 0.126 0.051 0.080
(0.021) (0.000) (0.020) (0.000)
  • Notes: Standard errors in parentheses. Panel A reports the estimation results of the homogeneous slope model. Panel B reports the estimation results of the heterogeneous slope model. The “uncorrected” and the “corrected” columns show respectively the MLE and the bias-corrected parameter estimates. Data source: PSID 1987 - 1997.
Figure 2: Heterogeneous Impact of Children on Single Mother Labor Force Participation
Acknowledgement
We would like to express our gratitude to valuable comments and suggestions from Geert Dhaene, Jinyong Hahn, Chen Hsiao, Whitney Newey, and Jun Yu. Our research assistant Yuchi Liu helped us with various tasks for the empirical analysis. Leng’s research is partially supported by the National Natural Science Foundation of China Grant 72001183. Sun acknowledges the support from the National Natural Science Foundation of China under Grant Number 72203032.
Appendix
[…]
AppendixAppendix
A Proofs
A.1 Lemmas
Proof of Lemma 1. Note that ${H}(\theta)=D\big{[}\partial_{\phi\phi^{\prime}}l(\theta,\phi)\mid_{\phi=\phi(\theta)}\big{]}D^{\prime},$ where $D$ is a $(N+T-2)K\times(N+T)K$ matix. Since $K$ is fixed, we assume without loss of generality that $K=1.$ We express $H(\theta)$ as:
$H(\theta)=\begin{pmatrix}H_{\breve{\alpha}\breve{\alpha}^{\prime}}\left(\theta\right),&H_{\breve{\alpha}\breve{\gamma}^{\prime}}\left(\theta\right)\\ H_{\breve{\gamma}\breve{\alpha}^{\prime}}\left(\theta\right),&H_{\breve{\gamma}\breve{\gamma}^{\prime}}\left(\theta\right)\end{pmatrix}.$
Consider any element $h_{i,j}^{\breve{\alpha}}(\theta)$ of $N\cdot H_{\breve{\alpha}\breve{\alpha}^{\prime}}(\theta),$ where $i,j=1,\ldots,N-1.$ Under Assumptions 1 and 1, by applying Davydov’s inequality, it follows that
$\sup_{i,j}\sup_{\theta\in\Theta}\text{Var}\big{(}h_{i,j}^{\breve{\alpha}}(\theta)\big{)}=o(1)$
as $N$ and $T\rightarrow\infty.$ Consequently, by the Markov inequality,
$\sup_{i,j}\sup_{\theta\in\Theta}\big{\{}h_{i,j}^{\breve{\alpha}}(\theta)-\mathbb{E}[h_{i,j}^{\breve{\alpha}}(\theta)]\big{\}}=o_{\mathbb{P}}(1).$
Let $h_{t,s}^{\breve{\gamma}}(\theta)$ denote the $(t,s)$-th element of $T\cdot H_{\breve{\gamma}\breve{\gamma}^{\prime}}(\theta)$ for $t,s=1,\ldots,T-1.$ By conditional cross-sectional independence, we also have
$\sup_{t,s}\sup_{\theta\in\Theta}\big{\{}h_{t,s}^{\breve{\gamma}}(\theta)-\mathbb{E}[h_{t,s}^{\breve{\gamma}}(\theta)]\big{\}}=o_{\mathbb{P}}(1).$
Moreover, for each $h_{i,t}^{\breve{\alpha}\breve{\gamma}}(\theta),$ the $(i,t)$-th element of $NT\cdot H_{\breve{\alpha}\breve{\gamma}^{\prime}}(\theta),$ with $i=1,\ldots,N-1$ and $t=1,\ldots,T-1,$ by Assumption 1 we have that
$\sup_{i,t}\sup_{\theta\in\Theta}\big{\{}h_{i,t}^{\breve{\alpha}\breve{\gamma}}(\theta)-\mathbb{E}[h_{i,t}^{\breve{\alpha}\breve{\gamma}}(\theta)]\big{\}}=o_{\mathbb{P}}(1).$
Therefore,
$N\left\|H_{\breve{\alpha}\breve{\alpha}^{\prime}}(\theta)-\mathbb{E}[H_{\breve{\alpha}\breve{\alpha}^{\prime}}(\theta)]\right\|_{\max}=o_{\mathbb{P}}(1),\quad T\left\|H_{\breve{\gamma}\breve{\gamma}^{\prime}}(\theta)-\mathbb{E}[H_{\breve{\gamma}\breve{\gamma}^{\prime}}(\theta)]\right\|_{\max}=o_{\mathbb{P}}(1),$
and
$NT\left\|H_{\breve{\alpha}\breve{\gamma}^{\prime}}(\theta)-\mathbb{E}[H_{\breve{\alpha}\breve{\gamma}^{\prime}}(\theta)]\right\|_{\max}=o_{\mathbb{P}}(1),$
uniformly for $\theta\in\Theta,$ where $\|A\|_{\max}$ denotes the max norm of a matrix $A$. Since $\mathbb{E}[s(\theta)]=0,$ $s(\theta)=s(\theta)-\mathbb{E}[s(\theta)]=:(\widetilde{s}_{\breve{\alpha}}^{\prime}(\theta),\widetilde{s}_{\breve{\gamma}}^{\prime}(\theta))^{\prime}.$ We now decompose the quadratic form on the right hand side of (2.7):
$s^{\prime}(\theta)[-{H}(\theta)]^{-1}s\left(\theta\right)$
$=$ $[s(\theta)-\mathbb{E}s(\theta)]^{\prime}[-{H}\left(\theta\right)]^{-1}[s(\theta)-\mathbb{E}s(\theta)]$
$=$ $[s(\theta)-\mathbb{E}s(\theta)]^{\prime}[-\mathbb{E}{H}\left(\theta\right)]^{-1}[s(\theta)-\mathbb{E}s(\theta)]^{\prime}(1+o_{\mathbb{P}}(1))$
$=$ $-\big{[}\widetilde{s}_{\breve{\alpha}}^{\prime}(\theta)H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\alpha}}(\theta)+\widetilde{s}_{\breve{\gamma}}^{\prime}(\theta)H_{\breve{\gamma}\breve{\gamma}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\gamma}}(\theta)+2\widetilde{s}_{\breve{\alpha}}^{\prime}(\theta)H_{\breve{\alpha}\breve{\gamma}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\gamma}}\left(\theta\right)\big{]}(1+o_{\mathbb{P}}(1)).$ (A.1)
By matrix inversion in block form, it follows that
$\sup_{\theta\in\Theta}\left\|H_{\breve{\alpha}\breve{\gamma}^{\prime}}^{\ast}(\theta)\right\|_{\max}=O(1),$
and
$\sup_{\theta\in\Theta}\left\|H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)\right\|_{\max}=O(T),\quad\sup_{\theta\in\Theta}\left\|H_{\breve{\gamma}\breve{\gamma}^{\prime}}^{\ast}\left(\theta\right)\right\|_{\max}=O(N).$
In the following, we show
$\sup_{\theta\in\Theta}\mathbb{E}\big{[}\widetilde{s}_{\breve{\alpha}}^{\prime}(\theta)H_{\breve{\alpha}\breve{\gamma}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\gamma}}(\theta)\big{]}=O\big{(}1/(NT)\big{)},$ (A.2)
and
$\sup_{\theta\in\Theta}\,\mathbb{E}\big{[}\widetilde{s}_{\breve{\alpha}}^{\prime}(\theta)H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}\left(\theta\right)\widetilde{s}_{\breve{\alpha}}(\theta)\big{]}=O(1/T),\quad\sup_{\theta\in\Theta}\,\mathbb{E}\big{[}\widetilde{s}_{\breve{\gamma}}^{\prime}(\theta)H_{\breve{\gamma}\breve{\gamma}^{\prime}}^{\ast}\left(\theta\right)\widetilde{s}_{\breve{\gamma}}(\theta)\big{]}=O(1/N).$ (A.3)
Therefore, $\mathbb{E}\big{[}\widetilde{s}_{\breve{\alpha}}^{\prime}\left(\theta\right)H_{\breve{\alpha}\breve{\gamma}^{\prime}}^{\ast}\left(\theta\right)\widetilde{s}_{\breve{\gamma}}(\theta)\big{]}$ is negligible in the expansion for $\mathbb{E}[\widehat{l}(\theta)]$ compared to the other two terms from the quadratic form.
Let
$\partial_{\alpha_{i}}\widetilde{l}_{i,t}(\theta,\phi(\theta)):=\{\partial_{\alpha_{i}}l_{i,t}(\theta,\phi)-\mathbb{E}[\partial_{\alpha_{i}}l_{i,t}(\theta,\phi)]\}|_{\phi=\phi(\theta)}$
and
$\partial_{\gamma_{t}}\widetilde{l}_{i,t}(\theta,\phi(\theta)):=\{\partial_{\gamma_{t}}l_{i,t}(\theta,\phi)-\mathbb{E}[\partial_{\gamma_{t}}l_{i,t}(\theta,\phi)]\}|_{\phi=\phi(\theta)}$
for $i=1,\ldots,N$ and $t=1,\ldots,T,$ respectively. Define $h_{i,t}^{\ast\breve{\alpha}\breve{\gamma}}(\theta)$ as the $(i,t)$-th element of $H_{\breve{\alpha}\breve{\gamma}^{\prime}}^{\ast}(\theta)$ where $i=1,\ldots,N-1$ and $t=1,\ldots,T-1.$ Under the assumption of conditional cross-sectional independence, we have
$\mathbb{E}\big{\{}(N\sqrt{T}\widetilde{s}_{\breve{\alpha}}^{\prime}\left(\theta\right))H_{\breve{\alpha}\breve{\gamma}^{\prime}}^{\ast}\left(\theta\right)(\sqrt{N}T\widetilde{s}_{\breve{\gamma}}(\theta))\big{\}}$
$=$ $\frac{1}{\sqrt{NT}}\mathbb{E}\bigg{\{}\sum_{i=1}^{N-1}\sum_{t=1}^{T-1}h_{i,t}^{\ast\breve{\alpha}\breve{\gamma}}(\theta)\cdot\bigg{[}\sum_{s=1}^{T}\big{(}\partial_{\alpha_{i}}\widetilde{l}_{i,s}(\theta,\phi(\theta))-\partial_{\alpha_{N}}\widetilde{l}_{N,s}(\theta,\phi(\theta))\big{)}\bigg{]}$
$\hskip 120.00018pt\cdot\bigg{[}\sum_{j=1}^{N}\big{(}\partial_{\gamma_{t}}\widetilde{l}_{j,t}(\theta,\phi(\theta))-\partial_{\gamma_{T}}\widetilde{l}_{j,T}(\theta,\phi(\theta))\big{)}\bigg{]}\bigg{\}}$
$=$ $\frac{1}{\sqrt{NT}}\sum_{i=1}^{N-1}\mathbb{E}\left[\bigg{(}\sum_{s=1}^{T}\partial_{\alpha_{i}}\widetilde{l}_{i,s}(\theta,\phi(\theta))\bigg{)}\bigg{(}\sum_{t=1}^{T-1}\partial_{\gamma_{t}}\widetilde{l}_{i,t}(\theta,\phi(\theta))\cdot h_{i,t}^{\ast\breve{\alpha}\breve{\gamma}}(\theta)\bigg{)}\right]$
$\quad-\frac{1}{\sqrt{NT}}\sum_{i=1}^{N-1}\left[\bigg{(}\sum_{t=1}^{T-1}h_{i,t}^{\ast\breve{\alpha}\breve{\gamma}}(\theta)\bigg{)}\mathbb{E}\bigg{(}\sum_{s=1}^{T}\partial_{\alpha_{i}}\widetilde{l}_{i,s}(\theta,\phi(\theta))\partial_{\gamma_{T}}\widetilde{l}_{i,T}(\theta,\phi(\theta))\bigg{)}\right]$
$\quad-\frac{1}{\sqrt{NT}}\mathbb{E}\left[\bigg{(}\sum_{s=1}^{T}\partial_{\alpha_{N}}\widetilde{l}_{N,s}(\theta,\phi(\theta))\bigg{)}\left(\sum_{t=1}^{T-1}\partial_{\gamma_{t}}\widetilde{l}_{N,t}(\theta,\phi(\theta))\bigg{(}\sum_{i=1}^{N-1}h_{i,t}^{\ast\breve{\alpha}\breve{\gamma}}(\theta)\bigg{)}\right)\right]$
$\quad+\frac{1}{\sqrt{NT}}\bigg{(}\sum_{i=1}^{N-1}\sum_{t=1}^{T-1}h_{i,t}^{\ast\breve{\alpha}\breve{\gamma}}(\theta)\bigg{)}\mathbb{E}\bigg{(}\sum_{s=1}^{T}\partial_{\alpha_{N}}\widetilde{l}_{N,s}(\theta,\phi(\theta))\partial_{\gamma_{T}}\widetilde{l}_{N,T}(\theta,\phi(\theta))\bigg{)}$
=:absent:=:= : $A_{NT,1}(\theta)-A_{NT,2}(\theta)-A_{NT,3}(\theta)+A_{NT,4}(\theta).$ (A.4)
By Assumptions 1 and 1, using Davydov’s inequality, we obtain for some $\eta>2$ that
$|A_{NT,1}(\theta)|$
$=$ $\frac{1}{\sqrt{NT}}\sum_{i=1}^{N-1}\left|\mathbb{E}\left[\bigg{(}\sum_{s=1}^{T-1}\partial_{\alpha_{i}}\widetilde{l}_{i,s}(\theta,\phi(\theta))\bigg{)}\bigg{(}\sum_{t=1}^{T-1}\partial_{\gamma_{t}}\widetilde{l}_{i,t}(\theta,\phi(\theta))\cdot h_{i,t}^{\ast\breve{\alpha}\breve{\gamma}}(\theta)\bigg{)}\right.\right.$
$\hskip 56.9055pt\left.\left.+\sum_{t=1}^{T-1}\partial_{\gamma_{t}}\widetilde{l}_{i,t}(\theta,\phi(\theta))\partial_{\alpha_{i}}\widetilde{l}_{i,T}(\theta,\phi(\theta))\cdot h_{i,t}^{\ast\breve{\alpha}\breve{\gamma}}(\theta)\right]\right|$
$\leq$ $\sqrt{NT}\sup_{\theta\in\Theta}\left\|H_{\breve{\alpha}\breve{\gamma}^{\prime}}^{\ast}(\theta)\right\|_{\max}\cdot\sup_{i,t}\{\mathbb{E}[g^{\eta}(w_{it})]\}^{2/\eta}$
$\quad\cdot\sup_{i}\left([a_{i}(0)]^{1-2/\eta}+2\sum_{m=1}^{T-2}\big{(}1-m/(T-1)\big{)}[a_{i}(m)]^{1-2/\eta}+\frac{1}{T-1}\sum_{m=1}^{T-1}[a_{i}(m)]^{1-2/\eta}\right)$
$=$ $O\big{(}\sqrt{NT}\big{)},$
$|A_{NT,2}(\theta)|$ $=\frac{T}{\sqrt{NT}}\sum_{i=1}^{N-1}\bigg{(}\frac{1}{T}\sum_{t=1}^{T-1}|h_{i,t}^{\ast\breve{\alpha}\breve{\gamma}}(\theta)|\bigg{)}\cdot\left|\mathbb{E}\bigg{(}\sum_{s=1}^{T}\partial_{\alpha_{i}}\widetilde{l}_{i,s}(\theta,\phi(\theta))\partial_{\gamma_{T}}\widetilde{l}_{i,T}(\theta,\phi(\theta))\bigg{)}\right|$
$\leq\sqrt{NT}\sup_{\theta\in\Theta}\left\|H_{\breve{\alpha}\breve{\gamma}^{\prime}}^{\ast}(\theta)\right\|_{\max}\cdot\sup_{i,t}\{\mathbb{E}[g^{\eta}(w_{it})]\}^{2/\eta}\cdot\sup_{i}\left(\sum_{m=0}^{T-1}[a_{i}(m)]^{1-2/\eta}\right)=O\big{(}\sqrt{NT}\big{)},$
$|A_{NT,3}(\theta)|$
$=$ $\frac{N}{\sqrt{NT}}\left|\mathbb{E}\left[\left(\sum_{s=1}^{T}\partial_{\alpha_{N}}\widetilde{l}_{N,s}(\theta,\phi(\theta))\right)\left(\sum_{t=1}^{T-1}\partial_{\gamma_{t}}\widetilde{l}_{N,t}(\theta,\phi(\theta))\bigg{(}\frac{1}{N}\sum_{i=1}^{N-1}h_{i,t}^{\ast\breve{\alpha}\breve{\gamma}}(\theta)\bigg{)}\right)\right]\right|$
$\leq$ $\sqrt{NT}\sup_{\theta\in\Theta}\left\|H_{\breve{\alpha}\breve{\gamma}^{\prime}}^{\ast}(\theta)\right\|_{\max}\cdot\sup_{i,t}\{\mathbb{E}[g^{\eta}(w_{it})]\}^{2/\eta}$
$\quad\cdot\sup_{i}\left([a_{i}(0)]^{1-2/\eta}+2\sum_{m=1}^{T-2}\big{(}1-m/(T-1)\big{)}[a_{i}(m)]^{1-2/\eta}+\frac{1}{T-1}\sum_{m=1}^{T-1}[a_{i}(m)]^{1-2/\eta}\right)$
$=$ $O\big{(}\sqrt{NT}\big{)},$
and
$|A_{NT,4}(\theta)|\leq\sqrt{NT}\sup_{\theta\in\Theta}\left\|H_{\breve{\alpha}\breve{\gamma}^{\prime}}^{\ast}(\theta)\right\|_{\max}\cdot\sup_{i,t}\{\mathbb{E}[g^{\eta}(w_{it})]\}^{2/\eta}\cdot\sup_{i}\left(\sum_{m=0}^{T-1}[a_{i}(m)]^{1-2/\eta}\right)=O\big{(}\sqrt{NT}\big{)}.$
Thus, we obtain (A.2) from (A.4).
Let $h_{i,j}^{\ast\breve{\alpha}}(\theta)$ denote the $(i,j)$-th element of $T^{-1}H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta),$ where $i,j=1,\ldots,N-1.$ By Assumption 1, we know that $\sup_{i\geq 1}h_{i,i}^{\ast\breve{\alpha}}(\theta)=O(1)$ uniformly in $\theta.$ Define $\alpha:=(\alpha_{1}^{\prime},\ldots,\alpha_{N}^{\prime})^{\prime}$ and $\gamma:=(\gamma_{1}^{\prime},\ldots,\gamma_{T}^{\prime})^{\prime}.$ Noting that
$N\cdot H_{\breve{\alpha}\breve{\alpha}^{\prime}}(\theta)=D_{1}[N\cdot\partial_{\alpha\alpha^{\prime}}l(\theta,\phi)|_{\phi=\phi(\theta)}]D_{1}^{\prime},$
$N\cdot H_{\breve{\alpha}\breve{\alpha}^{\prime}}(\theta)$ can be written as the sum of a diagonal matrix and a rank-one matrix. Thus, it follows from the lemma in Miller (1981) that $\sup_{i\neq j}h_{i,j}^{\ast\breve{\alpha}}(\theta)=O(1/N)$ uniformly in $\theta.$ Since $s(\theta)=D\partial_{\phi}l(\theta,\phi)|_{\phi=\phi(\theta)},$ we have
$T\mathbb{E}\big{[}\widetilde{s}_{\breve{\alpha}}^{\prime}(\theta)H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\alpha}}(\theta)\big{]}$
$=$ $\frac{1}{N}\mathbb{E}\big{\{}(N\sqrt{T}\widetilde{s}_{\breve{\alpha}}^{\prime}(\theta))[N^{-1}H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)](N\sqrt{T}\widetilde{s}_{\breve{\alpha}}(\theta))\big{\}}$
$=$ $\frac{1}{N}\sum_{i=1}^{N-1}h_{i,i}^{\ast\breve{\alpha}}(\theta)\,\mathbb{E}\bigg{(}\frac{1}{\sqrt{T}}\sum_{t=1}^{T}\big{[}\partial_{\alpha_{i}}\widetilde{l}_{i,t}(\theta,\phi(\theta))-\partial_{\alpha_{N}}\widetilde{l}_{N,t}(\theta,\phi(\theta))\big{]}\bigg{)}^{2}$
$+\frac{2}{N}\sum_{1\leq i<j\leq N-1}h_{i,j}^{\ast\breve{\alpha}}(\theta)\,\mathbb{E}\left[\bigg{(}\frac{1}{\sqrt{T}}\sum_{t=1}^{T}\big{[}\partial_{\alpha_{i}}\widetilde{l}_{i,t}(\theta,\phi(\theta))-\partial_{\alpha_{N}}\widetilde{l}_{N,t}(\theta,\phi(\theta))\big{]}\bigg{)}\right.$
$\hskip 116.65646pt\left.\cdot\bigg{(}\frac{1}{\sqrt{T}}\sum_{t=1}^{T}\big{(}\partial_{\alpha_{j}}\widetilde{l}_{j,t}(\theta,\phi(\theta))-\partial_{\alpha_{N}}\widetilde{l}_{N,t}(\theta,\phi(\theta))\big{)}\bigg{)}\right]$
=:absent:=:= : $I_{NT,1}(\theta)+I_{NT,2}(\theta).$ (A.5)
We now show that $I_{NT,1}(\theta)=O(1)$ uniformly for $\theta.$ By Assumptions 1 and 1, applying Davydov’s inequality gives that for some $\eta>2,$
$\sup_{\theta\in\Theta}|I_{NT,1}(\theta)|$
$=$ $\sup_{\theta\in\Theta}\frac{1}{NT}\sum_{i=1}^{N-1}|h_{i,i}^{\ast\breve{\alpha}}(\theta)|$
$\hskip 42.67912pt\cdot\bigg{|}\sum_{t=1}^{T}\sum_{s=1}^{T}\mathbb{E}\big{[}\partial_{\alpha_{i}}\widetilde{l}_{i,t}(\theta,\phi(\theta))\partial_{\alpha_{i}}\widetilde{l}_{i,s}(\theta,\phi(\theta))\big{]}+\sum_{t=1}^{T}\sum_{s=1}^{T}\mathbb{E}\big{[}\partial_{\alpha_{N}}\widetilde{l}_{N,t}(\theta,\phi(\theta))\partial_{\alpha_{N}}\widetilde{l}_{N,s}(\theta,\phi(\theta))\big{]}\bigg{|}$
$\leq\bigg{(}\frac{1}{N}\sum_{i=1}^{N-1}|h_{i,i}^{\ast\breve{\alpha}}(\theta)|\bigg{)}\cdot\sup_{i,t}\{\mathbb{E}[g^{\eta}(w_{it})]\}^{2/\eta}\cdot\sup_{i}\left([a_{i}(0)]^{1-2/\eta}+2\sum_{m=1}^{T-1}(1-m/T)\cdot[a_{i}(m)]^{1-2/\eta}\right)$
$=O(1)$
as $N$ and $T\rightarrow\infty,$ where the conditional cross-sectional independence leads to the first equality. Similarly, for $I_{NT,2}(\theta),$ we have
$\sup_{\theta\in\Theta}|I_{NT,2}(\theta)|$
$=$ $\frac{2}{N}\sum_{1\leq i<j\leq N-1}|h_{i,j}^{\ast\breve{\alpha}}(\theta)|\cdot\mathbb{E}\left[\bigg{(}\frac{1}{\sqrt{T}}\sum_{t=1}^{T}\partial_{\alpha_{N}}\widetilde{l}_{N,t}(\theta,\phi(\theta))\bigg{)}^{2}\right]$
$\leq$ $\bigg{(}\frac{2}{N}\sum_{1\leq i<j\leq N-1}|h_{i,j}^{\ast\breve{\alpha}}(\theta)|\bigg{)}\cdot\sup_{i,t}\{\mathbb{E}[g^{\eta}(w_{it})]\}^{2/\eta}\cdot\sup_{i}\left([a_{i}(0)]^{1-2/\eta}+2\sum_{m=1}^{T-1}(1-m/T)\cdot[a_{i}(m)]^{1-2/\eta}\right)$
$=$ $O(1),$
noting that $\sup_{i\neq j}h_{i,j}^{\ast\breve{\alpha}}(\theta)=O(1/N)$ uniformly in $\theta.$ Thus, from (A.5), we obtain the first result in (A.3). Similarly, we can derive the other result in (A.3). Now, combining (A.1), (A.2), and (A.3) yields that
$\mathbb{E}\big{\{}s^{\prime}(\theta)[-{H}(\theta)]^{-1}s(\theta)\big{\}}=$ $-\mathbb{E}\big{\{}s_{\breve{\alpha}}^{\prime}(\theta)H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)s_{\breve{\alpha}}(\theta)\big{\}}-\mathbb{E}\big{\{}s_{\breve{\gamma}}^{\prime}(\theta)H_{\breve{\gamma}\breve{\gamma}^{\prime}}^{\ast}(\theta)s_{\breve{\gamma}}(\theta)\big{\}}$
$-\mathbb{E}\left\{\big{[}s_{\breve{\alpha}}^{\prime}(\theta)H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)s_{\breve{\alpha}}(\theta)+s_{\breve{\gamma}}^{\prime}\left(\theta\right)H_{\breve{\gamma}\breve{\gamma}^{\prime}}^{\ast}\left(\theta\right)s_{\breve{\gamma}}(\theta)\big{]}\cdot o_{\mathbb{P}}(1)\right\}.$
Note that, by Assumption 1, for $S\in\{0,1,2\},$
$\sup_{i,t}\mathbb{E}\left(|\partial_{\phi_{1}\ldots\phi_{S}}l_{it}\left(\theta,\phi\right)|\mathds{1}\{|\partial_{\phi_{1}\ldots\phi_{S}}l_{it}\left(\theta,\phi\right)|>a\}\right)\rightarrow 0$
as $a\rightarrow\infty,$ which shows that $\{\partial_{\phi_{1}\ldots\phi_{S}}l_{it}\left(\theta,\phi\right)\}_{i,t}$ are uniformly integrable. By the $L_{1}$ Convergence Theorem, we hence have
$\sqrt{NT}\mathbb{E}\left\{\big{[}s_{\breve{\alpha}}^{\prime}(\theta)H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)s_{\breve{\alpha}}(\theta)+s_{\breve{\gamma}}^{\prime}\left(\theta\right)H_{\breve{\gamma}\breve{\gamma}^{\prime}}^{\ast}\left(\theta\right)s_{\breve{\gamma}}(\theta)\big{]}\cdot o_{\mathbb{P}}(1)\right\}\rightarrow 0,$
as $N,T\rightarrow\infty.$ This completes the proof of Lemma 1.
Lemma A.1
Under Assumption 1, $r^{s}(\theta,a)$ defined by (2.3) satisfies
$\sup_{\theta\in\Theta}\sup_{a}\left\|r^{s}(\theta,a)\right\|_{\max}=O_{\mathbb{P}}\left(1/\left(NT\right)\right),$
as $N$ and $T\rightarrow\infty$, where $a=\left(a_{1},\ldots,a_{\left(N+T-2\right)K}\right)^{\prime}$ for each component $a_{j}\in(0,1)$.
Proof. In the proof, we retain the notation and setup from Lemma 1. From equations (2.3) and (2.5), it is evident that the dominant term in the expansion of $\widehat{\psi}(\theta)-\psi\left(\theta\right)$ is $[-{H}(\theta)]^{-1}s(\theta).$ Specifically,
$[-{H}(\theta)]^{-1}s(\theta)$ $=[-\mathbb{E}{H}(\theta)]^{-1}[{s}(\theta)-\mathbb{E}{s}(\theta)](1+o_{\mathbb{P}}(1))$
$=-\begin{pmatrix}H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\alpha}}(\theta)+H_{\breve{\alpha}\breve{\gamma}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\gamma}}(\theta)\\ H_{\breve{\gamma}\breve{\alpha}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\alpha}}(\theta)+H_{\breve{\gamma}\breve{\gamma}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\gamma}}(\theta)\end{pmatrix}+[-\mathbb{E}{H}(\theta)]^{-1}[{s}(\theta)-\mathbb{E}{s}(\theta)]\cdot o_{\mathbb{P}}(1)$ (A.6)
where $o_{\mathbb{P}}(1)$ is uniform in both $i$ and $t.$ Let $u_{i}^{\breve{\alpha}}(\theta)$ denote the $i$-th element of $H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\alpha}}(\theta)$ for $i=1,\ldots,N-1.$ It follows that
$N\sqrt{T}u_{i}^{\breve{\alpha}}(\theta)=N\sum_{j=1}^{N-1}h_{i,j}^{\ast\breve{\alpha}}(\theta)\bigg{(}\frac{1}{\sqrt{T}}\sum_{t=1}^{T}\big{[}\partial_{\alpha_{j}}\widetilde{l}_{j,t}(\theta,\phi(\theta))-\partial_{\alpha_{N}}\widetilde{l}_{N,t}(\theta,\phi(\theta))\big{]}\bigg{)}.$ (A.7)
Using the result from Miller (1981), we know that $\sup_{i\geq 1}h_{i,i}^{\ast\breve{\alpha}}(\theta)=O(1)$ and $\sup_{i\neq j}h_{i,j}^{\ast\breve{\alpha}}(\theta)=O(1/N)$ uniformly in $\theta,$ and it also implies that $\sum_{j=1}^{N-1}h_{i,j}^{\ast\breve{\alpha}}(\theta)=O(1/N)$ uniformly in both $i$ and $\theta.$ Since $\frac{1}{\sqrt{T}}\sum_{t=1}^{T}\partial_{\alpha_{i}}\widetilde{l}_{i,t}(\theta,\phi(\theta))=O_{\mathbb{P}}(1)$ uniformly over $i$ by Assumptions 1 and 1, we obtain
$N\sum_{j=1}^{N-1}h_{i,j}^{\ast\breve{\alpha}}(\theta)\bigg{(}\frac{1}{\sqrt{T}}\sum_{t=1}^{T}\partial_{\alpha_{j}}\widetilde{l}_{j,t}(\theta,\phi(\theta))\bigg{)}=O_{\mathbb{P}}(1)$
and
$N\sum_{j=1}^{N-1}h_{i,j}^{\ast\breve{\alpha}}(\theta)\cdot\bigg{(}\frac{1}{\sqrt{T}}\sum_{t=1}^{T}\partial_{\alpha_{N}}\widetilde{l}_{N,t}(\theta,\phi(\theta))\bigg{)}=O_{\mathbb{P}}(1).$
Therefore, $N\sqrt{T}u_{i}^{\breve{\alpha}}(\theta)=O_{\mathbb{P}}(1)$ uniformly for $i.$ Let $u_{i}^{\breve{\alpha}\breve{\gamma}}(\theta)$ denote the $i$-th element of $H_{\breve{\alpha}\breve{\gamma}^{\prime}}^{\ast}\left(\theta\right)\widetilde{s}_{\breve{\gamma}}(\theta)$ for $i=1,\ldots,N-1.$ Given that $\sup_{\theta\in\Theta}\|H_{\breve{\alpha}\breve{\gamma}^{\prime}}^{\ast}(\theta)\|_{\max}=O(1),$ we have
$\sqrt{N}Tu_{i}^{\breve{\alpha}\breve{\gamma}}(\theta)=\sum_{t=1}^{T-1}h_{i,t}^{\ast\breve{\alpha}\breve{\gamma}}(\theta)\bigg{(}\frac{1}{\sqrt{N}}\sum_{j=1}^{N}\big{[}\partial_{\gamma_{t}}\widetilde{l}_{j,t}(\theta,\phi(\theta))-\partial_{\gamma_{T}}\widetilde{l}_{j,T}(\theta,\phi(\theta))\big{]}\bigg{)}.$ (A.8)
Since $\sum_{t=1}^{T-1}h_{i,t}^{\ast\breve{\alpha}\breve{\gamma}}(\theta)=O(1),$ which follows from $\mathbb{E}[\partial_{\breve{\alpha}\breve{\gamma}^{\prime}}l^{*}(\theta,\psi)|_{\psi=\psi(\theta)}]=\partial_{\breve{\gamma}^{\prime}}\mathbb{E}[\partial_{\breve{\alpha}}l^{*}(\theta,\psi)|_{\psi=\psi(\theta)}]=0,$ it holds that
$\sum_{t=1}^{T-1}h_{i,t}^{\ast\breve{\alpha}\breve{\gamma}}(\theta)\bigg{(}\frac{1}{\sqrt{N}}\sum_{j=1}^{N}\partial_{\gamma_{t}}\widetilde{l}_{j,t}(\theta,\phi(\theta))\bigg{)}=O_{\mathbb{P}}(1),$
and
$\bigg{(}\sum_{t=1}^{T-1}h_{i,t}^{\ast\breve{\alpha}\breve{\gamma}}(\theta)\bigg{)}\bigg{(}\frac{1}{\sqrt{N}}\sum_{j=1}^{N}\partial_{\gamma_{T}}\widetilde{l}_{j,T}(\theta,\phi(\theta))\bigg{)}=O_{\mathbb{P}}(1),$
uniformly in $\theta.$ From (A.7) and (A.8), we conclude that
$\sup_{\theta\in\Theta}\|H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\alpha}}(\theta)\|_{\max}=O_{\mathbb{P}}\big{(}1/(N\sqrt{T})\big{)},\quad\sup_{\theta\in\Theta}\|H_{\breve{\alpha}\breve{\gamma}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\gamma}}(\theta)\|_{\max}=O_{\mathbb{P}}\big{(}1/(\sqrt{N}T)\big{)}.$ (A.9)
Similarly, we can show that
$\sup_{\theta\in\Theta}\|H_{\breve{\gamma}\breve{\gamma}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\gamma}}(\theta)\|_{\max}=O_{\mathbb{P}}\big{(}1/(\sqrt{N}T)\big{)},\quad\sup_{\theta\in\Theta}\|H_{\breve{\gamma}\breve{\alpha}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\alpha}}(\theta)\|_{\max}=O_{\mathbb{P}}\big{(}1/(N\sqrt{T})\big{)}.$ (A.10)
Thus, combining (A.6), (A.9), and (A.10), we have
$\left\|[-{H}(\theta)]^{-1}s(\theta)+[\mathbb{E}{H}(\theta)]^{-1}[{s}(\theta)-\mathbb{E}{s}(\theta)]\right\|_{\max}=o_{\mathbb{P}}\big{(}1/(NT)^{3/4}\big{)},$
where $o_{\mathbb{P}}(\cdot)$ is uniform in $\theta.$ Noting that
$\widehat{\psi}(\theta)-\psi(\theta)=[-{H}\left(\theta\right)]^{-1}s(\theta)(1+o_{\mathbb{P}}(1)),$
we further have
$\left\|\widehat{\psi}(\theta)-\psi(\theta)+[\mathbb{E}{H}(\theta)]^{-1}[{s}(\theta)-\mathbb{E}{s}(\theta)]\right\|_{\max}=o_{\mathbb{P}}\big{(}1/(NT)^{3/4}\big{)}.$ (A.11)
From (A.11), we conclude that $\sup_{\theta\in\Theta}\|G_{\psi_{j}}\left(\theta,a_{j}\right)-\partial_{\psi_{j}}H(\theta)\|_{\max}=o_{\mathbb{P}}(1).$ Without loss of generality, let $\psi_{j}=\alpha_{i}$ for $i=1,\ldots,N-1.$ We can express $\partial_{\alpha_{i}}H(\theta)$ as
$\partial_{\alpha_{i}}H(\theta)=\begin{pmatrix}\partial_{\alpha_{i}}{H}_{\breve{\alpha}\breve{\alpha}^{\prime}}\left(\theta\right),&\partial_{\alpha_{i}}{H}_{\breve{\alpha}\breve{\gamma}^{\prime}}(\theta)\\ \partial_{\alpha_{i}}{H}_{\breve{\gamma}\breve{\alpha}^{\prime}}\left(\theta\right),&\partial_{\alpha_{i}}{H}_{\breve{\gamma}\breve{\gamma}^{\prime}}(\theta)\end{pmatrix}.$
In particular, let $\text{diag}\{a_{1},\ldots,a_{n}\}$ denote a diagonal matrix with elements $a_{1},\ldots,a_{n}$ on its diagonal. We then have
$N\cdot\partial_{\alpha_{i}}{H}_{\breve{\alpha}\breve{\alpha}^{\prime}}\left(\theta\right)=$ $\frac{1}{T}\sum_{t=1}^{T}\text{diag}\big{\{}0,\ldots,0,\partial_{\alpha_{i}\alpha_{i}\alpha_{i}}l_{i,t}(\theta,\phi)|_{\phi=\phi(\theta)},0,\ldots,0\big{\}}$
$\quad-\bigg{(}\frac{1}{T}\sum_{t=1}^{T}\partial_{\alpha_{N}\alpha_{N}\alpha_{N}}l_{N,t}(\theta,\phi)|_{\phi=\phi(\theta)}\bigg{)}\,\iota_{N-1}\iota_{N-1}^{\prime},$
$NT\cdot\partial_{\alpha_{i}}{H}_{\breve{\gamma}\breve{\gamma}^{\prime}}\left(\theta\right)=$ $\text{diag}\left\{\partial_{\gamma_{1}\gamma_{1}\alpha_{i}}l_{i,1}(\theta,\phi)|_{\phi=\phi(\theta)},\ldots,\partial_{\gamma_{T-1}\gamma_{T-1}\alpha_{i}}l_{i,T-1}(\theta,\phi)|_{\phi=\phi(\theta)}\right\}$
$\quad+\partial_{\gamma_{T}\gamma_{T}\alpha_{i}}l_{i,T}(\theta,\phi)|_{\phi=\phi(\theta)}\,\iota_{N-1}\iota_{N-1}^{\prime},$
$NT\cdot\partial_{\alpha_{i}}{H}_{\breve{\alpha}\breve{\gamma}^{\prime}}\left(\theta\right)=$ $\big{(}\partial_{\alpha_{i}\alpha_{i}\gamma_{1}}l_{i,1}(\theta,\phi)|_{\phi=\phi(\theta)},\ldots,\partial_{\alpha_{i}\alpha_{i}\gamma_{T-1}}l_{i,T-1}(\theta,\phi)|_{\phi=\phi(\theta)}\big{)}\otimes e_{i}$
$\quad+\big{(}\partial_{\alpha_{N}\alpha_{N}\gamma_{1}}l_{N,1}(\theta,\phi)|_{\phi=\phi(\theta)},\ldots,\partial_{\alpha_{N}\alpha_{N}\gamma_{T-1}}l_{N,T-1}(\theta,\phi)|_{\phi=\phi(\theta)}\big{)}\otimes\iota_{N-1},$
where $e_{i}$ is an $(N-1)\times 1$ vector with the $i$-th component equal to 1 and all other elements equal to 0. Using (A.11), for each $j=1,\ldots,N-1,$ we write
$\sup_{a_{j}\in(0,1)}r_{\psi_{j}}^{s}(\theta,a_{j})$
$=$ $\sup_{a_{j}\in(0,1)}\frac{1}{2}\big{(}\widehat{\psi}(\theta)-\psi(\theta)\big{)}^{\prime}G_{\psi_{j}}(\theta,a_{j})\big{(}\widehat{\psi}(\theta)-\psi\left(\theta\right)\big{)}$
$=$ $\frac{1}{2}\begin{pmatrix}H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\alpha}}(\theta)+H_{\breve{\alpha}\breve{\gamma}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\gamma}}(\theta)\\ H_{\breve{\gamma}\breve{\alpha}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\alpha}}(\theta)+H_{\breve{\gamma}\breve{\gamma}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\gamma}}(\theta)\end{pmatrix}^{\prime}\partial_{\alpha_{i}}H(\theta)\begin{pmatrix}H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\alpha}}(\theta)+H_{\breve{\alpha}\breve{\gamma}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\gamma}}(\theta)\\ H_{\breve{\gamma}\breve{\alpha}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\alpha}}(\theta)+H_{\breve{\gamma}\breve{\gamma}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\gamma}}(\theta)\end{pmatrix}\big{(}1+o_{\mathbb{P}}(1)\big{)}$
$=$ $\frac{1}{2}\left\{\big{(}H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\alpha}}(\theta)+H_{\breve{\alpha}\breve{\gamma}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\gamma}}(\theta)\big{)}^{\prime}\partial_{\alpha_{i}}{H}_{\breve{\alpha}\breve{\alpha}^{\prime}}(\theta)\big{(}H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\alpha}}(\theta)+H_{\breve{\alpha}\breve{\gamma}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\gamma}}(\theta)\big{)}\right.$
$\qquad+\big{(}H_{\breve{\gamma}\breve{\alpha}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\alpha}}(\theta)+H_{\breve{\gamma}\breve{\gamma}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\gamma}}(\theta)\big{)}^{\prime}\partial_{\alpha_{i}}{H}_{\breve{\gamma}\breve{\gamma}^{\prime}}(\theta)\big{(}H_{\breve{\gamma}\breve{\alpha}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\alpha}}(\theta)+H_{\breve{\gamma}\breve{\gamma}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\gamma}}(\theta)\big{)}$
$\qquad\left.+2\big{(}H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\alpha}}(\theta)+H_{\breve{\alpha}\breve{\gamma}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\gamma}}(\theta)\big{)}^{\prime}\partial_{\alpha_{i}}{H}_{\breve{\alpha}\breve{\gamma}^{\prime}}(\theta)\big{(}H_{\breve{\gamma}\breve{\alpha}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\alpha}}(\theta)+H_{\breve{\gamma}\breve{\gamma}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\gamma}}(\theta)\big{)}\right\}\big{(}1+o_{\mathbb{P}}(1)\big{)}$
=:absent:=:= : $\frac{1}{2}\left\{C_{NT,1}(\theta)+C_{NT,2}(\theta)+C_{NT,3}(\theta)\right\}\big{(}1+o_{\mathbb{P}}(1)\big{)}.$ (A.12)
Let $\widetilde{u}_{i}^{\breve{\alpha}}(\theta)$ denote the $i$-th element of $H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\alpha}}(\theta)+H_{\breve{\alpha}\breve{\gamma}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\gamma}}(\theta)$ for $i=1,\ldots,N-1.$ Using (A.7) and (A.8), we can show that
$C_{NT,1}(\theta)$
$=$ $\frac{1}{N}[\widetilde{u}_{i}^{\breve{\alpha}}(\theta)]^{2}\bigg{(}\frac{1}{T}\sum_{t=1}^{T}\partial_{\alpha_{i}\alpha_{i}\alpha_{i}}l_{i,t}(\theta,\phi)|_{\phi=\phi(\theta)}\bigg{)}-\frac{1}{N}\bigg{(}\sum_{i=1}^{N-1}\widetilde{u}_{i}^{\breve{\alpha}}(\theta)\bigg{)}^{2}\bigg{(}\frac{1}{T}\sum_{t=1}^{T}\partial_{\alpha_{N}\alpha_{N}\alpha_{N}}l_{N,t}(\theta,\phi)|_{\phi=\phi(\theta)}\bigg{)}$
$=$ $O_{\mathbb{P}}\big{(}1/(NT)\big{)},$
uniformly in $\theta.$ By similar arguments, we also show that
$\sup_{\theta\in\Theta}C_{NT,2}(\theta)=O_{\mathbb{P}}\big{(}1/(NT)^{3/2}\big{)},\quad\sup_{\theta\in\Theta}C_{NT,3}(\theta)=O_{\mathbb{P}}\big{(}1/(NT)^{3/2}\big{)}.$
Therefore, when $\psi_{j}=\alpha_{i},$ we have
$\sup_{\theta\in\Theta}\sup_{a_{j}\in(0,1)}r_{\psi_{j}}^{s}(\theta,a_{j})$
$=$ $-\frac{1}{2N}\bigg{(}\sum_{i=1}^{N-1}\widetilde{u}_{i}^{\breve{\alpha}}(\theta)\bigg{)}^{2}\bigg{(}\frac{1}{T}\sum_{t=1}^{T}\partial_{\alpha_{N}\alpha_{N}\alpha_{N}}l_{N,t}(\theta,\phi)|_{\phi=\phi(\theta)}\bigg{)}+o_{\mathbb{P}}\big{(}1/(NT)\big{)}$ (A.13)
$=$ $O_{\mathbb{P}}\big{(}1/(NT)\big{)}.$
Similarly, when $\psi_{j}=\gamma_{t},$ we have
$\sup_{\theta\in\Theta}\sup_{a_{j}\in(0,1)}r_{\psi_{j}}^{s}(\theta,a_{j})$
$=$ $-\frac{1}{2T}\bigg{(}\sum_{t=1}^{T-1}\widetilde{u}_{t}^{\breve{\gamma}}(\theta)\bigg{)}^{2}\bigg{(}\frac{1}{N}\sum_{i=1}^{N}\partial_{\gamma_{T}\gamma_{T}\gamma_{T}}l_{i,T}(\theta,\phi)|_{\phi=\phi(\theta)}\bigg{)}+o_{\mathbb{P}}\big{(}1/(NT)\big{)}$ (A.14)
$=$ $O_{\mathbb{P}}\big{(}1/(NT)\big{)},$
where $\widetilde{u}_{t}^{\breve{\gamma}}(\theta)$ denotes the $t$-th element of $H_{\breve{\gamma}\breve{\alpha}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\alpha}}(\theta)+H_{\breve{\gamma}\breve{\gamma}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\gamma}}(\theta)$ for $t=1,\ldots,T-1.$ This shows the desired result.
Lemma A.2
Under Assumption 1, $r(\theta)$ defined by (2.7) satisfies
$\sup_{\theta\in\Theta}\mathbb{E}[r(\theta)]=O\left(1/\left(NT\right)\right),$
as $N$ and $T\rightarrow\infty$.
Proof. In the proof, we retain the notation and setup from Lemmas 1 and A.1. From Lemma A.1, we know that $NT\cdot r^{s}\left(\theta,a\right)$ is comparable to $(NT)^{3/4}\cdot s(\theta).$ Following similar arguments as in Lemma 1, we can show that
$\mathbb{E}\big{[}(r^{s}(\theta,a))^{\prime}[{H}\left(\theta\right)]^{-1}r^{s}(\theta,a)\big{]}=O\big{(}1/(NT)\big{)}.$
Moreover, from (A.11), (A.13), and (A.14), we have
$\big{(}\widehat{\psi}(\theta)-\psi(\theta)\big{)}^{\prime}r^{s}(\theta,a_{0}\,\iota_{N+T-2})$
$=$ $\frac{1}{2N}\bigg{(}\sum_{i=1}^{N-1}\widetilde{u}_{i}^{\breve{\alpha}}(\theta)\bigg{)}^{2}\bigg{(}\frac{1}{T}\sum_{t=1}^{T}\partial_{\alpha_{N}\alpha_{N}\alpha_{N}}l_{N,t}(\theta,\phi)|_{\phi=\phi(\theta)}\bigg{)}$
$+\frac{1}{2T}\bigg{(}\sum_{t=1}^{T-1}\widetilde{u}_{t}^{\breve{\gamma}}(\theta)\bigg{)}^{2}\bigg{(}\frac{1}{N}\sum_{i=1}^{N}\partial_{\gamma_{T}\gamma_{T}\gamma_{T}}l_{i,T}(\theta,\phi)|_{\phi=\phi(\theta)}\bigg{)}$
$\quad+\bigg{(}\sum_{i=1}^{N-1}\widetilde{u}_{i}^{\breve{\alpha}}(\theta)+\sum_{t=1}^{T-1}\widetilde{u}_{t}^{\breve{\gamma}}(\theta)\bigg{)}\cdot o_{\mathbb{P}}\big{(}1/(NT)\big{)}$
$=$ $O_{\mathbb{P}}\big{(}1/(NT)^{5/4}\big{)}.$ (A.15)
By Assumption 1 and applying the $L_{1}$ Convergence Theorem to (A.15), we find that
$\sup_{\theta\in\Theta}\mathbb{E}\big{[}\big{(}\widehat{\psi}(\theta)-\psi(\theta)\big{)}^{\prime}r^{s}\left(\theta,a_{0}\,\iota_{N+T-2}\right)\big{]}=o\big{(}1/(NT)\big{)}.$
Thus,
$\mathbb{E}\big{[}r(\theta)\big{]}=\frac{1}{2}\mathbb{E}\big{[}(r^{s}\left(\theta,a\right))^{\prime}[{H}(\theta)]^{-1}r^{s}\left(\theta,a\right)\big{]}+\frac{1}{3}\mathbb{E}\big{[}\big{(}\widehat{\psi}(\theta)-\psi(\theta)\big{)}^{\prime}r^{s}\left(\theta,a_{0}\,\iota_{N+T-2}\right)\big{]}=O\big{(}1/(NT)\big{)}$
uniformly in $\theta.$
A.2 Proof of Theorem 1
From (2.7), we have
$\mathbb{E}[l(\theta)]=\mathbb{E}[\widehat{l}(\theta)]+\frac{1}{2}\mathbb{E}\big{\{}s^{\prime}(\theta)[H(\theta)]^{-1}s(\theta)\big{\}}-\mathbb{E}[r(\theta)].$
Thus, the desired result follows directly from Lemmas 1 and A.2.
A.3 Proof of Theorem 2
We retain the notation introduced in Lemmas 1-A.2. First, we show that
$\widehat{B}_{\breve{\alpha}}(\theta)-B_{\breve{\alpha}}(\theta)=o_{\mathbb{P}}\big{(}1/\sqrt{NT}\big{)},\quad\text{and}\quad\widehat{B}_{\breve{\gamma}}(\theta)-B_{\breve{\gamma}}(\theta)=o_{\mathbb{P}}\big{(}1/\sqrt{NT}\big{)}.$ (A.16)
Note that
$\widehat{B}_{\breve{\alpha}}(\theta)-B_{\breve{\alpha}}(\theta)=\frac{1}{2}\text{trace}\big{\{}\widehat{H}_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}\left(\theta\right)[\Sigma^{\breve{\alpha}}(\theta)-\widehat{\Sigma}^{\breve{\alpha}}_{\tau}(\theta)]\big{\}}+\frac{1}{2}\text{trace}\big{\{}[\widehat{H}_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)-H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)]\Sigma^{\breve{\alpha}}(\theta)\big{\}}.$
Define $\widetilde{s}_{i,t}^{\alpha}(\theta):=\{\partial_{\alpha_{i}}l_{i,t}(\theta,{\phi})-\mathbb{E}[\partial_{\alpha_{i}}l_{i,t}(\theta,{\phi})]\}|_{\phi={\phi}(\theta)}.$ Then we have
$N^{2}\big{[}{\Sigma}_{1}^{\breve{\alpha}}(\theta)-\widehat{\Sigma}_{1,\tau}^{\breve{\alpha}}(\theta)\big{]}$
$=$ $\frac{1}{T^{2}}\sum_{t=1}^{T}\sum_{s=1}^{T}\mathds{1}\left\{\left|t-s\right|\leq\tau\right\}$
[…]
$+\frac{1}{T^{2}}\sum_{t=1}^{T}\sum_{s=1}^{T}\mathds{1}\left\{\left|t-s\right|>\tau\right\}$
×$\qquad\times\text{diag}\big{\{}\mathbb{E}\big{[}\widetilde{s}_{1,t}^{\alpha}(\theta)\big{(}\widetilde{s}_{1,s}^{\alpha}(\theta)\big{)}^{\prime}\big{]},\ldots,\mathbb{E}\big{[}\widetilde{s}_{N-1,t}^{\alpha}(\theta)\big{(}\widetilde{s}_{N-1,s}^{\alpha}(\theta)\big{)}^{\prime}\big{]}\big{\}}.$ (A.17)
Under Assumptions 1 and 1, by Davydov’s inequality,
$\mathbb{E}\big{[}\widetilde{s}_{i,t}^{\alpha}(\theta)\big{(}\widetilde{s}_{i,s}^{\alpha}(\theta)\big{)}^{\prime}\big{]}=\mathbb{E}\big{[}\widetilde{s}_{i,t}^{\alpha}(\theta)\widetilde{s}_{i,s}^{\alpha}(\theta)\big{]}\leq 8[a_{i}(|t-s|)]^{1-2/\eta}\sup_{i,t}\{\mathbb{E}[g^{\eta}(w_{it})]\}^{2/\eta},$
where we assume without loss of generality that $K=1.$ Thus, as $T\rightarrow\infty,$
$\frac{1}{T}\sum_{t=1}^{T}\sum_{s=1}^{T}\mathds{1}\left\{\left|t-s\right|>\tau\right\}\mathbb{E}\big{[}\widetilde{s}_{i,t}^{\alpha}(\theta)\big{(}\widetilde{s}_{i,s}^{\alpha}(\theta)\big{)}^{\prime}\big{]}$
$\leq$ $8\sup_{i,t}\{\mathbb{E}[g^{\eta}(w_{it})]\}^{2/\eta}\cdot\frac{1}{T}\sum_{t=1}^{T}\sum_{s=1}^{T}\mathds{1}\left\{\left|t-s\right|>\tau\right\}[a_{i}(|t-s|)]^{1-2/\eta}$
$\leq$ $8\sup_{i,t}\{\mathbb{E}[g^{\eta}(w_{it})]\}^{2/\eta}\cdot\frac{T-\tau}{T}\sum_{m=\tau+1}^{\infty}[a_{i}(m)]^{1-2/\eta}$
$=$ $o(1)$
given that $\tau/T\to 0.$ Combining this with (A.17) implies
$\sup_{\theta\in\Theta}N^{2}\|{\Sigma}_{1}^{\breve{\alpha}}(\theta)-\widehat{\Sigma}_{1,\tau}^{\breve{\alpha}}(\theta)\|_{\max}=o_{\mathbb{P}}(1/T).$ (A.18)
From the proof of Lemma A.1, we also have $\sup_{\theta\in\Theta}\|\widehat{\psi}(\theta)-\psi(\theta)\|_{\max}=o_{\mathbb{P}}(1).$ Moreover, under the conditional strong mixing condition, each element of $T^{-1}\widehat{H}_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)$ converges in probability to its counterpart in $T^{-1}H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)$ by the law of large numbers. Therefore, from (A.18),
$\text{trace}\big{\{}\widehat{H}_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)[\Sigma_{1}^{\breve{\alpha}}(\theta)-\widehat{\Sigma}_{1,\tau}^{\breve{\alpha}}(\theta)]\big{\}}$
$=$ $\frac{T}{N^{2}}\text{trace}\big{\{}T^{-1}\widehat{H}_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)\,N^{2}[\Sigma^{\breve{\alpha}}_{1}(\theta)-\widehat{\Sigma}^{\breve{\alpha}}_{1,\tau}(\theta)]\big{\}}$
$=$ $\frac{T}{N}\cdot\frac{1}{N}\sum_{i=1}^{N-1}h_{i,i}^{\ast\breve{\alpha}}(\theta)\cdot o_{\mathbb{P}}(1/T)$
$=$ $o_{\mathbb{P}}(1/T).$
Similarly,
$\text{trace}\big{\{}\widehat{H}_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)[\Sigma_{2}^{\breve{\alpha}}(\theta)-\widehat{\Sigma}_{2,\tau}^{\breve{\alpha}}(\theta)]\big{\}}$
$=$ $\frac{T}{N^{2}}\text{trace}\big{\{}T^{-1}\widehat{H}_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)\,N^{2}[\Sigma_{2}^{\breve{\alpha}}(\theta)-\widehat{\Sigma}_{2,\tau}^{\breve{\alpha}}(\theta)]\big{\}}$
$=$ $\frac{T}{N^{2}}\cdot\sum_{i=1}^{N-1}\sum_{j=1}^{N-1}h_{i,j}^{\ast\breve{\alpha}}(\theta)\cdot o_{\mathbb{P}}(1/T)$
$=$ $o_{\mathbb{P}}(1/N^{2}),$
where the last equality follows from the fact that $\sum_{j=1}^{N-1}h_{i,j}^{\ast\breve{\alpha}}(\theta)=O(1/N)$ uniformly in both $i$ and $\theta,$ as shown in Lemma A.1. Therefore,
$\text{trace}\big{\{}\widehat{H}_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)[\Sigma^{\breve{\alpha}}(\theta)-\widehat{\Sigma}^{\breve{\alpha}}_{\tau}(\theta)]\big{\}}=o_{\mathbb{P}}\big{(}1/\sqrt{NT}\big{)}.$ (A.19)
Moreover, from Theorem 1, we also have
$\text{trace}\big{\{}[\widehat{H}_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)-H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)]\Sigma^{\breve{\alpha}}(\theta)\big{\}}=o_{\mathbb{P}}(1)\cdot\text{trace}\big{\{}H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)\Sigma^{\breve{\alpha}}(\theta)\big{\}}=o_{\mathbb{P}}(1)\cdot B_{\breve{\alpha}}(\theta)=o_{\mathbb{P}}\big{(}1/\sqrt{NT}\big{)}.$ (A.20)
Combining (A.19) and (A.20) shows the first part of (A.16). The second part of (A.16) can be shown similarly.
Noting that
$l(\theta)=\widehat{l}(\theta)+\frac{1}{2}s_{\breve{\alpha}}^{\prime}(\theta)H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)s_{\breve{\alpha}}(\theta)+\frac{1}{2}s_{\breve{\gamma}}^{\prime}(\theta)H_{\breve{\gamma}\breve{\gamma}^{\prime}}^{\ast}(\theta)s_{\breve{\gamma}}(\theta)+o_{\mathbb{P}}\big{(}1/\sqrt{NT}\big{)}$
as implied by the proof of Theorem 1, we have
$L(\theta)-l(\theta)$ $=\big{[}\widehat{B}_{\breve{\alpha}}(\theta)-B_{\breve{\alpha}}(\theta)\big{]}+\big{[}\widehat{B}_{\breve{\gamma}}(\theta)-B_{\breve{\gamma}}(\theta)\big{]}$
$\quad-\bigg{[}\frac{1}{2}s_{\breve{\alpha}}^{\prime}(\theta)H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)s_{\breve{\alpha}}(\theta)-B_{\breve{\alpha}}(\theta)\bigg{]}-\bigg{[}\frac{1}{2}s_{\breve{\gamma}}^{\prime}(\theta)H_{\breve{\gamma}\breve{\gamma}^{\prime}}^{\ast}(\theta)s_{\breve{\gamma}}(\theta)-B_{\breve{\gamma}}(\theta)\bigg{]}$
$+o_{\mathbb{P}}\big{(}1/\sqrt{NT}\big{)}.$
Moreover, from Lemma 1,
$\frac{1}{2}s_{\breve{\alpha}}^{\prime}(\theta)H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)s_{\breve{\alpha}}(\theta)-B_{\breve{\alpha}}(\theta)=o_{\mathbb{P}}\big{(}1/\sqrt{NT}\big{)},\quad\text{and}\quad\frac{1}{2}s_{\breve{\gamma}}^{\prime}\left(\theta\right)H_{\breve{\gamma}\breve{\gamma}^{\prime}}^{\ast}\left(\theta\right)s_{\breve{\gamma}}(\theta)-B_{\breve{\gamma}}(\theta)=o_{\mathbb{P}}\big{(}1/\sqrt{NT}\big{)}.$
Thus, combining this with (A.16) yields the desired result.
A.4 Proof of Corollary 1
Consider a mean value expansion of $\triangledown_{\theta}L(\widehat{\theta}_{L})$ around $\theta_{0}:$
$0=\triangledown_{\theta}L(\widehat{\theta}_{L})$ $=\triangledown_{\theta}L(\theta_{0})+\triangledown_{\theta\theta^{\prime}}L(\overline{\theta})(\widehat{\theta}_{L}-\theta_{0}),$ (A.21)
where each component of $\overline{\theta}$ lies between the corresponding components of $\widehat{\theta}_{L}$ and $\theta_{0}.$ From Theorem 2, we further have
$0=$ $\triangledown_{\theta}l(\theta_{0})+\triangledown_{\theta\theta^{\prime}}l(\overline{\theta})(\widehat{\theta}_{L}-\theta_{0})+o_{\mathbb{P}}\big{(}1/\sqrt{NT}\big{)}$ (A.22)
Under Assumption 2, $\theta_{0}$ is the unique maximizer of $\mathbb{E}[l(\theta)].$ Since $L(\theta)-\mathbb{E}[l(\theta)]=o_{\mathbb{P}}(1),$ under Assumption 2, we conclude that
$\widehat{\theta}_{L}\overset{\mathbb{P}}{\longrightarrow}\theta_{0}.$
This also implies $\overline{\theta}\overset{\mathbb{P}}{\longrightarrow}\theta_{0}.$ By Assumption 2, it follows that
$\triangledown_{\theta\theta^{\prime}}l(\overline{\theta})-\mathbb{E}[\triangledown_{\theta\theta^{\prime}}l(\theta_{0})]=o_{\mathbb{P}}(1).$
Thus, using Assumption 2 and Slutsky’s theorem, we obtain from (A.22) that
$\sqrt{NT}(\widehat{\theta}_{L}-\theta_{0})$ $=-\big{(}\mathbb{E}[\triangledown_{\theta\theta^{\prime}}l(\theta_{0})]\big{)}^{-1}\sqrt{NT}\triangledown_{\theta}l(\theta_{0})+o_{\mathbb{P}}(1).$ (A.23)
Given the conditional cross-sectional independence and the conditional strong mixing condition (Assumption 1), (A.23) shows that
$\sqrt{NT}(\widehat{\theta}_{L}-\theta_{0})\overset{\mathbb{D}}{\longrightarrow}\mathcal{N}\left(0,\Sigma^{-1}\Omega\Sigma^{-1}\right)$ (A.24)
where $\mathbb{E}[\triangledown_{\theta}l(\theta_{0})]=0,$ and
$\Omega=\lim_{N,T\rightarrow\infty}(NT)\mathbb{E}[\triangledown_{\theta}l(\theta_{0})\triangledown_{\theta^{\prime}}l(\theta_{0})].$
Under Assumption 2, the information matrix equality ($\Sigma^{-1}=\Omega$) holds, where the asymptotic variance in (A.24) simplifies to $\Sigma$.
A.5 Proof of Lemma A.3
Lemma A.3
Under Assumptions 1 and 2, and under the null hypothesis $H_{0}:R(\theta_{0})=0$, as $N$ and $T\to\infty,$
$\widehat{\theta}_{R}-\widehat{\theta}_{L}=\Xi(\theta_{0})\triangledown_{\theta}l(\theta_{0})+o_{\mathbb{P}}\big{(}1/\sqrt{NT}\big{)},$
where
$\Xi(\theta_{0}):=\big{(}\mathbb{E[}\triangledown_{\theta\theta^{\prime}}l(\theta_{0})]\big{)}^{-1}[J(\theta_{0})]^{\prime}\Big{(}J(\theta_{0})\big{(}\mathbb{E[}\triangledown_{\theta\theta^{\prime}}l(\theta_{0})]\big{)}^{-1}[J(\theta_{0})]^{\prime}\Big{)}^{-1}J(\theta_{0})\big{(}\mathbb{E[}\triangledown_{\theta\theta^{\prime}}l(\theta_{0})]\big{)}^{-1}.$
Proof. Consider the constrained maximization problem:
$\big{(}\widehat{\theta}_{R},\widehat{\lambda}\big{)}=\arg\max_{\theta\in\Theta,\lambda\in\mathbb{R}}\big{[}L(\theta)+\lambda^{\prime}R(\theta)\big{]}$
where $\lambda$ is the Lagrangian multiplier. The first-order conditions for this constrained maximization are
$\sqrt{NT}\triangledown_{\theta}L(\widehat{\theta}_{R})+[J(\widehat{\theta}_{R})]^{\prime}\sqrt{NT}\widehat{\lambda}$ $=0,$ (A.25)
$\sqrt{NT}R(\widehat{\theta}_{R})$ $=0.$ (A.26)
Following a similar proof as in Corollary 1, we have $\sqrt{NT}\big{(}\widehat{\theta}_{R}-\theta_{0}\big{)}=O_{\mathbb{P}}\left(1\right).$ Furthermore, $\sqrt{NT}\widehat{\lambda}=O_{\mathbb{P}}(1)$ as implied by (A.25) since $\sqrt{NT}\triangledown_{\theta}L(\theta_{0})=O_{\mathbb{P}}(1)$ and $J(\theta_{0})$ is of full row rank. By Theorem 2, we express the mean value expansion of $\triangledown_{\theta}L(\widehat{\theta}_{R})$ around $\theta_{0}$ as
$\triangledown_{\theta}L(\widehat{\theta}_{R})$ $=\triangledown_{\theta}L(\theta_{0})+\triangledown_{\theta\theta^{\prime}}L(\overline{\theta})\big{(}\widehat{\theta}_{R}-\theta_{0}\big{)}$
$=\triangledown_{\theta}l(\theta_{0})+\mathbb{E}[\triangledown_{\theta\theta^{\prime}}l(\theta_{0})]\big{(}\widehat{\theta}_{R}-\theta_{0}\big{)}+o_{\mathbb{P}}\big{(}1/\sqrt{NT}\big{)},$
where each component of $\overline{\theta}$ lies between the corresponding components of $\widehat{\theta}_{R}$ and $\theta_{0}.$ Combining this with (A.25) yields that
$\mathbb{E}[\triangledown_{\theta\theta^{\prime}}l(\theta_{0})]\sqrt{NT}\big{(}\widehat{\theta}_{R}-\theta_{0}\big{)}+[J(\theta_{0})]^{\prime}\sqrt{NT}\widehat{\lambda}=-\sqrt{NT}\triangledown_{\theta}l(\theta_{0})+o_{\mathbb{P}}(1).$ (A.27)
Under $H_{0},$ from (A.26), we also obtain
00 0 $=\sqrt{NT}R(\widehat{\theta}_{R})$
$=\sqrt{NT}R(\theta_{0})+J(\theta_{0})\sqrt{NT}\big{(}\widehat{\theta}_{R}-\theta_{0}\big{)}+o_{\mathbb{P}}(1)$
$=J(\theta_{0})\sqrt{NT}\big{(}\widehat{\theta}_{R}-\theta_{0}\big{)}+o_{\mathbb{P}}(1).$ (A.28)
Now, combining (A.27) and (A.28), we have
$\begin{pmatrix}\mathbb{E}[\triangledown_{\theta\theta^{\prime}}l(\theta_{0})]&\left[J(\theta_{0})\right]^{\prime}\\ J(\theta_{0})&0\end{pmatrix}\begin{pmatrix}\sqrt{NT}(\widehat{\theta}_{R}-\theta_{0})\\ \sqrt{NT}\widehat{\lambda}\end{pmatrix}=\begin{pmatrix}-\sqrt{NT}\triangledown_{\theta}l(\theta_{0})\\ 0\end{pmatrix}+o_{\mathbb{P}}(1)$
which has the solution
$\sqrt{NT}\big{(}\widehat{\theta}_{R}-\theta_{0}\big{)}=-\left(\big{(}\mathbb{E}[\triangledown_{\theta\theta^{\prime}}l(\theta_{0})]\big{)}^{-1}-\Xi\left(\theta_{0}\right)\right)\sqrt{NT}\triangledown_{\theta}l(\theta_{0})+o_{\mathbb{P}}\left(1\right).$ (A.29)
The conclusion follows by subtracting (A.23) in Corollary 1 from both sides of (A.29).
A.6 Proof of Corollary 2
Let $\widehat{\xi}:=-NT\triangledown_{\theta^{\prime}}l(\theta_{0})\Xi\left(\theta_{0}\right)\triangledown_{\theta}l(\theta_{0}).$ We first show that
$\widehat{\xi}_{LR}=\widehat{\xi}+o_{\mathbb{P}}(1),\qquad\widehat{\xi}_{LM}=\widehat{\xi}+o_{\mathbb{P}}(1),\qquad\widehat{\xi}_{Wald}=\widehat{\xi}+o_{\mathbb{P}}(1).$ (A.30)
A mean value expansion of $L(\widehat{\theta}_{R})$ around $\widehat{\theta}_{L}$ gives
${L}(\widehat{\theta}_{R})={L}(\widehat{\theta}_{L})+\triangledown_{\theta^{\prime}}{L}(\widehat{\theta}_{L})\big{(}\widehat{\theta}_{R}-\widehat{\theta}_{L}\big{)}+\frac{1}{2}\big{(}\widehat{\theta}_{R}-\widehat{\theta}_{L}\big{)}^{\prime}\triangledown_{\theta\theta^{\prime}}L(\theta^{\ast})\big{(}\widehat{\theta}_{R}-\widehat{\theta}_{L}\big{)}$
where each component of $\theta^{\ast}$ lies between the corresponding components of $\widehat{\theta}_{R}$ and $\widehat{\theta}_{L}$. Noting that $\triangledown_{\theta}{L}(\widehat{\theta}_{L})=0,$ we obtain
$\widehat{\xi}_{LR}=-NT\big{(}\widehat{\theta}_{R}-\widehat{\theta}_{L}\big{)}^{\prime}\triangledown_{\theta\theta^{\prime}}L(\theta^{\ast})\big{(}\widehat{\theta}_{R}-\widehat{\theta}_{L}\big{)}.$
Since $\triangledown_{\theta\theta^{\prime}}L(\theta^{\ast})-\mathbb{E}[\triangledown_{\theta\theta^{\prime}}L(\theta_{0})]=o_{\mathbb{P}}(1),$ by Lemma A.3, we further conclude that
$\widehat{\xi}_{LR}=\widehat{\xi}+o_{\mathbb{P}}(1).$
For $\widehat{\xi}_{LM},$ note that
$\triangledown_{\theta}L(\widehat{\theta}_{R})$ $=\triangledown_{\theta}L(\widehat{\theta}_{L})+\triangledown_{\theta\theta^{\prime}}L(\dot{\theta})\big{(}\widehat{\theta}_{R}-\widehat{\theta}_{L}\big{)}$
$=\mathbb{E}[\triangledown_{\theta\theta^{\prime}}L(\theta_{0})]\big{(}\widehat{\theta}_{R}-\widehat{\theta}_{L}\big{)}+o_{\mathbb{P}}\big{(}1/\sqrt{NT}\big{)},$
where each component of $\dot{\theta}$ lies between the corresponding components of $\widehat{\theta}_{R}$ and $\widehat{\theta}_{L}.$ Thus,
$\widehat{\xi}_{LM}$ $=-NT\big{(}\widehat{\theta}_{R}-\widehat{\theta}_{L}\big{)}^{\prime}\mathbb{E}[\triangledown_{\theta\theta^{\prime}}L(\theta_{0})][\triangledown_{\theta\theta^{\prime}}L(\widehat{\theta}_{R})]^{-1}\mathbb{E}[\triangledown_{\theta\theta^{\prime}}L(\theta_{0})]\big{(}\widehat{\theta}_{R}-\widehat{\theta}_{L}\big{)}+o_{\mathbb{P}}\left(1\right)$
$=-NT\big{(}\widehat{\theta}_{R}-\widehat{\theta}_{L}\big{)}^{\prime}\mathbb{E}[\triangledown_{\theta\theta^{\prime}}L(\theta_{0})]\big{(}\widehat{\theta}_{R}-\widehat{\theta}_{L}\big{)}+o_{\mathbb{P}}\left(1\right)$
$=\widehat{\xi}+o_{\mathbb{P}}\left(1\right).$
For $\widehat{\xi}_{Wald}$, a mean value expansion of $R(\widehat{\theta}_{L})$ gives
$R(\widehat{\theta}_{L})=R(\widehat{\theta}_{R})+J(\ddot{\theta})(\widehat{\theta}_{L}-\widehat{\theta}_{R})=J(\ddot{\theta})(\widehat{\theta}_{L}-\widehat{\theta}_{R})$
[…]
$\widehat{\xi}_{Wald}$ $=-NT(\widehat{\theta}_{L}-\widehat{\theta}_{R})^{\prime}J(\ddot{\theta})^{\prime}\left(J(\widehat{\theta}_{L})[\triangledown_{\theta\theta^{\prime}}L(\widehat{\theta}_{L})]^{-1}J(\widehat{\theta}_{L})^{\prime}\right)^{-1}J(\ddot{\theta})(\widehat{\theta}_{L}-\widehat{\theta}_{R})$
$=\widehat{\xi}+o_{\mathbb{P}}\left(1\right).$
Next we show
$\widehat{\xi}\overset{\mathbb{D}}{\longrightarrow}\chi^{2}\left(r\right).$ (A.31)
Since $\sqrt{NT}\triangledown_{\theta}l(\theta_{0})\overset{\mathbb{D}}{\longrightarrow}\mathcal{N}(0,\Sigma),$ by the delta method, we have
$J(\theta_{0})\big{(}\mathbb{E}[\triangledown_{\theta\theta^{\prime}}l(\theta_{0})]\big{)}^{-1}\sqrt{NT}\triangledown_{\theta}l(\theta_{0})\overset{\mathbb{D}}{\longrightarrow}\mathcal{N}\big{(}0,J(\theta_{0})\Sigma J(\theta_{0})^{\prime}\big{)}.$
This implies (A.31). Therefore, the desired result follows from (A.30) and (A.31). The proof is complete.
B Additional Simulation Results
B.1 Logit Model
Table B.1: Empirical Size of LM and Wald Tests for Logit Model
(a) LM Tests
Static Model Dynamic Model
$(N,T)$ Uncorrected Corrected Uncorrected Corrected
$(30,30)$ $15\%$ $9\%$ $18\%$ $10\%$
$(60,60)$ $15\%$ $5\%$ $17\%$ $7\%$
$(90,90)$ $18\%$ $6\%$ $16\%$ $5\%$
(b) Wald Tests
Static Model Dynamic Model
$(N,T)$ Uncorrected Corrected Uncorrected Corrected
$(30,30)$ $15\%$ $9\%$ $18\%$ $10\%$
$(60,60)$ $15\%$ $5\%$ $17\%$ $7\%$
$(90,90)$ $18\%$ $6\%$ $16\%$ $5\%$
[…]
Notes: The null hypothesis $H_{0}:(\beta_{0},\rho_{0})^{\prime}=(0.5,0.5)^{\prime}$ for the dynamic model, and $H_{0}:\beta_{0}=0.5$ for the static model, with a significance level $5\%$. The number of replications is $1000$.
Table B.2: Empirical Power of LM and Wald Tests for Logit Model using $L(\theta)$
(a) LM Tests
$\delta$
$(N,T)$ $-0.20$ $-0.15$ $-0.10$ $0.10$ $0.15$ $0.20$
Static Model
$(30,30)$ $44\%$ $32\%$ $22\%$ $10\%$ $14\%$ $21\%$
$(60,60)$ $86\%$ $62\%$ $37\%$ $25\%$ $50\%$ $74\%$
$(90,90)$ $100\%$ $93\%$ $65\%$ $50\%$ $84\%$ $98\%$
Dynamic Model
$(30,30)$ $48\%$ $32\%$ $20\%$ $18\%$ $28\%$ $43\%$
$(60,60)$ $94\%$ $73\%$ $40\%$ $45\%$ $75\%$ $94\%$
$(90,90)$ $100\%$ $98\%$ $76\%$ $78\%$ $98\%$ $100\%$
(b) Wald Tests
$\delta$
$(N,T)$ $-0.20$ $-0.15$ $-0.10$ $0.10$ $0.15$ $0.20$
Static Model
$(30,30)$ $44\%$ $32\%$ $22\%$ $10\%$ $14\%$ $21\%$
$(60,60)$ $86\%$ $62\%$ $37\%$ $25\%$ $50\%$ $74\%$
$(90,90)$ $100\%$ $93\%$ $65\%$ $50\%$ $84\%$ $98\%$
Dynamic Model
$(30,30)$ $47\%$ $31\%$ $20\%$ $18\%$ $28\%$ $43\%$
$(60,60)$ $94\%$ $73\%$ $39\%$ $45\%$ $75\%$ $94\%$
$(90,90)$ $100\%$ $98\%$ $76\%$ $78\%$ $98\%$ $100\%$
[…]
Notes: The null hypothesis $H_{0}:(\beta_{0},\rho_{0})^{\prime}=(0.5+\delta,0.5+\delta)^{\prime}$ for the dynamic model, and $H_{0}:\beta_{0}=0.5+\delta$ for the static model, with a significance level $5\%,$ where $\delta\in\{\pm 0.2,\pm 0.15,\pm 0.1\}$. The number of replications is $1000$.
B.2 Probit Model
Table B.3: Comparison between the Bias-corrected and Uncorrected MLE
$(N,T)$ $(30,30)$ $(60,60)$ $(90,90)$
Bias RMSE Bias RMSE Bias RMSE
Static Model
$\widehat{\beta}$ $0.150$ $0.197$ $0.063$ $0.082$ $0.042$ $0.054$
$\widehat{\beta}_{L}$ $0.055$ $0.124$ $0.013$ $0.049$ $0.007$ $0.032$
Dynamic Model
$\widehat{\beta}$ $0.159$ $0.247$ $0.067$ $0.086$ $0.043$ $0.055$
$\widehat{\beta}_{L}^{(1)}$ $0.059$ $0.134$ $0.015$ $0.052$ $0.008$ $0.032$
$\widehat{\beta}_{L}^{(2)}$ $0.062$ $0.136$ $0.016$ $0.053$ $0.008$ $0.032$
$\widehat{\rho}$ $-0.055$ $0.124$ $-0.022$ $0.054$ $-0.017$ $0.036$
$\widehat{\rho}_{L}^{(1)}$ $-0.030$ $0.105$ $-0.009$ $0.047$ $-0.007$ $0.031$
$\widehat{\rho}_{L}^{(2)}$ $-0.022$ $0.106$ $-0.003$ $0.047$ $-0.003$ $0.031$
[…]
Notes: The true structure parameters are $\theta_{0}=(\beta_{0},\rho_{0})^{\prime}=(0.5,0.5)^{\prime}$ for the dynamic model, and $\theta_{0}=\beta_{0}=0.5$ for the static model. In the static model, $\widehat{\beta}$ and $\widehat{\beta}_{L}$ represent the uncorrected MLE and our bias-corrected MLE, respectively. In the dynamic model, $(\widehat{\beta},\widehat{\rho})$ is the uncorrected MLE and $(\widehat{\beta}_{L}^{(\tau)},\widehat{\rho}_{L}^{(\tau)})^{\prime}$ represents our bias-corrected MLE with the truncation parameter $\tau$ set to $\tau=1$ or $2$. The number of replications is $1000$.
Table B.4: Empirical Size of LR, LM, and Wald Tests Based on Bias-corrected and Uncorrected Likelihoods for Probit Model
(a) LR Tests
Static Model Dynamic Model
$(N,T)$ Uncorrected Corrected Uncorrected Corrected
$(30,30)$ $29\%$ $13\%$ $27\%$ $13\%$
$(60,60)$ $26\%$ $8\%$ $23\%$ $7\%$
$(90,90)$ $30\%$ $6\%$ $24\%$ $6\%$
(b) LM Tests
Static Model Dynamic Model
$(N,T)$ Uncorrected Corrected Uncorrected Corrected
$(30,30)$ $28\%$ $13\%$ $26\%$ $13\%$
$(60,60)$ $26\%$ $8\%$ $22\%$ $7\%$
$(90,90)$ $30\%$ $6\%$ $24\%$ $6\%$
(c) Wald Tests
Static Model Dynamic Model
$(N,T)$ Uncorrected Corrected Uncorrected Corrected
$(30,30)$ $28\%$ $13\%$ $25\%$ $13\%$
$(60,60)$ $26\%$ $8\%$ $22\%$ $7\%$
$(90,90)$ $30\%$ $6\%$ $24\%$ $6\%$
[…]
Notes: The null hypothesis $H_{0}:(\beta_{0},\rho_{0})^{\prime}=(0.5,0.5)^{\prime}$ for the dynamic model, and $H_{0}:\beta_{0}=0.5$ for the static model, with a significance level $5\%$. The number of replications is $1000$.
Table B.5: Empirical Power of LR, LM, and Wald Tests Based on Bias-corrected Likelihoods for Probit Model
(a) LR Tests
$\delta$
$(N,T)$ $-0.20$ $-0.15$ $-0.10$ $0.10$ $0.15$ $0.20$
Static Model
$(30,30)$ $75\%$ $59\%$ $41\%$ $13\%$ $23\%$ $37\%$
$(60,60)$ $100\%$ $96\%$ $71\%$ $52\%$ $86\%$ $98\%$
$(90,90)$ $100\%$ $100\%$ $96\%$ $88\%$ $99\%$ $100\%$
Dynamic Model
$(30,30)$ $83\%$ $63\%$ $40\%$ $26\%$ $47\%$ $69\%$
$(60,60)$ $100\%$ $98\%$ $78\%$ $78\%$ $98\%$ $100\%$
$(90,90)$ $100\%$ $100\%$ $99\%$ $98\%$ $100\%$ $100\%$
(b) LM Tests
$\delta$
$(N,T)$ $-0.20$ $-0.15$ $-0.10$ $0.10$ $0.15$ $0.20$
Static Model
$(30,30)$ $75\%$ $59\%$ $40\%$ $13\%$ $23\%$ $37\%$
$(60,60)$ $100\%$ $95\%$ $71\%$ $52\%$ $86\%$ $98\%$
$(90,90)$ $100\%$ $100\%$ $96\%$ $88\%$ $99\%$ $100\%$
Dynamic Model
$(30,30)$ $83\%$ $62\%$ $38\%$ $26\%$ $47\%$ $69\%$
$(60,60)$ $100\%$ $98\%$ $78\%$ $78\%$ $98\%$ $100\%$
$(90,90)$ $100\%$ $100\%$ $99\%$ $98\%$ $100\%$ $100\%$
(c) Wald Tests
$\delta$
$(N,T)$ $-0.20$ $-0.15$ $-0.10$ $0.10$ $0.15$ $0.20$
Static Model
$(30,30)$ $75\%$ $58\%$ $40\%$ $13\%$ $23\%$ $37\%$
$(60,60)$ $100\%$ $95\%$ $72\%$ $52\%$ $86\%$ $98\%$
$(90,90)$ $100\%$ $100\%$ $95\%$ $88\%$ $99\%$ $100\%$
Dynamic Model
$(30,30)$ $82\%$ $62\%$ $38\%$ $26\%$ $47\%$ $69\%$
$(60,60)$ $100\%$ $98\%$ $78\%$ $78\%$ $98\%$ $100\%$
$(90,90)$ $100\%$ $100\%$ $99\%$ $98\%$ $100\%$ $100\%$
[…]
Notes: The null hypothesis $H_{0}:(\beta_{0},\rho_{0})^{\prime}=(0.5+\delta,0.5+\delta)^{\prime}$ for the dynamic model, and $H_{0}:\beta_{0}=0.5+\delta$ for the static model, with a significance level $5\%,$ where $\delta\in\{\pm 0.2,\pm 0.15,\pm 0.1\}$. The number of replications is $1000$.
C Additional Empirical Results
C.1 Sample
The data set is constructed from waves $20-30$ of PSID, covering the survey years 1987 to 1997. We selected this time span because 1997 was the last year PSID conducted annual surveys, switching to a biennial schedule in 1999. Each survey wave includes questions about employment status in the prior calendar year, such as, ‘About how many weeks did (he/she) work on that job last year?’ and ‘During the weeks that (he/she) worked, about how many hours did (he/she) usually work per week?’ We use responses to these questions to construct individual labor force participation (LFP) status for the prior year. Thus, wave 30 (1997) provides LFP data for 1996, wave 29 (1996) for 1995, and so on. In our final sample, labor force participation status is observed from 1987 to 1996, with lagged participation status observed from 1986 to 1995.
We also restrict our sample to individuals whose LFP status changed at least once during the sample period. Individuals whose labor supply status never changed during the sample period are not informative for our analysis, because the dependent variable could be perfectly predicted by setting the individual fixed effects in the intercept of our models to plus or negative infinity.
The final sample consists of a panel of $N=90$ individuals over $T=10$ years. Table C.1 presents the summary statistics. Figure C.1 shows the average percentage of individuals with given current and past LFP status. Figure C.2 plots the LFP transition rates by year.
Table C.1: Descriptive Statistics
Variable Mean Std. Dev. Min Max
Labor Force Participation 0.482 0.500 0.000 1.000
Lagged Participation 0.458 0.498 0.000 1.000
Number of Children 2.24 1.098 1.000 6.000
  • Notes: Summary statistics are computed based on a sample of 90 individuals observed from 1987 to 1996. All variables have 900 observations. Data source: PSID 1987 - 1997.
Figure C.1: Labor Force Participation Contingency Table. Each cell contains the percentage of individuals with given current and past labor force participation status.
Figure C.2: Labor Force Participation Status Transition Rates By Year
C.2 Extra Empirical Results
Table C.2: Estimated Heterogeneous Effects on Labor Force Participation
10th 25th 75th 90th
Mean Median Std. Dev. Percentile Percentile Percentile Percentile
$\alpha_{i}$ $0$0.975 6.203 -6.909 -2.443 4.292 6.774
$\gamma_{t}$ 0.000 0.146 0.773 -1.151 -0.466 0.638 0.979
$\beta_{it}$ 0.126 -0.504 6.250 -6.835 -2.628 4.050 6.982

### v4

@@
 Abstract
-We introduce a generic class of dynamic nonlinear heterogeneous parameter models that incorporate individual and time effects in both the intercept and slope.
-To address the incidental parameter problem inherent in this class of models, we develop an analytical bias correction procedure to construct a bias-corrected likelihood.
-The resulting maximum likelihood estimators are automatically bias-corrected.
-Moreover, likelihood-based tests statistics—including likelihood-ratio, Lagrange-multiplier, and Wald tests—follow the limiting chi-square distribution under the null hypothesis.
+We introduce a generic class of dynamic nonlinear heterogeneous parameter models that incorporate individual and time fixed effects in both the intercept and slope.
+These models are subject to the incidental parameter problem, in that the limiting distribution of the point estimator is not centered at zero, and that test statistics do not follow their standard asymptotic distributions as in the absence of the fixed effects.
+To address the problem, we develop an analytical bias correction procedure to construct a bias-corrected likelihood.
+The resulting estimator follows an asymptotic normal distribution with mean zero.
+Moreover, likelihood-based tests statistics—including likelihood-ratio, Lagrange-multiplier, and Wald tests—follow the limiting chi-squared distribution under the null hypothesis.
 Simulations demonstrate the effectiveness of the proposed correction method, and an empirical application on the labor force participation of single mothers underscores its practical importance.
 Keywords:
-incidental parameter problem; bias correction; fixed effects; maximum likelihood
+incidental parameter problem; bias correction; fixed effects; panel data models
 JEL Classification:
@@
 1 Introduction
-Panel data models with two-way fixed effects are widely used in empirical research to control for unobserved heterogeneity that may correlate with covariates.
+Panel data is common in empirical research.
+Panel data models with two-way fixed effects are widely used to control for unobserved heterogeneity that may correlate with covariates.
 Individual effects capture time-invariant, individual-specific characteristics, while time effects account for shifts that are common across individuals but vary over time.
@@
 This flexible framework accommodates multi-dimensional heterogeneities and encompasses many commonly used panel data models.
-To address the incidental parameter problem inherent in DN-HP models, we propose a likelihood-based analytical bias correction approach.
-This method ensures the validity of the maximum likelihood estimator and likelihood-based inference for the structural parameters of DN-HP models.
-As a preview, Figure 1 presents simulated boxplots of the estimation bias in a DN-HP logit model, comparing maximum likelihood estimators from our bias-corrected likelihoods and the original likelihoods.
-The results reveal that the maximum likelihood estimators obtained from the bias-corrected likelihoods significantly reduce estimation bias while maintaining comparable mean square errors.
+DN-HP models are generally subject to the incidental parameter problem, in that the limiting distribution of the point estimator is not centered at 00, and that the test statistics do not follow their standard asymptotic distributions (of $\chi^{2}$, for many test statistics).
+To address the incidental parameter problem, we propose an analytical bias correction approach.
+This method restores the standard asymptotic distributions for the point estimator and the test statistics.
 Figure 1: Comparison of Uncorrected and Corrected Maximum Likelihood Estimators
+{justify}
 Notes: Model.
 $Y_{it}=\mathds{1}\{\rho Y_{it-1}+(\beta+\alpha_{1,i}+\gamma_{1,t})X_{it}+\alpha_{2,i}+\gamma_{2,t}+\varepsilon_{it}>0\}$ where $\mathds{1}\{\cdot\}$ denotes the indicator function, $\varepsilon_{it}$ is standard logistic independent of $X_{it}$, and the true values $\rho_{0}=\beta_{0}=0.5$.
-($\widehat{\beta},\widehat{\rho})$ are the uncorrected estimators and $(\widehat{\beta}_{L},\widehat{\rho}_{L})$ are the corrected estimators.
+$\widehat{\beta}$ and $\widehat{\rho}$ are the uncorrected estimators while $\widehat{\beta}_{L}$ and $\widehat{\rho}_{L}$ are the corrected estimators.
 $1000$ replications.
-For details on the data-generating process, please refer to Section 4.
-1.1 Literature Review
+The detailed data-generating process is given in Section 4.
+As a preview, Figure 1 presents simulated boxplots of the estimation bias in a DN-HP logit model, comparing estimators from our bias correction procedure to the uncorrected ones.
+The results reveal that our bias correction procedure significantly reduces the estimation bias while maintaining comparable mean square errors.
+Literature Review.
 The estimation and inference of fixed effects models in the presence of the incidental parameter problem have been extensively studied.
@@
 Neither our DN-HP models nor our bias correction procedure have been discussed in the current literature.
-1.2 This Paper
-We assume that the multi-dimensional fixed effects enter the log-likelihood additively.
-Identification restrictions are imposed on the fixed effects, whose dimension grows with the sample size.
-Our corrected likelihood is derived using two Taylor expansions: the expansion of the score function and the expansion of the original likelihood, both with respect to the fixed effects.
-Under appropriate dominance conditions, we explicitly derive the two additive bias terms in the original likelihoods of DN-HP models.
-These bias terms are expressed as quadratic forms involving the score and Hessian of the fixed effects.
-We estimate the derived bias terms under weak serial dependence (conditional strong mixing) and conditional cross-sectional independence.
-We rigorously show that the maximum likelihood estimator (MLE) based on the bias-corrected likelihood of the DN-HP model retains a limiting normal distribution with zero mean (automatically bias-corrected), consistent with the limiting distribution of the classical MLE in the absence of incidental parameters.
-Beyond the bias-corrected MLE, this likelihood-based bias correction approach also facilitates hypothesis testing using the likelihood ratio (LR), Lagrange multiplier (LM), and Wald statistics.
-We demonstrate that these three test statistics, derived from the bias-corrected likelihood, are asymptotically equivalent, sharing the same asymptotic $\chi^{2}$ distribution under the null hypothesis.
-We demonstrate the finite-sample performance of our method through Monte Carlo simulations, comparing it to both the original likelihood and the Jackknife.
+This paper.
+We employ the maximum likelihood framework and focus on multi-dimensional additive two-way fixed effects, whose number grows with the sample size.
+We consider an arbitrary DN-HP model whose log-likelihood function is specified up to the unknown parameters.
+We construct a modified, or corrected, log-likelihood function by adding two bias correction terms to the original one.
+These two bias terms are analytically derived by combining i) the Taylor expansion of the original log-likelihood function, in terms of the fixed effects, and ii) an asymptotic expansion of the fixed effects themselves.
+We show that the corrected log-likelihood function does not suffer from the incidental parameter problem, under appropriate regularity conditions.
+One key benefit of our procedure is that we achieve bias corrections for the point estimators and the test statistics by modifying a single object, the log-likelihood function.
+We rigorously show that the estimators obtained by maximizing the corrected log-likelihood function retains a limiting normal distribution with zero mean, consistent with the asymptotic theory of the classical maximum likelihood estimators (MLE) in the absence of incidental parameters.
+Beyond the estimators, we also show that the likelihood ratio (LR), Lagrange multiplier (LM), and Wald statistics, derived from the same corrected log-likelihood function, are asymptotically equivalent, sharing the same asymptotic $\chi^{2}$ distribution under the null hypothesis.
+We demonstrate the finite-sample performance of our method through Monte Carlo simulations, comparing it to both the original likelihood and the Jackknife, which is a potential bias correction device [^1] It is worth mentioning that, to our knowledge, there currently has been no study on using the jackknife under the DN-HP models..
+We show that our correction procedure reduces the bias significantly, without increasing the root mean squared errors, and restores the test sizes of the LR test to its nominal level.
+We also find that our procedure outperforms the Jackknife.
 In the empirical application, we apply the DN-HP logit model to study the aforementioned question of how the number of children affects a mother’s labor force participation, with a focus on single-mother households.
@@
 The remainder of the paper is organized as follows.
-Section 2 describes our settings, the identification restriction, the incidental parameter problem under DN-HP, and gives an overview of our bias correction procedures.
-Section 3 provides the bias expansion of the likelihood and the formula for our corrected likelihood.
-The statistical properties of the corrected likelihood.
-We rigorously show that the estimator of the structural parameter, and the LR, LM, and Wald statistics all follow their standard asymptotic distribution.
-Section 4 presents simulation studies to demonstrate our method’s performance.
-Here we show that our correction procedure reduces the bias significantly, without increasing the root mean squared errors.
-In addition, we also show that our procedure restores the test size of the LR test to its nominal level.
-We also compare our procedure with the Jackknife and find that our procedure outperforms the Jackknife.
+Section 2 explains our settings, the identification restriction, and the incidental parameter problem of DN-HP models.
+Section 3 describes our bias correction procedure and provides relevant statistical properties, including those we claim above.
+Section 4 presents simulation studies to demonstrate the performance of method.
 Section 5 applies our method to study single mothers’ labor force participation.
+Finally in Section 6, we leave some closing remarks.
 Proofs, technical details, and additional results are provided in the appendix.
-Notations.
-In what follows, we denote $\iota_{n}$ to be an $n\times 1$ vector of ones and for $\mathbb{I}_{n}$ to be an $n\times n$ identity matrix.
-Let $\otimes$ be the kronecker product and $\mathds{1}\left\{\cdot\right\}$ the indicator function.
-In addition, for each square matrix $A_{i}$, $i=1,\ldots,n$, $\operatorname*{diag}\left\{A_{1},\ldots,A_{n}\right\}$ represent a block diagonal matrix, where each matrix $A_{i}$ is the $i$-th diagonal element.
-2 Bias in Heterogeneous Parameter Model
-2.1 DN-HP Model
-Consider the responses $Y_{it},$ which follow the conditional density $f\left(Y_{it};\theta_{0}+\alpha_{i}^{0}+\gamma_{t}^{0}\right),$ given covariates $X_{it}$ (omitted for simplicity).
-Here, $i=1,\ldots,N$ represents individuals, and $t=1,\ldots,T$ represents time periods.
-The density is known up to $\theta_{0}$, $\alpha_{i}^{0}$, and $\gamma_{t}^{0},$ and the covariates $X_{it}$ may include lagged $Y_{it}$.
-Specifically, $\theta_{0}$ is the $K$-dimensional structural parameter of interest, where $K$ is fixed and known.
-The nuisance heterogeneous effects $\alpha_{i}^{0}=\left(\alpha_{1,i}^{0},\ldots,\alpha_{K,i}^{0}\right)^{\prime}$ and $\gamma_{t}^{0}=\left(\gamma_{1,t}^{0},\ldots,\gamma_{K,t}^{0}\right)^{\prime}$ may be correlated with $X_{it}$.
-Throughout the paper, we treat $\alpha_{i}^{0}$ and $\gamma_{t}^{0}$ as fixed effects (unknown parameters) by conditioning on them in the density.
-Denote $\phi:=\left(\alpha_{1}^{\prime},\ldots,\alpha_{N}^{\prime},\gamma_{1}^{\prime},\ldots,\gamma_{T}^{\prime}\right)^{\prime}$ and define the quasi-likelihood function as
-$l(\theta,\phi):=\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}l_{i,t}\left(\theta,\phi\right),\qquad l_{i,t}(\theta,\phi):=\log f\left(Y_{it};\theta+\alpha_{i}+\gamma_{t}\right).$
-Note that if the data are independent across both $i$ and $t,$ $l(\theta,\phi)$ represents the exact likelihood.
-However, when observations for each unit $i$ exhibit serial correlation, $l(\theta,\phi)$ is misspecified even though the likelihood for each observation $Y_{it}\mid X_{it},$ $l_{i,t}(\theta,\phi),$ is correctly specified.
-Our asymptotic theory assumes conditional cross-sectional independence and weak time-serial dependence (Assumption 1).
-For simplicity, we refer to $l(\theta,\phi)$ as the likelihood function hereafter.
-Here is an example that illustrates how our model captures greater heterogeneity than conventional two-way effects panel data models.
+Notation.
+We denote $\mathbb{I}_{n}$ to be the $n\times n$ identity matrix and $\iota_{n}$ the $n\times 1$ vector of ones.
+$\otimes$ denotes the Kronecker product and $\mathds{1}\{\cdot\}$ is the indicator function.
+Next, for a sequence of square matrices $A_{i}$, $i=1,\ldots,n$, $\operatorname*{diag}\{A_{1},\ldots,A_{n}\}$ represents a block diagonal matrix where each $A_{i}$ is the $i$-th diagonal element.
+For a vector $v$ and a function $f(v)$, we write $\partial_{v}$ and $\partial_{vv^{\prime}}$ to represent the first and second partial derivatives of $f(v)$ with respect to (w.r.t.) $v$.
+2 Bias in Heterogeneous Parameter Models
+2.1 Models and Estimations
+Let $\{(Y_{it},X_{it}^{\prime})^{\prime}:i=1,\ldots,N;t=1,\ldots,T\}$ be a panel data set of $N$ individuals and $T$ time periods, where, $i$ represents the individual and $t$ represents the time period.
+Here $Y_{it}$ is a scalar response variable, and $X_{it}$ is a $K\times 1$ vector of regressors, with $K$ known and fixed.
+For each individual $i$, the response $Y_{it}$ is generated sequentially over $t$, by
+$Y_{it}|X_{i1},\ldots,X_{it},\phi^{0}\sim f(y|X_{it},\theta_{0}+\alpha_{i}^{0}+\gamma_{t}^{0}),$ (2.1)
+where $\theta_{0}$ is a $K\times 1$ vector of (non-random) unknown common parameters of interests, $\phi^{0}:=(\alpha_{1}^{0\prime},\ldots,\alpha_{N}^{0\prime},\gamma_{1}^{0\prime},\ldots,\gamma_{T}^{0\prime})^{\prime}$ consists of the $K\times 1$ vectors of random variables $\alpha_{i}^{0}$ and $\gamma_{t}^{0}$, and $f(y|\cdot)$ is a density known up to $\theta_{0}$, $\alpha_{i}^{0}$, and $\gamma_{t}^{0}$.
+We restrict our attention to the case where $f\left(\cdot\right)$ depends on $\theta_{0}$, $\alpha_{i}^{0}$, and $\gamma_{t}^{0}$ through the additive structure $\theta_{0}+\alpha_{i}^{0}+\gamma_{t}^{0}$, the “heterogeneous parameter”.
+The vectors $\alpha_{i}^{0}$ and $\gamma_{t}^{0}$ capture the individual-specific and time-specific heterogeneities, respectively.
+Generally, $\alpha_{i}^{0}$ and $\gamma_{t}^{0}$ may be correlated with the regressor $X_{it}$.
+However, throughout the paper, we employ the “fixed-effect” framework to condition on the realizations of $\alpha_{i}^{0}$ and $\gamma_{t}^{0}$ (denoted by the same symbols) in the density, and treat them as unknown nuisance parameters to be estimated together with $\theta_{0}$.
+Conditioning on $\alpha_{i}^{0}$ and $\gamma_{t}^{0}$, we assume $(Y_{it},X_{it}^{\prime})^{\prime}$ independent across $i$ but serially dependent over $t$.
+In particular, we allow $X_{it}$ to contain both strictly exogenous components and predetermined components w.r.t.
+$Y_{it}$.
+For instance, $X_{it}$ may contain lagged values of $Y_{it}$.
+In this paper, we focus on the case where only heterogeneous parameters appear in (2.1).
+This is without loss of generality.
+As illustrated in the example below, our approach accommodates settings where both heterogeneous parameters and “homogeneous parameters” are present.
+[^2] Another example is the classic Neyman-Scott model, where $Y_{it}$ follows a normal distribution with a homogeneous variance but heterogeneous means.
+See Example 3 below.
 Example 1 (heterogeneous parameter model)
-Consider a probit model
-$Y_{it}=\mathds{1}\left\{\pi_{it}\left(\theta,\alpha_{i},\gamma_{t}\right)+\varepsilon_{it}>0\right\}$
-where the linear index $\pi_{it}(\cdot)$ varies across both individuals and time periods.
-The error term $\varepsilon_{it}$ follows a standard normal distribution denoted by $\Phi$.
-Without of loss of generality, let $X_{it}\in\mathbb{R},$ and the parameter vectors $\theta=\left(\theta_{1},\theta_{2}\right)^{\prime},$ $\alpha_{i}=\left(\alpha_{1,i},\alpha_{2,i}\right)^{\prime},$ and $\gamma_{t}=\left(\gamma_{1,t},\gamma_{2,t}\right)^{\prime}.$ We accommodate the linear index of the form:
-$\pi_{it}\left(\theta,\alpha_{i},\gamma_{t}\right)=\left(\theta_{1}+\alpha_{1,i}+\gamma_{1,t}\right)+X_{it}\left(\theta_{2}+\alpha_{2,i}+\gamma_{2,t}\right).$ (2.1)
-In the special case where $\alpha_{i}=\alpha_{1,i}$ and $\gamma_{t}=\gamma_{1,t},$ (2.1) simplifies to
-$\pi_{it}\left(\theta,\alpha_{i},\gamma_{t}\right)=\left(\theta_{1}+\alpha_{1,i}+\gamma_{1,t}\right)+X_{it}\theta_{2},$
-which corresponds to the traditional two-way fixed effects model. Assume that $\varepsilon_{it}$ is independent of $X_{it}.$ The likelihood for individual $i$ at time $t$ is given by
-$l_{i,t}\left(\theta,\phi\right):=Y_{it}\log\Phi\left(\pi_{it}\left(\theta,\alpha_{i},\gamma_{t}\right)\right)+\left(1-Y_{it}\right)\log\Phi\left(-\pi_{it}\left(\theta,\alpha_{i},\gamma_{t}\right)\right).$
-Due to the additive form of the parameters, the structural parameters, $\theta,$ is not identified in the likelihood.
-To resolve this, we reparameterize the likelihood by setting $\alpha_{N}=-\sum_{i=1}^{N-1}\alpha_{i}$ and $\gamma_{T}=-\sum_{t=1}^{T-1}\gamma_{t},$ thereby imposing the constraints $\sum_{i=1}^{N}\alpha_{i}=0$ and $\sum_{t=1}^{T}\gamma_{t}=0.$ Let $\psi:=\left(\alpha_{1}^{\prime},\ldots,\alpha_{N-1}^{\prime},\gamma_{1}^{\prime},\ldots,\gamma_{T-1}^{\prime}\right)^{\prime}.$ The reparameterized likelihood is then $l\left(\theta,D^{\prime}\psi\right)=:l^{*}\left(\theta,\psi\right)$ for $D:=\operatorname*{diag}\{D_{1},D_{2}\}$ where $D_{1}:=(\mathbb{I}_{N-1},-\iota_{N-1})\otimes\mathbb{I}_{K}$ and $D_{2}:=(\mathbb{I}_{T-1},-\iota_{T-1})\otimes\mathbb{I}_{K}.$
+Consider a probit model with data generating process (DGP)
+$Y_{it}|X_{i1},\ldots,X_{it},\phi^{0}=\mathds{1}\{\pi(X_{it},\theta_{0},\alpha_{i}^{0},\gamma_{t}^{0})+\varepsilon_{it}>0\},$ (2.2)
+where $\varepsilon_{it}$ is a standard normal random variable independent of $X_{it}$, $\theta_{0}$, $\alpha_{i}^{0}$, and $\gamma_{t}^{0}$.
+For arguments $\theta$, $\alpha_{i}$, and $\gamma_{t}$ of dimension $K\times 1$, we allow the linear index [^3] We use a model with a linear index solely for the purpose of illustration.
+Our method applies to general models.
+$\pi(X_{it},\theta,\alpha_{i},\gamma_{t})$ to take the following general form:
+$\pi(X_{it},\theta,\alpha_{i},\gamma_{t})=\sum_{k=1}^{K}(\theta_{k}+\alpha_{k,i}+\gamma_{k,t})X_{k,it},$
+where $X_{k,it}$, $\theta_{k}$, $\alpha_{k,i}$, and $\gamma_{k,t}$ are the $k$-th components of $X_{it}$, $\theta$, $\alpha_{i}$, and $\gamma_{t}$, respectively, $k=1,\ldots,K$.
+In addition, when $X_{1,it}=1$, the DGP in (2.2) consists of the traditional individual- and time-effects in the intercept.
+When $X_{k,it}=Y_{it-1}$ for some $k$, the DGP is dynamic because each $Y_{it}$ depends on its past value.
+Moreover, we also allow a subset of the coefficients to be homogeneous, thereby accommodating both heterogeneous and homogeneous parameters.
+For instance, when it is known that $\alpha_{1,i}^{0}=\gamma_{1,t}^{0}=0$, the model can be estimated under constraints $\alpha_{1,i}=\gamma_{1,t}=0$ for all $(i,t)$.
+In this case, the linear index is effectively
+$\pi(X_{it},\theta,\alpha_{i},\gamma_{t})=\theta_{1}X_{1,it}+\sum_{k=2}^{K}(\theta_{k}+\alpha_{k,i}+\gamma_{k,t})X_{k,it},$
+where $X_{1,it}$ is associated with a “homogeneous parameter” $\theta_{1}$ while all other $X_{k,it}$ ($k=2,\ldots,K$) are associated with heterogeneous parameters $(\theta_{k}+\alpha_{k,i}+\gamma_{k,t})$.
+In a similar fashion, our approach also applies to, e.g.,
+$\pi(X_{it},\theta,\alpha_{i},\gamma_{t})=(\alpha_{1,i}+\gamma_{1,t})X_{1,it}+\sum_{k=2}^{K}(\theta_{k}+\alpha_{k,i}+\gamma_{k,t})X_{k,it},$
+with the constraint $\theta_{1}=0$. Other variants of $\pi(X_{it},\theta,\alpha_{i},\gamma_{t})$ include:
+$\pi(X_{it},\theta,\alpha_{i},\gamma_{t})$ $=(\alpha_{1,i}+\gamma_{1,t})+\theta_{2}Y_{it-1}+\sum_{k=3}^{K}\theta_{k}X_{k,it},$ (2.3)
+$\pi(X_{it},\theta,\alpha_{i},\gamma_{t})$ $=(\theta_{1}+\alpha_{1,i}+\gamma_{1,t})+\sum_{k=2}^{K}(\theta_{k}+\alpha_{k,i}+\gamma_{k,t})X_{k,it},$ (2.4)
+$\pi(X_{it},\theta,\alpha_{i},\gamma_{t})$ $=(\theta_{1}+\alpha_{1,i}+\gamma_{1,t})+(\theta_{2}+\alpha_{2,i}+\gamma_{2,t})Y_{it-1}$
+$\quad+\sum_{k=3}^{K_{1}}(\theta_{k}+\alpha_{k,i}+\gamma_{k,t})X_{k,it}+\sum_{k=K_{1}+1}^{K}\theta_{k}X_{k,it}$ (2.5)
+for some $3\leq K_{1}<K$.
+To our knowledge, except (2.3) considered by Fernández-Val and Weidner (2016), none of the above specifications are fully addressed under nonlinear models with predetermined regressors.
+The specification in (2.4) is the same to Keane and Neal (2020) and Lu and Su (2023), except that they consider only the linear regression model.
+Jochmans and Otsu (2019) provide a bias correction technique for general nonlinear models with (2.4) only for strictly exogenous $X_{k,it}$.
+Currently, no methods are available for nonlinear models with (2.5).
+We are interested in the estimation of and the inference about $\theta_{0}$ in (2.1), under the presence of the nuisance parameter $\phi^{0}$.
+Our approach is likelihood-based.
+Denote $\phi:=(\alpha_{1}^{\prime},\ldots,\alpha_{N}^{\prime},\gamma_{1}^{\prime},\ldots,\gamma_{T}^{\prime})^{\prime}$ and define the log-likelihood function (“likelihood” hereafter) as
+$\ell(\theta,\phi):=\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}l_{i,t}(\theta,\phi),\qquad l_{i,t}(\theta,\phi):=\log f(Y_{it}|X_{it},\theta+\alpha_{i}+\gamma_{t}).$ (2.6)
+Due to the additive structure $\theta+\alpha_{i}+\gamma_{t}$, $\theta$ is not identified because the likelihood function is invariant to transformations
+$\alpha_{i}\mapsto\alpha_{i}+c_{1},\qquad\gamma_{t}\mapsto\gamma_{t}+c_{2},\qquad\theta\mapsto\theta-c_{1}-c_{2}$ (2.7)
+for all pairs of constant $(c_{1},c_{2})$ and all $(i,t)$.
+To resolve this, we follow Lu and Su (2023) to reparameterize the likelihood by setting $\alpha_{N}=-\sum_{i=1}^{N-1}\alpha_{i}$ and $\gamma_{T}=-\sum_{t=1}^{T-1}\gamma_{t}$, thereby imposing the normalization $\sum_{i=1}^{N}\alpha_{i}=0$ and $\sum_{t=1}^{T}\gamma_{t}=0$.
+Any “average effect” is absorbed by $\theta$.
+Let $\psi:=\left(\alpha_{1}^{\prime},\ldots,\alpha_{N-1}^{\prime},\gamma_{1}^{\prime},\ldots,\gamma_{T-1}^{\prime}\right)^{\prime}$, the reparameterized likelihood can be constructed from the original $\ell\left(\theta,\phi\right)$ as
+$\ell\left(\theta,D^{\prime}\psi\right)=:l(\theta,\psi),\qquad D:=\operatorname*{diag}\{D_{1},D_{2}\},$
+where $D_{1}:=(\mathbb{I}_{N-1},-\iota_{N-1})\otimes\mathbb{I}_{K}$ and $D_{2}:=(\mathbb{I}_{T-1},-\iota_{T-1})\otimes\mathbb{I}_{K}$.
+This rules out the indeterminancy in (2.7) without explicitly using the Lagrangian.
+We use the following example to illustrate the reparameterization.
+Example 2 (reparameterization)
+For some scalar structural parameter $\mu_{0}$ of interests and $K=1$, consider $Y_{it}=\mu_{0}+\alpha_{i}^{0}+\gamma_{t}^{0}+\varepsilon_{it}$ where $\varepsilon_{it}$ is standard normal independent across both $i$ and $t$, and independent of $\alpha_{i}^{0}$ and $\gamma_{t}^{0}$.
+We restrict the variance of $\varepsilon_{it}$ to be $1$ because we use this example only to demonstrate the identification of $\mu_{0}$.
+In Example 3 below, we indeed allow $\varepsilon_{it}$ to have an unknown variance.
+Conditioning on $\phi^{0}$, the original likelihood function is
+$\ell(\mu,\phi)$ $=-\frac{1}{2}\log(2\pi)-\frac{1}{NT}\sum_{i=1}^{N-1}\sum_{t=1}^{T-1}\frac{1}{2}(Y_{it}-\mu-\alpha_{i}-\gamma_{t})^{2}-\frac{1}{NT}\sum_{t=1}^{T-1}\frac{1}{2}(Y_{Nt}-\mu-\alpha_{N}-\gamma_{t})^{2}$
+$\quad-\frac{1}{NT}\sum_{i=1}^{N-1}\frac{1}{2}(Y_{iT}-\mu-\alpha_{i}-\gamma_{T})^{2}-\frac{1}{NT}\frac{1}{2}(Y_{NT}-\mu-\alpha_{N}-\gamma_{T})^{2}.$
+Setting $\alpha_{N}=-\sum_{i=1}^{N-1}\alpha_{i}$ and $\gamma_{T}=-\sum_{t=1}^{T-1}\gamma_{t}$, the reparameterized likelihood is
+$l(\mu,\psi)$ $=-\frac{1}{2}\log(2\pi)-\frac{1}{NT}\sum_{i=1}^{N-1}\sum_{t=1}^{T-1}\frac{1}{2}(Y_{it}-\mu-\alpha_{i}-\gamma_{t})^{2}-\frac{1}{NT}\sum_{t=1}^{T-1}\frac{1}{2}\left(Y_{Nt}-\mu+\sum_{i=1}^{N-1}\alpha_{i}-\gamma_{t}\right)^{2}$
+$\quad-\frac{1}{NT}\sum_{i=1}^{N-1}\frac{1}{2}\left(Y_{iT}-\mu-\alpha_{i}+\sum_{t=1}^{T-1}\gamma_{t}\right)^{2}-\frac{1}{NT}\frac{1}{2}\left(Y_{NT}-\mu+\sum_{i=1}^{N-1}\alpha_{i}+\sum_{t=1}^{T-1}\gamma_{t}\right)^{2}.$
+Calculating the first derivative of $l(\mu,\psi)$ w.r.t. $\alpha_{i}$, we have
+$\partial_{\alpha_{i}}l(\mu,\psi)=\frac{1}{NT}\sum_{t=1}^{T}(Y_{it}-\mu-\alpha_{i})-\frac{1}{NT}\sum_{t=1}^{T}\left(Y_{Nt}-\mu+\sum_{i=1}^{N-1}\alpha_{i}\right)$
+for every $i=1,\ldots,N-1$. Equating $\partial_{\alpha_{i}}l(\mu,\psi)=0$, we obtain
+$\alpha_{i}=\frac{1}{T}\sum_{t=1}^{T}Y_{it}-\frac{1}{T}\sum_{t=1}^{T}Y_{Nt}-\sum_{i=1}^{N-1}\alpha_{i}.$ (2.8)
+Notice that (2.8) is independent of $\mu$ and all $\gamma_{t}$. Now, sum (2.8) over $i=1,\ldots,N-1$, we obtain
+$\sum_{i=1}^{N-1}\alpha_{i}=\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}Y_{it}-\frac{1}{T}\sum_{t=1}^{T}Y_{Nt}.$ (2.9)
+Combining (2.8) and (2.9), it follows that
+$\widehat{\alpha}_{i}=\frac{1}{T}\sum_{t=1}^{T}Y_{it}-\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}Y_{it},\qquad i=1,\ldots,N-1.$
+It can be calculated in the same way that
+$\widehat{\gamma}_{t}=\frac{1}{N}\sum_{i=1}^{N}Y_{it}-\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}Y_{it},\qquad t=1,\ldots,T-1.$
+Subsequently,
+$\widehat{\mu}=\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}Y_{it}.$ (2.10)
+In this simple example, $\widehat{\alpha}_{i}$ and $\widehat{\gamma}_{t}$ do not depend on $\mu$. This is generally not the case for complex models.
+In addition, maximizing $l(\mu,\psi)$ is identical to maximizing the original likelihood $\ell\left(\mu,\phi\right)$ under constraints $\sum_{i=1}^{N}\alpha_{i}=0$ and $\sum_{t=1}^{T}\gamma_{t}=0$.
+Specifically, it can be verified that the Lagrangian $\ell\left(\mu,\phi\right)+\lambda^{\alpha}\sum_{i=1}^{N}\alpha_{i}+\lambda^{\gamma}\sum_{t=1}^{T}\gamma_{t}$, where $\lambda^{\alpha}$ and $\lambda^{\gamma}$ are the two Lagrange multipliers, delivers upon maximization the same $\widehat{\alpha}_{i}$ and $\widehat{\gamma}_{t}$ for $i=1,\ldots,N-1$ and $t=1,\ldots,T$, respectively.
+For $\widehat{\alpha}_{N}$ and $\widehat{\gamma}_{T}$, the Lagrangian produces
+$\widehat{\alpha}_{N}=\frac{1}{T}\sum_{t=1}^{T}Y_{Nt}-\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}Y_{it},\qquad\widehat{\gamma}_{T}=\frac{1}{N}\sum_{i=1}^{N}Y_{iT}-\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}Y_{it}.$
+Indeed, $\widehat{\alpha}_{N}=-\sum_{i=1}^{N-1}\widehat{\alpha}_{i}$ as in Equation (2.9), and it can be verified that $\widehat{\gamma}_{T}=-\sum_{t=1}^{T-1}\widehat{\gamma}_{t}$ as well.
 2.2 Incidental Parameter Problem
-In this section, we illustrate the incidental parameter problem (IPP) that arises in our heterogeneous parameter model. For a given $\theta,$ define
-$\psi(\theta):=\arg\max_{\psi\in\Lambda_{\theta}}\mathbb{E}[l^{*}(\theta,\psi)],\quad\widehat{\psi}(\theta):=\arg\max_{\psi\in\Lambda_{\theta}}l^{*}(\theta,\psi)$ (2.2)
-where $\Lambda_{\theta}$ is a compact subset of $\mathbb{R}^{(N+T-2)K}$ for fixed $N$ and $T$ (Assumption 1) and $\mathbb{E}$ denotes expectations with respect to the true conditional density.
-Specifically, $\mathbb{E}$ represents the conditional expectation given the true values of fixed effects and the covariates.
-For notational simplicity, we use $\mathbb{E}$ when no confusion arises.
-Let the score and Hessian be
-$s\left(\theta,\psi\right):=\partial_{\psi}l^{*}\left(\theta,\psi\right),\quad H\left(\theta,\psi\right):=\partial_{\psi\psi^{\prime}}l^{*}\left(\theta,\psi\right),$
-where $\partial_{\psi}$ and $\partial_{\psi\psi^{\prime}}$ represent the first and second partial derivatives with respect to $\psi$.
-By (2.2), we know that $s(\theta,\widehat{\psi}\left(\theta\right))=0.$ Given $\theta,$ consider a Taylor expansion of $s(\theta,\widehat{\psi}\left(\theta\right))$ around $\psi(\theta)$:
-$0=s\left(\theta,\psi\left(\theta\right)\right)+H\left(\theta,\psi(\theta)\right)\big{(}\widehat{\psi}(\theta)-\psi(\theta)\big{)}+r^{s}(\theta,a),$ (2.3)
-where
-$r^{s}\left(\theta,a\right)$ $:=\big{(}r_{\psi_{1}}^{s}\left(\theta,a_{1}\right),\ldots,r_{\psi_{(N+T-2)K}}^{s}(\theta,a_{(N+T-2)K})\big{)}^{\prime},$
-$r_{\psi_{j}}^{s}\left(\theta,a_{j}\right)$ $:=\frac{1}{2}\big{(}\widehat{\psi}\left(\theta\right)-\psi(\theta)\big{)}^{\prime}G_{\psi_{j}}(\theta,a_{j})\big{(}\widehat{\psi}(\theta)-\psi(\theta)\big{)},$
-$G_{\psi_{j}}(\theta,a_{j})$ $:=\partial_{\psi_{j}}H(\theta,a_{j}\widehat{\psi}(\theta)+(1-a_{j})\psi\left(\theta\right))$
-with $a=(a_{1},\ldots,a_{(N+T-2)K})^{\prime}$ for each $a_{j}\in(0,1),$ and $\psi_{j}$ denotes the $j$th component of $\psi.$ Here $G_{\psi_{j}}(\theta,a_{j})$ is the derivative of each element of $H(\theta,\psi)$ with respect to $\psi_{j},$ evaluated at the point $a_{j}\widehat{\psi}\left(\theta\right)+\left(1-a_{j}\right)\psi\left(\theta\right).$ For notational simplicity, we write
-$s(\theta):=s\left(\theta,\psi\left(\theta\right)\right)\quad\text{and}\quad H\left(\theta\right):=H\left(\theta,\psi(\theta)\right).$ (2.4)
-From (2.3), by Assumption 1, we obtain that
-$\widehat{\psi}(\theta)-\psi(\theta)$ $=[-{H}(\theta)]^{-1}\big{(}s(\theta)+r^{s}\left(\theta,a\right)\big{)}.$ (2.5)
-Next, define the profile likelihoods as $\widehat{l}(\theta):=l^{*}(\theta,\psi)\mid_{\psi=\widehat{\psi}(\theta)}$ and $l(\theta):=l^{*}(\theta,\psi)\mid_{\psi={\psi}(\theta)}$.
-A Taylor expansion of $\widehat{l}(\theta)$ around $\psi(\theta)$ gives
-$\widehat{l}(\theta)=l(\theta)+s^{\prime}\left(\theta\right)\big{(}\widehat{\psi}(\theta)-\psi(\theta)\big{)}+\frac{1}{2}\big{(}\widehat{\psi}(\theta)-\psi(\theta)\big{)}^{\prime}{H}(\theta)\big{(}\widehat{\psi}(\theta)-\psi(\theta)\big{)}+r^{l}\left(\theta,a_{0}\right)$ (2.6)
-where $r^{l}\left(\theta,a_{0}\right):=1/3\big{(}\widehat{\psi}(\theta)-\psi(\theta)\big{)}^{\prime}r^{s}(\theta,a_{0}\,\iota_{(N+T-2)K})$ for some $a_{0}\in(0,1).$ Combining (2.5) and (2.6), we have
-$\widehat{l}(\theta)=l(\theta)-\frac{1}{2}s^{\prime}(\theta)[{H}(\theta)]^{-1}s(\theta)+r(\theta)$ (2.7)
-where $r(\theta):=1/2(r^{s}\left(\theta,a\right))^{\prime}[{H}(\theta)]^{-1}r^{s}\left(\theta,a\right)+r^{l}\left(\theta,a_{0}\right).$ Here we suppress $a$ and $a_{0}$ in $r(\theta)$ as they are constants.
-Under mild regularity conditions, we show that
-$\sqrt{NT}\mathbb{E}\big{\{}s^{\prime}(\theta)[{H}(\theta)]^{-1}s(\theta)\big{\}}\nrightarrow 0$
-and that $r(\theta)$ is a remainder term that is negligible for inference (see Section 3).
-This implies that the presence of $s^{\prime}(\theta)[{H}(\theta)]^{-1}s(\theta)$ in (2.7), due to the estimation of $\psi(\theta),$ contributes the MLE of $\theta_{0}$ based on $\widehat{l}(\theta).$ Consequently, the structural estimators and $\widehat{l}(\theta)$ based test statistics no longer follow the standard asymptotic distributions as in classical maximum likelihood theory.
-This is referred to as the IPP in this context.
-We use the following example to illustrate IPP.
-Example 2 (heterogeneous means)
-Assume that $Y_{it}$ follows a normal distribution with mean $\theta_{0}+\alpha_{i}^{0}+\gamma_{t}^{0}$ and unit variance, independent across both $i$ and $t.$ $\theta_{0}$ is the structural parameter of interest, while $\alpha_{i}^{0}$ and $\gamma_{t}^{0}$ are fixed effects.
-The likelihood function is
-$l\left(\theta,\phi\right)=\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}-\frac{1}{2}\left(Y_{it}-\theta-\alpha_{i}-\gamma_{t}\right)^{2}-\frac{1}{2}\log(2\pi).$
-Define $\widetilde{Y}_{it}:=Y_{it}-\mathbb{E}Y_{it}.$ Under the identification condition, the reparameterized likelihood $l^{\ast}(\theta,\psi)=l(\theta,D^{\prime}\psi).$ Given $\theta,$ maximizing $l^{\ast}(\theta,\psi)$ with respect to $\psi$ yields
-$\widehat{\alpha}_{i}(\theta)-\alpha_{i}^{0}$ $=$ $\frac{1}{T}\sum_{t=1}^{T}\widetilde{Y}_{it}-\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}\widetilde{Y}_{it},\quad i=1,\ldots,N-1,$ (2.8)
-$\widehat{\gamma}_{t}(\theta)-\gamma_{t}^{0}$ $=$ $\frac{1}{N}\sum_{i=1}^{N}\widetilde{Y}_{it}-\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}\widetilde{Y}_{it},\quad t=1,\ldots,T-1.$ (2.9)
-Let $\big{(}\widehat{\alpha}_{1}(\theta),\ldots,\widehat{\alpha}_{N-1}(\theta),\widehat{\gamma}_{1}(\theta),\ldots,\widehat{\gamma}_{T-1}(\theta)\big{)}^{\prime}=:\widehat{\psi}(\theta)$ and $\big{(}{\alpha}_{1}^{0},\ldots,{\alpha}_{N-1}^{0},{\gamma}_{1}^{0},\ldots,{\gamma}_{T-1}^{0}\big{)}^{\prime}=:\psi_{0}.$ Define $\widehat{\alpha}_{N}(\theta):=-\sum_{i=1}^{N-1}\widehat{\alpha}_{i}(\theta)$ and $\widehat{\gamma}_{T}(\theta):=-\sum_{t=1}^{T-1}\widehat{\gamma}_{t}(\theta).$ Note that (2.8) and (2.9) also hold for $i=N$ and $t=T$ respectively.
-Denote $\phi_{0}:=({\alpha}_{1}^{0},\ldots,{\alpha}_{N-1}^{0},{\alpha}_{N}^{0},{\gamma}_{1}^{0},\ldots,{\gamma}_{T-1}^{0},{\gamma}_{T}^{0})^{\prime}$ and $\widehat{\phi}(\theta):=\big{(}\widehat{\alpha}_{1}(\theta),\ldots,\widehat{\alpha}_{N-1}(\theta),\widehat{\alpha}_{N}(\theta),\widehat{\gamma}_{1}(\theta),\ldots,\widehat{\gamma}_{T-1}(\theta),\widehat{\gamma}_{T}(\theta)\big{)}^{\prime}.$ Note that
-$\widehat{l}(\theta)-l(\theta)$ $=$ $l(\theta,\widehat{\phi}(\theta))-l(\theta,\phi_{0})$
-$=$ $\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}\big{[}Y_{it}(\theta)-\frac{1}{2}(\widehat{\alpha}_{i}(\theta)-\alpha_{i}^{0})\big{]}(\widehat{\alpha}_{i}(\theta)-\alpha_{i}^{0})$
-$+\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}\big{[}Y_{it}(\theta)-\frac{1}{2}(\widehat{\gamma}_{t}(\theta)-\gamma_{t}^{0})\big{]}(\widehat{\gamma}_{t}(\theta)-\gamma_{t}^{0})$
-where $Y_{it}(\theta):=Y_{it}-\theta-\alpha_{i}^{0}-\gamma_{t}^{0}.$ Since $Y_{it}$ is independent across $i$ and $t,$ by substituting (2.8) and (2.9), we obtain
-$\mathbb{E}\big{[}\widehat{l}(\theta)-l(\theta)\big{]}$ $=\frac{1}{T}\cdot\frac{1}{2}\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}\mathbb{E}\big{[}\widetilde{Y}_{it}^{2}\big{]}+\frac{1}{N}\cdot\frac{1}{2}\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}\mathbb{E}\big{[}\widetilde{Y}_{it}^{2}\big{]}+o\big{(}1/\sqrt{NT}\big{)}$ (2.10)
-$=\frac{1}{2}\frac{1}{T}+\frac{1}{2}\frac{1}{N}+o\big{(}1/\sqrt{NT}\big{)}.$
-This leads to the conclusion that $\sqrt{NT}\mathbb{E}\big{[}\widehat{l}(\theta)-l(\theta)\big{]}\nrightarrow 0,$ revealing the IPP.
-2.3 Bias Correction
-The expansion in (2.10) from Example 2 motivates the correction of the likelihood $\widehat{l}(\theta)$ by subtracting the first two terms on the right-hand side of (2.10), ensuring that the corrected likelihood converges to $l(\theta)$ at a rate faster than $1/\sqrt{NT}.$ This paper proposes an analytical bias correction procedure by estimating the bias term $1/2\mathbb{E}\{s^{\prime}(\theta)[{H}(\theta)]^{-1}s(\theta)\}$ from (2.7) under $N/T\rightarrow\kappa\in(0,\infty)$.
-Specifically, let $s_{\breve{\alpha}}(\theta)$ represent the first $(N-1)K$ elements of $s(\theta),$ and $s_{\breve{\gamma}}(\theta)$ represent the last $(T-1)K$ elements of $s(\theta),$ where $\breve{\alpha}:=(\alpha_{1}^{\prime},\ldots,\alpha_{N-1}^{\prime})^{\prime}$ and $\breve{\gamma}:=(\gamma_{1}^{\prime},\ldots,\gamma_{T-1}^{\prime})^{\prime}.$ Accordingly, we partition $[\mathbb{E}H(\theta)]^{-1}$ as
-$[\mathbb{E}H(\theta)]^{-1}=:\begin{pmatrix}H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta),&H_{\breve{\alpha}\breve{\gamma}^{\prime}}^{\ast}(\theta)\\ H_{\breve{\gamma}\breve{\alpha}^{\prime}}^{\ast}(\theta),&H_{\breve{\gamma}\breve{\gamma}^{\prime}}^{\ast}(\theta)\end{pmatrix},$
-where $H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)$ and $H_{\breve{\gamma}\breve{\gamma}^{\prime}}^{\ast}\left(\theta\right)$ are $(N-1)K\times(N-1)K$ and $(T-1)K\times(T-1)K$ diagonal blocks, respectively, and $H_{\breve{\alpha}\breve{\gamma}^{\prime}}^{\ast}(\theta)=[H_{\breve{\gamma}\breve{\alpha}^{\prime}}^{\ast}(\theta)]^{\prime}.$ Lemma 1 (see Section 3) shows that $1/2\mathbb{E}\{s^{\prime}(\theta)[{H}(\theta)]^{-1}s(\theta)\}$ can be asymptotically decomposed into two bias components: $1/2\mathbb{E}\{s_{\breve{\alpha}}^{\prime}(\theta)[H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)]s_{\breve{\alpha}}(\theta)\}=:{B}_{\breve{\alpha}}(\theta)$ and $1/2\mathbb{E}\{s_{\breve{\gamma}}^{\prime}(\theta)[H_{\breve{\gamma}\breve{\gamma}^{\prime}}^{\ast}(\theta)]s_{\breve{\gamma}}(\theta)\}=:{B}_{\breve{\gamma}}(\theta),$ resulting from the estimation of the individual and time effects respectively.
-By estimating the two bias components, we construct the corrected likelihood as follows:
-$L(\theta):=\widehat{l}(\theta)+\widehat{B}_{\breve{\alpha}}(\theta)+\widehat{B}_{\breve{\gamma}}(\theta).$ (2.11)
-The explicit forms of $\widehat{B}_{\breve{\alpha}}(\theta)$ and $\widehat{B}_{\breve{\gamma}}(\theta)$ are defined in Section 3. Subsequently, Theorem 2 demonstrates that
-$L(\theta)-l(\theta)=o_{\mathbb{P}}\big{(}1/\sqrt{NT}\big{)},$
-implying that the corrected likelihood no longer suffers from the IPP. Define the MLE of $\theta_{0}$ as the maximizer of $L(\theta):$
-$\widehat{\theta}_{L}=\arg\max_{\theta\in\Theta}L(\theta),$ (2.12)
-where $\Theta$ is a compact subset of $\mathbb{R}^{K}.$ Under mild regularity conditions, this MLE is asymptotically unbiased, and its asymptotic properties align with the standard results for classical MLE in the independent and identically distributed (i.i.d.) case.
+The dynamic nonlinear heterogeneous parameter, DN-HP, models generally suffer from the incidental parameter problem (IPP).
+We now discuss this problem.
+We use $\mathbb{E}$ to denote the expectation w.r.t.
+$\prod_{i=1}^{N}\prod_{t=1}^{T}f(y_{it}|X_{it},\theta_{0}+\alpha_{i}^{0}+\gamma_{t}^{0}),$
+which is the true density evaluated at $\theta_{0}$, given: i) initial values of the predetermined regressors, ii) all strictly exogenous regressors, and iii) the unobserved effect $\phi^{0}$.
+Remark 1 (expectation)
+Our definition of $\mathbb{E}$ depends on the model specification and is the same to Arellano and Bonhomme (2009) and Fernández-Val and Weidner (2016).
+Consider the frequently used autoregressive model with $X_{it}=(Y_{it-1},Z_{it}^{\prime})^{\prime}$ for some strictly exogenous regressor $Z_{it}$.
+It is common that, for each $i=1,\ldots,N$, the density of $Y_{i1}$ is missing and (2.6) becomes a conditional likelihood as in Chapter 5 of Hamilton (1994).
+In this case, the expectation is taken w.r.t.
+$f_{Y_{i2},\ldots,Y_{iT}}\left(\cdot|Y_{i1},Z_{i1},\ldots,Z_{iT},\phi^{0};\theta_{0}\right)$, which is the density of only $Y_{i2},\ldots,Y_{iT}$, conditional on the initial value $Y_{i1}$, strictly exogenous regressors $Z_{i1},\ldots,Z_{iT}$, and the unobserved effects $\phi^{0}$.
+However, when the density of $Y_{i1}$ is indeed specified, so that (2.6) is an exact likelihood, the expectation is taken w.r.t.
+$f_{Y_{i1},\ldots,Y_{iT}}\left(\cdot|Z_{i1},\ldots,Z_{iT},\phi^{0};\theta_{0}\right)$, which is the joint density of all $Y_{i1},\ldots,Y_{iT}$, conditioning on the strictly exogenous regressors $Z_{i1},\ldots,Z_{iT}$ and the unobserved effects $\phi^{0}$ only.
+Following the profiled likelihood framework of Pace and Salvan (2006), we continue our discussion by defining, for a given $\theta$ and each pair of $(N,T)$,
+$\widehat{\psi}(\theta):=\arg\max_{\psi}l(\theta,\psi),\qquad\psi(\theta):=\arg\max_{\psi}\mathbb{E}[l(\theta,\psi)].$ (2.11)
+Here $\psi(\theta)$ is referred to as the pseudo-true value and $l(\theta,\psi(\theta))$ is not subject to the IPP.
+See, e.g., Hahn and Newey (2004) and Arellano and Hahn (2016).
+In practice, however, $\psi(\theta)$ is generally infeasible and the plug-in version $l(\theta,\widehat{\psi}(\theta))$ is used instead.
+Specifically, the MLE of $\theta_{0}$, $\widehat{\theta}:=\arg\max_{\theta}\ell(\theta,\widehat{\psi}(\theta))$.
+The LR and LM test statistics (for hypothesis about $\theta_{0}$) are also constructed using $l(\theta,\widehat{\psi}(\theta))$.
+A fundamental problem is that $l(\theta,\widehat{\psi}(\theta))$ contains a non-negligible asymptotic bias of order $O(1/\sqrt{NT})$, due to the estimation errors in $\widehat{\psi}(\theta)$.
+This is referred to as the IPP in the context of the DN-HP model.
+The consequence is that, $\widehat{\theta}$ and many likelihood-based test statistics deviate from their standard asymptotic distribution, leading to incorrect inferences.
+To demonstrate the non-negligible bias in $l(\theta,\widehat{\psi}(\theta))$, consider a Taylor expansion of $l(\theta,\widehat{\psi}(\theta))$ around $\psi(\theta)$:
+$l(\theta,\widehat{\psi}(\theta))=l(\theta,\psi(\theta))+s^{\prime}(\theta)[\widehat{\psi}(\theta)-\psi(\theta)]+\frac{1}{2}[\widehat{\psi}(\theta)-\psi(\theta)]^{\prime}H(\theta)[\widehat{\psi}(\theta)-\psi(\theta)]+r^{l}(\theta),$ (2.12)
+where $\widehat{\psi}(\theta)-\psi(\theta)$ represents the estimation errors of the fixed effects,
+$s(\theta)$ $:=s(\theta,\psi(\theta)),\qquad s(\theta,\psi):=\partial_{\psi}l(\theta,\psi),$
+$H\left(\theta\right)$ $:=H(\theta,\psi(\theta)),\qquad H(\theta,\psi):=\partial_{\psi\psi^{\prime}}l(\theta,\psi)$
+are the score and Hessian.
+The remainder term $r^{l}(\theta)$ is characterized by Equation (A.17) in Appendix A.
+Here the estimation errors $\widehat{\psi}(\theta)-\psi(\theta)$ can be derived by another Taylor expansion.
+Seeing $s(\theta,\widehat{\psi}(\theta))=0$ as $\widehat{\psi}(\theta)$ is the maximizer, a Taylor expansion of $s(\theta,\widehat{\psi}(\theta))$ around $\psi(\theta)$ gives
+00 0 $=s\left(\theta\right)+H\left(\theta\right)[\widehat{\psi}(\theta)-\psi(\theta)]+r^{s}(\theta),$
+$\widehat{\psi}(\theta)-\psi(\theta)$ $=-[H\left(\theta\right)]^{-1}s\left(\theta\right)-[H\left(\theta\right)]^{-1}r^{s}(\theta),$ (2.13)
+where $r^{s}(\theta)$ is the remainder term of the expansion, which is given by Equation (A.6) in Appendix A. Combining (2.13) and (2.12), and taking expectation, we obtain
+$\sqrt{NT}\mathbb{E}[l(\theta,\widehat{\psi}(\theta))-l(\theta,\psi(\theta))]=-\sqrt{NT}B(\theta)+\sqrt{NT}\mathbb{E}r(\theta),\qquad B(\theta):=\frac{1}{2}\mathbb{E\{}s^{\prime}(\theta)[H(\theta)]^{-1}s(\theta)\},$ (2.14)
+where $r(\theta)$ is a remainder term depending on $r^{l}(\theta)$ and $r^{s}(\theta)$, given by Equation (A.16) in Appendix A.
+By Lemma A.3 in Appendix A, we show that $\mathbb{E}r(\theta)$ is of order $o(1/\sqrt{NT})$, uniformly in $\theta$, as $N,T\rightarrow\infty$ with $N/T\rightarrow\kappa$ for some $0<\kappa<\infty$.
+This implies that $\sqrt{NT}\mathbb{E}r(\theta)=o(1)$ and is negligible.
+$B(\theta)$ is the IPP bias arising from the estimation errors of $\widehat{\psi}(\theta)$.
+As opposite to $\mathbb{E}r(\theta)$, the bias $B(\theta)$ is not negligible, in the sense that $\sqrt{NT}\mathbb{E}B(\theta)\not\rightarrow 0$ under the above asymptotic sequences $\left\{N,T\right\}$ (see Theorem 1).
+Consequently, $\widehat{\theta}$ and many likelihood-based test statistics inherit the bias from the likelihood, leading to invalid inferences.
+We consider the asymptotic sequences $N,T\rightarrow\infty$ with $N/T\rightarrow\kappa$ for some $0<\kappa<\infty$, because $\sqrt{NT}B(\theta)$ explodes under $N/T\rightarrow 0$ or $N/T\rightarrow\infty$.
+We illustrate the IPP and its consequence to the structural parameter estimator in the example below, together with the issue of $N/T\rightarrow\kappa$.
+Example 3 (heterogeneous mean model)
+Following the same setting in Example 2 but let $\varepsilon_{it}$ be normal with mean 00 and variance $\sigma_{0}>0$.
+In the current example, $\theta_{0}=(\mu_{0},\sigma_{0})^{\prime}$ is the structural parameter of interests.
+This example is a modified version of Neyman and Scott (1948) and is similar to Fernández-Val and Weidner (2016), except that the mean of $Y_{it}$ is $\theta_{0}+\alpha_{i}^{0}+\gamma_{t}^{0}$, conditional on $\phi^{0}$.
+Define [^4] In this simple example, $\varepsilon_{it}$ is independent across both $i$ and $t$.
+Therefore, $\mathbb{E}$ is simply w.r.t.
+the normal density having mean $\mu^{0}+\alpha_{i}^{0}+\gamma_{t}^{0}$ and variance $\sigma^{0}$.
+$\widetilde{Y}_{it}:=Y_{it}-\mathbb{E}Y_{it}$.
+Following the same steps in Example 2, we can derive
+$\widehat{\alpha}_{i}-\alpha_{i}^{0}$ $=\frac{1}{T}\sum_{t=1}^{T}\widetilde{Y}_{it}-\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}\widetilde{Y}_{it},\qquad i=1,\ldots,N-1,$
+$\widehat{\gamma}_{t}-\gamma_{t}^{0}$ $=\frac{1}{N}\sum_{i=1}^{N}\widetilde{Y}_{it}-\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}\widetilde{Y}_{it},\qquad t=1,\ldots,T-1,$ (2.15)
+where $\alpha_{i}^{0}$ and $\gamma_{t}^{0}$ can be obtained by maximizing $\mathbb{E}l(\mu,\psi)$ in the same way. Denote additionally
+$\widehat{\alpha}_{N}:=-\sum_{i=1}^{N-1}\widehat{\alpha}_{i},\qquad\alpha_{N}^{0}:=-\sum_{i=1}^{N-1}\alpha_{i}^{0},\qquad\widehat{\gamma}_{T}:=-\sum_{t=1}^{T-1}\widehat{\gamma}_{t},\qquad\gamma_{T}^{0}:=-\sum_{t=1}^{T-1}\gamma_{t}^{0}.$
+It follows that
+$\widehat{\alpha}_{N}-\alpha_{N}^{0}=\frac{1}{T}\sum_{t=1}^{T}\widetilde{Y}_{Nt}-\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}\widetilde{Y}_{it},\qquad\widehat{\gamma}_{T}-\gamma_{T}^{0}=\frac{1}{N}\sum_{i=1}^{N}\widetilde{Y}_{iT}-\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}\widetilde{Y}_{it}.$ (2.16)
+Now, for $\psi^{0}:=({\alpha}_{1}^{0},\ldots,{\alpha}_{N-1}^{0},{\gamma}_{1}^{0},\ldots,{\gamma}_{T-1}^{0})^{\prime}$, $\widehat{\psi}:=(\widehat{\alpha}_{1},\ldots,\widehat{\alpha}_{N-1},\widehat{\gamma}_{1},\ldots,\widehat{\gamma}_{T-1})^{\prime}$, and $Y_{it}(\mu):=Y_{it}-\mu-\alpha_{i}^{0}-\gamma_{t}^{0}$, we can calculate
+$l(\theta,\widehat{\psi})-l(\theta,\psi^{0})$ $=\frac{1}{\sigma}\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}\left(Y_{it}(\mu)-\frac{1}{2}(\widehat{\alpha}_{i}-\alpha_{i}^{0})\right)(\widehat{\alpha}_{i}-\alpha_{i}^{0})$
+$\quad\quad+\frac{1}{\sigma}\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}\left(Y_{it}(\mu)-\frac{1}{2}(\widehat{\gamma}_{t}-\gamma_{t}^{0})\right)(\widehat{\gamma}_{t}-\gamma_{t}^{0}),$ (2.17)
+where $\theta=(\mu,\sigma)^{\prime}$ with $\sigma>0$. Substituting (2.15) and (2.16) into (2.17), we obtain, after some algebra,
+$l(\theta,\widehat{\psi})-l(\theta,\psi^{0})$ $=\frac{1}{2\sigma}\frac{1}{NT^{2}}\sum_{i=1}^{N}\sum_{t=1}^{T}\sum_{s=1}^{T}\widetilde{Y}_{it}\widetilde{Y}_{is}+\frac{1}{2\sigma}\frac{1}{N^{2}T}\sum_{i=1}^{N}\sum_{j=1}^{N}\sum_{t=1}^{T}\widetilde{Y}_{it}\widetilde{Y}_{jt}$
+$\quad\quad-\frac{2}{\sigma}\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}\left(\mu-\mu^{0}\right)\widetilde{Y}_{it}+O_{p}(1/(NT)).$
+Here, $\mathbb{E}\widetilde{Y}_{it}=0$ by definition and $\mathbb{E}\widetilde{Y}_{it}\widetilde{Y}_{js}=0$, for $i\neq j$ or $t\neq s$, by independence.
+Therefore, calculating the expectation, we arrive at
+$\mathbb{E[}l(\theta,\widehat{\psi})-l(\theta,\psi^{0})]=-B(\theta)+O(1/(NT)),\qquad B(\theta)=-\left(\frac{1}{T}+\frac{1}{N}\right)\frac{\sigma^{0}}{2\sigma}.$
+The term $B(\theta)$ is the IPP bias. Apparently,
+$\sqrt{NT}B(\theta)=-\left(\sqrt{\frac{N}{T}}+\sqrt{\frac{T}{N}}\right)\frac{\sigma^{0}}{2\sigma}\not\rightarrow 0$
+as $N,T\rightarrow\infty$ and $N/T\rightarrow\kappa$.
+When $N/T\rightarrow 0$ (or $N/T\rightarrow\infty$), $\sqrt{T/N}\rightarrow\infty$ (or $\sqrt{N/T}\rightarrow\infty$) and $\sqrt{NT}B(\theta)$ explodes.
+The bias term $B(\theta)$ causes the variance estimator to be biased. We illustrate this consequence here. Observe that
+$l(\theta,\widehat{\psi})$ $=-\frac{1}{2}\log(2\pi)-\frac{1}{2}\log\sigma-\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}\frac{1}{2\sigma}(Y_{it}-\mu-\widehat{\alpha}_{i}-\widehat{\gamma}_{t})^{2},$
+$\partial_{\sigma}l(\theta,\widehat{\psi})$ $=-\frac{1}{2\sigma}+\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}\frac{1}{2\sigma^{2}}(Y_{it}-\mu-\widehat{\alpha}_{i}-\widehat{\gamma}_{t})^{2}.$
+Following the same steps in Example 2, we obtain exactly the same estimator of $\mu_{0}$, $\widehat{\mu}$, as in Equation (2.10); and it is independent of $\sigma$.
+Therefore, equating $\partial_{\sigma}l(\theta,\widehat{\psi})=0$, substituting $\mu$ with $\widehat{\mu}$, and solving for $\sigma$, the estimator of $\sigma^{0}$ is
+$\widehat{\sigma}$ $=\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}(Y_{it}-\widehat{\mu}-\widehat{\alpha}_{i}-\widehat{\gamma}_{t})^{2}$
+$=\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}[Y_{it}-\mu_{0}-\alpha_{i}^{0}-\gamma_{t}^{0}-(\widehat{\mu}-\mu_{0})-(\widehat{\alpha}_{i}-\alpha_{i}^{0})-(\widehat{\gamma}_{t}-\gamma_{t}^{0})]^{2},$
+where it is not difficult to see
+$\widehat{\mu}-\mu_{0}=\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}\widetilde{Y}_{it}.$
+Therefore, seeing $\widetilde{Y}_{it}=Y_{it}-\mu_{0}-\alpha_{i}^{0}-\gamma_{t}^{0}$ by definition, we finally obtain
+$\widehat{\sigma}$ $=\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}\widetilde{Y}_{it}^{2}-\frac{1}{NT^{2}}\sum_{i=1}^{N}\sum_{t=1}^{T}\sum_{s=1}^{T}\widetilde{Y}_{it}\widetilde{Y}_{is}-\frac{1}{N^{2}T}\sum_{i=1}^{N}\sum_{j=1}^{N}\sum_{t=1}^{T}\widetilde{Y}_{it}\widetilde{Y}_{jt}$
+$\quad\quad+\frac{1}{N^{2}T^{2}}\sum_{i=1}^{N}\sum_{j=1}^{N}\sum_{t=1}^{T}\sum_{s=1}^{T},\widetilde{Y}_{it}\widetilde{Y}_{js}$
+from where we conclude that
+$\mathbb{E}\widehat{\sigma}=\sigma_{0}-\frac{\sigma_{0}}{T}-\frac{\sigma_{0}}{N}+\frac{\sigma_{0}}{NT},$
+leading to
+$\sqrt{NT}(\mathbb{E}\widehat{\sigma}-\sigma_{0})\rightarrow-\kappa\sigma_{0}-\frac{\sigma_{0}}{\kappa}$
+as $N,T\rightarrow\infty$ with $N/T\rightarrow\kappa$.
+While $\widehat{\mu}$ is not affected by the IPP, the variance estimator $\widehat{\sigma}$ contains two non-negligible bias terms, $-\sigma_{0}/T$ and $-\sigma_{0}/N$.
+These bias terms cause $\sqrt{NT}(\widehat{\sigma}-\sigma_{0})$ to be centered at $-\kappa\sigma_{0}-\sigma_{0}/\kappa\neq 0$.
+3 Bias Correction and Asymptotic Theory
+The bias expansion of $l(\theta,\widehat{\psi}(\theta))$ in Equation (2.14) and Example 3 above motivate our approach of correcting the likelihood $l(\theta,\widehat{\psi}(\theta))$ by adding $B(\theta)$.
+The corrected likelihood can be viewed as an approximation to the IPP-free infeasible likelihood $l(\theta,\psi(\theta))$, with approximation error $o(1/\sqrt{NT})$.
+In Theorem 1 below we formalize the expansion in Equation (2.14).
+Specifically, we show that $B(\theta)$ can be further decomposed into $B(\theta)=B_{\alpha}(\theta)+B_{\gamma}(\theta)+o(1/\sqrt{NT})$, in which $B_{\alpha}(\theta)$ appears because of the estimation errors of the individual-specific heterogeneities $\alpha_{i}^{0}$, while $B_{\gamma}(\theta)$ is present due to the estimation errors of the time-specific heterogeneities $\gamma_{t}^{0}$.
+The exact formulas for both terms are given below.
+Theorem 2 follows to establish the corrected likelihood as
+$L(\theta):=\widehat{l}(\theta)+\widehat{B}_{\alpha}(\theta)+\widehat{B}_{\gamma}(\theta)$ (3.1)
+where we denote $\widehat{l}(\theta):=l(\theta,\widehat{\psi}(\theta))$ for simplicity from now on.
+$\widehat{B}_{\alpha}(\theta)$ and $\widehat{B}_{\gamma}(\theta)$ are estimates of $B_{\alpha}(\theta)$ and $B_{\gamma}(\theta)$, respectively.
+The analytical formulas for $\widehat{B}_{\alpha}(\theta)$ and $\widehat{B}_{\gamma}(\theta)$ are provided below.
+We then show that
+$L(\theta)-l(\theta)=o_{\mathbb{P}}(1/\sqrt{NT}),$
+where we use $l(\theta):=l(\theta,\psi(\theta))$ for simplicity below.
+This implies that $L(\theta)$ does not suffer from the IPP.
+The corrected estimator of $\theta_{0}$, and the LR, LM, and Wald test statistics, may subsequently be derived from $L\left(\theta\right)$, instead of $\widehat{l}(\theta)$.
+Specifically, the corrected estimator of $\theta_{0}$ is
+$\widehat{\theta}_{L}:=\arg\max_{\theta}L(\theta).$
+Under mild regularity conditions, $\widehat{\theta}_{L}$ is asymptotically unbiased, and its asymptotic properties align with the standard results of the classical MLE without the nuisance parameters.
 Moreover, the test statistics constructed from the corrected likelihood such as the Wald, LR, and LM also follow the limiting $\chi^{2}$ distribution under the null hypothesis.
-Since the corrected likelihood depends implicitly on $\widehat{\psi}(\theta),$ which maximizes the reparameterized likelihood function $l^{\ast}(\theta,\psi),$ obtaining $\widehat{\theta}_{L}$ from $L(\theta)$ involves a two-layer optimization.
-In the inner layer, $\widehat{\psi}(\theta)$ is found by maximizing $l^{\ast}(\theta,\psi)$ using the Newton-Raphson algorithm with analytical derivatives.
-In the outer layer, we maximize $L(\theta)$ with respect to $\theta,$ substituting $\widehat{\psi}(\theta)$ and using numerical derivatives.
-3 Asymptotic Theory
-In this section, we develop the asymptotic theory for our bias correction procedure, as well as for the estimation and testing based on the corrected likelihood function.
-Let $\phi(\theta):=D^{\prime}\psi(\theta)$ and $\widehat{\phi}(\theta):=D^{\prime}\widehat{\psi}(\theta).$ The following assumptions are introduced.
+These results are formally stated in Corollaries 1 and 2, respectively.
+Remark 2 (computation)
+The computation of $\widehat{\theta}_{L}$ involves a two-layer maximization but can be done efficiently.
+In the inner layer, $\widehat{\psi}(\theta)$ is obtained by maximizing $l(\theta,\psi)$, keeping $\theta$ fixed.
+This is large-scale maximization problem and a few strategies can be exploited to expedite the procedure:
+  1. 1.
+    For many commonly used models, the first and second derivatives of $l(\theta,\psi)$ have clear analytical expressions, so that the standard Newton-Raphson algorithm can be employed to obtain $\widehat{\psi}(\theta)$.
+  2. 2.
+    Many models (specifically, generalized linear models) are equipped with linear indices, as in Example 1.
+In that case, the chain-rule-based technique in Sun and Dhaene (2019) can be used to achieve a fast calculation of the derivatives.
+The computational complexity of such a technique only grows with the number of linear indices, not the dimension of $\psi$, even without analytical expressions for the derivatives of $l(\theta,\psi)$.
+  3. 3.
+    The derivatives of the reparameterized $l(\theta,\psi)$ can be identically constructed from the derivatives of the original likelihood $\ell(\theta,\phi)$.
+Specifically, the reparameterization implies
+    $\partial_{\psi}l(\theta,\psi)=D\partial_{\phi}\ell(\theta,\phi),\qquad\partial_{\psi\psi^{\prime}}l(\theta,\psi)=D\partial_{\phi\phi^{\prime}}\ell(\theta,\phi)D^{\prime},$
+    holding true for every $\phi=D^{\prime}\psi$ and every $\theta$.
+In the outer layer where $L(\theta)$ is maximized w.r.t.
+$\theta$, it can be tedious to obtain analytical formulas for the derivatives of the corrected likelihood $L(\theta)$.
+However, the dimension of $\theta$ is usually small, so that invoking the Newton-Raphson algorithm with numerical derivatives generally does not suffer from too much computational pressure, provided that the inner-layer maximization is done efficiently.
+Now, we proceed to our bias correction procedure on the likelihood and present the related asymptotic theory.
+Denote $\widehat{\phi}(\theta):=D^{\prime}\widehat{\psi}(\theta)$ and $\phi(\theta):=D^{\prime}\psi(\theta)$, which are essentially the “unparameterized” counterparts of $\widehat{\psi}(\theta)$ and $\psi(\theta)$, respectively.
+We impose the following assumption.
 Assumption 1
-  • (i) Suppose $N/T\rightarrow\kappa$ for some $0<\kappa<\infty$ as $N$ and $T\to\infty.$
-  • (ii) Let $\Theta$ be a compact subset of $\mathbb{R}^{K}$ with $\theta_{0}$ in its interior.
-For each $\theta\in\Theta,$ there exists a compact subset $\Lambda_{\theta}$ that contains both $\widehat{\phi}(\theta)$ and $\phi(\theta)$ in its interior, such that $l(\theta,\phi)$ is three-times continuously differentiable with respect to $\phi\in\Lambda_{\theta}$ almost surely.
-Let $w_{it}=(Y_{it},X_{it}^{\prime})^{\prime},$ and assume there exists a function $g(w_{it})>0$ independent of $\theta$ and $\phi,$ such that
-    $\sup_{\theta\in\Theta}\sup_{\phi\in\Lambda_{\theta}}\left|\partial_{\phi_{1}\ldots\phi_{S}}l_{i,t}(\theta,\phi)\right|<g(w_{it}),$
-    where $\phi_{s}$ represents any element of $\phi$ and $S\in\{0,1,2,3\}$ ($S=0$ understood as no differentiation), and $\sup_{N,T}\max_{1\leq i\leq N,1\leq t\leq T}\mathbb{E}\left[g^{\eta}(w_{it})\right]<\infty,$ for some $\eta>2.$
-  • (iii) For each $k=1,\ldots,K$, $\sum_{i=1}^{N}\alpha_{ik}=\sum_{t=1}^{T}\gamma_{tk}=0$.
-  • (iv) Given any fixed $\theta\in\Theta,$ for all $\phi\in\Lambda_{\theta}$ if $\phi\neq\phi(\theta),$ $\mathbb{P}\big{(}l(\theta,\phi)\neq l(\theta,\phi(\theta))\big{)}>0.$
-  • (v) Conditional on $\phi(\theta),$$\left\{w_{it}\right\}$ are independent across $i$.
-Denote $\mathcal{A}_{it}$ and $\mathcal{B}_{it}$ to be the $\sigma$-algebras generated, respectively, by $\left\{w_{is}:1\leq s\leq t\right\}$ and $\left\{w_{is}:t\leq s\leq T\right\}$.
-$\left\{w_{it}\right\}$ is conditional strong mixing with mixing coefficient
-    $a_{i}\left(m\right):=\sup_{\theta\in\Theta}\sup_{t\geq 1}\sup_{A\in\mathcal{A}_{it},B\in\mathcal{B}_{it+m}}\left|\mathbb{P}\left(A\cap B\mid\phi(\theta)\right)-\mathbb{P}\left(A\mid\phi(\theta)\right)\mathbb{P}\left(B\mid\phi(\theta)\right)\right|,$
-    and, $\sup_{N}\max_{1\leq i\leq N}\sum_{m=0}^{\infty}\left[a_{i}\left(m\right)\right]^{1-2/\eta}<\infty,$ for some $\eta>2.$
-  • (vi) For $\theta\in\Theta,$ $\sqrt{NT}\,\mathbb{E}\left[{H}(\theta)\right]$ is nonsingular when $N$ and $T$ are large.
-Assumption 1 requires that both $N$ and $T$ are large and proportional, which is standardis standard in the literature.
-Assumption 1 imposes smoothness and boundedness conditions on the likelihood.
-Specifically, it requires that the likelihood and its derivatives for unit $i$ at time $t$ are unformly bounded by a measurable function whose moments of order greater than 2 exist.
-Assumption 1 is the identification condition for the incidental parameters (fixed effects) as they enter the likelihood additively.
-Given Assumption 1, Assumption 1 ensures that for each fixed $\theta,$ $\psi(\theta)$ defined in (2.2) is the unique maximizer.
-Assumption 1 assumes that the data are independent across units $i$ and conditional strong mixing across time $t.$ Under Assumptions 1 and 1, we can apply Davydov’s inequality to control the convergence rate of certain expectations of summations.
-Assumption 1 requires that the identified incidental parameter Hessian is invertible.
-The following lemma states the bias induced by IPP, as shown in (2.7).
-Lemma 1
-Under Assumption 1, we have
-$\mathbb{E}\left\{s^{\prime}(\theta)[{H}\left(\theta\right)]^{-1}s(\theta)\right\}=\mathbb{E}\left\{s_{\breve{\alpha}}^{\prime}(\theta)H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)s_{\breve{\alpha}}(\theta)\right\}+\mathbb{E}\left\{s_{\breve{\gamma}}^{\prime}(\theta)H_{\breve{\gamma}\breve{\gamma}^{\prime}}^{\ast}(\theta)s_{\breve{\gamma}}(\theta)\right\}+o\big{(}1/\sqrt{NT}\big{)},$
-where $o(\cdot)$ is uniform in $\theta,$ $s_{\breve{\alpha}}(\theta)$ represent the first $(N-1)K$ elements of $s(\theta),$ and $s_{\breve{\gamma}}(\theta)$ represent the last $(T-1)K$ elements of $s(\theta).$
-The following theorem presents the bias terms of $\mathbb{E}[\widehat{l}(\theta)]$ to approximate $\mathbb{E}[l(\theta)].$
+  • (i) Suppose $N/T\rightarrow\kappa$ for some $0<\kappa<\infty$ as $N,T\rightarrow\infty$.
+[…]
+    $\sup_{\theta\in\Theta}\sup_{\phi\in\Phi}\left|\partial_{\phi_{1}\cdots\phi_{S}}l_{i,t}(\theta,\phi)\right|<g(w_{it}),$
+    where $\phi_{s}$ represents any element of $\phi$ and $S\in\{0,1,2,3\}$ ($S=0$ is understood as not taking derivatives), and
+    $\sup_{N,T}\max_{1\leq i\leq N,1\leq t\leq T}\mathbb{E}\{[g(w_{it})]^{\eta}\}<\infty,$
+    for some $\eta>2$.
+  • (iii) For each $k=1,\ldots,K$, $\sum_{i=1}^{N}\alpha_{k,i}^{0}=\sum_{t=1}^{T}\gamma_{k,t}^{0}=0$, where $\alpha_{k,i}^{0}$ and $\gamma_{k,t}^{0}$ are the $k$-th component of $\alpha_{i}^{0}$ and $\gamma_{t}^{0}$, respectively.
+  • (iv)  For each $\theta\in\Theta$, $\mathbb{P}[l(\theta,\psi)\neq l(\theta,\psi(\theta))]>0$ for every $\psi$ such that $D^{\prime}\psi\in\Phi$ and $\psi\neq\psi(\theta)$.
+  • (v) Conditional on $\phi^{0}$, $\{w_{it}\}$ are independent across $i$ and conditionally strong mixing with mixing coefficient
+    $a_{i}(m):=\sup_{t\geq 1}\sup_{A\in\mathcal{A}_{it},B\in\mathcal{B}_{it+m}}\left|\mathbb{P}(A\cap B|\phi^{0})-\mathbb{P}(A|\phi^{0})\mathbb{P}(B|\phi^{0})\right|$
+    such that, for $\eta>2$ as above,
+    $\sup_{N}\max_{1\leq i\leq N}\sum_{m=0}^{\infty}[a_{i}(m)]^{1-2/\eta}<\infty,$
+    where $\mathcal{A}_{it}$ and $\mathcal{B}_{it}$ are the $\sigma$-algebras generated by $\{w_{is}:1\leq s\leq t\}$ and $\{w_{is}:t\leq s\leq T\}$, respectively.
+  • (vi) The matrix has $\sqrt{NT}\mathbb{E}H(\theta)$ has eigenvalues $\lambda_{p}(\theta)$ for $p=1,\ldots,(N+T-2)K$ which satisfy
+    $\sup_{N>N_{0},T>T_{0}}\max_{1\leq p\leq(N+T-2)K}\sup_{\theta\in\Theta}\lambda_{p}(\theta)<0$
+    for some integers $N_{0}$ and $T_{0}$.
+Here Assumption 1 is the same to Fernández-Val and Weidner (2016).
+It is used to ensure that the bias $\sqrt{NT}B(\theta)$ does not explode as $N,T\rightarrow\infty$, as discussed above.
+Assumption 1 is standard in the literature.
+See, e.g., Hahn and Newey (2004), Hahn and Kuersteiner (2011), and Arellano and Hahn (2016).
+It simply requires that the likelihood function $l_{i,t}(\theta,\phi)$ is differentiable in $\phi$ at least three times.
+The term $B(\theta)$ in (2.14) depends on the first and second derivatives, and the remainder $r(\theta)$ depends on the third.
+We require $l_{i,t}(\theta,\phi)$ and its derivatives to be uniformly bounded by a function $g(w_{it})$ whose moment exists to an order greater than the second.
+This is used to ensure, amongst other things, i) the interchangeability of $\mathbb{E}$ and $\partial$; and ii) the existence of moments, such as $\mathbb{E\{}s^{\prime}(\theta)[H(\theta)]^{-1}s(\theta)\}$ in (2.14), for every $\theta\in\Theta$.
+It is also used to establish the uniform rate of convergence, over $\theta\in\Theta$, of the corrected likelihood $L(\theta)$.
+The uniform convergence of $L(\theta)$ guarantees that the bias correction effect carries over to both the estimators and likelihood-based test statistics.
+Assumption 1 states that the true effects $\alpha_{i}^{0}$ and $\gamma_{t}^{0}$ sums up to 00 over $i=1,\ldots,N$ and $t=1,\ldots,T$.
+Given the additive structure of $\theta_{0}+\alpha_{i}^{0}+\gamma_{t}^{0}$ in (2.1), this condition is without loss of generality because we merely let $\theta_{0}$ capture the average effects of $\alpha_{i}^{0}$ and $\gamma_{t}^{0}$.
+Assumption 1 is an identification condition for $\psi(\theta)$.
+It is standard for maximum likelihood models.
+See, e.g., Newey and McFadden (1994).
+Assumption 1 is common in the field of dynamic panels.
+See, for instance, Hahn and Kuersteiner (2007), Fernández-Val and Weidner (2016), and Dhaene and Jochmans (2015).
+Notice that, the argument is conditional on the true heterogeneous effects $\phi^{0}$.
+Without conditioning, $w_{it}$ may not be, for instance, independent over $i$, because $w_{it}$ and $w_{jt}$ may depend on the same $\gamma_{t}^{0}$.
+In addition, we only require the mixing coefficient $a_{i}(m)$ to decay at a polynomial rate such that $[a_{i}(m)]^{1-2/\eta}$ is summable as $m\rightarrow\infty$.
+The mixing condition is used to guarantee that certain autocovariances which appear in our proof are summable by Davydov’s inequality.
+Notice that we are not requiring $w_{it}$ to be identically distributed over $i$ or stationary over $t$.
+These are generally not possible because $w_{it}$ depends on $\alpha_{i}^{0}$ and $\gamma_{t}^{0}$.
+Finally, Assumption 1 simply requires that the normalized expected Hessian $\sqrt{NT}\mathbb{E}H(\theta)$ has eigenvalues $\lambda_{p}(\theta)$ that are strictly less than 00 for all $N>N_{0}$ and $T>T_{0}$.
+This ensures that $\mathbb{E\{}s^{\prime}(\theta)[H(\theta)]^{-1}s(\theta)\}$ is well defined for large $N$ and $T$.
+This is a mild condition in that we only impose it on the expected Hessian $\mathbb{E}H(\theta)$ of the reparameterized likelihood.
+We state the condition with “less than 00” because the Hessian in a maximum likelihood model is usually negative definite.
+Next, partition [^5] Note that $s(\theta)=s(\theta)-\mathbb{E}s\left(\theta\right)$ because $\mathbb{E}s\left(\theta\right)=0$ by definition.
+We keep $\mathbb{E}s\left(\theta\right)$ here for the convenience of constructing the corrected likelihood.
+$s(\theta)-\mathbb{E}s\left(\theta\right)$ and $[\mathbb{E}H(\theta)]^{-1}$ as
+$s(\theta)-\mathbb{E}s\left(\theta\right)=:\left(\begin{array}[c]{c}\widetilde{s}_{\alpha}(\theta)\\ \widetilde{s}_{\gamma}(\theta)\end{array}\right),\qquad[\mathbb{E}H(\theta)]^{-1}=:\left(\begin{array}[c]{cc}H_{\alpha\alpha}^{\mathcal{\ast}}(\theta)&H_{\alpha\gamma}^{\mathcal{\ast}}(\theta)\\ H_{\gamma\alpha}^{\mathcal{\ast}}(\theta)&H_{\gamma\gamma}^{\mathcal{\ast}}(\theta)\end{array}\right),$ (3.2)
+where $\widetilde{s}_{\alpha}(\theta)$ is $(N-1)K\times 1$ and $\widetilde{s}_{\gamma}(\theta)$ is $(T-1)K\times 1$ while $H_{\alpha\alpha}^{\mathcal{\ast}}(\theta)$ is $(N-1)K\times(N-1)K$ and $H_{\gamma\gamma}^{\ast}(\theta)$ is $(T-1)K\times(T-1)K$.
+The off-diagonal blocks $H_{\alpha\gamma}^{\ast}(\theta)$ and $H_{\gamma\alpha}^{\mathcal{\ast}}(\theta)$ are defined accordingly.
+Further denote
+$S_{\alpha\alpha}(\theta):=\mathbb{E\{}\widetilde{s}_{\alpha}(\theta)\mathbb{[}\widetilde{s}_{\alpha}(\theta)]^{\prime}\},\qquad S_{\gamma\gamma}(\theta):=\mathbb{E\{}\widetilde{s}_{\gamma}(\theta)\mathbb{[}\widetilde{s}_{\gamma}(\theta)]^{\prime}\}.$
+The next theorem formally states the asymptotic expansion of $\widehat{l}(\theta)$.
 Theorem 1
-Under Assumption 1, as $N$ and $T\rightarrow\infty,$ we have
-$\mathbb{E}[l(\theta)]=\mathbb{E}[\widehat{l}(\theta)]+B_{\breve{\alpha}}(\theta)+B_{\breve{\gamma}}(\theta)+o\big{(}1/\sqrt{NT}\big{)},$
-where $o(\cdot)$ is uniform over $\theta\in\Theta.$ The bias terms $B_{\breve{\alpha}}(\theta)$ and $B_{\breve{\gamma}}(\theta)$ are defined as follows:
-$B_{\breve{\alpha}}(\theta):=\frac{1}{2}\mathbb{E}\big{[}{s}_{\breve{\alpha}}^{\prime}(\theta)H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta){s}_{\breve{\alpha}}(\theta)\big{]},\quad B_{\breve{\gamma}}(\theta):=\frac{1}{2}\mathbb{E}\big{[}{s}_{\breve{\gamma}}^{\prime}(\theta){H}_{\breve{\gamma}\breve{\gamma}^{\prime}}^{\ast}(\theta){s}_{\breve{\gamma}}(\theta)\big{]}.$
-Since $B_{\breve{\alpha}}(\theta)=O\big{(}1/\sqrt{NT}\big{)}$ and $B_{\breve{\gamma}}(\theta)=O\big{(}1/\sqrt{NT}\big{)}$ as $N$ and $T\to\infty,$ these bias terms introduce a non-zero limiting center to the MLE based on $\widehat{l}(\theta).$ Therefore, it is necessary to eliminate the effect of this bias.
-In the following, we estimate these bias terms and add them back to $\widehat{l}(\theta).$ Let $\psi(\theta)=:(\alpha^{\prime}_{1}(\theta),\ldots,\alpha^{\prime}_{N-1}(\theta),\gamma^{\prime}_{1}(\theta),\ldots,\gamma^{\prime}_{T-1}(\theta))^{\prime},$ $-\sum_{i=1}^{N-1}\alpha_{i}(\theta)=:\alpha_{N}(\theta),$ and $-\sum_{t=1}^{T-1}\gamma_{t}(\theta)=:\gamma_{T}(\theta).$ Then, $\phi(\theta)=(\alpha^{\prime}_{1}(\theta),\ldots,\alpha^{\prime}_{N}(\theta),\gamma^{\prime}_{1}(\theta),\ldots,\gamma^{\prime}_{T}(\theta))^{\prime}.$ Noting that $\mathbb{E}[s_{\breve{\alpha}}(\theta)]=0$ and $\mathbb{E}[s_{\breve{\gamma}}(\theta)]=0,$ we have
-$B_{\breve{\alpha}}(\theta)$ $=\frac{1}{2}\mathbb{E}\big{\{}[s_{\breve{\alpha}}(\theta)-\mathbb{E}s_{\breve{\alpha}}(\theta)]^{\prime}H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}\left(\theta\right)[s_{\breve{\alpha}}(\theta)-\mathbb{E}s_{\breve{\alpha}}(\theta)]\big{\}}$
-$=\frac{1}{2}\text{{trace}}\big{\{}H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)\Sigma^{\breve{\alpha}}(\theta)\big{\}},$ (3.1)
-where
-$\Sigma^{\breve{\alpha}}(\theta)$ $:=\Sigma_{1}^{\breve{\alpha}}(\theta)+\Sigma_{2}^{\breve{\alpha}}(\theta),$
-$\Sigma_{1}^{\breve{\alpha}}(\theta)$ $:=\text{{diag}}\big{\{}\mathbb{E}\big{[}\widetilde{s}_{1}^{\alpha}\left(\theta\right)\left(\widetilde{s}_{1}^{\alpha}\left(\theta\right)\right)^{\prime}\big{]},\ldots,\mathbb{E}\big{[}\widetilde{s}_{N-1}^{\alpha}\left(\theta\right)\left(\widetilde{s}_{N-1}^{\alpha}\left(\theta\right)\right)^{\prime}\big{]}\big{\}},$
-$\Sigma_{2}^{\breve{\alpha}}(\theta)$ $:=(\iota_{N-1}\iota_{N-1}^{\prime})\otimes\mathbb{E}\big{[}\widetilde{s}_{N}^{\alpha}\left(\theta\right)\left(\widetilde{s}_{N}^{\alpha}\left(\theta\right)\right)^{\prime}\big{]},$
-$\widetilde{s}_{i}^{\alpha}\left(\theta\right)$ :=$:=\big{\{}\partial_{\alpha_{i}}l\left(\theta,\phi\right)-\mathbb{E}[\partial_{\alpha_{i}}l\left(\theta,\phi\right)]\big{\}}\big{|}_{\phi=\phi\left(\theta\right)},\quad\text{$i=1,\ldots,N.$}$
-Similarly, we have
-$B_{\breve{\gamma}}(\theta)$ $=\frac{1}{2}\mathbb{E}\big{\{}[s_{\breve{\gamma}}(\theta)-\mathbb{E}s_{\breve{\gamma}}(\theta)]^{\prime}H_{\breve{\gamma}\breve{\gamma}^{\prime}}^{\ast}\left(\theta\right)[s_{\breve{\gamma}}(\theta)-\mathbb{E}s_{\breve{\gamma}}(\theta)]\big{\}}$
-$=\frac{1}{2}\text{{trace}}\big{\{}H_{\breve{\gamma}\breve{\gamma}^{\prime}}^{\ast}(\theta)\Sigma^{\breve{\gamma}}(\theta)\big{\}},$ (3.2)
-where
-$\Sigma^{\breve{\gamma}}(\theta)$ $:=\mathbb{E}\big{\{}[s_{\breve{\gamma}}(\theta)-\mathbb{E}s_{\breve{\gamma}}(\theta)][s_{\breve{\gamma}}(\theta)-\mathbb{E}s_{\breve{\gamma}}(\theta)]^{\prime}\big{\}}=\mathbb{E}\big{[}\widetilde{S}_{\breve{\gamma}}(\theta)(\widetilde{S}_{\breve{\gamma}}(\theta))^{\prime}\big{]}$
-$\widetilde{S}_{\breve{\gamma}}(\theta)$ $:=\big{(}(\widetilde{s}_{1}^{\gamma}(\theta))^{\prime},\ldots,(\widetilde{s}_{T-1}^{\gamma}(\theta))^{\prime}\big{)}^{\prime}-\iota_{T-1}\otimes\widetilde{s}_{T}^{\gamma}(\theta)$
-$\widetilde{s}_{t}^{\gamma}(\theta)$ :=$:=\big{\{}\partial_{\gamma_{t}}l\left(\theta,\phi\right)-\mathbb{E}[\partial_{\gamma_{t}}l\left(\theta,\phi\right)]\big{\}}\big{|}_{\phi=\phi(\theta)}.$
-Next, we estimate the two bias terms.
-We first estimate the Hessian matrix $H(\theta)=H(\theta,\psi(\theta))$ by substituting $\widehat{\psi}(\theta).$ Define $\widehat{H}(\theta):=H(\theta,\widehat{\psi}(\theta)).$ We have $\widehat{H}(\theta)=D\big{[}\partial_{\phi\phi^{\prime}}l(\theta,\phi)\mid_{\phi=\widehat{\phi}(\theta)}\big{]}D^{\prime}$ with $\widehat{\phi}(\theta)=D^{\prime}\widehat{\psi}(\theta).$ Following (3) and (3), we estimate ${B}_{\breve{\alpha}}(\theta)$ and ${B}_{\breve{\gamma}}(\theta)$ by
-$\widehat{B}_{\breve{\alpha}}(\theta):=\frac{1}{2}\operatorname*{trace}\left(\widehat{H}_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}\left(\theta\right)\widehat{\Sigma}_{\tau}^{\breve{\alpha}}(\theta)\right)\quad\text{and}\quad\widehat{B}_{\breve{\gamma}}(\theta):=\frac{1}{2}\operatorname*{trace}\left(\widehat{H}_{\breve{\gamma}\breve{\gamma}^{\prime}}^{\ast}(\theta)\widehat{\Sigma}_{\nu}^{\breve{\gamma}}(\theta)\right),$
-where $\widehat{H}_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)$ and $\widehat{H}_{\breve{\gamma}\breve{\gamma}^{\prime}}^{\ast}(\theta)$ are the first $(N-1)K\times(N-1)K$ and last $(T-1)K\times(T-1)K$ blocks of
-$[\widehat{H}(\theta)]^{-1}=:\begin{pmatrix}\widehat{H}_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}\left(\theta\right),&\widehat{H}_{\breve{\alpha}\breve{\gamma}^{\prime}}^{\ast}(\theta)\\ \widehat{H}_{\breve{\gamma}\breve{\alpha}^{\prime}}^{\ast}\left(\theta\right),&\widehat{H}_{\breve{\gamma}\breve{\gamma}^{\prime}}^{\ast}(\theta)\end{pmatrix},$
-and $\widehat{\Sigma}_{\tau}^{\breve{\alpha}}(\theta)$ and $\widehat{\Sigma}_{\nu}^{\breve{\gamma}}(\theta)$ are the estimators of $\Sigma^{\breve{\alpha}}(\theta)$ and $\Sigma^{\breve{\gamma}}(\theta),$ defined as follows.
-For some $\tau\rightarrow\infty$ such that $\tau/T\rightarrow 0$,
-$\widehat{\Sigma}_{\tau}^{\breve{\alpha}}(\theta)$ $:=\widehat{\Sigma}_{1,\tau}^{\breve{\alpha}}(\theta)+\widehat{\Sigma}_{2,\tau}^{\breve{\alpha}}(\theta),$
-$\widehat{\Sigma}_{1,\tau}^{\breve{\alpha}}(\theta)$ $:=\frac{1}{N^{2}T^{2}}\sum_{t=1}^{T}\sum_{s=1}^{T}\mathds{1}\left\{\left|t-s\right|\leq\tau\right\}\operatorname*{diag}\big{\{}\widehat{s}_{1,t}^{\alpha}(\theta)\big{(}\widehat{s}_{1,s}^{\alpha}(\theta)\big{)}^{\prime},\ldots,\widehat{s}_{N-1,t}^{\alpha}(\theta)\big{(}\widehat{s}_{N-1,s}^{\alpha}(\theta)\big{)}^{\prime}\big{\}},$
-$\widehat{\Sigma}_{2,\tau}^{\breve{\alpha}}(\theta)$ $:=\left(\iota_{N-1}\iota_{N-1}^{\prime}\right)\otimes\frac{1}{N^{2}T^{2}}\sum_{t=1}^{T}\sum_{s=1}^{T}\mathds{1}\left\{\left|t-s\right|\leq\tau\right\}\widehat{s}_{N,t}^{\alpha}(\theta)\big{(}\widehat{s}_{N,s}^{\alpha}(\theta)\big{)}^{\prime},$
-$\widehat{s}_{i,t}^{\alpha}(\theta)$ $:=\bigg{[}\partial_{\alpha_{i}}l_{i,t}(\theta,\phi)-\frac{1}{T}\sum_{t=1}^{T}\partial_{\alpha_{i}}l_{i,t}(\theta,\phi)\bigg{]}\bigg{|}_{\phi=\widehat{\phi}(\theta)}.$
-Here, $\tau$ is a truncation parameter commonly used in the literature (see, e.g., Hahn and Kuersteiner (2011)).
-When the data exhibit serial dependence, we set $\tau=1$ or $\tau=2$ as in Fernández-Val and Weidner (2016).
-Our simulations show that the results for $\tau=1$ and $\tau=2$ are similar.
-For data without serial dependence, we set $\tau=0.$ Additionally, we define
-$\widehat{\Sigma}^{\breve{\gamma}}_{\nu}(\theta)$ $:=\frac{1}{N^{2}T^{2}}\sum_{i=1}^{N}\sum_{j=1}^{N}\mathds{1}\left\{\left|i-j\right|\leq\upsilon\right\}\widehat{S}^{\breve{\gamma}}_{i}(\theta)\big{(}\widehat{S}^{\breve{\gamma}}_{j}(\theta)\big{)}^{\prime}$
-$\widehat{S}^{\breve{\gamma}}_{i}(\theta)$ $:=\big{(}(\widehat{s}_{i,1}^{\gamma}(\theta))^{\prime},\ldots,(\widehat{s}_{i,T-1}^{\gamma}(\theta))^{\prime}\big{)}^{\prime}-\iota_{T-1}\otimes\widehat{s}_{i,T}^{\gamma}(\theta)$
-$\widehat{s}_{i,t}^{\gamma}(\theta)$ $:=\bigg{[}\partial_{\gamma_{t}}l_{i,t}\left(\theta,\phi\right)-\frac{1}{N}\sum_{i=1}^{N}\partial_{\gamma_{t}}l_{i,t}\left(\theta,\phi\right)\bigg{]}\bigg{|}_{\phi=\widehat{\phi}(\theta)},\quad i=1,\ldots,N,\ t=1,\ldots,T.$
-Specifically, when we set $\upsilon=0$ given conditional cross-sectional independence, $\widehat{\Sigma}^{\breve{\gamma}}_{\nu}(\theta)$ simplifies to
-$\widehat{\Sigma}^{\breve{\gamma}}_{0}(\theta)=\frac{1}{N^{2}T^{2}}\sum_{i=1}^{N}\widehat{S}^{\breve{\gamma}}_{i}(\theta)\big{(}\widehat{S}^{\breve{\gamma}}_{i}(\theta)\big{)}^{\prime}.$
-Thus, using the estimated bias, we obtain the corrected likelihood $L(\theta):$
-$L(\theta)=\widehat{l}(\theta)+\widehat{B}_{\breve{\alpha}}(\theta)+\widehat{B}_{\breve{\gamma}}(\theta)$
-as defined in (2.11).
-The following theorem demonstrates that the difference between the corrected likelihood $L(\theta)$ and and the infeasible likelihood $l(\theta)$ is of order $o_{\mathbb{P}}\big{(}1/\sqrt{NT}\big{)},$ implying that $L(\theta)$ no longer suffers from IPP.
+Under Assumption 1 and uniformly over $\theta\in\Theta$,
+$\mathbb{E}l(\theta)=\mathbb{E}\widehat{l}(\theta)+B_{\alpha}(\theta)+B_{\gamma}(\theta)+o(1/\sqrt{NT}),$
+as $N,T\rightarrow\infty$, where
+$B_{\alpha}(\theta):=\frac{1}{2}\operatorname*{trace}[S_{\alpha\alpha}(\theta)H_{\alpha\alpha}^{\mathcal{\ast}}(\theta)],\qquad B_{\gamma}(\theta):=\frac{1}{2}\operatorname*{trace}\mathbb{[}S_{\gamma\gamma}(\theta)H_{\gamma\gamma}^{\ast}(\theta)],$
+and $B_{\alpha}(\theta)+B_{\gamma}(\theta)=O(1/\sqrt{NT})$.
+[…]
+The bias terms $B_{\alpha}(\theta)$ and $B_{\gamma}(\theta)$ involves $\psi(\theta)$ and $\mathbb{E}$ which are infeasible to compute.
+However, $B_{\alpha}(\theta)$ and $B_{\gamma}(\theta)$ can be estimated by plug-in estimates of $H_{\alpha\alpha}^{\mathcal{\ast}}(\theta)$, $H_{\gamma\gamma}^{\ast}(\theta)$, $S_{\alpha\alpha}(\theta)$, and $S_{\gamma\gamma}(\theta)$.
+The two Hessian terms can be estimated in a straightforward way.
+Denote $\widehat{H}(\theta):=H(\theta,\widehat{\psi}(\theta))$ and partition $[\widehat{H}(\theta)]^{-1}$ into
+$[\widehat{H}(\theta)]^{-1}=:\left(\begin{array}[c]{cc}\widehat{H}_{\alpha\alpha}^{\mathcal{\ast}}(\theta)&\widehat{H}_{\alpha\gamma}^{\mathcal{\ast}}(\theta)\\ \widehat{H}_{\gamma\alpha}^{\mathcal{\ast}}(\theta)&\widehat{H}_{\gamma\gamma}^{\mathcal{\ast}}(\theta)\end{array}\right)$ (3.3)
+in the same manner as in (3.2).
+The term $H_{\alpha\alpha}^{\mathcal{\ast}}(\theta)$ can be estimated by the $(N-1)K\times(N-1)K$ block $\widehat{H}_{\alpha\alpha}^{\mathcal{\ast}}(\theta)$, and $H_{\gamma\gamma}^{\mathcal{\ast}}(\theta)$ by the $(T-1)K\times(T-1)K$ block $\widehat{H}_{\gamma\gamma}^{\mathcal{\ast}}(\theta)$.
+Estimators of $S_{\alpha\alpha}(\theta)$ and $S_{\gamma\gamma}(\theta)$ are constructed from the derivatives of $l_{i,t}(\theta,\phi)$, the original likelihood for individual $i$ at time $t$.
+This procedure needs some explanation.
+The reparameterization ensures the following identity:
+$S_{\alpha\alpha}(\theta)=D_{1}\mathcal{S}_{\alpha\alpha}(\theta)D_{1}^{\prime},$
+where $\mathcal{S}_{\alpha\alpha}(\theta)$ is the outer product of the $NK\times 1$ vector of $\partial_{\alpha}\ell(\theta,\phi)-\mathbb{E}\partial_{\alpha}\ell(\theta,\phi)$ evaluated at $\phi=\phi(\theta)=D^{\prime}\psi(\theta)$, where $\partial_{\alpha}\ell(\theta,\phi)$ is the derivative of the original likelihood $\ell(\theta,\phi)$ w.r.t.
+$\alpha:=(\alpha_{1}^{\prime},\ldots,\alpha_{N}^{\prime})^{\prime}$.
+Specifically,
+$\mathcal{S}_{\alpha\alpha}(\theta)$ $:=\frac{1}{N^{2}T^{2}}\sum_{t=1}^{T}\sum_{s=1}^{T}\mathbb{E\{}\widetilde{s}_{t}^{\alpha}(\theta)[\widetilde{s}_{s}^{\alpha}(\theta)]^{\prime}\},\qquad\widetilde{s}_{t}^{\alpha}(\theta):=\{[\widetilde{s}_{1,t}^{\alpha}(\theta)]^{\prime},\ldots,[\widetilde{s}_{N,t}^{\alpha}(\theta)]^{\prime}\}^{\prime},$
+$\widetilde{s}_{i,t}^{\alpha}(\theta)$ $:=s_{i,t}^{\alpha}(\theta,\phi(\theta))-\mathbb{E}s_{i,t}^{\alpha}(\theta,\phi(\theta)),\qquad s_{i,t}^{\alpha}(\theta,\phi):=\partial_{\alpha_{i}}l_{i,t}(\theta,\phi).$
+Here, $\widetilde{s}_{i,t}^{\alpha}(\theta)$ is the centered version of $s_{i,t}^{\alpha}(\theta,\phi(\theta))$, the derivative, w.r.t.
+$\alpha_{i}$ and evaluated at $\phi(\theta)$, of the original likelihood $l_{i,t}(\theta,\phi)$ for individual $i$ at time $t$.
+Under cross-sectional independence, $\mathcal{S}_{\alpha\alpha}(\theta)$ simplifies to
+$\mathcal{S}_{\alpha\alpha}(\theta)=\frac{1}{N^{2}T^{2}}\sum_{t=1}^{T}\sum_{s=1}^{T}\mathbb{E}\operatorname*{diag}\{\widetilde{s}_{1,t}^{\alpha}(\theta)[\widetilde{s}_{1,s}^{\alpha}(\theta)]^{\prime},\ldots,\widetilde{s}_{N,t}^{\alpha}(\theta)[\widetilde{s}_{N,s}^{\alpha}(\theta)]^{\prime}\}.$
+Naturally, the estimator of $S_{\alpha\alpha}(\theta)$ is
+$\widehat{S}_{\alpha\alpha}(\theta)$ $:=D_{1}\widehat{\mathcal{S}}_{\alpha\alpha}(\theta)D_{1}^{\prime},$ (3.4)
+$\widehat{\mathcal{S}}_{\alpha\alpha}(\theta)$ $:=\frac{1}{N^{2}T^{2}}\sum_{t=1}^{T}\sum_{s=1}^{T}\mathds{1}\{\left|t-s\right|\leq\tau\}\operatorname*{diag}\{\widehat{s}_{1,t}^{\alpha}(\theta)[\widehat{s}_{1,s}^{\alpha}(\theta)]^{\prime},\ldots,\widehat{s}_{N,t}^{\alpha}(\theta)[\widehat{s}_{N,s}^{\alpha}(\theta)]^{\prime}\},$
+$\widehat{s}_{i,t}^{\alpha}(\theta)$ $:=s_{i,t}^{\alpha}(\theta,\widehat{\phi}(\theta))-\frac{1}{T}\sum_{t=1}^{T}s_{i,t}^{\alpha}(\theta,\widehat{\phi}(\theta)).$
+The indicator $\mathds{1}\{\left|t-s\right|\leq\tau\}$ is a truncation mechanism with truncation parameter $\tau$ (see Remark 3 below).
+Similarly for $S_{\gamma\gamma}(\theta)$, the reparameterization produces the identity:
+$S_{\gamma\gamma}(\theta)=D_{2}\mathcal{S}_{\gamma\gamma}(\theta)D_{2}^{\prime},$
+where $\mathcal{S}_{\gamma\gamma}(\theta)$ is the outer product of the $TK\times 1$ vector of $\partial_{\gamma}\ell(\theta,\phi)-\mathbb{E}\partial_{\gamma}\ell(\theta,\phi)$, for $\gamma:=(\gamma_{1}^{\prime},\ldots,\gamma_{T}^{\prime})^{\prime}$, evaluated at $\phi=\phi(\theta)$:
+$\mathcal{S}_{\gamma\gamma}(\theta)$ $=\frac{1}{N^{2}T^{2}}\sum_{i=1}^{N}\sum_{j=1}^{N}\mathbb{E\{}\widetilde{s}_{i}^{\gamma}(\theta)[\widetilde{s}_{j}^{\gamma}(\theta)]^{\prime}\},\qquad\widetilde{s}_{i}^{\gamma}(\theta):=\{[\widetilde{s}_{i,1}^{\gamma}(\theta)]^{\prime},\ldots,[\widetilde{s}_{i,T}^{\gamma}(\theta)]^{\prime}\}^{\prime},$
+$\widetilde{s}_{i,t}^{\gamma}(\theta)$ $:=s_{i,t}^{\gamma}(\theta,\phi(\theta))-\mathbb{E}s_{i,t}^{\gamma}(\theta,\phi(\theta)),\qquad s_{i,t}^{\gamma}(\theta,\phi):=\partial_{\gamma_{t}}l_{i,t}(\theta,\phi),$
+where $\widetilde{s}_{i,t}^{\gamma}(\theta)$ is the centered version of $s_{i,t}^{\gamma}(\theta,\phi(\theta))$, the derivative of $l_{i,t}(\theta,\phi)$ w.r.t.
+$\gamma_{t}$ and evaluated at $\phi(\theta)$.
+By cross-sectional independence, $\mathcal{S}_{\gamma\gamma}(\theta)$ similarly reduces to
+$\mathcal{S}_{\gamma\gamma}(\theta)=\frac{1}{N^{2}T^{2}}\sum_{i=1}^{N}\mathbb{E\{}s_{i}^{\gamma}(\theta)[s_{i}^{\gamma}(\theta)]^{\prime}\}.$ (3.5)
+Intuitively, the estimator for $S_{\gamma\gamma}(\theta)$ is
+$\widehat{S}_{\gamma\gamma}(\theta)$ $=D_{2}\widehat{\mathcal{S}}_{\gamma\gamma}(\theta)D_{2}^{\prime},$ (3.6)
+$\widehat{\mathcal{S}}_{\gamma\gamma}(\theta)$ $:=\frac{1}{N^{2}T^{2}}\sum_{i=1}^{N}\mathbb{E\{}\widehat{s}_{i}^{\gamma}(\theta)[\widehat{s}_{i}^{\gamma}(\theta)]^{\prime}\},\qquad\widehat{s}_{i}^{\gamma}(\theta):=\{[\widehat{s}_{i,1}^{\gamma}(\theta)]^{\prime},\ldots,[\widehat{s}_{i,T}^{\gamma}(\theta)]^{\prime}\}^{\prime},$
+$\widehat{s}_{i,t}^{\gamma}(\theta)$ $:=s_{i,t}^{\gamma}(\theta,\widehat{\phi}(\theta))-\frac{1}{T}\sum_{t=1}^{T}s_{i,t}^{\gamma}(\theta,\widehat{\phi}(\theta)).$
+Remark 3 (truncation)
+In the estimator $\widehat{S}_{\alpha\alpha}(\theta)$, we use the truncation technique commonly found in e.g., Hahn and Kuersteiner (2011) (see also Hahn and Kuersteiner (2007)), Fernández-Val and Weidner (2016), and Galvao and Kato (2016).
+This is related to the literature on consistent estimations of autocovariances, $\mathbb{E\{}\widetilde{s}_{i,t}^{\alpha}(\theta)[\widetilde{s}_{i,s}^{\alpha}(\theta)]\}$ in our case.
+Detailed discussions in this regard can be found in Andrews (1991).
+Without truncation, we would have $\widehat{S}_{\alpha\alpha}(\theta)=\widehat{s}_{\alpha}(\theta)[\widehat{s}_{\alpha}(\theta)]^{\prime}=0$ for every $\theta$ because $\widehat{s}_{\alpha}(\theta)=0$ by definition, where $\widehat{s}_{\alpha}(\theta)$ is defined from $\widehat{s}(\theta):=s(\theta,\widehat{\psi}(\theta))$ under the same partition in (3.2).
+Here, the truncation parameter $\tau\rightarrow\infty$ but sufficiently slower relative to $T$ (see Theorem 2 below).
+Hahn and Kuersteiner (2011) suggests the default value of $\tau=1$ for small $T$.
+For intermediate $T$, we recommend following the advices from Fernández-Val and Weidner (2016) to conduct a sensitivity test for different values of $\tau$ and to avoid using very large $\tau$ values.
+In our simulation, we use $\tau=1$ and $2$ and find the difference relatively insignificant.
+On the other hand, the term $\widehat{S}_{\gamma\gamma}(\theta)$ does not require any truncation, because we have already eliminated $\mathbb{E\{}\widetilde{s}_{i}^{\gamma}(\theta)[\widetilde{s}_{j}^{\gamma}(\theta)]^{\prime}\}$ for $i\neq j$ from Equation (3.5) using cross-sectional independence.
+One may still opt to truncate $\widehat{\mathcal{S}}_{\gamma\gamma}(\theta)$, such that all elements beyond the $\tau$-th sub- and super-diagonal are symmetrically set to 00.
+We have experimented with such truncations but found little impact on $\widehat{\theta}_{L}$ and test statistics.
+The comparison is, hence, not reported in the paper.
+Using Equations (3.3), (3.4), and (3.6), we can construct the corrected likelihood $L(\theta)$, which is stated in the theorem below with its rate of convergence.
 Theorem 2
-Under Assumption 1, the corrected likelihood satisfies
-$L(\theta)=l(\theta)+o_{\mathbb{P}}\big{(}1/\sqrt{NT}\big{)},$
-as $N$ and $T\to\infty.$
-Compared to $\widehat{l}(\theta),$ the difference between the corrected likelihood and the infeasible likelihood is negligible.
-The corrected likelihood is derived from the expansion (2.7).
-Specifically, (2.7) is based on two Taylor expansions: the expansion of the score function for incidental parameters in (2.3) and the expansion of the original likelihood in (2.6).
-It is seen that the bias terms originate from the quadratic form on the right-hand side of (2.7), expressed in terms of the score and Hessian of the incidental parameters.
-Under Assumption 1, the quadratic form can be further decomposed into two additive sub-quadratic forms (Lemma 1), capturing the bias contributions from the individual and time dimensions, respectively.
-Consequently, our bias-corrected likelihood represents the original likelihood augmented by the estimation of these two additive bias terms.
-Last, we demonstrate that the MLE obtained from the corrected likelihood, $\widehat{\theta}_{L}=\arg\max_{\theta}L(\theta),$ follows an limiting normal distribution with zero mean.
-Additional regularity conditions are required.
+Under Assumption 1 and for some $\tau\rightarrow\infty$ such that $\tau/T\rightarrow 0$,
+$L(\theta)-l(\theta)=o_{\mathbb{P}}(1/\sqrt{NT}),$
+uniformly over $\theta\in\Theta$ as $N,T\rightarrow\infty$, where
+$L(\theta):=\widehat{l}(\theta)+\widehat{B}_{\alpha}(\theta)+\widehat{B}_{\gamma}(\theta),$
+as in (3.1), with
+$\widehat{B}_{\alpha}(\theta):=\frac{1}{2}\operatorname*{trace}[\widehat{S}_{\alpha\alpha}(\theta)\widehat{H}_{\alpha\alpha}^{\mathcal{\ast}}(\theta)],\qquad\widehat{B}_{\gamma}(\theta):=\frac{1}{2}\operatorname*{trace}\mathbb{[}\widehat{S}_{\gamma\gamma}(\theta)\widehat{H}_{\gamma\gamma}^{\ast}(\theta)].$
+As mentioned above, Theorem 2 states that $L(\theta)$ does not suffer from the IPP.
+Showing 2 essentially amounts to establishing $\widehat{B}_{\alpha}(\theta)+\widehat{B}_{\gamma}(\theta)=B_{\alpha}(\theta)+B_{\gamma}(\theta)+o_{\mathbb{P}}(1/\sqrt{NT})$ under $\tau/T\rightarrow 0$ and Assumption 1.
+Here, the difference between $\widehat{B}_{\alpha}(\theta)$ and $B_{\alpha}(\theta)$ is due to the truncation and the estimation error.
+For the truncation, we show that $1/\left(N^{2}T^{2}\right)\sum_{t=1}^{T}\sum_{s=1}^{T}\left(\mathds{1}\{\left|t-s\right|>\tau\}\cdot\right)$, which is the truncated (i.e., the removed) part in (3.4), is negligible relative to $1/\sqrt{NT}$ under the aforementioned conditions.
+Subsequently, the uniform law of large numbers guarantees that the estimation error is also negligible relative to $1/\sqrt{NT}$.
+Next, we turn our attention to the asymptotic properties of the $L(\theta)$-based estimators $\widehat{\theta}_{L}=\arg\max_{\theta}L(\theta)$ and of the $L(\theta)$-based LR, LM, Wald test statistics.
+We enforce the following assumption.
 Assumption 2
-  • (i) The profile likelihood function $l(\theta)$ is twice differentiable with respect to $\theta\in\Theta,$ where $\Theta$ is a compact subset of $\mathbb{R}^{K}$ with $\theta_{0}$ in its interior.
-  • (ii) For all $\theta\in\Theta$ and $\theta\neq\theta_{0},$ $\mathbb{P}\big{(}l(\theta)\neq l(\theta_{0})\big{)}>0.$
+  • (i) The likelihood $l_{i,t}(\theta,\phi)$ is four-time continuously differentiable w.r.t. to both $\theta\in\Theta$ and $\phi\in\Phi$ almost surely, such that
+    $\sup_{\theta\in\Theta}\sup_{\phi\in\Phi}\left|\partial_{\nu_{1}\cdots\nu_{S}}l_{i,t}(\theta,\phi)\right|<g(w_{it}),$
+    where $g(w_{it})$ is defined in Assumption 1, $\nu_{s}$ represents any element of $(\theta^{\prime},\phi^{\prime})^{\prime}$, and $S\in\{1,2,3,4\}$.
+  • (ii) For every $\theta\in\Theta$ and $\theta\neq\theta_{0}$, $\mathbb{P}[l(\theta)\neq l(\theta_{0})]>0$.
   • (iii) For large $N$ and $T$,
-    $\mathbb{E}\big{[}\sup_{\theta\in\Theta}|l(\theta)|\big{]}<\infty.$
-  • (iv) For large $N$ and $T,$$\{\|\triangledown_{\theta}l(\theta)\|,\theta\in\Theta\}$ is uniformly integrable, and
-    $\mathbb{E}\big{[}\sup_{\theta\in\Theta}\|\triangledown_{\theta\theta^{\prime}}l(\theta)\|\big{]}<\infty$
-    where $\triangledown_{\theta}$ and $\triangledown_{\theta\theta^{\prime}}$ denote the first and second total derivatives with respect to $\theta.$ Here, $\|\triangledown_{\theta\theta^{\prime}}l(\theta)\|$ defines the Euclidean norm of $\triangledown_{\theta\theta^{\prime}}l(\theta).$
-  • (v) The Hessian matrix $\mathbb{E}[\triangledown_{\theta\theta^{\prime}}l(\theta_{0})]$ is nonsingular for large $N$ and $T$.
-Assumption 2 requires the differentiability likelihood on a compact parameter space of $\theta.$ Assumption 2 serves as an identification condition for $\theta$ ensuring that $\mathbb{E}[l(\theta)]$ is uniquely maximized at $\theta_{0}.$ Assumption 2 shows a dominance condition implying that $1/(NT)\sum_{i=1}^{N}\sum_{t=1}^{T}l_{i,t}(\theta,\phi(\theta))$ converges uniformly in probability to $\mathbb{E}[l(\theta)]$ over $\theta.$ Consequently, Assumptions 2-2 guarantee that $\widehat{\theta}_{L}$ is an asymptotically consistent estimator of $\theta_{0}.$ Assumption 2 establishes a dominance condition for the Hessian, so that for any consistent estimator $\widetilde{\theta}$ of $\theta_{0},$
-$1/(NT)\sum_{i=1}^{N}\sum_{t=1}^{T}l_{i,t}(\widetilde{\theta},\phi(\widetilde{\theta}))\stackrel{{\scriptstyle\mathbb{P}}}{{\longrightarrow}}\mathbb{E}[\triangledown_{\theta\theta^{\prime}}l(\theta_{0})].$
-Assumption 2 ensures that the Hessian for $\theta$ is invertible.
+    $\mathbb{E}\sup_{\theta\in\Theta}\left|l(\theta)\right|<\infty.$
+  • (iv) The sequence of functions $\{\left\|\nabla_{\theta}l(\theta)\right\|:\theta\in\Theta\}$ is uniformly integrable, and
+    $\mathbb{E}\sup_{\theta\in\Theta}\left\|\nabla_{\theta\theta^{\prime}}l(\theta)\right\|<\infty$
+[…]
+  • (v) The profiled Hessian $\mathbb{E}\nabla_{\theta\theta^{\prime}}l(\theta)$ is strictly negative definite for large $N$ and $T$.
+Here Assumption 2 is similar to Assumption 1, except that we impose that $l_{i,t}(\theta,\phi)$ is four-time differentiable, and require that the derivatives of $l_{i,t}(\theta,\phi)$ w.r.t.
+to both $\theta$ and $\phi$ to be uniformly dominated by $g(w_{it})$.
+This assumption ensures that the corrected likelihood $L(\theta)$, which depends on second derivatives of $l_{i,t}(\theta,\phi)$, is twice differentiable.
+This is required by standard techniques used to show, e.g., the asymptotic normality of $\widehat{\theta}_{L}$.
+See., e.g., Hayashi (2000).
+Similar to Assumption 1, Assumption 2 is an identification condition for $\theta_{0}$.
+It guarantees that $\operatorname*{plim}_{N,T\rightarrow\infty}l(\theta)$ is uniquely maximized at $\theta_{0}$.
+Assumption 2 is a dominance condition implying that $l(\theta)-\mathbb{E}l(\theta)$ converges to 00 uniformly as $N,T\rightarrow\infty$.
+Together, Assumptions 2 to 2 are used to establish the consistency of $\widehat{\theta}_{L}$.
+In addition to these, Assumption 2 establishes a dominance condition for the profiled Hessian, so that $\triangledown_{\theta\theta^{\prime}}l(\overline{\theta})\overset{\mathbb{P}}{\longrightarrow}\mathbb{E}[\triangledown_{\theta\theta^{\prime}}l(\theta_{0})]$ for any consistent estimator $\overline{\theta}$ of $\theta_{0}$.
+Assumption 2 requires that the expected profiled Hessian is strictly negative definite for sufficiently large $N$ and $T$.
+Assumptions 2 and 2 are additionally imposed to establish the asymptotic distributions for the quantities of interests.
+The corollary below states the asymptotic properties of $\widehat{\theta}_{L}$.
 Corollary 1
-Under Assumptions 1 and 2, we have
-$\sqrt{NT}(\widehat{\theta}_{L}-\theta_{0})\stackrel{{\scriptstyle\mathbb{D}}}{{\longrightarrow}}\mathcal{N}(0,\Sigma)$
-as $N$ and $T\to\infty,$ where $\Sigma=-\lim_{N,T\to\infty}\big{(}\mathbb{E}[\triangledown_{\theta\theta^{\prime}}l(\theta_{0})]\big{)}^{-1},$ and $\mathcal{N}(0,\Sigma)$ denotes the normal distribution with mean 0 and variance $\Sigma.$
-We consider a generic nonlinear null hypothesis $H_{0}:R(\theta_{0})=0,$ where $R(\theta)$ is an $r\times 1$ vector-valued function independent of the data for some $r\leq K.$ Assume that $R(\theta)$ is continuously differentiable, with its Jacobian denoted as $J(\theta):=\triangledown_{\theta^{\prime}}R(\theta),$ and that $\operatorname*{rank}[J(\theta)]=r.$ Let $\widehat{\theta}_{R}:=\arg\max_{\theta}L(\theta)$ subject to $R(\theta)=0.$ Define the test statistics:
-$\widehat{\xi}_{LR}:=-2NT\big{\{}L(\widehat{\theta}_{R})-L(\widehat{\theta}_{L})\big{\}},$
-$\widehat{\xi}_{LM}:=-NT\big{\{}\triangledown_{\theta^{\prime}}L(\widehat{\theta}_{R})[\triangledown_{\theta\theta^{\prime}}L(\widehat{\theta}_{R})]^{-1}\triangledown_{\theta}L(\widehat{\theta}_{R})\big{\}},$
-$\widehat{\xi}_{Wald}:=-NT\big{\{}R(\widehat{\theta}_{L})^{\prime}\big{(}J(\widehat{\theta}_{L})[\triangledown_{\theta\theta^{\prime}}L(\widehat{\theta}_{L})]^{-1}J(\widehat{\theta}_{L})^{\prime}\big{)}^{-1}R(\widehat{\theta}_{L})\big{\}}.$
-The following corollary states that the hypothesis testing based on the corrected likelihood aligns with standard results in the context of MLE.
+Under Assumptions 1 and 2, $\widehat{\theta}_{L}\overset{\mathbb{P}}{\rightarrow}\theta_{0}$ and
+$\sqrt{NT}[-\mathbb{E}\triangledown_{\theta\theta^{\prime}}l(\theta_{0})]^{1/2}(\widehat{\theta}_{L}-\theta_{0})\overset{\mathbb{D}}{\longrightarrow}\mathcal{N}(0,\mathbb{I}_{K})$
+as $N,T\rightarrow\infty$.
+Here, $[-\mathbb{E}\triangledown_{\theta\theta^{\prime}}l(\theta_{0})]^{1/2}$ is the square root of $-\mathbb{E}\triangledown_{\theta\theta^{\prime}}l(\theta_{0})$ which can be obtained by square-rooting the eigenvalues in the eigendecomposition of $-\mathbb{E}\triangledown_{\theta\theta^{\prime}}l(\theta_{0})$.
+By Assumption 2, $\mathbb{E}\triangledown_{\theta\theta^{\prime}}l(\theta_{0})$ is strictly negative definite so that $-\mathbb{E}\triangledown_{\theta\theta^{\prime}}l(\theta_{0})$ only has strictly positive eigenvalues.
+In addition, since we focus on maximum likelihood models, we use the information matrix equality to simplify the presentation of Corollary 1.
+In case the information matrix equality does not hold [^6] The information matrix equality generally does not hold for, e.g., non-likelihood M-estimation models, misspecified models, and quasi-likelihood models., $-\mathbb{E}\triangledown_{\theta\theta^{\prime}}l(\theta_{0})$ shall be replaced by its sandwich form and $[-\mathbb{E}\triangledown_{\theta\theta^{\prime}}l(\theta_{0})]^{1/2}$ shall be constructed thereof.
+See, e.g., White (1982) for detailed discussions.
+For inference procedures, we consider a generic null hypothesis $H_{0}:R(\theta_{0})=0$, where $R(\theta)$ is a known $r\times 1$ ($r\leq K$) vector-valued non-random function with Jacobian $J(\theta):=\partial_{\theta^{\prime}}R(\theta)$ satisfying $\operatorname*{rank}[J(\theta)]=r$.
+The corollary below gives the asymptotic distribution, under $H_{0}$, of $\widehat{\xi}_{LR}$ (the LR statistic), $\widehat{\xi}_{LM}$ (the LM statistic), and $\widehat{\xi}_{Wald}$ (the Wald statistic), defined respectively as
+$\widehat{\xi}_{LR}$ $:=-2NT\{L(\widehat{\theta}_{R})-L(\widehat{\theta}_{L})\},$
+$\widehat{\xi}_{LM}$ $:=-NT\{\triangledown_{\theta^{\prime}}L(\widehat{\theta}_{R})[\triangledown_{\theta\theta^{\prime}}L(\widehat{\theta}_{R})]^{-1}\triangledown_{\theta}L(\widehat{\theta}_{R})\},$
+$\widehat{\xi}_{Wald}$ $:=-NT\{R^{\prime}(\widehat{\theta}_{L})[J(\widehat{\theta}_{L})[\triangledown_{\theta\theta^{\prime}}L(\widehat{\theta}_{L})]^{-1}J^{\prime}(\widehat{\theta}_{L})]^{-1}R(\widehat{\theta}_{L})\},$
+where $\widehat{\theta}_{R}:=\arg\max_{\theta}L(\theta)$ subject to $R(\theta)=0$.
 Corollary 2
-Suppose Assumptions 1 and 2.
-Under the null hypothesis $H_{0}:R(\theta_{0})=0,$ the statistics $\widehat{\xi}_{LR},$ $\widehat{\xi}_{LM},$ and $\widehat{\xi}_{Wald}$ are asymptotically equivalent, following a common limiting $\chi^{2}(r)$ distribution with degrees of freedom $r$.
-4 Simulations
-In the simulation study, we consider data generating process (DGP) from a dynamic logit model as follows:
+Under Assumptions 1 to 2 and $H_{0}$,
+$\widehat{\xi}_{LR},\widehat{\xi}_{LM},\widehat{\xi}_{Wald}\overset{\mathbb{D}}{\longrightarrow}\chi^{2}(r)$
+as $N,T\rightarrow\infty$, where $\chi^{2}(r)$ is the $\chi^{2}$-distribution with degrees of freedom $r$.
+Similarly, the test statistics $\widehat{\xi}_{LM}$ and $\widehat{\xi}_{Wald}$ are constructed under the information matrix equality.
+Adjustments are generally required if it does not hold.
+For instance, White (1982) shows that $-[\triangledown_{\theta\theta^{\prime}}L(\widehat{\theta}_{L})]^{-1}$ involved in $\widehat{\xi}_{Wald}$ shall be replaced by its sandwich form for quasi-likelihood models.
+Given Theorem 2, Corollaries 1 and 2 are established using standard techniques.
+Particularly, Theorem 2 guarantees $L(\theta)\rightarrow_{p}\mathbb{E}l(\theta)$, so that the consistency of $\widehat{\theta}_{L}$ follows from Newey and McFadden (1994).
+Theorem 2 also implies that $\nabla_{\theta}L(\theta)-\nabla_{\theta}l(\theta)$ is $o_{\mathbb{P}}(1/\sqrt{NT})$, from which the asymptotic normality is established using techniques in, e.g., Hayashi (2000).
+The proof of Corollary 2 also closely resembles the standard techniques in, e.g., Hayashi (2000), where it is shown that all $\widehat{\xi}_{LR}$, $\widehat{\xi}_{LM}$, and $\widehat{\xi}_{Wald}$ are equivalent, converging in probability to some $\widehat{\xi}$ which is asymptotically $\chi^{2}(r)$.
+4 Simulation
+We conduct the simulation study in this section.
+We first illustrate that our procedure is capable of reducing the bias of the MLE as well as delivering likelihood-based tests with an empirical size very close to the nominal level.
+Followed by this, we then consider a different design in which the MLEs and the tests suffer from severe biases and size-distortions.
+We demonstrate that our approach leads to dramatic improvements on the estimators and substantial reductions of the Type-I error risks of the tests, even in a severely biased setting.
+Finally in this section, we compare our method with the Jackknife [^7] Since there has currently been no Jackknife procedures for DN-HP models.
+We compared our method to the Jackknife of Fernández-Val and Weidner (2016), who consider models with fixed effects only in the intercept.
+and show that our method out-performs.
+In Design 1, we consider a dynamic binary response model as:
 $Y_{it}=\mathds{1}\{\rho Y_{it-1}+(\beta+\alpha_{1,i}+\gamma_{1,t})X_{it}+\alpha_{2,i}+\gamma_{2,t}+\varepsilon_{it}>0\}$ (4.1)
-where the disturbance term $\varepsilon_{it}$ follows a standard logistic distribution, independent across both $i$ and $t,$ and also independent of the regressor $X_{it}.$ According to the identification restriction, we generate true values for the incidental parameters $\{\alpha_{1,i}^{0},\gamma_{1,t}^{0}\}$ from $\mathcal{N}(0,0.01)$, after demeaning.
-$\{\alpha_{2,i}^{0},\gamma_{2,t}^{0}\}$ are also generated from $\mathcal{N}(0,0.01)$ where we only demean $\gamma_{2,t}^{0}$.
-Given these parameters, $X_{it}$ is independently generated from $\mathcal{N}\big{(}\sum_{k=1}^{2}(\alpha_{k,i}^{0}+\gamma_{k,t}^{0}),0.25\big{)}$ across both $i$ and $t.$ The true structural parameter is set as $\theta_{0}=(\rho_{0},\beta_{0})^{\prime}=(0.5,0.5)^{\prime}.$ The initial $Y_{i0}=\mathds{1}\{(\beta_{0}+\alpha_{1,i}^{0})X_{i0}+\alpha_{2,i}^{0}+\varepsilon_{i0}>0\}$ with $X_{i0}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0},0.25)$ and $\varepsilon_{i0}$ following a standard logistic distribution.
-Model (4.1) also encompasses the static logit model as a specific case where $\rho\equiv 0.$ We simulate this static case with the same DGP as the dynamic model, except for $\rho_{0}=0$ ($\theta_{0}=\beta_{0}$).
-We demonstrate the finite sample performance of the MLE of $\theta_{0}$ by maximizing our bias-corrected likelihood $L(\theta),$ and compare it with the uncorrected (original) MLE obtained from $\widehat{l}(\theta)$ without bias correction.
-In addition, we compare hypothesis testing in terms of size and power based on both the corrected and uncorrected likelihoods.
-We consider the sample size $(N,T)\in\{(30,30),(60,60),(90,90)\},$ with 1000 replications for each DGP.
-Table 1 reports the estimation bias and RMSE of our bias-corrected and the uncorrected MLEs for both the static ($\theta_{0}=\beta_{0}$) and dynamic ($\theta_{0}=(\rho_{0},\beta_{0})^{\prime}$) logit models.
-In the static model, $\widehat{\beta}$ represents the uncorrected MLE from $\widehat{l}(\theta)$, and $\widehat{\beta}_{L}$ is our bias-corrected MLE.
-For the dynamic model, $(\widehat{\beta},\widehat{\rho})$ denotes the uncorrected MLEs, while $(\widehat{\beta}_{L}^{(\tau)},\widehat{\rho}_{L}^{(\tau)})^{\prime}$ represents our bias-corrected estimates, with the truncation parameter $\tau$ set to $\tau=1$ or $2$.
-In the static model, ${Y_{it}}$ are generated independently across time $t$ where we set $\tau=0$ for the estimation, while the data in the dynamic model exhibit serial dependence.
-Table 1: Comparison between Bias-corrected and Uncorrected MLEs
+where the true value is $\theta_{0}=(\rho_{0},\beta_{0})^{\prime}=(0.5,0.5)^{\prime}$ and the disturbance $\varepsilon_{it}$ is standard logistic, independent and identically distributed (i.i.d.) over $(i,t)$ and independent from $X_{it}$.
+The true incidental parameter $\{\alpha_{1,i}^{0},\gamma_{1,t}^{0}\}$ are generated from $\mathcal{N}(0,0.01)$, and demeaned after they are generated.
+$\{\alpha_{2,i}^{0},\gamma_{2,t}^{0}\}$ are also generated from $\mathcal{N}(0,0.01)$ but only $\gamma_{2,t}^{0}$ is demeaned.
+Conditional on the generated $\alpha_{k,i}^{0}$ and $\gamma_{k,t}^{0}$ ($k=1,2$), the regressor $X_{it}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0}+\gamma_{1,t}^{0}+\gamma_{2,t}^{0},0.25)$ i.i.d.
+over $(i,t)$ and the initial $Y_{i0}=\mathds{1}\{(\beta_{0}+\alpha_{1,i}^{0})X_{i0}+\alpha_{2,i}^{0}+\varepsilon_{i0}>0\}$ with $X_{i0}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0},0.25)$ i.i.d.
+over $(i,t)$.
+We also simulate a static model, where we set $\rho\equiv 0$ in (4.1), using data generated in the same way but setting $\rho_{0}=0$ (hence $\theta_{0}=\beta_{0}$).
+Table 1 reports the estimation bias and RMSE, under Design 1, of $\widehat{\theta}_{L}$ and $\widehat{\theta}$ for both the static ($\theta_{0}=\beta_{0}$) and dynamic ($\theta_{0}=(\rho_{0},\beta_{0})^{\prime}$) logit models.
+For the static model, $\widehat{\beta}$ denotes the uncorrected MLE obtained by maximing $\widehat{l}(\theta)$, and $\widehat{\beta}_{L}$ is the corrected estimator, obtained by maximing $L(\theta)$.
+For the dynamic model, $(\widehat{\beta},\widehat{\rho})^{\prime}$ is the uncorrected estimator for $(\rho_{0},\beta_{0})^{\prime}$, while $(\widehat{\beta}_{L}^{(\tau)},\widehat{\rho}_{L}^{(\tau)})^{\prime}$ represents the bias-corrected estimator, with the truncation parameter $\tau\in(1,2)$.
+For the static model (with $\rho_{0}\equiv 0$), we set $\tau=0$ since the data are independent over $i$, conditional on $\{\alpha_{k,i}^{0},\gamma_{k,t}^{0}\}$.
+The results reveal that the uncorrected estimators are severely biased in both models, while our bias-corrected estimators substantially reduce the bias.
+For instance, in the static model with $(N,T)=(30,30)$, the bias for $\widehat{\beta}_{L}$ is only $0.05$, which is less than $1/3$ of the significant bias of $0.158$ for $\widehat{\beta}$.
+As the sample size increases to $(N,T)=(90,90)$, $\widehat{\beta}_{L}$ is nearly unbiased, while the bias of $\widehat{\beta}$ remaining five times larger than that of $\widehat{\beta}_{L}$.
+Furthermore, the RMSEs from our corrected estimators are consistently lower than those of the uncorrected estimators, especially with smaller sample sizes.
+This indicates that our procedure is capable of reducing the IPP bias without increasing the RMSE.
+For large $(N,T)$, the choice of $\tau$ has negligible impact on the performance of our bias-corrected estimators.
+Table 1: Comparisons of Uncorrected and Corrected Logit Estimates, Design 1
 $(N,T)$ $(30,30)$ $(60,60)$ $(90,90)$
@@
 […]
-    Notes: The true structure parameters are $\theta_{0}=(\beta_{0},\rho_{0})^{\prime}=(0.5,0.5)^{\prime}$ for the dynamic model, and $\theta_{0}=\beta_{0}=0.5$ for the static model.
-In the static model, $\widehat{\beta}$ and $\widehat{\beta}_{L}$ represent the uncorrected MLE and our bias-corrected MLE, respectively.
-In the dynamic model, $(\widehat{\beta},\widehat{\rho})$ is the uncorrected MLE and $(\widehat{\beta}_{L}^{(\tau)},\widehat{\rho}_{L}^{(\tau)})^{\prime}$ represents our bias-corrected MLE with the truncation parameter $\tau$ set to $\tau=1$ or $2$.
+Notes: Logit model with $\varepsilon_{it}$ standard logistic.
+Design 1: $\{\alpha_{k,i}^{0},\gamma_{k,t}^{0}\}\sim\mathcal{N}(0,0.01)$, $X_{it}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0}+\gamma_{1,t}^{0}+\gamma_{2,t}^{0},0.25)$, and $X_{i0}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0},0.25)$.
+For the static model, $\widehat{\beta}$ and $\widehat{\beta}_{L}$ are, respectively, the uncorrected and corrected estimates of $\beta_{0}=0.5$.
+For the dynamic model $(\widehat{\beta},\widehat{\rho})^{\prime}$ is the uncorrected estimate of $(\rho_{0},\beta_{0})^{\prime}=(0.5,0.5)^{\prime}$ from and $(\widehat{\beta}_{L}^{(\tau)},\widehat{\rho}_{L}^{(\tau)})^{\prime}$ is the corrected estimate with truncation parameter $\tau\in(1,2)$.
 The number of replications is $1000$.
-The results in Table 1 reveal that the uncorrected MLEs are severely biased in both models, while our bias-corrected estimators substantially reduce the bias.
-For instance, in the static model with $(N,T)=(30,30)$, the bias for $\widehat{\beta}_{L}$ is only $0.05$, which is less than $1/3$ of the significant bias of $0.158$ for $\widehat{\beta}$.
-As the sample size increases to $(N,T)=(90,90)$, $\widehat{\beta}_{L}$ is nearly unbiased, while the bias for $\widehat{\beta}$ remaining five times larger than that for $\widehat{\beta}_{L}$.
-Furthermore, the RMSEs form our corrected MLEs are consistently lower than those of the uncorrected MLEs, especially with smaller sample sizes, underscoring the advantage of our bias correction method.
-Additionally, in the dynamic model, the choice of $\tau$ has little impact on the performance of our bias-corrected estimators.
-We also conduct LR tests using both bias-corrected and uncorrected likelihoods.
-Consider the null hypothesis $H_{0}:\theta_{0}=\theta_{0}+\delta$ with $\delta\in\{0,\pm 0.2,\pm 0.15,\pm 0.1\}.$ We report the empirical size ($\delta=0$) and power ($\delta\neq 0$) at a $5\%$ significance level.
-Table 2 presents the empirical size based on corrected and uncorrected likelihoods, where the corrected likelihoods use truncation $\tau=0$ for the static model and $\tau=1$ for the dynamic model.
-The results show that the LR tests using uncorrected likelihoods suffer from severe size distortion, while the empirical size from our bias-corrected likelihoods maintains the significance level as the sample size increases.
-Additionally, our bias-corrected tests significantly reduce Type-I error even for small sample size.
-Table 2: Empirical Size of LR Tests Based on Bias-corrected and Uncorrected Likelihoods
+Next, under Design 1, we conduct LR tests using both the corrected and the uncorrected likelihoods for the null hypothesis $H_{0}:\theta_{0}=(0.5,0.5)^{\prime}+\delta$ with $\delta\in\{0,\pm 0.2,\pm 0.15,\pm 0.1\}$.
+We report the empirical size ($\delta=0$) and the empirical power ($\delta\neq 0$) at the $5\%$ significance level.
+For the corrected likelihood, we set the truncation parameter $\tau=0$ for the static model and $\tau=1$ for the dynamic model.
+Table 2 presents the empirical size.
+The results show that the LR test using the uncorrected likelihood suffers from a severe size distortion.
+On the opposite, the empirical size from our bias-corrected likelihood maintains the significance level as the sample size increases.
+Table 2: Empirical Size of LR Tests Based on Corrected and Uncorrected Likelihoods for Logit Model, Design 1
 Static Model Dynamic Model
@@
 […]
-    Notes: The null hypothesis $H_{0}:(\beta_{0},\rho_{0})^{\prime}=(0.5,0.5)^{\prime}$ for the dynamic model, and $H_{0}:\beta_{0}=0.5$ for the static model, with a significance level $5\%$.
+Notes: Logit model with $\varepsilon_{it}$ standard logistic.
+Design 1: $\{\alpha_{k,i}^{0},\gamma_{k,t}^{0}\}\sim\mathcal{N}(0,0.01)$, $X_{it}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0}+\gamma_{1,t}^{0}+\gamma_{2,t}^{0},0.25)$, and $X_{i0}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0},0.25)$.
+The null hypothesis $H_{0}:(\beta_{0},\rho_{0})^{\prime}=(0.5,0.5)^{\prime}$ for the dynamic model, and $H_{0}:\beta_{0}=0.5$ for the static model, with a significance level of $5\%$.
 The number of replications is $1000$.
-Table 3 presents the empirical power of LR tests based on our bias-corrected likelihoods.
+Table 3 presents the empirical power of LR tests based on our corrected likelihood under Design 1.
 The empirical power increases as both the value of $\delta$ and the sample size grow.
-For $(N,T)=(90,90)$, when $\delta=\pm 0.2$, the rejection rates reach nearly 100% for both the static and dynamic models.
-We also perform LM and Wald tests, reporting empirical size and power under the same DGPs in the appendix.
-The results align with those of the LR tests.
-Additionally, we perform simulations using probit models with results provided in the appendix.
-Table 3: Empirical Power of LR Tests Based on Bias-corrected Likelihoods
+For $(N,T)=(90,90)$, when $\delta=\pm 0.2$, the rejection rates reach nearly $100\%$ for both the static and dynamic models.
+We also perform LM and Wald test, reporting the empirical size and power under the same setting in Appendix B.1.
+The results align with those of the LR test.
+We additionally perform the simulations for a probit model with $\varepsilon_{it}\sim\mathcal{N}(0,1)$ i.i.d.
+over $(i,t)$ and independent from $X_{it}$.
+The results are provided in Appendix B.2.
+Table 3: Empirical Power of LR Test for Logit Model using $L(\theta)$, Design 1
 $\delta$
@@
 […]
-    Notes: The null hypothesis $H_{0}:(\beta_{0},\rho_{0})^{\prime}=(0.5+\delta,0.5+\delta)^{\prime}$ for the dynamic model, and $H_{0}:\beta_{0}=0.5+\delta$ for the static model, with a significance level $5\%,$ where $\delta\in\{\pm 0.2,\pm 0.15,\pm 0.1\}$.
+Notes: Logit model with $\varepsilon_{it}$ standard logistic.
+Design 1: $\{\alpha_{k,i}^{0},\gamma_{k,t}^{0}\}\sim\mathcal{N}(0,0.01)$, $X_{it}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0}+\gamma_{1,t}^{0}+\gamma_{2,t}^{0},0.25)$, and $X_{i0}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0},0.25)$.
+The null hypothesis $H_{0}:(\beta_{0},\rho_{0})^{\prime}=(0.5+\delta,0.5+\delta)^{\prime}$ for the dynamic model, and $H_{0}:\beta_{0}=0.5+\delta$ for the static model, with a significance level $5\%,$ where $\delta\in\{\pm 0.2,\pm 0.15,\pm 0.1\}$.
 The number of replications is $1000$.
-Finally, we compare our bias correction method to the Jackknife approach by Fernández-Val and Weidner (2016), who consider a probit model in their simulation.
-We consider a similar setup, using the same settings as in (4.1), except that the disturbance term $\varepsilon_{it}$ follows a standard normal distribution.
-Table 4 displays the estimation bias and RMSE of the uncorrected MLEs, our bias-corrected MLEs, and the Jackknife bias-corrected estimators, both for the static ($\theta_{0}=\beta_{0}$) and dynamic ($\theta_{0}=(\rho_{0},\beta_{0})^{\prime}$) probit models.
-In the static model, $\widehat{\beta}_{L}$ is our bias-corrected MLE with $\tau=0$, and $\widehat{\beta}_{J}$ represents the Jackknife-corrected MLE.
-In the dynamic model, $(\widehat{\beta}_{L},\widehat{\rho}_{L})^{\prime}$ is our bias-corrected MLE with $\tau=1$, and $(\widehat{\beta}_{J},\widehat{\rho}_{J})^{\prime}$ is the Jackknife-corrected MLE.
-Table 4 shows that our bias-corrected MLEs outperform both the uncorrected and Jackknife bias-corrected estimators, with consistently smaller RMSEs across all DGPs.
-The bias correction by our method is significantly greater than that of the Jackknife approach, especially when the sample size is small.
-For example, in the static model with $(N,T)=(30,30)$, the bias of our corrected MLE is only $0.048$, while the bias of the Jackknife estimator is $-1.228,$ over twenty-five times larger.
+Next, we conduct the same experiment under Design 2, where we generate $\{\alpha_{k,i}^{0},\gamma_{k,t}^{0}\}\sim\mathcal{N}(0,1/4)$, $X_{it}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0}+\gamma_{1,t}^{0}+\gamma_{2,t}^{0},1)$, and $X_{i0}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0},1)$, keeping everything else intact as in (4.1).
+We increase the volatility of $\alpha_{k,i}^{0}$, $\gamma_{k,t}^{0}$, $X_{it}$, and $X_{i0}$ to induce larger biases and size distortions.
+This time, we present the results for the probit model and leave that for the logit model in Appendix B.1.
+Table 4 presents the comaprison of the bias and RMSE of the corrected and the uncorrected estimators for the probit model, under Design 2.
+First of all, the probit estimator here is much more biased than the previous logit counterparts.
+For the static model with $(N,T)=(30,30)$, $\widehat{\beta}$ contains a bias of $0.943$.
+Recall that the true value is only $\beta_{0}=0.5$, implying that the bias is almost $200\%$.
+Our corrected estimator $\widehat{\beta}_{L}$ successfully reduces this bias to $0.328$, which is one third of the uncorrected and roughly $60\%$ biased.
+As the sample sizes increase, the bias systematically becomes smaller both for the uncorrected and corrected estimators, with dramatic improvements observed from the corrected ones both for the stiatic and dynamic models.
+We indeed observe that the RMSE of $\widehat{\beta}_{L}$ is twice as large as that of $\widehat{\beta}$ for the static model under $N,T=(30,30)$.
+However, it is the only incidence here and we attribute this to some sort of numerical coincidence [^8] Below, when we compare our method with the Jackknife under the same setting, it will be seen that the RMSE of our corrected estimator is indeed smaller than the uncorrected.
+See also Footnote 9 below..
+It can be checked that the RMSE of our corrected estimator is no worse the uncorrected one in every other case.
+Table 4: Comparisons of Uncorrected and Corrected Probit Estimates, Design 2
+$(N,T)$ $(30,30)$ $(60,60)$ $(90,90)$
+Bias RMSE Bias RMSE Bias RMSE
+Static Model
+$\widehat{\beta}$ $0.943$ $2.272$ $0.151$ $0.252$ $0.073$ $0.081$
+$\widehat{\beta}_{L}$ $0.328$ $5.170$ $0.041$ $0.062$ $0.017$ $0.040$
+Dynamic Model
+$\widehat{\beta}$ $0.909$ $1.842$ $0.166$ $0.248$ $0.084$ $0.100$
+$\widehat{\beta}_{L}^{(1)}$ $0.207$ $0.396$ $0.048$ $0.078$ $0.022$ $0.042$
+$\widehat{\beta}_{L}^{(2)}$ $0.202$ $0.282$ $0.049$ $0.080$ $0.022$ $0.042$
+$\widehat{\rho}$ $-0.030$ $0.162$ $-0.016$ $0.073$ $-0.011$ $0.045$
+$\widehat{\rho}_{L}^{(1)}$ $-0.015$ $0.140$ $-0.005$ $0.066$ $-0.004$ $0.044$
+$\widehat{\rho}_{L}^{(2)}$ $-0.009$ $0.141$ $-0.001$ $0.067$ $-0.001$ $0.045$
+[…]
+Notes: Probit model with $\varepsilon_{it}\sim\mathcal{N}(0,1)$.
+Design 2: $\{\alpha_{k,i}^{0},\gamma_{k,t}^{0}\}\sim\mathcal{N}(0,1/4)$, $X_{it}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0}+\gamma_{1,t}^{0}+\gamma_{2,t}^{0},1)$, and $X_{i0}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0},1)$.
+For the static model, $\widehat{\beta}$ and $\widehat{\beta}_{L}$ are, respectively, the uncorrected and corrected estimates of $\beta_{0}=0.5$.
+For the dynamic model $(\widehat{\beta},\widehat{\rho})^{\prime}$ is the uncorrected estimate of $(\rho_{0},\beta_{0})^{\prime}=(0.5,0.5)^{\prime}$ from and $(\widehat{\beta}_{L}^{(\tau)},\widehat{\rho}_{L}^{(\tau)})^{\prime}$ is the corrected estimate with truncation parameter $\tau\in(1,2)$.
+The number of replications is $1000$.
+Table 5 below reports the empirical size of the LR test for the probit model under Design 2.
+As before, the null hypothesis is $H_{0}:(\beta_{0},\rho_{0})^{\prime}=(0.5,0.5)^{\prime}$ for the dynamic model and $H_{0}:\beta_{0}=0.5$ for the static.
+We set the truncation parameter $\tau=0$ for the static model and $\tau=1$ for the dynamic model.
+Here the size distortion is much more fatal than in the logit case above.
+For the static model with $(N,T)=(30,30)$, the uncorrected LR test produces an empirical size of $77\%$.
+Even under $(N,T)=(90,90)$, the empirical size is still $72\%$, indicating a prohibitively high risk of committing a Type-I error.
+On the other hand, the $L(\theta)$-based LR test delivers a $14\%$ empirical size only, which is almost a $60\%$ decrease in magnitude.
+Similar performance is observed in other cases.
+The results here demonstrate that our method can reduce the type-I error risks by a major margin.
+Indeed, our test still over-rejects under $(N,T)=(90,90)$.
+This is anticipated, because the IPP under Design 2 is much more drastic and $(N,T)=(90,90)$ may not be sufficient for the test statistic to be well-approximated by $\chi^{2}$.
+However, it is clear that, as $(N,T)$ increase, the empirical size does become smaller.
+We have tried additional small-scale simulations (not reported though) with even larger $(N,T)$ and found that the empirical size still decreases with $(N,T)$.
+We similarly report the results from the LM and the Wald test in Appendix B.2.
+We report the empirical power under Design 2 in the Appendix B, since they do not add too much value in the discussion here.
+A quick summary for those results is that our procedure produces tests with stronger power in Design 2.
+Table 5: Empirical Size of LR Tests Based on Corrected and Uncorrected Likelihoods for Logit Model, Design 2
+Static Model Dynamic Model
+$(N,T)$ Uncorrected Corrected Uncorrected Corrected
+$(30,30)$ $77\%$ $39\%$ $71\%$ $37\%$
+$(60,60)$ $76\%$ $20\%$ $66\%$ $19\%$
+$(90,90)$ $72\%$ $14\%$ $63\%$ $14\%$
+[…]
+Notes: Probit model with $\varepsilon_{it}\sim\mathcal{N}(0,1)$.
+Design 2: $\{\alpha_{k,i}^{0},\gamma_{k,t}^{0}\}\sim\mathcal{N}(0,1/4)$, $X_{it}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0}+\gamma_{1,t}^{0}+\gamma_{2,t}^{0},1)$, and $X_{i0}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0},1)$.
+The null hypothesis $H_{0}:(\beta_{0},\rho_{0})^{\prime}=(0.5,0.5)^{\prime}$ for the dynamic model, and $H_{0}:\beta_{0}=0.5$ for the static model, with a significance level of $5\%$.
+The number of replications is $1000$.
+Finally, we compare our method with the Jackknife of Fernández-Val and Weidner (2016).
+We consider the same probit model and continue the same setting and design as above.
+Table 6 displays the estimation bias and the RMSE of the uncorrected estimators $\{\widehat{\rho},\widehat{\beta}\}$, our corrected estimator $\{\widehat{\rho}_{L},\widehat{\beta}_{L}\}$, and the Jackknife-corrected estimator $\{\widehat{\rho}_{J},\widehat{\beta}_{J}\}$ for the dynamic model.
+For the static model, only $\widehat{\beta}$, $\widehat{\beta}_{L}$, and $\widehat{\beta}_{J}$ apply and are reported.
+We set the truncation parameter $\tau=0$ for static model and $\tau=1$ for dynamic model.
+Across all pairs of $(N,T)$, the Jackknife-corrected $\widehat{\beta}_{J}$ consistently exhibits larger bias and RMSE than our $\widehat{\beta}_{L}$.
+Even worse, $\widehat{\beta}_{J}$ shows a larger bias and RMSE than the uncorrected $\widehat{\beta}$ across all pairs of $(N,T)$.
+For instance, with $(N,T)=(30,30)$ and the static model, $\widehat{\beta}_{J}$ shows a bias of $-4.929$, which is roughly $900\%$ biased (in magnitude) relative to the true value $\beta_{0}=0.5$.
+Observe that the uncorrected estimator $\widehat{\beta}$ only shows a roughly $170\%$ bias relative to the true value.
+Our corrected $\widehat{\beta}_{L}$ performs the best among the three, showing a $36\%$ bias [^9] Also notice that the RMSE of $\widehat{\beta}_{L}$ in Table 6 is much smaller than that of $\widehat{\beta}_{L}$ in Table 4, where both tables are essentially under the same setting.
+This motivates our attribution of the large RMSE in Table 4 to numerical coincidence..
+To summarize, we find that the Jackknife is not a suitable candidate for DN-HP models.
+Although it does reduce the bias for the homogeneous coefficient for large $N,T$, the bias the Jackknife causes to the heterogeneous parameter is dangerously large, especially in small sample.
 Table 6: Comparison with the Jackknife Bias Correction Method
@@
 Static Model
-$\widehat{\beta}$ $0.144$ $0.197$ $0.061$ $0.081$ $0.041$ $0.052$
-$\widehat{\beta}_{L}$ $0.048$ $0.123$ $0.011$ $0.051$ $0.007$ $0.031$
-$\widehat{\beta}_{J}$ $-1.228$ $5.930$ $-0.022$ $0.058$ $-0.007$ $0.032$
+$\widehat{\beta}$ $0.846$ $2.168$ $0.151$ $0.261$ $0.074$ $0.082$
+$\widehat{\beta}_{L}$ $0.182$ $0.339$ $0.041$ $0.068$ $0.019$ $0.033$
+$\widehat{\beta}_{J}$ $-4.929$ $14.383$ $-0.504$ $0.981$ $-0.105$ $0.214$
 Dynamic Model
-$\widehat{\beta}$ $0.172$ $0.262$ $0.069$ $0.088$ $0.042$ $0.054$
-$\widehat{\beta}_{L}$ $0.072$ $0.207$ $0.017$ $0.053$ $0.007$ $0.032$
-$\widehat{\beta}_{J}$ $-1.339$ $3.873$ $-0.029$ $0.101$ $-0.009$ $0.034$
-$\widehat{\rho}$ $-0.050$ $0.122$ $-0.028$ $0.056$ $-0.017$ $0.037$
-$\widehat{\rho}_{L}$ $-0.025$ $0.104$ $-0.014$ $0.049$ $-0.007$ $0.032$
-$\widehat{\rho}_{J}$ $0.003$ $0.115$ $-0.001$ $0.050$ $0.000$ $0.033$
+$\widehat{\beta}$ $1.040$ $2.343$ $0.184$ $0.287$ $0.080$ $0.090$
+$\widehat{\beta}_{L}$ $0.209$ $0.332$ $0.052$ $0.082$ $0.020$ $0.049$
+$\widehat{\beta}_{J}$ $-3.352$ $7.708$ $-0.558$ $1.125$ $-0.118$ $0.217$
+$\widehat{\rho}$ $-0.046$ $0.169$ $-0.022$ $0.072$ $-0.014$ $0.046$
+$\widehat{\rho}_{L}$ $-0.028$ $0.145$ $-0.011$ $0.067$ $-0.009$ $0.044$
+$\widehat{\rho}_{J}$ $-0.029$ $0.169$ $-0.006$ $0.068$ $-0.002$ $0.044$
 […]
-    Notes: $(\beta_{0},\rho_{0})^{\prime}=(0.5,0.5)^{\prime}$ for the dynamic model, and $\beta_{0}=0$ for the static model.
-Accordingly, $(\widehat{\beta},\widehat{\rho})^{\prime}$ (or $\widehat{\beta}$) denotes the uncorrected MLE; $(\widehat{\beta}_{L},\widehat{\rho}_{L})^{\prime}$ (or $\widehat{\beta}_{L}$) is our bias-corrected MLE with truncation parameter $\tau=1$ for the dynamic model or $\tau=0$ for the static model; and $(\widehat{\beta}_{J},\widehat{\rho}_{J})^{\prime}$ (or $\widehat{\beta}_{J}$) represents the Jackknife bias-corrected estimators.
+Notes: Probit model with $\varepsilon_{it}\sim\mathcal{N}(0,1)$.
+Design 2: $\{\alpha_{k,i}^{0},\gamma_{k,t}^{0}\}\sim\mathcal{N}(0,1/4)$, $X_{it}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0}+\gamma_{1,t}^{0}+\gamma_{2,t}^{0},1)$, and $X_{i0}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0},1)$.
+The true value is $(\beta_{0},\rho_{0})^{\prime}=(0.5,0.5)^{\prime}$ for the dynamic model, and $\beta_{0}=0$ for the static model.
+Accordingly, $(\widehat{\beta},\widehat{\rho})^{\prime}$ (or $\widehat{\beta}$) denotes the uncorrected estimator; $(\widehat{\beta}_{L},\widehat{\rho}_{L})^{\prime}$ (or $\widehat{\beta}_{L}$) is our bias-corrected MLE with truncation parameter $\tau=1$ for the dynamic model or $\tau=0$ for the static model; and $(\widehat{\beta}_{J},\widehat{\rho}_{J})^{\prime}$ (or $\widehat{\beta}_{J}$) represents the Jackknife bias-corrected estimators.
 The number of replications is $1000$.
@@
 Figure 2: Heterogeneous Impact of Children on Single Mother Labor Force Participation
+6 Final Remarks
+In this paper, we propose a likelihood-based analytical bias correction procedure for a general class of dynamic nonlinear heterogeneous parameter models subject to the incidental parameter problem.
+We give the analytical form of a corrected likelihood and show that such a corrected likelihood delivers point estimates that are asymptotically unbiased and test statistics that are asymptotically $\chi^{2}$-distributed.
+Simulation studies and empirical analyses support our claims.
+Several issues are, however, not formally addressed in this paper.
+We conjecture, but have not formally investigated, that the average partial effects associated these models are asymptotically unbiased.
+Specifically, we conjecture that the rate of convergence of the average partial effects are only $O(1/\sqrt{T})$ or $O(1/\sqrt{N})$, slower than the potential bias, indicating that bias corrections may have no benefit in the limit.
+This result is formally shown by Fernández-Val and Weidner (2016) for models with additive fixed effects in the intercept.
+Second, we have not applied our procedure to general M-estimation models.
+We intuitively believe that our approach applies to them as long as the objective function is differentiable, only with minor modifications regarding the information matrix equality.
+We would like to dedicate these extensions to future research.
 Acknowledgement
-We would like to express our gratitude to valuable comments and suggestions from Geert Dhaene, Jinyong Hahn, Chen Hsiao, Whitney Newey, and Jun Yu.
+We would like to express our gratitude to valuable comments and suggestions from Geert Dhaene, Yanqin Fan, Jinyong Hahn, Chen Hsiao, Whitney Newey, and Jun Yu.
 Our research assistant Yuchi Liu helped us with various tasks for the empirical analysis.
@@
 A Proofs
-A.1 Lemmas
-Proof of Lemma 1.
-Note that ${H}(\theta)=D\big{[}\partial_{\phi\phi^{\prime}}l(\theta,\phi)\mid_{\phi=\phi(\theta)}\big{]}D^{\prime},$ where $D$ is a $(N+T-2)K\times(N+T)K$ matix.
-Since $K$ is fixed, we assume without loss of generality that $K=1.$ We express $H(\theta)$ as:
-$H(\theta)=\begin{pmatrix}H_{\breve{\alpha}\breve{\alpha}^{\prime}}\left(\theta\right),&H_{\breve{\alpha}\breve{\gamma}^{\prime}}\left(\theta\right)\\ H_{\breve{\gamma}\breve{\alpha}^{\prime}}\left(\theta\right),&H_{\breve{\gamma}\breve{\gamma}^{\prime}}\left(\theta\right)\end{pmatrix}.$
-Consider any element $h_{i,j}^{\breve{\alpha}}(\theta)$ of $N\cdot H_{\breve{\alpha}\breve{\alpha}^{\prime}}(\theta),$ where $i,j=1,\ldots,N-1.$ Under Assumptions 1 and 1, by applying Davydov’s inequality, it follows that
-$\sup_{i,j}\sup_{\theta\in\Theta}\text{Var}\big{(}h_{i,j}^{\breve{\alpha}}(\theta)\big{)}=o(1)$
-as $N$ and $T\rightarrow\infty.$ Consequently, by the Markov inequality,
-$\sup_{i,j}\sup_{\theta\in\Theta}\big{\{}h_{i,j}^{\breve{\alpha}}(\theta)-\mathbb{E}[h_{i,j}^{\breve{\alpha}}(\theta)]\big{\}}=o_{\mathbb{P}}(1).$
-Let $h_{t,s}^{\breve{\gamma}}(\theta)$ denote the $(t,s)$-th element of $T\cdot H_{\breve{\gamma}\breve{\gamma}^{\prime}}(\theta)$ for $t,s=1,\ldots,T-1.$ By conditional cross-sectional independence, we also have
-$\sup_{t,s}\sup_{\theta\in\Theta}\big{\{}h_{t,s}^{\breve{\gamma}}(\theta)-\mathbb{E}[h_{t,s}^{\breve{\gamma}}(\theta)]\big{\}}=o_{\mathbb{P}}(1).$
-Moreover, for each $h_{i,t}^{\breve{\alpha}\breve{\gamma}}(\theta),$ the $(i,t)$-th element of $NT\cdot H_{\breve{\alpha}\breve{\gamma}^{\prime}}(\theta),$ with $i=1,\ldots,N-1$ and $t=1,\ldots,T-1,$ by Assumption 1 we have that
-$\sup_{i,t}\sup_{\theta\in\Theta}\big{\{}h_{i,t}^{\breve{\alpha}\breve{\gamma}}(\theta)-\mathbb{E}[h_{i,t}^{\breve{\alpha}\breve{\gamma}}(\theta)]\big{\}}=o_{\mathbb{P}}(1).$
-Therefore,
-$N\left\|H_{\breve{\alpha}\breve{\alpha}^{\prime}}(\theta)-\mathbb{E}[H_{\breve{\alpha}\breve{\alpha}^{\prime}}(\theta)]\right\|_{\max}=o_{\mathbb{P}}(1),\quad T\left\|H_{\breve{\gamma}\breve{\gamma}^{\prime}}(\theta)-\mathbb{E}[H_{\breve{\gamma}\breve{\gamma}^{\prime}}(\theta)]\right\|_{\max}=o_{\mathbb{P}}(1),$
-and
-$NT\left\|H_{\breve{\alpha}\breve{\gamma}^{\prime}}(\theta)-\mathbb{E}[H_{\breve{\alpha}\breve{\gamma}^{\prime}}(\theta)]\right\|_{\max}=o_{\mathbb{P}}(1),$
-uniformly for $\theta\in\Theta,$ where $\|A\|_{\max}$ denotes the max norm of a matrix $A$.
-Since $\mathbb{E}[s(\theta)]=0,$ $s(\theta)=s(\theta)-\mathbb{E}[s(\theta)]=:(\widetilde{s}_{\breve{\alpha}}^{\prime}(\theta),\widetilde{s}_{\breve{\gamma}}^{\prime}(\theta))^{\prime}.$ We now decompose the quadratic form on the right hand side of (2.7):
-$s^{\prime}(\theta)[-{H}(\theta)]^{-1}s\left(\theta\right)$
+In this appendix, we prove the asymptotic properties of $L(\theta)$, the corrected likelihood, $\widehat{\theta}_{L}$ the corrected estimator, and $\{\widehat{\xi}_{LR},\widehat{\xi}_{LM},\widehat{\xi}_{Wald}\}$, the three test statistics we consider in this paper.
+To continue, we first introduce three Lemmas.
+These lemmas are useful in establishing Theorems 1 (asymptotic bias expansion) and 2 (corrected likelihood).
+The proof of Corollaries 1 (estimators) and 2 (tests) then follow.
+Define the partition
+$H(\theta)=:\begin{pmatrix}H_{\alpha\alpha^{\prime}}\left(\theta\right),&H_{\alpha\gamma^{\prime}}\left(\theta\right)\\ H_{\gamma\alpha^{\prime}}\left(\theta\right),&H_{\gamma\gamma^{\prime}}\left(\theta\right)\end{pmatrix},$
+where $H_{\alpha\alpha^{\prime}}\left(\theta\right)$ is $\left(N-1\right)K\times\left(N-1\right)K$, $H_{\gamma\gamma^{\prime}}\left(\theta\right)$ is $\left(T-1\right)K\times\left(T-1\right)K$, $H_{\alpha\gamma^{\prime}}\left(\theta\right)$ is $\left(N-1\right)K\times\left(T-1\right)K$, and $H_{\gamma\alpha^{\prime}}\left(\theta\right)=\left[H_{\alpha\gamma^{\prime}}\left(\theta\right)\right]^{\prime}$.
+The next lemma states that the quadratic form $\mathbb{E}\left\{s^{\prime}(\theta)[{H}\left(\theta\right)]^{-1}s(\theta)\right\}$ can be decomposed into two dominating terms of order $O(1/\sqrt{NT})$.
+They establish the two additive bias terms $B_{\alpha}(\theta)$ and $B_{\gamma}(\theta)$, arising from estimating individual effects and time effects, respectively.
+Lemma A.1
+Under Assumption 1, we have
+$\mathbb{E}\left\{s^{\prime}(\theta)[{H}\left(\theta\right)]^{-1}s(\theta)\right\}=\mathbb{E}\left\{\widetilde{s}_{\alpha}^{\prime}(\theta)H_{\alpha\alpha^{\prime}}^{\ast}(\theta)\widetilde{s}_{\alpha}(\theta)\right\}+\mathbb{E\{}\widetilde{s}_{\gamma}^{\prime}(\theta)H_{\gamma\gamma^{\prime}}^{\ast}(\theta)\widetilde{s}_{\gamma}(\theta)\}+o(1/\sqrt{NT}),$
+where $o(\cdot)$ is uniform in $\theta$, and
+$\sup_{\theta\in\Theta}\mathbb{E}\left\{\widetilde{s}_{\alpha}^{\prime}(\theta)H_{\alpha\alpha^{\prime}}^{\ast}(\theta)\widetilde{s}_{\alpha}(\theta)\right\}=O\left(1/T\right),\qquad\sup_{\theta\in\Theta}\mathbb{E}\left[\widetilde{s}_{\gamma}^{\prime}(\theta)H_{\gamma\gamma^{\prime}}^{\ast}(\theta)\widetilde{s}_{\gamma}(\theta)\right]=O(1/N).$
+Proof. Seeing $\mathbb{E}s(\theta)=0$, the claim without $\mathbb{E}$ is
+$\quad\quad s^{\prime}(\theta)[-{H}(\theta)]^{-1}s\left(\theta\right)$
 $=[s(\theta)-\mathbb{E}s(\theta)]^{\prime}[-{H}\left(\theta\right)]^{-1}[s(\theta)-\mathbb{E}s(\theta)]$
 $=[s(\theta)-\mathbb{E}s(\theta)]^{\prime}[-\mathbb{E}{H}\left(\theta\right)]^{-1}[s(\theta)-\mathbb{E}s(\theta)]^{\prime}(1+o_{\mathbb{P}}(1))$
-$=$ $-\big{[}\widetilde{s}_{\breve{\alpha}}^{\prime}(\theta)H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\alpha}}(\theta)+\widetilde{s}_{\breve{\gamma}}^{\prime}(\theta)H_{\breve{\gamma}\breve{\gamma}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\gamma}}(\theta)+2\widetilde{s}_{\breve{\alpha}}^{\prime}(\theta)H_{\breve{\alpha}\breve{\gamma}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\gamma}}\left(\theta\right)\big{]}(1+o_{\mathbb{P}}(1)).$ (A.1)
-By matrix inversion in block form, it follows that
-$\sup_{\theta\in\Theta}\left\|H_{\breve{\alpha}\breve{\gamma}^{\prime}}^{\ast}(\theta)\right\|_{\max}=O(1),$
+$=-\left[\widetilde{s}_{\alpha}^{\prime}(\theta)H_{\alpha\alpha^{\prime}}^{\ast}(\theta)\widetilde{s}_{\alpha}(\theta)+\widetilde{s}_{\gamma}^{\prime}(\theta)H_{\gamma\gamma^{\prime}}^{\ast}(\theta)\widetilde{s}_{\gamma}(\theta)+2\widetilde{s}_{\alpha}^{\prime}(\theta)H_{\alpha\gamma^{\prime}}^{\ast}(\theta)\widetilde{s}_{\gamma}\left(\theta\right)\right](1+o_{\mathbb{P}}(1)).$ (A.1)
+Since $K$ is fixed, we assume without loss of generality that $K=1$. In what follows, we show
+$\sup_{\theta\in\Theta}\mathbb{E}\left[\widetilde{s}_{\alpha}^{\prime}(\theta)H_{\alpha\gamma^{\prime}}^{\ast}(\theta)\widetilde{s}_{\gamma}\left(\theta\right)\right]=O\left(1/(NT)\right),$ (A.2)
 and
-$\sup_{\theta\in\Theta}\left\|H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)\right\|_{\max}=O(T),\quad\sup_{\theta\in\Theta}\left\|H_{\breve{\gamma}\breve{\gamma}^{\prime}}^{\ast}\left(\theta\right)\right\|_{\max}=O(N).$
-In the following, we show
-$\sup_{\theta\in\Theta}\mathbb{E}\big{[}\widetilde{s}_{\breve{\alpha}}^{\prime}(\theta)H_{\breve{\alpha}\breve{\gamma}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\gamma}}(\theta)\big{]}=O\big{(}1/(NT)\big{)},$ (A.2)
-and
-$\sup_{\theta\in\Theta}\,\mathbb{E}\big{[}\widetilde{s}_{\breve{\alpha}}^{\prime}(\theta)H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}\left(\theta\right)\widetilde{s}_{\breve{\alpha}}(\theta)\big{]}=O(1/T),\quad\sup_{\theta\in\Theta}\,\mathbb{E}\big{[}\widetilde{s}_{\breve{\gamma}}^{\prime}(\theta)H_{\breve{\gamma}\breve{\gamma}^{\prime}}^{\ast}\left(\theta\right)\widetilde{s}_{\breve{\gamma}}(\theta)\big{]}=O(1/N).$ (A.3)
-Therefore, $\mathbb{E}\big{[}\widetilde{s}_{\breve{\alpha}}^{\prime}\left(\theta\right)H_{\breve{\alpha}\breve{\gamma}^{\prime}}^{\ast}\left(\theta\right)\widetilde{s}_{\breve{\gamma}}(\theta)\big{]}$ is negligible in the expansion for $\mathbb{E}[\widehat{l}(\theta)]$ compared to the other two terms from the quadratic form.
-Let
-$\partial_{\alpha_{i}}\widetilde{l}_{i,t}(\theta,\phi(\theta)):=\{\partial_{\alpha_{i}}l_{i,t}(\theta,\phi)-\mathbb{E}[\partial_{\alpha_{i}}l_{i,t}(\theta,\phi)]\}|_{\phi=\phi(\theta)}$
-and
-$\partial_{\gamma_{t}}\widetilde{l}_{i,t}(\theta,\phi(\theta)):=\{\partial_{\gamma_{t}}l_{i,t}(\theta,\phi)-\mathbb{E}[\partial_{\gamma_{t}}l_{i,t}(\theta,\phi)]\}|_{\phi=\phi(\theta)}$
-for $i=1,\ldots,N$ and $t=1,\ldots,T,$ respectively.
-Define $h_{i,t}^{\ast\breve{\alpha}\breve{\gamma}}(\theta)$ as the $(i,t)$-th element of $H_{\breve{\alpha}\breve{\gamma}^{\prime}}^{\ast}(\theta)$ where $i=1,\ldots,N-1$ and $t=1,\ldots,T-1.$ Under the assumption of conditional cross-sectional independence, we have
-$\mathbb{E}\big{\{}(N\sqrt{T}\widetilde{s}_{\breve{\alpha}}^{\prime}\left(\theta\right))H_{\breve{\alpha}\breve{\gamma}^{\prime}}^{\ast}\left(\theta\right)(\sqrt{N}T\widetilde{s}_{\breve{\gamma}}(\theta))\big{\}}$
-$=$ $\frac{1}{\sqrt{NT}}\mathbb{E}\bigg{\{}\sum_{i=1}^{N-1}\sum_{t=1}^{T-1}h_{i,t}^{\ast\breve{\alpha}\breve{\gamma}}(\theta)\cdot\bigg{[}\sum_{s=1}^{T}\big{(}\partial_{\alpha_{i}}\widetilde{l}_{i,s}(\theta,\phi(\theta))-\partial_{\alpha_{N}}\widetilde{l}_{N,s}(\theta,\phi(\theta))\big{)}\bigg{]}$
-$\hskip 120.00018pt\cdot\bigg{[}\sum_{j=1}^{N}\big{(}\partial_{\gamma_{t}}\widetilde{l}_{j,t}(\theta,\phi(\theta))-\partial_{\gamma_{T}}\widetilde{l}_{j,T}(\theta,\phi(\theta))\big{)}\bigg{]}\bigg{\}}$
-$=$ $\frac{1}{\sqrt{NT}}\sum_{i=1}^{N-1}\mathbb{E}\left[\bigg{(}\sum_{s=1}^{T}\partial_{\alpha_{i}}\widetilde{l}_{i,s}(\theta,\phi(\theta))\bigg{)}\bigg{(}\sum_{t=1}^{T-1}\partial_{\gamma_{t}}\widetilde{l}_{i,t}(\theta,\phi(\theta))\cdot h_{i,t}^{\ast\breve{\alpha}\breve{\gamma}}(\theta)\bigg{)}\right]$
-$\quad-\frac{1}{\sqrt{NT}}\sum_{i=1}^{N-1}\left[\bigg{(}\sum_{t=1}^{T-1}h_{i,t}^{\ast\breve{\alpha}\breve{\gamma}}(\theta)\bigg{)}\mathbb{E}\bigg{(}\sum_{s=1}^{T}\partial_{\alpha_{i}}\widetilde{l}_{i,s}(\theta,\phi(\theta))\partial_{\gamma_{T}}\widetilde{l}_{i,T}(\theta,\phi(\theta))\bigg{)}\right]$
-$\quad-\frac{1}{\sqrt{NT}}\mathbb{E}\left[\bigg{(}\sum_{s=1}^{T}\partial_{\alpha_{N}}\widetilde{l}_{N,s}(\theta,\phi(\theta))\bigg{)}\left(\sum_{t=1}^{T-1}\partial_{\gamma_{t}}\widetilde{l}_{N,t}(\theta,\phi(\theta))\bigg{(}\sum_{i=1}^{N-1}h_{i,t}^{\ast\breve{\alpha}\breve{\gamma}}(\theta)\bigg{)}\right)\right]$
-$\quad+\frac{1}{\sqrt{NT}}\bigg{(}\sum_{i=1}^{N-1}\sum_{t=1}^{T-1}h_{i,t}^{\ast\breve{\alpha}\breve{\gamma}}(\theta)\bigg{)}\mathbb{E}\bigg{(}\sum_{s=1}^{T}\partial_{\alpha_{N}}\widetilde{l}_{N,s}(\theta,\phi(\theta))\partial_{\gamma_{T}}\widetilde{l}_{N,T}(\theta,\phi(\theta))\bigg{)}$
-=:absent:=:= : $A_{NT,1}(\theta)-A_{NT,2}(\theta)-A_{NT,3}(\theta)+A_{NT,4}(\theta).$ (A.4)
-By Assumptions 1 and 1, using Davydov’s inequality, we obtain for some $\eta>2$ that
-$|A_{NT,1}(\theta)|$
-$=$ $\frac{1}{\sqrt{NT}}\sum_{i=1}^{N-1}\left|\mathbb{E}\left[\bigg{(}\sum_{s=1}^{T-1}\partial_{\alpha_{i}}\widetilde{l}_{i,s}(\theta,\phi(\theta))\bigg{)}\bigg{(}\sum_{t=1}^{T-1}\partial_{\gamma_{t}}\widetilde{l}_{i,t}(\theta,\phi(\theta))\cdot h_{i,t}^{\ast\breve{\alpha}\breve{\gamma}}(\theta)\bigg{)}\right.\right.$
-$\hskip 56.9055pt\left.\left.+\sum_{t=1}^{T-1}\partial_{\gamma_{t}}\widetilde{l}_{i,t}(\theta,\phi(\theta))\partial_{\alpha_{i}}\widetilde{l}_{i,T}(\theta,\phi(\theta))\cdot h_{i,t}^{\ast\breve{\alpha}\breve{\gamma}}(\theta)\right]\right|$
-$\leq$ $\sqrt{NT}\sup_{\theta\in\Theta}\left\|H_{\breve{\alpha}\breve{\gamma}^{\prime}}^{\ast}(\theta)\right\|_{\max}\cdot\sup_{i,t}\{\mathbb{E}[g^{\eta}(w_{it})]\}^{2/\eta}$
-$\quad\cdot\sup_{i}\left([a_{i}(0)]^{1-2/\eta}+2\sum_{m=1}^{T-2}\big{(}1-m/(T-1)\big{)}[a_{i}(m)]^{1-2/\eta}+\frac{1}{T-1}\sum_{m=1}^{T-1}[a_{i}(m)]^{1-2/\eta}\right)$
-$=$ $O\big{(}\sqrt{NT}\big{)},$
-$|A_{NT,2}(\theta)|$ $=\frac{T}{\sqrt{NT}}\sum_{i=1}^{N-1}\bigg{(}\frac{1}{T}\sum_{t=1}^{T-1}|h_{i,t}^{\ast\breve{\alpha}\breve{\gamma}}(\theta)|\bigg{)}\cdot\left|\mathbb{E}\bigg{(}\sum_{s=1}^{T}\partial_{\alpha_{i}}\widetilde{l}_{i,s}(\theta,\phi(\theta))\partial_{\gamma_{T}}\widetilde{l}_{i,T}(\theta,\phi(\theta))\bigg{)}\right|$
-$\leq\sqrt{NT}\sup_{\theta\in\Theta}\left\|H_{\breve{\alpha}\breve{\gamma}^{\prime}}^{\ast}(\theta)\right\|_{\max}\cdot\sup_{i,t}\{\mathbb{E}[g^{\eta}(w_{it})]\}^{2/\eta}\cdot\sup_{i}\left(\sum_{m=0}^{T-1}[a_{i}(m)]^{1-2/\eta}\right)=O\big{(}\sqrt{NT}\big{)},$
-$|A_{NT,3}(\theta)|$
-$=$ $\frac{N}{\sqrt{NT}}\left|\mathbb{E}\left[\left(\sum_{s=1}^{T}\partial_{\alpha_{N}}\widetilde{l}_{N,s}(\theta,\phi(\theta))\right)\left(\sum_{t=1}^{T-1}\partial_{\gamma_{t}}\widetilde{l}_{N,t}(\theta,\phi(\theta))\bigg{(}\frac{1}{N}\sum_{i=1}^{N-1}h_{i,t}^{\ast\breve{\alpha}\breve{\gamma}}(\theta)\bigg{)}\right)\right]\right|$
-$\leq$ $\sqrt{NT}\sup_{\theta\in\Theta}\left\|H_{\breve{\alpha}\breve{\gamma}^{\prime}}^{\ast}(\theta)\right\|_{\max}\cdot\sup_{i,t}\{\mathbb{E}[g^{\eta}(w_{it})]\}^{2/\eta}$
-$\quad\cdot\sup_{i}\left([a_{i}(0)]^{1-2/\eta}+2\sum_{m=1}^{T-2}\big{(}1-m/(T-1)\big{)}[a_{i}(m)]^{1-2/\eta}+\frac{1}{T-1}\sum_{m=1}^{T-1}[a_{i}(m)]^{1-2/\eta}\right)$
-$=$ $O\big{(}\sqrt{NT}\big{)},$
-and
-$|A_{NT,4}(\theta)|\leq\sqrt{NT}\sup_{\theta\in\Theta}\left\|H_{\breve{\alpha}\breve{\gamma}^{\prime}}^{\ast}(\theta)\right\|_{\max}\cdot\sup_{i,t}\{\mathbb{E}[g^{\eta}(w_{it})]\}^{2/\eta}\cdot\sup_{i}\left(\sum_{m=0}^{T-1}[a_{i}(m)]^{1-2/\eta}\right)=O\big{(}\sqrt{NT}\big{)}.$
-Thus, we obtain (A.2) from (A.4).
-Let $h_{i,j}^{\ast\breve{\alpha}}(\theta)$ denote the $(i,j)$-th element of $T^{-1}H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta),$ where $i,j=1,\ldots,N-1.$ By Assumption 1, we know that $\sup_{i\geq 1}h_{i,i}^{\ast\breve{\alpha}}(\theta)=O(1)$ uniformly in $\theta.$ Define $\alpha:=(\alpha_{1}^{\prime},\ldots,\alpha_{N}^{\prime})^{\prime}$ and $\gamma:=(\gamma_{1}^{\prime},\ldots,\gamma_{T}^{\prime})^{\prime}.$ Noting that
-$N\cdot H_{\breve{\alpha}\breve{\alpha}^{\prime}}(\theta)=D_{1}[N\cdot\partial_{\alpha\alpha^{\prime}}l(\theta,\phi)|_{\phi=\phi(\theta)}]D_{1}^{\prime},$
-$N\cdot H_{\breve{\alpha}\breve{\alpha}^{\prime}}(\theta)$ can be written as the sum of a diagonal matrix and a rank-one matrix.
-Thus, it follows from the lemma in Miller (1981) that $\sup_{i\neq j}h_{i,j}^{\ast\breve{\alpha}}(\theta)=O(1/N)$ uniformly in $\theta.$ Since $s(\theta)=D\partial_{\phi}l(\theta,\phi)|_{\phi=\phi(\theta)},$ we have
-$T\mathbb{E}\big{[}\widetilde{s}_{\breve{\alpha}}^{\prime}(\theta)H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\alpha}}(\theta)\big{]}$
-$=$ $\frac{1}{N}\mathbb{E}\big{\{}(N\sqrt{T}\widetilde{s}_{\breve{\alpha}}^{\prime}(\theta))[N^{-1}H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)](N\sqrt{T}\widetilde{s}_{\breve{\alpha}}(\theta))\big{\}}$
-$=$ $\frac{1}{N}\sum_{i=1}^{N-1}h_{i,i}^{\ast\breve{\alpha}}(\theta)\,\mathbb{E}\bigg{(}\frac{1}{\sqrt{T}}\sum_{t=1}^{T}\big{[}\partial_{\alpha_{i}}\widetilde{l}_{i,t}(\theta,\phi(\theta))-\partial_{\alpha_{N}}\widetilde{l}_{N,t}(\theta,\phi(\theta))\big{]}\bigg{)}^{2}$
-$+\frac{2}{N}\sum_{1\leq i<j\leq N-1}h_{i,j}^{\ast\breve{\alpha}}(\theta)\,\mathbb{E}\left[\bigg{(}\frac{1}{\sqrt{T}}\sum_{t=1}^{T}\big{[}\partial_{\alpha_{i}}\widetilde{l}_{i,t}(\theta,\phi(\theta))-\partial_{\alpha_{N}}\widetilde{l}_{N,t}(\theta,\phi(\theta))\big{]}\bigg{)}\right.$
-$\hskip 116.65646pt\left.\cdot\bigg{(}\frac{1}{\sqrt{T}}\sum_{t=1}^{T}\big{(}\partial_{\alpha_{j}}\widetilde{l}_{j,t}(\theta,\phi(\theta))-\partial_{\alpha_{N}}\widetilde{l}_{N,t}(\theta,\phi(\theta))\big{)}\bigg{)}\right]$
-=:absent:=:= : $I_{NT,1}(\theta)+I_{NT,2}(\theta).$ (A.5)
-We now show that $I_{NT,1}(\theta)=O(1)$ uniformly for $\theta.$ By Assumptions 1 and 1, applying Davydov’s inequality gives that for some $\eta>2,$
-$\sup_{\theta\in\Theta}|I_{NT,1}(\theta)|$
-$=$ $\sup_{\theta\in\Theta}\frac{1}{NT}\sum_{i=1}^{N-1}|h_{i,i}^{\ast\breve{\alpha}}(\theta)|$
-$\hskip 42.67912pt\cdot\bigg{|}\sum_{t=1}^{T}\sum_{s=1}^{T}\mathbb{E}\big{[}\partial_{\alpha_{i}}\widetilde{l}_{i,t}(\theta,\phi(\theta))\partial_{\alpha_{i}}\widetilde{l}_{i,s}(\theta,\phi(\theta))\big{]}+\sum_{t=1}^{T}\sum_{s=1}^{T}\mathbb{E}\big{[}\partial_{\alpha_{N}}\widetilde{l}_{N,t}(\theta,\phi(\theta))\partial_{\alpha_{N}}\widetilde{l}_{N,s}(\theta,\phi(\theta))\big{]}\bigg{|}$
-$\leq\bigg{(}\frac{1}{N}\sum_{i=1}^{N-1}|h_{i,i}^{\ast\breve{\alpha}}(\theta)|\bigg{)}\cdot\sup_{i,t}\{\mathbb{E}[g^{\eta}(w_{it})]\}^{2/\eta}\cdot\sup_{i}\left([a_{i}(0)]^{1-2/\eta}+2\sum_{m=1}^{T-1}(1-m/T)\cdot[a_{i}(m)]^{1-2/\eta}\right)$
-$=O(1)$
-as $N$ and $T\rightarrow\infty,$ where the conditional cross-sectional independence leads to the first equality. Similarly, for $I_{NT,2}(\theta),$ we have
-$\sup_{\theta\in\Theta}|I_{NT,2}(\theta)|$
-$=$ $\frac{2}{N}\sum_{1\leq i<j\leq N-1}|h_{i,j}^{\ast\breve{\alpha}}(\theta)|\cdot\mathbb{E}\left[\bigg{(}\frac{1}{\sqrt{T}}\sum_{t=1}^{T}\partial_{\alpha_{N}}\widetilde{l}_{N,t}(\theta,\phi(\theta))\bigg{)}^{2}\right]$
-$\leq$ $\bigg{(}\frac{2}{N}\sum_{1\leq i<j\leq N-1}|h_{i,j}^{\ast\breve{\alpha}}(\theta)|\bigg{)}\cdot\sup_{i,t}\{\mathbb{E}[g^{\eta}(w_{it})]\}^{2/\eta}\cdot\sup_{i}\left([a_{i}(0)]^{1-2/\eta}+2\sum_{m=1}^{T-1}(1-m/T)\cdot[a_{i}(m)]^{1-2/\eta}\right)$
-$=$ $O(1),$
-noting that $\sup_{i\neq j}h_{i,j}^{\ast\breve{\alpha}}(\theta)=O(1/N)$ uniformly in $\theta.$ Thus, from (A.5), we obtain the first result in (A.3).
-Similarly, we can derive the other result in (A.3).
-Now, combining (A.1), (A.2), and (A.3) yields that
-$\mathbb{E}\big{\{}s^{\prime}(\theta)[-{H}(\theta)]^{-1}s(\theta)\big{\}}=$ $-\mathbb{E}\big{\{}s_{\breve{\alpha}}^{\prime}(\theta)H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)s_{\breve{\alpha}}(\theta)\big{\}}-\mathbb{E}\big{\{}s_{\breve{\gamma}}^{\prime}(\theta)H_{\breve{\gamma}\breve{\gamma}^{\prime}}^{\ast}(\theta)s_{\breve{\gamma}}(\theta)\big{\}}$
-$-\mathbb{E}\left\{\big{[}s_{\breve{\alpha}}^{\prime}(\theta)H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)s_{\breve{\alpha}}(\theta)+s_{\breve{\gamma}}^{\prime}\left(\theta\right)H_{\breve{\gamma}\breve{\gamma}^{\prime}}^{\ast}\left(\theta\right)s_{\breve{\gamma}}(\theta)\big{]}\cdot o_{\mathbb{P}}(1)\right\}.$
-Note that, by Assumption 1, for $S\in\{0,1,2\},$
-$\sup_{i,t}\mathbb{E}\left(|\partial_{\phi_{1}\ldots\phi_{S}}l_{it}\left(\theta,\phi\right)|\mathds{1}\{|\partial_{\phi_{1}\ldots\phi_{S}}l_{it}\left(\theta,\phi\right)|>a\}\right)\rightarrow 0$
+$\sup_{\theta\in\Theta}\mathbb{E}\left[\widetilde{s}_{\alpha}^{\prime}(\theta)H_{\alpha\alpha^{\prime}}^{\ast}(\theta)\widetilde{s}_{\alpha}(\theta)\right]=O(1/T),\qquad\sup_{\theta\in\Theta}\mathbb{E}\left[\widetilde{s}_{\gamma}^{\prime}(\theta)H_{\gamma\gamma^{\prime}}^{\ast}(\theta)\widetilde{s}_{\gamma}(\theta)\right]=O(1/N).$ (A.3)
+Thus, $\mathbb{E}\left[\widetilde{s}_{\alpha}^{\prime}(\theta)H_{\alpha\gamma^{\prime}}^{\ast}(\theta)\widetilde{s}_{\gamma}\left(\theta\right)\right]$ is negligible in the expansion relative to the other two terms from the quadratic form.
+First, under Assumptions 1 and 1, and by Davydov’s inequality, it follows that $\sup_{i,j}\sup_{\theta\in\Theta}\operatorname*{Var}[h_{i,j}^{\alpha}(\theta)]=o(1)$ as $N,T\rightarrow\infty$, where $\operatorname*{Var}$ denotes the variance and $h_{i,j}^{\alpha}(\theta)$ is the $\left(i,j\right)$-th element of $N\cdot H_{\alpha\alpha^{\prime}}\left(\theta\right)$ with $i,j=1,\ldots,N-1$.
+Invoking the Markov’s inequality,
+$\sup_{i,j}\sup_{\theta\in\Theta}\left\{h_{i,j}^{\alpha}(\theta)-\mathbb{E}[h_{i,j}^{\alpha}(\theta)]\right\}=o_{\mathbb{P}}(1).$
+Similarly, we have
+$\sup_{t,s}\sup_{\theta\in\Theta}\left\{h_{t,s}^{\gamma}(\theta)-\mathbb{E}[h_{t,s}^{\gamma}(\theta)]\right\}$ $=o_{\mathbb{P}}(1),$
+$\sup_{i,t}\sup_{\theta\in\Theta}\left\{h_{i,t}^{\alpha\gamma}(\theta)-\mathbb{E}[h_{i,t}^{\alpha\gamma}(\theta)]\right\}$ $=o_{\mathbb{P}}(1),$
+where $h_{t,s}^{\gamma}(\theta)$ denote the $(t,s)$-th element of $T\cdot H_{\gamma\gamma^{\prime}}(\theta)$, with $t,s=1,\ldots,T-1$, and $h_{i,t}^{\alpha\gamma}(\theta)$ the $(i,t)$-th element of $NT\cdot H_{\alpha\gamma^{\prime}}(\theta)$, for $i=1,\ldots,N-1$ and $t=1,\ldots,T-1$.
+Therefore, we have
+$N\left\|H_{\alpha\alpha^{\prime}}(\theta)-\mathbb{E}[H_{\alpha\alpha^{\prime}}(\theta)]\right\|_{\max}$ $=o_{\mathbb{P}}(1),\qquad T\left\|H_{\gamma\gamma^{\prime}}(\theta)-\mathbb{E}[H_{\gamma\gamma^{\prime}}(\theta)]\right\|_{\max}=o_{\mathbb{P}}(1),$
+$NT\left\|H_{\alpha\gamma^{\prime}}(\theta)-\mathbb{E}[H_{\alpha\gamma^{\prime}}(\theta)]\right\|_{\max}$ $=o_{\mathbb{P}}(1),$
+uniformly for $\theta\in\Theta$, where $\|A\|_{\max}$ denotes the max norm of a matrix $A$. In addition, by matrix inversion in block form, it follows that
+$\sup_{\theta\in\Theta}\left\|H_{\alpha\alpha^{\prime}}^{\ast}(\theta)\right\|_{\max}$ $=O(T),\qquad\sup_{\theta\in\Theta}\left\|H_{\gamma\gamma^{\prime}}^{\ast}\left(\theta\right)\right\|_{\max}=O(N),$
+$\sup_{\theta\in\Theta}\left\|H_{\alpha\gamma^{\prime}}^{\ast}(\theta)\right\|_{\max}$ $=O(1).$
+For $i=1,\ldots,N-1$ and $t=1,\ldots,T-1$, define $h_{i,t}^{\ast\alpha\gamma}(\theta)$ as the $(i,t)$-th element of $H_{\alpha\gamma^{\prime}}^{\ast}(\theta)$. We first show (A.2). We have
+$\quad\quad\mathbb{E}\left\{(N\sqrt{T}\widetilde{s}_{\alpha}^{\prime}\left(\theta\right))H_{\alpha\gamma^{\prime}}^{\ast}\left(\theta\right)(\sqrt{N}T\widetilde{s}_{\gamma}(\theta))\right\}$
+$=\frac{1}{\sqrt{NT}}\mathbb{E}\left\{\sum_{i=1}^{N-1}\sum_{t=1}^{T-1}h_{i,t}^{\ast\alpha\gamma}(\theta)\cdot\left[\sum_{s=1}^{T}\widetilde{s}_{i,s}^{\alpha}(\theta)-\widetilde{s}_{N,s}^{\alpha}(\theta)\right]\left[\sum_{j=1}^{N}\widetilde{s}_{j,t}^{\gamma}(\theta)-\widetilde{s}_{j,T}^{\gamma}(\theta)\right]\right\}$
+$=\frac{1}{\sqrt{NT}}\sum_{i=1}^{N-1}\mathbb{E}\left[\left(\sum_{s=1}^{T}\widetilde{s}_{i,s}^{\alpha}(\theta)\right)\left(\sum_{t=1}^{T-1}\widetilde{s}_{i,t}^{\gamma}(\theta)h_{i,t}^{\ast\alpha\gamma}(\theta)\right)\right]$
+$\quad\quad-\frac{1}{\sqrt{NT}}\sum_{i=1}^{N-1}\left[\left(\sum_{t=1}^{T-1}h_{i,t}^{\ast\alpha\gamma}(\theta)\right)\mathbb{E}\left(\sum_{s=1}^{T}\widetilde{s}_{i,s}^{\alpha}(\theta)\widetilde{s}_{j,T}^{\gamma}(\theta)\right)\right]$
+$\quad\quad-\frac{1}{\sqrt{NT}}\mathbb{E}\left\{\left(\sum_{s=1}^{T}\widetilde{s}_{N,s}^{\alpha}(\theta)\right)\left[\sum_{t=1}^{T-1}\widetilde{s}_{N,t}^{\gamma}(\theta)\left(\sum_{i=1}^{N-1}h_{i,t}^{\ast\alpha\gamma}(\theta)\right)\right]\right\}$
+$\quad\quad+\frac{1}{\sqrt{NT}}\left(\sum_{i=1}^{N-1}\sum_{t=1}^{T-1}h_{i,t}^{\ast\alpha\gamma}(\theta)\right)\mathbb{E}\left(\sum_{s=1}^{T}\widetilde{s}_{N,s}^{\alpha}(\theta)\widetilde{s}_{N,T}^{\gamma}(\theta)\right)$
+$=:A_{NT,1}(\theta)-A_{NT,2}(\theta)-A_{NT,3}(\theta)+A_{NT,4}(\theta).$ (A.4)
+By Assumptions 1 and 1, and using Davydov’s inequality, we obtain for some $\eta>2$ that
+$\left|A_{NT,1}(\theta)\right|$ $=\frac{1}{\sqrt{NT}}\sum_{i=1}^{N-1}\left|\mathbb{E}\left[\left(\sum_{s=1}^{T-1}\widetilde{s}_{i,s}^{\alpha}(\theta)\right)\left(\sum_{t=1}^{T-1}\widetilde{s}_{i,t}^{\gamma}(\theta)h_{i,t}^{\ast\alpha\gamma}(\theta)\right)+\sum_{t=1}^{T-1}\widetilde{s}_{i,T}^{\alpha}(\theta)\widetilde{s}_{i,t}^{\gamma}(\theta)h_{i,t}^{\ast\alpha\gamma}(\theta)\right]\right|$
+$\leq\sqrt{NT}\left(\sup_{\theta\in\Theta}\left\|H_{\alpha\gamma^{\prime}}^{\ast}(\theta)\right\|_{\max}\right)\left(\sup_{i,t}\{\mathbb{E}[g^{\eta}(w_{it})]\}^{2/\eta}\right)$
+×$\quad\quad\times\sup\left([a_{i}(0)]^{1-2/\eta}+2\sum_{m=1}^{T-2}\left[\left(1-\frac{m}{T-1}\right)[a_{i}(m)]^{1-2/\eta}\right]+\frac{1}{T-1}\sum_{m=1}^{T-1}[a_{i}(m)]^{1-2/\eta}\right)$
+$=O(\sqrt{NT}).$
+Similarly,
+$\left|A_{NT,2}(\theta)\right|$ $=\frac{T}{\sqrt{NT}}\sum_{i=1}^{N-1}\left(\frac{1}{T}\sum_{t=1}^{T-1}\left|h_{i,t}^{\ast\alpha\gamma}(\theta)\right|\right)\left|\mathbb{E}\left(\sum_{s=1}^{T}\widetilde{s}_{i,s}^{\alpha}(\theta)\widetilde{s}_{i,T}^{\gamma}(\theta)\right)\right|$
+$\leq\sqrt{NT}\left(\sup_{\theta\in\Theta}\left\|H_{\alpha\gamma^{\prime}}^{\ast}(\theta)\right\|_{\max}\right)\left(\sup_{i,t}\{\mathbb{E}[g^{\eta}(w_{it})]\}^{2/\eta}\right)\left(\sup_{i}\sum_{m=0}^{T-1}[a_{i}(m)]^{1-2/\eta}\right)$
+$=O(\sqrt{NT});$
+$\left|A_{NT,3}(\theta)\right|$ $=\frac{N}{\sqrt{NT}}\left|\mathbb{E}\left[\left(\sum_{s=1}^{T}\widetilde{s}_{N,s}^{\alpha}(\theta)\right)\left(\sum_{t=1}^{T-1}\widetilde{s}_{N,t}^{\gamma}(\theta)\left(\frac{1}{N}\sum_{i=1}^{N-1}h_{i,t}^{\ast\alpha\gamma}(\theta)\right)\right)\right]\right|$
+$=\sqrt{NT}\left(\sup_{\theta\in\Theta}\left\|H_{\alpha\gamma^{\prime}}^{\ast}(\theta)\right\|_{\max}\right)\left(\sup_{i,t}\{\mathbb{E}[g^{\eta}(w_{it})]\}^{2/\eta}\right)$
+×$\quad\quad\times\sup_{i}\left([a_{i}(0)]^{1-2/\eta}+2\sum_{m=1}^{T-2}\left[\left(1-\frac{m}{T-1}\right)[a_{i}(m)]^{1-2/\eta}\right]+\frac{1}{T-1}\sum_{m=1}^{T-1}[a_{i}(m)]^{1-2/\eta}\right)$
+$=O(\sqrt{NT});$
+$\left|A_{NT,4}(\theta)\right|$ $\leq\sqrt{NT}\left(\sup_{\theta\in\Theta}\left\|H_{\alpha\gamma^{\prime}}^{\ast}(\theta)\right\|_{\max}\right)\left(\sup_{i,t}\{\mathbb{E}[g^{\eta}(w_{it})]\}^{2/\eta}\right)\sup_{i}\left(\sum_{m=0}^{T-1}[a_{i}(m)]^{1-2/\eta}\right)$
+$=O(\sqrt{NT}).$
+Therefore, (A.2) follows from (A.4).
+Let $h_{i,j}^{\ast\alpha}(\theta)$ denote the $(i,j)$-th element of $T^{-1}H_{\alpha\alpha^{\prime}}^{\ast}(\theta)$, $i,j=1,\ldots,N-1$.
+By Assumption 1, we know that $\sup_{i\geq 1}h_{i,i}^{\ast\alpha}(\theta)=O(1)$ uniformly in $\theta$.
+Recall, $\alpha:=(\alpha_{1}^{\prime},\ldots,\alpha_{N}^{\prime})^{\prime}$ and $\gamma:=(\gamma_{1}^{\prime},\ldots,\gamma_{T}^{\prime})^{\prime}$.
+Note that
+$N\cdot H_{\alpha\alpha^{\prime}}(\theta)=D_{1}[N\cdot\partial_{\alpha\alpha^{\prime}}l(\theta,\phi)|_{\phi=\phi(\theta)}]D_{1}^{\prime},$
+implying $N\cdot H_{\alpha\alpha^{\prime}}(\theta)$ can be written as the sum of a diagonal matrix and a rank-one matrix.
+Thus, following from Miller (1981), we have $\sup_{i\neq j}h_{i,j}^{\ast\alpha}(\theta)=O(1/N)$ uniformly in $\theta$.
+Now,
+$\quad\quad T\mathbb{E}\left[\widetilde{s}_{\alpha}^{\prime}(\theta)H_{\alpha\alpha^{\prime}}^{\ast}(\theta)\widetilde{s}_{\alpha}(\theta)\right]$
+$=\frac{1}{N}\mathbb{E}\left\{(N\sqrt{T}\widetilde{s}_{\alpha}^{\prime}(\theta))[N^{-1}H_{\alpha\alpha^{\prime}}^{\ast}(\theta)](N\sqrt{T}\widetilde{s}_{\alpha}(\theta))\right\}$
+$=\frac{1}{N}\sum_{i=1}^{N-1}h_{i,i}^{\ast\alpha}(\theta)\mathbb{E}\left(\frac{1}{\sqrt{T}}\sum_{t=1}^{T}\left[\widetilde{s}_{i,t}^{\alpha}(\theta)-\widetilde{s}_{N,t}^{\alpha}(\theta)\right]\right)^{2}$
+$\quad\quad+\frac{2}{N}\sum_{1\leq i<j\leq N-1}h_{i,j}^{\ast\alpha}(\theta)\mathbb{E}\left[\left(\frac{1}{\sqrt{T}}\sum_{t=1}^{T}\left[\widetilde{s}_{i,t}^{\alpha}(\theta)-\widetilde{s}_{N,t}^{\alpha}(\theta)\right]\right)\left(\frac{1}{\sqrt{T}}\sum_{t=1}^{T}\left[\widetilde{s}_{j,t}^{\alpha}(\theta)-\widetilde{s}_{N,t}^{\alpha}(\theta)\right]\right)\right]$
+$=:I_{NT,1}(\theta)+I_{NT,2}(\theta).$ (A.5)
+Here, by Assumptions 1 and 1, applying Davydov’s inequality gives, for some $\eta>2$,
+$\quad\quad\sup_{\theta\in\Theta}\left|I_{NT,1}(\theta)\right|$
+$=\sup_{\theta\in\Theta}\frac{1}{NT}\sum_{i=1}^{N-1}\left|h_{i,i}^{\ast\alpha}(\theta)\right|\left|\sum_{t=1}^{T}\sum_{s=1}^{T}\mathbb{E}\left[\widetilde{s}_{i,t}^{\alpha}(\theta)\widetilde{s}_{i,s}^{\alpha}(\theta)\right]+\sum_{t=1}^{T}\sum_{s=1}^{T}\mathbb{E}\left[\widetilde{s}_{N,t}^{\alpha}(\theta)\widetilde{s}_{N,s}^{\alpha}(\theta)\right]\right|$
+$\leq\left(\frac{1}{N}\sum_{i=1}^{N-1}|h_{i,i}^{\ast\breve{\alpha}}(\theta)|\right)\left(\sup_{i,t}\{\mathbb{E}[g^{\eta}(w_{it})]\}^{2/\eta}\right)\sup_{i}\left([a_{i}(0)]^{1-2/\eta}+2\sum_{m=1}^{T-1}\left(1-\frac{m}{T}\right)[a_{i}(m)]^{1-2/\eta}\right)$
+$=O\left(1\right)$
+as $N,T\rightarrow\infty$, where the conditional cross-sectional independence leads to the first equality. Similarly
+$\quad\quad\sup_{\theta\in\Theta}\left|I_{NT,2}(\theta)\right|$
+$=\frac{2}{N}\sum_{1\leq i<j\leq N-1}\left\{\left|h_{i,j}^{\ast\alpha}(\theta)\right|\mathbb{E}\left[\left(\frac{1}{\sqrt{T}}\sum_{t=1}^{T}\widetilde{s}_{N,t}^{\alpha}(\theta)\right)^{2}\right]\right\}$
+$\leq\left(\frac{2}{N}\sum_{1\leq i<j\leq N-1}|h_{i,j}^{\ast\alpha}(\theta)|\right)\left(\sup_{i,t}\{\mathbb{E}[g^{\eta}(w_{it})]\}^{2/\eta}\right)\sup_{i}\left([a_{i}(0)]^{1-2/\eta}+2\sum_{m=1}^{T-1}\left(1-\frac{m}{T}\right)[a_{i}(m)]^{1-2/\eta}\right)$
+$=O\left(1\right),$
+noting that $\sup_{i\neq j}h_{i,j}^{\ast\alpha}(\theta)=O(1/N)$ uniformly in $\theta$. Thus, the first part of (A.3) follows from (A.5). The second part of (A.3) follows from a similar procedure.
+Combining (A.1), (A.2), and (A.3),
+$\mathbb{E}\left\{s^{\prime}(\theta)[-{H}(\theta)]^{-1}s(\theta)\right\}$ $=\mathbb{E}\left\{\widetilde{s}_{\alpha}^{\prime}(\theta)H_{\alpha\alpha^{\prime}}^{\ast}(\theta)\widetilde{s}_{\alpha}(\theta)\right\}+\mathbb{E}\left\{\widetilde{s}_{\gamma}^{\prime}(\theta)H_{\gamma\gamma^{\prime}}^{\ast}(\theta)\widetilde{s}_{\gamma}(\theta)\right\}$
+$\quad\quad+\mathbb{E}\left\{\left[\widetilde{s}_{\alpha}^{\prime}(\theta)H_{\alpha\alpha^{\prime}}^{\ast}(\theta)\widetilde{s}_{\alpha}(\theta)+\widetilde{s}_{\gamma}^{\prime}\left(\theta\right)H_{\gamma\gamma^{\prime}}^{\ast}\left(\theta\right)\widetilde{s}_{\gamma}(\theta)\right]\cdot o_{\mathbb{P}}(1)\right\}.$
+By Assumption 1, for $S\in\{0,1,2\}$, we have
+$\sup_{i,t}\mathbb{E}\left(|\partial_{\phi_{1}\ldots\phi_{S}}l_{i,t}\left(\theta,\phi\right)|\mathds{1}\{|\partial_{\phi_{1}\ldots\phi_{S}}l_{i,t}\left(\theta,\phi\right)|>a\}\right)\rightarrow 0$
 as $a\rightarrow\infty$, which shows that $\{\partial_{\phi_{1}\ldots\phi_{S}}l_{it}\left(\theta,\phi\right)\}_{i,t}$ are uniformly integrable. By the $L_{1}$ convergence theorem, we hence have
-$\sqrt{NT}\mathbb{E}\left\{\big{[}s_{\breve{\alpha}}^{\prime}(\theta)H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)s_{\breve{\alpha}}(\theta)+s_{\breve{\gamma}}^{\prime}\left(\theta\right)H_{\breve{\gamma}\breve{\gamma}^{\prime}}^{\ast}\left(\theta\right)s_{\breve{\gamma}}(\theta)\big{]}\cdot o_{\mathbb{P}}(1)\right\}\rightarrow 0,$
-as $N,T\rightarrow\infty.$ This completes the proof of Lemma 1.
+$\sqrt{NT}\mathbb{E}\left\{\left[\widetilde{s}_{\alpha}^{\prime}(\theta)H_{\alpha\alpha^{\prime}}^{\ast}(\theta)\widetilde{s}_{\alpha}(\theta)+\widetilde{s}_{\gamma}^{\prime}\left(\theta\right)H_{\gamma\gamma^{\prime}}^{\ast}\left(\theta\right)\widetilde{s}_{\gamma}(\theta)\right]\cdot o_{\mathbb{P}}(1)\right\}\rightarrow 0,$
+as $N,T\rightarrow\infty$. This completes the proof.
+The next lemma shows that the remainder term in Equation (2.13), $r^{s}\left(\theta\right):=r^{s}\left(\theta,a\right)$ is $O_{\mathbb{P}}\left(1/\left(NT\right)\right)$ uniformly, where $r^{s}\left(\theta,a\right)$ is defined as [^10] Feng et al.
+(2014) point out that the expansion in (2.13) needs to be done component-by-component.
+This leads to a different $a_{j}$ for each component of $r^{s}\left(\theta,a\right)$.
+$r^{s}\left(\theta,a\right)$ $:=\left(r_{\psi_{1}}^{s}\left(\theta,a_{1}\right),\ldots,r_{\psi_{(N+T-2)K}}^{s}(\theta,a_{(N+T-2)K})\right)^{\prime},$ (A.6)
+$r_{\psi_{j}}^{s}\left(\theta,a_{j}\right)$ $:=\frac{1}{2}\left(\widehat{\psi}\left(\theta\right)-\psi(\theta)\right)^{\prime}G_{\psi_{j}}(\theta,a_{j})\left(\widehat{\psi}(\theta)-\psi(\theta)\right),$
+$G_{\psi_{j}}(\theta,a_{j})$ $:=\partial_{\psi_{j}}H(\theta,a_{j}\widehat{\psi}(\theta)+(1-a_{j})\psi\left(\theta\right))$
+with $a:=(a_{1},\ldots,a_{(N+T-2)K})^{\prime}$ for each $a_{j}\in(0,1)$, and $\psi_{j}$ denotes the $j$-th component of $\psi$.
+Here $G_{\psi_{j}}(\theta,a_{j})$ is the derivative of each element of $H(\theta,\psi)$ with respect to $\psi_{j}$, evaluated at the point $a_{j}\widehat{\psi}\left(\theta\right)+\left(1-a_{j}\right)\psi\left(\theta\right)$.
 Lemma A.2
-Under Assumption 1, $r^{s}(\theta,a)$ defined by (2.3) satisfies
+Under Assumption 1,
 $\sup_{\theta\in\Theta}\sup_{a}\left\|r^{s}(\theta,a)\right\|_{\max}=O_{\mathbb{P}}\left(1/\left(NT\right)\right),$
-as $N$ and $T\rightarrow\infty$, where $a=\left(a_{1},\ldots,a_{\left(N+T-2\right)K}\right)^{\prime}$ for each component $a_{j}\in(0,1)$.
+as $N$ and $T\rightarrow\infty$, where $\sup_{a}$ is taken over $a_{j}\in(0,1)$ for all $j=1,\ldots,(N+T-2)K$.
 Proof.
-In the proof, we retain the notation and setup from Lemma 1.
-From equations (2.3) and (2.5), it is evident that the dominant term in the expansion of $\widehat{\psi}(\theta)-\psi\left(\theta\right)$ is $[-{H}(\theta)]^{-1}s(\theta).$ Specifically,
+In the proof, we retain the notation and setup from Lemma A.1.
+From (2.13), it is evident that the dominant term in the expansion of $\widehat{\psi}(\theta)-\psi\left(\theta\right)$ is $[-{H}(\theta)]^{-1}s(\theta)$.
+Specifically,
 $[-{H}(\theta)]^{-1}s(\theta)$ $=[-\mathbb{E}{H}(\theta)]^{-1}[{s}(\theta)-\mathbb{E}{s}(\theta)](1+o_{\mathbb{P}}(1))$
-$=-\begin{pmatrix}H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\alpha}}(\theta)+H_{\breve{\alpha}\breve{\gamma}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\gamma}}(\theta)\\ H_{\breve{\gamma}\breve{\alpha}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\alpha}}(\theta)+H_{\breve{\gamma}\breve{\gamma}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\gamma}}(\theta)\end{pmatrix}+[-\mathbb{E}{H}(\theta)]^{-1}[{s}(\theta)-\mathbb{E}{s}(\theta)]\cdot o_{\mathbb{P}}(1)$ (A.6)
-where $o_{\mathbb{P}}(1)$ is uniform in both $i$ and $t.$ Let $u_{i}^{\breve{\alpha}}(\theta)$ denote the $i$-th element of $H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\alpha}}(\theta)$ for $i=1,\ldots,N-1.$ It follows that
-$N\sqrt{T}u_{i}^{\breve{\alpha}}(\theta)=N\sum_{j=1}^{N-1}h_{i,j}^{\ast\breve{\alpha}}(\theta)\bigg{(}\frac{1}{\sqrt{T}}\sum_{t=1}^{T}\big{[}\partial_{\alpha_{j}}\widetilde{l}_{j,t}(\theta,\phi(\theta))-\partial_{\alpha_{N}}\widetilde{l}_{N,t}(\theta,\phi(\theta))\big{]}\bigg{)}.$ (A.7)
-Using the result from Miller (1981), we know that $\sup_{i\geq 1}h_{i,i}^{\ast\breve{\alpha}}(\theta)=O(1)$ and $\sup_{i\neq j}h_{i,j}^{\ast\breve{\alpha}}(\theta)=O(1/N)$ uniformly in $\theta,$ and it also implies that $\sum_{j=1}^{N-1}h_{i,j}^{\ast\breve{\alpha}}(\theta)=O(1/N)$ uniformly in both $i$ and $\theta.$ Since $\frac{1}{\sqrt{T}}\sum_{t=1}^{T}\partial_{\alpha_{i}}\widetilde{l}_{i,t}(\theta,\phi(\theta))=O_{\mathbb{P}}(1)$ uniformly over $i$ by Assumptions 1 and 1, we obtain
-$N\sum_{j=1}^{N-1}h_{i,j}^{\ast\breve{\alpha}}(\theta)\bigg{(}\frac{1}{\sqrt{T}}\sum_{t=1}^{T}\partial_{\alpha_{j}}\widetilde{l}_{j,t}(\theta,\phi(\theta))\bigg{)}=O_{\mathbb{P}}(1)$
-and
-$N\sum_{j=1}^{N-1}h_{i,j}^{\ast\breve{\alpha}}(\theta)\cdot\bigg{(}\frac{1}{\sqrt{T}}\sum_{t=1}^{T}\partial_{\alpha_{N}}\widetilde{l}_{N,t}(\theta,\phi(\theta))\bigg{)}=O_{\mathbb{P}}(1).$
-Therefore, $N\sqrt{T}u_{i}^{\breve{\alpha}}(\theta)=O_{\mathbb{P}}(1)$ uniformly for $i.$ Let $u_{i}^{\breve{\alpha}\breve{\gamma}}(\theta)$ denote the $i$-th element of $H_{\breve{\alpha}\breve{\gamma}^{\prime}}^{\ast}\left(\theta\right)\widetilde{s}_{\breve{\gamma}}(\theta)$ for $i=1,\ldots,N-1.$ Given that $\sup_{\theta\in\Theta}\|H_{\breve{\alpha}\breve{\gamma}^{\prime}}^{\ast}(\theta)\|_{\max}=O(1),$ we have
-$\sqrt{N}Tu_{i}^{\breve{\alpha}\breve{\gamma}}(\theta)=\sum_{t=1}^{T-1}h_{i,t}^{\ast\breve{\alpha}\breve{\gamma}}(\theta)\bigg{(}\frac{1}{\sqrt{N}}\sum_{j=1}^{N}\big{[}\partial_{\gamma_{t}}\widetilde{l}_{j,t}(\theta,\phi(\theta))-\partial_{\gamma_{T}}\widetilde{l}_{j,T}(\theta,\phi(\theta))\big{]}\bigg{)}.$ (A.8)
-Since $\sum_{t=1}^{T-1}h_{i,t}^{\ast\breve{\alpha}\breve{\gamma}}(\theta)=O(1),$ which follows from $\mathbb{E}[\partial_{\breve{\alpha}\breve{\gamma}^{\prime}}l^{*}(\theta,\psi)|_{\psi=\psi(\theta)}]=\partial_{\breve{\gamma}^{\prime}}\mathbb{E}[\partial_{\breve{\alpha}}l^{*}(\theta,\psi)|_{\psi=\psi(\theta)}]=0,$ it holds that
-$\sum_{t=1}^{T-1}h_{i,t}^{\ast\breve{\alpha}\breve{\gamma}}(\theta)\bigg{(}\frac{1}{\sqrt{N}}\sum_{j=1}^{N}\partial_{\gamma_{t}}\widetilde{l}_{j,t}(\theta,\phi(\theta))\bigg{)}=O_{\mathbb{P}}(1),$
-and
-$\bigg{(}\sum_{t=1}^{T-1}h_{i,t}^{\ast\breve{\alpha}\breve{\gamma}}(\theta)\bigg{)}\bigg{(}\frac{1}{\sqrt{N}}\sum_{j=1}^{N}\partial_{\gamma_{T}}\widetilde{l}_{j,T}(\theta,\phi(\theta))\bigg{)}=O_{\mathbb{P}}(1),$
+$=-\begin{pmatrix}H_{\alpha\alpha^{\prime}}^{\ast}(\theta)\widetilde{s}_{\alpha}(\theta)+H_{\alpha\gamma^{\prime}}^{\ast}(\theta)\widetilde{s}_{\gamma}(\theta)\\ H_{\gamma\alpha^{\prime}}^{\ast}(\theta)\widetilde{s}_{\alpha}(\theta)+H_{\gamma\gamma^{\prime}}^{\ast}(\theta)\widetilde{s}_{\gamma}(\theta)\end{pmatrix}+[-\mathbb{E}{H}(\theta)]^{-1}[{s}(\theta)-\mathbb{E}{s}(\theta)]\cdot o_{\mathbb{P}}(1)$ (A.7)
+where $o_{\mathbb{P}}(1)$ is uniform in both $i$ and $t$.
+Let $u_{i}^{\alpha}(\theta)$ denote the $i$-th element of $H_{\alpha\alpha^{\prime}}^{\ast}(\theta)\widetilde{s}_{\alpha}(\theta)$.
+It follows that
+$N\sqrt{T}u_{i}^{\alpha}(\theta)=N\sum_{j=1}^{N-1}\left[h_{i,j}^{\ast\alpha}(\theta)\left(\frac{1}{\sqrt{T}}\sum_{t=1}^{T}\left[\widetilde{s}_{j,t}^{\alpha}(\theta)-\widetilde{s}_{N,t}^{\alpha}(\theta)\right]\right)\right].$ (A.8)
+Using the result from Miller (1981), we know that $\sup_{i\geq 1}h_{i,i}^{\ast\alpha}(\theta)=O(1)$ and $\sup_{i\neq j}h_{i,j}^{\ast\alpha}(\theta)=O(1/N)$ uniformly in $\theta$, implying $\sum_{j=1}^{N-1}h_{i,j}^{\ast\alpha}(\theta)=O(1/N)$ uniformly in both $i$ and $\theta$.
+Since $1/\sqrt{T}\sum_{t=1}^{T}\widetilde{s}_{i,t}^{\alpha}(\theta)=O_{\mathbb{P}}(1)$ uniformly over $i$ by Assumptions 1 and 1, we obtain
+$N\sum_{j=1}^{N-1}\left[h_{i,j}^{\ast\alpha}(\theta)\left(\frac{1}{\sqrt{T}}\sum_{t=1}^{T}\widetilde{s}_{j,t}^{\alpha}(\theta)\right)\right]=O_{\mathbb{P}}(1),\qquad N\left(\sum_{j=1}^{N-1}h_{i,j}^{\ast\breve{\alpha}}(\theta)\right)\left(\frac{1}{\sqrt{T}}\sum_{t=1}^{T}\widetilde{s}_{N,t}^{\alpha}(\theta)\right)=O_{\mathbb{P}}(1).$
+Therefore, $N\sqrt{T}u_{i}^{\alpha}(\theta)=O_{\mathbb{P}}(1)$ uniformly over $i$.
+Let $u_{i}^{\alpha\gamma}(\theta)$ denote the $i$-th element of $H_{\alpha\gamma^{\prime}}^{\ast}\left(\theta\right)\widetilde{s}_{\gamma}(\theta)$.
+Given that $\sup_{\theta\in\Theta}\|H_{\alpha\gamma^{\prime}}^{\ast}(\theta)\|_{\max}=O(1)$, we have
+$T\sqrt{N}u_{i}^{\alpha\gamma}(\theta)=\sum_{t=1}^{T-1}\left[h_{i,t}^{\ast\alpha\gamma}(\theta)\left(\frac{1}{\sqrt{N}}\sum_{j=1}^{N}\left[\widetilde{s}_{j,t}^{\gamma}(\theta)-\widetilde{s}_{j,T}^{\gamma}(\theta)\right]\right)\right].$ (A.9)
+Since $\sum_{t=1}^{T-1}h_{i,t}^{\ast\alpha\gamma}(\theta)=O(1)$, which follows from $\mathbb{E}[\partial_{\breve{\alpha}\breve{\gamma}^{\prime}}\ell(\theta,\psi)|_{\psi=\psi(\theta)}]=\partial_{\breve{\gamma}^{\prime}}\mathbb{E}[\partial_{\breve{\alpha}}\ell(\theta,\psi)|_{\psi=\psi(\theta)}]=0$, for $\breve{\alpha}:=\left(\alpha_{1}^{\prime},\ldots,\alpha_{N-1}^{\prime}\right)^{\prime}$ and $\breve{\gamma}:=\left(\gamma_{1}^{\prime},\ldots,\gamma_{T-1}^{\prime}\right)^{\prime}$, it holds that
+$\sum_{t=1}^{T-1}\left[h_{i,t}^{\ast\alpha\gamma}(\theta)\left(\frac{1}{\sqrt{N}}\sum_{j=1}^{N}\widetilde{s}_{j,t}^{\gamma}(\theta)\right)\right]=O_{\mathbb{P}}(1),\qquad\left(\sum_{t=1}^{T-1}h_{i,t}^{\ast\alpha\gamma}(\theta)\right)\left(\frac{1}{\sqrt{N}}\sum_{j=1}^{N}\widetilde{s}_{j,T}^{\gamma}(\theta)\right)=O_{\mathbb{P}}(1)$
 uniformly in $\theta$. From (A.8) and (A.9), we conclude that
-$\sup_{\theta\in\Theta}\|H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\alpha}}(\theta)\|_{\max}=O_{\mathbb{P}}\big{(}1/(N\sqrt{T})\big{)},\quad\sup_{\theta\in\Theta}\|H_{\breve{\alpha}\breve{\gamma}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\gamma}}(\theta)\|_{\max}=O_{\mathbb{P}}\big{(}1/(\sqrt{N}T)\big{)}.$ (A.9)
-Similarly, we can show that
-$\sup_{\theta\in\Theta}\|H_{\breve{\gamma}\breve{\gamma}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\gamma}}(\theta)\|_{\max}=O_{\mathbb{P}}\big{(}1/(\sqrt{N}T)\big{)},\quad\sup_{\theta\in\Theta}\|H_{\breve{\gamma}\breve{\alpha}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\alpha}}(\theta)\|_{\max}=O_{\mathbb{P}}\big{(}1/(N\sqrt{T})\big{)}.$ (A.10)
+$\sup_{\theta\in\Theta}\|H_{\alpha\alpha^{\prime}}^{\ast}(\theta)\widetilde{s}_{\alpha}(\theta)\|_{\max}=O_{\mathbb{P}}\left(1/(N\sqrt{T})\right),\qquad\sup_{\theta\in\Theta}\|H_{\alpha\gamma^{\prime}}^{\ast}(\theta)\widetilde{s}_{\gamma}(\theta)\|_{\max}=O_{\mathbb{P}}\left(1/(T\sqrt{N})\right).$ (A.10)
+Following a similar proof, we also have
+$\sup_{\theta\in\Theta}\|H_{\gamma\gamma^{\prime}}^{\ast}(\theta)\widetilde{s}_{\gamma}(\theta)\|_{\max}=O_{\mathbb{P}}\left(1/(T\sqrt{N})\right),\qquad\sup_{\theta\in\Theta}\|H_{\gamma\alpha^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\alpha}}(\theta)\|_{\max}=O_{\mathbb{P}}\left(1/(N\sqrt{T})\right).$ (A.11)
 Thus, combining (A.7), (A.10), and (A.11), we have
-$\left\|[-{H}(\theta)]^{-1}s(\theta)+[\mathbb{E}{H}(\theta)]^{-1}[{s}(\theta)-\mathbb{E}{s}(\theta)]\right\|_{\max}=o_{\mathbb{P}}\big{(}1/(NT)^{3/4}\big{)},$
+$\left\|[-{H}(\theta)]^{-1}s(\theta)+[\mathbb{E}{H}(\theta)]^{-1}[{s}(\theta)-\mathbb{E}{s}(\theta)]\right\|_{\max}=o_{\mathbb{P}}\left(1/(NT)^{3/4}\right),$
 where $o_{\mathbb{P}}(\cdot)$ is uniform in $\theta$. Noting that
@@
 we further have
-$\left\|\widehat{\psi}(\theta)-\psi(\theta)+[\mathbb{E}{H}(\theta)]^{-1}[{s}(\theta)-\mathbb{E}{s}(\theta)]\right\|_{\max}=o_{\mathbb{P}}\big{(}1/(NT)^{3/4}\big{)}.$ (A.11)
-From (A.11), we conclude that $\sup_{\theta\in\Theta}\|G_{\psi_{j}}\left(\theta,a_{j}\right)-\partial_{\psi_{j}}H(\theta)\|_{\max}=o_{\mathbb{P}}(1).$ Without loss of generality, let $\psi_{j}=\alpha_{i}$ for $i=1,\ldots,N-1.$ We can express $\partial_{\alpha_{i}}H(\theta)$ as
-$\partial_{\alpha_{i}}H(\theta)=\begin{pmatrix}\partial_{\alpha_{i}}{H}_{\breve{\alpha}\breve{\alpha}^{\prime}}\left(\theta\right),&\partial_{\alpha_{i}}{H}_{\breve{\alpha}\breve{\gamma}^{\prime}}(\theta)\\ \partial_{\alpha_{i}}{H}_{\breve{\gamma}\breve{\alpha}^{\prime}}\left(\theta\right),&\partial_{\alpha_{i}}{H}_{\breve{\gamma}\breve{\gamma}^{\prime}}(\theta)\end{pmatrix}.$
-In particular, let $\text{diag}\{a_{1},\ldots,a_{n}\}$ denote a diagonal matrix with elements $a_{1},\ldots,a_{n}$ on its diagonal. We then have
-$N\cdot\partial_{\alpha_{i}}{H}_{\breve{\alpha}\breve{\alpha}^{\prime}}\left(\theta\right)=$ $\frac{1}{T}\sum_{t=1}^{T}\text{diag}\big{\{}0,\ldots,0,\partial_{\alpha_{i}\alpha_{i}\alpha_{i}}l_{i,t}(\theta,\phi)|_{\phi=\phi(\theta)},0,\ldots,0\big{\}}$
-$\quad-\bigg{(}\frac{1}{T}\sum_{t=1}^{T}\partial_{\alpha_{N}\alpha_{N}\alpha_{N}}l_{N,t}(\theta,\phi)|_{\phi=\phi(\theta)}\bigg{)}\,\iota_{N-1}\iota_{N-1}^{\prime},$
-$NT\cdot\partial_{\alpha_{i}}{H}_{\breve{\gamma}\breve{\gamma}^{\prime}}\left(\theta\right)=$ $\text{diag}\left\{\partial_{\gamma_{1}\gamma_{1}\alpha_{i}}l_{i,1}(\theta,\phi)|_{\phi=\phi(\theta)},\ldots,\partial_{\gamma_{T-1}\gamma_{T-1}\alpha_{i}}l_{i,T-1}(\theta,\phi)|_{\phi=\phi(\theta)}\right\}$
-$\quad+\partial_{\gamma_{T}\gamma_{T}\alpha_{i}}l_{i,T}(\theta,\phi)|_{\phi=\phi(\theta)}\,\iota_{N-1}\iota_{N-1}^{\prime},$
-$NT\cdot\partial_{\alpha_{i}}{H}_{\breve{\alpha}\breve{\gamma}^{\prime}}\left(\theta\right)=$ $\big{(}\partial_{\alpha_{i}\alpha_{i}\gamma_{1}}l_{i,1}(\theta,\phi)|_{\phi=\phi(\theta)},\ldots,\partial_{\alpha_{i}\alpha_{i}\gamma_{T-1}}l_{i,T-1}(\theta,\phi)|_{\phi=\phi(\theta)}\big{)}\otimes e_{i}$
-$\quad+\big{(}\partial_{\alpha_{N}\alpha_{N}\gamma_{1}}l_{N,1}(\theta,\phi)|_{\phi=\phi(\theta)},\ldots,\partial_{\alpha_{N}\alpha_{N}\gamma_{T-1}}l_{N,T-1}(\theta,\phi)|_{\phi=\phi(\theta)}\big{)}\otimes\iota_{N-1},$
-where $e_{i}$ is an $(N-1)\times 1$ vector with the $i$-th component equal to 1 and all other elements equal to 0. Using (A.11), for each $j=1,\ldots,N-1,$ we write
-$\sup_{a_{j}\in(0,1)}r_{\psi_{j}}^{s}(\theta,a_{j})$
-$=$ $\sup_{a_{j}\in(0,1)}\frac{1}{2}\big{(}\widehat{\psi}(\theta)-\psi(\theta)\big{)}^{\prime}G_{\psi_{j}}(\theta,a_{j})\big{(}\widehat{\psi}(\theta)-\psi\left(\theta\right)\big{)}$
-$=$ $\frac{1}{2}\begin{pmatrix}H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\alpha}}(\theta)+H_{\breve{\alpha}\breve{\gamma}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\gamma}}(\theta)\\ H_{\breve{\gamma}\breve{\alpha}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\alpha}}(\theta)+H_{\breve{\gamma}\breve{\gamma}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\gamma}}(\theta)\end{pmatrix}^{\prime}\partial_{\alpha_{i}}H(\theta)\begin{pmatrix}H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\alpha}}(\theta)+H_{\breve{\alpha}\breve{\gamma}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\gamma}}(\theta)\\ H_{\breve{\gamma}\breve{\alpha}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\alpha}}(\theta)+H_{\breve{\gamma}\breve{\gamma}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\gamma}}(\theta)\end{pmatrix}\big{(}1+o_{\mathbb{P}}(1)\big{)}$
-$=$ $\frac{1}{2}\left\{\big{(}H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\alpha}}(\theta)+H_{\breve{\alpha}\breve{\gamma}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\gamma}}(\theta)\big{)}^{\prime}\partial_{\alpha_{i}}{H}_{\breve{\alpha}\breve{\alpha}^{\prime}}(\theta)\big{(}H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\alpha}}(\theta)+H_{\breve{\alpha}\breve{\gamma}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\gamma}}(\theta)\big{)}\right.$
-$\qquad+\big{(}H_{\breve{\gamma}\breve{\alpha}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\alpha}}(\theta)+H_{\breve{\gamma}\breve{\gamma}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\gamma}}(\theta)\big{)}^{\prime}\partial_{\alpha_{i}}{H}_{\breve{\gamma}\breve{\gamma}^{\prime}}(\theta)\big{(}H_{\breve{\gamma}\breve{\alpha}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\alpha}}(\theta)+H_{\breve{\gamma}\breve{\gamma}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\gamma}}(\theta)\big{)}$
-$\qquad\left.+2\big{(}H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\alpha}}(\theta)+H_{\breve{\alpha}\breve{\gamma}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\gamma}}(\theta)\big{)}^{\prime}\partial_{\alpha_{i}}{H}_{\breve{\alpha}\breve{\gamma}^{\prime}}(\theta)\big{(}H_{\breve{\gamma}\breve{\alpha}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\alpha}}(\theta)+H_{\breve{\gamma}\breve{\gamma}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\gamma}}(\theta)\big{)}\right\}\big{(}1+o_{\mathbb{P}}(1)\big{)}$
-=:absent:=:= : $\frac{1}{2}\left\{C_{NT,1}(\theta)+C_{NT,2}(\theta)+C_{NT,3}(\theta)\right\}\big{(}1+o_{\mathbb{P}}(1)\big{)}.$ (A.12)
-Let $\widetilde{u}_{i}^{\breve{\alpha}}(\theta)$ denote the $i$-th element of $H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\alpha}}(\theta)+H_{\breve{\alpha}\breve{\gamma}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\gamma}}(\theta)$ for $i=1,\ldots,N-1.$ Using (A.7) and (A.8), we can show that
-$C_{NT,1}(\theta)$
-$=$ $\frac{1}{N}[\widetilde{u}_{i}^{\breve{\alpha}}(\theta)]^{2}\bigg{(}\frac{1}{T}\sum_{t=1}^{T}\partial_{\alpha_{i}\alpha_{i}\alpha_{i}}l_{i,t}(\theta,\phi)|_{\phi=\phi(\theta)}\bigg{)}-\frac{1}{N}\bigg{(}\sum_{i=1}^{N-1}\widetilde{u}_{i}^{\breve{\alpha}}(\theta)\bigg{)}^{2}\bigg{(}\frac{1}{T}\sum_{t=1}^{T}\partial_{\alpha_{N}\alpha_{N}\alpha_{N}}l_{N,t}(\theta,\phi)|_{\phi=\phi(\theta)}\bigg{)}$
-$=$ $O_{\mathbb{P}}\big{(}1/(NT)\big{)},$
-uniformly in $\theta.$ By similar arguments, we also show that
-$\sup_{\theta\in\Theta}C_{NT,2}(\theta)=O_{\mathbb{P}}\big{(}1/(NT)^{3/2}\big{)},\quad\sup_{\theta\in\Theta}C_{NT,3}(\theta)=O_{\mathbb{P}}\big{(}1/(NT)^{3/2}\big{)}.$
+$\left\|\widehat{\psi}(\theta)-\psi(\theta)+[\mathbb{E}{H}(\theta)]^{-1}[{s}(\theta)-\mathbb{E}{s}(\theta)]\right\|_{\max}=o_{\mathbb{P}}\left(1/(NT)^{3/4}\right).$ (A.12)
+From (A.12), we conclude that $\sup_{\theta\in\Theta}\|G_{\psi_{j}}\left(\theta,a_{j}\right)-\partial_{\psi_{j}}H(\theta)\|_{\max}=o_{\mathbb{P}}(1)$.
+Without loss of generality, let $\psi_{j}=\alpha_{i}$ for $i=1,\ldots,N-1$.
+We can express $\partial_{\alpha_{i}}H(\theta)$ as
+$\partial_{\alpha_{i}}H(\theta)=\begin{pmatrix}\partial_{\alpha_{i}}{H}_{\alpha\alpha^{\prime}}\left(\theta\right),&\partial_{\alpha_{i}}{H}_{\alpha\gamma^{\prime}}(\theta)\\ \partial_{\alpha_{i}}{H}_{\gamma\alpha^{\prime}}\left(\theta\right),&\partial_{\alpha_{i}}{H}_{\gamma\gamma^{\prime}}(\theta)\end{pmatrix}.$
+We then have
+$N\cdot\partial_{\alpha_{i}}{H}_{\breve{\alpha}\breve{\alpha}^{\prime}}\left(\theta\right)$ $=\frac{1}{T}\sum_{t=1}^{T}\operatorname*{diag}\left(0,\ldots,0,\partial_{\alpha_{i}\alpha_{i}\alpha_{i}}l_{i,t}(\theta,\phi)|_{\phi=\phi(\theta)},0,\ldots,0\right)$
+$\quad\quad-\left(\frac{1}{T}\sum_{t=1}^{T}\partial_{\alpha_{N}\alpha_{N}\alpha_{N}}l_{N,t}(\theta,\phi)|_{\phi=\phi(\theta)}\right)\iota_{N-1}\iota_{N-1}^{\prime},$
+$NT\cdot\partial_{\alpha_{i}}{H}_{\breve{\gamma}\breve{\gamma}^{\prime}}\left(\theta\right)$ $=\operatorname*{diag}\left\{\partial_{\gamma_{1}\gamma_{1}\alpha_{i}}l_{i,1}(\theta,\phi)|_{\phi=\phi(\theta)},\ldots,\partial_{\gamma_{T-1}\gamma_{T-1}\alpha_{i}}l_{i,T-1}(\theta,\phi)|_{\phi=\phi(\theta)}\right\}$
+$\quad\quad+\partial_{\gamma_{T}\gamma_{T}\alpha_{i}}l_{i,T}(\theta,\phi)|_{\phi=\phi(\theta)}\,\iota_{N-1}\iota_{N-1}^{\prime},$
+$NT\cdot\partial_{\alpha_{i}}{H}_{\breve{\alpha}\breve{\gamma}^{\prime}}\left(\theta\right)$ $=\left(\partial_{\alpha_{i}\alpha_{i}\gamma_{1}}l_{i,1}(\theta,\phi)|_{\phi=\phi(\theta)},\ldots,\partial_{\alpha_{i}\alpha_{i}\gamma_{T-1}}l_{i,T-1}(\theta,\phi)|_{\phi=\phi(\theta)}\right)\otimes e_{i}$
+$\quad\quad+\left(\partial_{\alpha_{N}\alpha_{N}\gamma_{1}}l_{N,1}(\theta,\phi)|_{\phi=\phi(\theta)},\ldots,\partial_{\alpha_{N}\alpha_{N}\gamma_{T-1}}l_{N,T-1}(\theta,\phi)|_{\phi=\phi(\theta)}\right)\otimes\iota_{N-1},$
+where $e_{i}$ is an $(N-1)\times 1$ vector with the $i$-th component equal to $1$ and all other elements equal to 00. Using (A.12), for each $j=1,\ldots,N-1$, we write
+$\quad\quad\sup_{a_{j}\in(0,1)}r_{\psi_{j}}^{s}(\theta,a_{j})$
+$=\sup_{a_{j}\in(0,1)}\frac{1}{2}\left(\widehat{\psi}(\theta)-\psi(\theta)\right)^{\prime}G_{\psi_{j}}(\theta,a_{j})\left(\widehat{\psi}(\theta)-\psi\left(\theta\right)\right)$
+$=\frac{1}{2}\begin{pmatrix}H_{\alpha\alpha^{\prime}}^{\ast}(\theta)\widetilde{s}_{\alpha}(\theta)+H_{\alpha\gamma^{\prime}}^{\ast}(\theta)\widetilde{s}_{\gamma}(\theta)\\ H_{\gamma\alpha^{\prime}}^{\ast}(\theta)\widetilde{s}_{\alpha}(\theta)+H_{\gamma\gamma^{\prime}}^{\ast}(\theta)\widetilde{s}_{\gamma}(\theta)\end{pmatrix}^{\prime}\partial_{\alpha_{i}}H(\theta)\begin{pmatrix}H_{\alpha\alpha^{\prime}}^{\ast}(\theta)\widetilde{s}_{\alpha}(\theta)+H_{\alpha\gamma^{\prime}}^{\ast}(\theta)\widetilde{s}_{\gamma}(\theta)\\ H_{\gamma\alpha^{\prime}}^{\ast}(\theta)\widetilde{s}_{\alpha}(\theta)+H_{\gamma\gamma^{\prime}}^{\ast}(\theta)\widetilde{s}_{\gamma}(\theta)\end{pmatrix}\left[1+o_{\mathbb{P}}(1)\right]$
+$=\frac{1}{2}\left\{\left(H_{\alpha\alpha^{\prime}}^{\ast}(\theta)\widetilde{s}_{\alpha}(\theta)+H_{\alpha\gamma^{\prime}}^{\ast}(\theta)\widetilde{s}_{\gamma}(\theta)\right)^{\prime}\partial_{\alpha_{i}}{H}_{\alpha\alpha^{\prime}}(\theta)\left(H_{\alpha\alpha^{\prime}}^{\ast}(\theta)\widetilde{s}_{\alpha}(\theta)+H_{\alpha\gamma^{\prime}}^{\ast}(\theta)\widetilde{s}_{\gamma}(\theta)\right)\right.$
+$\quad\quad+\left(H_{\gamma\alpha^{\prime}}^{\ast}(\theta)\widetilde{s}_{\alpha}(\theta)+H_{\gamma\gamma^{\prime}}^{\ast}(\theta)\widetilde{s}_{\gamma}(\theta)\right)^{\prime}\partial_{\alpha_{i}}{H}_{\gamma\gamma^{\prime}}(\theta)\left(H_{\gamma\alpha^{\prime}}^{\ast}(\theta)\widetilde{s}_{\alpha}(\theta)+H_{\gamma\gamma^{\prime}}^{\ast}(\theta)\widetilde{s}_{\gamma}(\theta)\right)$
+$\quad\quad\left.+2\left(H_{\alpha\alpha^{\prime}}^{\ast}(\theta)\widetilde{s}_{\alpha}(\theta)+H_{\alpha\gamma^{\prime}}^{\ast}(\theta)\widetilde{s}_{\gamma}(\theta)\right)^{\prime}\partial_{\alpha_{i}}{H}_{\breve{\alpha}\breve{\gamma}^{\prime}}(\theta)\left(H_{\gamma\alpha^{\prime}}^{\ast}(\theta)\widetilde{s}_{\alpha}(\theta)+H_{\gamma\gamma^{\prime}}^{\ast}(\theta)\widetilde{s}_{\gamma}(\theta)\right)\right\}\left[1+o_{\mathbb{P}}(1)\right]$
+$=:\frac{1}{2}\left\{C_{NT,1}(\theta)+C_{NT,2}(\theta)+C_{NT,3}(\theta)\right\}\left[1+o_{\mathbb{P}}(1)\right].$ (A.13)
+Let $v_{i}^{\alpha}(\theta)$ denote the $i$-th element of $H_{\alpha\alpha^{\prime}}^{\ast}(\theta)\widetilde{s}_{\alpha}(\theta)+H_{\alpha\gamma^{\prime}}^{\ast}(\theta)\widetilde{s}_{\gamma}(\theta)$, for $i=1,\ldots,N-1$.
+Using (A.8) and (A.9), it can be shown that
+$\quad\quad C_{NT,1}(\theta)$
+$=\frac{1}{N}\left(\sum_{i=1}^{N-1}v_{i}^{\alpha}(\theta)\right)^{2}\left(\frac{1}{T}\sum_{t=1}^{T}\partial_{\alpha_{i}\alpha_{i}\alpha_{i}}l_{i,t}(\theta,\phi)|_{\phi=\phi(\theta)}\right)-\frac{1}{N}\left(\sum_{i=1}^{N-1}v_{i}^{\alpha}(\theta)\right)^{2}\left(\frac{1}{T}\sum_{t=1}^{T}\partial_{\alpha_{N}\alpha_{N}\alpha_{N}}l_{N,t}(\theta,\phi)|_{\phi=\phi(\theta)}\right)$
+$=O_{\mathbb{P}}\left(1/(NT)\right),$
+uniformly in $\theta$. By similar arguments, we also have
+$\sup_{\theta\in\Theta}C_{NT,2}(\theta)=O_{\mathbb{P}}\left(1/(NT)^{3/2}\right),\qquad\sup_{\theta\in\Theta}C_{NT,3}(\theta)=O_{\mathbb{P}}\left(1/(NT)^{3/2}\right).$
 Therefore, when $\psi_{j}=\alpha_{i}$, we have
-$\sup_{\theta\in\Theta}\sup_{a_{j}\in(0,1)}r_{\psi_{j}}^{s}(\theta,a_{j})$
-$=$ $-\frac{1}{2N}\bigg{(}\sum_{i=1}^{N-1}\widetilde{u}_{i}^{\breve{\alpha}}(\theta)\bigg{)}^{2}\bigg{(}\frac{1}{T}\sum_{t=1}^{T}\partial_{\alpha_{N}\alpha_{N}\alpha_{N}}l_{N,t}(\theta,\phi)|_{\phi=\phi(\theta)}\bigg{)}+o_{\mathbb{P}}\big{(}1/(NT)\big{)}$ (A.13)
-$=$ $O_{\mathbb{P}}\big{(}1/(NT)\big{)}.$
-Similarly, when $\psi_{j}=\gamma_{t},$ we have
-$\sup_{\theta\in\Theta}\sup_{a_{j}\in(0,1)}r_{\psi_{j}}^{s}(\theta,a_{j})$
-$=$ $-\frac{1}{2T}\bigg{(}\sum_{t=1}^{T-1}\widetilde{u}_{t}^{\breve{\gamma}}(\theta)\bigg{)}^{2}\bigg{(}\frac{1}{N}\sum_{i=1}^{N}\partial_{\gamma_{T}\gamma_{T}\gamma_{T}}l_{i,T}(\theta,\phi)|_{\phi=\phi(\theta)}\bigg{)}+o_{\mathbb{P}}\big{(}1/(NT)\big{)}$ (A.14)
-$=$ $O_{\mathbb{P}}\big{(}1/(NT)\big{)},$
-where $\widetilde{u}_{t}^{\breve{\gamma}}(\theta)$ denotes the $t$-th element of $H_{\breve{\gamma}\breve{\alpha}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\alpha}}(\theta)+H_{\breve{\gamma}\breve{\gamma}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\gamma}}(\theta)$ for $t=1,\ldots,T-1.$ This shows the desired result.
+$\quad\quad\sup_{\theta\in\Theta}\sup_{a_{j}\in(0,1)}r_{\psi_{j}}^{s}(\theta,a_{j})$
+$=-\frac{1}{2N}\left(\sum_{i=1}^{N-1}v_{i}^{\alpha}(\theta)\right)^{2}\left(\frac{1}{T}\sum_{t=1}^{T}\partial_{\alpha_{N}\alpha_{N}\alpha_{N}}l_{N,t}(\theta,\phi)|_{\phi=\phi(\theta)}\right)+o_{\mathbb{P}}\left(1/(NT)\right)$
+$=O_{\mathbb{P}}\left(1/(NT)\right).$ (A.14)
+Similarly, when $\psi_{j}=\gamma_{t}$ ($t=1,\ldots,T-1$),
+$\quad\quad\sup_{\theta\in\Theta}\sup_{a_{j}\in(0,1)}r_{\psi_{j}}^{s}(\theta,a_{j})$
+$=-\frac{1}{2T}\left(\sum_{t=1}^{T-1}v_{t}^{\gamma}(\theta)\right)^{2}\left(\frac{1}{N}\sum_{i=1}^{N}\partial_{\gamma_{T}\gamma_{T}\gamma_{T}}l_{i,T}(\theta,\phi)|_{\phi=\phi(\theta)}\right)+o_{\mathbb{P}}\left(1/(NT)\right)$
+$=O_{\mathbb{P}}\left(1/(NT)\right),$ (A.15)
+where $v_{t}^{\gamma}(\theta)$ denotes the $t$-th element of $H_{\gamma\alpha^{\prime}}^{\ast}(\theta)\widetilde{s}_{\alpha}(\theta)+H_{\gamma\gamma^{\prime}}^{\ast}(\theta)\widetilde{s}_{\gamma}(\theta)$ for $t=1,\ldots,T-1$.
+This leads to the desired result.
+The next lemma states that the final remainder, $r(\theta)$ in Equation (2.14), is $O\left(1/\left(NT\right)\right)$ uniformly in $\theta$, where
+$r(\theta)$ $:=\frac{1}{2}(r^{s}(\theta,a))^{\prime}[H\left(\theta\right)]^{-1}r^{s}(\theta,a)+r^{l}(\theta,a_{0}),$ (A.16)
+$r^{l}\left(\theta,a_{0}\right)$ $:=\frac{1}{3}\left(\widehat{\psi}(\theta)-\psi(\theta)\right)^{\prime}r^{s}(\theta,a_{0}\times\iota_{(N+T-2)K})$ (A.17)
+for some $a_{0}\in(0,1)$.
+Here $r^{l}\left(\theta,a_{0}\right)=:r^{l}(\theta)$ as in Equation (2.12) from the main text.
+The dependency of $r(\theta)$ on $a$ and $a_{0}$ is suppress for simplicity here.
 Lemma A.3
-Under Assumption 1, $r(\theta)$ defined by (2.7) satisfies
+Under Assumption 1, $r(\theta)$ defined by (A.16) satisfies
 $\sup_{\theta\in\Theta}\mathbb{E}[r(\theta)]=O\left(1/\left(NT\right)\right),$
@@
 Proof.
-In the proof, we retain the notation and setup from Lemmas 1 and A.1.
-From Lemma A.1, we know that $NT\cdot r^{s}\left(\theta,a\right)$ is comparable to $(NT)^{3/4}\cdot s(\theta).$ Following similar arguments as in Lemma 1, we can show that
-$\mathbb{E}\big{[}(r^{s}(\theta,a))^{\prime}[{H}\left(\theta\right)]^{-1}r^{s}(\theta,a)\big{]}=O\big{(}1/(NT)\big{)}.$
+In the proof, we retain the notation and setup from Lemmas A.1 and A.2.
+From Lemma A.2, we know that $NT\cdot r^{s}\left(\theta,a\right)$ is comparable to $(NT)^{3/4}\cdot s(\theta)$.
+Following similar arguments as in Lemma A.1, we can show that
+$\mathbb{E}\left\{(r^{s}(\theta,a))^{\prime}[{H}\left(\theta\right)]^{-1}r^{s}(\theta,a)\right\}=O\left(1/(NT)\right).$
 Moreover, from (A.12), (A.14), and (A.14), we have
-$\big{(}\widehat{\psi}(\theta)-\psi(\theta)\big{)}^{\prime}r^{s}(\theta,a_{0}\,\iota_{N+T-2})$
-$=$ $\frac{1}{2N}\bigg{(}\sum_{i=1}^{N-1}\widetilde{u}_{i}^{\breve{\alpha}}(\theta)\bigg{)}^{2}\bigg{(}\frac{1}{T}\sum_{t=1}^{T}\partial_{\alpha_{N}\alpha_{N}\alpha_{N}}l_{N,t}(\theta,\phi)|_{\phi=\phi(\theta)}\bigg{)}$
-$+\frac{1}{2T}\bigg{(}\sum_{t=1}^{T-1}\widetilde{u}_{t}^{\breve{\gamma}}(\theta)\bigg{)}^{2}\bigg{(}\frac{1}{N}\sum_{i=1}^{N}\partial_{\gamma_{T}\gamma_{T}\gamma_{T}}l_{i,T}(\theta,\phi)|_{\phi=\phi(\theta)}\bigg{)}$
-$\quad+\bigg{(}\sum_{i=1}^{N-1}\widetilde{u}_{i}^{\breve{\alpha}}(\theta)+\sum_{t=1}^{T-1}\widetilde{u}_{t}^{\breve{\gamma}}(\theta)\bigg{)}\cdot o_{\mathbb{P}}\big{(}1/(NT)\big{)}$
-$=$ $O_{\mathbb{P}}\big{(}1/(NT)^{5/4}\big{)}.$ (A.15)
-By Assumption 1 and applying the $L_{1}$ Convergence Theorem to (A.15), we find that
-$\sup_{\theta\in\Theta}\mathbb{E}\big{[}\big{(}\widehat{\psi}(\theta)-\psi(\theta)\big{)}^{\prime}r^{s}\left(\theta,a_{0}\,\iota_{N+T-2}\right)\big{]}=o\big{(}1/(NT)\big{)}.$
+$\quad\quad\left(\widehat{\psi}(\theta)-\psi(\theta)\right)^{\prime}r^{s}(\theta,a_{0}\,\iota_{N+T-2})$
+$=\frac{1}{2N}\left(\sum_{i=1}^{N-1}v_{i}^{\alpha}(\theta)\right)^{2}\left(\frac{1}{T}\sum_{t=1}^{T}\partial_{\alpha_{N}\alpha_{N}\alpha_{N}}l_{N,t}(\theta,\phi)|_{\phi=\phi(\theta)}\right)$
+$\quad\quad+\frac{1}{2T}\left(\sum_{t=1}^{T-1}v_{t}^{\gamma}(\theta)\right)^{2}\left(\frac{1}{N}\sum_{i=1}^{N}\partial_{\gamma_{T}\gamma_{T}\gamma_{T}}l_{i,T}(\theta,\phi)|_{\phi=\phi(\theta)}\right)$
+$\quad\quad+\left(\sum_{i=1}^{N-1}v_{i}^{\alpha}(\theta)+\sum_{t=1}^{T-1}v_{t}^{\gamma}(\theta)\right)\cdot o_{\mathbb{P}}\left(1/(NT)\right)$
+$=O_{\mathbb{P}}\left(1/(NT)^{5/4}\right).$ (A.18)
+By Assumption 1 and applying the $L_{1}$ convergence theorem to A.18, we find that
+$\sup_{\theta\in\Theta}\mathbb{E}\left[\left(\widehat{\psi}(\theta)-\psi(\theta)\right)^{\prime}r^{s}\left(\theta,a_{0}\,\iota_{N+T-2}\right)\right]=o\left(1/(NT)\right).$
 Thus,
-$\mathbb{E}\big{[}r(\theta)\big{]}=\frac{1}{2}\mathbb{E}\big{[}(r^{s}\left(\theta,a\right))^{\prime}[{H}(\theta)]^{-1}r^{s}\left(\theta,a\right)\big{]}+\frac{1}{3}\mathbb{E}\big{[}\big{(}\widehat{\psi}(\theta)-\psi(\theta)\big{)}^{\prime}r^{s}\left(\theta,a_{0}\,\iota_{N+T-2}\right)\big{]}=O\big{(}1/(NT)\big{)}$
+$\mathbb{E}\left[r(\theta)\right]=\frac{1}{2}\mathbb{E}\left[(r^{s}\left(\theta,a\right))^{\prime}[{H}(\theta)]^{-1}r^{s}\left(\theta,a\right)\right]+\frac{1}{3}\mathbb{E}\left[\left(\widehat{\psi}(\theta)-\psi(\theta)\right)^{\prime}r^{s}\left(\theta,a_{0}\,\iota_{N+T-2}\right)\right]=O\left(1/(NT)\right)$
 uniformly in $\theta$.
-A.2 Proof of Theorem 1
-From (2.7), we have
-$\mathbb{E}[l(\theta)]=\mathbb{E}[\widehat{l}(\theta)]+\frac{1}{2}\mathbb{E}\big{\{}s^{\prime}(\theta)[H(\theta)]^{-1}s(\theta)\big{\}}-\mathbb{E}[r(\theta)].$
-Thus, the desired result follows directly from Lemmas 1 and A.2.
-A.3 Proof of Theorem 2
-We retain the notation introduced in Lemmas 1-A.2. First, we show that
-$\widehat{B}_{\breve{\alpha}}(\theta)-B_{\breve{\alpha}}(\theta)=o_{\mathbb{P}}\big{(}1/\sqrt{NT}\big{)},\quad\text{and}\quad\widehat{B}_{\breve{\gamma}}(\theta)-B_{\breve{\gamma}}(\theta)=o_{\mathbb{P}}\big{(}1/\sqrt{NT}\big{)}.$ (A.16)
-Note that
-$\widehat{B}_{\breve{\alpha}}(\theta)-B_{\breve{\alpha}}(\theta)=\frac{1}{2}\text{trace}\big{\{}\widehat{H}_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}\left(\theta\right)[\Sigma^{\breve{\alpha}}(\theta)-\widehat{\Sigma}^{\breve{\alpha}}_{\tau}(\theta)]\big{\}}+\frac{1}{2}\text{trace}\big{\{}[\widehat{H}_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)-H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)]\Sigma^{\breve{\alpha}}(\theta)\big{\}}.$
-Define $\widetilde{s}_{i,t}^{\alpha}(\theta):=\{\partial_{\alpha_{i}}l_{i,t}(\theta,{\phi})-\mathbb{E}[\partial_{\alpha_{i}}l_{i,t}(\theta,{\phi})]\}|_{\phi={\phi}(\theta)}.$ Then we have
-$N^{2}\big{[}{\Sigma}_{1}^{\breve{\alpha}}(\theta)-\widehat{\Sigma}_{1,\tau}^{\breve{\alpha}}(\theta)\big{]}$
+Next, we prove Theorem 1 and 2 below.
+Proof of Theorem 1. Following the expansion in (2.12) and (2.13), we have
+$\mathbb{E[}\widehat{l}(\theta)-l(\theta)]=-\frac{1}{2}\mathbb{E\{}s^{\prime}(\theta)[H(\theta)]^{-1}s(\theta)\}+\mathbb{E}r(\theta)$
+for $r\left(\theta\right)$ defined in (A.16). The claim in Theorem 1 follows directly from Lemmas A.1 and A.3.
+The proof of Theorem 2 is as follows.
+Proof of Theorem 2. We retain the notation introduced in Lemmas A.1-A.3. We first show
+$\widehat{B}_{\alpha}(\theta)-B_{\alpha}(\theta)=o_{\mathbb{P}}(1/\sqrt{NT}),\qquad\widehat{B}_{\gamma}(\theta)-B_{\gamma}(\theta)=o_{\mathbb{P}}(1/\sqrt{NT}).$ (A.19)
+To show the first part of (A.19), note that
+$\widehat{B}_{\alpha}(\theta)-B_{\alpha}(\theta)=\frac{1}{2}\operatorname*{trace}\left\{\widehat{H}_{\alpha\alpha^{\prime}}^{\ast}\left(\theta\right)\left[S_{\alpha\alpha}\left(\theta\right)-\widehat{S}_{\alpha\alpha}\left(\theta\right)\right]\right\}+\frac{1}{2}\operatorname*{trace}\left\{\left[\widehat{H}_{\alpha\alpha^{\prime}}^{\ast}(\theta)-H_{\alpha\alpha^{\prime}}^{\ast}(\theta)\right]S_{\alpha\alpha}\left(\theta\right)\right\},$
+where
+$S_{\alpha\alpha}\left(\theta\right)$ $=S_{\alpha\alpha}^{1}\left(\theta\right)+S_{\alpha\alpha}^{2}\left(\theta\right),$
+$S_{\alpha\alpha}^{1}\left(\theta\right)$ $:=\operatorname*{diag}\left\{\mathbb{E}\left[s_{1}^{\alpha}\left(\theta\right)\left(s_{1}^{\alpha}\left(\theta\right)\right)^{\prime}\right],\ldots,\mathbb{E}\left[s_{N-1}^{\alpha}\left(\theta\right)\left(s_{N-1}^{\alpha}\left(\theta\right)\right)^{\prime}\right]\right\},$
+$S_{\alpha\alpha}^{2}\left(\theta\right)$ $:=(\iota_{N-1}\iota_{N-1}^{\prime})\otimes\mathbb{E}\left[s_{N}^{\alpha}\left(\theta\right)\left(s_{N}^{\alpha}\left(\theta\right)\right)^{\prime}\right];$
+$\widehat{S}_{\alpha\alpha}\left(\theta\right)$ $=\widehat{S}_{\alpha\alpha}^{1}\left(\theta\right)+\widehat{S}_{\alpha\alpha}^{2}\left(\theta\right),$
+$S_{\alpha\alpha}^{1}\left(\theta\right)$ $:=\operatorname*{diag}\left\{\widehat{s}_{1}^{\alpha}\left(\theta\right)\left(\widehat{s}_{1}^{\alpha}\left(\theta\right)\right)^{\prime},\ldots,\widehat{s}_{N-1}^{\alpha}\left(\theta\right)\left(\widehat{s}_{N-1}^{\alpha}\left(\theta\right)\right)^{\prime}\right\},$
+(${}_{\alpha\alpha}^{2}\left(\theta\right)$ $:=(\iota_{N-1}\iota_{N-1}^{\prime})\otimes\left[\widehat{s}_{N}^{\alpha}\left(\theta\right)\left(\widehat{s}_{N}^{\alpha}\left(\theta\right)\right)^{\prime}\right].$
+Here we have
+$\quad\quad N^{2}\left[S_{\alpha\alpha}^{1}\left(\theta\right)-\widehat{S}_{\alpha\alpha}^{1}\left(\theta\right)\right]$
 $=\frac{1}{T^{2}}\sum_{t=1}^{T}\sum_{s=1}^{T}\mathds{1}\left\{\left|t-s\right|\leq\tau\right\}$
-[…]
-$+\frac{1}{T^{2}}\sum_{t=1}^{T}\sum_{s=1}^{T}\mathds{1}\left\{\left|t-s\right|>\tau\right\}$
-×$\qquad\times\text{diag}\big{\{}\mathbb{E}\big{[}\widetilde{s}_{1,t}^{\alpha}(\theta)\big{(}\widetilde{s}_{1,s}^{\alpha}(\theta)\big{)}^{\prime}\big{]},\ldots,\mathbb{E}\big{[}\widetilde{s}_{N-1,t}^{\alpha}(\theta)\big{(}\widetilde{s}_{N-1,s}^{\alpha}(\theta)\big{)}^{\prime}\big{]}\big{\}}.$ (A.17)
+×$\quad\quad\times\operatorname*{diag}\left\{\widetilde{s}_{1,t}^{\alpha}(\theta)\left(\widetilde{s}_{1,s}^{\alpha}(\theta)\right)^{\prime}-\widehat{s}_{1,t}^{\alpha}(\theta)\left(\widehat{s}_{1,s}^{\alpha}(\theta)\right)^{\prime},\ldots,\widetilde{s}_{N-1,t}^{\alpha}(\theta)\left(\widetilde{s}_{N-1,s}^{\alpha}(\theta)\right)^{\prime}-\widehat{s}_{N-1,t}^{\alpha}(\theta)\left(\widehat{s}_{N-1,s}^{\alpha}(\theta)\right)^{\prime}\right\}$
+$\quad\quad+\frac{1}{T^{2}}\sum_{t=1}^{T}\sum_{s=1}^{T}\mathds{1}\left\{\left|t-s\right|>\tau\right\}$
+×$\quad\quad\times\operatorname*{diag}\left\{\mathbb{E}\left[\widetilde{s}_{1,t}^{\alpha}(\theta)\left(\widetilde{s}_{1,s}^{\alpha}(\theta)\right)^{\prime}\right],\ldots,\mathbb{E}\left[\widetilde{s}_{N-1,t}^{\alpha}(\theta)\left(\widetilde{s}_{N-1,s}^{\alpha}(\theta)\right)^{\prime}\right]\right\}.$ (A.20)
 Under Assumptions 1 and 1, by Davydov’s inequality,
-$\mathbb{E}\big{[}\widetilde{s}_{i,t}^{\alpha}(\theta)\big{(}\widetilde{s}_{i,s}^{\alpha}(\theta)\big{)}^{\prime}\big{]}=\mathbb{E}\big{[}\widetilde{s}_{i,t}^{\alpha}(\theta)\widetilde{s}_{i,s}^{\alpha}(\theta)\big{]}\leq 8[a_{i}(|t-s|)]^{1-2/\eta}\sup_{i,t}\{\mathbb{E}[g^{\eta}(w_{it})]\}^{2/\eta},$
+$\mathbb{E}\left[\widetilde{s}_{i,t}^{\alpha}(\theta)\left(\widetilde{s}_{i,s}^{\alpha}(\theta)\right)^{\prime}\right]\leq 8[a_{i}(|t-s|)]^{1-2/\eta}\sup_{i,t}\{\mathbb{E}[g^{\eta}(w_{it})]\}^{2/\eta},$
 where we assume without loss of generality that $K=1$. Thus, as $T\rightarrow\infty$,
-$\frac{1}{T}\sum_{t=1}^{T}\sum_{s=1}^{T}\mathds{1}\left\{\left|t-s\right|>\tau\right\}\mathbb{E}\big{[}\widetilde{s}_{i,t}^{\alpha}(\theta)\big{(}\widetilde{s}_{i,s}^{\alpha}(\theta)\big{)}^{\prime}\big{]}$
-$\leq$ $8\sup_{i,t}\{\mathbb{E}[g^{\eta}(w_{it})]\}^{2/\eta}\cdot\frac{1}{T}\sum_{t=1}^{T}\sum_{s=1}^{T}\mathds{1}\left\{\left|t-s\right|>\tau\right\}[a_{i}(|t-s|)]^{1-2/\eta}$
-$\leq$ $8\sup_{i,t}\{\mathbb{E}[g^{\eta}(w_{it})]\}^{2/\eta}\cdot\frac{T-\tau}{T}\sum_{m=\tau+1}^{\infty}[a_{i}(m)]^{1-2/\eta}$
+$\quad\quad\frac{1}{T}\sum_{t=1}^{T}\sum_{s=1}^{T}\mathds{1}\left\{\left|t-s\right|>\tau\right\}\mathbb{E}\left[\widetilde{s}_{i,t}^{\alpha}(\theta)\left(\widetilde{s}_{i,s}^{\alpha}(\theta)\right)^{\prime}\right]$
+$\leq 8\sup_{i,t}\{\mathbb{E}[g^{\eta}(w_{it})]\}^{2/\eta}\left(\frac{1}{T}\sum_{t=1}^{T}\sum_{s=1}^{T}\mathds{1}\left\{\left|t-s\right|>\tau\right\}[a_{i}(|t-s|)]^{1-2/\eta}\right)$
+$\leq 8\sup_{i,t}\{\mathbb{E}[g^{\eta}(w_{it})]\}^{2/\eta}\left(\frac{T-\tau}{T}\sum_{m=\tau+1}^{\infty}[a_{i}(m)]^{1-2/\eta}\right)$
 $=o(1)$
-given that $\tau/T\to 0.$ Combining this with (A.17) implies
-$\sup_{\theta\in\Theta}N^{2}\|{\Sigma}_{1}^{\breve{\alpha}}(\theta)-\widehat{\Sigma}_{1,\tau}^{\breve{\alpha}}(\theta)\|_{\max}=o_{\mathbb{P}}(1/T).$ (A.18)
-From the proof of Lemma A.1, we also have $\sup_{\theta\in\Theta}\|\widehat{\psi}(\theta)-\psi(\theta)\|_{\max}=o_{\mathbb{P}}(1).$ Moreover, under the conditional strong mixing condition, each element of $T^{-1}\widehat{H}_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)$ converges in probability to its counterpart in $T^{-1}H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)$ by the law of large numbers.
-Therefore, from (A.18),
-$\text{trace}\big{\{}\widehat{H}_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)[\Sigma_{1}^{\breve{\alpha}}(\theta)-\widehat{\Sigma}_{1,\tau}^{\breve{\alpha}}(\theta)]\big{\}}$
-$=$ $\frac{T}{N^{2}}\text{trace}\big{\{}T^{-1}\widehat{H}_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)\,N^{2}[\Sigma^{\breve{\alpha}}_{1}(\theta)-\widehat{\Sigma}^{\breve{\alpha}}_{1,\tau}(\theta)]\big{\}}$
-$=$ $\frac{T}{N}\cdot\frac{1}{N}\sum_{i=1}^{N-1}h_{i,i}^{\ast\breve{\alpha}}(\theta)\cdot o_{\mathbb{P}}(1/T)$
+given that $\tau/T\rightarrow 0$. Combining this with (A.20), we have
+$\sup_{\theta\in\Theta}N^{2}\left\|S_{\alpha\alpha}^{1}\left(\theta\right)-\widehat{S}_{\alpha\alpha}^{1}\left(\theta\right)\right\|_{\max}=o_{\mathbb{P}}(1/T).$ (A.21)
+From the proof of Lemma A.2, we also have $\sup_{\theta\in\Theta}\|\widehat{\psi}(\theta)-\psi(\theta)\|_{\max}=o_{\mathbb{P}}(1)$.
+Moreover, under the conditional strong mixing assumption, each element of $T^{-1}\widehat{H}_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)$ converges in probability to the corresponding element in $T^{-1}H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)$ by the law of large numbers.
+Therefore (A.21) further leads to
+$\quad\quad\operatorname*{trace}\left\{\widehat{H}_{\alpha\alpha^{\prime}}^{\ast}\left(\theta\right)\left[S_{\alpha\alpha}^{1}\left(\theta\right)-\widehat{S}_{\alpha\alpha}^{1}\left(\theta\right)\right]\right\}$
+$=\frac{T}{N^{2}}\operatorname*{trace}\left\{T^{-1}\widehat{H}_{\alpha\alpha^{\prime}}^{\ast}(\theta)\cdot N^{2}\left[S_{\alpha\alpha}^{1}\left(\theta\right)-\widehat{S}_{\alpha\alpha}^{1}\left(\theta\right)\right]\right\}$
+$=\frac{T}{N}\frac{1}{N}\sum_{i=1}^{N-1}h_{i,i}^{\ast\alpha}(\theta)\cdot o_{\mathbb{P}}(1/T)$
 $=o_{\mathbb{P}}(1/T).$
-Similarly,
-$\text{trace}\big{\{}\widehat{H}_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)[\Sigma_{2}^{\breve{\alpha}}(\theta)-\widehat{\Sigma}_{2,\tau}^{\breve{\alpha}}(\theta)]\big{\}}$
-$=$ $\frac{T}{N^{2}}\text{trace}\big{\{}T^{-1}\widehat{H}_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)\,N^{2}[\Sigma_{2}^{\breve{\alpha}}(\theta)-\widehat{\Sigma}_{2,\tau}^{\breve{\alpha}}(\theta)]\big{\}}$
-$=$ $\frac{T}{N^{2}}\cdot\sum_{i=1}^{N-1}\sum_{j=1}^{N-1}h_{i,j}^{\ast\breve{\alpha}}(\theta)\cdot o_{\mathbb{P}}(1/T)$
+A similar procedure can be followed to show
+$\quad\quad\operatorname*{trace}\left\{\widehat{H}_{\alpha\alpha^{\prime}}^{\ast}\left(\theta\right)\left[S_{\alpha\alpha}^{2}\left(\theta\right)-\widehat{S}_{\alpha\alpha}^{2}\left(\theta\right)\right]\right\}$
+$=\frac{T}{N^{2}}\operatorname*{trace}\left\{T^{-1}\widehat{H}_{\alpha\alpha^{\prime}}^{\ast}(\theta)\cdot N^{2}\left[S_{\alpha\alpha}^{2}\left(\theta\right)-\widehat{S}_{\alpha\alpha}^{2}\left(\theta\right)\right]\right\}$
+$=\frac{T}{N^{2}}\sum_{i=1}^{N-1}\sum_{j=1}^{N-1}h_{i,j}^{\ast\alpha}(\theta)\cdot o_{\mathbb{P}}(1/T)$
 $=o_{\mathbb{P}}(1/N^{2}),$
-where the last equality follows from the fact that $\sum_{j=1}^{N-1}h_{i,j}^{\ast\breve{\alpha}}(\theta)=O(1/N)$ uniformly in both $i$ and $\theta,$ as shown in Lemma A.1. Therefore,
-$\text{trace}\big{\{}\widehat{H}_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)[\Sigma^{\breve{\alpha}}(\theta)-\widehat{\Sigma}^{\breve{\alpha}}_{\tau}(\theta)]\big{\}}=o_{\mathbb{P}}\big{(}1/\sqrt{NT}\big{)}.$ (A.19)
-Moreover, from Theorem 1, we also have
-$\text{trace}\big{\{}[\widehat{H}_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)-H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)]\Sigma^{\breve{\alpha}}(\theta)\big{\}}=o_{\mathbb{P}}(1)\cdot\text{trace}\big{\{}H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)\Sigma^{\breve{\alpha}}(\theta)\big{\}}=o_{\mathbb{P}}(1)\cdot B_{\breve{\alpha}}(\theta)=o_{\mathbb{P}}\big{(}1/\sqrt{NT}\big{)}.$ (A.20)
-Combining (A.19) and (A.20) shows the first part of (A.16). The second part of (A.16) can be shown similarly.
+where the last equality follows from, as shown in Lemma A.2, $\sum_{j=1}^{N-1}h_{i,j}^{\ast\alpha}(\theta)=O(1/N)$ uniformly in both $i$ and $\theta$. The above leads to
+$\operatorname*{trace}\left\{\widehat{H}_{\alpha\alpha^{\prime}}^{\ast}(\theta)\left[S_{\alpha\alpha}\left(\theta\right)-\widehat{S}_{\alpha\alpha}\left(\theta\right)\right]\right\}=o_{\mathbb{P}}(1/\sqrt{NT}).$ (A.22)
+Second, from Theorem 1, we also have
+$\operatorname*{trace}\left\{\left[\widehat{H}_{\alpha\alpha^{\prime}}^{\ast}(\theta)-H_{\alpha\alpha^{\prime}}^{\ast}(\theta)\right]S_{\alpha\alpha}\left(\theta\right)\right\}=o_{\mathbb{P}}(1)\cdot\operatorname*{trace}\left\{H_{\alpha\alpha^{\prime}}^{\ast}(\theta)S_{\alpha\alpha}\left(\theta\right)\right\}=o_{\mathbb{P}}(1)\cdot B_{\alpha}(\theta)=o_{\mathbb{P}}(1/\sqrt{NT}).$ (A.23)
+Combining A.22 and A.23 leads to the first part of A.19. The second part of A.19 can be shown similarly.
 Noting that
-$l(\theta)=\widehat{l}(\theta)+\frac{1}{2}s_{\breve{\alpha}}^{\prime}(\theta)H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)s_{\breve{\alpha}}(\theta)+\frac{1}{2}s_{\breve{\gamma}}^{\prime}(\theta)H_{\breve{\gamma}\breve{\gamma}^{\prime}}^{\ast}(\theta)s_{\breve{\gamma}}(\theta)+o_{\mathbb{P}}\big{(}1/\sqrt{NT}\big{)}$
+$l(\theta)=\widehat{l}(\theta)+\frac{1}{2}s_{\alpha}^{\prime}(\theta)H_{\alpha\alpha^{\prime}}^{\ast}(\theta)s_{\alpha}(\theta)+\frac{1}{2}s_{\gamma}^{\prime}(\theta)H_{\gamma\gamma^{\prime}}^{\ast}(\theta)s_{\gamma}(\theta)+o_{\mathbb{P}}(1/\sqrt{NT})$
 as implied by the proof of Theorem 1, we have
-$L(\theta)-l(\theta)$ $=\big{[}\widehat{B}_{\breve{\alpha}}(\theta)-B_{\breve{\alpha}}(\theta)\big{]}+\big{[}\widehat{B}_{\breve{\gamma}}(\theta)-B_{\breve{\gamma}}(\theta)\big{]}$
-$\quad-\bigg{[}\frac{1}{2}s_{\breve{\alpha}}^{\prime}(\theta)H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)s_{\breve{\alpha}}(\theta)-B_{\breve{\alpha}}(\theta)\bigg{]}-\bigg{[}\frac{1}{2}s_{\breve{\gamma}}^{\prime}(\theta)H_{\breve{\gamma}\breve{\gamma}^{\prime}}^{\ast}(\theta)s_{\breve{\gamma}}(\theta)-B_{\breve{\gamma}}(\theta)\bigg{]}$
-$+o_{\mathbb{P}}\big{(}1/\sqrt{NT}\big{)}.$
-Moreover, from Lemma 1,
-$\frac{1}{2}s_{\breve{\alpha}}^{\prime}(\theta)H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)s_{\breve{\alpha}}(\theta)-B_{\breve{\alpha}}(\theta)=o_{\mathbb{P}}\big{(}1/\sqrt{NT}\big{)},\quad\text{and}\quad\frac{1}{2}s_{\breve{\gamma}}^{\prime}\left(\theta\right)H_{\breve{\gamma}\breve{\gamma}^{\prime}}^{\ast}\left(\theta\right)s_{\breve{\gamma}}(\theta)-B_{\breve{\gamma}}(\theta)=o_{\mathbb{P}}\big{(}1/\sqrt{NT}\big{)}.$
-Thus, combining this with (A.16) yields the desired result.
-A.4 Proof of Corollary 1
-Consider a mean value expansion of $\triangledown_{\theta}L(\widehat{\theta}_{L})$ around $\theta_{0}:$
-$0=\triangledown_{\theta}L(\widehat{\theta}_{L})$ $=\triangledown_{\theta}L(\theta_{0})+\triangledown_{\theta\theta^{\prime}}L(\overline{\theta})(\widehat{\theta}_{L}-\theta_{0}),$ (A.21)
+$L(\theta)-l(\theta)$ $=\left[\widehat{B}_{\alpha}(\theta)-B_{\alpha}(\theta)\right]+\left[\widehat{B}_{\gamma}(\theta)-B_{\gamma}(\theta)\right]$
+$\quad\quad-\left[\frac{1}{2}\widetilde{s}_{\alpha}^{\prime}(\theta)H_{\alpha\alpha^{\prime}}^{\ast}(\theta)\widetilde{s}_{\alpha}(\theta)-B_{\alpha}(\theta)\right]-\left[\frac{1}{2}\widetilde{s}_{\gamma}^{\prime}(\theta)H_{\gamma\gamma^{\prime}}^{\ast}(\theta)\widetilde{s}_{\gamma}(\theta)-B_{\gamma}(\theta)\right]$
+$\quad\quad+o_{\mathbb{P}}(1/\sqrt{NT}).$
+As from Lemma A.1, we have
+$\frac{1}{2}\widetilde{s}_{\alpha}^{\prime}(\theta)H_{\alpha\alpha^{\prime}}^{\ast}(\theta)\widetilde{s}_{\alpha}(\theta)-B_{\alpha}(\theta)=o_{\mathbb{P}}(1/\sqrt{NT}),\qquad\frac{1}{2}\widetilde{s}_{\gamma}^{\prime}(\theta)H_{\gamma\gamma^{\prime}}^{\ast}(\theta)\widetilde{s}_{\gamma}(\theta)-B_{\gamma}(\theta)=o_{\mathbb{P}}(1/\sqrt{NT}).$
+The conclusion follows by combining this and (A.19).
+We continue to establishing the asymptotitc properties of $\{\widehat{\theta}_{L},\widehat{\xi}_{LR},\widehat{\xi}_{LM},\widehat{\xi}_{Wald}\}$.
+We first prove Corollary 1 below which states the consistency and asymptotic normality of $\widehat{\theta}_{L}$.
+Proof of Corollary 1. Consider a mean value expansion of $\triangledown_{\theta}L(\widehat{\theta}_{L})$ around $\theta_{0}$:
+$0=\triangledown_{\theta}L(\widehat{\theta}_{L})=\triangledown_{\theta}L(\theta_{0})+\triangledown_{\theta\theta^{\prime}}L(\overline{\theta})(\widehat{\theta}_{L}-\theta_{0}),$
 where each component of $\overline{\theta}$ lies between the corresponding components of $\widehat{\theta}_{L}$ and $\theta_{0}$. From Theorem 2, we further have
-$0=$ $\triangledown_{\theta}l(\theta_{0})+\triangledown_{\theta\theta^{\prime}}l(\overline{\theta})(\widehat{\theta}_{L}-\theta_{0})+o_{\mathbb{P}}\big{(}1/\sqrt{NT}\big{)}$ (A.22)
+$0=\triangledown_{\theta}l(\theta_{0})+\triangledown_{\theta\theta^{\prime}}l(\overline{\theta})(\widehat{\theta}_{L}-\theta_{0})+o_{\mathbb{P}}(1/\sqrt{NT})$ (A.24)
 Under Assumption 2, $\theta_{0}$ is the unique maximizer of $\mathbb{E}[l(\theta)]$. Since $L(\theta)-\mathbb{E}[l(\theta)]=o_{\mathbb{P}}(1)$, under Assumption 2, we conclude that
@@
 Thus, using Assumption 2 and Slutsky’s theorem, we obtain from (A.24) that
-$\sqrt{NT}(\widehat{\theta}_{L}-\theta_{0})$ $=-\big{(}\mathbb{E}[\triangledown_{\theta\theta^{\prime}}l(\theta_{0})]\big{)}^{-1}\sqrt{NT}\triangledown_{\theta}l(\theta_{0})+o_{\mathbb{P}}(1).$ (A.23)
-Given the conditional cross-sectional independence and the conditional strong mixing condition (Assumption 1), (A.23) shows that
-$\sqrt{NT}(\widehat{\theta}_{L}-\theta_{0})\overset{\mathbb{D}}{\longrightarrow}\mathcal{N}\left(0,\Sigma^{-1}\Omega\Sigma^{-1}\right)$ (A.24)
-where $\mathbb{E}[\triangledown_{\theta}l(\theta_{0})]=0,$ and
-$\Omega=\lim_{N,T\rightarrow\infty}(NT)\mathbb{E}[\triangledown_{\theta}l(\theta_{0})\triangledown_{\theta^{\prime}}l(\theta_{0})].$
-Under Assumption 2, the information matrix equality ($\Sigma^{-1}=\Omega$) holds, where the asymptotic variance in (A.24) simplifies to $\Sigma$.
-A.5 Proof of Lemma A.3
+$\sqrt{NT}(\widehat{\theta}_{L}-\theta_{0})=-\left\{\mathbb{E}[\triangledown_{\theta\theta^{\prime}}l(\theta_{0})]\right\}^{-1}\sqrt{NT}\triangledown_{\theta}l(\theta_{0})+o_{\mathbb{P}}(1).$ (A.25)
+Given the conditional cross-sectional independence and the conditional strong mixing condition (Assumption 1), (A.25) leads to
+$\left[\Sigma^{-1}\Omega\Sigma^{-1}\right]^{-1/2}\sqrt{NT}(\widehat{\theta}_{L}-\theta_{0})\overset{\mathbb{D}}{\longrightarrow}\mathcal{N}\left(0,\mathbb{I}_{K}\right)$ (A.26)
+as $N,T\rightarrow\infty$, where $\mathbb{E}[\triangledown_{\theta}l(\theta_{0})]=0$ and
+$\Sigma:=-\mathbb{E}\left[\triangledown_{\theta\theta^{\prime}}l(\theta_{0})\right],\qquad\Omega:=\left(NT\right)\mathbb{E}[\triangledown_{\theta}l(\theta_{0})\triangledown_{\theta^{\prime}}l(\theta_{0})].$
+Since we consider a likelihood model, $\Sigma^{-1}\Omega\Sigma^{-1}$ in (A.26) simplifies to $\Sigma^{-1}$ under the information matrix equality ($\Sigma=\Omega$).
+To show the asymptotic distribution of $\widehat{\xi}_{LR},\widehat{\xi}_{LM},\widehat{\xi}_{Wald}$, we find the following lemma useful.
 Lemma A.4
-Under Assumptions 1 and 2, and under the null hypothesis $H_{0}:R(\theta_{0})=0$, as $N$ and $T\to\infty,$
-$\widehat{\theta}_{R}-\widehat{\theta}_{L}=\Xi(\theta_{0})\triangledown_{\theta}l(\theta_{0})+o_{\mathbb{P}}\big{(}1/\sqrt{NT}\big{)},$
+Under Assumptions 1 and 2, and under the null hypothesis $H_{0}:R(\theta_{0})=0$, as $N$ and $T\rightarrow\infty,$
+$\widehat{\theta}_{R}-\widehat{\theta}_{L}=\Xi(\theta_{0})\triangledown_{\theta}l(\theta_{0})+o_{\mathbb{P}}(1/\sqrt{NT}),$
 where
-$\Xi(\theta_{0}):=\big{(}\mathbb{E[}\triangledown_{\theta\theta^{\prime}}l(\theta_{0})]\big{)}^{-1}[J(\theta_{0})]^{\prime}\Big{(}J(\theta_{0})\big{(}\mathbb{E[}\triangledown_{\theta\theta^{\prime}}l(\theta_{0})]\big{)}^{-1}[J(\theta_{0})]^{\prime}\Big{)}^{-1}J(\theta_{0})\big{(}\mathbb{E[}\triangledown_{\theta\theta^{\prime}}l(\theta_{0})]\big{)}^{-1}.$
+$\Xi(\theta_{0}):=\left\{\mathbb{E[}\triangledown_{\theta\theta^{\prime}}l(\theta_{0})]\right\}^{-1}[J(\theta_{0})]^{\prime}\left\{J(\theta_{0})\left(\mathbb{E[}\triangledown_{\theta\theta^{\prime}}l(\theta_{0})]\right)^{-1}[J(\theta_{0})]^{\prime}\right\}^{-1}J(\theta_{0})\left\{\mathbb{E[}\triangledown_{\theta\theta^{\prime}}l(\theta_{0})]\right\}^{-1}.$
 Proof. Consider the constrained maximization problem:
-$\big{(}\widehat{\theta}_{R},\widehat{\lambda}\big{)}=\arg\max_{\theta\in\Theta,\lambda\in\mathbb{R}}\big{[}L(\theta)+\lambda^{\prime}R(\theta)\big{]}$
+$\left\{\widehat{\theta}_{R},\widehat{\lambda}\right\}=\arg\max_{\theta\in\Theta,\lambda\in\mathbb{R}}\left[L(\theta)+\lambda^{\prime}R(\theta)\right]$
 where $\lambda$ is the Lagrangian multiplier. The first-order conditions for this constrained maximization are
@@
 $\sqrt{NT}R(\widehat{\theta}_{R})$ $=0.$ (A.28)
-Following a similar proof as in Corollary 1, we have $\sqrt{NT}\big{(}\widehat{\theta}_{R}-\theta_{0}\big{)}=O_{\mathbb{P}}\left(1\right).$ Furthermore, $\sqrt{NT}\widehat{\lambda}=O_{\mathbb{P}}(1)$ as implied by (A.25) since $\sqrt{NT}\triangledown_{\theta}L(\theta_{0})=O_{\mathbb{P}}(1)$ and $J(\theta_{0})$ is of full row rank.
-By Theorem 2, we express the mean value expansion of $\triangledown_{\theta}L(\widehat{\theta}_{R})$ around $\theta_{0}$ as
-$\triangledown_{\theta}L(\widehat{\theta}_{R})$ $=\triangledown_{\theta}L(\theta_{0})+\triangledown_{\theta\theta^{\prime}}L(\overline{\theta})\big{(}\widehat{\theta}_{R}-\theta_{0}\big{)}$
-$=\triangledown_{\theta}l(\theta_{0})+\mathbb{E}[\triangledown_{\theta\theta^{\prime}}l(\theta_{0})]\big{(}\widehat{\theta}_{R}-\theta_{0}\big{)}+o_{\mathbb{P}}\big{(}1/\sqrt{NT}\big{)},$
+Following a similar proof as in Corollary 1, we have $\sqrt{NT}(\widehat{\theta}_{R}-\theta_{0})=O_{\mathbb{P}}\left(1\right)$.
+Furthermore, $\sqrt{NT}\widehat{\lambda}=O_{\mathbb{P}}(1)$ as implied by (A.27) since $\sqrt{NT}\triangledown_{\theta}L(\theta_{0})=O_{\mathbb{P}}(1)$ and $J(\theta_{0})$ is of full row rank.
+By Theorem 2, a mean value expansion of $\triangledown_{\theta}L(\widehat{\theta}_{R})$ around $\theta_{0}$ is
+$\triangledown_{\theta}L(\widehat{\theta}_{R})$ $=\triangledown_{\theta}L(\theta_{0})+\triangledown_{\theta\theta^{\prime}}L(\overline{\theta})(\widehat{\theta}_{R}-\theta_{0})$
+$=\triangledown_{\theta}l(\theta_{0})+\mathbb{E}[\triangledown_{\theta\theta^{\prime}}l(\theta_{0})](\widehat{\theta}_{R}-\theta_{0})+o_{\mathbb{P}}(1/\sqrt{NT}),$
 where each component of $\overline{\theta}$ lies between the corresponding components of $\widehat{\theta}_{R}$ and $\theta_{0}.$ Combining this with (A.27) yields that
-$\mathbb{E}[\triangledown_{\theta\theta^{\prime}}l(\theta_{0})]\sqrt{NT}\big{(}\widehat{\theta}_{R}-\theta_{0}\big{)}+[J(\theta_{0})]^{\prime}\sqrt{NT}\widehat{\lambda}=-\sqrt{NT}\triangledown_{\theta}l(\theta_{0})+o_{\mathbb{P}}(1).$ (A.27)
-Under $H_{0},$ from (A.26), we also obtain
+$\mathbb{E}[\triangledown_{\theta\theta^{\prime}}l(\theta_{0})]\sqrt{NT}(\widehat{\theta}_{R}-\theta_{0})+[J(\theta_{0})]^{\prime}\sqrt{NT}\widehat{\lambda}=-\sqrt{NT}\triangledown_{\theta}l(\theta_{0})+o_{\mathbb{P}}(1).$ (A.29)
+Under $H_{0}$ and from (A.28), we also obtain
 00 0 $=\sqrt{NT}R(\widehat{\theta}_{R})$
-$=\sqrt{NT}R(\theta_{0})+J(\theta_{0})\sqrt{NT}\big{(}\widehat{\theta}_{R}-\theta_{0}\big{)}+o_{\mathbb{P}}(1)$
-$=J(\theta_{0})\sqrt{NT}\big{(}\widehat{\theta}_{R}-\theta_{0}\big{)}+o_{\mathbb{P}}(1).$ (A.28)
+$=\sqrt{NT}R(\theta_{0})+J(\theta_{0})\sqrt{NT}(\widehat{\theta}_{R}-\theta_{0})+o_{\mathbb{P}}(1)$
+$=J(\theta_{0})\sqrt{NT}(\widehat{\theta}_{R}-\theta_{0})+o_{\mathbb{P}}(1).$ (A.30)
 Now, combining (A.29) and (A.30), we have
@@
 which has the solution
-$\sqrt{NT}\big{(}\widehat{\theta}_{R}-\theta_{0}\big{)}=-\left(\big{(}\mathbb{E}[\triangledown_{\theta\theta^{\prime}}l(\theta_{0})]\big{)}^{-1}-\Xi\left(\theta_{0}\right)\right)\sqrt{NT}\triangledown_{\theta}l(\theta_{0})+o_{\mathbb{P}}\left(1\right).$ (A.29)
+$\sqrt{NT}(\widehat{\theta}_{R}-\theta_{0})=-\left(\{\mathbb{E}[\triangledown_{\theta\theta^{\prime}}l(\theta_{0})]\}^{-1}-\Xi\left(\theta_{0}\right)\right)\sqrt{NT}\triangledown_{\theta}l(\theta_{0})+o_{\mathbb{P}}\left(1\right).$ (A.31)
 The conclusion follows by subtracting (A.25) in Corollary 1 from both sides of (A.31).
-A.6 Proof of Corollary 2
-Let $\widehat{\xi}:=-NT\triangledown_{\theta^{\prime}}l(\theta_{0})\Xi\left(\theta_{0}\right)\triangledown_{\theta}l(\theta_{0}).$ We first show that
+Finally, we proof the asymptotic distribution of $\widehat{\xi}_{LR},\widehat{\xi}_{LM},\widehat{\xi}_{Wald}$ below.
+Proof of Corollary 2.
+Let $\widehat{\xi}:=-NT\triangledown_{\theta^{\prime}}l(\theta_{0})\Xi\left(\theta_{0}\right)\triangledown_{\theta}l(\theta_{0})$ for $\Xi\left(\theta_{0}\right)$ defined in Lemma A.4.
+We first show that
 $\widehat{\xi}_{LR}=\widehat{\xi}+o_{\mathbb{P}}(1),\qquad\widehat{\xi}_{LM}=\widehat{\xi}+o_{\mathbb{P}}(1),\qquad\widehat{\xi}_{Wald}=\widehat{\xi}+o_{\mathbb{P}}(1).$ (A.32)
 A mean value expansion of $L(\widehat{\theta}_{R})$ around $\widehat{\theta}_{L}$ gives
-${L}(\widehat{\theta}_{R})={L}(\widehat{\theta}_{L})+\triangledown_{\theta^{\prime}}{L}(\widehat{\theta}_{L})\big{(}\widehat{\theta}_{R}-\widehat{\theta}_{L}\big{)}+\frac{1}{2}\big{(}\widehat{\theta}_{R}-\widehat{\theta}_{L}\big{)}^{\prime}\triangledown_{\theta\theta^{\prime}}L(\theta^{\ast})\big{(}\widehat{\theta}_{R}-\widehat{\theta}_{L}\big{)}$
-where each component of $\theta^{\ast}$ lies between the corresponding components of $\widehat{\theta}_{R}$ and $\widehat{\theta}_{L}$.
+${L}(\widehat{\theta}_{R})={L}(\widehat{\theta}_{L})+\triangledown_{\theta^{\prime}}{L}(\widehat{\theta}_{L})(\widehat{\theta}_{R}-\widehat{\theta}_{L})+\frac{1}{2}(\widehat{\theta}_{R}-\widehat{\theta}_{L})^{\prime}\triangledown_{\theta\theta^{\prime}}L(\dot{\theta})(\widehat{\theta}_{R}-\widehat{\theta}_{L})$
+where each component of $\dot{\theta}$ lies between the corresponding components of $\widehat{\theta}_{R}$ and $\widehat{\theta}_{L}$.
 Noting that $\triangledown_{\theta}{L}(\widehat{\theta}_{L})=0$, we obtain
-$\widehat{\xi}_{LR}=-NT\big{(}\widehat{\theta}_{R}-\widehat{\theta}_{L}\big{)}^{\prime}\triangledown_{\theta\theta^{\prime}}L(\theta^{\ast})\big{(}\widehat{\theta}_{R}-\widehat{\theta}_{L}\big{)}.$
-Since $\triangledown_{\theta\theta^{\prime}}L(\theta^{\ast})-\mathbb{E}[\triangledown_{\theta\theta^{\prime}}L(\theta_{0})]=o_{\mathbb{P}}(1),$ by Lemma A.3, we further conclude that
+$\widehat{\xi}_{LR}=-NT(\widehat{\theta}_{R}-\widehat{\theta}_{L})^{\prime}\triangledown_{\theta\theta^{\prime}}L(\dot{\theta})(\widehat{\theta}_{R}-\widehat{\theta}_{L}).$
+Since $\triangledown_{\theta\theta^{\prime}}L(\dot{\theta})-\mathbb{E}[\triangledown_{\theta\theta^{\prime}}L(\theta_{0})]=o_{\mathbb{P}}(1)$, by Lemma A.4, we further conclude that
 $\widehat{\xi}_{LR}=\widehat{\xi}+o_{\mathbb{P}}(1).$
 For $\widehat{\xi}_{LM},$ note that
-$\triangledown_{\theta}L(\widehat{\theta}_{R})$ $=\triangledown_{\theta}L(\widehat{\theta}_{L})+\triangledown_{\theta\theta^{\prime}}L(\dot{\theta})\big{(}\widehat{\theta}_{R}-\widehat{\theta}_{L}\big{)}$
-$=\mathbb{E}[\triangledown_{\theta\theta^{\prime}}L(\theta_{0})]\big{(}\widehat{\theta}_{R}-\widehat{\theta}_{L}\big{)}+o_{\mathbb{P}}\big{(}1/\sqrt{NT}\big{)},$
-where each component of $\dot{\theta}$ lies between the corresponding components of $\widehat{\theta}_{R}$ and $\widehat{\theta}_{L}.$ Thus,
-$\widehat{\xi}_{LM}$ $=-NT\big{(}\widehat{\theta}_{R}-\widehat{\theta}_{L}\big{)}^{\prime}\mathbb{E}[\triangledown_{\theta\theta^{\prime}}L(\theta_{0})][\triangledown_{\theta\theta^{\prime}}L(\widehat{\theta}_{R})]^{-1}\mathbb{E}[\triangledown_{\theta\theta^{\prime}}L(\theta_{0})]\big{(}\widehat{\theta}_{R}-\widehat{\theta}_{L}\big{)}+o_{\mathbb{P}}\left(1\right)$
-$=-NT\big{(}\widehat{\theta}_{R}-\widehat{\theta}_{L}\big{)}^{\prime}\mathbb{E}[\triangledown_{\theta\theta^{\prime}}L(\theta_{0})]\big{(}\widehat{\theta}_{R}-\widehat{\theta}_{L}\big{)}+o_{\mathbb{P}}\left(1\right)$
+$\triangledown_{\theta}L(\widehat{\theta}_{R})$ $=\triangledown_{\theta}L(\widehat{\theta}_{L})+\triangledown_{\theta\theta^{\prime}}L(\ddot{\theta})(\widehat{\theta}_{R}-\widehat{\theta}_{L})$
+$=\mathbb{E}[\triangledown_{\theta\theta^{\prime}}L(\theta_{0})](\widehat{\theta}_{R}-\widehat{\theta}_{L})+o_{\mathbb{P}}(1/\sqrt{NT}),$
+[…]
+$\widehat{\xi}_{LM}$ $=-NT(\widehat{\theta}_{R}-\widehat{\theta}_{L})^{\prime}\mathbb{E}[\triangledown_{\theta\theta^{\prime}}L(\theta_{0})][\triangledown_{\theta\theta^{\prime}}L(\widehat{\theta}_{R})]^{-1}\mathbb{E}[\triangledown_{\theta\theta^{\prime}}L(\theta_{0})](\widehat{\theta}_{R}-\widehat{\theta}_{L})+o_{\mathbb{P}}\left(1\right)$
+$=-NT(\widehat{\theta}_{R}-\widehat{\theta}_{L})^{\prime}\mathbb{E}[\triangledown_{\theta\theta^{\prime}}L(\theta_{0})](\widehat{\theta}_{R}-\widehat{\theta}_{L})+o_{\mathbb{P}}\left(1\right)$
 $=\widehat{\xi}+o_{\mathbb{P}}\left(1\right).$
 For $\widehat{\xi}_{Wald}$, a mean value expansion of $R(\widehat{\theta}_{L})$ gives
-$R(\widehat{\theta}_{L})=R(\widehat{\theta}_{R})+J(\ddot{\theta})(\widehat{\theta}_{L}-\widehat{\theta}_{R})=J(\ddot{\theta})(\widehat{\theta}_{L}-\widehat{\theta}_{R})$
-[…]
-$\widehat{\xi}_{Wald}$ $=-NT(\widehat{\theta}_{L}-\widehat{\theta}_{R})^{\prime}J(\ddot{\theta})^{\prime}\left(J(\widehat{\theta}_{L})[\triangledown_{\theta\theta^{\prime}}L(\widehat{\theta}_{L})]^{-1}J(\widehat{\theta}_{L})^{\prime}\right)^{-1}J(\ddot{\theta})(\widehat{\theta}_{L}-\widehat{\theta}_{R})$
+$R(\widehat{\theta}_{L})=R(\widehat{\theta}_{R})+J(\dddot{\theta})(\widehat{\theta}_{L}-\widehat{\theta}_{R})=J(\dddot{\theta})(\widehat{\theta}_{L}-\widehat{\theta}_{R})$
+where each component of $\dddot{\theta}$ lies between the corresponding components of $\widehat{\theta}_{L}$ and $\widehat{\theta}_{R}$.
+Since both $\widehat{\theta}_{L}$ and $\widehat{\theta}_{R}$(under $H_{0}$) are consistent estimators for $\theta_{0},$ we obtain that
+$\widehat{\xi}_{Wald}$ $=-NT(\widehat{\theta}_{L}-\widehat{\theta}_{R})^{\prime}J(\dddot{\theta})^{\prime}\left\{J(\widehat{\theta}_{L})[\triangledown_{\theta\theta^{\prime}}L(\widehat{\theta}_{L})]^{-1}J(\widehat{\theta}_{L})^{\prime}\right\}^{-1}J(\dddot{\theta})(\widehat{\theta}_{L}-\widehat{\theta}_{R})$
 $=\widehat{\xi}+o_{\mathbb{P}}\left(1\right).$
-Next we show
-$\widehat{\xi}\overset{\mathbb{D}}{\longrightarrow}\chi^{2}\left(r\right).$ (A.31)
-Since $\sqrt{NT}\triangledown_{\theta}l(\theta_{0})\overset{\mathbb{D}}{\longrightarrow}\mathcal{N}(0,\Sigma),$ by the delta method, we have
-$J(\theta_{0})\big{(}\mathbb{E}[\triangledown_{\theta\theta^{\prime}}l(\theta_{0})]\big{)}^{-1}\sqrt{NT}\triangledown_{\theta}l(\theta_{0})\overset{\mathbb{D}}{\longrightarrow}\mathcal{N}\big{(}0,J(\theta_{0})\Sigma J(\theta_{0})^{\prime}\big{)}.$
-This implies (A.31). Therefore, the desired result follows from (A.30) and (A.31). The proof is complete.
-B Additional Simulation Results
+Next, since $\Sigma^{-1/2}\sqrt{NT}\triangledown_{\theta}l(\theta_{0})\overset{\mathbb{D}}{\longrightarrow}\mathcal{N}(0,\mathbb{I}_{K})$, applying the delta method delivers $\widehat{\xi}\overset{\mathbb{D}}{\longrightarrow}\chi^{2}\left(r\right)$.
+The desired results follows from here.
+B Extra Simulation Results
 B.1 Logit Model
-Table B.1: Empirical Size of LM and Wald Tests for Logit Model
+Table B.1: Empirical Size of LM and Wald Tests for Logit Model, Design 1
 (a) LM Tests
@@
 […]
-Notes: The null hypothesis $H_{0}:(\beta_{0},\rho_{0})^{\prime}=(0.5,0.5)^{\prime}$ for the dynamic model, and $H_{0}:\beta_{0}=0.5$ for the static model, with a significance level $5\%$.
+Notes: Logit model with $\varepsilon_{it}$ standard logistic.
+Design 1: $\{\alpha_{k,i}^{0},\gamma_{k,t}^{0}\}\sim\mathcal{N}(0,0.01)$, $X_{it}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0}+\gamma_{1,t}^{0}+\gamma_{2,t}^{0},0.25)$, and $X_{i0}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0},0.25)$.
+The null hypothesis $H_{0}:(\beta_{0},\rho_{0})^{\prime}=(0.5,0.5)^{\prime}$ for the dynamic model, and $H_{0}:\beta_{0}=0.5$ for the static model, with a significance level $5\%$.
 The number of replications is $1000$.
-Table B.2: Empirical Power of LM and Wald Tests for Logit Model using $L(\theta)$
+Table B.2: Empirical Power of LM and Wald Tests for Logit Model using $L(\theta)$, Design 1
 (a) LM Tests
@@
 […]
-Notes: The null hypothesis $H_{0}:(\beta_{0},\rho_{0})^{\prime}=(0.5+\delta,0.5+\delta)^{\prime}$ for the dynamic model, and $H_{0}:\beta_{0}=0.5+\delta$ for the static model, with a significance level $5\%,$ where $\delta\in\{\pm 0.2,\pm 0.15,\pm 0.1\}$.
+Notes: Logit model with $\varepsilon_{it}$ standard logistic.
+Design 1: $\{\alpha_{k,i}^{0},\gamma_{k,t}^{0}\}\sim\mathcal{N}(0,0.01)$, $X_{it}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0}+\gamma_{1,t}^{0}+\gamma_{2,t}^{0},0.25)$, and $X_{i0}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0},0.25)$.
+The null hypothesis $H_{0}:(\beta_{0},\rho_{0})^{\prime}=(0.5+\delta,0.5+\delta)^{\prime}$ for the dynamic model, and $H_{0}:\beta_{0}=0.5+\delta$ for the static model, with a significance level $5\%,$ where $\delta\in\{\pm 0.2,\pm 0.15,\pm 0.1\}$.
+The number of replications is $1000$.
+Table B.3: Comparisons of Uncorrected and Corrected Logit Estimates, Design 2
+$(N,T)$ $(30,30)$ $(60,60)$ $(90,90)$
+Bias RMSE Bias RMSE Bias RMSE
+Static Model
+$\widehat{\beta}$ $0.454$ $1.201$ $0.100$ $0.132$ $0.060$ $0.071$
+$\widehat{\beta}_{L}$ $0.101$ $0.199$ $0.020$ $0.056$ $0.010$ $0.035$
+Dynamic Model
+$\widehat{\beta}$ $0.527$ $1.583$ $0.109$ $0.212$ $0.059$ $0.071$
+$\widehat{\beta}_{L}^{(1)}$ $0.117$ $0.204$ $0.024$ $0.058$ $0.009$ $0.038$
+$\widehat{\beta}_{L}^{(2)}$ $0.126$ $0.278$ $0.025$ $0.059$ $0.009$ $0.041$
+$\widehat{\rho}$ $-0.110$ $0.241$ $-0.055$ $0.110$ $-0.037$ $0.071$
+$\widehat{\rho}_{L}^{(1)}$ $-0.041$ $0.192$ $-0.017$ $0.090$ $-0.010$ $0.059$
+$\widehat{\rho}_{L}^{(2)}$ $-0.035$ $0.196$ $-0.012$ $0.091$ $-0.007$ $0.059$
+[…]
+Notes: Logit model with $\varepsilon_{it}$ standard logistic.
+Design 2: $\{\alpha_{k,i}^{0},\gamma_{k,t}^{0}\}\sim\mathcal{N}(0,1/4)$, $X_{it}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0}+\gamma_{1,t}^{0}+\gamma_{2,t}^{0},1)$, and $X_{i0}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0},1)$.
+For the static model, $\widehat{\beta}$ and $\widehat{\beta}_{L}$ are, respectively, the uncorrected and corrected estimates of $\beta_{0}=0.5$.
+For the dynamic model $(\widehat{\beta},\widehat{\rho})^{\prime}$ is the uncorrected estimate of $(\rho_{0},\beta_{0})^{\prime}=(0.5,0.5)^{\prime}$ from and $(\widehat{\beta}_{L}^{(\tau)},\widehat{\rho}_{L}^{(\tau)})^{\prime}$ is the corrected estimate with truncation parameter $\tau\in(1,2)$.
+The number of replications is $1000$.
+Table B.4: Empirical Size of LR, LM, and Wald Tests for Logit Model, Design 2
+(a) LR Tests
+Static Model Dynamic Model
+$(N,T)$ Uncorrected Corrected Uncorrected Corrected
+$(30,30)$ $47\%$ $20\%$ $43\%$ $18\%$
+$(60,60)$ $43\%$ $10\%$ $38\%$ $9\%$
+$(90,90)$ $42\%$ $9\%$ $34\%$ $7\%$
+(b) LM Tests
+Static Model Dynamic Model
+$(N,T)$ Uncorrected Corrected Uncorrected Corrected
+$(30,30)$ $42\%$ $18\%$ $38\%$ $16\%$
+$(60,60)$ $42\%$ $9\%$ $37\%$ $9\%$
+$(90,90)$ $41\%$ $9\%$ $34\%$ $7\%$
+(c) Wald Tests
+Static Model Dynamic Model
+$(N,T)$ Uncorrected Corrected Uncorrected Corrected
+$(30,30)$ $32\%$ $15\%$ $27\%$ $15\%$
+$(60,60)$ $41\%$ $9\%$ $36\%$ $8\%$
+$(90,90)$ $42\%$ $9\%$ $34\%$ $7\%$
+[…]
+Notes: Logit model with $\varepsilon_{it}$ standard logistic.
+Design 2: $\{\alpha_{k,i}^{0},\gamma_{k,t}^{0}\}\sim\mathcal{N}(0,1/4)$, $X_{it}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0}+\gamma_{1,t}^{0}+\gamma_{2,t}^{0},1)$, and $X_{i0}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0},1)$.
+The null hypothesis $H_{0}:(\beta_{0},\rho_{0})^{\prime}=(0.5,0.5)^{\prime}$ for the dynamic model, and $H_{0}:\beta_{0}=0.5$ for the static model, with a significance level $5\%$.
+The number of replications is $1000$.
+Table B.5: Empirical Power of LR, LM, and Wald Tests for Logit Model using $L(\theta)$, Design 2
+(a) LR Tests
+$\delta$
+$(N,T)$ $-0.20$ $-0.15$ $-0.10$ $0.10$ $0.15$ $0.20$
+Static Model
+$(30,30)$ $79\%$ $62\%$ $45\%$ $12\%$ $16\%$ $26\%$
+$(60,60)$ $100\%$ $95\%$ $73\%$ $40\%$ $75\%$ $93\%$
+$(90,90)$ $100\%$ $100\%$ $94\%$ $81\%$ $98\%$ $100\%$
+Dynamic Model
+$(30,30)$ $75\%$ $58\%$ $41\%$ $19\%$ $26\%$ $35\%$
+$(60,60)$ $100\%$ $94\%$ $69\%$ $44\%$ $80\%$ $96\%$
+$(90,90)$ $100\%$ $100\%$ $93\%$ $86\%$ $99\%$ $100\%$
+(b) LM Tests
+$\delta$
+$(N,T)$ $-0.20$ $-0.15$ $-0.10$ $0.10$ $0.15$ $0.20$
+Static Model
+$(30,30)$ $77\%$ $60\%$ $42\%$ $12\%$ $17\%$ $28\%$
+$(60,60)$ $99\%$ $95\%$ $72\%$ $41\%$ $77\%$ $93\%$
+$(90,90)$ $100\%$ $100\%$ $94\%$ $82\%$ $98\%$ $100\%$
+Dynamic Model
+$(30,30)$ $73\%$ $56\%$ $38\%$ $19\%$ $27\%$ $36\%$
+$(60,60)$ $100\%$ $93\%$ $68\%$ $45\%$ $80\%$ $94\%$
+$(90,90)$ $100\%$ $100\%$ $93\%$ $86\%$ $99\%$ $99\%$
+(c) Wald Tests
+$\delta$
+$(N,T)$ $-0.20$ $-0.15$ $-0.10$ $0.10$ $0.15$ $0.20$
+Static Model
+$(30,30)$ $75\%$ $57\%$ $39\%$ $10\%$ $17\%$ $29\%$
+$(60,60)$ $100\%$ $95\%$ $72\%$ $41\%$ $77\%$ $95\%$
+$(90,90)$ $100\%$ $100\%$ $94\%$ $81\%$ $98\%$ $100\%$
+Dynamic Model
+$(30,30)$ $69\%$ $50\%$ $34\%$ $18\%$ $26\%$ $36\%$
+$(60,60)$ $100\%$ $93\%$ $68\%$ $45\%$ $81\%$ $97\%$
+$(90,90)$ $100\%$ $100\%$ $92\%$ $86\%$ $99\%$ $100\%$
+[…]
+Notes: Logit model with $\varepsilon_{it}$ standard logistic.
+Design 2: $\{\alpha_{k,i}^{0},\gamma_{k,t}^{0}\}\sim\mathcal{N}(0,1/4)$, $X_{it}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0}+\gamma_{1,t}^{0}+\gamma_{2,t}^{0},1)$, and $X_{i0}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0},1)$.
+The null hypothesis $H_{0}:(\beta_{0},\rho_{0})^{\prime}=(0.5+\delta,0.5+\delta)^{\prime}$ for the dynamic model, and $H_{0}:\beta_{0}=0.5+\delta$ for the static model, with a significance level $5\%,$ where $\delta\in\{\pm 0.2,\pm 0.15,\pm 0.1\}$.
 The number of replications is $1000$.
 B.2 Probit Model
-Table B.3: Comparison between the Bias-corrected and Uncorrected MLE
+Table B.6: Comparisons of Uncorrected and Corrected Probit Estimates, Design 1
 $(N,T)$ $(30,30)$ $(60,60)$ $(90,90)$
@@
 […]
-Notes: The true structure parameters are $\theta_{0}=(\beta_{0},\rho_{0})^{\prime}=(0.5,0.5)^{\prime}$ for the dynamic model, and $\theta_{0}=\beta_{0}=0.5$ for the static model.
-In the static model, $\widehat{\beta}$ and $\widehat{\beta}_{L}$ represent the uncorrected MLE and our bias-corrected MLE, respectively.
-In the dynamic model, $(\widehat{\beta},\widehat{\rho})$ is the uncorrected MLE and $(\widehat{\beta}_{L}^{(\tau)},\widehat{\rho}_{L}^{(\tau)})^{\prime}$ represents our bias-corrected MLE with the truncation parameter $\tau$ set to $\tau=1$ or $2$.
+Notes: Probit model with $\varepsilon_{it}\sim\mathcal{N}(0,1)$.
+Design 1: $\{\alpha_{k,i}^{0},\gamma_{k,t}^{0}\}\sim\mathcal{N}(0,0.01)$, $X_{it}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0}+\gamma_{1,t}^{0}+\gamma_{2,t}^{0},0.25)$, and $X_{i0}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0},0.25)$.
+For the static model, $\widehat{\beta}$ and $\widehat{\beta}_{L}$ are, respectively, the uncorrected and corrected estimates of $\beta_{0}=0.5$.
+For the dynamic model $(\widehat{\beta},\widehat{\rho})^{\prime}$ is the uncorrected estimate of $(\rho_{0},\beta_{0})^{\prime}=(0.5,0.5)^{\prime}$ from and $(\widehat{\beta}_{L}^{(\tau)},\widehat{\rho}_{L}^{(\tau)})^{\prime}$ is the corrected estimate with truncation parameter $\tau\in(1,2)$.
 The number of replications is $1000$.
-Table B.4: Empirical Size of LR, LM, and Wald Tests Based on Bias-corrected and Uncorrected Likelihoods for Probit Model
+Table B.7: Empirical Size of LR, LM, and Wald Tests Based on Corrected and Uncorrected Likelihoods for Probit Model, Design 1
 (a) LR Tests
@@
 […]
-Notes: The null hypothesis $H_{0}:(\beta_{0},\rho_{0})^{\prime}=(0.5,0.5)^{\prime}$ for the dynamic model, and $H_{0}:\beta_{0}=0.5$ for the static model, with a significance level $5\%$.
+Notes: Probit model with $\varepsilon_{it}\sim\mathcal{N}(0,1)$.
+Design 1: $\{\alpha_{k,i}^{0},\gamma_{k,t}^{0}\}\sim\mathcal{N}(0,0.01)$, $X_{it}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0}+\gamma_{1,t}^{0}+\gamma_{2,t}^{0},0.25)$, and $X_{i0}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0},0.25)$.
+The null hypothesis $H_{0}:(\beta_{0},\rho_{0})^{\prime}=(0.5,0.5)^{\prime}$ for the dynamic model, and $H_{0}:\beta_{0}=0.5$ for the static model, with a significance level $5\%$.
 The number of replications is $1000$.
-Table B.5: Empirical Power of LR, LM, and Wald Tests Based on Bias-corrected Likelihoods for Probit Model
+Table B.8: Empirical Power of LR, LM, and Wald Tests Based on Bias-corrected Likelihoods for Probit Model, Design 1
 (a) LR Tests
@@
 […]
-Notes: The null hypothesis $H_{0}:(\beta_{0},\rho_{0})^{\prime}=(0.5+\delta,0.5+\delta)^{\prime}$ for the dynamic model, and $H_{0}:\beta_{0}=0.5+\delta$ for the static model, with a significance level $5\%,$ where $\delta\in\{\pm 0.2,\pm 0.15,\pm 0.1\}$.
+Notes: Probit model with $\varepsilon_{it}\sim\mathcal{N}(0,1)$.
+Design 1: $\{\alpha_{k,i}^{0},\gamma_{k,t}^{0}\}\sim\mathcal{N}(0,0.01)$, $X_{it}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0}+\gamma_{1,t}^{0}+\gamma_{2,t}^{0},0.25)$, and $X_{i0}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0},0.25)$.
+The null hypothesis $H_{0}:(\beta_{0},\rho_{0})^{\prime}=(0.5+\delta,0.5+\delta)^{\prime}$ for the dynamic model, and $H_{0}:\beta_{0}=0.5+\delta$ for the static model, with a significance level $5\%,$ where $\delta\in\{\pm 0.2,\pm 0.15,\pm 0.1\}$.
+The number of replications is $1000$.
+Table B.9: Empirical Size of LM and Wald Tests Based on Corrected and Uncorrected Likelihoods for Probit Model, Design 2
+(a) LM Tests
+Static Model Dynamic Model
+$(N,T)$ Uncorrected Corrected Uncorrected Corrected
+$(30,30)$ $69\%$ $34\%$ $59\%$ $30\%$
+$(60,60)$ $74\%$ $18\%$ $62\%$ $17\%$
+$(90,90)$ $71\%$ $14\%$ $60\%$ $13\%$
+(b) Wald Tests
+Static Model Dynamic Model
+$(N,T)$ Uncorrected Corrected Uncorrected Corrected
+$(30,30)$ $37\%$ $29\%$ $25\%$ $26\%$
+$(60,60)$ $71\%$ $18\%$ $56\%$ $16\%$
+$(90,90)$ $72\%$ $14\%$ $60\%$ $13\%$
+[…]
+Notes: Probit model with $\varepsilon_{it}\sim\mathcal{N}(0,1)$.
+Design 2: $\{\alpha_{k,i}^{0},\gamma_{k,t}^{0}\}\sim\mathcal{N}(0,1/4)$, $X_{it}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0}+\gamma_{1,t}^{0}+\gamma_{2,t}^{0},1)$, and $X_{i0}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0},1)$.
+The null hypothesis $H_{0}:(\beta_{0},\rho_{0})^{\prime}=(0.5,0.5)^{\prime}$ for the dynamic model, and $H_{0}:\beta_{0}=0.5$ for the static model, with a significance level $5\%$.
+The number of replications is $1000$.
+Table B.10: Empirical Power of LR, LM, and Wald Tests Based on Bias-corrected Likelihoods for Probit Model, Design 2
+(a) LR Tests
+$\delta$
+$(N,T)$ $-0.20$ $-0.15$ $-0.10$ $0.10$ $0.15$ $0.20$
+Static Model
+$(30,30)$ $97\%$ $90\%$ $77\%$ $14\%$ $13\%$ $15\%$
+$(60,60)$ $100\%$ $100\%$ $96\%$ $34\%$ $70\%$ $83\%$
+$(90,90)$ $100\%$ $100\%$ $99\%$ $85\%$ $98\%$ $96\%$
+Dynamic Model
+$(30,30)$ $95\%$ $86\%$ $74\%$ $25\%$ $28\%$ $38\%$
+$(60,60)$ $100\%$ $100\%$ $93\%$ $47\%$ $84\%$ $96\%$
+$(90,90)$ $100\%$ $100\%$ $100\%$ $90\%$ $99\%$ $100\%$
+(b) LM Tests
+$\delta$
+$(N,T)$ $-0.20$ $-0.15$ $-0.10$ $0.10$ $0.15$ $0.20$
+Static Model
+$(30,30)$ $96\%$ $87\%$ $73\%$ $12\%$ $13\%$ $17\%$
+$(60,60)$ $100\%$ $100\%$ $96\%$ $37\%$ $69\%$ $76\%$
+$(90,90)$ $100\%$ $100\%$ $99\%$ $86\%$ $91\%$ $88\%$
+Dynamic Model
+$(30,30)$ $94\%$ $83\%$ $70\%$ $21\%$ $27\%$ $38\%$
+$(60,60)$ $100\%$ $99\%$ $92\%$ $48\%$ $81\%$ $91\%$
+$(90,90)$ $100\%$ $99\%$ $99\%$ $90\%$ $94\%$ $94\%$
+(c) Wald Tests
+$\delta$
+$(N,T)$ $-0.20$ $-0.15$ $-0.10$ $0.10$ $0.15$ $0.20$
+Static Model
+$(30,30)$ $91\%$ $83\%$ $68\%$ $10\%$ $13\%$ $19\%$
+$(60,60)$ $99\%$ $99\%$ $95\%$ $39\%$ $76\%$ $92\%$
+$(90,90)$ $99\%$ $99\%$ $99\%$ $88\%$ $99\%$ $99\%$
+Dynamic Model
+$(30,30)$ $88\%$ $75\%$ $60\%$ $21\%$ $27\%$ $39\%$
+$(60,60)$ $100\%$ $99\%$ $91\%$ $49\%$ $85\%$ $98\%$
+$(90,90)$ $99\%$ $99\%$ $99\%$ $91\%$ $99\%$ $100\%$
+[…]
+Notes: Probit model with $\varepsilon_{it}\sim\mathcal{N}(0,1)$.
+Design 2: $\{\alpha_{k,i}^{0},\gamma_{k,t}^{0}\}\sim\mathcal{N}(0,1/4)$, $X_{it}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0}+\gamma_{1,t}^{0}+\gamma_{2,t}^{0},1)$, and $X_{i0}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0},1)$.
+The null hypothesis $H_{0}:(\beta_{0},\rho_{0})^{\prime}=(0.5+\delta,0.5+\delta)^{\prime}$ for the dynamic model, and $H_{0}:\beta_{0}=0.5+\delta$ for the static model, with a significance level $5\%,$ where $\delta\in\{\pm 0.2,\pm 0.15,\pm 0.1\}$.
 The number of replications is $1000$.

### v5

@@
 The resulting estimator follows an asymptotic normal distribution with mean zero.
-Moreover, likelihood-based tests statistics—including likelihood-ratio, Lagrange-multiplier, and Wald tests—follow the limiting chi-squared distribution under the null hypothesis.
+Moreover, likelihood-based test statistics—including likelihood-ratio, Lagrange-multiplier, and Wald tests—follow the limiting chi-squared distribution under the null hypothesis.
 Simulations demonstrate the effectiveness of the proposed correction method, and an empirical application on the labor force participation of single mothers underscores its practical importance.
@@
 1 Introduction
-Panel data is common in empirical research.
+Panel data are common in empirical research.
 Panel data models with two-way fixed effects are widely used to control for unobserved heterogeneity that may correlate with covariates.
@@
 Traditionally, these models address level heterogeneity by incorporating fixed effects only in the intercept, adjusting for baseline outcome differences across individuals and time periods.
-However, many empirical applications reveal substantial variation in how outcomes respond to covariates (Browning and Carro, 2007).
-When such response heterogeneity are present and correlated with covariates, failing to account for them can lead to biased estimates and invalid inference (Su and Chen, 2013).
+However, both economic theory and empirical evidence indicate substantial variation in how outcomes respond to covariates across individuals and over time (Browning and Carro, 2007).
+Cross-sectional response heterogeneity reflects individual-specific factors such as preferences, productivity, and access to resources (Heckman, 2001), whereas time-varying response heterogeneity captures changes in behavior driven by evolving circumstances—including policy reforms, business cycles, and technological progress (Okui and Wang, 2021).
+When such response heterogeneities are correlated with covariates, failing to account for them can lead to biased estimates and invalid inference (Su and Chen, 2013).
 To illustrate, consider the classic problem of estimating the impact of children on a mother’s labor force participation.
-Research indicates that mothers with more children may differ systematically from those with fewer or none (Blau and Robins 1988; Angrist and Evans 1998; Kleven et al.
-2019).
+Research indicates that mothers with more children may differ systematically from those with fewer or none (Blau and Robins, 1988; Angrist and Evans, 1998; Kleven et al., 2019).
 First, mothers who choose to have more children might be inherently less inclined to work.
-Second, mothers with more children may be those whose labor supply is less sensitive to family size, perhaps due to better financial resources or access to childcare.
-In a panel dataset tracking households over time, we can control for the first type of level heterogeneity by including fixed effects in the intercept.
-However, to address the second type of response heterogeneity, we must include fixed effects in the slope.
+Second, mothers with more children may be those whose labor supply is less sensitive to family size, perhaps due to better financial resources or childcare access.
+Moreover, over time, declining fertility and rising female labor force participation suggest a shifting baseline, while advances in home technology and welfare reforms may lessen the burden of additional children, enhancing labor supply responsiveness.
+In a panel dataset tracking households over time, we can control for the first type of heterogeneity—level differences in baseline participation—by including fixed effects in the intercept.
+However, to address the second type of heterogeneity—variations in response to children across individuals and over time—we must include fixed effects in the slope.
 Given that labor force participation is a binary outcome and typically exhibits serial correlation, this example underscores the need for a dynamic nonlinear panel data model with fixed effects in both the intercept and slope.
-In this paper, we introduce a general class of dynamic nonlinear heterogeneous parameter models (DN-HP) that incorporate individual and time fixed effects in both the intercept and slope.
-This flexible framework accommodates multi-dimensional heterogeneities and encompasses many commonly used panel data models.
-DN-HP models are generally subject to the incidental parameter problem, in that the limiting distribution of the point estimator is not centered at 00, and that the test statistics do not follow their standard asymptotic distributions (of $\chi^{2}$, for many test statistics).
-To address the incidental parameter problem, we propose an analytical bias correction approach.
-This method restores the standard asymptotic distributions for the point estimator and the test statistics.
+In this paper, we introduce a general class of dynamic nonlinear heterogeneous-parameter (DN-HP) models that incorporate individual and time fixed effects in both the intercept and the slope.
+This flexible framework accommodates multi-dimensional heterogeneities and encompasses many commonly used panel-data models, including both static and dynamic specifications for linear and limited-dependent-variable outcomes.
+It nests individual-specific slope models that capture heterogeneous responses across individuals, as well as time-varying-coefficient models that allow response parameters to evolve over time due to shifts in the underlying economic or structural environment.
+DN-HP models are generally subject to the incidental-parameter problem: the limiting distribution of the estimator is not centered at zero, and standard test statistics (e.g., Wald, LM, and LR) no longer follow their conventional asymptotic $\chi^{2}$ distributions.
+To address this, we propose an analytical bias-correction procedure that restores valid large-sample inference for both parameter estimates and test statistics.
 Figure 1: Comparison of Uncorrected and Corrected Maximum Likelihood Estimators
@@
 Notes: Model.
-$Y_{it}=\mathds{1}\{\rho Y_{it-1}+(\beta+\alpha_{1,i}+\gamma_{1,t})X_{it}+\alpha_{2,i}+\gamma_{2,t}+\varepsilon_{it}>0\}$ where $\mathds{1}\{\cdot\}$ denotes the indicator function, $\varepsilon_{it}$ is standard logistic independent of $X_{it}$, and the true values $\rho_{0}=\beta_{0}=0.5$.
+$Y_{it}=\mathds{1}\{\rho Y_{it-1}+(\beta+\alpha_{1,i}+\gamma_{1,t})Z_{it}+\alpha_{2,i}+\gamma_{2,t}+\varepsilon_{it}>0\}$ where $\mathds{1}\{\cdot\}$ denotes the indicator function, $\varepsilon_{it}$ is standard logistic independent of the exogenous regressor $Z_{it}$, and the true values $\rho_{0}=\beta_{0}=0.5$.
 $\widehat{\beta}$ and $\widehat{\rho}$ are the uncorrected estimators while $\widehat{\beta}_{L}$ and $\widehat{\rho}_{L}$ are the corrected estimators.
@@
 The detailed data-generating process is given in Section 4.
-As a preview, Figure 1 presents simulated boxplots of the estimation bias in a DN-HP logit model, comparing estimators from our bias correction procedure to the uncorrected ones.
-The results reveal that our bias correction procedure significantly reduces the estimation bias while maintaining comparable mean square errors.
+As a preview, Figure 1 presents simulated boxplots of the estimation bias in a two-way heterogeneous parameter logit model, comparing estimators from our bias correction procedure to the uncorrected ones.
+The results reveal that our bias correction procedure significantly reduces the estimation bias while maintaining comparable mean squared errors.
 Literature Review.
-The estimation and inference of fixed effects models in the presence of the incidental parameter problem have been extensively studied.
-Early work, such as Chamberlain (1980), Arellano and Bond (1991), and Lancaster (2002), focused on frameworks with only individual effects in the intercept, addressing fixed-$T$ consistency (short panels) for structural parameter estimators in specific models.
-In recent decades, the availability of long panel datasets has motivated a large-$N,T$ framework, where the cross-sectional size $N$ and the time dimension $T$ grow proportionally to infinity.
-Under this rectangular asymptotic, estimators remain consistent but typically exhibit a non-negligible bias of order $O(1/T)$ due to the presence of individual effects in the intercept.
-To enable valid inference, researchers have developed various bias correction methods for a broad range of models in a maximum likelihood framework.
-For instance, Hahn and Newey (2004) and Hahn and Kuersteiner (2011) propose “parameter-based” bias correction methods to eliminate the bias from the maximum likelihood estimator.
-Two alternative approaches have also been explored: “score-based” and “likelihood-based” methods.
-Woutersen (2002) and Li et al.
-(2003) modify the profiled score to remove the bias, while Bester and Hansen (2009) and Arellano and Hahn (2016) identify and correct the bias directly for the likelihood.
-Most of these methods are analytical, estimating the bias using explicit formulas.
-Alternatively, numerical procedures are available, estimating the bias through methods such as the Jackknife (Dhaene and Jochmans, 2015), the bootstrap (Bellio et al., 2023), and integrated likelihoods (Arellano and Bonhomme, 2009).
-For two-way fixed effects models, incorporating time effects in the intercept generates an additional bias of $O(1/N)$.
-Moon and Weidner (2017) propose a parameter-based analytical correction for dynamic linear models with two-way interactive effects.
-For nonlinear models, Fernández-Val and Weidner (2016) develop parameter-based bias correction techniques with additive fixed effects, while Chen et al.
-(2021) focus on interactive fixed effects in static models.
-Alternatively, Jochmans and Otsu (2019) propose a likelihood-based analytical bias correction for static models, allowing fixed effects to enter the model in an arbitrary but known form.
-Other model-specific solutions include Bai (2009) and Charbonneau (2017), among others.
-For an extensive review of the incidental parameter problem for models with two-way fixed effects in the intercept, see Fernández-Val and Weidner (2018).
-To our knowledge, our paper represents one of the first efforts to extend intercept-only two-way fixed effects models into multi-dimensional two-way fixed effects models, accounting for both level and response heterogeneities.
-Two prior studies in this line of research are Keane and Neal (2020) and Lu and Su (2023).
-Both examine linear heterogeneous parameter models with additive effects in both the intercept and slope.
-Keane and Neal (2020) propose an iterative procedure, which they refer to as the “mean observation OLS,” to eliminate the bias, while Lu and Su (2023) rely on a parameter-based Jackknife approach and provide a uniform inference theory for structural parameters.
-Our study significantly expands the model framework and introduces a likelihood-based analytical bias correction technique to ensure valid estimation and inference.
-Neither our DN-HP models nor our bias correction procedure have been discussed in the current literature.
+The estimation and inference of fixed-effects models in the presence of the incidental-parameter problem have been extensively studied.
+Early work, such as Chamberlain (1980), Arellano and Bond (1991), and Lancaster (2002), focused on frameworks with only individual effects in the intercept, establishing fixed-$T$ consistency (short panels) for structural parameter estimators in specific models.
+In recent decades, the availability of long panel datasets has motivated a large–$N,T$ framework, where the cross-sectional size $N$ and the time dimension $T$ grow at similar rates under rectangular asymptotics.
+In this setting, estimators remain consistent but exhibit a non-negligible bias of order $O\left(1/T\right)$ because each individual effect is estimated from only $T$ observations.
+This incidental-parameter bias is especially pronounced in nonlinear or dynamic models.
+To address it, researchers have developed a variety of bias correction techniques within a maximum likelihood framework.
+Hahn and Newey (2004) and Hahn and Kuersteiner (2011) propose “parameter-based” corrections that remove the bias from the likelihood estimator; Woutersen (2002) and Li et al.
+(2003) propose “score-based” methods that modify the profiled score; while “likelihood-based” approaches such as Bester and Hansen (2009) and Arellano and Hahn (2016) adjust the log-likelihood directly.
+Most of these procedures are analytical, relying on closed-form approximations to the bias.
+Alternatively, numerical corrections estimate the bias through resampling or re-estimation, including the jackknife (Dhaene and Jochmans, 2015), the bootstrap (Bellio et al., 2023; Higgins and Jochmans, 2024), and integrated-likelihood methods (Arellano and Bonhomme, 2009).
+For two-way fixed effects models, incorporating time effects in the intercept generates an additional bias of order $O\left(1/N\right)$ on top of the $O\left(1/T\right)$ term from individual effects.
+Several studies extend bias-correction methods to this setting.
+Moon and Weidner (2017) develop a parameter-based analytical correction for dynamic linear models with interactive effects.
+For nonlinear models, Fernández-Val and Weidner (2016) propose parameter-based techniques for additive fixed effects, while Chen et al.
+(2021) study interactive effects in static models.
+Alternatively, Jochmans and Otsu (2019) provide a likelihood-based approach for static models with an arbitrary but known fixed-effect structure.
+Other model-specific contributions include Bai (2009) and Charbonneau (2017).
+For a comprehensive overview of these developments, see Fernández-Val and Weidner (2018).
+Despite this progress, the two-way literature primarily addresses heterogeneity in levels—that is, intercept effects—rather than heterogeneity in responses (slopes).
+Extending bias-correction methods to settings with slope heterogeneity remains largely unexplored and is the focus of our contribution.
+A substantial related literature explores heterogeneity in slope coefficients within panel models, emphasizing individual-specific parameters (Robertson and Symons, 1992; Pesaran and Smith, 1995) and time-varying coefficients (Robinson, 1989; Stock and Watson, 1996), most often in linear or static contexts (Hsiao, 2014; Pesaran, 2015).
+Recent research in dynamic nonlinear panels has advanced these ideas further.
+Chernozhukov et al.
+(2013) derive partial identification with uniform inference for nonseparable panels under time-homogeneity.
+Browning and Carro (2014) provide mixture-based point identification for dynamic binary models with maximal cross-sectional heterogeneity, while Browning and Carro (2010) analyze short-panel dynamic binary models with heterogeneous transition probabilities and propose a mean-integrated-MSE (MIMSE) estimator that better balances bias–variance trade-offs in small–$T$ settings than analytical bias correction.
+While these studies focus on identification or fixed–$T$ settings within semiparametric or nonparametric frameworks, we adopt a parametric, likelihood-based approach that achieves point identification and enables bias-corrected inference for dynamic nonlinear models with two-way slope and intercept effects under large–$N,T$ asymptotics.
+Two recent studies are closely aligned with our framework.
+Keane and Neal (2020) examine linear models with two-way slope and intercept effects, employing an iterative “mean-observation OLS” estimator to address bias.
+Their analysis of U.S.
+agricultural data reveals pronounced regional differences in heat-yield sensitivity, alongside temporal adaptation as farmers adopt new technologies over time.
+Lu and Su (2023) study a similar two-way heterogeneous-slope specification and implement a parameter-based jackknife estimator that enables uniform inference for structural parameters.
+Their cross-country analysis of the Feldstein–Horioka relation uncovers variation across nations and periods, driven by financial integration and evolving policy regimes.
+Both studies underscore the importance of slope heterogeneity across individuals and time but remain confined to linear settings.
+Our study advances this literature by extending the framework to a dynamic nonlinear context and developing a likelihood-based analytical bias-correction procedure that ensures valid estimation and inference.
+To our knowledge, neither our DN-HP model nor its associated bias correction has been previously explored.
+Together, the model and method provide a unified approach to estimation and inference in dynamic panels with multi-dimensional heterogeneity.
 This paper.
@@
 We show that the corrected log-likelihood function does not suffer from the incidental parameter problem, under appropriate regularity conditions.
+Our method can be viewed as an extension of Arellano and Hahn (2016) and Bester and Hansen (2009) to the two-way DN-HP models.
 One key benefit of our procedure is that we achieve bias corrections for the point estimators and the test statistics by modifying a single object, the log-likelihood function.
 We rigorously show that the estimators obtained by maximizing the corrected log-likelihood function retains a limiting normal distribution with zero mean, consistent with the asymptotic theory of the classical maximum likelihood estimators (MLE) in the absence of incidental parameters.
-Beyond the estimators, we also show that the likelihood ratio (LR), Lagrange multiplier (LM), and Wald statistics, derived from the same corrected log-likelihood function, are asymptotically equivalent, sharing the same asymptotic $\chi^{2}$ distribution under the null hypothesis.
-We demonstrate the finite-sample performance of our method through Monte Carlo simulations, comparing it to both the original likelihood and the Jackknife, which is a potential bias correction device [^1] It is worth mentioning that, to our knowledge, there currently has been no study on using the jackknife under the DN-HP models..
+Beyond the estimators, we also show that the likelihood ratio (LR), Lagrange-multiplier (LM), and Wald statistics, derived from the same corrected log-likelihood function, are asymptotically equivalent, sharing the same asymptotic $\chi^{2}$ distribution under the null hypothesis.
+We demonstrate the finite-sample performance of our method through Monte Carlo simulations, comparing it to the original likelihood and potential bias correction devices of the bootstrap and the jackknife.
 We show that our correction procedure reduces the bias significantly, without increasing the root mean squared errors, and restores the test sizes of the LR test to its nominal level.
-We also find that our procedure outperforms the Jackknife.
-In the empirical application, we apply the DN-HP logit model to study the aforementioned question of how the number of children affects a mother’s labor force participation, with a focus on single-mother households.
+We find that our procedure outperforms the jackknife when it comes to heterogeneous slope coefficients.
+We also find that the LR test based on our approach has a stronger power near the true hypothesis than the LR test based on the bootstrap.
+In the empirical application, we apply our approach with a probit model to study the aforementioned question of how the number of children affects a mother’s labor force participation, with a focus on single-mother households.
 Our analysis reveals that this impact varies significantly across individuals and over time.
@@
 Section 2 explains our settings, the identification restriction, and the incidental parameter problem of DN-HP models.
-Section 3 describes our bias correction procedure and provides relevant statistical properties, including those we claim above.
-Section 4 presents simulation studies to demonstrate the performance of method.
+Section 3 describes our bias correction procedure and provides relevant statistical properties.
+Section 4 presents simulation studies to demonstrate the performance of the method.
 Section 5 applies our method to study single mothers’ labor force participation.
 Finally in Section 6, we leave some closing remarks.
-Proofs, technical details, and additional results are provided in the appendix.
+Proofs, technical details, additional results, elaborated discussions, etc.
+are provided in the appendix.
+These items have references starting with letters.
 Notation.
@@
 For each individual $i$, the response $Y_{it}$ is generated sequentially over $t$, by
-$Y_{it}|X_{i1},\ldots,X_{it},\phi^{0}\sim f(y|X_{it},\theta_{0}+\alpha_{i}^{0}+\gamma_{t}^{0}),$ (2.1)
-where $\theta_{0}$ is a $K\times 1$ vector of (non-random) unknown common parameters of interests, $\phi^{0}:=(\alpha_{1}^{0\prime},\ldots,\alpha_{N}^{0\prime},\gamma_{1}^{0\prime},\ldots,\gamma_{T}^{0\prime})^{\prime}$ consists of the $K\times 1$ vectors of random variables $\alpha_{i}^{0}$ and $\gamma_{t}^{0}$, and $f(y|\cdot)$ is a density known up to $\theta_{0}$, $\alpha_{i}^{0}$, and $\gamma_{t}^{0}$.
+$Y_{it}|X_{it},\ldots,X_{i1},\phi^{0}\sim f(y|X_{it},\theta_{0}+\alpha_{i}^{0}+\gamma_{t}^{0}),$ (2.1)
+where $\theta_{0}$ is a $K\times 1$ vector of (non-random) unknown structural parameters of interests, $\phi^{0}:=(\alpha_{1}^{0\prime},\ldots,\alpha_{N}^{0\prime},\gamma_{1}^{0\prime},\ldots,\gamma_{T}^{0\prime})^{\prime}$ consists of the $K\times 1$ random vectors $\alpha_{i}^{0}$ and $\gamma_{t}^{0}$, and $f(y|\cdot)$ is a density known up to $\theta_{0}$, $\alpha_{i}^{0}$, and $\gamma_{t}^{0}$.
 We restrict our attention to the case where $f\left(\cdot\right)$ depends on $\theta_{0}$, $\alpha_{i}^{0}$, and $\gamma_{t}^{0}$ through the additive structure $\theta_{0}+\alpha_{i}^{0}+\gamma_{t}^{0}$, the “heterogeneous parameter”.
@@
 Generally, $\alpha_{i}^{0}$ and $\gamma_{t}^{0}$ may be correlated with the regressor $X_{it}$.
-However, throughout the paper, we employ the “fixed-effect” framework to condition on the realizations of $\alpha_{i}^{0}$ and $\gamma_{t}^{0}$ (denoted by the same symbols) in the density, and treat them as unknown nuisance parameters to be estimated together with $\theta_{0}$.
+Throughout the paper, we employ the “fixed-effect” framework to condition on the realizations of $\alpha_{i}^{0}$ and $\gamma_{t}^{0}$ (denoted by the same symbols) in the density, and treat them as unknown nuisance parameters to be estimated together with $\theta_{0}$.
 Conditioning on $\alpha_{i}^{0}$ and $\gamma_{t}^{0}$, we assume $(Y_{it},X_{it}^{\prime})^{\prime}$ independent across $i$ but serially dependent over $t$.
@@
 For instance, $X_{it}$ may contain lagged values of $Y_{it}$.
-In this paper, we focus on the case where only heterogeneous parameters appear in (2.1).
-This is without loss of generality.
-As illustrated in the example below, our approach accommodates settings where both heterogeneous parameters and “homogeneous parameters” are present.
-[^2] Another example is the classic Neyman-Scott model, where $Y_{it}$ follows a normal distribution with a homogeneous variance but heterogeneous means.
-See Example 3 below.
-Example 1 (heterogeneous parameter model)
-Consider a probit model with data generating process (DGP)
-$Y_{it}|X_{i1},\ldots,X_{it},\phi^{0}=\mathds{1}\{\pi(X_{it},\theta_{0},\alpha_{i}^{0},\gamma_{t}^{0})+\varepsilon_{it}>0\},$ (2.2)
-where $\varepsilon_{it}$ is a standard normal random variable independent of $X_{it}$, $\theta_{0}$, $\alpha_{i}^{0}$, and $\gamma_{t}^{0}$.
-For arguments $\theta$, $\alpha_{i}$, and $\gamma_{t}$ of dimension $K\times 1$, we allow the linear index [^3] We use a model with a linear index solely for the purpose of illustration.
-Our method applies to general models.
-$\pi(X_{it},\theta,\alpha_{i},\gamma_{t})$ to take the following general form:
-$\pi(X_{it},\theta,\alpha_{i},\gamma_{t})=\sum_{k=1}^{K}(\theta_{k}+\alpha_{k,i}+\gamma_{k,t})X_{k,it},$
-where $X_{k,it}$, $\theta_{k}$, $\alpha_{k,i}$, and $\gamma_{k,t}$ are the $k$-th components of $X_{it}$, $\theta$, $\alpha_{i}$, and $\gamma_{t}$, respectively, $k=1,\ldots,K$.
-In addition, when $X_{1,it}=1$, the DGP in (2.2) consists of the traditional individual- and time-effects in the intercept.
-When $X_{k,it}=Y_{it-1}$ for some $k$, the DGP is dynamic because each $Y_{it}$ depends on its past value.
-Moreover, we also allow a subset of the coefficients to be homogeneous, thereby accommodating both heterogeneous and homogeneous parameters.
-For instance, when it is known that $\alpha_{1,i}^{0}=\gamma_{1,t}^{0}=0$, the model can be estimated under constraints $\alpha_{1,i}=\gamma_{1,t}=0$ for all $(i,t)$.
-In this case, the linear index is effectively
-$\pi(X_{it},\theta,\alpha_{i},\gamma_{t})=\theta_{1}X_{1,it}+\sum_{k=2}^{K}(\theta_{k}+\alpha_{k,i}+\gamma_{k,t})X_{k,it},$
-where $X_{1,it}$ is associated with a “homogeneous parameter” $\theta_{1}$ while all other $X_{k,it}$ ($k=2,\ldots,K$) are associated with heterogeneous parameters $(\theta_{k}+\alpha_{k,i}+\gamma_{k,t})$.
-In a similar fashion, our approach also applies to, e.g.,
-$\pi(X_{it},\theta,\alpha_{i},\gamma_{t})=(\alpha_{1,i}+\gamma_{1,t})X_{1,it}+\sum_{k=2}^{K}(\theta_{k}+\alpha_{k,i}+\gamma_{k,t})X_{k,it},$
-with the constraint $\theta_{1}=0$. Other variants of $\pi(X_{it},\theta,\alpha_{i},\gamma_{t})$ include:
-$\pi(X_{it},\theta,\alpha_{i},\gamma_{t})$ $=(\alpha_{1,i}+\gamma_{1,t})+\theta_{2}Y_{it-1}+\sum_{k=3}^{K}\theta_{k}X_{k,it},$ (2.3)
-$\pi(X_{it},\theta,\alpha_{i},\gamma_{t})$ $=(\theta_{1}+\alpha_{1,i}+\gamma_{1,t})+\sum_{k=2}^{K}(\theta_{k}+\alpha_{k,i}+\gamma_{k,t})X_{k,it},$ (2.4)
-$\pi(X_{it},\theta,\alpha_{i},\gamma_{t})$ $=(\theta_{1}+\alpha_{1,i}+\gamma_{1,t})+(\theta_{2}+\alpha_{2,i}+\gamma_{2,t})Y_{it-1}$
-$\quad+\sum_{k=3}^{K_{1}}(\theta_{k}+\alpha_{k,i}+\gamma_{k,t})X_{k,it}+\sum_{k=K_{1}+1}^{K}\theta_{k}X_{k,it}$ (2.5)
-for some $3\leq K_{1}<K$.
-To our knowledge, except (2.3) considered by Fernández-Val and Weidner (2016), none of the above specifications are fully addressed under nonlinear models with predetermined regressors.
-The specification in (2.4) is the same to Keane and Neal (2020) and Lu and Su (2023), except that they consider only the linear regression model.
-Jochmans and Otsu (2019) provide a bias correction technique for general nonlinear models with (2.4) only for strictly exogenous $X_{k,it}$.
-Currently, no methods are available for nonlinear models with (2.5).
-We are interested in the estimation of and the inference about $\theta_{0}$ in (2.1), under the presence of the nuisance parameter $\phi^{0}$.
+In this paper, we focus on the case where all parameters in (2.1) are heterogeneous.
+This is without loss of generality, because a “homogeneous parameter” can be accommodated by setting the relevant component of $\alpha_{i}^{0}+\gamma_{t}^{0}$ to 0 for all $\left(i,t\right)$.
+Our specification accommodates a general class of models that is not yet covered in the current literature.
+We motivate our specification in Example A.1 in Section A of the supplementary appendix.
+We are interested in the estimation of and the inference about the structural parameter $\theta_{0}$ in (2.1), under the presence of the nuisance parameter $\phi^{0}$.
 Our approach is likelihood-based.
@@
 where $D_{1}:=(\mathbb{I}_{N-1},-\iota_{N-1})\otimes\mathbb{I}_{K}$ and $D_{2}:=(\mathbb{I}_{T-1},-\iota_{T-1})\otimes\mathbb{I}_{K}$.
-This rules out the indeterminancy in (2.7) without explicitly using the Lagrangian.
-We use the following example to illustrate the reparameterization.
-Example 2 (reparameterization)
-For some scalar structural parameter $\mu_{0}$ of interests and $K=1$, consider $Y_{it}=\mu_{0}+\alpha_{i}^{0}+\gamma_{t}^{0}+\varepsilon_{it}$ where $\varepsilon_{it}$ is standard normal independent across both $i$ and $t$, and independent of $\alpha_{i}^{0}$ and $\gamma_{t}^{0}$.
-We restrict the variance of $\varepsilon_{it}$ to be $1$ because we use this example only to demonstrate the identification of $\mu_{0}$.
-In Example 3 below, we indeed allow $\varepsilon_{it}$ to have an unknown variance.
-Conditioning on $\phi^{0}$, the original likelihood function is
-$\ell(\mu,\phi)$ $=-\frac{1}{2}\log(2\pi)-\frac{1}{NT}\sum_{i=1}^{N-1}\sum_{t=1}^{T-1}\frac{1}{2}(Y_{it}-\mu-\alpha_{i}-\gamma_{t})^{2}-\frac{1}{NT}\sum_{t=1}^{T-1}\frac{1}{2}(Y_{Nt}-\mu-\alpha_{N}-\gamma_{t})^{2}$
-$\quad-\frac{1}{NT}\sum_{i=1}^{N-1}\frac{1}{2}(Y_{iT}-\mu-\alpha_{i}-\gamma_{T})^{2}-\frac{1}{NT}\frac{1}{2}(Y_{NT}-\mu-\alpha_{N}-\gamma_{T})^{2}.$
-Setting $\alpha_{N}=-\sum_{i=1}^{N-1}\alpha_{i}$ and $\gamma_{T}=-\sum_{t=1}^{T-1}\gamma_{t}$, the reparameterized likelihood is
-$l(\mu,\psi)$ $=-\frac{1}{2}\log(2\pi)-\frac{1}{NT}\sum_{i=1}^{N-1}\sum_{t=1}^{T-1}\frac{1}{2}(Y_{it}-\mu-\alpha_{i}-\gamma_{t})^{2}-\frac{1}{NT}\sum_{t=1}^{T-1}\frac{1}{2}\left(Y_{Nt}-\mu+\sum_{i=1}^{N-1}\alpha_{i}-\gamma_{t}\right)^{2}$
-$\quad-\frac{1}{NT}\sum_{i=1}^{N-1}\frac{1}{2}\left(Y_{iT}-\mu-\alpha_{i}+\sum_{t=1}^{T-1}\gamma_{t}\right)^{2}-\frac{1}{NT}\frac{1}{2}\left(Y_{NT}-\mu+\sum_{i=1}^{N-1}\alpha_{i}+\sum_{t=1}^{T-1}\gamma_{t}\right)^{2}.$
-Calculating the first derivative of $l(\mu,\psi)$ w.r.t. $\alpha_{i}$, we have
-$\partial_{\alpha_{i}}l(\mu,\psi)=\frac{1}{NT}\sum_{t=1}^{T}(Y_{it}-\mu-\alpha_{i})-\frac{1}{NT}\sum_{t=1}^{T}\left(Y_{Nt}-\mu+\sum_{i=1}^{N-1}\alpha_{i}\right)$
-for every $i=1,\ldots,N-1$. Equating $\partial_{\alpha_{i}}l(\mu,\psi)=0$, we obtain
-$\alpha_{i}=\frac{1}{T}\sum_{t=1}^{T}Y_{it}-\frac{1}{T}\sum_{t=1}^{T}Y_{Nt}-\sum_{i=1}^{N-1}\alpha_{i}.$ (2.8)
-Notice that (2.8) is independent of $\mu$ and all $\gamma_{t}$. Now, sum (2.8) over $i=1,\ldots,N-1$, we obtain
-$\sum_{i=1}^{N-1}\alpha_{i}=\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}Y_{it}-\frac{1}{T}\sum_{t=1}^{T}Y_{Nt}.$ (2.9)
-Combining (2.8) and (2.9), it follows that
-$\widehat{\alpha}_{i}=\frac{1}{T}\sum_{t=1}^{T}Y_{it}-\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}Y_{it},\qquad i=1,\ldots,N-1.$
-It can be calculated in the same way that
-$\widehat{\gamma}_{t}=\frac{1}{N}\sum_{i=1}^{N}Y_{it}-\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}Y_{it},\qquad t=1,\ldots,T-1.$
-Subsequently,
-$\widehat{\mu}=\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}Y_{it}.$ (2.10)
-In this simple example, $\widehat{\alpha}_{i}$ and $\widehat{\gamma}_{t}$ do not depend on $\mu$. This is generally not the case for complex models.
-In addition, maximizing $l(\mu,\psi)$ is identical to maximizing the original likelihood $\ell\left(\mu,\phi\right)$ under constraints $\sum_{i=1}^{N}\alpha_{i}=0$ and $\sum_{t=1}^{T}\gamma_{t}=0$.
-Specifically, it can be verified that the Lagrangian $\ell\left(\mu,\phi\right)+\lambda^{\alpha}\sum_{i=1}^{N}\alpha_{i}+\lambda^{\gamma}\sum_{t=1}^{T}\gamma_{t}$, where $\lambda^{\alpha}$ and $\lambda^{\gamma}$ are the two Lagrange multipliers, delivers upon maximization the same $\widehat{\alpha}_{i}$ and $\widehat{\gamma}_{t}$ for $i=1,\ldots,N-1$ and $t=1,\ldots,T$, respectively.
-For $\widehat{\alpha}_{N}$ and $\widehat{\gamma}_{T}$, the Lagrangian produces
-$\widehat{\alpha}_{N}=\frac{1}{T}\sum_{t=1}^{T}Y_{Nt}-\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}Y_{it},\qquad\widehat{\gamma}_{T}=\frac{1}{N}\sum_{i=1}^{N}Y_{iT}-\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}Y_{it}.$
-Indeed, $\widehat{\alpha}_{N}=-\sum_{i=1}^{N-1}\widehat{\alpha}_{i}$ as in Equation (2.9), and it can be verified that $\widehat{\gamma}_{T}=-\sum_{t=1}^{T-1}\widehat{\gamma}_{t}$ as well.
+This rules out the indeterminacy in (2.3) without explicitly using the Lagrangian.
+We use Example A.2 to illustrate the reparameterization strategy.
 2.2 Incidental Parameter Problem
@@
 We use $\mathbb{E}$ to denote the expectation w.r.t.
-$\prod_{i=1}^{N}\prod_{t=1}^{T}f(y_{it}|X_{it},\theta_{0}+\alpha_{i}^{0}+\gamma_{t}^{0}),$
-which is the true density evaluated at $\theta_{0}$, given: i) initial values of the predetermined regressors, ii) all strictly exogenous regressors, and iii) the unobserved effect $\phi^{0}$.
-Remark 1 (expectation)
-Our definition of $\mathbb{E}$ depends on the model specification and is the same to Arellano and Bonhomme (2009) and Fernández-Val and Weidner (2016).
-Consider the frequently used autoregressive model with $X_{it}=(Y_{it-1},Z_{it}^{\prime})^{\prime}$ for some strictly exogenous regressor $Z_{it}$.
-It is common that, for each $i=1,\ldots,N$, the density of $Y_{i1}$ is missing and (2.6) becomes a conditional likelihood as in Chapter 5 of Hamilton (1994).
-In this case, the expectation is taken w.r.t.
-$f_{Y_{i2},\ldots,Y_{iT}}\left(\cdot|Y_{i1},Z_{i1},\ldots,Z_{iT},\phi^{0};\theta_{0}\right)$, which is the density of only $Y_{i2},\ldots,Y_{iT}$, conditional on the initial value $Y_{i1}$, strictly exogenous regressors $Z_{i1},\ldots,Z_{iT}$, and the unobserved effects $\phi^{0}$.
-However, when the density of $Y_{i1}$ is indeed specified, so that (2.6) is an exact likelihood, the expectation is taken w.r.t.
-$f_{Y_{i1},\ldots,Y_{iT}}\left(\cdot|Z_{i1},\ldots,Z_{iT},\phi^{0};\theta_{0}\right)$, which is the joint density of all $Y_{i1},\ldots,Y_{iT}$, conditioning on the strictly exogenous regressors $Z_{i1},\ldots,Z_{iT}$ and the unobserved effects $\phi^{0}$ only.
+$\prod_{i=1}^{N}\prod_{t=1}^{T}f(y_{it}|X_{it},\theta_{0}+\alpha_{i}^{0}+\gamma_{t}^{0})$, which is the true density evaluated at $\theta_{0}$, given: i) initial values of the predetermined regressors, ii) all strictly exogenous regressors, and iii) the unobserved effect $\phi^{0}$.
+(See Remark 5 for an elaborated discussion).
 Following the profiled likelihood framework of Pace and Salvan (2006), we continue our discussion by defining, for a given $\theta$ and each pair of $(N,T)$,
@@
 are the score and Hessian.
-The remainder term $r^{l}(\theta)$ is characterized by Equation (A.17) in Appendix A.
+The remainder term $r^{l}(\theta)$ is characterized by Equation (E.17) in Section E of the supplementary appendix.
 Here the estimation errors $\widehat{\psi}(\theta)-\psi(\theta)$ can be derived by another Taylor expansion.
 Seeing $s(\theta,\widehat{\psi}(\theta))=0$ as $\widehat{\psi}(\theta)$ is the maximizer, a Taylor expansion of $s(\theta,\widehat{\psi}(\theta))$ around $\psi(\theta)$ gives
-00 0 $=s\left(\theta\right)+H\left(\theta\right)[\widehat{\psi}(\theta)-\psi(\theta)]+r^{s}(\theta),$
+$0$ $=s\left(\theta\right)+H\left(\theta\right)[\widehat{\psi}(\theta)-\psi(\theta)]+r^{s}(\theta),$
 $\widehat{\psi}(\theta)-\psi(\theta)$ $=-[H\left(\theta\right)]^{-1}s\left(\theta\right)-[H\left(\theta\right)]^{-1}r^{s}(\theta),$ (2.6)
-where $r^{s}(\theta)$ is the remainder term of the expansion, which is given by Equation (A.6) in Appendix A. Combining (2.13) and (2.12), and taking expectation, we obtain
-$\sqrt{NT}\mathbb{E}[l(\theta,\widehat{\psi}(\theta))-l(\theta,\psi(\theta))]=-\sqrt{NT}B(\theta)+\sqrt{NT}\mathbb{E}r(\theta),\qquad B(\theta):=\frac{1}{2}\mathbb{E\{}s^{\prime}(\theta)[H(\theta)]^{-1}s(\theta)\},$ (2.14)
-where $r(\theta)$ is a remainder term depending on $r^{l}(\theta)$ and $r^{s}(\theta)$, given by Equation (A.16) in Appendix A.
-By Lemma A.3 in Appendix A, we show that $\mathbb{E}r(\theta)$ is of order $o(1/\sqrt{NT})$, uniformly in $\theta$, as $N,T\rightarrow\infty$ with $N/T\rightarrow\kappa$ for some $0<\kappa<\infty$.
-This implies that $\sqrt{NT}\mathbb{E}r(\theta)=o(1)$ and is negligible.
-$B(\theta)$ is the IPP bias arising from the estimation errors of $\widehat{\psi}(\theta)$.
-As opposite to $\mathbb{E}r(\theta)$, the bias $B(\theta)$ is not negligible, in the sense that $\sqrt{NT}\mathbb{E}B(\theta)\not\rightarrow 0$ under the above asymptotic sequences $\left\{N,T\right\}$ (see Theorem 1).
-Consequently, $\widehat{\theta}$ and many likelihood-based test statistics inherit the bias from the likelihood, leading to invalid inferences.
-We consider the asymptotic sequences $N,T\rightarrow\infty$ with $N/T\rightarrow\kappa$ for some $0<\kappa<\infty$, because $\sqrt{NT}B(\theta)$ explodes under $N/T\rightarrow 0$ or $N/T\rightarrow\infty$.
-We illustrate the IPP and its consequence to the structural parameter estimator in the example below, together with the issue of $N/T\rightarrow\kappa$.
-Example 3 (heterogeneous mean model)
-Following the same setting in Example 2 but let $\varepsilon_{it}$ be normal with mean 00 and variance $\sigma_{0}>0$.
+where $r^{s}(\theta)$ is the remainder term of the expansion, which is given by Equation (E.6) in Section E of the supplementary appendix. Combining (2.6) and (2.5), and taking expectation, we obtain
+$\sqrt{NT}\mathbb{E}[l(\theta,\widehat{\psi}(\theta))-l(\theta,\psi(\theta))]$ $=-\sqrt{NT}B(\theta)+\sqrt{NT}\mathbb{E}r(\theta),$
+$B(\theta)$ $:=\frac{1}{2}\mathbb{E\{}s^{\prime}(\theta)[H(\theta)]^{-1}s(\theta)\},$ (2.7)
+[…]
+3 Bias Correction and Asymptotic Theory
+The bias expansion of $\widehat{l}(\theta):=l(\theta,\widehat{\psi}(\theta))$ in Equation (2.7) and Example A.3 above motivate our approach of correcting the likelihood $l(\theta,\widehat{\psi}(\theta))$ by adding $B(\theta)$.
+Our corrected likelihood can be viewed as an approximation to the IPP-free infeasible likelihood $l(\theta):=l(\theta,\psi(\theta))$, with approximation error $o_{\mathbb{P}}(1/\sqrt{NT})$.
+That is, it does not suffer from the IPP to the first order.
+Intuitively, the estimator of $\theta_{0}$, and the LR, LM, and Wald test statistics, will follow their standard asymptotic distributions when they are derived from the corrected likelihood (instead of $\widehat{l}(\theta)$).
+In what follows, we explain our correction procedure and discuss relevant asymptotic results.
+We start by formalizing the expansion in Equation (2.7) to show that $B(\theta)$ can be further decomposed into $B(\theta)=B_{\alpha}(\theta)+B_{\gamma}(\theta)+o(1/\sqrt{NT})$, so that
+$\mathbb{E}l(\theta)=\mathbb{E}\widehat{l}(\theta)+B_{\alpha}(\theta)+B_{\gamma}(\theta)+o(1/\sqrt{NT}),$ (3.1)
+where
+$B_{\alpha}(\theta):=\frac{1}{2}\operatorname*{trace}[S_{\alpha\alpha}(\theta)H_{\alpha\alpha}^{\mathcal{\ast}}(\theta)],\qquad B_{\gamma}(\theta):=\frac{1}{2}\operatorname*{trace}\mathbb{[}S_{\gamma\gamma}(\theta)H_{\gamma\gamma}^{\ast}(\theta)]$
+appear because of the estimation errors of, respectively, the individual-specific heterogeneities $\alpha_{i}^{0}$ and the time-specific heterogeneities $\gamma_{t}^{0}$.
+This result is formally given by Theorem 1 in Section C.3 of the supplementary appendix.
+The terms $H_{\alpha\alpha}^{\mathcal{\ast}}(\theta)$, $H_{\gamma\gamma}^{\ast}(\theta)$, $S_{\alpha\alpha}(\theta)$, and $S_{\gamma\gamma}(\theta)$ are defined from partitioning [^2] Note that $\mathbb{E}s\left(\theta\right)=0$ by definition.
+We keep $\mathbb{E}s\left(\theta\right)$ here for the convenience of constructing the corrected likelihood.
+$\mathbb{E}H(\theta)]^{-1}$ and $s(\theta)-\mathbb{E}s\left(\theta\right)$ as
+$\mathbb{E}H(\theta)]^{-1}=:\left(\begin{array}[c]{cc}H_{\alpha\alpha}^{\mathcal{\ast}}(\theta)&H_{\alpha\gamma}^{\mathcal{\ast}}(\theta)\\ H_{\gamma\alpha}^{\mathcal{\ast}}(\theta)&H_{\gamma\gamma}^{\mathcal{\ast}}(\theta)\end{array}\right),\qquad s(\theta)-\mathbb{E}s\left(\theta\right)=:\left(\begin{array}[c]{c}\widetilde{s}_{\alpha}(\theta)\\ \widetilde{s}_{\gamma}(\theta)\end{array}\right).$ (3.2)
+Here $H_{\alpha\alpha}^{\mathcal{\ast}}(\theta)$ is $(N-1)K\times(N-1)K$, which corresponds to the inverse expected Hessian w.r.t.
+$\left(\alpha_{1}^{\prime},\ldots,\alpha_{N-1}^{\prime}\right)^{\prime}$, and $H_{\gamma\gamma}^{\ast}(\theta)$ is $(T-1)K\times(T-1)K$, corresponding to the inverse expected Hessian w.r.t.
+$\left(\gamma_{1}^{\prime},\ldots,\gamma_{T-1}^{\prime}\right)^{\prime}$.
+The off-diagonal blocks $H_{\alpha\gamma}^{\mathcal{\ast}}(\theta)$ and $H_{\gamma\alpha}^{\mathcal{\ast}}(\theta)$ are defined accordingly (but are irrelevant to the construction of the corrected likelihood).
+Similarly, $\widetilde{s}_{\alpha}(\theta)$, the score w.r.t.
+$\left(\alpha_{1}^{\prime},\ldots,\alpha_{N-1}^{\prime}\right)^{\prime}$, is $(N-1)K\times 1$ and $\widetilde{s}_{\gamma}(\theta)$, the score w.r.t.
+$\left(\gamma_{1}^{\prime},\ldots,\gamma_{T-1}^{\prime}\right)^{\prime}$ is $(T-1)K\times 1$.
+Subsequently, $S_{\alpha\alpha}(\theta)$ and $S_{\gamma\gamma}(\theta)$ are constructed as
+$S_{\alpha\alpha}(\theta):=\mathbb{E\{}\widetilde{s}_{\alpha}(\theta)\mathbb{[}\widetilde{s}_{\alpha}(\theta)]^{\prime}\},\qquad S_{\gamma\gamma}(\theta):=\mathbb{E\{}\widetilde{s}_{\gamma}(\theta)\mathbb{[}\widetilde{s}_{\gamma}(\theta)]^{\prime}\},$
+which are essentially the covariance matrices of the scores $\widetilde{s}_{\alpha}(\theta)$ and $\widetilde{s}_{\gamma}(\theta)$, respectively.
+Estimating $B_{\alpha}(\theta)$ and $B_{\gamma}(\theta)$ by plug-in estimates, the corrected likelihood is established as
+$L(\theta)$ $:=\widehat{l}(\theta)+\widehat{B}_{\alpha}(\theta)+\widehat{B}_{\gamma}(\theta),$ (3.3)
+$\widehat{B}_{\alpha}(\theta)$ $:=\frac{1}{2}\operatorname*{trace}[\widehat{S}_{\alpha\alpha}(\theta)\widehat{H}_{\alpha\alpha}^{\mathcal{\ast}}(\theta)],\qquad\widehat{B}_{\gamma}(\theta):=\frac{1}{2}\operatorname*{trace}\mathbb{[}\widehat{S}_{\gamma\gamma}(\theta)\widehat{H}_{\gamma\gamma}^{\ast}(\theta)].$
+Intuitively, by adding back $\widehat{B}_{\alpha}(\theta)$ and $\widehat{B}_{\gamma}(\theta)$, we have
+$L(\theta)-l(\theta)=o_{\mathbb{P}}(1/\sqrt{NT})$ (3.4)
+as $N,T\rightarrow\infty$.
+This result is stated formally by Theorem 2 in Section C.3 of the supplementary appendix.
+Here $\widehat{H}_{\alpha\alpha}^{\mathcal{\ast}}(\theta)$, $\widehat{H}_{\gamma\gamma}^{\ast}(\theta)$, $\widehat{S}_{\alpha\alpha}(\theta)$, and $\widehat{S}_{\gamma\gamma}(\theta)$ are the plug-in versions of the corresponding quantities in $B_{\alpha}(\theta)$ and $B_{\gamma}(\theta)$.
+They are constructed as follows.
+First, denoting $\widehat{H}(\theta):=H(\theta,\widehat{\psi}(\theta))$, the two Hessian terms $\widehat{H}_{\alpha\alpha}^{\mathcal{\ast}}(\theta)$ and $\widehat{H}_{\gamma\gamma}^{\ast}(\theta)$ are defined from the partition of $\widehat{H}(\theta)]^{-1}$ as
+$\widehat{H}(\theta)]^{-1}=:\left(\begin{array}[c]{cc}\widehat{H}_{\alpha\alpha}^{\mathcal{\ast}}(\theta)&\widehat{H}_{\alpha\gamma}^{\mathcal{\ast}}(\theta)\\ \widehat{H}_{\gamma\alpha}^{\mathcal{\ast}}(\theta)&\widehat{H}_{\gamma\gamma}^{\mathcal{\ast}}(\theta)\end{array}\right)$ (3.5)
+in the same manner as in (3.2).
+Here $\widehat{H}_{\alpha\alpha}^{\mathcal{\ast}}(\theta)$ and $\widehat{H}_{\gamma\gamma}^{\mathcal{\ast}}(\theta)$ estimate $H_{\alpha\alpha}^{\mathcal{\ast}}(\theta)$ and $H_{\gamma\gamma}^{\mathcal{\ast}}(\theta)$, respectively.
+Second, denote
+$s_{i,t}^{\gamma}(\theta,\phi):=\partial_{\gamma_{t}}l_{i,t}(\theta,\phi),\qquad s_{i,t}^{\alpha}(\theta,\phi):=\partial_{\alpha_{i}}l_{i,t}(\theta,\phi),$
+which are the derivatives of the original likelihood $l_{i,t}(\theta,\phi)$ w.r.t.
+$\alpha_{i}$ and $\gamma_{t}$, respectively, evaluated at $\phi(\theta)$.
+Omitting some technical details into Section C.1 of the supplementary appendix, $\widehat{S}_{\alpha\alpha}(\theta)$ and $\widehat{S}_{\gamma\gamma}(\theta)$ are defined as:
+$\widehat{S}_{\alpha\alpha}(\theta)$ $:=D_{1}\widehat{\mathcal{S}}_{\alpha\alpha}(\theta)D_{1}^{\prime},$ (3.6)
+$\widehat{\mathcal{S}}_{\alpha\alpha}(\theta)$ $:=\frac{1}{N^{2}T^{2}}\sum_{t=1}^{T}\sum_{s=1}^{T}\mathds{1}\{\left|t-s\right|\leq\tau\}\operatorname*{diag}\{\widehat{s}_{1,t}^{\alpha}(\theta)[\widehat{s}_{1,s}^{\alpha}(\theta)]^{\prime},\ldots,\widehat{s}_{N,t}^{\alpha}(\theta)[\widehat{s}_{N,s}^{\alpha}(\theta)]^{\prime}\},$
+$\widehat{s}_{i,t}^{\alpha}(\theta)$ $:=s_{i,t}^{\alpha}(\theta,\widehat{\phi}(\theta))-\frac{1}{T}\sum_{t=1}^{T}s_{i,t}^{\alpha}(\theta,\widehat{\phi}(\theta)),$
+where $\widehat{\phi}(\theta)):=D^{\prime}\widehat{\psi}\left(\theta\right)$ is the “unparameterized” counterpart of $\widehat{\psi}\left(\theta\right)$; and
+$\widehat{S}_{\gamma\gamma}(\theta)$ $:=D_{2}\widehat{\mathcal{S}}_{\gamma\gamma}(\theta)D_{2}^{\prime},$ (3.7)
+$\widehat{\mathcal{S}}_{\gamma\gamma}(\theta)$ $:=\frac{1}{N^{2}T^{2}}\sum_{i=1}^{N}\widehat{s}_{i}^{\gamma}(\theta)[\widehat{s}_{i}^{\gamma}(\theta)]^{\prime},\qquad\widehat{s}_{i}^{\gamma}(\theta):=\{[\widehat{s}_{i,1}^{\gamma}(\theta)]^{\prime},\ldots,[\widehat{s}_{i,T}^{\gamma}(\theta)]^{\prime}\}^{\prime},$
+$\widehat{s}_{i,t}^{\gamma}(\theta)$ $:=s_{i,t}^{\gamma}(\theta,\widehat{\phi}(\theta))-\frac{1}{N}\sum_{i=1}^{N}s_{i,t}^{\gamma}(\theta,\widehat{\phi}(\theta)).$
+The indicator $\mathds{1}\{\left|t-s\right|\leq\tau\}$ is a truncation mechanism, with truncation parameter $\tau$, which is common in the relevant literature.
+See Remark 6 for detailed discussions.
+In our simulation, we use $\tau=1$ and $2$ and find the difference relatively insignificant.
+We use the following remark to provide some advice for an efficient computation of $\widehat{B}_{\alpha}(\theta)$ and $\widehat{B}_{\gamma}(\theta)$.
+Remark 1 (efficient computation of bias terms)
+The terms $\widehat{S}_{\alpha\alpha}(\theta)$ and $\widehat{S}_{\gamma\gamma}(\theta)$ are constructed using the derivatives of the original (i.e., “unparameterized”) likelihood.
+This is valid because the reparameterization produces
+$\partial_{\psi}l(\theta,\psi)=D\partial_{\phi}\ell(\theta,\phi),\qquad\partial_{\psi\psi^{\prime}}l(\theta,\psi)=D\partial_{\phi\phi^{\prime}}\ell(\theta,\phi)D^{\prime},$
+holding true for every $\phi=D^{\prime}\psi$ and every $\theta$.
+The second identity above implies that the Hessian in (3.5) may also be constructed similarly from the original likelihood.
+Generally, it may be easier to construct $\widehat{S}_{\alpha\alpha}(\theta)$, $\widehat{S}_{\gamma\gamma}(\theta)$, and $\widehat{H}(\theta)$ from the original likelihood, because analytical expressions for $\partial_{\phi}\ell(\theta,\phi)$ and $\partial_{\phi\phi^{\prime}}\ell(\theta,\phi)$ are well-known for many frequently used models (e.g., the probit, the logit, and the Poisson).
+In addition, if the underlying model is equipped with linear indices [^3] See also Example A.
+Note that many frequently used models including the probit, the logit, the Poisson, etc.
+have linear indices., these derivatives may be calculated efficiently, making use of the chain rule.
+In particular, the likelihood $l_{i,t}(\theta,\phi)$ can be viewed as a compound function, as $l_{i,t}(\theta,\phi):=l_{i,t}(\pi_{i,t}\left(\theta,\phi\right))$ with $\pi_{i,t}\left(\theta,\phi\right):=X_{it}^{\prime}(\theta+\alpha_{i}+\gamma_{t})$ being the linear index.
+Using the chain rules, derivatives can be obtained as
+$\partial_{\phi}l_{i,t}(\theta,\phi)$ $=\partial_{\pi}l_{i,t}(\pi_{i,t}\left(\theta,\phi\right))\partial_{\phi}\pi_{i,t}\left(\theta,\phi\right),$
+$\partial_{\phi\phi^{\prime}}l_{i,t}(\theta,\phi)$ $=\partial_{\pi^{2}}l_{i,t}(\pi_{i,t}\left(\theta,\phi\right))\partial_{\phi}\pi_{i,t}\left(\theta,\phi\right)\left[\partial_{\phi}\pi_{i,t}\left(\theta,\phi\right)\right]^{\prime}.$
+Here $\partial_{\phi}\pi_{i,t}\left(\theta,\phi\right)$ is high-dimensional but can be calculated easily, because $\pi_{i,t}\left(\theta,\phi\right)$ is only linear.
+The more complex components $\partial_{\pi}l_{i,t}(\cdot)$ and $\partial_{\pi^{2}}l_{i,t}(\cdot)$ are only scalars and, therefore, can be calculated afforably using numerical differentiation (if their analytical expressions are too complex).
+This procedure extends easily to models with multiple linear indices (see, e.g., Sun and Dhaene (2019)).
+Its complexity only grows with the number of linear indices, but not the dimension of $\phi$, implying computational efficiency even for large $N,T$.
+Finally, we would like to mention that, indeed, our method requires inverting the Hessian $\widehat{H}(\theta)$.
+However, this is computationally affordable with modern computers for intermediate $\left(N,T\right)$.
+For instance for $K=3$ and $\left(N,T\right)=\left(90,90\right)$, $\widehat{H}(\theta)$ is $540\times 540$.
+The time spent in the inversion is typically below 0.01 seconds with an AMD Ryzen 9 5950X CPU.
+Remark 2 (bias correction)
+As opposite to fixed-$T$ methods including, e.g., the conditional logit of Chamberlain (1980), our approach is a large-$T$, “first-order” bias correction procedure.
+Similar to Arellano and Hahn (2016), Hahn and Newey (2004), Bester and Hansen (2009), Arellano and Hahn (2016), Fernández-Val and Weidner (2016), our approach only removes the bias up to $O(1/\sqrt{NT})$, making $L(\theta)$ (hence the estimators and test statistics) asymptotically unbiased under $N/T\rightarrow\kappa$ as $N,T\rightarrow\infty$.
+For fixed $T$, our approach generally does not eliminate the IPP.
+However, one advantage of our method is that it mitigates the IPP for a wide range of models, especially for those without fixed-$T$ consistent estimation techniques.
+For instance, there has been no fixed-$T$ consistent procedure for the two-way heterogeneous parameter logit model, to the best of our knowledge.
+Indeed, since our approach is only first-order, caution shall be observed when the sample size is small, in which case higher-order bias terms could be non-negligible, requiring higher-order corrections.
+For models with individual effects, available methods include Dhaene and Jochmans (2015), Dhaene and Sun (2021), and Schumann (2023).
+However, we are not aware of any higher-order methods for two-way DN-HP models.
+To keep the main text compact, we provide by Assumption 1 of Section C.2 (in the supplementary appendix) rigorous regularity conditions for the validity of the expansion in (3.1) and the (uniform) rate of convergence of $L\left(\theta\right)$ as in (3.4).
+Here, we would like to mention that most of our conditions are standard in the literature.
+See, e.g., Hahn and Kuersteiner (2007), Fernández-Val and Weidner (2016), and Dhaene and Jochmans (2015).
+Detailed motivations for these assumptions are provided in Section C.2.
+In addition, we also provide low-level conditions for a logit model by Lemma C.1 in Section C.2.
+The proposed $L(\theta)$ can be viewed as an extension of the corrected likelihood of Bester and Hansen (2009) and Arellano and Hahn (2016) to two-way DN-HP models.
+This extension is, however, non-trivial.
+We provide an elaborated discussion in Section D of the supplementary appendix for this matter.
+When the model only contains individual effects in the intercept, our method reduces to, e.g., Arellano and Hahn (2016), applied on the original likelihood $\ell\left(\theta,\phi\right)$, except that we use a truncation-based (instead of a bandwidth-based) approach to guarantee the positive definiteness of $\widehat{\mathcal{S}}_{\alpha\alpha}(\theta)$.
+We illustrate this in Example D.1 of Section D of the supplementary appendix, using a dynamic logit model.
+Note that, for the logit model, there are fixed-$T$ consistent estimators when fixed effects only appear in the intercept (but not in the slope).
+See Remark 7.
+Using the corrected likelihood $L(\theta)$, we may obtain the corrected estimator of $\theta_{0}$ as
+$\widehat{\theta}_{L}:=\arg\max_{\theta}L(\theta).$
+Intuitively, since $L(\theta)$ approximates $l(\theta)$ with an error negligible relative to $1/\sqrt{NT}$, the estimator $\widehat{\theta}_{L}$ inherits this rate and is consistent and asymptotically normal with mean zero under $N/T\rightarrow\kappa$ as $N,T\rightarrow\infty$ and other mild regularity conditions.
+In particular,
+$\sqrt{NT}[-\mathbb{E}\triangledown_{\theta\theta^{\prime}}l(\theta_{0})]^{1/2}(\widehat{\theta}_{L}-\theta_{0})\overset{\mathbb{D}}{\longrightarrow}\mathcal{N}(0,\mathbb{I}_{K}),$ (3.8)
+where $\mathcal{N}(\mu,\Sigma)$ stands for the normal distribution with mean $\mu$ and covariance matrix $\Sigma$, $\nabla_{\theta\theta^{\prime}}$ denotes the second total derivative w.r.t.
+to $\theta$.
+The proof of this claim is straightforward and largely relies only on standard techniques.
+We give this result formally in Corollary 1 in Section C.3 of the supplementary appendix and provide regularity conditions thereof.
+The result in (3.8) makes use of the information matrix equality to simplify the presentation.
+We refer the reader to, e.g., White (1982) when such an equality does not hold.
+See also the discussion of Corollary 1.
+Before continuing, we leave the remark below for an efficient calculation of $\widehat{\theta}_{L}$.
+Remark 3 (two-layer maximization)
+The computation of $\widehat{\theta}_{L}$ involves a two-layer maximization: In the inner layer, $\widehat{\psi}(\theta)$ is obtained by maximizing $l(\theta,\psi)$, keeping $\theta$ fixed; while in the outer layer, $L(\theta)$ is maximized w.r.t.
+$\theta$.
+This maximization problem can be done efficiently using the Newton-Raphson algorithm in both layers.
+In the inner layer, we recommend following Remark 1 to calculate the first and second derivatives efficiently.
+In the outer layer, it is generally affordable to use numerical derivatives, because the dimension of $\theta$ is usually small, provided that the inner-layer maximization is done efficiently.
+As a reference, with an AMD Ryzen 9 5950X CPU, the calculation of $\widehat{\theta}_{L}$ for the dynamic logit model considered in our simulation (see Section 4 below) with $\left(N,T\right)=\left(90,90\right)$ (note, $NT=8100$) takes less than $50$ seconds.
+For hypothesis testing procedures, we consider a generic null hypothesis $H_{0}:R(\theta_{0})=0$, where $R(\theta)$ is a known $r\times 1$ ($r\leq K$) vector-valued non-random function with Jacobian $J(\theta):=\partial_{\theta^{\prime}}R(\theta)$ satisfying $\operatorname*{rank}[J(\theta)]=r$.
+The LR ($\widehat{\xi}_{LR}$), LM ($\widehat{\xi}_{LM}$), and Wald ($\widehat{\xi}_{LM}$) test statistics can be constructed as
+$\widehat{\xi}_{LR}$ $:=-2NT\{L(\widehat{\theta}_{R})-L(\widehat{\theta}_{L})\},$
+$\widehat{\xi}_{LM}$ $:=-NT\{\triangledown_{\theta^{\prime}}L(\widehat{\theta}_{R})[\triangledown_{\theta\theta^{\prime}}L(\widehat{\theta}_{R})]^{-1}\triangledown_{\theta}L(\widehat{\theta}_{R})\},$
+$\widehat{\xi}_{Wald}$ $:=-NT\{R^{\prime}(\widehat{\theta}_{L})[J(\widehat{\theta}_{L})[\triangledown_{\theta\theta^{\prime}}L(\widehat{\theta}_{L})]^{-1}J^{\prime}(\widehat{\theta}_{L})]^{-1}R(\widehat{\theta}_{L})\},$
+where $\nabla_{\theta}$ denotes the first total derivative w.r.t.
+to $\theta$, and $\widehat{\theta}_{R}:=\arg\max_{\theta}L(\theta)$ subject to $R(\theta)=0$.
+The same argument above intuitively indicates that
+$\widehat{\xi}_{LR},\widehat{\xi}_{LM},\widehat{\xi}_{Wald}\overset{\mathbb{D}}{\longrightarrow}\chi^{2}(r),$
+under $H_{0}$ as $N,T\rightarrow\infty$, where $\chi^{2}(r)$ is the $\chi^{2}$-distribution with degrees of freedom $r$.
+The proof is also quite textbook-standard.
+We state this result formally by Corollary 2 in Section C.3 of the supplementary appendix.
+It is worth to mention that the inference procedure based on the corrected likelihood $L\left(\theta\right)$ is uniformly valid regardless of whether the slope coefficient is heterogeneous or homogeneous.
+We elaborate on this in Remark 4 in Section B of the supplementary appendix.
+Finally, although we focus on the structural parameters and the hypothesis tests in this paper, we include a brief discussion about the average partial effects (APEs) in Section H of the supplementary appendix, where we i) discuss that the APEs of a DN-HP model may be asymptotically unbiased, following from the same arguments as in Fernández-Val and Weidner (2016); and ii) show using simulation studies that calculating the APEs using our corrected estimator $\widehat{\theta}_{L}$ may already lead to a considerable reduction of the finite-sample bias.
+4 Simulation
+In this section, we present a simulation study. We consider dynamic binary response models specified according to the data-generating process (Design 1):
+$Y_{it}=\mathds{1}\{\rho_{0}Y_{it-1}+(\beta_{0}+\alpha_{1,i}^{0}+\gamma_{1,t}^{0})Z_{it}+\alpha_{2,i}^{0}+\gamma_{2,t}^{0}+\varepsilon_{it}>0\},$ (4.1)
+[…]
+Summarizing important insights in Tables 1 and 2, our findings are as follows.
+  1. 1.
+    The MLEs $\widehat{\beta}$ and $\widehat{\rho}$ may exhibit significant bias, especially when the sample size is small.
+Our procedure reduces the bias considerably without inflating the RMSE, even at $\left(N,T\right)=\left(30,30\right)$.
+  2. 2.
+    Our corrected estimators show smaller biases than the jackknife and the bootstrap for $\beta$ when the sample size is small.
+The biases of all candidate estimators, except the MLEs, become similar as the sample size increases.
+For $\left(N,T\right)=\left(30,30\right)$, the jackknife estimators may show larger biases than the MLEs for structural parameters associated with two-way heterogeneities.
+We discuss the possible cause in Section I of the supplementary appendix.
+  3. 3.
+    The LR test based on $\widehat{l}\left(\theta\right)$ has severe size-distortions even with $\left(N,T\right)=\left(90,90\right)$.
+On the contrary, the LR test based on our $L\left(\theta\right)$ is able to deliver an empirical size close to the nominal level of $5\%$ as the sample size increases, maintaining relatively strong powers.
+For small sample sizes, $L\left(\theta\right)$ reduces the type-I error risk considerably.
+The empirical sizes from the bootstrap are close to the nominal level with small sample sizes.
+However, the powers are relatively low, especially near the true null hypothesis.
+Table 1: Comparisons of Uncorrected and Corrected Logit Estimates, Design 1
+$(N,T)$ $(30,30)$ $(60,60)$ $(90,90)$
+Bias RMSE Bias RMSE Bias RMSE
+Static Logit Model
+$\widehat{\beta}$ $0.165$ $0.197$ $0.073$ $0.086$ $0.045$ $0.053$
+$\widehat{\beta}_{L}$ $0.055$ $0.106$ $0.015$ $0.044$ $0.006$ $0.027$
+$\widehat{\beta}_{J}$ 0.776-0.776 $1.716$ 0.022-0.022 $0.049$ 0.008-0.008 $0.028$
+$\widehat{\beta}_{B}$ 0.069-0.069 $0.098$ 0.008-0.008 $0.040$ 0.004-0.004 $0.026$
+Dynamic Logit Model
+$\widehat{\beta}$ $0.170$ $0.210$ $0.072$ $0.087$ $0.046$ $0.054$
+$\widehat{\beta}_{L}^{(1)}$ $0.061$ $0.120$ $0.016$ $0.046$ $0.007$ $0.027$
+$\widehat{\beta}_{L}^{(2)}$ $0.063$ $0.122$ $0.017$ $0.046$ $0.007$ $0.027$
+$\widehat{\beta}_{J}$ 1.014-1.014 $2.120$ 0.029-0.029 $0.111$ 0.009-0.009 $0.029$
+$\widehat{\beta}_{B}$ 0.079-0.079 $0.110$ 0.010-0.010 $0.042$ 0.004-0.004 $0.026$
+$\widehat{\rho}$ 0.109-0.109 $0.209$ 0.051-0.051 $0.097$ 0.030-0.030 $0.059$
+$\widehat{\rho}_{L}^{(1)}$ 0.047-0.047 $0.165$ 0.018-0.018 $0.079$ 0.007-0.007 $0.049$
+$\widehat{\rho}_{L}^{(2)}$ 0.042-0.042 $0.168$ 0.013-0.013 $0.079$ 0.003-0.003 $0.049$
+$\widehat{\rho}_{J}$ $0.010$ $0.177$ $0.002$ $0.081$ $0.003$ $0.051$
+$\widehat{\rho}_{B}$ $0.009$ $0.153$ $0.000$ $0.077$ $0.003$ $0.049$
+[…]
+Table 2: Empirical Size and Power from LR Test for Logit Model, Design 1
+Size Power at $\delta$ (Analytical) Power at $\delta$ (Bootstrap)
+$(N,T)$ Uncorrected Analytical Bootstrap 0.2-0.2 0.1-0.1 $0.1$ $0.2$ 0.2-0.2 0.1-0.1 $0.1$ $0.2$
+Static Logit Model
+$(30,30)$ $38$ $13$ 0 $88$ $48$ $12$ $45$ $53$ $12$ 0 0
+$(60,60)$ $40$ 88 22 $100$ $85$ $59$ $99$ $99$ $74$ 0 $32$
+$(90,90)$ $38$ 55 33 $100$ $98$ $95$ $100$ $100$ $98$ 99 $97$
+Dynamic Logit Model
+$(30,30)$ $37$ $13$ 11 $84$ $42$ 1919 $55$ $42$ 88 11 44
+$(60,60)$ $36$ $10$ 33 $100$ $79$ $64$ $99$ $99$ $65$ 77 $68$
+$(90,90)$ $33$ 66 33 $100$ $98$ $95$ $100$ $100$ $96$ $34$ $99$
+[…]
+In Section F.1 of the supplementary appendix, we report the empirical sizes and powers from the LM and the Wald test, which are similar to the LR here.
+We also report results from the probit model in Section F.1.
+In Table F.6, we present simulation results for $\left(N,T\right)=\left(90,10\right)$ showing the performance of our approach for very small $T$.
+Second, we present additional simulation results in Section F.2 of the supplementary appendix for models with heterogeneous autoregressive coefficient.
+Next, in Section F.3 of the supplementary appendix, we present results from an alternative design (Design 2), where the biases of the MLEs are larger.
+Our bias correction procedure is effective under both Designs 1 and 2.
+Finally, in Section F.4 of the supplementary appendix, we simulate a dynamic panel probit model with additive two-way fixed effects in the intercept, under Design 1 of Fernández-Val and Weidner (2016).
+5 Empirical Analysis
+In this section we apply our likelihood-based bias correction method to examining the determinants of the labor force participation (LFP) decision of single mothers.
+In particular, we look at the impact of the number of children on the decision of the mother to engage in paid employment.
+An extensive literature in labor economics has studied the labor supply decisions of married women (Killingsworth and Heckman, 1986; Angrist and Evans, 1998; Blau and Kahn, 2007).
+These studies have uncovered the impacts of a variety of economic variables, including female market wage and husband income (Mincer, 1962), education (Heath and Jayachandran, 2018), childcare costs (Connelly, 1992), the cost of home technology (Greenwood et al., 2016), and culture norms (Fernández, 2013).
+Among these variables, the number of children consistently emerges as one of the most important determinants of female labor supply (Nakamura and Nakamura, 1992).
+This is unsurprising since women continue to bear a disproportionate share of child-rearing responsibilities (Aguero and Marks, 2008).
+In contrast to the substantial body of research on the labor supply behavior of married women, the labor supply decisions of single mothers have received relatively limited attention, with only a few exceptions (Kimmel, 1998; Blundell et al., 2016).
+Single mothers, however, face distinctive challenges when it comes to balancing work and child-rearing responsibilities due to the absence of a second earner in the household.
+Moreover, their employment decisions may have a more pronounced impact on their children’s well-being compared to the decisions of married women and should thus be of great importance to economists and policymakers.
+To study the labor supply decisions of single mothers, we compile a dataset from waves 20–30 of the Panel Study of Income Dynamics (PSID), which span the period of 1987 to 1997.
+Our sample includes only single mothers, defined as unmarried female household heads with at least one child.
+The dependent variable is labor force participation status, defined as whether the individual worked any hours during the interview year.
+Our main explanatory variable is the number of children under 18 in the household.
+Following Dhaene and Jochmans (2015), we focus on the informative sample of individuals aged 18–60 whose current and lagged LFP status each changed at least once during the period.
+This restriction ensures within-individual variation in both variables, which is necessary for identification of the dynamic model.
+After applying this criterion, the estimation sample comprises $N=86$ single mothers observed for $T=10$ consecutive years (1987–1996), with lagged participation status measured from 1986 to 1995.
+Further details on data construction are provided in Section G of the supplementary appendix.
+On average, 28 percent of these individuals switched into or out of the labor force in a given year (Figure G.1G.2).
+Table G.1 reports the summary statistics.
+Let $Y_{it}\in\{0,1\}$ denote the labor-force-participation status of single mother $i$ in year $t$, and let $X_{it}$ denote her number of children.
+To assess the impact of fertility on labor supply, we estimate the following dynamic probit models:
+$Y_{it}$ $=\mathds{1}\!\left\{\rho Y_{i,t-1}+\beta X_{it}+c_{i}+\tau_{t}+\varepsilon_{it}>0\right\},$ (5.1)
+$Y_{it}$ $=\mathds{1}\!\left\{(\rho+\zeta_{i}+\eta_{t})Y_{i,t-1}+(\beta+\alpha_{i}+\gamma_{t})X_{it}+c_{i}+\tau_{t}+\varepsilon_{it}>0\right\},$ (5.2)
+where $\varepsilon_{it}$ follows the standard normal distribution, and $\sum_{i}c_{i}=\sum_{i}\zeta_{i}=\sum_{i}\alpha_{i}=\sum_{t}\eta_{t}=\sum_{t}\gamma_{t}=0$.
+Model (5.1) is the dynamic homogeneous-slope model, which includes two-way fixed effects in the intercept.
+Model (5.2) is the dynamic heterogeneous-slope model, which extends the specification by allowing the slope coefficients—that is, the coefficients on both the lagged dependent variable and the number of children—to vary across individuals and over time: $\rho_{it}=\rho+\zeta_{i}+\eta_{t}$ and $\beta_{it}=\beta+\alpha_{i}+\gamma_{t}$.
+The parameter $\beta_{it}$ measures the heterogeneous impact of the number of children on a mother’s latent propensity to work, while $\rho_{it}$ captures heterogeneous state dependence in labor force participation.
+Model (5.2) nests Model (5.1) as a special case when $\rho_{it}$ and $\beta_{it}$ are constant across $i$ and $t$.
+As discussed in Section 1, when the true effect of $X_{it}$ on the outcome, $\beta_{it}$, varies across the population, it is important to account for its potential correlation with $X_{it}$.
+In practice, such correlation often arises from self-selection.
+For instance, single mothers who choose to have more children might be those with greater financial resources or better access to childcare, such that their labor supply is less affected by additional children.
+Alternatively, correlation between $\beta_{it}$ and $X_{it}$ may stem from common time trends: over time, fertility rates declined while concurrent factors such as rising childcare and schooling costs, advances in home technology, and evolving welfare policies (e.g., child tax credits) altered the effect of each additional child on their mother’s LFP.
+In both scenarios, Model (5.2) is the appropriate specification.
+It controls for unobserved heterogeneity in $\beta_{it}$ by incorporating individual and time fixed effects into the slope and allowing them to correlate with $X_{it}$.
+Controlling for heterogeneous state dependence further improves robustness by allowing the persistence of labor-force participation to differ across individuals and periods, capturing additional sources of dynamic heterogeneity.
+Given our small sample size, direct estimation of both models could suffer from severe incidental parameter problems, which we address using our bias correction procedure.
+Table 3 presents the estimation results.
+Panel A reports results for Model (5.1) under the homogeneous slope assumption.
+For each parameter, we provide the MLE and bias-corrected estimates, along with their standard errors.
+Examining the bias-corrected estimates, we observe that an increase in the number of children appears to reduce a single mother’s likelihood of labor force participation, while past LFP status strongly predicts current LFP.
+However, this relationship shifts dramatically once we account for slope heterogeneity.
+Panel B presents the results for Model (5.2), allowing for heterogeneous slopes.
+Here, the bias-corrected estimates reveal a positive average impact of additional children on a mother’s labor supply: each additional child increases the latent propensity to participate in the labor force by 0.436, which corresponds to an average increase of about 7 percent in the probability of employment.
+This indicates that, on average, single mothers with more children are more likely to work, likely driven by the increased financial demands of supporting multiple children independently.
+Additionally, unlike the homogeneous-slope model, the serial correlation in labor-force participation remains statistically significant but diminishes in magnitude, suggesting that what appears to be strong state dependence in the homogeneous model partly reflects unobserved heterogeneity in individuals’ persistence of employment rather than uniform structural dynamics.
+Together, these findings underscore the importance of controlling for unobserved slope heterogeneity, as failing to do so can lead to substantially different conclusions.
+In addition to highlighting the difference between homogeneous and heterogeneous slope models, Table 3 demonstrates the importance of bias-correction in estimating dynamic nonlinear panel data models.
+Examining the results in Panel B, we observe that the MLE overestimates the impact of the number of children while underestimating the degree of serial correlation in LFP.
+Bias correction results in a 38 percent lower estimate of the former and a 32 percent higher estimate of the latter.
+Finally, in Figure 2(a), we plot the distribution of the estimated $\beta_{it}$ from Model (5.2), with summary statistics in Table G.2 (supplementary appendix).
+Each $\beta_{it}$ estimate is constructed as the sum of the bias-corrected average effect $\beta$ and the estimated individual and time effects $\alpha_{i}$ and $\gamma_{t}$.
+These estimates reveal substantial heterogeneity in how the number of children influences single mothers’ labor force participation.
+While the mean effect is positive, about 46 percent of the individual-time observations show a decrease in LFP with each additional child, suggesting that the average effect alone inadequately captures the diverse impacts across individuals and over time.
+Figure 2(b) illustrates the temporal evolution of these heterogeneous effects.
+The median $\beta_{it}$ rises steadily from 1987 to 1994—implying an increasingly positive labor-supply response to additional children—before declining modestly thereafter.
+The interquartile and decile bands remain wide throughout, highlighting persistent dispersion across individuals.
+Together, these results reinforce the importance of accounting for both individual and temporal variations in slope coefficients, as the impact of children on labor force participation can vary significantly based on unobserved factors unique to each mother and time period.
+Table 3: Single Mother Labor Force Participation: Estimation Results
+   Number of Children    Lagged Participation
+Uncorrected Corrected Uncorrected Corrected
+      Panel A. Homogeneous Slope
+Model parameter -0.923 -0.851 0.589 0.540
+(0.232) (0.228) (0.108) (0.107)
+      Panel B. Heterogeneous Slope
+Model parameter 0.704 0.436 0.191 0.253
+(0.141) (0.144) (0.106) (0.105)
+  • Notes: Standard errors in parentheses.
+Panel A reports the estimation results of the dynamic homogeneous-slope model.
+Panel B reports the estimation results of the dynamic heterogeneous-slope model.
+The “uncorrected” and the “corrected” columns show respectively the MLE and the bias-corrected parameter estimates.
+Data source: PSID 1987–1997.
+Figure 2: Estimated Heterogeneous Effects of Children on Single Mother Labor Force Participation
+(a) Distribution of Estimated Effects
+(b) Temporal Evolution of Estimated Effects
+[…]
+Notes: Panel (a) shows the distribution of the estimated individual-time effects $\beta_{it}$ from the dynamic heterogeneous-slope model.
+Panel (b) plots the median (solid), interquartile range (25th–75th, dashed), and decile range (10th–90th, dotted) of $\beta_{it}$ over time.
+In conclusion, our analysis demonstrates the efficacy of our likelihood-based bias correction method for estimating dynamic nonlinear models of labor supply with multi-dimensional heterogeneities.
+We show the importance of controlling for unobserved slope heterogeneity when it may correlate with regressors and the value of panel data models that accommodate two-way fixed effects in both the intercept and the slope.
+Our findings reveal that the number of children has, on average, a positive influence on a single mother’s labor force participation, but such effect is highly heterogeneous across individuals and time periods.
+This heterogeneity underscores the complex nature of labor supply decisions and the unique challenges each single mother faces in balancing work and child-rearing responsibilities.
+6 Final Remarks
+In this paper, we propose a likelihood-based analytical bias correction procedure for a general class of two-way dynamic nonlinear heterogeneous parameter models subject to the incidental parameter problem.
+We give the analytical form of a corrected likelihood and show that it delivers point estimators that are asymptotically unbiased and test statistics that are asymptotically $\chi^{2}$-distributed.
+Simulation studies and empirical analyses support our claims.
+Several issues deserve further studies.
+First, we have not rigorously investigated the asymptotic properties of the average partial effects.
+Admittedly, average partial effect is very important, especially for nonlinear models.
+It is therefore worthwhile to formally establish its asymptotic theory.
+Second, there has been various fixed-$T$ consistent solutions for the logit model with fixed effects.
+However, to the best of our knowledge, we have not seen any such developments for the two-way heterogeneous parameter logit model.
+It could be useful to push such a development, because the logit model is widely used in many studies.
+Finally, certain micro panels tend to have a small $T$, in which case the higher-order bias terms may be non-negligible.
+For this reason, it may be potentially important to develop high-order bias correction approaches for two-way dynamic nonlinear heterogeneous parameter models.
+Acknowledgement
+We would like to thank the editor Iván Fernández-Val and two anonymous referees for their valuable comments and suggestions.
+We would like to express our gratitude to comments and suggestions from Geert Dhaene, Yanqin Fan, Jinyong Hahn, Chen Hsiao, Whitney Newey, and Jun Yu.
+Our research assistants Yuchi Liu and Xinrui Zhang helped us with various tasks.
+Leng’s research is partially supported by the National Natural Science Foundation of China (72573136) and the Fujian Natural Science Foundation in China (2024J08013).
+Sun acknowledges the support from the National Natural Science Foundation of China under Grant Number 72203032.
+Appendix
+[…]
+AppendixAppendix
+A Examples
+Example A.1 (heterogeneous parameter model)
+Consider a probit model with the data-generating process
+$Y_{it}=\mathds{1}\{\pi(X_{it},\theta_{0},\alpha_{i}^{0},\gamma_{t}^{0})+\varepsilon_{it}>0\},$ (A.1)
+where $\varepsilon_{it}$ is a standard normal random variable independent of $X_{it}$, $\theta_{0}$, $\alpha_{i}^{0}$, and $\gamma_{t}^{0}$.
+For arguments $\theta$, $\alpha_{i}$, and $\gamma_{t}$ of dimension $K\times 1$, we allow the linear index [^5] We use a model with a linear index solely for the purpose of illustration.
+Our method applies to general models.
+One example is the modified Neyman-Scott model in Example A.3.
+$\pi(X_{it},\theta,\alpha_{i},\gamma_{t})$ to take the following general form:
+$\pi(X_{it},\theta,\alpha_{i},\gamma_{t})=\sum_{k=1}^{K}(\theta_{k}+\alpha_{k,i}+\gamma_{k,t})X_{k,it},$
+where $X_{k,it}$, $\theta_{k}$, $\alpha_{k,i}$, and $\gamma_{k,t}$ are the $k$-th components of $X_{it}$, $\theta$, $\alpha_{i}$, and $\gamma_{t}$, respectively, $k=1,\ldots,K$.
+In addition, when $X_{1,it}=1$, the data-generating process in (A.1) consists of the traditional individual- and time-effects in the intercept.
+When $X_{k,it}=Y_{it-1}$ for some $k$, the data-generating process is dynamic because each $Y_{it}$ depends on its past value.
+Moreover, we also allow a subset of the coefficients to be homogeneous, thereby accommodating both heterogeneous and homogeneous parameters.
+For instance, when it is known that $\alpha_{1,i}^{0}=\gamma_{1,t}^{0}=0$, the model can be estimated under constraints $\alpha_{1,i}=\gamma_{1,t}=0$ for all $(i,t)$.
+In this case, the linear index is effectively
+$\pi(X_{it},\theta,\alpha_{i},\gamma_{t})=\theta_{1}X_{1,it}+\sum_{k=2}^{K}(\theta_{k}+\alpha_{k,i}+\gamma_{k,t})X_{k,it},$
+where $X_{1,it}$ is associated with a “homogeneous parameter” $\theta_{1}$ while all other $X_{k,it}$ ($k=2,\ldots,K$) are associated with heterogeneous parameters $(\theta_{k}+\alpha_{k,i}+\gamma_{k,t})$.
+In a similar fashion, our approach also applies to, e.g.,
+$\pi(X_{it},\theta,\alpha_{i},\gamma_{t})=(\alpha_{1,i}+\gamma_{1,t})X_{1,it}+\sum_{k=2}^{K}(\theta_{k}+\alpha_{k,i}+\gamma_{k,t})X_{k,it},$
+with the constraint $\theta_{1}=0$. Other variants of $\pi(X_{it},\theta,\alpha_{i},\gamma_{t})$ include:
+$\pi(X_{it},\theta,\alpha_{i},\gamma_{t})$ $=(\alpha_{1,i}+\gamma_{1,t})+\theta_{2}Y_{it-1}+\sum_{k=3}^{K}\theta_{k}X_{k,it},$ (A.2)
+$\pi(X_{it},\theta,\alpha_{i},\gamma_{t})$ $=(\theta_{1}+\alpha_{1,i}+\gamma_{1,t})+\sum_{k=2}^{K}(\theta_{k}+\alpha_{k,i}+\gamma_{k,t})X_{k,it},$ (A.3)
+$\pi(X_{it},\theta,\alpha_{i},\gamma_{t})$ $=(\theta_{1}+\alpha_{1,i}+\gamma_{1,t})+(\theta_{2}+\alpha_{2,i}+\gamma_{2,t})Y_{it-1}$
+$\quad+\sum_{k=3}^{K_{1}}(\theta_{k}+\alpha_{k,i}+\gamma_{k,t})X_{k,it}+\sum_{k=K_{1}+1}^{K}\theta_{k}X_{k,it}$ (A.4)
+for some $3\leq K_{1}<K$.
+To our knowledge, except (A.2 ) considered by Fernández-Val and Weidner (2016), none of the above specifications are fully addressed under nonlinear models with predetermined regressors.
+The specification in (A.3) is the same to Keane and Neal (2020) and Lu and Su (2023), except that they consider only the linear regression model.
+Jochmans and Otsu (2019) provide a bias correction technique for general nonlinear models with (A.3) only for strictly exogenous $X_{k,it}$.
+Currently, no methods are available for nonlinear models with (A.4).
+Example A.2 (reparameterization)
+For some scalar structural parameter $\mu_{0}$ of interests and $K=1$, consider $Y_{it}=\mu_{0}+\alpha_{i}^{0}+\gamma_{t}^{0}+\varepsilon_{it}$ where $\varepsilon_{it}$ is standard normal independent across both $i$ and $t$, and independent of $\alpha_{i}^{0}$ and $\gamma_{t}^{0}$.
+We restrict the variance of $\varepsilon_{it}$ to be $1$ because we use this example only to demonstrate the identification of $\mu_{0}$.
+In Example A.3 below, we indeed allow $\varepsilon_{it}$ to have an unknown variance.
+Conditioning on $\phi^{0}$, the original likelihood function is
+$\ell(\mu,\phi)$ $=-\frac{1}{2}\log(2\pi)-\frac{1}{NT}\sum_{i=1}^{N-1}\sum_{t=1}^{T-1}\frac{1}{2}(Y_{it}-\mu-\alpha_{i}-\gamma_{t})^{2}-\frac{1}{NT}\sum_{t=1}^{T-1}\frac{1}{2}(Y_{Nt}-\mu-\alpha_{N}-\gamma_{t})^{2}$
+$\quad-\frac{1}{NT}\sum_{i=1}^{N-1}\frac{1}{2}(Y_{iT}-\mu-\alpha_{i}-\gamma_{T})^{2}-\frac{1}{NT}\frac{1}{2}(Y_{NT}-\mu-\alpha_{N}-\gamma_{T})^{2}.$
+Setting $\alpha_{N}=-\sum_{i=1}^{N-1}\alpha_{i}$ and $\gamma_{T}=-\sum_{t=1}^{T-1}\gamma_{t}$, the reparameterized likelihood is
+$l(\mu,\psi)$ $=-\frac{1}{2}\log(2\pi)-\frac{1}{NT}\sum_{i=1}^{N-1}\sum_{t=1}^{T-1}\frac{1}{2}(Y_{it}-\mu-\alpha_{i}-\gamma_{t})^{2}-\frac{1}{NT}\sum_{t=1}^{T-1}\frac{1}{2}\left(Y_{Nt}-\mu+\sum_{i=1}^{N-1}\alpha_{i}-\gamma_{t}\right)^{2}$
+$\quad-\frac{1}{NT}\sum_{i=1}^{N-1}\frac{1}{2}\left(Y_{iT}-\mu-\alpha_{i}+\sum_{t=1}^{T-1}\gamma_{t}\right)^{2}-\frac{1}{NT}\frac{1}{2}\left(Y_{NT}-\mu+\sum_{i=1}^{N-1}\alpha_{i}+\sum_{t=1}^{T-1}\gamma_{t}\right)^{2}.$
+Calculating the first derivative of $l(\mu,\psi)$ w.r.t. $\alpha_{i}$, we have
+$\partial_{\alpha_{i}}l(\mu,\psi)=\frac{1}{NT}\sum_{t=1}^{T}(Y_{it}-\mu-\alpha_{i})-\frac{1}{NT}\sum_{t=1}^{T}\left(Y_{Nt}-\mu+\sum_{i=1}^{N-1}\alpha_{i}\right)$
+for every $i=1,\ldots,N-1$. Equating $\partial_{\alpha_{i}}l(\mu,\psi)=0$, we obtain
+$\alpha_{i}=\frac{1}{T}\sum_{t=1}^{T}Y_{it}-\frac{1}{T}\sum_{t=1}^{T}Y_{Nt}-\sum_{i=1}^{N-1}\alpha_{i}.$ (A.5)
+Notice that (A.5) is independent of $\mu$ and all $\gamma_{t}$. Now, sum (A.5) over $i=1,\ldots,N-1$, we obtain
+$\sum_{i=1}^{N-1}\alpha_{i}=\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}Y_{it}-\frac{1}{T}\sum_{t=1}^{T}Y_{Nt}.$ (A.6)
+Combining (A.5) and (A.6), it follows that
+$\widehat{\alpha}_{i}=\frac{1}{T}\sum_{t=1}^{T}Y_{it}-\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}Y_{it},\qquad i=1,\ldots,N-1.$
+It can be calculated in the same way that
+$\widehat{\gamma}_{t}=\frac{1}{N}\sum_{i=1}^{N}Y_{it}-\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}Y_{it},\qquad t=1,\ldots,T-1.$
+Subsequently,
+$\widehat{\mu}=\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}Y_{it}.$ (A.7)
+In this simple example, $\widehat{\alpha}_{i}$ and $\widehat{\gamma}_{t}$ do not depend on $\mu$. This is generally not the case for complex models.
+In addition, maximizing $l(\mu,\psi)$ is identical to maximizing the original likelihood $\ell\left(\mu,\phi\right)$ under constraints $\sum_{i=1}^{N}\alpha_{i}=0$ and $\sum_{t=1}^{T}\gamma_{t}=0$.
+Specifically, it can be verified that the Lagrangian $\ell\left(\mu,\phi\right)+\lambda^{\alpha}\sum_{i=1}^{N}\alpha_{i}+\lambda^{\gamma}\sum_{t=1}^{T}\gamma_{t}$, where $\lambda^{\alpha}$ and $\lambda^{\gamma}$ are the two Lagrange-multipliers, delivers upon maximization the same $\widehat{\alpha}_{i}$ and $\widehat{\gamma}_{t}$ for $i=1,\ldots,N-1$ and $t=1,\ldots,T$, respectively.
+For $\widehat{\alpha}_{N}$ and $\widehat{\gamma}_{T}$, the Lagrangian produces
+$\widehat{\alpha}_{N}=\frac{1}{T}\sum_{t=1}^{T}Y_{Nt}-\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}Y_{it},\qquad\widehat{\gamma}_{T}=\frac{1}{N}\sum_{i=1}^{N}Y_{iT}-\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}Y_{it}.$
+Indeed, $\widehat{\alpha}_{N}=-\sum_{i=1}^{N-1}\widehat{\alpha}_{i}$ as in Equation (A.6), and it can be verified that $\widehat{\gamma}_{T}=-\sum_{t=1}^{T-1}\widehat{\gamma}_{t}$ as well.
+Example A.3 (heterogeneous mean model)
+Following the same setting in Example A.2 but let $\varepsilon_{it}$ be normal with mean 0 and variance $\sigma_{0}>0$.
 In the current example, $\theta_{0}=(\mu_{0},\sigma_{0})^{\prime}$ is the structural parameter of interests.
@@
 $\widetilde{Y}_{it}:=Y_{it}-\mathbb{E}Y_{it}$.
-Following the same steps in Example 2, we can derive
+Following the same steps in Example A.2, we can derive
 $\widehat{\alpha}_{i}-\alpha_{i}^{0}$ $=\frac{1}{T}\sum_{t=1}^{T}\widetilde{Y}_{it}-\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}\widetilde{Y}_{it},\qquad i=1,\ldots,N-1,$
-$\widehat{\gamma}_{t}-\gamma_{t}^{0}$ $=\frac{1}{N}\sum_{i=1}^{N}\widetilde{Y}_{it}-\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}\widetilde{Y}_{it},\qquad t=1,\ldots,T-1,$ (2.15)
+$\widehat{\gamma}_{t}-\gamma_{t}^{0}$ $=\frac{1}{N}\sum_{i=1}^{N}\widetilde{Y}_{it}-\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}\widetilde{Y}_{it},\qquad t=1,\ldots,T-1,$ (A.8)
 where $\alpha_{i}^{0}$ and $\gamma_{t}^{0}$ can be obtained by maximizing $\mathbb{E}l(\mu,\psi)$ in the same way. Denote additionally
@@
 It follows that
-$\widehat{\alpha}_{N}-\alpha_{N}^{0}=\frac{1}{T}\sum_{t=1}^{T}\widetilde{Y}_{Nt}-\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}\widetilde{Y}_{it},\qquad\widehat{\gamma}_{T}-\gamma_{T}^{0}=\frac{1}{N}\sum_{i=1}^{N}\widetilde{Y}_{iT}-\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}\widetilde{Y}_{it}.$ (2.16)
+$\widehat{\alpha}_{N}-\alpha_{N}^{0}=\frac{1}{T}\sum_{t=1}^{T}\widetilde{Y}_{Nt}-\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}\widetilde{Y}_{it},\qquad\widehat{\gamma}_{T}-\gamma_{T}^{0}=\frac{1}{N}\sum_{i=1}^{N}\widetilde{Y}_{iT}-\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}\widetilde{Y}_{it}.$ (A.9)
 Now, for $\psi^{0}:=({\alpha}_{1}^{0},\ldots,{\alpha}_{N-1}^{0},{\gamma}_{1}^{0},\ldots,{\gamma}_{T-1}^{0})^{\prime}$, $\widehat{\psi}:=(\widehat{\alpha}_{1},\ldots,\widehat{\alpha}_{N-1},\widehat{\gamma}_{1},\ldots,\widehat{\gamma}_{T-1})^{\prime}$, and $Y_{it}(\mu):=Y_{it}-\mu-\alpha_{i}^{0}-\gamma_{t}^{0}$, we can calculate
 $l(\theta,\widehat{\psi})-l(\theta,\psi^{0})$ $=\frac{1}{\sigma}\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}\left(Y_{it}(\mu)-\frac{1}{2}(\widehat{\alpha}_{i}-\alpha_{i}^{0})\right)(\widehat{\alpha}_{i}-\alpha_{i}^{0})$
-$\quad\quad+\frac{1}{\sigma}\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}\left(Y_{it}(\mu)-\frac{1}{2}(\widehat{\gamma}_{t}-\gamma_{t}^{0})\right)(\widehat{\gamma}_{t}-\gamma_{t}^{0}),$ (2.17)
-where $\theta=(\mu,\sigma)^{\prime}$ with $\sigma>0$. Substituting (2.15) and (2.16) into (2.17), we obtain, after some algebra,
+$\quad\quad+\frac{1}{\sigma}\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}\left(Y_{it}(\mu)-\frac{1}{2}(\widehat{\gamma}_{t}-\gamma_{t}^{0})\right)(\widehat{\gamma}_{t}-\gamma_{t}^{0}),$ (A.10)
+where $\theta=(\mu,\sigma)^{\prime}$ with $\sigma>0$. Substituting (A.8) and (A.9) into (A.10), we obtain, after some algebra,
 $l(\theta,\widehat{\psi})-l(\theta,\psi^{0})$ $=\frac{1}{2\sigma}\frac{1}{NT^{2}}\sum_{i=1}^{N}\sum_{t=1}^{T}\sum_{s=1}^{T}\widetilde{Y}_{it}\widetilde{Y}_{is}+\frac{1}{2\sigma}\frac{1}{N^{2}T}\sum_{i=1}^{N}\sum_{j=1}^{N}\sum_{t=1}^{T}\widetilde{Y}_{it}\widetilde{Y}_{jt}$
@@
 The bias term $B(\theta)$ causes the variance estimator to be biased. We illustrate this consequence here. Observe that
-$l(\theta,\widehat{\psi})$ $=-\frac{1}{2}\log(2\pi)-\frac{1}{2}\log\sigma-\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}\frac{1}{2\sigma}(Y_{it}-\mu-\widehat{\alpha}_{i}-\widehat{\gamma}_{t})^{2},$
-$\partial_{\sigma}l(\theta,\widehat{\psi})$ $=-\frac{1}{2\sigma}+\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}\frac{1}{2\sigma^{2}}(Y_{it}-\mu-\widehat{\alpha}_{i}-\widehat{\gamma}_{t})^{2}.$
-Following the same steps in Example 2, we obtain exactly the same estimator of $\mu_{0}$, $\widehat{\mu}$, as in Equation (2.10); and it is independent of $\sigma$.
+[…]
+[…]
+Following the same steps in Example A.2, we obtain exactly the same estimator of $\mu_{0}$, $\widehat{\mu}$, as in Equation (A.7); and it is independent of $\sigma$.
 Therefore, equating $\partial_{\sigma}l(\theta,\widehat{\psi})=0$, substituting $\mu$ with $\widehat{\mu}$, and solving for $\sigma$, the estimator of $\sigma^{0}$ is
@@
 as $N,T\rightarrow\infty$ with $N/T\rightarrow\kappa$.
-While $\widehat{\mu}$ is not affected by the IPP, the variance estimator $\widehat{\sigma}$ contains two non-negligible bias terms, $-\sigma_{0}/T$ and $-\sigma_{0}/N$.
-These bias terms cause $\sqrt{NT}(\widehat{\sigma}-\sigma_{0})$ to be centered at $-\kappa\sigma_{0}-\sigma_{0}/\kappa\neq 0$.
-3 Bias Correction and Asymptotic Theory
-The bias expansion of $l(\theta,\widehat{\psi}(\theta))$ in Equation (2.14) and Example 3 above motivate our approach of correcting the likelihood $l(\theta,\widehat{\psi}(\theta))$ by adding $B(\theta)$.
-The corrected likelihood can be viewed as an approximation to the IPP-free infeasible likelihood $l(\theta,\psi(\theta))$, with approximation error $o(1/\sqrt{NT})$.
-In Theorem 1 below we formalize the expansion in Equation (2.14).
-Specifically, we show that $B(\theta)$ can be further decomposed into $B(\theta)=B_{\alpha}(\theta)+B_{\gamma}(\theta)+o(1/\sqrt{NT})$, in which $B_{\alpha}(\theta)$ appears because of the estimation errors of the individual-specific heterogeneities $\alpha_{i}^{0}$, while $B_{\gamma}(\theta)$ is present due to the estimation errors of the time-specific heterogeneities $\gamma_{t}^{0}$.
-The exact formulas for both terms are given below.
-Theorem 2 follows to establish the corrected likelihood as
-$L(\theta):=\widehat{l}(\theta)+\widehat{B}_{\alpha}(\theta)+\widehat{B}_{\gamma}(\theta)$ (3.1)
-where we denote $\widehat{l}(\theta):=l(\theta,\widehat{\psi}(\theta))$ for simplicity from now on.
-$\widehat{B}_{\alpha}(\theta)$ and $\widehat{B}_{\gamma}(\theta)$ are estimates of $B_{\alpha}(\theta)$ and $B_{\gamma}(\theta)$, respectively.
-The analytical formulas for $\widehat{B}_{\alpha}(\theta)$ and $\widehat{B}_{\gamma}(\theta)$ are provided below.
-We then show that
-$L(\theta)-l(\theta)=o_{\mathbb{P}}(1/\sqrt{NT}),$
-where we use $l(\theta):=l(\theta,\psi(\theta))$ for simplicity below.
-This implies that $L(\theta)$ does not suffer from the IPP.
-The corrected estimator of $\theta_{0}$, and the LR, LM, and Wald test statistics, may subsequently be derived from $L\left(\theta\right)$, instead of $\widehat{l}(\theta)$.
-Specifically, the corrected estimator of $\theta_{0}$ is
-$\widehat{\theta}_{L}:=\arg\max_{\theta}L(\theta).$
-Under mild regularity conditions, $\widehat{\theta}_{L}$ is asymptotically unbiased, and its asymptotic properties align with the standard results of the classical MLE without the nuisance parameters.
-Moreover, the test statistics constructed from the corrected likelihood such as the Wald, LR, and LM also follow the limiting $\chi^{2}$ distribution under the null hypothesis.
-These results are formally stated in Corollaries 1 and 2, respectively.
-Remark 2 (computation)
-The computation of $\widehat{\theta}_{L}$ involves a two-layer maximization but can be done efficiently.
-In the inner layer, $\widehat{\psi}(\theta)$ is obtained by maximizing $l(\theta,\psi)$, keeping $\theta$ fixed.
-This is large-scale maximization problem and a few strategies can be exploited to expedite the procedure:
-  1. 1.
-    For many commonly used models, the first and second derivatives of $l(\theta,\psi)$ have clear analytical expressions, so that the standard Newton-Raphson algorithm can be employed to obtain $\widehat{\psi}(\theta)$.
-  2. 2.
-    Many models (specifically, generalized linear models) are equipped with linear indices, as in Example 1.
-In that case, the chain-rule-based technique in Sun and Dhaene (2019) can be used to achieve a fast calculation of the derivatives.
-The computational complexity of such a technique only grows with the number of linear indices, not the dimension of $\psi$, even without analytical expressions for the derivatives of $l(\theta,\psi)$.
-  3. 3.
-    The derivatives of the reparameterized $l(\theta,\psi)$ can be identically constructed from the derivatives of the original likelihood $\ell(\theta,\phi)$.
-Specifically, the reparameterization implies
-    $\partial_{\psi}l(\theta,\psi)=D\partial_{\phi}\ell(\theta,\phi),\qquad\partial_{\psi\psi^{\prime}}l(\theta,\psi)=D\partial_{\phi\phi^{\prime}}\ell(\theta,\phi)D^{\prime},$
-    holding true for every $\phi=D^{\prime}\psi$ and every $\theta$.
-In the outer layer where $L(\theta)$ is maximized w.r.t.
-$\theta$, it can be tedious to obtain analytical formulas for the derivatives of the corrected likelihood $L(\theta)$.
-However, the dimension of $\theta$ is usually small, so that invoking the Newton-Raphson algorithm with numerical derivatives generally does not suffer from too much computational pressure, provided that the inner-layer maximization is done efficiently.
-Now, we proceed to our bias correction procedure on the likelihood and present the related asymptotic theory.
-Denote $\widehat{\phi}(\theta):=D^{\prime}\widehat{\psi}(\theta)$ and $\phi(\theta):=D^{\prime}\psi(\theta)$, which are essentially the “unparameterized” counterparts of $\widehat{\psi}(\theta)$ and $\psi(\theta)$, respectively.
-We impose the following assumption.
+While $\widehat{\mu}$ is not affected by the IPP, the variance estimator $\widehat{\sigma}$ contains two non-negligible bias terms, $\sigma_{0}/T$ and $\sigma_{0}/N$.
+These bias terms cause $\sqrt{NT}(\widehat{\sigma}-\sigma_{0})$ to be centered at $\kappa\sigma_{0}-\sigma_{0}/\kappa\neq 0$.
+B Remarks
+Remark 4 (uniform inference)
+The inference based on our corrected likelihood is uniformly valid regardless of whether the underlying data-generating process has a homogeneous or heterogeneous slope coefficient.
+This is similar to Lu and Su (2023), except that they consider a linear regression.
+For instance, consider the data-generating process
+$Y_{it}=\mathds{1}\{\left(\beta_{0}+\alpha_{1,i}^{0}+\gamma_{1,t}^{0}\right)X_{it}+\alpha_{2,i}^{0}+\gamma_{2,t}^{0}+\varepsilon_{it}>0\},$
+where the slope coefficient of $X_{it}$ can be homogeneous, if $\alpha_{1,i}^{0}=\gamma_{1,t}^{0}=0$ for all $i$ and $t$, or heterogeneous, if otherwise.
+In either case, inference based on the corrected likelihood under the model
+$Y_{it}=\mathds{1}\{(\beta+\alpha_{1,i}+\gamma_{1,t})X_{it}+\alpha_{2,i}+\gamma_{2,t}+\varepsilon_{it}>0\},\ \alpha_{1,i}\neq 0\ \text{and }\gamma_{1,t}\neq 0,$
+remains valid.
+This uniform validity shares a similar spirit as the uniform inference in Lu and Su (2023).
+In fact, the homogeneous data-generating process is just a special case of the heterogeneous one.
+By employing the bias-corrected likelihood for the heterogeneous model, inference on $\beta$ remains valid, albeit with a possible loss of efficiency when the true data-generating process is homogeneous.
+In addition, we treat the heterogeneous variables $\alpha_{k,i}^{0}$ and $\gamma_{k,t}^{0}$ as parameters by conditioning on them, which is common for panel maximum likelihood methods, we do not further distinguish between locally heterogeneous and fully heterogeneous cases as in Lu and Su (2023).
+Remark 5 (expectation)
+Our definition of $\mathbb{E}$ depends on the model specification and is the same to Arellano and Bonhomme (2009) and Fernández-Val and Weidner (2016) .
+Consider the frequently used autoregressive model with $X_{it}=(Y_{it-1},Z_{it}^{\prime})^{\prime}$ for some strictly exogenous regressor $Z_{it}$.
+It is common that, for each $i=1,\ldots,N$, the density of $Y_{i1}$ is missing and (2.2) becomes a conditional likelihood as in Chapter 5 of Hamilton (1994).
+In this case, the expectation is taken w.r.t.
+$f_{Y_{i2},\ldots,Y_{iT}}\left(\cdot|Y_{i1},Z_{i1},\ldots,Z_{iT},\phi^{0};\theta_{0}\right)$, which is the density of only $Y_{i2},\ldots,Y_{iT}$, conditional on the initial value $Y_{i1}$, strictly exogenous regressors $Z_{i1},\ldots,Z_{iT}$, and the unobserved effects $\phi^{0}$.
+However, when the density of $Y_{i1}$ is indeed specified, so that (2.2) is an exact likelihood, the expectation is taken w.r.t.
+$f_{Y_{i1},\ldots,Y_{iT}}\left(\cdot|Z_{i1},\ldots,Z_{iT},\phi^{0};\theta_{0}\right)$, which is the joint density of all $Y_{i1},\ldots,Y_{iT}$, conditioning on the strictly exogenous regressors $Z_{i1},\ldots,Z_{iT}$ and the unobserved effects $\phi^{0}$ only.
+Remark 6 (truncation)
+In the estimator $\widehat{S}_{\alpha\alpha}(\theta)$, we use the truncation technique commonly found in e.g., Hahn and Kuersteiner (2011) (see also Hahn and Kuersteiner (2007)), Fernández-Val and Weidner (2016), and Galvao and Kato (2016).
+This is related to the literature on consistent estimations of autocovariances, $\mathbb{E\{}\widetilde{s}_{i,t}^{\alpha}(\theta)[\widetilde{s}_{i,s}^{\alpha}(\theta)]\}$ in our case.
+Detailed discussions in this regard can be found in Andrews (1991).
+Without truncation, we would have $\widehat{S}_{\alpha\alpha}(\theta)=\widehat{s}_{\alpha}(\theta)[\widehat{s}_{\alpha}(\theta)]^{\prime}=0$ for every $\theta$ because $\widehat{s}_{\alpha}(\theta)=0$ by definition, where $\widehat{s}_{\alpha}(\theta)$ is defined from $\widehat{s}(\theta):=s(\theta,\widehat{\psi}(\theta))$ under the same partition in (3.2).
+Here, the truncation parameter $\tau\rightarrow\infty$ but sufficiently slower relative to $T$ (see Theorem 2).
+Hahn and Kuersteiner (2011) suggests the default value of $\tau=1$ for small $T$.
+For intermediate $T$, we recommend following the advice from Fernández-Val and Weidner (2016) to conduct a sensitivity test for different values of $\tau$ and to avoid using very large $\tau$ values.
+In our simulation, we use $\tau=1$ and $2$ and find the difference relatively insignificant.
+On the other hand, the term $\widehat{S}_{\gamma\gamma}(\theta)$ does not require any truncation, because we have already eliminated $\mathbb{E\{}\widetilde{s}_{i}^{\gamma}(\theta)[\widetilde{s}_{j}^{\gamma}(\theta)]^{\prime}\}$ for $i\neq j$ from Equation (C.1) using cross-sectional independence.
+One may still opt to truncate $\widehat{\mathcal{S}}_{\gamma\gamma}(\theta)$, such that all elements beyond the $\tau$-th sub- and super-diagonal are symmetrically set to 0.
+We have experimented with such truncations but found little impact on $\widehat{\theta}_{L}$ and test statistics.
+The comparison is, hence, not reported in the paper.
+C Technical Details
+C.1 Bias Terms
+The reparameterization ensures the following identity:
+$S_{\alpha\alpha}(\theta)=D_{1}\mathcal{S}_{\alpha\alpha}(\theta)D_{1}^{\prime},$
+where $\mathcal{S}_{\alpha\alpha}(\theta)$ is the outer product of the $NK\times 1$ vector of $\partial_{\alpha}\ell(\theta,\phi)-\mathbb{E}\partial_{\alpha}\ell(\theta,\phi)$ evaluated at $\phi=\phi(\theta)$, where $\partial_{\alpha}\ell(\theta,\phi)$ is the derivative of the original likelihood $\ell(\theta,\phi)$ w.r.t.
+$\alpha:=(\alpha_{1}^{\prime},\ldots,\alpha_{N}^{\prime})^{\prime}$.
+Specifically,
+$\mathcal{S}_{\alpha\alpha}(\theta)$ $:=\frac{1}{N^{2}T^{2}}\sum_{t=1}^{T}\sum_{s=1}^{T}\mathbb{E\{}\widetilde{s}_{t}^{\alpha}(\theta)[\widetilde{s}_{s}^{\alpha}(\theta)]^{\prime}\},\qquad\widetilde{s}_{t}^{\alpha}(\theta):=\{[\widetilde{s}_{1,t}^{\alpha}(\theta)]^{\prime},\ldots,[\widetilde{s}_{N,t}^{\alpha}(\theta)]^{\prime}\}^{\prime},$
+$\widetilde{s}_{i,t}^{\alpha}(\theta)$ $:=s_{i,t}^{\alpha}(\theta,\phi(\theta))-\mathbb{E}s_{i,t}^{\alpha}(\theta,\phi(\theta)).$
+Under cross-sectional independence, $\mathcal{S}_{\alpha\alpha}(\theta)$ simplifies to
+$\mathcal{S}_{\alpha\alpha}(\theta)=\frac{1}{N^{2}T^{2}}\sum_{t=1}^{T}\sum_{s=1}^{T}\mathbb{E}\operatorname*{diag}\{\widetilde{s}_{1,t}^{\alpha}(\theta)[\widetilde{s}_{1,s}^{\alpha}(\theta)]^{\prime},\ldots,\widetilde{s}_{N,t}^{\alpha}(\theta)[\widetilde{s}_{N,s}^{\alpha}(\theta)]^{\prime}\}.$
+Intuitively, $\widehat{S}_{\alpha\alpha}(\theta)$ in (3.6) is a natural plug-in estimator of $S_{\alpha\alpha}(\theta)$.
+Similarly for $S_{\gamma\gamma}(\theta)$, the reparameterization produces the identity:
+$S_{\gamma\gamma}(\theta)=D_{2}\mathcal{S}_{\gamma\gamma}(\theta)D_{2}^{\prime},$
+where $\mathcal{S}_{\gamma\gamma}(\theta)$ is the outer product of the $TK\times 1$ vector of $\partial_{\gamma}\ell(\theta,\phi)-\mathbb{E}\partial_{\gamma}\ell(\theta,\phi)$, for $\gamma:=(\gamma_{1}^{\prime},\ldots,\gamma_{T}^{\prime})^{\prime}$, evaluated at $\phi=\phi(\theta)$:
+$\mathcal{S}_{\gamma\gamma}(\theta)$ $=\frac{1}{N^{2}T^{2}}\sum_{i=1}^{N}\sum_{j=1}^{N}\mathbb{E\{}\widetilde{s}_{i}^{\gamma}(\theta)[\widetilde{s}_{j}^{\gamma}(\theta)]^{\prime}\},\qquad\widetilde{s}_{i}^{\gamma}(\theta):=\{[\widetilde{s}_{i,1}^{\gamma}(\theta)]^{\prime},\ldots,[\widetilde{s}_{i,T}^{\gamma}(\theta)]^{\prime}\}^{\prime},$
+$\widetilde{s}_{i,t}^{\gamma}(\theta)$ $:=s_{i,t}^{\gamma}(\theta,\phi(\theta))-\mathbb{E}s_{i,t}^{\gamma}(\theta,\phi(\theta)).$
+By cross-sectional independence, $\mathcal{S}_{\gamma\gamma}(\theta)$ similarly reduces to
+$\mathcal{S}_{\gamma\gamma}(\theta)=\frac{1}{N^{2}T^{2}}\sum_{i=1}^{N}\mathbb{E\{}s_{i}^{\gamma}(\theta)[s_{i}^{\gamma}(\theta)]^{\prime}\}.$ (C.1)
+Naturally, $\widehat{S}_{\gamma\gamma}(\theta)$ in (3.7) is also an intuitive plug-in estimator of $S_{\gamma\gamma}(\theta)$.
+C.2 Regularity Conditions
+Denote $\phi(\theta):=D^{\prime}\psi(\theta)$, which is the “unparameterized” counterpart of $\psi(\theta)$.
+We impose the following assumption to establish the asymptotic theory for the expansion in (3.1) and for the corrected likelihood $L(\theta)$.
 Assumption 1
@@
     where $\phi_{s}$ represents any element of $\phi$ and $S\in\{0,1,2,3\}$ ($S=0$ is understood as not taking derivatives), and
-    $\sup_{N,T}\max_{1\leq i\leq N,1\leq t\leq T}\mathbb{E}\{[g(w_{it})]^{\eta}\}<\infty,$
-    for some $\eta>2$.
+    $\sup_{N,T}\max_{1\leq i\leq N,1\leq t\leq T}\mathbb{E}_{\phi}\{[g(w_{it})]^{\eta}\}<\infty$
+    for some $\eta>2,$ where $\mathbb{E}_{\phi}$ denotes the conditional expectation w.r.t. the joint distribution of $w_{it}$, given the heterogeneous effects $\phi^{0}$.
   • (iii) For each $k=1,\ldots,K$, $\sum_{i=1}^{N}\alpha_{k,i}^{0}=\sum_{t=1}^{T}\gamma_{k,t}^{0}=0$, where $\alpha_{k,i}^{0}$ and $\gamma_{k,t}^{0}$ are the $k$-th component of $\alpha_{i}^{0}$ and $\gamma_{t}^{0}$, respectively.
@@
     where $\mathcal{A}_{it}$ and $\mathcal{B}_{it}$ are the $\sigma$-algebras generated by $\{w_{is}:1\leq s\leq t\}$ and $\{w_{is}:t\leq s\leq T\}$, respectively.
-  • (vi) The matrix has $\sqrt{NT}\mathbb{E}H(\theta)$ has eigenvalues $\lambda_{p}(\theta)$ for $p=1,\ldots,(N+T-2)K$ which satisfy
+  • (vi) The matrix $\sqrt{NT}\mathbb{E}[H(\theta)]$ has eigenvalues $\lambda_{p}(\theta)$ for $p=1,\ldots,(N+T-2)K$ which satisfy
     $\sup_{N>N_{0},T>T_{0}}\max_{1\leq p\leq(N+T-2)K}\sup_{\theta\in\Theta}\lambda_{p}(\theta)<0$
-    for some integers $N_{0}$ and $T_{0}$.
-Here Assumption 1 is the same to Fernández-Val and Weidner (2016).
+    a.s. for some integers $N_{0}$ and $T_{0}$.
+Assumption 1 is the same to Fernández-Val and Weidner (2016).
 It is used to ensure that the bias $\sqrt{NT}B(\theta)$ does not explode as $N,T\rightarrow\infty$, as discussed above.
@@
 The uniform convergence of $L(\theta)$ guarantees that the bias correction effect carries over to both the estimators and likelihood-based test statistics.
-Assumption 1 states that the true effects $\alpha_{i}^{0}$ and $\gamma_{t}^{0}$ sums up to 00 over $i=1,\ldots,N$ and $t=1,\ldots,T$.
+Assumption 1 states that the true effects $\alpha_{i}^{0}$ and $\gamma_{t}^{0}$ sums up to 0 over $i=1,\ldots,N$ and $t=1,\ldots,T$.
 Given the additive structure of $\theta_{0}+\alpha_{i}^{0}+\gamma_{t}^{0}$ in (2.1), this condition is without loss of generality because we merely let $\theta_{0}$ capture the average effects of $\alpha_{i}^{0}$ and $\gamma_{t}^{0}$.
@@
 Without conditioning, $w_{it}$ may not be, for instance, independent over $i$, because $w_{it}$ and $w_{jt}$ may depend on the same $\gamma_{t}^{0}$.
-In addition, we only require the mixing coefficient $a_{i}(m)$ to decay at a polynomial rate such that $[a_{i}(m)]^{1-2/\eta}$ is summable as $m\rightarrow\infty$.
+In addition, we only require the mixing coefficient $a_{i}(m)$ to decay at a polynomial rate such that $a_{i}(m)]^{1-2/\eta}$ is summable as $m\rightarrow\infty$.
 The mixing condition is used to guarantee that certain autocovariances which appear in our proof are summable by Davydov’s inequality.
@@
 These are generally not possible because $w_{it}$ depends on $\alpha_{i}^{0}$ and $\gamma_{t}^{0}$.
-Finally, Assumption 1 simply requires that the normalized expected Hessian $\sqrt{NT}\mathbb{E}H(\theta)$ has eigenvalues $\lambda_{p}(\theta)$ that are strictly less than 00 for all $N>N_{0}$ and $T>T_{0}$.
+Finally, Assumption 1 simply requires that the normalized expected Hessian $\sqrt{NT}\mathbb{E}H(\theta)$ has eigenvalues $\lambda_{p}(\theta)$ that are strictly less than 0 for all $N>N_{0}$ and $T>T_{0}$.
 This ensures that $\mathbb{E\{}s^{\prime}(\theta)[H(\theta)]^{-1}s(\theta)\}$ is well defined for large $N$ and $T$.
-This is a mild condition in that we only impose it on the expected Hessian $\mathbb{E}H(\theta)$ of the reparameterized likelihood.
-We state the condition with “less than 00” because the Hessian in a maximum likelihood model is usually negative definite.
-Next, partition [^5] Note that $s(\theta)=s(\theta)-\mathbb{E}s\left(\theta\right)$ because $\mathbb{E}s\left(\theta\right)=0$ by definition.
-We keep $\mathbb{E}s\left(\theta\right)$ here for the convenience of constructing the corrected likelihood.
-$s(\theta)-\mathbb{E}s\left(\theta\right)$ and $[\mathbb{E}H(\theta)]^{-1}$ as
-$s(\theta)-\mathbb{E}s\left(\theta\right)=:\left(\begin{array}[c]{c}\widetilde{s}_{\alpha}(\theta)\\ \widetilde{s}_{\gamma}(\theta)\end{array}\right),\qquad[\mathbb{E}H(\theta)]^{-1}=:\left(\begin{array}[c]{cc}H_{\alpha\alpha}^{\mathcal{\ast}}(\theta)&H_{\alpha\gamma}^{\mathcal{\ast}}(\theta)\\ H_{\gamma\alpha}^{\mathcal{\ast}}(\theta)&H_{\gamma\gamma}^{\mathcal{\ast}}(\theta)\end{array}\right),$ (3.2)
-where $\widetilde{s}_{\alpha}(\theta)$ is $(N-1)K\times 1$ and $\widetilde{s}_{\gamma}(\theta)$ is $(T-1)K\times 1$ while $H_{\alpha\alpha}^{\mathcal{\ast}}(\theta)$ is $(N-1)K\times(N-1)K$ and $H_{\gamma\gamma}^{\ast}(\theta)$ is $(T-1)K\times(T-1)K$.
-The off-diagonal blocks $H_{\alpha\gamma}^{\ast}(\theta)$ and $H_{\gamma\alpha}^{\mathcal{\ast}}(\theta)$ are defined accordingly.
-Further denote
-$S_{\alpha\alpha}(\theta):=\mathbb{E\{}\widetilde{s}_{\alpha}(\theta)\mathbb{[}\widetilde{s}_{\alpha}(\theta)]^{\prime}\},\qquad S_{\gamma\gamma}(\theta):=\mathbb{E\{}\widetilde{s}_{\gamma}(\theta)\mathbb{[}\widetilde{s}_{\gamma}(\theta)]^{\prime}\}.$
-The next theorem formally states the asymptotic expansion of $\widehat{l}(\theta)$.
-Theorem 1
-Under Assumption 1 and uniformly over $\theta\in\Theta$,
-$\mathbb{E}l(\theta)=\mathbb{E}\widehat{l}(\theta)+B_{\alpha}(\theta)+B_{\gamma}(\theta)+o(1/\sqrt{NT}),$
-as $N,T\rightarrow\infty$, where
-$B_{\alpha}(\theta):=\frac{1}{2}\operatorname*{trace}[S_{\alpha\alpha}(\theta)H_{\alpha\alpha}^{\mathcal{\ast}}(\theta)],\qquad B_{\gamma}(\theta):=\frac{1}{2}\operatorname*{trace}\mathbb{[}S_{\gamma\gamma}(\theta)H_{\gamma\gamma}^{\ast}(\theta)],$
-and $B_{\alpha}(\theta)+B_{\gamma}(\theta)=O(1/\sqrt{NT})$.
-[…]
-The bias terms $B_{\alpha}(\theta)$ and $B_{\gamma}(\theta)$ involves $\psi(\theta)$ and $\mathbb{E}$ which are infeasible to compute.
-However, $B_{\alpha}(\theta)$ and $B_{\gamma}(\theta)$ can be estimated by plug-in estimates of $H_{\alpha\alpha}^{\mathcal{\ast}}(\theta)$, $H_{\gamma\gamma}^{\ast}(\theta)$, $S_{\alpha\alpha}(\theta)$, and $S_{\gamma\gamma}(\theta)$.
-The two Hessian terms can be estimated in a straightforward way.
-Denote $\widehat{H}(\theta):=H(\theta,\widehat{\psi}(\theta))$ and partition $[\widehat{H}(\theta)]^{-1}$ into
-$[\widehat{H}(\theta)]^{-1}=:\left(\begin{array}[c]{cc}\widehat{H}_{\alpha\alpha}^{\mathcal{\ast}}(\theta)&\widehat{H}_{\alpha\gamma}^{\mathcal{\ast}}(\theta)\\ \widehat{H}_{\gamma\alpha}^{\mathcal{\ast}}(\theta)&\widehat{H}_{\gamma\gamma}^{\mathcal{\ast}}(\theta)\end{array}\right)$ (3.3)
-in the same manner as in (3.2).
-The term $H_{\alpha\alpha}^{\mathcal{\ast}}(\theta)$ can be estimated by the $(N-1)K\times(N-1)K$ block $\widehat{H}_{\alpha\alpha}^{\mathcal{\ast}}(\theta)$, and $H_{\gamma\gamma}^{\mathcal{\ast}}(\theta)$ by the $(T-1)K\times(T-1)K$ block $\widehat{H}_{\gamma\gamma}^{\mathcal{\ast}}(\theta)$.
-Estimators of $S_{\alpha\alpha}(\theta)$ and $S_{\gamma\gamma}(\theta)$ are constructed from the derivatives of $l_{i,t}(\theta,\phi)$, the original likelihood for individual $i$ at time $t$.
-This procedure needs some explanation.
-The reparameterization ensures the following identity:
-$S_{\alpha\alpha}(\theta)=D_{1}\mathcal{S}_{\alpha\alpha}(\theta)D_{1}^{\prime},$
-where $\mathcal{S}_{\alpha\alpha}(\theta)$ is the outer product of the $NK\times 1$ vector of $\partial_{\alpha}\ell(\theta,\phi)-\mathbb{E}\partial_{\alpha}\ell(\theta,\phi)$ evaluated at $\phi=\phi(\theta)=D^{\prime}\psi(\theta)$, where $\partial_{\alpha}\ell(\theta,\phi)$ is the derivative of the original likelihood $\ell(\theta,\phi)$ w.r.t.
-$\alpha:=(\alpha_{1}^{\prime},\ldots,\alpha_{N}^{\prime})^{\prime}$.
-Specifically,
-$\mathcal{S}_{\alpha\alpha}(\theta)$ $:=\frac{1}{N^{2}T^{2}}\sum_{t=1}^{T}\sum_{s=1}^{T}\mathbb{E\{}\widetilde{s}_{t}^{\alpha}(\theta)[\widetilde{s}_{s}^{\alpha}(\theta)]^{\prime}\},\qquad\widetilde{s}_{t}^{\alpha}(\theta):=\{[\widetilde{s}_{1,t}^{\alpha}(\theta)]^{\prime},\ldots,[\widetilde{s}_{N,t}^{\alpha}(\theta)]^{\prime}\}^{\prime},$
-$\widetilde{s}_{i,t}^{\alpha}(\theta)$ $:=s_{i,t}^{\alpha}(\theta,\phi(\theta))-\mathbb{E}s_{i,t}^{\alpha}(\theta,\phi(\theta)),\qquad s_{i,t}^{\alpha}(\theta,\phi):=\partial_{\alpha_{i}}l_{i,t}(\theta,\phi).$
-Here, $\widetilde{s}_{i,t}^{\alpha}(\theta)$ is the centered version of $s_{i,t}^{\alpha}(\theta,\phi(\theta))$, the derivative, w.r.t.
-$\alpha_{i}$ and evaluated at $\phi(\theta)$, of the original likelihood $l_{i,t}(\theta,\phi)$ for individual $i$ at time $t$.
-Under cross-sectional independence, $\mathcal{S}_{\alpha\alpha}(\theta)$ simplifies to
-$\mathcal{S}_{\alpha\alpha}(\theta)=\frac{1}{N^{2}T^{2}}\sum_{t=1}^{T}\sum_{s=1}^{T}\mathbb{E}\operatorname*{diag}\{\widetilde{s}_{1,t}^{\alpha}(\theta)[\widetilde{s}_{1,s}^{\alpha}(\theta)]^{\prime},\ldots,\widetilde{s}_{N,t}^{\alpha}(\theta)[\widetilde{s}_{N,s}^{\alpha}(\theta)]^{\prime}\}.$
-Naturally, the estimator of $S_{\alpha\alpha}(\theta)$ is
-$\widehat{S}_{\alpha\alpha}(\theta)$ $:=D_{1}\widehat{\mathcal{S}}_{\alpha\alpha}(\theta)D_{1}^{\prime},$ (3.4)
-$\widehat{\mathcal{S}}_{\alpha\alpha}(\theta)$ $:=\frac{1}{N^{2}T^{2}}\sum_{t=1}^{T}\sum_{s=1}^{T}\mathds{1}\{\left|t-s\right|\leq\tau\}\operatorname*{diag}\{\widehat{s}_{1,t}^{\alpha}(\theta)[\widehat{s}_{1,s}^{\alpha}(\theta)]^{\prime},\ldots,\widehat{s}_{N,t}^{\alpha}(\theta)[\widehat{s}_{N,s}^{\alpha}(\theta)]^{\prime}\},$
-$\widehat{s}_{i,t}^{\alpha}(\theta)$ $:=s_{i,t}^{\alpha}(\theta,\widehat{\phi}(\theta))-\frac{1}{T}\sum_{t=1}^{T}s_{i,t}^{\alpha}(\theta,\widehat{\phi}(\theta)).$
-The indicator $\mathds{1}\{\left|t-s\right|\leq\tau\}$ is a truncation mechanism with truncation parameter $\tau$ (see Remark 3 below).
-Similarly for $S_{\gamma\gamma}(\theta)$, the reparameterization produces the identity:
-$S_{\gamma\gamma}(\theta)=D_{2}\mathcal{S}_{\gamma\gamma}(\theta)D_{2}^{\prime},$
-where $\mathcal{S}_{\gamma\gamma}(\theta)$ is the outer product of the $TK\times 1$ vector of $\partial_{\gamma}\ell(\theta,\phi)-\mathbb{E}\partial_{\gamma}\ell(\theta,\phi)$, for $\gamma:=(\gamma_{1}^{\prime},\ldots,\gamma_{T}^{\prime})^{\prime}$, evaluated at $\phi=\phi(\theta)$:
-$\mathcal{S}_{\gamma\gamma}(\theta)$ $=\frac{1}{N^{2}T^{2}}\sum_{i=1}^{N}\sum_{j=1}^{N}\mathbb{E\{}\widetilde{s}_{i}^{\gamma}(\theta)[\widetilde{s}_{j}^{\gamma}(\theta)]^{\prime}\},\qquad\widetilde{s}_{i}^{\gamma}(\theta):=\{[\widetilde{s}_{i,1}^{\gamma}(\theta)]^{\prime},\ldots,[\widetilde{s}_{i,T}^{\gamma}(\theta)]^{\prime}\}^{\prime},$
-$\widetilde{s}_{i,t}^{\gamma}(\theta)$ $:=s_{i,t}^{\gamma}(\theta,\phi(\theta))-\mathbb{E}s_{i,t}^{\gamma}(\theta,\phi(\theta)),\qquad s_{i,t}^{\gamma}(\theta,\phi):=\partial_{\gamma_{t}}l_{i,t}(\theta,\phi),$
-where $\widetilde{s}_{i,t}^{\gamma}(\theta)$ is the centered version of $s_{i,t}^{\gamma}(\theta,\phi(\theta))$, the derivative of $l_{i,t}(\theta,\phi)$ w.r.t.
-$\gamma_{t}$ and evaluated at $\phi(\theta)$.
-By cross-sectional independence, $\mathcal{S}_{\gamma\gamma}(\theta)$ similarly reduces to
-$\mathcal{S}_{\gamma\gamma}(\theta)=\frac{1}{N^{2}T^{2}}\sum_{i=1}^{N}\mathbb{E\{}s_{i}^{\gamma}(\theta)[s_{i}^{\gamma}(\theta)]^{\prime}\}.$ (3.5)
-Intuitively, the estimator for $S_{\gamma\gamma}(\theta)$ is
-$\widehat{S}_{\gamma\gamma}(\theta)$ $=D_{2}\widehat{\mathcal{S}}_{\gamma\gamma}(\theta)D_{2}^{\prime},$ (3.6)
-$\widehat{\mathcal{S}}_{\gamma\gamma}(\theta)$ $:=\frac{1}{N^{2}T^{2}}\sum_{i=1}^{N}\mathbb{E\{}\widehat{s}_{i}^{\gamma}(\theta)[\widehat{s}_{i}^{\gamma}(\theta)]^{\prime}\},\qquad\widehat{s}_{i}^{\gamma}(\theta):=\{[\widehat{s}_{i,1}^{\gamma}(\theta)]^{\prime},\ldots,[\widehat{s}_{i,T}^{\gamma}(\theta)]^{\prime}\}^{\prime},$
-$\widehat{s}_{i,t}^{\gamma}(\theta)$ $:=s_{i,t}^{\gamma}(\theta,\widehat{\phi}(\theta))-\frac{1}{T}\sum_{t=1}^{T}s_{i,t}^{\gamma}(\theta,\widehat{\phi}(\theta)).$
-Remark 3 (truncation)
-In the estimator $\widehat{S}_{\alpha\alpha}(\theta)$, we use the truncation technique commonly found in e.g., Hahn and Kuersteiner (2011) (see also Hahn and Kuersteiner (2007)), Fernández-Val and Weidner (2016), and Galvao and Kato (2016).
-This is related to the literature on consistent estimations of autocovariances, $\mathbb{E\{}\widetilde{s}_{i,t}^{\alpha}(\theta)[\widetilde{s}_{i,s}^{\alpha}(\theta)]\}$ in our case.
-Detailed discussions in this regard can be found in Andrews (1991).
-Without truncation, we would have $\widehat{S}_{\alpha\alpha}(\theta)=\widehat{s}_{\alpha}(\theta)[\widehat{s}_{\alpha}(\theta)]^{\prime}=0$ for every $\theta$ because $\widehat{s}_{\alpha}(\theta)=0$ by definition, where $\widehat{s}_{\alpha}(\theta)$ is defined from $\widehat{s}(\theta):=s(\theta,\widehat{\psi}(\theta))$ under the same partition in (3.2).
-Here, the truncation parameter $\tau\rightarrow\infty$ but sufficiently slower relative to $T$ (see Theorem 2 below).
-Hahn and Kuersteiner (2011) suggests the default value of $\tau=1$ for small $T$.
-For intermediate $T$, we recommend following the advices from Fernández-Val and Weidner (2016) to conduct a sensitivity test for different values of $\tau$ and to avoid using very large $\tau$ values.
-In our simulation, we use $\tau=1$ and $2$ and find the difference relatively insignificant.
-On the other hand, the term $\widehat{S}_{\gamma\gamma}(\theta)$ does not require any truncation, because we have already eliminated $\mathbb{E\{}\widetilde{s}_{i}^{\gamma}(\theta)[\widetilde{s}_{j}^{\gamma}(\theta)]^{\prime}\}$ for $i\neq j$ from Equation (3.5) using cross-sectional independence.
-One may still opt to truncate $\widehat{\mathcal{S}}_{\gamma\gamma}(\theta)$, such that all elements beyond the $\tau$-th sub- and super-diagonal are symmetrically set to 00.
-We have experimented with such truncations but found little impact on $\widehat{\theta}_{L}$ and test statistics.
-The comparison is, hence, not reported in the paper.
-Using Equations (3.3), (3.4), and (3.6), we can construct the corrected likelihood $L(\theta)$, which is stated in the theorem below with its rate of convergence.
-Theorem 2
-Under Assumption 1 and for some $\tau\rightarrow\infty$ such that $\tau/T\rightarrow 0$,
-$L(\theta)-l(\theta)=o_{\mathbb{P}}(1/\sqrt{NT}),$
-uniformly over $\theta\in\Theta$ as $N,T\rightarrow\infty$, where
-$L(\theta):=\widehat{l}(\theta)+\widehat{B}_{\alpha}(\theta)+\widehat{B}_{\gamma}(\theta),$
-as in (3.1), with
-$\widehat{B}_{\alpha}(\theta):=\frac{1}{2}\operatorname*{trace}[\widehat{S}_{\alpha\alpha}(\theta)\widehat{H}_{\alpha\alpha}^{\mathcal{\ast}}(\theta)],\qquad\widehat{B}_{\gamma}(\theta):=\frac{1}{2}\operatorname*{trace}\mathbb{[}\widehat{S}_{\gamma\gamma}(\theta)\widehat{H}_{\gamma\gamma}^{\ast}(\theta)].$
-As mentioned above, Theorem 2 states that $L(\theta)$ does not suffer from the IPP.
-Showing 2 essentially amounts to establishing $\widehat{B}_{\alpha}(\theta)+\widehat{B}_{\gamma}(\theta)=B_{\alpha}(\theta)+B_{\gamma}(\theta)+o_{\mathbb{P}}(1/\sqrt{NT})$ under $\tau/T\rightarrow 0$ and Assumption 1.
-Here, the difference between $\widehat{B}_{\alpha}(\theta)$ and $B_{\alpha}(\theta)$ is due to the truncation and the estimation error.
-For the truncation, we show that $1/\left(N^{2}T^{2}\right)\sum_{t=1}^{T}\sum_{s=1}^{T}\left(\mathds{1}\{\left|t-s\right|>\tau\}\cdot\right)$, which is the truncated (i.e., the removed) part in (3.4), is negligible relative to $1/\sqrt{NT}$ under the aforementioned conditions.
-Subsequently, the uniform law of large numbers guarantees that the estimation error is also negligible relative to $1/\sqrt{NT}$.
-Next, we turn our attention to the asymptotic properties of the $L(\theta)$-based estimators $\widehat{\theta}_{L}=\arg\max_{\theta}L(\theta)$ and of the $L(\theta)$-based LR, LM, Wald test statistics.
-We enforce the following assumption.
+This is a mild condition in that we only impose it on the expected Hessian $\mathbb{E}H(\theta)$ of the reparameterized likelihood, which does not suffer from the indeterminacy in Equation (2.3).
+We state the condition with “less than 0” because the Hessian in a maximum likelihood model is usually negative definite.
+The following Lemma provides low-level sufficient conditions for Assumptions 1 and 1 under a logit model
+$Y_{it}=\mathds{1}\{X_{it}^{\prime}(\theta+\alpha_{i}+\gamma_{t})+\varepsilon_{it}>0\},$ (C.2)
+where $\varepsilon_{it}$ follows a standard logistic distribution independent of $X_{it}$, $\alpha_{i}$, and $\gamma_{t}.$
+Lemma C.1
+Under model (C.2), a sufficient condition of Assumption 1 is that $\sup_{i,t}\mathbb{E}_{\phi}\|X_{it}\|_{\max}^{3\eta}$$\infty$ for some $\eta>2.$ Let $X_{i\cdot}:=\left(X_{i1},\ldots,X_{iT}\right)^{\prime}$ and $X_{\cdot t}:=\left(X_{1t},\ldots,X_{Nt}\right)^{\prime}$.
+Assumption 1 is satisfied if, for all $i$, the columns of $X_{i\cdot}$ are linearly independent and, for all $t$, the columns of $X_{\cdot t}$ are linearly independent.
+Proof. Under the logit model (C.2) the log-likelihood for individual $i$ at time $t$ is
+$l_{i,t}\left(\theta,\alpha_{i},\gamma_{t}\right)=\left(Y_{it}-1\right)X_{it}^{\prime}(\theta+\alpha_{i}+\gamma_{t})-\log\big(1+\exp\left(-X_{it}^{\prime}(\theta+\alpha_{i}+\gamma_{t})\right)\big).$
+We assume that $X_{it}$ is scalar regressor for simplicity. The derivatives of $l_{i,t}\left(\theta,\alpha_{i},\gamma_{t}\right)$ w.r.t. the heterogeneous parameters are, without loss of generality,
+$\partial_{\alpha_{i}}l_{i,t}\left(\theta,\alpha_{i},\gamma_{t}\right)=\left(Y_{it}-\frac{1}{1+\exp(-X_{it}(\theta+\alpha_{i}+\gamma_{t}))}\right)X_{it},$
+$\partial_{\alpha_{i}\gamma_{t}}l_{i,t}\left(\theta,\alpha_{i},\gamma_{t}\right)=\frac{-\exp(-X_{it}(\theta+\alpha_{i}+\gamma_{t}))}{[1+\exp(-X_{it}(\theta+\alpha_{i}+\gamma_{t}))]^{2}}X_{it}^{2},$
+and
+$\partial_{\alpha_{i}\gamma_{t}\gamma_{t}}l_{i,t}\left(\theta,\alpha_{i},\gamma_{t}\right)=\frac{\exp(-X_{it}(\theta+\alpha_{i}+\gamma_{t}))\left[1-\exp(-X_{it}(\theta+\alpha_{i}+\gamma_{t}))\right]}{[1+\exp(-X_{it}(\theta+\alpha_{i}+\gamma_{t}))]^{3}}X_{it}^{3}.$
+Note that the cross derivatives w.r.t. $\alpha_{i}$ and $\alpha_{j},$ or $\gamma_{t}$ and $\gamma_{s},$ are all 0 for $i\neq j$ or $t\neq s$.
+Assumption 1. We have
+$\sup_{\theta,\alpha_{i},\gamma_{t}}\left|l_{i,t}\left(\theta,\alpha_{i},\gamma_{t}\right)\right|$ $\leq\sup_{\theta,\alpha_{i},\gamma_{t}}\left|X_{it}(\theta+\alpha_{i}+\gamma_{t})\right|+\sup_{\theta,\alpha_{i},\gamma_{t}}\log\big(1+\exp\left(-X_{it}(\theta+\alpha_{i}+\gamma_{t})\right)\big)$
+$\leq 2\sup_{\theta,\alpha_{i},\gamma_{t}}\left|X_{it}(\theta+\alpha_{i}+\gamma_{t})\right|+\log 2$
+$\leq C|X_{it}|+\log 2,$ (C.3)
+where $C<\infty$ is a positive constant depending on the compact parameter spaces $\Theta$ and $\Phi.$ It is easy to see that
+$\sup_{\theta,\alpha_{i},\gamma_{t}}\left|\partial_{\alpha_{i}}l_{i,t}\left(\theta,\alpha_{i},\gamma_{t}\right)\right|\leq\left|X_{it}\right|.$ (C.4)
+For the second derivative, observe that
+$\sup_{\theta,\alpha_{i},\gamma_{t}}\left|\partial_{\alpha_{i}\gamma_{t}}l_{i,t}\left(\theta,\alpha_{i},\gamma_{t}\right)\right|\leq\sup_{\pi\in\mathbb{R}}\frac{\exp(-\pi)}{(\exp(-\pi)+1)^{2}}X_{it}^{2}\leq\frac{1}{4}X_{it}^{2}.$
+For the third derivative, we have
+$\sup_{\theta,\alpha_{i},\gamma_{t}}\left|\partial_{\alpha_{i}\gamma_{t}\gamma_{t}}l_{i,t}\left(\theta,\alpha_{i},\gamma_{t}\right)\right|$ $\leq\sup_{\pi\in\mathbb{R}}\left|\frac{\exp(-\pi)\left(1-\exp(-\pi)\right)}{(\exp(-\pi)+1)^{3}}\right|\left|X_{it}\right|^{3}$
+$\leq\sup_{s\geq 0}\frac{s+s^{2}}{(1+s)^{3}}\left|X_{it}\right|^{3}$
+$\leq\left(\frac{1}{5}+\frac{1}{3}\right)\left|X_{it}\right|^{3}.$ (C.5)
+Combining (C.3)–(C.5), Assumption 1 holds with
+$g\left(w_{it}\right)=C(\left\|X_{it}\right\|_{\max}+1)^{3}+\log 2,$
+where $C>1.$ Therefore, the condition $\sup_{i,t}\mathbb{E}_{\phi}\left\|X_{it}\right\|_{\max}^{3\eta}<\infty$ is sufficient for Assumption 1.
+Assumption 1. Under the logit model (C.2), the Hessian can be expressed as
+$\sqrt{NT}H(\theta)=\frac{1}{\sqrt{NT}}DX^{\prime}H^{\pi}(\theta)XD^{\prime},$ (C.6)
+where $D$ is the deterministic matrix defined in Section 2.1, and $X=(X^{\alpha},X^{\gamma})$ with
+$X^{\gamma}=\left(\begin{array}[c]{ccc}X_{11}^{\prime}&&\\ &\ddots&\\ &&X_{1T}^{\prime}\\ X_{21}^{\prime}&&\\ &\ddots&\\ &&X_{2T}^{\prime}\\ \vdots&\ddots&\vdots\\ X_{N1}^{\prime}&&\\ &\ddots&\\ &&X_{NT}^{\prime}\end{array}\right),\qquad X^{\alpha}=\left(\begin{array}[c]{cccc}X_{11}^{\prime}&&&\\ \vdots&&&\\ X_{1T}^{\prime}&&&\\ &X_{21}^{\prime}&&\\ &\vdots&&\\ &X_{2T}^{\prime}&&\\ &&\ddots&\\ &&&X_{N1}^{\prime}\\ &&&\\ &&&X_{NT}^{\prime}\end{array}\right).$
+Moreover, $H^{\pi}(\theta)=\operatorname*{diag}\left\{H_{1}^{\pi}(\theta),\ldots,H_{N}^{\pi}(\theta)\right\}$ is a diagnal matrix, where block $H_{i}^{\pi}(\theta)=\operatorname*{diag}\left\{h_{i,1}^{\pi}(\theta),\ldots,h_{i,T}^{\pi}(\theta)\right\}$ and each element is given by
+$h_{i,t}^{\pi}(\theta)=-\frac{\exp(-\pi_{i,t}(\theta))}{[1+\exp(-\pi_{i,t}(\theta))]^{2}}<0,\quad\pi_{i,t}(\theta)=X_{it}^{\prime}(\theta+\alpha_{i}(\theta)+\gamma_{t}(\theta)).$
+Here, $\alpha_{i}(\theta)$ and $\gamma_{t}(\theta)$ denote the pseudo-true values defined by (2.4). Let $\lambda_{\max}(A)$ denote the maximum eigenvalue of a matrix $A.$ Using (C.6), we have
+$\lambda_{\max}(\sqrt{NT}\mathbb{E}H(\theta))\leq\sup_{i,t}h_{i,t}^{\pi}(\theta)\lambda_{\max}(DX^{\prime}XD^{\prime}/\sqrt{NT}).$
+Therefore, Assumption 1 is implied if $XD^{\prime}$ has full column rank.
+Given the structure of $X$, this requires that, for all $i$, the columns of $X_{i\cdot}$ are linearly independent and that, for all $t$, the columns of $X_{\cdot t}$ are linearly independent.
+The first condition requires that certain higher moment of $\left\|X_{it}\right\|_{\max}$ are bounded away from the infinity uniformly.
+Note that $X_{it}$ has a fixed dimension, so that the choice of norm is irrelevant.
+This is relatively a mild condition.
+The second condition states that the regressors are linearly independent in both the individual and time dimension.
+This condition is standard for models with two-way heterogeneities, and it is usually satisfied if the regressors have sufficient variations across both the cross-sectional and time dimension.
+See Remark 1 of Fernández-Val and Weidner (2016) for a similar argument.
+Notably, it rules out the possibility of $Y_{it-1}$ being constant across $i$ or $t$.
+To establish the asymptotic properties of $\widehat{\theta}_{L}$ and $L(\theta)$-based LR, LM, Wald test statistics, we enforce the following assumption.
 Assumption 2
@@
     $\mathbb{E}\sup_{\theta\in\Theta}\left\|\nabla_{\theta\theta^{\prime}}l(\theta)\right\|<\infty$
-[…]
+    where $\left\|\cdot\right\|$ is the matrix Frobenius norm.
   • (v) The profiled Hessian $\mathbb{E}\nabla_{\theta\theta^{\prime}}l(\theta)$ is strictly negative definite for large $N$ and $T$.
@@
 It guarantees that $\operatorname*{plim}_{N,T\rightarrow\infty}l(\theta)$ is uniquely maximized at $\theta_{0}$.
-Assumption 2 is a dominance condition implying that $l(\theta)-\mathbb{E}l(\theta)$ converges to 00 uniformly as $N,T\rightarrow\infty$.
+Assumption 2 is a dominance condition implying that $l(\theta)-\mathbb{E}l(\theta)$ converges to 0 uniformly as $N,T\rightarrow\infty$.
 Together, Assumptions 2 to 2 are used to establish the consistency of $\widehat{\theta}_{L}$.
@@
 Assumptions 2 and 2 are additionally imposed to establish the asymptotic distributions for the quantities of interests.
+C.3 Asymptotic Theory
+The next theorem formally states the asymptotic expansion of $\widehat{l}(\theta)$.
+Theorem 1
+Under Assumption 1 and uniformly over $\theta\in\Theta$,
+$\mathbb{E}l(\theta)=\mathbb{E}\widehat{l}(\theta)+B_{\alpha}(\theta)+B_{\gamma}(\theta)+o(1/\sqrt{NT}),$
+as $N,T\rightarrow\infty$, where
+$B_{\alpha}(\theta)=\frac{1}{2}\operatorname*{trace}[S_{\alpha\alpha}(\theta)H_{\alpha\alpha}^{\mathcal{\ast}}(\theta)],\qquad B_{\gamma}(\theta)=\frac{1}{2}\operatorname*{trace}\mathbb{[}S_{\gamma\gamma}(\theta)H_{\gamma\gamma}^{\ast}(\theta)],$
+and $B_{\alpha}(\theta)+B_{\gamma}(\theta)=O(1/\sqrt{NT})$.
+Proof. In Section E.
+[…]
+The next theorem, which shows that approximation error of the corrected likelihood $L(\theta)$ to the infeasible IPP-free likelihood $l(\theta)$ is of order $1/\sqrt{NT}$ uniformly over $\theta$.
+This is our claim in the beginning of this section.
+Theorem 2
+Under Assumption 1 and for some $\tau\rightarrow\infty$ such that $\tau/T\rightarrow 0$, we have
+$L(\theta)-l(\theta)=o_{\mathbb{P}}(1/\sqrt{NT}),$
+uniformly over $\theta\in\Theta$ as $N,T\rightarrow\infty$, where $L(\theta)$ is the corrected likelihood defined in Equation (3.3).
+Showing Theorem 2 essentially amounts to establishing $\widehat{B}_{\alpha}(\theta)+\widehat{B}_{\gamma}(\theta)=B_{\alpha}(\theta)+B_{\gamma}(\theta)+o_{\mathbb{P}}(1/\sqrt{NT})$ under $\tau/T\rightarrow 0$ and Assumption 1.
+Here, the difference between $\widehat{B}_{\alpha}(\theta)$ and $B_{\alpha}(\theta)$ is due to the truncation and the estimation error.
+For the truncation, we show that
+$\frac{1}{N^{2}T^{2}}\sum_{t=1}^{T}\sum_{s=1}^{T}\left(\mathds{1}\{\left|t-s\right|>\tau\}\operatorname*{diag}\{\widehat{s}_{1,t}^{\alpha}(\theta)[\widehat{s}_{1,s}^{\alpha}(\theta)]^{\prime},\ldots,\widehat{s}_{N,t}^{\alpha}(\theta)[\widehat{s}_{N,s}^{\alpha}(\theta)]^{\prime}\right),$
+which is the truncated (i.e., the removed) part in (3.6), is negligible relative to $1/\sqrt{NT}$ under the aforementioned conditions.
+Subsequently, the uniform law of large numbers guarantees that the estimation error is also negligible relative to $1/\sqrt{NT}$.
 The corollary below states the asymptotic properties of $\widehat{\theta}_{L}$.
@@
 as $N,T\rightarrow\infty$.
-Here, $[-\mathbb{E}\triangledown_{\theta\theta^{\prime}}l(\theta_{0})]^{1/2}$ is the square root of $-\mathbb{E}\triangledown_{\theta\theta^{\prime}}l(\theta_{0})$ which can be obtained by square-rooting the eigenvalues in the eigendecomposition of $-\mathbb{E}\triangledown_{\theta\theta^{\prime}}l(\theta_{0})$.
-By Assumption 2, $\mathbb{E}\triangledown_{\theta\theta^{\prime}}l(\theta_{0})$ is strictly negative definite so that $-\mathbb{E}\triangledown_{\theta\theta^{\prime}}l(\theta_{0})$ only has strictly positive eigenvalues.
+Proof. In Section E.
+Here, $\mathbb{E}\triangledown_{\theta\theta^{\prime}}l(\theta_{0})]^{1/2}$ is the square root of $\mathbb{E}\triangledown_{\theta\theta^{\prime}}l(\theta_{0})$ which can be obtained by square-rooting the eigenvalues in the eigendecomposition of $\mathbb{E}\triangledown_{\theta\theta^{\prime}}l(\theta_{0})$.
+By Assumption 2, $\mathbb{E}\triangledown_{\theta\theta^{\prime}}l(\theta_{0})$ is strictly negative definite so that $\mathbb{E}\triangledown_{\theta\theta^{\prime}}l(\theta_{0})$ only has strictly positive eigenvalues.
 In addition, since we focus on maximum likelihood models, we use the information matrix equality to simplify the presentation of Corollary 1.
-In case the information matrix equality does not hold [^6] The information matrix equality generally does not hold for, e.g., non-likelihood M-estimation models, misspecified models, and quasi-likelihood models., $-\mathbb{E}\triangledown_{\theta\theta^{\prime}}l(\theta_{0})$ shall be replaced by its sandwich form and $[-\mathbb{E}\triangledown_{\theta\theta^{\prime}}l(\theta_{0})]^{1/2}$ shall be constructed thereof.
+In case the information matrix equality does not hold [^7] The information matrix equality generally does not hold for, e.g., non-likelihood M-estimation models, misspecified models, and quasi-likelihood models., $\mathbb{E}\triangledown_{\theta\theta^{\prime}}l(\theta_{0})$ shall be replaced by its sandwich form and $\mathbb{E}\triangledown_{\theta\theta^{\prime}}l(\theta_{0})]^{1/2}$ shall be constructed thereof.
 See, e.g., White (1982) for detailed discussions.
-For inference procedures, we consider a generic null hypothesis $H_{0}:R(\theta_{0})=0$, where $R(\theta)$ is a known $r\times 1$ ($r\leq K$) vector-valued non-random function with Jacobian $J(\theta):=\partial_{\theta^{\prime}}R(\theta)$ satisfying $\operatorname*{rank}[J(\theta)]=r$.
-The corollary below gives the asymptotic distribution, under $H_{0}$, of $\widehat{\xi}_{LR}$ (the LR statistic), $\widehat{\xi}_{LM}$ (the LM statistic), and $\widehat{\xi}_{Wald}$ (the Wald statistic), defined respectively as
-$\widehat{\xi}_{LR}$ $:=-2NT\{L(\widehat{\theta}_{R})-L(\widehat{\theta}_{L})\},$
-$\widehat{\xi}_{LM}$ $:=-NT\{\triangledown_{\theta^{\prime}}L(\widehat{\theta}_{R})[\triangledown_{\theta\theta^{\prime}}L(\widehat{\theta}_{R})]^{-1}\triangledown_{\theta}L(\widehat{\theta}_{R})\},$
-$\widehat{\xi}_{Wald}$ $:=-NT\{R^{\prime}(\widehat{\theta}_{L})[J(\widehat{\theta}_{L})[\triangledown_{\theta\theta^{\prime}}L(\widehat{\theta}_{L})]^{-1}J^{\prime}(\widehat{\theta}_{L})]^{-1}R(\widehat{\theta}_{L})\},$
-where $\widehat{\theta}_{R}:=\arg\max_{\theta}L(\theta)$ subject to $R(\theta)=0$.
+The next corollary states the asymptotic distribution of the LR ($\widehat{\xi}_{LR}$), the LM ($\widehat{\xi}_{LM}$), and the Wald ($\widehat{\xi}_{Wald}$) statistics.
 Corollary 2
@@
 $\widehat{\xi}_{LR},\widehat{\xi}_{LM},\widehat{\xi}_{Wald}\overset{\mathbb{D}}{\longrightarrow}\chi^{2}(r)$
-as $N,T\rightarrow\infty$, where $\chi^{2}(r)$ is the $\chi^{2}$-distribution with degrees of freedom $r$.
+as $N,T\rightarrow\infty$.
+Proof. In Section E.
 Similarly, the test statistics $\widehat{\xi}_{LM}$ and $\widehat{\xi}_{Wald}$ are constructed under the information matrix equality.
 Adjustments are generally required if it does not hold.
-For instance, White (1982) shows that $-[\triangledown_{\theta\theta^{\prime}}L(\widehat{\theta}_{L})]^{-1}$ involved in $\widehat{\xi}_{Wald}$ shall be replaced by its sandwich form for quasi-likelihood models.
+For instance, White (1982) shows that $\triangledown_{\theta\theta^{\prime}}L(\widehat{\theta}_{L})]^{-1}$ involved in $\widehat{\xi}_{Wald}$ shall be replaced by its sandwich form for quasi-likelihood models.
 Given Theorem 2, Corollaries 1 and 2 are established using standard techniques.
@@
 The proof of Corollary 2 also closely resembles the standard techniques in, e.g., Hayashi (2000), where it is shown that all $\widehat{\xi}_{LR}$, $\widehat{\xi}_{LM}$, and $\widehat{\xi}_{Wald}$ are equivalent, converging in probability to some $\widehat{\xi}$ which is asymptotically $\chi^{2}(r)$.
-4 Simulation
-We conduct the simulation study in this section.
-We first illustrate that our procedure is capable of reducing the bias of the MLE as well as delivering likelihood-based tests with an empirical size very close to the nominal level.
-Followed by this, we then consider a different design in which the MLEs and the tests suffer from severe biases and size-distortions.
-We demonstrate that our approach leads to dramatic improvements on the estimators and substantial reductions of the Type-I error risks of the tests, even in a severely biased setting.
-Finally in this section, we compare our method with the Jackknife [^7] Since there has currently been no Jackknife procedures for DN-HP models.
-We compared our method to the Jackknife of Fernández-Val and Weidner (2016), who consider models with fixed effects only in the intercept.
-and show that our method out-performs.
-In Design 1, we consider a dynamic binary response model as:
-$Y_{it}=\mathds{1}\{\rho Y_{it-1}+(\beta+\alpha_{1,i}+\gamma_{1,t})X_{it}+\alpha_{2,i}+\gamma_{2,t}+\varepsilon_{it}>0\}$ (4.1)
-where the true value is $\theta_{0}=(\rho_{0},\beta_{0})^{\prime}=(0.5,0.5)^{\prime}$ and the disturbance $\varepsilon_{it}$ is standard logistic, independent and identically distributed (i.i.d.) over $(i,t)$ and independent from $X_{it}$.
-The true incidental parameter $\{\alpha_{1,i}^{0},\gamma_{1,t}^{0}\}$ are generated from $\mathcal{N}(0,0.01)$, and demeaned after they are generated.
-$\{\alpha_{2,i}^{0},\gamma_{2,t}^{0}\}$ are also generated from $\mathcal{N}(0,0.01)$ but only $\gamma_{2,t}^{0}$ is demeaned.
-Conditional on the generated $\alpha_{k,i}^{0}$ and $\gamma_{k,t}^{0}$ ($k=1,2$), the regressor $X_{it}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0}+\gamma_{1,t}^{0}+\gamma_{2,t}^{0},0.25)$ i.i.d.
-over $(i,t)$ and the initial $Y_{i0}=\mathds{1}\{(\beta_{0}+\alpha_{1,i}^{0})X_{i0}+\alpha_{2,i}^{0}+\varepsilon_{i0}>0\}$ with $X_{i0}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0},0.25)$ i.i.d.
-over $(i,t)$.
-We also simulate a static model, where we set $\rho\equiv 0$ in (4.1), using data generated in the same way but setting $\rho_{0}=0$ (hence $\theta_{0}=\beta_{0}$).
-Table 1 reports the estimation bias and RMSE, under Design 1, of $\widehat{\theta}_{L}$ and $\widehat{\theta}$ for both the static ($\theta_{0}=\beta_{0}$) and dynamic ($\theta_{0}=(\rho_{0},\beta_{0})^{\prime}$) logit models.
-For the static model, $\widehat{\beta}$ denotes the uncorrected MLE obtained by maximing $\widehat{l}(\theta)$, and $\widehat{\beta}_{L}$ is the corrected estimator, obtained by maximing $L(\theta)$.
-For the dynamic model, $(\widehat{\beta},\widehat{\rho})^{\prime}$ is the uncorrected estimator for $(\rho_{0},\beta_{0})^{\prime}$, while $(\widehat{\beta}_{L}^{(\tau)},\widehat{\rho}_{L}^{(\tau)})^{\prime}$ represents the bias-corrected estimator, with the truncation parameter $\tau\in(1,2)$.
-For the static model (with $\rho_{0}\equiv 0$), we set $\tau=0$ since the data are independent over $i$, conditional on $\{\alpha_{k,i}^{0},\gamma_{k,t}^{0}\}$.
-The results reveal that the uncorrected estimators are severely biased in both models, while our bias-corrected estimators substantially reduce the bias.
-For instance, in the static model with $(N,T)=(30,30)$, the bias for $\widehat{\beta}_{L}$ is only $0.05$, which is less than $1/3$ of the significant bias of $0.158$ for $\widehat{\beta}$.
-As the sample size increases to $(N,T)=(90,90)$, $\widehat{\beta}_{L}$ is nearly unbiased, while the bias of $\widehat{\beta}$ remaining five times larger than that of $\widehat{\beta}_{L}$.
-Furthermore, the RMSEs from our corrected estimators are consistently lower than those of the uncorrected estimators, especially with smaller sample sizes.
-This indicates that our procedure is capable of reducing the IPP bias without increasing the RMSE.
-For large $(N,T)$, the choice of $\tau$ has negligible impact on the performance of our bias-corrected estimators.
-Table 1: Comparisons of Uncorrected and Corrected Logit Estimates, Design 1
-$(N,T)$ $(30,30)$ $(60,60)$ $(90,90)$
-Bias RMSE Bias RMSE Bias RMSE
-Static Model
-$\widehat{\beta}$ $0.158$ $0.253$ $0.075$ $0.109$ $0.051$ $0.072$
-$\widehat{\beta}_{L}$ $0.050$ $0.175$ $0.018$ $0.073$ $0.010$ $0.049$
-Dynamic Model
-$\widehat{\beta}$ $0.165$ $0.261$ $0.074$ $0.111$ $0.046$ $0.071$
-$\widehat{\beta}_{L}^{(1)}$ $0.059$ $0.181$ $0.017$ $0.077$ $0.006$ $0.051$
-$\widehat{\beta}_{L}^{(2)}$ $0.061$ $0.183$ $0.018$ $0.078$ $0.007$ $0.051$
-$\widehat{\rho}$ $-0.107$ $0.201$ $-0.048$ $0.093$ $-0.033$ $0.058$
-$\widehat{\rho}_{L}^{(1)}$ $-0.044$ $0.159$ $-0.014$ $0.076$ $-0.010$ $0.047$
-$\widehat{\rho}_{L}^{(2)}$ $-0.038$ $0.161$ $-0.009$ $0.077$ $-0.006$ $0.046$
-[…]
-Notes: Logit model with $\varepsilon_{it}$ standard logistic.
-Design 1: $\{\alpha_{k,i}^{0},\gamma_{k,t}^{0}\}\sim\mathcal{N}(0,0.01)$, $X_{it}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0}+\gamma_{1,t}^{0}+\gamma_{2,t}^{0},0.25)$, and $X_{i0}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0},0.25)$.
-For the static model, $\widehat{\beta}$ and $\widehat{\beta}_{L}$ are, respectively, the uncorrected and corrected estimates of $\beta_{0}=0.5$.
-For the dynamic model $(\widehat{\beta},\widehat{\rho})^{\prime}$ is the uncorrected estimate of $(\rho_{0},\beta_{0})^{\prime}=(0.5,0.5)^{\prime}$ from and $(\widehat{\beta}_{L}^{(\tau)},\widehat{\rho}_{L}^{(\tau)})^{\prime}$ is the corrected estimate with truncation parameter $\tau\in(1,2)$.
-The number of replications is $1000$.
-Next, under Design 1, we conduct LR tests using both the corrected and the uncorrected likelihoods for the null hypothesis $H_{0}:\theta_{0}=(0.5,0.5)^{\prime}+\delta$ with $\delta\in\{0,\pm 0.2,\pm 0.15,\pm 0.1\}$.
-We report the empirical size ($\delta=0$) and the empirical power ($\delta\neq 0$) at the $5\%$ significance level.
-For the corrected likelihood, we set the truncation parameter $\tau=0$ for the static model and $\tau=1$ for the dynamic model.
-Table 2 presents the empirical size.
-The results show that the LR test using the uncorrected likelihood suffers from a severe size distortion.
-On the opposite, the empirical size from our bias-corrected likelihood maintains the significance level as the sample size increases.
-Table 2: Empirical Size of LR Tests Based on Corrected and Uncorrected Likelihoods for Logit Model, Design 1
-Static Model Dynamic Model
-$(N,T)$ Uncorrected Corrected Uncorrected Corrected
-$(30,30)$ $16\%$ $9\%$ $18\%$ $10\%$
-$(60,60)$ $15\%$ $5\%$ $17\%$ $7\%$
-$(90,90)$ $18\%$ $6\%$ $16\%$ $5\%$
-[…]
-Notes: Logit model with $\varepsilon_{it}$ standard logistic.
-Design 1: $\{\alpha_{k,i}^{0},\gamma_{k,t}^{0}\}\sim\mathcal{N}(0,0.01)$, $X_{it}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0}+\gamma_{1,t}^{0}+\gamma_{2,t}^{0},0.25)$, and $X_{i0}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0},0.25)$.
-The null hypothesis $H_{0}:(\beta_{0},\rho_{0})^{\prime}=(0.5,0.5)^{\prime}$ for the dynamic model, and $H_{0}:\beta_{0}=0.5$ for the static model, with a significance level of $5\%$.
-The number of replications is $1000$.
-Table 3 presents the empirical power of LR tests based on our corrected likelihood under Design 1.
-The empirical power increases as both the value of $\delta$ and the sample size grow.
-For $(N,T)=(90,90)$, when $\delta=\pm 0.2$, the rejection rates reach nearly $100\%$ for both the static and dynamic models.
-We also perform LM and Wald test, reporting the empirical size and power under the same setting in Appendix B.1.
-The results align with those of the LR test.
-We additionally perform the simulations for a probit model with $\varepsilon_{it}\sim\mathcal{N}(0,1)$ i.i.d.
-over $(i,t)$ and independent from $X_{it}$.
-The results are provided in Appendix B.2.
-Table 3: Empirical Power of LR Test for Logit Model using $L(\theta)$, Design 1
-$\delta$
-$(N,T)$ $-0.20$ $-0.15$ $-0.10$ $0.10$ $0.15$ $0.20$
-Static Model
-$(30,30)$ $45\%$ $32\%$ $23\%$ $10\%$ $14\%$ $21\%$
-$(60,60)$ $86\%$ $63\%$ $37\%$ $25\%$ $49\%$ $74\%$
-$(90,90)$ $100\%$ $93\%$ $65\%$ $50\%$ $84\%$ $98\%$
-Dynamic Model
-$(30,30)$ $48\%$ $33\%$ $21\%$ $18\%$ $28\%$ $43\%$
-$(60,60)$ $95\%$ $74\%$ $40\%$ $45\%$ $75\%$ $94\%$
-$(90,90)$ $100\%$ $98\%$ $76\%$ $78\%$ $98\%$ $100\%$
-[…]
-Notes: Logit model with $\varepsilon_{it}$ standard logistic.
-Design 1: $\{\alpha_{k,i}^{0},\gamma_{k,t}^{0}\}\sim\mathcal{N}(0,0.01)$, $X_{it}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0}+\gamma_{1,t}^{0}+\gamma_{2,t}^{0},0.25)$, and $X_{i0}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0},0.25)$.
-The null hypothesis $H_{0}:(\beta_{0},\rho_{0})^{\prime}=(0.5+\delta,0.5+\delta)^{\prime}$ for the dynamic model, and $H_{0}:\beta_{0}=0.5+\delta$ for the static model, with a significance level $5\%,$ where $\delta\in\{\pm 0.2,\pm 0.15,\pm 0.1\}$.
-The number of replications is $1000$.
-Next, we conduct the same experiment under Design 2, where we generate $\{\alpha_{k,i}^{0},\gamma_{k,t}^{0}\}\sim\mathcal{N}(0,1/4)$, $X_{it}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0}+\gamma_{1,t}^{0}+\gamma_{2,t}^{0},1)$, and $X_{i0}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0},1)$, keeping everything else intact as in (4.1).
-We increase the volatility of $\alpha_{k,i}^{0}$, $\gamma_{k,t}^{0}$, $X_{it}$, and $X_{i0}$ to induce larger biases and size distortions.
-This time, we present the results for the probit model and leave that for the logit model in Appendix B.1.
-Table 4 presents the comaprison of the bias and RMSE of the corrected and the uncorrected estimators for the probit model, under Design 2.
-First of all, the probit estimator here is much more biased than the previous logit counterparts.
-For the static model with $(N,T)=(30,30)$, $\widehat{\beta}$ contains a bias of $0.943$.
-Recall that the true value is only $\beta_{0}=0.5$, implying that the bias is almost $200\%$.
-Our corrected estimator $\widehat{\beta}_{L}$ successfully reduces this bias to $0.328$, which is one third of the uncorrected and roughly $60\%$ biased.
-As the sample sizes increase, the bias systematically becomes smaller both for the uncorrected and corrected estimators, with dramatic improvements observed from the corrected ones both for the stiatic and dynamic models.
-We indeed observe that the RMSE of $\widehat{\beta}_{L}$ is twice as large as that of $\widehat{\beta}$ for the static model under $N,T=(30,30)$.
-However, it is the only incidence here and we attribute this to some sort of numerical coincidence [^8] Below, when we compare our method with the Jackknife under the same setting, it will be seen that the RMSE of our corrected estimator is indeed smaller than the uncorrected.
-See also Footnote 9 below..
-It can be checked that the RMSE of our corrected estimator is no worse the uncorrected one in every other case.
-Table 4: Comparisons of Uncorrected and Corrected Probit Estimates, Design 2
-$(N,T)$ $(30,30)$ $(60,60)$ $(90,90)$
-Bias RMSE Bias RMSE Bias RMSE
-Static Model
-$\widehat{\beta}$ $0.943$ $2.272$ $0.151$ $0.252$ $0.073$ $0.081$
-$\widehat{\beta}_{L}$ $0.328$ $5.170$ $0.041$ $0.062$ $0.017$ $0.040$
-Dynamic Model
-$\widehat{\beta}$ $0.909$ $1.842$ $0.166$ $0.248$ $0.084$ $0.100$
-$\widehat{\beta}_{L}^{(1)}$ $0.207$ $0.396$ $0.048$ $0.078$ $0.022$ $0.042$
-$\widehat{\beta}_{L}^{(2)}$ $0.202$ $0.282$ $0.049$ $0.080$ $0.022$ $0.042$
-$\widehat{\rho}$ $-0.030$ $0.162$ $-0.016$ $0.073$ $-0.011$ $0.045$
-$\widehat{\rho}_{L}^{(1)}$ $-0.015$ $0.140$ $-0.005$ $0.066$ $-0.004$ $0.044$
-$\widehat{\rho}_{L}^{(2)}$ $-0.009$ $0.141$ $-0.001$ $0.067$ $-0.001$ $0.045$
-[…]
-Notes: Probit model with $\varepsilon_{it}\sim\mathcal{N}(0,1)$.
-Design 2: $\{\alpha_{k,i}^{0},\gamma_{k,t}^{0}\}\sim\mathcal{N}(0,1/4)$, $X_{it}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0}+\gamma_{1,t}^{0}+\gamma_{2,t}^{0},1)$, and $X_{i0}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0},1)$.
-For the static model, $\widehat{\beta}$ and $\widehat{\beta}_{L}$ are, respectively, the uncorrected and corrected estimates of $\beta_{0}=0.5$.
-For the dynamic model $(\widehat{\beta},\widehat{\rho})^{\prime}$ is the uncorrected estimate of $(\rho_{0},\beta_{0})^{\prime}=(0.5,0.5)^{\prime}$ from and $(\widehat{\beta}_{L}^{(\tau)},\widehat{\rho}_{L}^{(\tau)})^{\prime}$ is the corrected estimate with truncation parameter $\tau\in(1,2)$.
-The number of replications is $1000$.
-Table 5 below reports the empirical size of the LR test for the probit model under Design 2.
-As before, the null hypothesis is $H_{0}:(\beta_{0},\rho_{0})^{\prime}=(0.5,0.5)^{\prime}$ for the dynamic model and $H_{0}:\beta_{0}=0.5$ for the static.
-We set the truncation parameter $\tau=0$ for the static model and $\tau=1$ for the dynamic model.
-Here the size distortion is much more fatal than in the logit case above.
-For the static model with $(N,T)=(30,30)$, the uncorrected LR test produces an empirical size of $77\%$.
-Even under $(N,T)=(90,90)$, the empirical size is still $72\%$, indicating a prohibitively high risk of committing a Type-I error.
-On the other hand, the $L(\theta)$-based LR test delivers a $14\%$ empirical size only, which is almost a $60\%$ decrease in magnitude.
-Similar performance is observed in other cases.
-The results here demonstrate that our method can reduce the type-I error risks by a major margin.
-Indeed, our test still over-rejects under $(N,T)=(90,90)$.
-This is anticipated, because the IPP under Design 2 is much more drastic and $(N,T)=(90,90)$ may not be sufficient for the test statistic to be well-approximated by $\chi^{2}$.
-However, it is clear that, as $(N,T)$ increase, the empirical size does become smaller.
-We have tried additional small-scale simulations (not reported though) with even larger $(N,T)$ and found that the empirical size still decreases with $(N,T)$.
-We similarly report the results from the LM and the Wald test in Appendix B.2.
-We report the empirical power under Design 2 in the Appendix B, since they do not add too much value in the discussion here.
-A quick summary for those results is that our procedure produces tests with stronger power in Design 2.
-Table 5: Empirical Size of LR Tests Based on Corrected and Uncorrected Likelihoods for Logit Model, Design 2
-Static Model Dynamic Model
-$(N,T)$ Uncorrected Corrected Uncorrected Corrected
-$(30,30)$ $77\%$ $39\%$ $71\%$ $37\%$
-$(60,60)$ $76\%$ $20\%$ $66\%$ $19\%$
-$(90,90)$ $72\%$ $14\%$ $63\%$ $14\%$
-[…]
-Notes: Probit model with $\varepsilon_{it}\sim\mathcal{N}(0,1)$.
-Design 2: $\{\alpha_{k,i}^{0},\gamma_{k,t}^{0}\}\sim\mathcal{N}(0,1/4)$, $X_{it}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0}+\gamma_{1,t}^{0}+\gamma_{2,t}^{0},1)$, and $X_{i0}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0},1)$.
-The null hypothesis $H_{0}:(\beta_{0},\rho_{0})^{\prime}=(0.5,0.5)^{\prime}$ for the dynamic model, and $H_{0}:\beta_{0}=0.5$ for the static model, with a significance level of $5\%$.
-The number of replications is $1000$.
-Finally, we compare our method with the Jackknife of Fernández-Val and Weidner (2016).
-We consider the same probit model and continue the same setting and design as above.
-Table 6 displays the estimation bias and the RMSE of the uncorrected estimators $\{\widehat{\rho},\widehat{\beta}\}$, our corrected estimator $\{\widehat{\rho}_{L},\widehat{\beta}_{L}\}$, and the Jackknife-corrected estimator $\{\widehat{\rho}_{J},\widehat{\beta}_{J}\}$ for the dynamic model.
-For the static model, only $\widehat{\beta}$, $\widehat{\beta}_{L}$, and $\widehat{\beta}_{J}$ apply and are reported.
-We set the truncation parameter $\tau=0$ for static model and $\tau=1$ for dynamic model.
-Across all pairs of $(N,T)$, the Jackknife-corrected $\widehat{\beta}_{J}$ consistently exhibits larger bias and RMSE than our $\widehat{\beta}_{L}$.
-Even worse, $\widehat{\beta}_{J}$ shows a larger bias and RMSE than the uncorrected $\widehat{\beta}$ across all pairs of $(N,T)$.
-For instance, with $(N,T)=(30,30)$ and the static model, $\widehat{\beta}_{J}$ shows a bias of $-4.929$, which is roughly $900\%$ biased (in magnitude) relative to the true value $\beta_{0}=0.5$.
-Observe that the uncorrected estimator $\widehat{\beta}$ only shows a roughly $170\%$ bias relative to the true value.
-Our corrected $\widehat{\beta}_{L}$ performs the best among the three, showing a $36\%$ bias [^9] Also notice that the RMSE of $\widehat{\beta}_{L}$ in Table 6 is much smaller than that of $\widehat{\beta}_{L}$ in Table 4, where both tables are essentially under the same setting.
-This motivates our attribution of the large RMSE in Table 4 to numerical coincidence..
-To summarize, we find that the Jackknife is not a suitable candidate for DN-HP models.
-Although it does reduce the bias for the homogeneous coefficient for large $N,T$, the bias the Jackknife causes to the heterogeneous parameter is dangerously large, especially in small sample.
-Table 6: Comparison with the Jackknife Bias Correction Method
-$(N,T)$ $(30,30)$ $(60,60)$ $(90,90)$
-Bias RMSE Bias RMSE Bias RMSE
-Static Model
-$\widehat{\beta}$ $0.846$ $2.168$ $0.151$ $0.261$ $0.074$ $0.082$
-$\widehat{\beta}_{L}$ $0.182$ $0.339$ $0.041$ $0.068$ $0.019$ $0.033$
-$\widehat{\beta}_{J}$ $-4.929$ $14.383$ $-0.504$ $0.981$ $-0.105$ $0.214$
-Dynamic Model
-$\widehat{\beta}$ $1.040$ $2.343$ $0.184$ $0.287$ $0.080$ $0.090$
-$\widehat{\beta}_{L}$ $0.209$ $0.332$ $0.052$ $0.082$ $0.020$ $0.049$
-$\widehat{\beta}_{J}$ $-3.352$ $7.708$ $-0.558$ $1.125$ $-0.118$ $0.217$
-$\widehat{\rho}$ $-0.046$ $0.169$ $-0.022$ $0.072$ $-0.014$ $0.046$
-$\widehat{\rho}_{L}$ $-0.028$ $0.145$ $-0.011$ $0.067$ $-0.009$ $0.044$
-$\widehat{\rho}_{J}$ $-0.029$ $0.169$ $-0.006$ $0.068$ $-0.002$ $0.044$
-[…]
-Notes: Probit model with $\varepsilon_{it}\sim\mathcal{N}(0,1)$.
-Design 2: $\{\alpha_{k,i}^{0},\gamma_{k,t}^{0}\}\sim\mathcal{N}(0,1/4)$, $X_{it}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0}+\gamma_{1,t}^{0}+\gamma_{2,t}^{0},1)$, and $X_{i0}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0},1)$.
-The true value is $(\beta_{0},\rho_{0})^{\prime}=(0.5,0.5)^{\prime}$ for the dynamic model, and $\beta_{0}=0$ for the static model.
-Accordingly, $(\widehat{\beta},\widehat{\rho})^{\prime}$ (or $\widehat{\beta}$) denotes the uncorrected estimator; $(\widehat{\beta}_{L},\widehat{\rho}_{L})^{\prime}$ (or $\widehat{\beta}_{L}$) is our bias-corrected MLE with truncation parameter $\tau=1$ for the dynamic model or $\tau=0$ for the static model; and $(\widehat{\beta}_{J},\widehat{\rho}_{J})^{\prime}$ (or $\widehat{\beta}_{J}$) represents the Jackknife bias-corrected estimators.
-The number of replications is $1000$.
-5 Empirical Analysis
-In this section we apply our likelihood-based bias correction method to examining the determinants of the labor force participation (LFP) decision of single mothers.
-In particular, we look at the impact of the number of children on the decision of the mother to engage in paid employment.
-An extensive literature in labor economics has studied the labor supply decisions of married women (Killingsworth and Heckman, 1986; Angrist and Evans, 1998; Blau and Kahn, 2007).
-These studies have uncovered the impacts of a variety of economic variables, including female market wage and husband income (Mincer, 1962), education (Heath and Jayachandran, 2018), childcare costs (Connelly, 1992), the cost of home technology (Greenwood et al., 2016), and culture norms (Fernández, 2013).
-Among these variables, the number of children consistently emerges as one of the most important determinants of female labor supply (Nakamura and Nakamura, 1992).
-This is unsurprising since women continue to bear a disproportionate share of child-rearing responsibilities (Aguero and Marks, 2008).
-In contrast to the substantial body of research on the labor supply behavior of married women, the labor supply decisions of single mothers have received relatively limited attention, with only a few exceptions (Kimmel, 1998; Blundell et al., 2016).
-Single mothers, however, face distinctive challenges when it comes to balancing work and child-rearing responsibilities due to the absence of a second earner in the household.
-Moreover, their employment decisions may have a more pronounced impact on their children’s well-being compared to the decisions of married women and should thus be of great importance to economists and policymakers.
-To study the labor supply decisions of single mothers, we compile a data set from waves $20-30$ of the Panel Study of Income Dynamics (PSID), which span the period of 1987 to 1997.
-Our sample includes only single mothers, defined as unmarried female household heads with at least one child.
-The dependent variable is labor force participation status, defined as whether the individual worked any hours during the interview year.
-Our main explanatory variable is the number of children under 18 in the household.
-Following Dhaene and Jochmans (2015), we restrict our sample to individuals aged $18-60$ whose LFP status changed at least once during the period.
-The final sample comprises $N=90$ single mothers whose labor supply status is continuously observed over $T=10$ years from 1987 to 1996.
-Further details on data construction are provided in Appendix C.
-On average, 27 percent of these individuals switched in and out of the labor force in each year (Figure C.1--C.2).
-Table C.1 presents the summary statistics.
-Let $Y_{it}\in\left\{0,1\right\}$ denote the labor force participation status of single mother $i$ in year $t$.
-Let $X_{it}$ denote her number of children.
-To assess the impact of $X_{it}$ on $Y_{it}$, we estimate the following two logit regressions:
-$Y_{it}$ $=\mathds{1}\left\{\rho Y_{it-1}+\beta X_{it}+\mu_{i}+\varphi_{t}+\varepsilon_{it}>0\right\}$ (5.1)
-$Y_{it}$ $=\mathds{1}\left\{\rho Y_{it-1}+\left(\beta+\alpha_{i}+\gamma_{t}\right)X_{it}+\mu_{i}+\varphi_{t}+\varepsilon_{it}>0\right\},$ (5.2)
-where $\varepsilon_{it}$ follows a standard logistic distribution, and $\sum_{i}\mu_{i}=\sum_{i}\alpha_{i}=\sum_{i}\gamma_{t}=0$.
-Model (5.1) is a dynamic homogeneous slope model where $\beta$ represents the average effect of an additional child on the log odds of a single mother’s LFP.
-Model (5.2) is a dynamic heterogeneous slope model, where $\beta_{it}=\beta+\alpha_{i}+\gamma_{t}$ captures the heterogeneous impact $X_{it}$ may have on the log odds of $Y_{it}$.
-Model (5.2) nests model (5.1) as a special case when $\beta_{it}$ does not depend on $i$ and $t$.
-As discussed in section 1 , when the true impact of $X_{it}$ varies among the population, model (5.1) is appropriate only if $\beta_{it}$ and $X_{it}$ are uncorrelated.
-In practice, they may correlate due to self-selection.
-For instance, single mothers who choose to have more children might be those with sufficient financial resources or access to childcare services, such that their labor force participation is less affected by additional children.
-Alternatively, a correlation between $\beta_{it}$ and $X_{it}$ could arise from common time trends.
-As we track single mother households from 1987 to 1996, more children were born while factors such as rising childcare and schooling costs, improvements in home technology, and evolving welfare policies changed the impact of each additional child on their mother’s LFP.
-In both scenarios, model (5.2) would be the appropriate model to use as it controls for unobserved heterogeneity in $\beta_{it}$ by incorporating individual and time fixed effects into the slope and allowing them to correlate with $X_{it}$.
-Given our small sample size, direct estimation of both models could suffer from severe incidental parameter problems, which we address using our bias correction procedure.
-Table 7 presents the estimation results.
-Panel A reports results for model (5.1) under the homogeneous slope assumption.
-For each parameter, we provide the MLE and bias-corrected estimates, along with their standard errors.
-Examining the bias-corrected estimates, we observe that an increase in the number of children appears to reduce a single mother’s likelihood of labor force participation, while past LFP status strongly predicts current LFP.
-However, this relationship shifts dramatically once we account for slope heterogeneity.
-Panel B presents the results for model (5.2), allowing for heterogeneous slopes.
-Here, the bias-corrected estimates reveal a positive average impact of additional children on a mother’s labor supply: each additional child on average increases the log odds of labor force participation by 0.126, which corresponds to an average increase of 5.5 percent in the likelihood of employment.
-This indicates that, on average, single mothers with more children are indeed more likely to work, likely due to the financial demands of supporting multiple children independently.
-Additionally, unlike the homogeneous slope model, the serial correlation in labor force participation remains statistically significant but diminishes in magnitude, suggesting that what appears to be strong state dependence in LFP under the homogeneous slope model may in part be driven by unobserved heterogeneity in how children affect labor supply decisions across individuals.
-Together, these findings underscore the importance of controlling for unobserved slope heterogeneity, as failing to do so can lead to substantially different conclusions.
-In addition to highlighting the difference between homogeneous and heterogeneous slope models, Table 7 demonstrates the importance of bias-correction in estimating dynamic nonlinear panel data models.
-Examining the results in Panel B, we observe that the MLE overestimates the impact of the number of children while underestimating the degree of serial correlation in LFP.
-Bias correction results in an 18 percent lower estimate of the former and a 57 percent higher estimate of the latter.
-Moreover, the corrected estimates yield smaller standard errors, improving statistical significance across all parameters.
-Finally, in Figure 2, we plot the distribution of the estimated $\beta_{it}$ from model (5.2), with summary statistics in Table C.2.
-Each $\beta_{it}$ estimate is constructed as the sum of the bias-corrected common effect $\beta$ and the estimated individual and time effects $\alpha_{i}$ and $\gamma_{t}$.
-These estimates reveal substantial heterogeneity in how the number of children influences single mothers’ labor force participation.
-While the mean effect is positive, over half of the individual-time observations show a decrease in LFP with each additional child, suggesting that the average effect alone may inadequately capture the diverse impacts across individuals and over time.
-This reinforces the importance of accounting for both individual and temporal variations in slope coefficients, as the impact of children on labor force participation can vary significantly based on unobserved factors unique to each mother and time period.
-In conclusion, our analysis demonstrates the efficacy of our likelihood-based bias correction method for estimating dynamic nonlinear models of labor supply with multidimensional heterogeneities.
-We show the importance of controlling for unobserved slope heterogeneity when it may correlate with regressors and the value of panel data models that accommodate two-way fixed effects in both the intercept and the slope.
-Our findings reveal that the number of children has, on average, a positive influence on a single mother’s labor force participation, but such effect is highly heterogeneous across individuals and time periods.
-This heterogeneity underscores the complex nature of labor supply decisions and the unique challenges each single mother faces in balancing work and child-rearing responsibilities.
-Table 7: Single Mother Labor Force Participation: Estimation Results
-         Number of Children          Lagged Participation
-Uncorrected Corrected Uncorrected Corrected
-                Panel A. Homogeneous Slope
-Model parameter -0.954 -0.880 0.584 0.535
-(0.240) (0.218) (0.113) (0.103)
-                Panel B. Heterogeneous Slope
-Model parameter 0.153 0.126 0.051 0.080
-(0.021) (0.000) (0.020) (0.000)
-  • Notes: Standard errors in parentheses.
-Panel A reports the estimation results of the homogeneous slope model.
-Panel B reports the estimation results of the heterogeneous slope model.
-The “uncorrected” and the “corrected” columns show respectively the MLE and the bias-corrected parameter estimates.
-Data source: PSID 1987 - 1997.
-Figure 2: Heterogeneous Impact of Children on Single Mother Labor Force Participation
-6 Final Remarks
-In this paper, we propose a likelihood-based analytical bias correction procedure for a general class of dynamic nonlinear heterogeneous parameter models subject to the incidental parameter problem.
-We give the analytical form of a corrected likelihood and show that such a corrected likelihood delivers point estimates that are asymptotically unbiased and test statistics that are asymptotically $\chi^{2}$-distributed.
-Simulation studies and empirical analyses support our claims.
-Several issues are, however, not formally addressed in this paper.
-We conjecture, but have not formally investigated, that the average partial effects associated these models are asymptotically unbiased.
-Specifically, we conjecture that the rate of convergence of the average partial effects are only $O(1/\sqrt{T})$ or $O(1/\sqrt{N})$, slower than the potential bias, indicating that bias corrections may have no benefit in the limit.
-This result is formally shown by Fernández-Val and Weidner (2016) for models with additive fixed effects in the intercept.
-Second, we have not applied our procedure to general M-estimation models.
-We intuitively believe that our approach applies to them as long as the objective function is differentiable, only with minor modifications regarding the information matrix equality.
-We would like to dedicate these extensions to future research.
-Acknowledgement
-We would like to express our gratitude to valuable comments and suggestions from Geert Dhaene, Yanqin Fan, Jinyong Hahn, Chen Hsiao, Whitney Newey, and Jun Yu.
-Our research assistant Yuchi Liu helped us with various tasks for the empirical analysis.
-Leng’s research is partially supported by the National Natural Science Foundation of China Grant 72001183.
-Sun acknowledges the support from the National Natural Science Foundation of China under Grant Number 72203032.
-Appendix
-[…]
-AppendixAppendix
-A Proofs
-In this appendix, we prove the asymptotic properties of $L(\theta)$, the corrected likelihood, $\widehat{\theta}_{L}$ the corrected estimator, and $\{\widehat{\xi}_{LR},\widehat{\xi}_{LM},\widehat{\xi}_{Wald}\}$, the three test statistics we consider in this paper.
+D Relation to Existing Approaches
+As discussed above, our method is a non-trivial extension of Bester and Hansen (2009) and Arellano and Hahn (2016) to two-way DN-HP models.
+In this section, we explain the technical challenges in developing a likelihood-based bias correction approach for the DN-HP models and provide Example D.1 to illustrate how our method relates to the aforecited.
+We also leave Remark 7 to briefly discuss the difference of our procedure to some existing methods providing fixed-$T$ consistency for the logit model
+The primary technical challenge lies in the high dimensionality of the two-way DN-HP models, due to incidental parameter vector $\phi$ (or $\psi$) having a dimension increasing with $N$ and $T$.
+In what follows, we briefly re-visit the procedure of Arellano and Hahn (2016), pointing out the key difference, and explain the challenges.
+For simplicity, we assume that the fixed effects are only scalars [^8] We will briefly discuss multi-dimensional fixed effects later..
+Consider an arbitrary panel data model with individual effects only, having likelihood function
+$\ell\left(\theta,\alpha\right):=\frac{1}{N}\sum_{i=1}^{N}l_{i}\left(\theta,\alpha_{i}\right),\qquad l_{i}\left(\theta,\alpha_{i}\right):=\frac{1}{T}\sum_{i=1}^{T}\log f\left(Y_{it}|X_{it};\theta,\alpha_{i}\right),$
+where $f\left(Y_{it}|X_{it};\theta,\alpha_{i}\right)$ is a known density and $\alpha:=\left(\alpha_{1},\ldots,\alpha_{N}\right)^{\prime}$.
+It is important to note that $l_{i}\left(\theta,\alpha_{i}\right)$ are independent across $i$, in the sense that $l_{i}\left(\theta,\alpha_{i}\right)$ and $l_{j}\left(\theta,\alpha_{j}\right)$ share no common arguments except $\theta$.
+Because of such independence, an expansion similar to (2.5) can be established individual by individual, as
+$\quad\quad l_{i}\left(\theta,\widehat{\alpha}_{i}\left(\theta\right)\right)$
+$=l_{i}\left(\theta,\alpha_{i}\left(\theta\right)\right)+\partial_{\alpha_{i}}l_{i}\left(\theta,\alpha_{i}\left(\theta\right)\right)\left(\widehat{\alpha}_{i}\left(\theta\right)-\alpha_{i}\left(\theta\right)\right)+\frac{1}{2}\partial_{\alpha_{i}^{2}}l_{i}\left(\theta,\alpha_{i}\left(\theta\right)\right)\left(\widehat{\alpha}_{i}\left(\theta\right)-\alpha_{i}\left(\theta\right)\right)^{2}+r_{i}^{l}\left(\theta\right),$ (D.1)
+where, for some $c_{i}\in\left(0,1\right)$,
+$r_{i}^{l}\left(\theta\right)$ $:=\frac{1}{6}\partial_{\alpha_{i}^{3}}l_{i}\left(\theta,c_{i}\alpha_{i}\left(\theta\right)+\left(1-c_{i}\right)\alpha_{i}\left(\theta\right)\right)\left(\widehat{\alpha}_{i}\left(\theta\right)-\alpha_{i}\left(\theta\right)\right)^{3},$
+$\widehat{\alpha}_{i}\left(\theta\right)$ $:=\arg\max_{\alpha_{i}}l_{i}\left(\theta,\alpha_{i}\right),\qquad\alpha_{i}\left(\theta\right):=\arg\max_{\alpha_{i}}\mathbb{E}l_{i}\left(\theta,\alpha_{i}\right),$ (D.2)
+$\partial_{\alpha_{i}}l_{i}\left(\theta,\alpha_{i}\left(\theta\right)\right)$ denotes $\left[\partial_{\alpha_{i}}l_{i}\left(\theta,\alpha_{i}\right)\right]|_{\alpha_{i}=\alpha_{i}\left(\theta\right)}$, and $\partial_{\alpha_{i}^{2}}l_{i}\left(\cdot,\cdot\right)$ and $\partial_{\alpha_{i}^{3}}l_{i}\left(\cdot,\cdot\right)$ are defined analogously.
+Here, notice importantly that $\widehat{\alpha}_{i}\left(\theta\right)$ and $\alpha_{i}\left(\theta\right)$ are obtained by solving, respectively, $\partial_{\alpha_{i}}l_{i}\left(\theta,\alpha_{i}\right)=0$ and $\mathbb{E}\partial_{\alpha_{i}}l_{i}\left(\theta,\alpha_{i}\right)=0$ individually.
+This is possible due to the independence of $l_{i}\left(\theta,\alpha_{i}\right)$ across $i$.
+Subsequently, a similar expansion to (2.6) can be constructed, again individually, as
+$0$ $=\partial_{\alpha_{i}}l_{i}\left(\theta,\alpha_{i}\left(\theta\right)\right)+\partial_{\alpha_{i}^{2}}l_{i}\left(\theta,\alpha_{i}\left(\theta\right)\right)\left(\widehat{\alpha}_{i}\left(\theta\right)-\alpha_{i}\left(\theta\right)\right)+r_{i}^{s}\left(\theta\right)$ (D.3)
+$\widehat{\alpha}_{i}\left(\theta\right)-\alpha_{i}\left(\theta\right)$ $=-\frac{1}{\partial_{\alpha_{i}^{2}}l_{i}\left(\theta,\alpha_{i}\left(\theta\right)\right)}\left(\partial_{\alpha_{i}}l_{i}\left(\theta,\alpha_{i}\left(\theta\right)\right)+r_{i}^{s}\left(\theta\right)\right),$
+where, for some $d_{i}\in\left(0,1\right)$,
+$r_{i}^{s}\left(\theta\right)=\frac{1}{2}\partial_{\alpha_{i}^{3}}l_{i}\left(\theta,d_{i}\alpha_{i}\left(\theta\right)+\left(1-d_{i}\right)\alpha_{i}\left(\theta\right)\right)\left(\widehat{\alpha}_{i}\left(\theta\right)-\alpha_{i}\left(\theta\right)\right)^{2}.$
+Similarly combining (D.3) and (D.1), we have
+$l_{i}\left(\theta,\widehat{\alpha}_{i}\left(\theta\right)\right)=l_{i}\left(\theta,\alpha_{i}\left(\theta\right)\right)-\frac{1}{2}\frac{\left[\partial_{\alpha_{i}}l_{i}\left(\theta,\alpha_{i}\left(\theta\right)\right)\right]^{2}}{\partial_{\alpha_{i}^{2}}l_{i}\left(\theta,\alpha_{i}\left(\theta\right)\right)}-r_{i}\left(\theta\right)$
+for a remainder term $r_{i}\left(\theta\right)$ depending on $r_{i}^{s}\left(\theta\right)$ and $r_{i}^{l}\left(\theta\right)$, amongst other things.
+The subsequent steps are omitted because they are similar to their DN-HP counterparts.
+The steps above are all followed only individually.
+This is possible due to the independence of $l_{i}\left(\theta,\alpha_{i}\right)$; and the benefit is that $r_{i}\left(\theta\right)$, $\partial_{\alpha_{i}^{2}}l_{i}\left(\theta,\alpha_{i}\left(\theta\right)\right)$, etc.
+are only scalars.
+For two-way DN-HP models, the likelihood $l_{i,t}(\theta,\phi)=\log f(Y_{it}|X_{it},\theta+\alpha_{i}+\gamma_{t})$ (defined in 2.2) depends on both $\alpha_{i}$ and $\gamma_{t}$.
+This is unlike models with individual effects only, and it implies that $\widehat{\psi}\left(\theta\right)$ must be obtained by solving the system of equations $\partial_{\psi}l\left(\theta,\psi\right)=0$ jointly, unlike $\widehat{\alpha}_{i}\left(\theta\right)$.
+It further implies that the expansions in (2.5) and (2.6) must be carried out jointly w.r.t.
+the entire vector $\psi$.
+In other words, these expansions are high-dimenional for DN-HP models.
+Consequently, it is much more involved to develop a likelihood-based bias correction for two-way DN-HP models, in the follows notable ways, all due to the high-dimensionality.
+  1. 1.
+    For models with individuals effects only, $r_{i}^{s}\left(\theta\right)$ is only a scalar and it is relatively straightforward to establish its rate of convergence.
+On the other hand, its DN-HP counterpart, given in (E.6), is a high-dimensional vector, with each component itself being a quadratic form of $\widehat{\psi}\left(\theta\right)-\psi(\theta)$ and $G_{\psi_{j}}(\theta,a_{j})$ that are already high-dimensional objects.
+It is technically more challenging to control the rate of convergence of (E.6).
+For instance, one obvious issue is that the choice of norm becomes relevant for (E.6) but not $r_{i}^{s}\left(\theta\right)$.
+The same arguments apply to other remainder terms.
+  2. 2.
+    When the model only contains (scalar) individual effects, the incidental parameter Hessian $\partial_{\alpha_{i}^{2}}l_{i}\left(\theta,\alpha_{i}\left(\theta\right)\right)$ is only a scalar and the rate of convergence of its inverse, $1/\partial_{\alpha_{i}^{2}}l_{i}\left(\theta,\alpha_{i}\left(\theta\right)\right)$, is relatively straightforward.
+For the DN-HP models, however, the incidental parameter Hessian $H\left(\theta\right)$ is a high-dimensional matrix.
+It is more complex to control the rate of its inverse, $\left[H\left(\theta\right)\right]^{-1}$.
+  3. 3.
+    When only individual effects are present, there is only one scalar bias term.
+It is relatively straightforward to derive its rate of convergence and propose an estimator for it.
+On the contrary, for the DN-HP models, the bias correction term $B(\theta)$, defined in (2.7), is a quadratic form of $s\left(\theta\right)$ and $H\left(\theta\right)$ which are high-dimensional objects.
+In addition to this, we show that $B(\theta)\$can be approximated by $B_{\alpha}(\theta)+B_{\gamma}(\theta)$, which is not needed for models with individual effects only.
+Other challenges include, e.g., the lack of stationarity across $t$ (due to the existence of $\gamma_{t}^{0}$), the need for the normalization and reparameterization, etc.
+Note that, for models with individual effects only, even if $\alpha_{i}$ is multi-dimensional, the steps above can still be carried out individually, implying that the above-discussed high-dimensionality problems still would not happen, as long as the dimension of $\alpha_{i}$ is fixed.
+Next, in Example D.1 below, we illustrate how our approach reduces to Arellano and Hahn (2016) and apply it to a dynamic logit model with only scalar individual effects in the intercept.
+Example D.1 (dynamic logit with individual effects only)
+Consider a model with only scalar individual effects in the intercept, whose likelihood for individual $i$ at time $t$ is $l_{i,t}\left(\theta,\alpha_{i}\right)$, and assume that the data is independent over $i$ while stationary across $t$.
+Since, reparameterization is not needed, we work with the original likelihood $\ell\left(\theta,\phi\right)\equiv N^{-1}T^{-1}\sum_{i=1}^{N}\sum_{t=1}^{T}l_{i,t}\left(\theta,\alpha_{i}\right)$ for $\phi\equiv\left(\alpha_{1},\ldots,\alpha_{N}\right)^{\prime}$.
+Given that $l_{i,t}\left(\theta,\alpha_{i}\right)$ are independent across $i$, $\widehat{\phi}(\theta)$ and $\phi(\theta)$ can be constructed individually, as
+$\widehat{\phi}(\theta)\equiv\left(\widehat{\alpha}_{1}\left(\theta\right),\ldots,\widehat{\alpha}_{N}\left(\theta\right)\right)^{\prime},\qquad\phi(\theta)\equiv\left(\alpha_{1}\left(\theta\right),\ldots,\alpha_{N}\left(\theta\right)\right)^{\prime},$
+where, $\widehat{\alpha}_{i}\left(\theta\right)$ and $\alpha_{i}\left(\theta\right)$ are defined in (D.2).
+Next, notice that $\mathbb{E}s_{i,t}^{\alpha}(\theta,\alpha_{i}(\theta))=\mathbb{E}T^{-1}\sum_{t=1}^{T}s_{i,t}^{\alpha}(\theta,\alpha_{i}(\theta))=0$, where the first equality is because of stationarity and the second because $\alpha_{i}(\theta)$ maximizes $\mathbb{E}l_{i}\left(\theta,\alpha_{i}\right)$.
+This implies that $S_{\alpha\alpha}(\theta)$ is a diagonal matrix defined as
+$S_{\alpha\alpha}(\theta)=\frac{1}{N^{2}T}\operatorname*{diag}\left\{S_{i}^{\alpha\alpha}\left(\theta\right)\right\},\qquad S_{i}^{\alpha\alpha}\left(\theta\right):=\frac{1}{T}\sum_{t=1}^{T}\sum_{s=1}^{T}\mathbb{E}\left\{s_{i,t}^{\alpha}(\theta,\alpha_{i}(\theta))s_{i,s}^{\alpha}(\theta,\alpha_{i}(\theta))\right\}.$
+Similarly, the expected Hessian is also diagonal, as
+$\mathbb{E}H(\theta)=\frac{1}{N}\operatorname*{diag}\left\{\mathbb{E}h_{i}^{\alpha\alpha}(\theta)\right\},\qquad h_{i}^{\alpha\alpha}(\theta):=\frac{1}{T}\sum_{t=1}^{T}\partial_{\alpha_{i}^{2}}l_{i,t}\left(\theta,\alpha_{i}\left(\theta\right)\right).$
+By stationarity, the inverse of $\mathbb{E}H(\theta)$ is simply
+$H_{\alpha\alpha}^{\mathcal{\ast}}(\theta)\equiv[\mathbb{E}H(\theta)]^{-1}=N\operatorname*{diag}\left\{[\mathbb{E}\partial_{\alpha_{i}^{2}}l_{i,t}\left(\theta,\alpha_{i}\left(\theta\right)\right)]^{-1}\right\}.$
+Finally, seeing $B_{\gamma}(\theta)\equiv 0$ without time effects and using the properties of $\operatorname*{trace}$, we have
+$B(\theta)\equiv B_{\alpha}(\theta)=\frac{1}{T}\left(\frac{1}{N}\sum_{i=1}^{N}B_{i}^{\alpha}(\theta)\right),\qquad B_{i}^{\alpha}(\theta):=\frac{1}{2}\frac{T^{-1}\sum_{t=1}^{T}\sum_{s=1}^{T}\mathbb{E}\left\{s_{i,t}^{\alpha}(\theta,\alpha_{i}(\theta))s_{i,s}^{\alpha}(\theta,\alpha_{i}(\theta))\right\}}{\mathbb{E}\partial_{\alpha_{i}^{2}}l_{i,t}\left(\theta,\alpha_{i}\left(\theta\right)\right)}.$
+This is identical to the bias term of Arellano and Hahn (2016). The term $B_{i}^{\alpha}(\theta)$ is estimated by
+$\widehat{B}_{i}^{\alpha}(\theta):=\frac{1}{2}\frac{\sum_{t=1}^{T}\sum_{s=1}^{T}\left(\mathds{1}\{\left|t-s\right|\leq\tau\}s_{i,t}^{\alpha}(\theta,\widehat{\alpha}_{i}(\theta))s_{i,s}^{\alpha}(\theta,\widehat{\alpha}_{i}(\theta))\right)}{\sum_{t=1}^{T}\partial_{\alpha_{i}^{2}}l_{i,t}\left(\theta,\alpha_{i}\left(\theta\right)\right)},$ (D.4)
+which is the same to the trace-based approach of Arellano and Hahn (2016), except that we use a truncation here.
+Next, consider a dynamic logit model $Y_{it}=\mathds{1}\left\{\theta_{0}Y_{it-1}+\alpha_{i}^{0}+\varepsilon_{it}>0\right\}$, where $Y_{it}$ is independent over $i$ and stationary across $t$, and $\varepsilon_{it}$ is standard logistic.
+The likelihood function for individual $i$ at time $t$ is
+$l_{i,t}\left(\theta,\alpha_{i}\right)=\left(Y_{i,t}-1\right)\pi_{i,t}\left(\theta,\alpha_{i}\right)-\log\left(1+\exp\left(-\pi_{i,t}\left(\theta,\alpha_{i}\right)\right)\right),\qquad\pi_{i,t}\left(\theta,\alpha_{i}\right):=\theta Y_{it-1}+\alpha_{i};$
+while the first and second derivatives of $l_{i,t}\left(\theta,\alpha_{i}\right)$ w.r.t. $\alpha_{i}$ are, respectively,
+$s_{i,t}^{\alpha}\left(\theta,\alpha_{i}\right)=Y_{it}-\frac{1}{\exp\left(-\pi_{i,t}\left(\theta,\alpha_{i}\right)\right)+1},\qquad\partial_{\alpha_{i}^{2}}l_{i,t}\left(\theta,\alpha_{i}\left(\theta\right)\right)=\frac{-\exp\left(-\pi_{i,t}\left(\theta,\alpha_{i}\right)\right)}{\left[\exp\left(-\pi_{i,t}\left(\theta,\alpha_{i}\right)\right)+1\right]^{2}}.$ (D.5)
+Substituting (D.5) into (D.4) and letting $\widehat{l}(\theta)\equiv\ell(\theta,\widehat{\phi}(\theta))$, we have
+$L(\theta)\equiv\widehat{l}(\theta)+\frac{1}{T}\left(\frac{1}{N}\sum_{i=1}^{N}\widehat{B}_{i}^{\alpha}(\theta)\right),$
+where, denoting $\widehat{\pi}_{i,t}\left(\theta\right):=\pi_{i,t}\left(\theta,\widehat{\alpha}_{i}\left(\theta\right)\right)$,
+$\widehat{B}_{i}^{\alpha}(\theta):=\frac{1}{2}\left(\sum_{t=1}^{T}\frac{-\exp\left(-\widehat{\pi}_{i,t}\left(\theta\right)\right)}{\left[\exp\left(-\widehat{\pi}_{i,t}\left(\theta\right)\right)+1\right]^{2}}\right)^{-1}\left(\sum_{t=2}^{T}\sum_{s=2}^{T}\mathds{1}\{\left|t-s\right|\leq\tau\}[Y_{it}Y_{is}+\widehat{b}_{i,t,s}^{\alpha,1}(\theta)+\widehat{b}_{i,t,s}^{\alpha,2}(\theta)]\right),$
+$\widehat{b}_{i,t,s}^{\alpha,1}(\theta):=-2Y_{it}\left[\exp\left(-\widehat{\pi}_{i,s}\left(\theta\right)\right)+1\right]^{-1},\qquad\widehat{b}_{i,t,s}^{\alpha,2}(\theta):=\left[\exp\left(-\widehat{\pi}_{i,t}\left(\theta\right)\right)+1\right]^{-1}\left[\exp\left(-\widehat{\pi}_{i,s}\left(\theta\right)\right)+1\right]^{-1}.$
+We use the following remark to briefly discuss the difference between our bias correction method and several existing approaches that remove the IPP for logit models.
+Remark 7 (panel logit models)
+For static logit models with individual effects in the intercept, Chamberlain (1980) develops a conditional maximum likelihood, on some sufficient statistics, that produces a fixed-$T$ consistent estimator for $\theta_{0}$.
+For dynamic models, Chamberlain (1985) shows that such sufficient statistics can be found when the only regressor is the lagged dependent variable.
+Honoré and Kyriazidou (2000) and Honoré and Weidner (2024) provide methods for dynamic logit models with both lagged and exogenous regressors.
+When an additional time effect is included in the model, Charbonneau (2017) and Dano et al.
+(2025) provide consistent estimation procedures.
+To our knowledge, there has been no fixed-$T$ consistent estimator available for logit models two-way heterogeneities.
+As discussed in Remark 2, our approach (and that of Bester and Hansen (2009) and Arellano and Hahn (2016)) does not generally eliminate the IPP bias.
+This is true even for the logit model: the simulation results from Dhaene and Sun (2021) show that a “second-order” corrected logit estimator is less biased than the first-order corrected.
+This implies that the first-order corrected estimator is likely not fixed-$T$ consistent.
+E Proofs
+In this section, we prove the asymptotic properties of $L(\theta)$, the corrected likelihood, $\widehat{\theta}_{L}$ the corrected estimator, and $\{\widehat{\xi}_{LR},\widehat{\xi}_{LM},\widehat{\xi}_{Wald}\}$, the three test statistics we consider in this paper.
 To continue, we first introduce three Lemmas.
@@
 where $H_{\alpha\alpha^{\prime}}\left(\theta\right)$ is $\left(N-1\right)K\times\left(N-1\right)K$, $H_{\gamma\gamma^{\prime}}\left(\theta\right)$ is $\left(T-1\right)K\times\left(T-1\right)K$, $H_{\alpha\gamma^{\prime}}\left(\theta\right)$ is $\left(N-1\right)K\times\left(T-1\right)K$, and $H_{\gamma\alpha^{\prime}}\left(\theta\right)=\left[H_{\alpha\gamma^{\prime}}\left(\theta\right)\right]^{\prime}$.
-The next lemma states that the quadratic form $\mathbb{E}\left\{s^{\prime}(\theta)[{H}\left(\theta\right)]^{-1}s(\theta)\right\}$ can be decomposed into two dominating terms of order $O(1/\sqrt{NT})$.
+The next lemma states that the quadratic form can be decomposed into two dominating terms of order $O(1/\sqrt{NT})$.
 They establish the two additive bias terms $B_{\alpha}(\theta)$ and $B_{\gamma}(\theta)$, arising from estimating individual effects and time effects, respectively.
-Lemma A.1
+Lemma E.1
 Under Assumption 1, we have
@@
 $=[s(\theta)-\mathbb{E}s(\theta)]^{\prime}[-\mathbb{E}{H}\left(\theta\right)]^{-1}[s(\theta)-\mathbb{E}s(\theta)]^{\prime}(1+o_{\mathbb{P}}(1))$
-$=-\left[\widetilde{s}_{\alpha}^{\prime}(\theta)H_{\alpha\alpha^{\prime}}^{\ast}(\theta)\widetilde{s}_{\alpha}(\theta)+\widetilde{s}_{\gamma}^{\prime}(\theta)H_{\gamma\gamma^{\prime}}^{\ast}(\theta)\widetilde{s}_{\gamma}(\theta)+2\widetilde{s}_{\alpha}^{\prime}(\theta)H_{\alpha\gamma^{\prime}}^{\ast}(\theta)\widetilde{s}_{\gamma}\left(\theta\right)\right](1+o_{\mathbb{P}}(1)).$ (A.1)
+$=-\left[\widetilde{s}_{\alpha}^{\prime}(\theta)H_{\alpha\alpha^{\prime}}^{\ast}(\theta)\widetilde{s}_{\alpha}(\theta)+\widetilde{s}_{\gamma}^{\prime}(\theta)H_{\gamma\gamma^{\prime}}^{\ast}(\theta)\widetilde{s}_{\gamma}(\theta)+2\widetilde{s}_{\alpha}^{\prime}(\theta)H_{\alpha\gamma^{\prime}}^{\ast}(\theta)\widetilde{s}_{\gamma}\left(\theta\right)\right](1+o_{\mathbb{P}}(1)).$ (E.1)
 Since $K$ is fixed, we assume without loss of generality that $K=1$. In what follows, we show
-$\sup_{\theta\in\Theta}\mathbb{E}\left[\widetilde{s}_{\alpha}^{\prime}(\theta)H_{\alpha\gamma^{\prime}}^{\ast}(\theta)\widetilde{s}_{\gamma}\left(\theta\right)\right]=O\left(1/(NT)\right),$ (A.2)
+$\sup_{\theta\in\Theta}\mathbb{E}\left[\widetilde{s}_{\alpha}^{\prime}(\theta)H_{\alpha\gamma^{\prime}}^{\ast}(\theta)\widetilde{s}_{\gamma}\left(\theta\right)\right]=O\left(1/(NT)\right),$ (E.2)
 and
-$\sup_{\theta\in\Theta}\mathbb{E}\left[\widetilde{s}_{\alpha}^{\prime}(\theta)H_{\alpha\alpha^{\prime}}^{\ast}(\theta)\widetilde{s}_{\alpha}(\theta)\right]=O(1/T),\qquad\sup_{\theta\in\Theta}\mathbb{E}\left[\widetilde{s}_{\gamma}^{\prime}(\theta)H_{\gamma\gamma^{\prime}}^{\ast}(\theta)\widetilde{s}_{\gamma}(\theta)\right]=O(1/N).$ (A.3)
+$\sup_{\theta\in\Theta}\mathbb{E}\left[\widetilde{s}_{\alpha}^{\prime}(\theta)H_{\alpha\alpha^{\prime}}^{\ast}(\theta)\widetilde{s}_{\alpha}(\theta)\right]=O(1/T),\qquad\sup_{\theta\in\Theta}\mathbb{E}\left[\widetilde{s}_{\gamma}^{\prime}(\theta)H_{\gamma\gamma^{\prime}}^{\ast}(\theta)\widetilde{s}_{\gamma}(\theta)\right]=O(1/N).$ (E.3)
 Thus, $\mathbb{E}\left[\widetilde{s}_{\alpha}^{\prime}(\theta)H_{\alpha\gamma^{\prime}}^{\ast}(\theta)\widetilde{s}_{\gamma}\left(\theta\right)\right]$ is negligible in the expansion relative to the other two terms from the quadratic form.
-First, under Assumptions 1 and 1, and by Davydov’s inequality, it follows that $\sup_{i,j}\sup_{\theta\in\Theta}\operatorname*{Var}[h_{i,j}^{\alpha}(\theta)]=o(1)$ as $N,T\rightarrow\infty$, where $\operatorname*{Var}$ denotes the variance and $h_{i,j}^{\alpha}(\theta)$ is the $\left(i,j\right)$-th element of $N\cdot H_{\alpha\alpha^{\prime}}\left(\theta\right)$ with $i,j=1,\ldots,N-1$.
+Under Assumptions 1 and 1, using Davydov’s inequality, we have $\sup_{i,j}\sup_{\theta\in\Theta}\operatorname*{Var}[h_{i,j}^{\alpha}(\theta)]=o(1)$ as $N,T\rightarrow\infty$, where $\operatorname*{Var}$ denotes the variance and $h_{i,j}^{\alpha}(\theta)$ is the $\left(i,j\right)$-th element of $N\cdot H_{\alpha\alpha^{\prime}}\left(\theta\right)$ with $i,j=1,\ldots,N-1$.
 Invoking the Markov’s inequality,
@@
 $NT\left\|H_{\alpha\gamma^{\prime}}(\theta)-\mathbb{E}[H_{\alpha\gamma^{\prime}}(\theta)]\right\|_{\max}$ $=o_{\mathbb{P}}(1),$
-uniformly for $\theta\in\Theta$, where $\|A\|_{\max}$ denotes the max norm of a matrix $A$. In addition, by matrix inversion in block form, it follows that
+uniformly for $\theta\in\Theta$, where $\|A\|_{\max}$ denotes the $\max$-norm of a matrix $A$. In addition, by matrix inversion in block form, it follows that
 $\sup_{\theta\in\Theta}\left\|H_{\alpha\alpha^{\prime}}^{\ast}(\theta)\right\|_{\max}$ $=O(T),\qquad\sup_{\theta\in\Theta}\left\|H_{\gamma\gamma^{\prime}}^{\ast}\left(\theta\right)\right\|_{\max}=O(N),$
 $\sup_{\theta\in\Theta}\left\|H_{\alpha\gamma^{\prime}}^{\ast}(\theta)\right\|_{\max}$ $=O(1).$
-For $i=1,\ldots,N-1$ and $t=1,\ldots,T-1$, define $h_{i,t}^{\ast\alpha\gamma}(\theta)$ as the $(i,t)$-th element of $H_{\alpha\gamma^{\prime}}^{\ast}(\theta)$. We first show (A.2). We have
+For $i=1,\ldots,N-1$ and $t=1,\ldots,T-1$, define $h_{i,t}^{\ast\alpha\gamma}(\theta)$ as the $(i,t)$-th element of $H_{\alpha\gamma^{\prime}}^{\ast}(\theta)$. We first show (E.2). We have
 $\quad\quad\mathbb{E}\left\{(N\sqrt{T}\widetilde{s}_{\alpha}^{\prime}\left(\theta\right))H_{\alpha\gamma^{\prime}}^{\ast}\left(\theta\right)(\sqrt{N}T\widetilde{s}_{\gamma}(\theta))\right\}$
@@
 $\quad\quad+\frac{1}{\sqrt{NT}}\left(\sum_{i=1}^{N-1}\sum_{t=1}^{T-1}h_{i,t}^{\ast\alpha\gamma}(\theta)\right)\mathbb{E}\left(\sum_{s=1}^{T}\widetilde{s}_{N,s}^{\alpha}(\theta)\widetilde{s}_{N,T}^{\gamma}(\theta)\right)$
-$=:A_{NT,1}(\theta)-A_{NT,2}(\theta)-A_{NT,3}(\theta)+A_{NT,4}(\theta).$ (A.4)
+$=:A_{NT,1}(\theta)-A_{NT,2}(\theta)-A_{NT,3}(\theta)+A_{NT,4}(\theta).$ (E.4)
 By Assumptions 1 and 1, and using Davydov’s inequality, we obtain for some $\eta>2$ that
@@
 $\leq\sqrt{NT}\left(\sup_{\theta\in\Theta}\left\|H_{\alpha\gamma^{\prime}}^{\ast}(\theta)\right\|_{\max}\right)\left(\sup_{i,t}\{\mathbb{E}[g^{\eta}(w_{it})]\}^{2/\eta}\right)$
-×$\quad\quad\times\sup\left([a_{i}(0)]^{1-2/\eta}+2\sum_{m=1}^{T-2}\left[\left(1-\frac{m}{T-1}\right)[a_{i}(m)]^{1-2/\eta}\right]+\frac{1}{T-1}\sum_{m=1}^{T-1}[a_{i}(m)]^{1-2/\eta}\right)$
+$\quad\quad\times\sup\left([a_{i}(0)]^{1-2/\eta}+2\sum_{m=1}^{T-2}\left[\left(1-\frac{m}{T-1}\right)[a_{i}(m)]^{1-2/\eta}\right]+\frac{1}{T-1}\sum_{m=1}^{T-1}[a_{i}(m)]^{1-2/\eta}\right)$
 $=O(\sqrt{NT}).$
@@
 $=\sqrt{NT}\left(\sup_{\theta\in\Theta}\left\|H_{\alpha\gamma^{\prime}}^{\ast}(\theta)\right\|_{\max}\right)\left(\sup_{i,t}\{\mathbb{E}[g^{\eta}(w_{it})]\}^{2/\eta}\right)$
-×$\quad\quad\times\sup_{i}\left([a_{i}(0)]^{1-2/\eta}+2\sum_{m=1}^{T-2}\left[\left(1-\frac{m}{T-1}\right)[a_{i}(m)]^{1-2/\eta}\right]+\frac{1}{T-1}\sum_{m=1}^{T-1}[a_{i}(m)]^{1-2/\eta}\right)$
+$\quad\quad\times\sup_{i}\left([a_{i}(0)]^{1-2/\eta}+2\sum_{m=1}^{T-2}\left[\left(1-\frac{m}{T-1}\right)[a_{i}(m)]^{1-2/\eta}\right]+\frac{1}{T-1}\sum_{m=1}^{T-1}[a_{i}(m)]^{1-2/\eta}\right)$
 $=O(\sqrt{NT});$
@@
 $=O(\sqrt{NT}).$
-Therefore, (A.2) follows from (A.4).
+Therefore, (E.2) follows from (E.4).
 Let $h_{i,j}^{\ast\alpha}(\theta)$ denote the $(i,j)$-th element of $T^{-1}H_{\alpha\alpha^{\prime}}^{\ast}(\theta)$, $i,j=1,\ldots,N-1$.
@@
 $\quad\quad+\frac{2}{N}\sum_{1\leq i<j\leq N-1}h_{i,j}^{\ast\alpha}(\theta)\mathbb{E}\left[\left(\frac{1}{\sqrt{T}}\sum_{t=1}^{T}\left[\widetilde{s}_{i,t}^{\alpha}(\theta)-\widetilde{s}_{N,t}^{\alpha}(\theta)\right]\right)\left(\frac{1}{\sqrt{T}}\sum_{t=1}^{T}\left[\widetilde{s}_{j,t}^{\alpha}(\theta)-\widetilde{s}_{N,t}^{\alpha}(\theta)\right]\right)\right]$
-$=:I_{NT,1}(\theta)+I_{NT,2}(\theta).$ (A.5)
+$=:I_{NT,1}(\theta)+I_{NT,2}(\theta).$ (E.5)
 Here, by Assumptions 1 and 1, applying Davydov’s inequality gives, for some $\eta>2$,
@@
 $=O\left(1\right),$
-noting that $\sup_{i\neq j}h_{i,j}^{\ast\alpha}(\theta)=O(1/N)$ uniformly in $\theta$. Thus, the first part of (A.3) follows from (A.5). The second part of (A.3) follows from a similar procedure.
-Combining (A.1), (A.2), and (A.3),
+noting that $\sup_{i\neq j}h_{i,j}^{\ast\alpha}(\theta)=O(1/N)$ uniformly in $\theta$. Thus, the first part of (E.3) follows from (E.5). The second part of (E.3) follows from a similar procedure.
+Combining (E.1), (E.2), and (E.3),
 $\mathbb{E}\left\{s^{\prime}(\theta)[-{H}(\theta)]^{-1}s(\theta)\right\}$ $=\mathbb{E}\left\{\widetilde{s}_{\alpha}^{\prime}(\theta)H_{\alpha\alpha^{\prime}}^{\ast}(\theta)\widetilde{s}_{\alpha}(\theta)\right\}+\mathbb{E}\left\{\widetilde{s}_{\gamma}^{\prime}(\theta)H_{\gamma\gamma^{\prime}}^{\ast}(\theta)\widetilde{s}_{\gamma}(\theta)\right\}$
@@
 This leads to a different $a_{j}$ for each component of $r^{s}\left(\theta,a\right)$.
-$r^{s}\left(\theta,a\right)$ $:=\left(r_{\psi_{1}}^{s}\left(\theta,a_{1}\right),\ldots,r_{\psi_{(N+T-2)K}}^{s}(\theta,a_{(N+T-2)K})\right)^{\prime},$ (A.6)
+$r^{s}\left(\theta,a\right)$ $:=\left(r_{\psi_{1}}^{s}\left(\theta,a_{1}\right),\ldots,r_{\psi_{(N+T-2)K}}^{s}(\theta,a_{(N+T-2)K})\right)^{\prime},$ (E.6)
 $r_{\psi_{j}}^{s}\left(\theta,a_{j}\right)$ $:=\frac{1}{2}\left(\widehat{\psi}\left(\theta\right)-\psi(\theta)\right)^{\prime}G_{\psi_{j}}(\theta,a_{j})\left(\widehat{\psi}(\theta)-\psi(\theta)\right),$
@@
 Here $G_{\psi_{j}}(\theta,a_{j})$ is the derivative of each element of $H(\theta,\psi)$ with respect to $\psi_{j}$, evaluated at the point $a_{j}\widehat{\psi}\left(\theta\right)+\left(1-a_{j}\right)\psi\left(\theta\right)$.
-Lemma A.2
+Lemma E.2
 Under Assumption 1,
@@
 Proof.
-In the proof, we retain the notation and setup from Lemma A.1.
-From (2.13), it is evident that the dominant term in the expansion of $\widehat{\psi}(\theta)-\psi\left(\theta\right)$ is $[-{H}(\theta)]^{-1}s(\theta)$.
+In the proof, we retain the notation and setup from Lemma E.1.
+From (2.6), it is evident that the dominant term in the expansion of $\widehat{\psi}(\theta)-\psi\left(\theta\right)$ is ${H}(\theta)]^{-1}s(\theta)$.
 Specifically,
 $[-{H}(\theta)]^{-1}s(\theta)$ $=[-\mathbb{E}{H}(\theta)]^{-1}[{s}(\theta)-\mathbb{E}{s}(\theta)](1+o_{\mathbb{P}}(1))$
-$=-\begin{pmatrix}H_{\alpha\alpha^{\prime}}^{\ast}(\theta)\widetilde{s}_{\alpha}(\theta)+H_{\alpha\gamma^{\prime}}^{\ast}(\theta)\widetilde{s}_{\gamma}(\theta)\\ H_{\gamma\alpha^{\prime}}^{\ast}(\theta)\widetilde{s}_{\alpha}(\theta)+H_{\gamma\gamma^{\prime}}^{\ast}(\theta)\widetilde{s}_{\gamma}(\theta)\end{pmatrix}+[-\mathbb{E}{H}(\theta)]^{-1}[{s}(\theta)-\mathbb{E}{s}(\theta)]\cdot o_{\mathbb{P}}(1)$ (A.7)
+$=-\begin{pmatrix}H_{\alpha\alpha^{\prime}}^{\ast}(\theta)\widetilde{s}_{\alpha}(\theta)+H_{\alpha\gamma^{\prime}}^{\ast}(\theta)\widetilde{s}_{\gamma}(\theta)\\ H_{\gamma\alpha^{\prime}}^{\ast}(\theta)\widetilde{s}_{\alpha}(\theta)+H_{\gamma\gamma^{\prime}}^{\ast}(\theta)\widetilde{s}_{\gamma}(\theta)\end{pmatrix}+[-\mathbb{E}{H}(\theta)]^{-1}[{s}(\theta)-\mathbb{E}{s}(\theta)]\cdot o_{\mathbb{P}}(1)$ (E.7)
 where $o_{\mathbb{P}}(1)$ is uniform in both $i$ and $t$.
@@
 It follows that
-$N\sqrt{T}u_{i}^{\alpha}(\theta)=N\sum_{j=1}^{N-1}\left[h_{i,j}^{\ast\alpha}(\theta)\left(\frac{1}{\sqrt{T}}\sum_{t=1}^{T}\left[\widetilde{s}_{j,t}^{\alpha}(\theta)-\widetilde{s}_{N,t}^{\alpha}(\theta)\right]\right)\right].$ (A.8)
+$N\sqrt{T}u_{i}^{\alpha}(\theta)=N\sum_{j=1}^{N-1}\left[h_{i,j}^{\ast\alpha}(\theta)\left(\frac{1}{\sqrt{T}}\sum_{t=1}^{T}\left[\widetilde{s}_{j,t}^{\alpha}(\theta)-\widetilde{s}_{N,t}^{\alpha}(\theta)\right]\right)\right].$ (E.8)
 Using the result from Miller (1981), we know that $\sup_{i\geq 1}h_{i,i}^{\ast\alpha}(\theta)=O(1)$ and $\sup_{i\neq j}h_{i,j}^{\ast\alpha}(\theta)=O(1/N)$ uniformly in $\theta$, implying $\sum_{j=1}^{N-1}h_{i,j}^{\ast\alpha}(\theta)=O(1/N)$ uniformly in both $i$ and $\theta$.
@@
 Given that $\sup_{\theta\in\Theta}\|H_{\alpha\gamma^{\prime}}^{\ast}(\theta)\|_{\max}=O(1)$, we have
-$T\sqrt{N}u_{i}^{\alpha\gamma}(\theta)=\sum_{t=1}^{T-1}\left[h_{i,t}^{\ast\alpha\gamma}(\theta)\left(\frac{1}{\sqrt{N}}\sum_{j=1}^{N}\left[\widetilde{s}_{j,t}^{\gamma}(\theta)-\widetilde{s}_{j,T}^{\gamma}(\theta)\right]\right)\right].$ (A.9)
+$T\sqrt{N}u_{i}^{\alpha\gamma}(\theta)=\sum_{t=1}^{T-1}\left[h_{i,t}^{\ast\alpha\gamma}(\theta)\left(\frac{1}{\sqrt{N}}\sum_{j=1}^{N}\left[\widetilde{s}_{j,t}^{\gamma}(\theta)-\widetilde{s}_{j,T}^{\gamma}(\theta)\right]\right)\right].$ (E.9)
 Since $\sum_{t=1}^{T-1}h_{i,t}^{\ast\alpha\gamma}(\theta)=O(1)$, which follows from $\mathbb{E}[\partial_{\breve{\alpha}\breve{\gamma}^{\prime}}\ell(\theta,\psi)|_{\psi=\psi(\theta)}]=\partial_{\breve{\gamma}^{\prime}}\mathbb{E}[\partial_{\breve{\alpha}}\ell(\theta,\psi)|_{\psi=\psi(\theta)}]=0$, for $\breve{\alpha}:=\left(\alpha_{1}^{\prime},\ldots,\alpha_{N-1}^{\prime}\right)^{\prime}$ and $\breve{\gamma}:=\left(\gamma_{1}^{\prime},\ldots,\gamma_{T-1}^{\prime}\right)^{\prime}$, it holds that
 $\sum_{t=1}^{T-1}\left[h_{i,t}^{\ast\alpha\gamma}(\theta)\left(\frac{1}{\sqrt{N}}\sum_{j=1}^{N}\widetilde{s}_{j,t}^{\gamma}(\theta)\right)\right]=O_{\mathbb{P}}(1),\qquad\left(\sum_{t=1}^{T-1}h_{i,t}^{\ast\alpha\gamma}(\theta)\right)\left(\frac{1}{\sqrt{N}}\sum_{j=1}^{N}\widetilde{s}_{j,T}^{\gamma}(\theta)\right)=O_{\mathbb{P}}(1)$
-uniformly in $\theta$. From (A.8) and (A.9), we conclude that
-$\sup_{\theta\in\Theta}\|H_{\alpha\alpha^{\prime}}^{\ast}(\theta)\widetilde{s}_{\alpha}(\theta)\|_{\max}=O_{\mathbb{P}}\left(1/(N\sqrt{T})\right),\qquad\sup_{\theta\in\Theta}\|H_{\alpha\gamma^{\prime}}^{\ast}(\theta)\widetilde{s}_{\gamma}(\theta)\|_{\max}=O_{\mathbb{P}}\left(1/(T\sqrt{N})\right).$ (A.10)
+uniformly in $\theta$. From (E.8) and (E.9), we conclude that
+$\sup_{\theta\in\Theta}\|H_{\alpha\alpha^{\prime}}^{\ast}(\theta)\widetilde{s}_{\alpha}(\theta)\|_{\max}=O_{\mathbb{P}}\left(1/(N\sqrt{T})\right),\qquad\sup_{\theta\in\Theta}\|H_{\alpha\gamma^{\prime}}^{\ast}(\theta)\widetilde{s}_{\gamma}(\theta)\|_{\max}=O_{\mathbb{P}}\left(1/(T\sqrt{N})\right).$ (E.10)
 Following a similar proof, we also have
-$\sup_{\theta\in\Theta}\|H_{\gamma\gamma^{\prime}}^{\ast}(\theta)\widetilde{s}_{\gamma}(\theta)\|_{\max}=O_{\mathbb{P}}\left(1/(T\sqrt{N})\right),\qquad\sup_{\theta\in\Theta}\|H_{\gamma\alpha^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\alpha}}(\theta)\|_{\max}=O_{\mathbb{P}}\left(1/(N\sqrt{T})\right).$ (A.11)
-Thus, combining (A.7), (A.10), and (A.11), we have
+$\sup_{\theta\in\Theta}\|H_{\gamma\gamma^{\prime}}^{\ast}(\theta)\widetilde{s}_{\gamma}(\theta)\|_{\max}=O_{\mathbb{P}}\left(1/(T\sqrt{N})\right),\qquad\sup_{\theta\in\Theta}\|H_{\gamma\alpha^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\alpha}}(\theta)\|_{\max}=O_{\mathbb{P}}\left(1/(N\sqrt{T})\right).$ (E.11)
+Thus, combining (E.7), (E.10 ), and (E.11), we have
 $\left\|[-{H}(\theta)]^{-1}s(\theta)+[\mathbb{E}{H}(\theta)]^{-1}[{s}(\theta)-\mathbb{E}{s}(\theta)]\right\|_{\max}=o_{\mathbb{P}}\left(1/(NT)^{3/4}\right),$
@@
 we further have
-$\left\|\widehat{\psi}(\theta)-\psi(\theta)+[\mathbb{E}{H}(\theta)]^{-1}[{s}(\theta)-\mathbb{E}{s}(\theta)]\right\|_{\max}=o_{\mathbb{P}}\left(1/(NT)^{3/4}\right).$ (A.12)
-From (A.12), we conclude that $\sup_{\theta\in\Theta}\|G_{\psi_{j}}\left(\theta,a_{j}\right)-\partial_{\psi_{j}}H(\theta)\|_{\max}=o_{\mathbb{P}}(1)$.
+$\left\|\widehat{\psi}(\theta)-\psi(\theta)+[\mathbb{E}{H}(\theta)]^{-1}[{s}(\theta)-\mathbb{E}{s}(\theta)]\right\|_{\max}=o_{\mathbb{P}}\left(1/(NT)^{3/4}\right).$ (E.12)
+From (E.12), we conclude that $\sup_{\theta\in\Theta}\|G_{\psi_{j}}\left(\theta,a_{j}\right)-\partial_{\psi_{j}}H(\theta)\|_{\max}=o_{\mathbb{P}}(1)$.
 Without loss of generality, let $\psi_{j}=\alpha_{i}$ for $i=1,\ldots,N-1$.
@@
 $\quad\quad+\left(\partial_{\alpha_{N}\alpha_{N}\gamma_{1}}l_{N,1}(\theta,\phi)|_{\phi=\phi(\theta)},\ldots,\partial_{\alpha_{N}\alpha_{N}\gamma_{T-1}}l_{N,T-1}(\theta,\phi)|_{\phi=\phi(\theta)}\right)\otimes\iota_{N-1},$
-where $e_{i}$ is an $(N-1)\times 1$ vector with the $i$-th component equal to $1$ and all other elements equal to 00. Using (A.12), for each $j=1,\ldots,N-1$, we write
+where $e_{i}$ is an $(N-1)\times 1$ vector with the $i$-th component equal to $1$ and all other elements equal to 0. Using (E.12), for each $j=1,\ldots,N-1$, we write
 $\quad\quad\sup_{a_{j}\in(0,1)}r_{\psi_{j}}^{s}(\theta,a_{j})$
@@
 $\quad\quad\left.+2\left(H_{\alpha\alpha^{\prime}}^{\ast}(\theta)\widetilde{s}_{\alpha}(\theta)+H_{\alpha\gamma^{\prime}}^{\ast}(\theta)\widetilde{s}_{\gamma}(\theta)\right)^{\prime}\partial_{\alpha_{i}}{H}_{\breve{\alpha}\breve{\gamma}^{\prime}}(\theta)\left(H_{\gamma\alpha^{\prime}}^{\ast}(\theta)\widetilde{s}_{\alpha}(\theta)+H_{\gamma\gamma^{\prime}}^{\ast}(\theta)\widetilde{s}_{\gamma}(\theta)\right)\right\}\left[1+o_{\mathbb{P}}(1)\right]$
-$=:\frac{1}{2}\left\{C_{NT,1}(\theta)+C_{NT,2}(\theta)+C_{NT,3}(\theta)\right\}\left[1+o_{\mathbb{P}}(1)\right].$ (A.13)
+$=:\frac{1}{2}\left\{C_{NT,1}(\theta)+C_{NT,2}(\theta)+C_{NT,3}(\theta)\right\}\left[1+o_{\mathbb{P}}(1)\right].$ (E.13)
 Let $v_{i}^{\alpha}(\theta)$ denote the $i$-th element of $H_{\alpha\alpha^{\prime}}^{\ast}(\theta)\widetilde{s}_{\alpha}(\theta)+H_{\alpha\gamma^{\prime}}^{\ast}(\theta)\widetilde{s}_{\gamma}(\theta)$, for $i=1,\ldots,N-1$.
-Using (A.8) and (A.9), it can be shown that
+Using (E.8) and (E.9), it can be shown that
 $\quad\quad C_{NT,1}(\theta)$
@@
 $=-\frac{1}{2N}\left(\sum_{i=1}^{N-1}v_{i}^{\alpha}(\theta)\right)^{2}\left(\frac{1}{T}\sum_{t=1}^{T}\partial_{\alpha_{N}\alpha_{N}\alpha_{N}}l_{N,t}(\theta,\phi)|_{\phi=\phi(\theta)}\right)+o_{\mathbb{P}}\left(1/(NT)\right)$
-$=O_{\mathbb{P}}\left(1/(NT)\right).$ (A.14)
+$=O_{\mathbb{P}}\left(1/(NT)\right).$ (E.14)
 Similarly, when $\psi_{j}=\gamma_{t}$ ($t=1,\ldots,T-1$),
@@
 $=-\frac{1}{2T}\left(\sum_{t=1}^{T-1}v_{t}^{\gamma}(\theta)\right)^{2}\left(\frac{1}{N}\sum_{i=1}^{N}\partial_{\gamma_{T}\gamma_{T}\gamma_{T}}l_{i,T}(\theta,\phi)|_{\phi=\phi(\theta)}\right)+o_{\mathbb{P}}\left(1/(NT)\right)$
-$=O_{\mathbb{P}}\left(1/(NT)\right),$ (A.15)
+$=O_{\mathbb{P}}\left(1/(NT)\right),$ (E.15)
 where $v_{t}^{\gamma}(\theta)$ denotes the $t$-th element of $H_{\gamma\alpha^{\prime}}^{\ast}(\theta)\widetilde{s}_{\alpha}(\theta)+H_{\gamma\gamma^{\prime}}^{\ast}(\theta)\widetilde{s}_{\gamma}(\theta)$ for $t=1,\ldots,T-1$.
@@
 The next lemma states that the final remainder, $r(\theta)$ in Equation (2.7), is $O\left(1/\left(NT\right)\right)$ uniformly in $\theta$, where
-$r(\theta)$ $:=\frac{1}{2}(r^{s}(\theta,a))^{\prime}[H\left(\theta\right)]^{-1}r^{s}(\theta,a)+r^{l}(\theta,a_{0}),$ (A.16)
-$r^{l}\left(\theta,a_{0}\right)$ $:=\frac{1}{3}\left(\widehat{\psi}(\theta)-\psi(\theta)\right)^{\prime}r^{s}(\theta,a_{0}\times\iota_{(N+T-2)K})$ (A.17)
+$r(\theta)$ $:=\frac{1}{2}(r^{s}(\theta,a))^{\prime}[H\left(\theta\right)]^{-1}r^{s}(\theta,a)+r^{l}(\theta,a_{0}),$ (E.16)
+$r^{l}\left(\theta,a_{0}\right)$ $:=\frac{1}{3}\left(\widehat{\psi}(\theta)-\psi(\theta)\right)^{\prime}r^{s}(\theta,a_{0}\times\iota_{(N+T-2)K})$ (E.17)
 for some $a_{0}\in(0,1)$.
@@
 The dependency of $r(\theta)$ on $a$ and $a_{0}$ is suppress for simplicity here.
-Lemma A.3
-Under Assumption 1, $r(\theta)$ defined by (A.16) satisfies
+Lemma E.3
+Under Assumption 1, $r(\theta)$ defined by (E.16) satisfies
 $\sup_{\theta\in\Theta}\mathbb{E}[r(\theta)]=O\left(1/\left(NT\right)\right),$
@@
 Proof.
-In the proof, we retain the notation and setup from Lemmas A.1 and A.2.
-From Lemma A.2, we know that $NT\cdot r^{s}\left(\theta,a\right)$ is comparable to $(NT)^{3/4}\cdot s(\theta)$.
-Following similar arguments as in Lemma A.1, we can show that
+In the proof, we retain the notation and setup from Lemmas E.1 and E.2.
+From Lemma E.2, we know that $NT\cdot r^{s}\left(\theta,a\right)$ is comparable to $(NT)^{3/4}\cdot s(\theta)$.
+Following similar arguments as in Lemma E.1, we can show that
 $\mathbb{E}\left\{(r^{s}(\theta,a))^{\prime}[{H}\left(\theta\right)]^{-1}r^{s}(\theta,a)\right\}=O\left(1/(NT)\right).$
-Moreover, from (A.12), (A.14), and (A.14), we have
+Moreover, from (E.12), (E.14), and (E.14), we have
 $\quad\quad\left(\widehat{\psi}(\theta)-\psi(\theta)\right)^{\prime}r^{s}(\theta,a_{0}\,\iota_{N+T-2})$
@@
 $\quad\quad+\left(\sum_{i=1}^{N-1}v_{i}^{\alpha}(\theta)+\sum_{t=1}^{T-1}v_{t}^{\gamma}(\theta)\right)\cdot o_{\mathbb{P}}\left(1/(NT)\right)$
-$=O_{\mathbb{P}}\left(1/(NT)^{5/4}\right).$ (A.18)
-By Assumption 1 and applying the $L_{1}$ convergence theorem to A.18, we find that
+$=O_{\mathbb{P}}\left(1/(NT)^{5/4}\right).$ (E.18)
+By Assumption 1 and applying the $L_{1}$ convergence theorem to E.18, we find that
 $\sup_{\theta\in\Theta}\mathbb{E}\left[\left(\widehat{\psi}(\theta)-\psi(\theta)\right)^{\prime}r^{s}\left(\theta,a_{0}\,\iota_{N+T-2}\right)\right]=o\left(1/(NT)\right).$
@@
 $\mathbb{E[}\widehat{l}(\theta)-l(\theta)]=-\frac{1}{2}\mathbb{E\{}s^{\prime}(\theta)[H(\theta)]^{-1}s(\theta)\}+\mathbb{E}r(\theta)$
-for $r\left(\theta\right)$ defined in (A.16). The claim in Theorem 1 follows directly from Lemmas A.1 and A.3.
+for $r\left(\theta\right)$ defined in (E.16). The claim in Theorem 1 follows directly from Lemmas E.1 and E.3.
 The proof of Theorem 2 is as follows.
-Proof of Theorem 2. We retain the notation introduced in Lemmas A.1-A.3. We first show
-$\widehat{B}_{\alpha}(\theta)-B_{\alpha}(\theta)=o_{\mathbb{P}}(1/\sqrt{NT}),\qquad\widehat{B}_{\gamma}(\theta)-B_{\gamma}(\theta)=o_{\mathbb{P}}(1/\sqrt{NT}).$ (A.19)
-To show the first part of (A.19), note that
+Proof of Theorem 2. We retain the notation introduced in Lemmas E.1-E.3. We first show
+$\widehat{B}_{\alpha}(\theta)-B_{\alpha}(\theta)=o_{\mathbb{P}}(1/\sqrt{NT}),\qquad\widehat{B}_{\gamma}(\theta)-B_{\gamma}(\theta)=o_{\mathbb{P}}(1/\sqrt{NT}).$ (E.19)
+To show the first part of (E.19), note that
 $\widehat{B}_{\alpha}(\theta)-B_{\alpha}(\theta)=\frac{1}{2}\operatorname*{trace}\left\{\widehat{H}_{\alpha\alpha^{\prime}}^{\ast}\left(\theta\right)\left[S_{\alpha\alpha}\left(\theta\right)-\widehat{S}_{\alpha\alpha}\left(\theta\right)\right]\right\}+\frac{1}{2}\operatorname*{trace}\left\{\left[\widehat{H}_{\alpha\alpha^{\prime}}^{\ast}(\theta)-H_{\alpha\alpha^{\prime}}^{\ast}(\theta)\right]S_{\alpha\alpha}\left(\theta\right)\right\},$
@@
 $\widehat{S}_{\alpha\alpha}\left(\theta\right)$ $=\widehat{S}_{\alpha\alpha}^{1}\left(\theta\right)+\widehat{S}_{\alpha\alpha}^{2}\left(\theta\right),$
-$S_{\alpha\alpha}^{1}\left(\theta\right)$ $:=\operatorname*{diag}\left\{\widehat{s}_{1}^{\alpha}\left(\theta\right)\left(\widehat{s}_{1}^{\alpha}\left(\theta\right)\right)^{\prime},\ldots,\widehat{s}_{N-1}^{\alpha}\left(\theta\right)\left(\widehat{s}_{N-1}^{\alpha}\left(\theta\right)\right)^{\prime}\right\},$
-(${}_{\alpha\alpha}^{2}\left(\theta\right)$ $:=(\iota_{N-1}\iota_{N-1}^{\prime})\otimes\left[\widehat{s}_{N}^{\alpha}\left(\theta\right)\left(\widehat{s}_{N}^{\alpha}\left(\theta\right)\right)^{\prime}\right].$
+$\widehat{S}_{\alpha\alpha}^{1}\left(\theta\right)$ $:=\operatorname*{diag}\left\{\widehat{s}_{1}^{\alpha}\left(\theta\right)\left(\widehat{s}_{1}^{\alpha}\left(\theta\right)\right)^{\prime},\ldots,\widehat{s}_{N-1}^{\alpha}\left(\theta\right)\left(\widehat{s}_{N-1}^{\alpha}\left(\theta\right)\right)^{\prime}\right\},$
+$\widehat{S}_{\alpha\alpha}^{2}\left(\theta\right)$ $:=(\iota_{N-1}\iota_{N-1}^{\prime})\otimes\left[\widehat{s}_{N}^{\alpha}\left(\theta\right)\left(\widehat{s}_{N}^{\alpha}\left(\theta\right)\right)^{\prime}\right].$
 Here we have
@@
 $=\frac{1}{T^{2}}\sum_{t=1}^{T}\sum_{s=1}^{T}\mathds{1}\left\{\left|t-s\right|\leq\tau\right\}$
-×$\quad\quad\times\operatorname*{diag}\left\{\widetilde{s}_{1,t}^{\alpha}(\theta)\left(\widetilde{s}_{1,s}^{\alpha}(\theta)\right)^{\prime}-\widehat{s}_{1,t}^{\alpha}(\theta)\left(\widehat{s}_{1,s}^{\alpha}(\theta)\right)^{\prime},\ldots,\widetilde{s}_{N-1,t}^{\alpha}(\theta)\left(\widetilde{s}_{N-1,s}^{\alpha}(\theta)\right)^{\prime}-\widehat{s}_{N-1,t}^{\alpha}(\theta)\left(\widehat{s}_{N-1,s}^{\alpha}(\theta)\right)^{\prime}\right\}$
+$\quad\quad\times\operatorname*{diag}\left\{\widetilde{s}_{1,t}^{\alpha}(\theta)\left(\widetilde{s}_{1,s}^{\alpha}(\theta)\right)^{\prime}-\widehat{s}_{1,t}^{\alpha}(\theta)\left(\widehat{s}_{1,s}^{\alpha}(\theta)\right)^{\prime},\ldots,\widetilde{s}_{N-1,t}^{\alpha}(\theta)\left(\widetilde{s}_{N-1,s}^{\alpha}(\theta)\right)^{\prime}-\widehat{s}_{N-1,t}^{\alpha}(\theta)\left(\widehat{s}_{N-1,s}^{\alpha}(\theta)\right)^{\prime}\right\}$
 $\quad\quad+\frac{1}{T^{2}}\sum_{t=1}^{T}\sum_{s=1}^{T}\mathds{1}\left\{\left|t-s\right|>\tau\right\}$
-×$\quad\quad\times\operatorname*{diag}\left\{\mathbb{E}\left[\widetilde{s}_{1,t}^{\alpha}(\theta)\left(\widetilde{s}_{1,s}^{\alpha}(\theta)\right)^{\prime}\right],\ldots,\mathbb{E}\left[\widetilde{s}_{N-1,t}^{\alpha}(\theta)\left(\widetilde{s}_{N-1,s}^{\alpha}(\theta)\right)^{\prime}\right]\right\}.$ (A.20)
+$\quad\quad\times\operatorname*{diag}\left\{\mathbb{E}\left[\widetilde{s}_{1,t}^{\alpha}(\theta)\left(\widetilde{s}_{1,s}^{\alpha}(\theta)\right)^{\prime}\right],\ldots,\mathbb{E}\left[\widetilde{s}_{N-1,t}^{\alpha}(\theta)\left(\widetilde{s}_{N-1,s}^{\alpha}(\theta)\right)^{\prime}\right]\right\}.$ (E.20)
 Under Assumptions 1 and 1, by Davydov’s inequality,
@@
 $=o(1)$
-given that $\tau/T\rightarrow 0$. Combining this with (A.20), we have
-$\sup_{\theta\in\Theta}N^{2}\left\|S_{\alpha\alpha}^{1}\left(\theta\right)-\widehat{S}_{\alpha\alpha}^{1}\left(\theta\right)\right\|_{\max}=o_{\mathbb{P}}(1/T).$ (A.21)
-From the proof of Lemma A.2, we also have $\sup_{\theta\in\Theta}\|\widehat{\psi}(\theta)-\psi(\theta)\|_{\max}=o_{\mathbb{P}}(1)$.
+given that $\tau/T\rightarrow 0$. Combining this with (E.20), we have
+$\sup_{\theta\in\Theta}N^{2}\left\|S_{\alpha\alpha}^{1}\left(\theta\right)-\widehat{S}_{\alpha\alpha}^{1}\left(\theta\right)\right\|_{\max}=o_{\mathbb{P}}(1/T).$ (E.21)
+From the proof of Lemma E.2, we also have $\sup_{\theta\in\Theta}\|\widehat{\psi}(\theta)-\psi(\theta)\|_{\max}=o_{\mathbb{P}}(1)$.
 Moreover, under the conditional strong mixing assumption, each element of $T^{-1}\widehat{H}_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)$ converges in probability to the corresponding element in $T^{-1}H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)$ by the law of large numbers.
-Therefore (A.21) further leads to
+Therefore (E.21) further leads to
 $\quad\quad\operatorname*{trace}\left\{\widehat{H}_{\alpha\alpha^{\prime}}^{\ast}\left(\theta\right)\left[S_{\alpha\alpha}^{1}\left(\theta\right)-\widehat{S}_{\alpha\alpha}^{1}\left(\theta\right)\right]\right\}$
@@
 $=o_{\mathbb{P}}(1/N^{2}),$
-where the last equality follows from, as shown in Lemma A.2, $\sum_{j=1}^{N-1}h_{i,j}^{\ast\alpha}(\theta)=O(1/N)$ uniformly in both $i$ and $\theta$. The above leads to
-$\operatorname*{trace}\left\{\widehat{H}_{\alpha\alpha^{\prime}}^{\ast}(\theta)\left[S_{\alpha\alpha}\left(\theta\right)-\widehat{S}_{\alpha\alpha}\left(\theta\right)\right]\right\}=o_{\mathbb{P}}(1/\sqrt{NT}).$ (A.22)
+where the last equality follows from, as shown in Lemma E.2, $\sum_{j=1}^{N-1}h_{i,j}^{\ast\alpha}(\theta)=O(1/N)$ uniformly in both $i$ and $\theta$. The above leads to
+$\operatorname*{trace}\left\{\widehat{H}_{\alpha\alpha^{\prime}}^{\ast}(\theta)\left[S_{\alpha\alpha}\left(\theta\right)-\widehat{S}_{\alpha\alpha}\left(\theta\right)\right]\right\}=o_{\mathbb{P}}(1/\sqrt{NT}).$ (E.22)
 Second, from Theorem 1, we also have
-$\operatorname*{trace}\left\{\left[\widehat{H}_{\alpha\alpha^{\prime}}^{\ast}(\theta)-H_{\alpha\alpha^{\prime}}^{\ast}(\theta)\right]S_{\alpha\alpha}\left(\theta\right)\right\}=o_{\mathbb{P}}(1)\cdot\operatorname*{trace}\left\{H_{\alpha\alpha^{\prime}}^{\ast}(\theta)S_{\alpha\alpha}\left(\theta\right)\right\}=o_{\mathbb{P}}(1)\cdot B_{\alpha}(\theta)=o_{\mathbb{P}}(1/\sqrt{NT}).$ (A.23)
-Combining A.22 and A.23 leads to the first part of A.19. The second part of A.19 can be shown similarly.
+$\operatorname*{trace}\left\{\left[\widehat{H}_{\alpha\alpha^{\prime}}^{\ast}(\theta)-H_{\alpha\alpha^{\prime}}^{\ast}(\theta)\right]S_{\alpha\alpha}\left(\theta\right)\right\}=o_{\mathbb{P}}(1)\cdot\operatorname*{trace}\left\{H_{\alpha\alpha^{\prime}}^{\ast}(\theta)S_{\alpha\alpha}\left(\theta\right)\right\}=o_{\mathbb{P}}(1)\cdot B_{\alpha}(\theta)=o_{\mathbb{P}}(1/\sqrt{NT}).$ (E.23)
+Combining E.22 and E.23 leads to the first part of E.19. The second part of E.19 can be shown similarly.
 Noting that
@@
 $\quad\quad+o_{\mathbb{P}}(1/\sqrt{NT}).$
-As from Lemma A.1, we have
+As from Lemma E.1, we have
 $\frac{1}{2}\widetilde{s}_{\alpha}^{\prime}(\theta)H_{\alpha\alpha^{\prime}}^{\ast}(\theta)\widetilde{s}_{\alpha}(\theta)-B_{\alpha}(\theta)=o_{\mathbb{P}}(1/\sqrt{NT}),\qquad\frac{1}{2}\widetilde{s}_{\gamma}^{\prime}(\theta)H_{\gamma\gamma^{\prime}}^{\ast}(\theta)\widetilde{s}_{\gamma}(\theta)-B_{\gamma}(\theta)=o_{\mathbb{P}}(1/\sqrt{NT}).$
-The conclusion follows by combining this and (A.19).
+The conclusion follows by combining this and (E.19).
 We continue to establishing the asymptotitc properties of $\{\widehat{\theta}_{L},\widehat{\xi}_{LR},\widehat{\xi}_{LM},\widehat{\xi}_{Wald}\}$.
@@
 where each component of $\overline{\theta}$ lies between the corresponding components of $\widehat{\theta}_{L}$ and $\theta_{0}$. From Theorem 2, we further have
-$0=\triangledown_{\theta}l(\theta_{0})+\triangledown_{\theta\theta^{\prime}}l(\overline{\theta})(\widehat{\theta}_{L}-\theta_{0})+o_{\mathbb{P}}(1/\sqrt{NT})$ (A.24)
+$0=\triangledown_{\theta}l(\theta_{0})+\triangledown_{\theta\theta^{\prime}}l(\overline{\theta})(\widehat{\theta}_{L}-\theta_{0})+o_{\mathbb{P}}(1/\sqrt{NT})$ (E.24)
 Under Assumption 2, $\theta_{0}$ is the unique maximizer of $\mathbb{E}[l(\theta)]$. Since $L(\theta)-\mathbb{E}[l(\theta)]=o_{\mathbb{P}}(1)$, under Assumption 2, we conclude that
@@
 $\triangledown_{\theta\theta^{\prime}}l(\overline{\theta})-\mathbb{E}[\triangledown_{\theta\theta^{\prime}}l(\theta_{0})]=o_{\mathbb{P}}(1).$
-Thus, using Assumption 2 and Slutsky’s theorem, we obtain from (A.24) that
-$\sqrt{NT}(\widehat{\theta}_{L}-\theta_{0})=-\left\{\mathbb{E}[\triangledown_{\theta\theta^{\prime}}l(\theta_{0})]\right\}^{-1}\sqrt{NT}\triangledown_{\theta}l(\theta_{0})+o_{\mathbb{P}}(1).$ (A.25)
-Given the conditional cross-sectional independence and the conditional strong mixing condition (Assumption 1), (A.25) leads to
-$\left[\Sigma^{-1}\Omega\Sigma^{-1}\right]^{-1/2}\sqrt{NT}(\widehat{\theta}_{L}-\theta_{0})\overset{\mathbb{D}}{\longrightarrow}\mathcal{N}\left(0,\mathbb{I}_{K}\right)$ (A.26)
+Thus, using Assumption 2 and Slutsky’s theorem, we obtain from (E.24) that
+$\sqrt{NT}(\widehat{\theta}_{L}-\theta_{0})=-\left\{\mathbb{E}[\triangledown_{\theta\theta^{\prime}}l(\theta_{0})]\right\}^{-1}\sqrt{NT}\triangledown_{\theta}l(\theta_{0})+o_{\mathbb{P}}(1).$ (E.25)
+Given the conditional cross-sectional independence and the conditional strong mixing condition (Assumption 1), (E.25) leads to
+$\left[\Sigma^{-1}\Omega\Sigma^{-1}\right]^{-1/2}\sqrt{NT}(\widehat{\theta}_{L}-\theta_{0})\overset{\mathbb{D}}{\longrightarrow}\mathcal{N}\left(0,\mathbb{I}_{K}\right)$ (E.26)
 as $N,T\rightarrow\infty$, where $\mathbb{E}[\triangledown_{\theta}l(\theta_{0})]=0$ and
 $\Sigma:=-\mathbb{E}\left[\triangledown_{\theta\theta^{\prime}}l(\theta_{0})\right],\qquad\Omega:=\left(NT\right)\mathbb{E}[\triangledown_{\theta}l(\theta_{0})\triangledown_{\theta^{\prime}}l(\theta_{0})].$
-Since we consider a likelihood model, $\Sigma^{-1}\Omega\Sigma^{-1}$ in (A.26) simplifies to $\Sigma^{-1}$ under the information matrix equality ($\Sigma=\Omega$).
+Since we consider a likelihood model, $\Sigma^{-1}\Omega\Sigma^{-1}$ in (E.26) simplifies to $\Sigma^{-1}$ under the information matrix equality ($\Sigma=\Omega$).
 To show the asymptotic distribution of $\widehat{\xi}_{LR},\widehat{\xi}_{LM},\widehat{\xi}_{Wald}$, we find the following lemma useful.
-Lemma A.4
+Lemma E.4
 Under Assumptions 1 and 2, and under the null hypothesis $H_{0}:R(\theta_{0})=0$, as $N$ and $T\rightarrow\infty,$
@@
 where $\lambda$ is the Lagrangian multiplier. The first-order conditions for this constrained maximization are
-$\sqrt{NT}\triangledown_{\theta}L(\widehat{\theta}_{R})+[J(\widehat{\theta}_{R})]^{\prime}\sqrt{NT}\widehat{\lambda}$ $=0,$ (A.27)
-$\sqrt{NT}R(\widehat{\theta}_{R})$ $=0.$ (A.28)
+$\sqrt{NT}\triangledown_{\theta}L(\widehat{\theta}_{R})+[J(\widehat{\theta}_{R})]^{\prime}\sqrt{NT}\widehat{\lambda}$ $=0,$ (E.27)
+$\sqrt{NT}R(\widehat{\theta}_{R})$ $=0.$ (E.28)
 Following a similar proof as in Corollary 1, we have $\sqrt{NT}(\widehat{\theta}_{R}-\theta_{0})=O_{\mathbb{P}}\left(1\right)$.
-Furthermore, $\sqrt{NT}\widehat{\lambda}=O_{\mathbb{P}}(1)$ as implied by (A.27) since $\sqrt{NT}\triangledown_{\theta}L(\theta_{0})=O_{\mathbb{P}}(1)$ and $J(\theta_{0})$ is of full row rank.
+Furthermore, $\sqrt{NT}\widehat{\lambda}=O_{\mathbb{P}}(1)$ as implied by (E.27) since $\sqrt{NT}\triangledown_{\theta}L(\theta_{0})=O_{\mathbb{P}}(1)$ and $J(\theta_{0})$ is of full row rank.
 By Theorem 2, a mean value expansion of $\triangledown_{\theta}L(\widehat{\theta}_{R})$ around $\theta_{0}$ is
@@
 $=\triangledown_{\theta}l(\theta_{0})+\mathbb{E}[\triangledown_{\theta\theta^{\prime}}l(\theta_{0})](\widehat{\theta}_{R}-\theta_{0})+o_{\mathbb{P}}(1/\sqrt{NT}),$
-where each component of $\overline{\theta}$ lies between the corresponding components of $\widehat{\theta}_{R}$ and $\theta_{0}.$ Combining this with (A.27) yields that
-$\mathbb{E}[\triangledown_{\theta\theta^{\prime}}l(\theta_{0})]\sqrt{NT}(\widehat{\theta}_{R}-\theta_{0})+[J(\theta_{0})]^{\prime}\sqrt{NT}\widehat{\lambda}=-\sqrt{NT}\triangledown_{\theta}l(\theta_{0})+o_{\mathbb{P}}(1).$ (A.29)
-Under $H_{0}$ and from (A.28), we also obtain
-00 0 $=\sqrt{NT}R(\widehat{\theta}_{R})$
+where each component of $\overline{\theta}$ lies between the corresponding components of $\widehat{\theta}_{R}$ and $\theta_{0}.$ Combining this with (E.27) yields that
+$\mathbb{E}[\triangledown_{\theta\theta^{\prime}}l(\theta_{0})]\sqrt{NT}(\widehat{\theta}_{R}-\theta_{0})+[J(\theta_{0})]^{\prime}\sqrt{NT}\widehat{\lambda}=-\sqrt{NT}\triangledown_{\theta}l(\theta_{0})+o_{\mathbb{P}}(1).$ (E.29)
+Under $H_{0}$ and from (E.28), we also obtain
+$0$ $=\sqrt{NT}R(\widehat{\theta}_{R})$
 $=\sqrt{NT}R(\theta_{0})+J(\theta_{0})\sqrt{NT}(\widehat{\theta}_{R}-\theta_{0})+o_{\mathbb{P}}(1)$
-$=J(\theta_{0})\sqrt{NT}(\widehat{\theta}_{R}-\theta_{0})+o_{\mathbb{P}}(1).$ (A.30)
-Now, combining (A.29) and (A.30), we have
+$=J(\theta_{0})\sqrt{NT}(\widehat{\theta}_{R}-\theta_{0})+o_{\mathbb{P}}(1).$ (E.30)
+Now, combining (E.29) and (E.30), we have
 $\begin{pmatrix}\mathbb{E}[\triangledown_{\theta\theta^{\prime}}l(\theta_{0})]&\left[J(\theta_{0})\right]^{\prime}\\ J(\theta_{0})&0\end{pmatrix}\begin{pmatrix}\sqrt{NT}(\widehat{\theta}_{R}-\theta_{0})\\ \sqrt{NT}\widehat{\lambda}\end{pmatrix}=\begin{pmatrix}-\sqrt{NT}\triangledown_{\theta}l(\theta_{0})\\ 0\end{pmatrix}+o_{\mathbb{P}}(1)$
 which has the solution
-$\sqrt{NT}(\widehat{\theta}_{R}-\theta_{0})=-\left(\{\mathbb{E}[\triangledown_{\theta\theta^{\prime}}l(\theta_{0})]\}^{-1}-\Xi\left(\theta_{0}\right)\right)\sqrt{NT}\triangledown_{\theta}l(\theta_{0})+o_{\mathbb{P}}\left(1\right).$ (A.31)
-The conclusion follows by subtracting (A.25) in Corollary 1 from both sides of (A.31).
-Finally, we proof the asymptotic distribution of $\widehat{\xi}_{LR},\widehat{\xi}_{LM},\widehat{\xi}_{Wald}$ below.
+$\sqrt{NT}(\widehat{\theta}_{R}-\theta_{0})=-\left(\{\mathbb{E}[\triangledown_{\theta\theta^{\prime}}l(\theta_{0})]\}^{-1}-\Xi\left(\theta_{0}\right)\right)\sqrt{NT}\triangledown_{\theta}l(\theta_{0})+o_{\mathbb{P}}\left(1\right).$ (E.31)
+The conclusion follows by subtracting (E.25) in Corollary 1 from both sides of (E.31).
+Below, we proof the asymptotic distribution of $\widehat{\xi}_{LR},\widehat{\xi}_{LM},\widehat{\xi}_{Wald}$ below.
 Proof of Corollary 2.
-Let $\widehat{\xi}:=-NT\triangledown_{\theta^{\prime}}l(\theta_{0})\Xi\left(\theta_{0}\right)\triangledown_{\theta}l(\theta_{0})$ for $\Xi\left(\theta_{0}\right)$ defined in Lemma A.4.
+Let $\widehat{\xi}:=-NT\triangledown_{\theta^{\prime}}l(\theta_{0})\Xi\left(\theta_{0}\right)\triangledown_{\theta}l(\theta_{0})$ for $\Xi\left(\theta_{0}\right)$ defined in Lemma E.4.
 We first show that
-$\widehat{\xi}_{LR}=\widehat{\xi}+o_{\mathbb{P}}(1),\qquad\widehat{\xi}_{LM}=\widehat{\xi}+o_{\mathbb{P}}(1),\qquad\widehat{\xi}_{Wald}=\widehat{\xi}+o_{\mathbb{P}}(1).$ (A.32)
+$\widehat{\xi}_{LR}=\widehat{\xi}+o_{\mathbb{P}}(1),\qquad\widehat{\xi}_{LM}=\widehat{\xi}+o_{\mathbb{P}}(1),\qquad\widehat{\xi}_{Wald}=\widehat{\xi}+o_{\mathbb{P}}(1).$ (E.32)
 A mean value expansion of $L(\widehat{\theta}_{R})$ around $\widehat{\theta}_{L}$ gives
@@
 $\widehat{\xi}_{LR}=-NT(\widehat{\theta}_{R}-\widehat{\theta}_{L})^{\prime}\triangledown_{\theta\theta^{\prime}}L(\dot{\theta})(\widehat{\theta}_{R}-\widehat{\theta}_{L}).$
-Since $\triangledown_{\theta\theta^{\prime}}L(\dot{\theta})-\mathbb{E}[\triangledown_{\theta\theta^{\prime}}L(\theta_{0})]=o_{\mathbb{P}}(1)$, by Lemma A.4, we further conclude that
+Since $\triangledown_{\theta\theta^{\prime}}L(\dot{\theta})-\mathbb{E}[\triangledown_{\theta\theta^{\prime}}L(\theta_{0})]=o_{\mathbb{P}}(1)$, by Lemma E.4, we further conclude that
 $\widehat{\xi}_{LR}=\widehat{\xi}+o_{\mathbb{P}}(1).$
@@
 $=\mathbb{E}[\triangledown_{\theta\theta^{\prime}}L(\theta_{0})](\widehat{\theta}_{R}-\widehat{\theta}_{L})+o_{\mathbb{P}}(1/\sqrt{NT}),$
-[…]
+where each component of $\ddot{\theta}$ lies between the corresponding components of $\widehat{\theta}_{R}$ and $\widehat{\theta}_{L}$. Thus,
 $\widehat{\xi}_{LM}$ $=-NT(\widehat{\theta}_{R}-\widehat{\theta}_{L})^{\prime}\mathbb{E}[\triangledown_{\theta\theta^{\prime}}L(\theta_{0})][\triangledown_{\theta\theta^{\prime}}L(\widehat{\theta}_{R})]^{-1}\mathbb{E}[\triangledown_{\theta\theta^{\prime}}L(\theta_{0})](\widehat{\theta}_{R}-\widehat{\theta}_{L})+o_{\mathbb{P}}\left(1\right)$
@@
 The desired results follows from here.
-B Extra Simulation Results
-B.1 Logit Model
-Table B.1: Empirical Size of LM and Wald Tests for Logit Model, Design 1
-(a) LM Tests
-Static Model Dynamic Model
-$(N,T)$ Uncorrected Corrected Uncorrected Corrected
-$(30,30)$ $15\%$ $9\%$ $18\%$ $10\%$
-$(60,60)$ $15\%$ $5\%$ $17\%$ $7\%$
-$(90,90)$ $18\%$ $6\%$ $16\%$ $5\%$
-(b) Wald Tests
-Static Model Dynamic Model
-$(N,T)$ Uncorrected Corrected Uncorrected Corrected
-$(30,30)$ $15\%$ $9\%$ $18\%$ $10\%$
-$(60,60)$ $15\%$ $5\%$ $17\%$ $7\%$
-$(90,90)$ $18\%$ $6\%$ $16\%$ $5\%$
-[…]
-Notes: Logit model with $\varepsilon_{it}$ standard logistic.
-Design 1: $\{\alpha_{k,i}^{0},\gamma_{k,t}^{0}\}\sim\mathcal{N}(0,0.01)$, $X_{it}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0}+\gamma_{1,t}^{0}+\gamma_{2,t}^{0},0.25)$, and $X_{i0}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0},0.25)$.
-The null hypothesis $H_{0}:(\beta_{0},\rho_{0})^{\prime}=(0.5,0.5)^{\prime}$ for the dynamic model, and $H_{0}:\beta_{0}=0.5$ for the static model, with a significance level $5\%$.
-The number of replications is $1000$.
-Table B.2: Empirical Power of LM and Wald Tests for Logit Model using $L(\theta)$, Design 1
-(a) LM Tests
-$\delta$
-$(N,T)$ $-0.20$ $-0.15$ $-0.10$ $0.10$ $0.15$ $0.20$
-Static Model
-$(30,30)$ $44\%$ $32\%$ $22\%$ $10\%$ $14\%$ $21\%$
-$(60,60)$ $86\%$ $62\%$ $37\%$ $25\%$ $50\%$ $74\%$
-$(90,90)$ $100\%$ $93\%$ $65\%$ $50\%$ $84\%$ $98\%$
-Dynamic Model
-$(30,30)$ $48\%$ $32\%$ $20\%$ $18\%$ $28\%$ $43\%$
-$(60,60)$ $94\%$ $73\%$ $40\%$ $45\%$ $75\%$ $94\%$
-$(90,90)$ $100\%$ $98\%$ $76\%$ $78\%$ $98\%$ $100\%$
-(b) Wald Tests
-$\delta$
-$(N,T)$ $-0.20$ $-0.15$ $-0.10$ $0.10$ $0.15$ $0.20$
-Static Model
-$(30,30)$ $44\%$ $32\%$ $22\%$ $10\%$ $14\%$ $21\%$
-$(60,60)$ $86\%$ $62\%$ $37\%$ $25\%$ $50\%$ $74\%$
-$(90,90)$ $100\%$ $93\%$ $65\%$ $50\%$ $84\%$ $98\%$
-Dynamic Model
-$(30,30)$ $47\%$ $31\%$ $20\%$ $18\%$ $28\%$ $43\%$
-$(60,60)$ $94\%$ $73\%$ $39\%$ $45\%$ $75\%$ $94\%$
-$(90,90)$ $100\%$ $98\%$ $76\%$ $78\%$ $98\%$ $100\%$
-[…]
-Notes: Logit model with $\varepsilon_{it}$ standard logistic.
-Design 1: $\{\alpha_{k,i}^{0},\gamma_{k,t}^{0}\}\sim\mathcal{N}(0,0.01)$, $X_{it}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0}+\gamma_{1,t}^{0}+\gamma_{2,t}^{0},0.25)$, and $X_{i0}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0},0.25)$.
-The null hypothesis $H_{0}:(\beta_{0},\rho_{0})^{\prime}=(0.5+\delta,0.5+\delta)^{\prime}$ for the dynamic model, and $H_{0}:\beta_{0}=0.5+\delta$ for the static model, with a significance level $5\%,$ where $\delta\in\{\pm 0.2,\pm 0.15,\pm 0.1\}$.
-The number of replications is $1000$.
-Table B.3: Comparisons of Uncorrected and Corrected Logit Estimates, Design 2
+F Additional Simulation Results
+F.1 Additional Results under Design 1
+Table F.1: Comparisons of Uncorrected and Corrected Probit Estimates, Design 1
 $(N,T)$ $(30,30)$ $(60,60)$ $(90,90)$
 Bias RMSE Bias RMSE Bias RMSE
-Static Model
-$\widehat{\beta}$ $0.454$ $1.201$ $0.100$ $0.132$ $0.060$ $0.071$
-$\widehat{\beta}_{L}$ $0.101$ $0.199$ $0.020$ $0.056$ $0.010$ $0.035$
-Dynamic Model
-$\widehat{\beta}$ $0.527$ $1.583$ $0.109$ $0.212$ $0.059$ $0.071$
-$\widehat{\beta}_{L}^{(1)}$ $0.117$ $0.204$ $0.024$ $0.058$ $0.009$ $0.038$
-$\widehat{\beta}_{L}^{(2)}$ $0.126$ $0.278$ $0.025$ $0.059$ $0.009$ $0.041$
-$\widehat{\rho}$ $-0.110$ $0.241$ $-0.055$ $0.110$ $-0.037$ $0.071$
-$\widehat{\rho}_{L}^{(1)}$ $-0.041$ $0.192$ $-0.017$ $0.090$ $-0.010$ $0.059$
-$\widehat{\rho}_{L}^{(2)}$ $-0.035$ $0.196$ $-0.012$ $0.091$ $-0.007$ $0.059$
+Static Probit Model
+$\widehat{\beta}$ $0.175$ $0.208$ $0.069$ $0.076$ $0.044$ $0.048$
+$\widehat{\beta}_{L}$ $0.067$ $0.101$ $0.016$ $0.034$ $0.008$ $0.019$
+$\widehat{\beta}_{J}$ 1.966-1.966 $3.942$ 0.040-0.040 $0.101$ 0.010-0.010 $0.022$
+$\widehat{\beta}_{B}$ 0.114-0.114 $0.153$ 0.013-0.013 $0.030$ 0.005-0.005 $0.018$
+Dynamic Probit Model
+$\widehat{\beta}$ $0.209$ $0.335$ $0.072$ $0.080$ $0.045$ $0.049$
+$\widehat{\beta}_{L}^{(1)}$ $0.079$ $0.116$ $0.019$ $0.036$ $0.008$ $0.021$
+$\widehat{\beta}_{L}^{(2)}$ $0.082$ $0.118$ $0.020$ $0.037$ $0.009$ $0.021$
+$\widehat{\beta}_{J}$ 2.161-2.161 $3.956$ 0.077-0.077 $0.218$ 0.013-0.013 $0.025$
+$\widehat{\beta}_{B}$ 0.155-0.155 $0.482$ 0.015-0.015 $0.032$ 0.006-0.006 $0.019$
+$\widehat{\rho}$ 0.044-0.044 $0.126$ 0.021-0.021 $0.059$ 0.015-0.015 $0.036$
+$\widehat{\rho}_{L}^{(1)}$ 0.024-0.024 $0.109$ 0.009-0.009 $0.053$ 0.007-0.007 $0.032$
+$\widehat{\rho}_{L}^{(2)}$ 0.016-0.016 $0.109$ 0.004-0.004 $0.052$ 0.003-0.003 $0.032$
+$\widehat{\rho}_{J}$ 0.008-0.008 $0.116$ $0.001$ $0.055$ $0.000$ $0.032$
+$\widehat{\rho}_{B}$ 0.004-0.004 $0.100$ $0.000$ $0.052$ 0.001-0.001 $0.031$
 […]
-Notes: Logit model with $\varepsilon_{it}$ standard logistic.
-Design 2: $\{\alpha_{k,i}^{0},\gamma_{k,t}^{0}\}\sim\mathcal{N}(0,1/4)$, $X_{it}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0}+\gamma_{1,t}^{0}+\gamma_{2,t}^{0},1)$, and $X_{i0}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0},1)$.
-For the static model, $\widehat{\beta}$ and $\widehat{\beta}_{L}$ are, respectively, the uncorrected and corrected estimates of $\beta_{0}=0.5$.
-For the dynamic model $(\widehat{\beta},\widehat{\rho})^{\prime}$ is the uncorrected estimate of $(\rho_{0},\beta_{0})^{\prime}=(0.5,0.5)^{\prime}$ from and $(\widehat{\beta}_{L}^{(\tau)},\widehat{\rho}_{L}^{(\tau)})^{\prime}$ is the corrected estimate with truncation parameter $\tau\in(1,2)$.
-The number of replications is $1000$.
-Table B.4: Empirical Size of LR, LM, and Wald Tests for Logit Model, Design 2
-(a) LR Tests
-Static Model Dynamic Model
-$(N,T)$ Uncorrected Corrected Uncorrected Corrected
-$(30,30)$ $47\%$ $20\%$ $43\%$ $18\%$
-$(60,60)$ $43\%$ $10\%$ $38\%$ $9\%$
-$(90,90)$ $42\%$ $9\%$ $34\%$ $7\%$
-(b) LM Tests
-Static Model Dynamic Model
-$(N,T)$ Uncorrected Corrected Uncorrected Corrected
-$(30,30)$ $42\%$ $18\%$ $38\%$ $16\%$
-$(60,60)$ $42\%$ $9\%$ $37\%$ $9\%$
-$(90,90)$ $41\%$ $9\%$ $34\%$ $7\%$
-(c) Wald Tests
-Static Model Dynamic Model
-$(N,T)$ Uncorrected Corrected Uncorrected Corrected
-$(30,30)$ $32\%$ $15\%$ $27\%$ $15\%$
-$(60,60)$ $41\%$ $9\%$ $36\%$ $8\%$
-$(90,90)$ $42\%$ $9\%$ $34\%$ $7\%$
+Table F.2: Empirical Size and Power from LR Test for Probit Model, Design 1
+(a) Size
+Size
+$(N,T)$ Uncorrected Analytical Bootstrap
+Static Probit Model
+$(30,30)$ $66$ $26$ 22
+$(60,60)$ $64$ $13$ 22
+$(90,90)$ $66$ 88 33
+Dynamic Probit Model
+$(30,30)$ $61$ $25$ 22
+$(60,60)$ $60$ $12$ 33
+$(90,90)$ $57$ 99 22
+(b) Power
+Power at $\delta$ (Analytical) Power at $\delta$ (Bootstrap)
+$(N,T)$ 0.2-0.2 0.1-0.1 $0.1$ $0.2$ 0.2-0.2 0.1-0.1 $0.1$ $0.2$
+Static Probit Model
+$(30,30)$ $99$ $80$ $16$ $58$ $85$ $24$ 0 0
+$(60,60)$ $100$ $99$ $83$ $100$ $100$ $96$ 0 $56$
+$(90,90)$ $100$ $100$ $100$ $100$ $100$ $100$ $15$ $100$
+Dynamic Probit Model
+$(30,30)$ $99$ $74$ $28$ $73$ $78$ $18$ 11 22
+$(60,60)$ $100$ $99$ $88$ $100$ $100$ $94$ 44 $86$
+$(90,90)$ $100$ $100$ $100$ $100$ $100$ $100$ $50$ $100$
 […]
-Notes: Logit model with $\varepsilon_{it}$ standard logistic.
-Design 2: $\{\alpha_{k,i}^{0},\gamma_{k,t}^{0}\}\sim\mathcal{N}(0,1/4)$, $X_{it}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0}+\gamma_{1,t}^{0}+\gamma_{2,t}^{0},1)$, and $X_{i0}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0},1)$.
-The null hypothesis $H_{0}:(\beta_{0},\rho_{0})^{\prime}=(0.5,0.5)^{\prime}$ for the dynamic model, and $H_{0}:\beta_{0}=0.5$ for the static model, with a significance level $5\%$.
-The number of replications is $1000$.
-Table B.5: Empirical Power of LR, LM, and Wald Tests for Logit Model using $L(\theta)$, Design 2
-(a) LR Tests
-$\delta$
-$(N,T)$ $-0.20$ $-0.15$ $-0.10$ $0.10$ $0.15$ $0.20$
-Static Model
-$(30,30)$ $79\%$ $62\%$ $45\%$ $12\%$ $16\%$ $26\%$
-$(60,60)$ $100\%$ $95\%$ $73\%$ $40\%$ $75\%$ $93\%$
-$(90,90)$ $100\%$ $100\%$ $94\%$ $81\%$ $98\%$ $100\%$
-Dynamic Model
-$(30,30)$ $75\%$ $58\%$ $41\%$ $19\%$ $26\%$ $35\%$
-$(60,60)$ $100\%$ $94\%$ $69\%$ $44\%$ $80\%$ $96\%$
-$(90,90)$ $100\%$ $100\%$ $93\%$ $86\%$ $99\%$ $100\%$
-(b) LM Tests
-$\delta$
-$(N,T)$ $-0.20$ $-0.15$ $-0.10$ $0.10$ $0.15$ $0.20$
-Static Model
-$(30,30)$ $77\%$ $60\%$ $42\%$ $12\%$ $17\%$ $28\%$
-$(60,60)$ $99\%$ $95\%$ $72\%$ $41\%$ $77\%$ $93\%$
-$(90,90)$ $100\%$ $100\%$ $94\%$ $82\%$ $98\%$ $100\%$
-Dynamic Model
-$(30,30)$ $73\%$ $56\%$ $38\%$ $19\%$ $27\%$ $36\%$
-$(60,60)$ $100\%$ $93\%$ $68\%$ $45\%$ $80\%$ $94\%$
-$(90,90)$ $100\%$ $100\%$ $93\%$ $86\%$ $99\%$ $99\%$
-(c) Wald Tests
-$\delta$
-$(N,T)$ $-0.20$ $-0.15$ $-0.10$ $0.10$ $0.15$ $0.20$
-Static Model
-$(30,30)$ $75\%$ $57\%$ $39\%$ $10\%$ $17\%$ $29\%$
-$(60,60)$ $100\%$ $95\%$ $72\%$ $41\%$ $77\%$ $95\%$
-$(90,90)$ $100\%$ $100\%$ $94\%$ $81\%$ $98\%$ $100\%$
-Dynamic Model
-$(30,30)$ $69\%$ $50\%$ $34\%$ $18\%$ $26\%$ $36\%$
-$(60,60)$ $100\%$ $93\%$ $68\%$ $45\%$ $81\%$ $97\%$
-$(90,90)$ $100\%$ $100\%$ $92\%$ $86\%$ $99\%$ $100\%$
+Notes: See also Table F.5 where we show that the sizes for the probit model with $\left(N,T\right)=\left(120,120\right)$ and $\left(150,150\right)$ are close to $5\%$.
+Table F.3: Empirical Size and Power from LM Test for Logit and Probit Model, Design 1
+Size Power at $\delta$
+$(N,T)$ Uncorrected Analytical 0.2-0.2 0.1-0.1 $0.1$ $0.2$
+Static Logit Model
+$(30,30)$ $37$ $13$ $87$ $47$ $13$ $46$
+$(60,60)$ $40$ 88 $100$ $85$ $59$ $99$
+$(90,90)$ $38$ 55 $100$ $98$ $95$ $100$
+Static Probit Model
+$(30,30)$ $64$ $25$ $99$ $79$ $17$ $60$
+$(60,60)$ $63$ $13$ $100$ $99$ $84$ $100$
+$(90,90)$ $66$ 88 $100$ $100$ $100$ $100$
+Dynamic Logit Model
+$(30,30)$ $35$ $13$ $83$ $41$ 2020 $56$
+$(60,60)$ $35$ 99 $100$ $79$ $64$ $99$
+$(90,90)$ $33$ 66 $100$ $98$ $95$ $100$
+Dynamic Probit Model
+$(30,30)$ $58$ $22$ $99$ $73$ $28$ $75$
+$(60,60)$ $58$ $12$ $100$ $99$ $89$ $99$
+$(90,90)$ $56$ 99 $100$ $100$ $100$ $99$
 […]
-Notes: Logit model with $\varepsilon_{it}$ standard logistic.
-Design 2: $\{\alpha_{k,i}^{0},\gamma_{k,t}^{0}\}\sim\mathcal{N}(0,1/4)$, $X_{it}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0}+\gamma_{1,t}^{0}+\gamma_{2,t}^{0},1)$, and $X_{i0}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0},1)$.
-The null hypothesis $H_{0}:(\beta_{0},\rho_{0})^{\prime}=(0.5+\delta,0.5+\delta)^{\prime}$ for the dynamic model, and $H_{0}:\beta_{0}=0.5+\delta$ for the static model, with a significance level $5\%,$ where $\delta\in\{\pm 0.2,\pm 0.15,\pm 0.1\}$.
-The number of replications is $1000$.
-B.2 Probit Model
-Table B.6: Comparisons of Uncorrected and Corrected Probit Estimates, Design 1
+Notes: See also Table F.5 where we show that the sizes for the probit model with $\left(N,T\right)=\left(120,120\right)$ and $\left(150,150\right)$ are close to $5\%$.
+Table F.4: Empirical Size and Power from Wald Test for Logit and Probit Model, Design 1
+Size Power at $\delta$
+$(N,T)$ Uncorrected Analytical 0.2-0.2 0.1-0.1 $0.1$ $0.2$
+Static Logit Model
+$(30,30)$ $37$ $13$ $87$ $46$ $12$ $45$
+$(60,60)$ $40$ 88 $100$ $85$ $59$ $99$
+$(90,90)$ $38$ 55 $100$ $98$ $95$ $100$
+Static Probit Model
+$(30,30)$ $63$ $23$ $98$ $78$ $17$ $59$
+$(60,60)$ $63$ $13$ $100$ $99$ $84$ $100$
+$(90,90)$ $66$ 88 $100$ $100$ $100$ $100$
+Dynamic Logit Model
+$(30,30)$ $34$ $12$ $82$ $40$ 2020 $56$
+$(60,60)$ $35$ 99 $100$ $79$ $64$ $99$
+$(90,90)$ $32$ 66 $100$ $98$ $95$ $100$
+Dynamic Probit Model
+$(30,30)$ $55$ 2121 $98$ $72$ $29$ $75$
+$(60,60)$ $58$ $12$ $99$ $98$ $89$ $99$
+$(90,90)$ $56$ 99 $100$ $100$ $100$ $100$
+[…]
+Notes: See also Table F.5 where we show that the sizes for the probit model with $\left(N,T\right)=\left(120,120\right)$ and $\left(150,150\right)$ are close to $5\%$.
+Table F.5: Empirical Size from LR, LM, and Wald Test for Probit Model, Design 1 with $N,T=(120,120)$ and $(150,150)$
+LR Test LM Test Wald Test
+$(N,T)$ Uncorrected Analytical Uncorrected Analytical Uncorrected Analytical
+Static Probit Model
+$(120,120)$ $63$ 66 $63$ 66 $63$ 66
+$(150,150)$ $64$ 55 $63$ 55 $63$ 55
+Dynamic Probit Model
+$(120,120)$ $59$ 88 $59$ 88 $59$ 88
+$(150,150)$ $56$ 77 $56$ 77 $56$ 77
+[…]
+Table F.6: Comparisons of Uncorrected and Corrected Logit and Probit Estimates, Design 1 with $N=90$ and $T=10$
+Probit Logit Probit Logit
+Bias RMSE Bias RMSE Bias RMSE Bias RMSE
+Static Model, $\theta_{0}=\beta_{0}=0$ Static Model, $\theta_{0}=\beta_{0}=0.5$
+$\widehat{\beta}$ $0.056$ $2.821$ $0.074$ $1.029$ $4.048$ $6.310$ $1.742$ $2.523$
+$\widehat{\beta}_{L}$ $0.039$ $1.048$ $0.084$ $1.030$ $0.623$ $1.577$ $0.299$ $1.029$
+Dynamic Model, $\theta_{0}=\left(0,0\right)$ Dynamic Model, $\theta_{0}=\left(0.5,0.5\right)$
+$\widehat{\beta}$ $0.104$ $2.501$ $0.081$ $1.342$ $3.931$ $6.008$ $1.953$ $2.982$
+$\widehat{\beta}_{L}$ $0.045$ $0.733$ $0.039$ $0.467$ $0.520$ $1.324$ $0.326$ $1.348$
+$\widehat{\rho}$ 0.331-0.331 $0.354$ 0.537-0.537 $0.572$ 0.266-0.266 $0.298$ 0.465-0.465 $0.507$
+$\widehat{\rho}_{L}$ 0.161-0.161 $0.197$ 0.230-0.230 $0.284$ 0.161-0.161 $0.222$ 0.240-0.240 $0.305$
+[…]
+Notes: $\left(N,T\right)=\left(90,10\right)$.
+For the dynamic model we set $\tau=1$, producing corrected estimators $\widehat{\rho}_{L}$ and $\widehat{\beta}_{L}$ for $\rho_{0}$ and $\beta_{0}$ respectively.
+F.2 Heterogeneous Autoregressive Parameters
+In this section, we present simulation studies for models with two-way heterogeneous autoregressive coefficient.
+We consider three models here: a Poission model, a logit model, and a probit model.
+For the Poisson, we consider a model specified according to the data-generating process $Y_{it}\sim\operatorname*{Poisson}\left(\lambda_{it}\right)$ for
+$\lambda_{it}=\exp\left\{\left(\rho_{0}+\alpha_{1,i}^{0}+\gamma_{1,t}^{0}\right)Y_{it-1}+(\beta_{0}+\alpha_{2,i}^{0}+\gamma_{2,t}^{0})Z_{it}+\alpha_{3,i}^{0}+\gamma_{3,t}^{0}\right\},$
+with $Y_{i0}\sim\operatorname*{Poisson}\left(\exp\left\{(\beta_{0}+\alpha_{1,i}^{0})Z_{i0}+\alpha_{2,i}^{0}\right\}\right)$, where the parameter of interests is $\theta_{0}=(\rho_{0},\beta_{0})^{\prime}=(-0.5,0.5)^{\prime}$.
+We set $\rho_{0}=-0.5$ to keep $Y_{it}$ from exploding as $t$ progresses.
+Similar to Section 4, we generate $\{\alpha_{k,i}^{0},\gamma_{k,t}^{0}\}\sim\mathcal{N}(0,0.04)$ for $k=1,\ldots,3$ and demean them afterwards [^10] As in Section 4, the individual effects are not normalized when the model is being estimated..
+The rest of the experiment is the same as Design 1 in Section 4.
+Table F.7 presents the simulation results.
+We see that the bias of the MLE $\widehat{\beta}$ of a static Poisson model is relative small, compared to the probit and logit models above.
+The biases in $\widehat{\beta}_{L}$, $\widehat{\beta}_{J}$, and $\widehat{\beta}_{B}$ are all very small, with $\widehat{\beta}_{L}$ and $\widehat{\beta}_{B}$ very similar and outperforming $\widehat{\beta}_{J}$.
+For the dynamic Poisson model, however, the MLE $\widehat{\rho}$ is severely biased, showing a bias of 0.204-0.204 for $\left(N,T\right)=\left(30,30\right)$.
+On the other hand, our bias correction procedure reduces the bias considerably, delivering $\widehat{\rho}_{L}^{(1)}$ which contains a bias of 0.077-0.077, roughly $1/3$ of the bias of the MLE.
+Our procedure also outperforms the jackknife and the bootstrap on $\rho$ for small samples.
+As the sample size increases, the biases of all estimators decrease, as expected, with the performance of our procedure and that of the bootstrap similar, outperforming the jackknife.
+For $\left(N,T\right)=\left(30,30\right)$, we notice that the bias of $\widehat{\rho}_{B}$ is $1.323$, which is very severe.
+We have double checked and this seems to be a valid result.
+We think this might be a coincidence and would like to emphasize that, as the sample size increases, the bootstrap is an effective bias correction device.
+For the hypothesis tests, we only report the results from the LR test for the dynamic model in Table F.8.
+For the static model, the sizes of the uncorrected, corrected, and the bootstrap LR test are all very close to the nominal level of $5\%$.
+The patterns of the LR test are similar to those from the probit and logit models in Section 4.
+For the LM and Wald test, the results are comparable to the LR test, and are, therefore, omitted here.
+Table F.7: Comparisons of Uncorrected and Corrected Poisson Estimates
 $(N,T)$ $(30,30)$ $(60,60)$ $(90,90)$
 Bias RMSE Bias RMSE Bias RMSE
-Static Model
-$\widehat{\beta}$ $0.150$ $0.197$ $0.063$ $0.082$ $0.042$ $0.054$
-$\widehat{\beta}_{L}$ $0.055$ $0.124$ $0.013$ $0.049$ $0.007$ $0.032$
-Dynamic Model
-$\widehat{\beta}$ $0.159$ $0.247$ $0.067$ $0.086$ $0.043$ $0.055$
-$\widehat{\beta}_{L}^{(1)}$ $0.059$ $0.134$ $0.015$ $0.052$ $0.008$ $0.032$
-$\widehat{\beta}_{L}^{(2)}$ $0.062$ $0.136$ $0.016$ $0.053$ $0.008$ $0.032$
-$\widehat{\rho}$ $-0.055$ $0.124$ $-0.022$ $0.054$ $-0.017$ $0.036$
-$\widehat{\rho}_{L}^{(1)}$ $-0.030$ $0.105$ $-0.009$ $0.047$ $-0.007$ $0.031$
-$\widehat{\rho}_{L}^{(2)}$ $-0.022$ $0.106$ $-0.003$ $0.047$ $-0.003$ $0.031$
+Static Poisson Model
+$\widehat{\beta}$ $0.016$ $0.043$ $0.005$ $0.018$ $0.003$ $0.012$
+$\widehat{\beta}_{L}$ 0.002-0.002 $0.039$ 0.002-0.002 $0.018$ 0.001-0.001 $0.011$
+$\widehat{\beta}_{J}$ 0.014-0.014 $0.051$ 0.004-0.004 $0.019$ 0.002-0.002 $0.012$
+$\widehat{\beta}_{B}$ $0.001$ $0.039$ $0.000$ $0.017$ $0.000$ $0.011$
+Dynamic Poisson Model
+$\widehat{\beta}$ $0.025$ $0.058$ $0.008$ $0.022$ $0.005$ $0.014$
+$\widehat{\beta}_{L}^{(1)}$ 0.005-0.005 $0.051$ 0.004-0.004 $0.021$ 0.002-0.002 $0.014$
+$\widehat{\beta}_{L}^{(2)}$ 0.005-0.005 $0.051$ 0.004-0.004 $0.021$ 0.002-0.002 $0.014$
+$\widehat{\beta}_{J}$ 0.060-0.060 $0.218$ 0.010-0.010 $0.025$ 0.004-0.004 $0.015$
+$\widehat{\beta}_{B}$ 0.003-0.003 $0.054$ $0.000$ $0.021$ $0.000$ $0.014$
+$\widehat{\rho}$ 0.204-0.204 $0.250$ 0.067-0.067 $0.075$ 0.041-0.041 $0.045$
+$\widehat{\rho}_{L}^{(1)}$ 0.077-0.077 $0.110$ 0.017-0.017 $0.036$ 0.006-0.006 $0.023$
+$\widehat{\rho}_{L}^{(2)}$ 0.086-0.086 $0.118$ 0.022-0.022 $0.038$ 0.010-0.010 $0.022$
+$\widehat{\rho}_{J}$ $0.588$ $0.821$ $0.060$ $0.083$ $0.014$ $0.027$
+$\widehat{\rho}_{B}$ $1.323$ $1.548$ $0.024$ $0.056$ $0.006$ $0.020$
 […]
-Notes: Probit model with $\varepsilon_{it}\sim\mathcal{N}(0,1)$.
-Design 1: $\{\alpha_{k,i}^{0},\gamma_{k,t}^{0}\}\sim\mathcal{N}(0,0.01)$, $X_{it}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0}+\gamma_{1,t}^{0}+\gamma_{2,t}^{0},0.25)$, and $X_{i0}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0},0.25)$.
-For the static model, $\widehat{\beta}$ and $\widehat{\beta}_{L}$ are, respectively, the uncorrected and corrected estimates of $\beta_{0}=0.5$.
-For the dynamic model $(\widehat{\beta},\widehat{\rho})^{\prime}$ is the uncorrected estimate of $(\rho_{0},\beta_{0})^{\prime}=(0.5,0.5)^{\prime}$ from and $(\widehat{\beta}_{L}^{(\tau)},\widehat{\rho}_{L}^{(\tau)})^{\prime}$ is the corrected estimate with truncation parameter $\tau\in(1,2)$.
-The number of replications is $1000$.
-Table B.7: Empirical Size of LR, LM, and Wald Tests Based on Corrected and Uncorrected Likelihoods for Probit Model, Design 1
-(a) LR Tests
-Static Model Dynamic Model
-$(N,T)$ Uncorrected Corrected Uncorrected Corrected
-$(30,30)$ $29\%$ $13\%$ $27\%$ $13\%$
-$(60,60)$ $26\%$ $8\%$ $23\%$ $7\%$
-$(90,90)$ $30\%$ $6\%$ $24\%$ $6\%$
-(b) LM Tests
-Static Model Dynamic Model
-$(N,T)$ Uncorrected Corrected Uncorrected Corrected
-$(30,30)$ $28\%$ $13\%$ $26\%$ $13\%$
-$(60,60)$ $26\%$ $8\%$ $22\%$ $7\%$
-$(90,90)$ $30\%$ $6\%$ $24\%$ $6\%$
-(c) Wald Tests
-Static Model Dynamic Model
-$(N,T)$ Uncorrected Corrected Uncorrected Corrected
-$(30,30)$ $28\%$ $13\%$ $25\%$ $13\%$
-$(60,60)$ $26\%$ $8\%$ $22\%$ $7\%$
-$(90,90)$ $30\%$ $6\%$ $24\%$ $6\%$
+Table F.8: Empirical Size and Power from LR Test for Two-Way Fixed-Effect Probit Estimates
+(a) Empirical Size
+Size
+$(N,T)$ Uncorrected Analytical Bootstrap
+$(30,30)$ $53$ $25$ $12$
+$(60,60)$ $48$ $12$ 55
+$(90,90)$ $46$ 88 44
+(b) Empirical Power
+Power at $\delta$ (Analytical) Power at $\delta$ (Bootstrap)
+$(N,T)$ 0.2-0.2 0.1-0.1 $0.1$ $0.2$ 0.2-0.2 0.1-0.1 $0.1$ $0.2$
+$(30,30)$ $99$ $57$ $93$ $100$ $76$ 1919 $56$ $98$
+$(60,60)$ $99$ $99$ $99$ $99$ $100$ $89$ $100$ $100$
+$(90,90)$ $99$ $99$ $99$ $99$ $99$ $99$ $99$ $99$
 […]
-Notes: Probit model with $\varepsilon_{it}\sim\mathcal{N}(0,1)$.
-Design 1: $\{\alpha_{k,i}^{0},\gamma_{k,t}^{0}\}\sim\mathcal{N}(0,0.01)$, $X_{it}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0}+\gamma_{1,t}^{0}+\gamma_{2,t}^{0},0.25)$, and $X_{i0}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0},0.25)$.
-The null hypothesis $H_{0}:(\beta_{0},\rho_{0})^{\prime}=(0.5,0.5)^{\prime}$ for the dynamic model, and $H_{0}:\beta_{0}=0.5$ for the static model, with a significance level $5\%$.
-The number of replications is $1000$.
-Table B.8: Empirical Power of LR, LM, and Wald Tests Based on Bias-corrected Likelihoods for Probit Model, Design 1
-(a) LR Tests
-$\delta$
-$(N,T)$ $-0.20$ $-0.15$ $-0.10$ $0.10$ $0.15$ $0.20$
-Static Model
-$(30,30)$ $75\%$ $59\%$ $41\%$ $13\%$ $23\%$ $37\%$
-$(60,60)$ $100\%$ $96\%$ $71\%$ $52\%$ $86\%$ $98\%$
-$(90,90)$ $100\%$ $100\%$ $96\%$ $88\%$ $99\%$ $100\%$
-Dynamic Model
-$(30,30)$ $83\%$ $63\%$ $40\%$ $26\%$ $47\%$ $69\%$
-$(60,60)$ $100\%$ $98\%$ $78\%$ $78\%$ $98\%$ $100\%$
-$(90,90)$ $100\%$ $100\%$ $99\%$ $98\%$ $100\%$ $100\%$
-(b) LM Tests
-$\delta$
-$(N,T)$ $-0.20$ $-0.15$ $-0.10$ $0.10$ $0.15$ $0.20$
-Static Model
-$(30,30)$ $75\%$ $59\%$ $40\%$ $13\%$ $23\%$ $37\%$
-$(60,60)$ $100\%$ $95\%$ $71\%$ $52\%$ $86\%$ $98\%$
-$(90,90)$ $100\%$ $100\%$ $96\%$ $88\%$ $99\%$ $100\%$
-Dynamic Model
-$(30,30)$ $83\%$ $62\%$ $38\%$ $26\%$ $47\%$ $69\%$
-$(60,60)$ $100\%$ $98\%$ $78\%$ $78\%$ $98\%$ $100\%$
-$(90,90)$ $100\%$ $100\%$ $99\%$ $98\%$ $100\%$ $100\%$
-(c) Wald Tests
-$\delta$
-$(N,T)$ $-0.20$ $-0.15$ $-0.10$ $0.10$ $0.15$ $0.20$
-Static Model
-$(30,30)$ $75\%$ $58\%$ $40\%$ $13\%$ $23\%$ $37\%$
-$(60,60)$ $100\%$ $95\%$ $72\%$ $52\%$ $86\%$ $98\%$
-$(90,90)$ $100\%$ $100\%$ $95\%$ $88\%$ $99\%$ $100\%$
-Dynamic Model
-$(30,30)$ $82\%$ $62\%$ $38\%$ $26\%$ $47\%$ $69\%$
-$(60,60)$ $100\%$ $98\%$ $78\%$ $78\%$ $98\%$ $100\%$
-$(90,90)$ $100\%$ $100\%$ $99\%$ $98\%$ $100\%$ $100\%$
+For the two binary response models, we find that adding heterogeneities to the autoregressive coefficient may induce some numerical difficulty.
+To deal with it, we apply certain restrictions to the data-generating process.
+In Remark 8 below, we first explain the numerical issue, followed by introducing the data-generating process and the restrictions we apply.
+Remark 8 (binary response model with heterogeneous autoregressive coefficient)
+For binary response models with heterogeneous autogressive coefficient, numerical difficulty may arise if the (lagged) dependent variable contains insufficient variations.
+Consider a simple binary response model
+$Y_{it}=\mathds{1}\{\left(\rho_{0}+\alpha_{1,i}+\gamma_{1,t}\right)Y_{it-1}+\alpha_{2,i}+\gamma_{2,t}+\varepsilon_{it}>0\}.$
+Suppose, for an individual $i$, there are too many time periods $t$ such that $Y_{it-1}=1$.
+Then, the regressor vector $\left(Y_{i1},\ldots,Y_{iT-1}\right)^{\prime}$ could be nearly collinear to the constant of $1$.
+This renders the separate identification of $\alpha_{1,i}$ and $\alpha_{2,i}$ numerically difficult.
+Similar situations also happen if, for some $t$, there are too many individuals with $Y_{it-1}=1$.
+When there are excessive zeros in $Y_{it-1}$ for some $i$ or $t$, the identification of $\alpha_{1,i}\$or $\gamma_{1,t}$ is also very difficult.
+For the simulation, we consider a binary response model specified according to the data-generating process
+$Y_{it}=\mathds{1}\{\left(\rho_{0}+\alpha_{1,i}^{0}+\gamma_{1,t}^{0}\right)Y_{it-1}+(\beta_{0}+\alpha_{2,i}^{0}+\gamma_{2,t}^{0})Z_{it}+\delta_{0}+\alpha_{2,i}^{0}+\gamma_{2,t}^{0}+\varepsilon_{it}>0\},$
+where, for $k=1,2,3$, $\{\alpha_{k,i}^{0},\gamma_{k,t}^{0}\}\sim\mathcal{N}(0,0.04)$ and are demeaned; $\varepsilon_{it}$ are either standard normal (probit) or standard logistic (logit); and $\delta_{0}$ is explained below.
+All other settings are the same as in Section 4, except that we restrict both $0.1\leq T^{-1}\sum_{t=1}^{T}Y_{it}\leq 0.6$ for all $i$ and $0.1\leq N^{-1}\sum_{i=1}^{N}Y_{it}\leq 0.6$ for all $t$.
+An additional common parameter $\delta_{0}=-1$ is included in the intercept to make it easier to generate a data set which satisfies the two restrictions [^11] The response $Y_{it}$ tends to contain too many ones without $\delta_{0}=-1$.
+This is because $\left(\rho_{0}+\alpha_{1,i}^{0}+\gamma_{1,t}^{0}\right)Y_{it-1}$ is positive on average..
+Note that $\delta_{0}$ is considered nuisance and is not a part of $\theta_{0}$.
+These measures seem to alleviate the numerical difficulty considerably, although they are, admittedly, somewhat ad-hoc.
+Table F.9 presents the simulation results for both the probit and logit models with heterogeneous autoregressive coefficient.
+For dynamic models, we see that $\widehat{\rho}$ exhibits larger biases, and our bias correction procedure reduces the bias considerably.
+Under the current setting, neither the jackknife nor the bootstrap perform better than our bias correction procedure.
+For static models, the conclusion is the same as in Section 4, except that the biases of all estimators are slightly larger (which is likely caused by having $\delta_{0}=-1$).
+Table F.10 presents the empirical sizes and powers from the LR tests.
+We see that the LR test based on the uncorrected likelihood is fatally size-distorted.
+This is anticipated because the IPP is much more severe with an additional set of heterogeneities.
+Our bias correction procedure reduces the empirical type-I error risk considerably.
+Similar to the results in Section 4, the bootstrap reduces the empirical type-I error risk at the cost of reducing the power of the test as well.
+Table F.9: Comparisons of Uncorrected and Corrected Probit and Logit Estimates with Heterogeneous Autoregressive Coefficient, Design 1
+$(N,T)$ $(30,30)$ $(60,60)$ $(90,90)$
+Bias RMSE Bias RMSE Bias RMSE
+Static Logit Model
+$\widehat{\beta}$ $0.196$ $0.255$ $0.081$ $0.097$ $0.050$ $0.059$
+$\widehat{\beta}_{L}$ $0.065$ $0.128$ $0.018$ $0.051$ $0.007$ $0.030$
+$\widehat{\beta}_{J}$ 1.344-1.344 $1.939$ 0.049-0.049 $0.108$ 0.013-0.013 $0.033$
+$\widehat{\beta}_{B}$ 0.143-0.143 $0.179$ 0.014-0.014 $0.046$ 0.006-0.006 $0.029$
+Static Probit Model
+$\widehat{\beta}$ $0.374$ $0.566$ $0.105$ $0.113$ $0.062$ $0.066$
+$\widehat{\beta}_{L}$ $0.125$ $0.164$ $0.038$ $0.053$ $0.017$ $0.027$
+$\widehat{\beta}_{J}$ 5.362-5.362 $7.774$ 0.350-0.350 $1.101$ 0.037-0.037 $0.059$
+$\widehat{\beta}_{B}$ 0.704-0.704 $0.923$ 0.032-0.032 $0.043$ 0.009-0.009 $0.022$
+Dynamic Logit Model
+$\widehat{\beta}$ $0.279$ $0.332$ $0.106$ $0.119$ $0.065$ $0.073$
+$\widehat{\beta}_{L}^{(1)}$ $0.101$ $0.160$ $0.026$ $0.054$ $0.011$ $0.032$
+$\widehat{\beta}_{L}^{(2)}$ $0.105$ $0.164$ $0.027$ $0.055$ $0.012$ $0.032$
+$\widehat{\beta}_{J}$ 2.345-2.345 $3.017$ 0.085-0.085 $0.138$ 0.020-0.020 $0.037$
+$\widehat{\beta}_{B}$ 0.369-0.369 $0.407$ 0.029-0.029 $0.051$ 0.011-0.011 $0.030$
+$\widehat{\rho}$ 0.546-0.546 $0.725$ 0.120-0.120 $0.169$ 0.059-0.059 $0.089$
+$\widehat{\rho}_{L}^{(1)}$ 0.194-0.194 $0.396$ 0.050-0.050 $0.127$ 0.022-0.022 $0.063$
+$\widehat{\rho}_{L}^{(2)}$ 0.193-0.193 $0.404$ 0.047-0.047 $0.127$ 0.019-0.019 $0.063$
+$\widehat{\rho}_{J}$ 0.363-0.363 $6.295$ $0.305$ $0.577$ $0.099$ $0.267$
+$\widehat{\rho}_{B}$ $0.551$ $0.734$ $0.160$ $0.186$ $0.043$ $0.071$
+Dynamic Probit Model
+$\widehat{\beta}$ $0.500$ $0.700$ $0.130$ $0.137$ $0.079$ $0.083$
+$\widehat{\beta}_{L}^{(1)}$ $0.179$ $0.215$ $0.046$ $0.059$ $0.023$ $0.032$
+$\widehat{\beta}_{L}^{(2)}$ $0.189$ $0.261$ $0.046$ $0.077$ $0.023$ $0.032$
+$\widehat{\beta}_{J}$ 7.236-7.236 $9.499$ 0.499-0.499 $0.932$ 0.057-0.057 $0.104$
+$\widehat{\beta}_{B}$ 2.386-2.386 $4.786$ 0.062-0.062 $0.069$ 0.018-0.018 $0.027$
+$\widehat{\rho}$ 0.518-0.518 $0.781$ 0.173-0.173 $0.208$ 0.083-0.083 $0.106$
+$\widehat{\rho}_{L}^{(1)}$ 0.226-0.226 $0.485$ 0.039-0.039 $0.259$ 0.024-0.024 $0.155$
+$\widehat{\rho}_{L}^{(2)}$ 0.213-0.213 $0.533$ 0.037-0.037 $0.261$ 0.023-0.023 $0.167$
+$\widehat{\rho}_{J}$ 4.299-4.299 $12.866$ 0.066-0.066 $2.459$ $0.116$ $1.033$
+$\widehat{\rho}_{B}$ 1.621-1.621 $3.122$ $0.126$ $0.186$ $0.091$ $0.105$
 […]
-Notes: Probit model with $\varepsilon_{it}\sim\mathcal{N}(0,1)$.
-Design 1: $\{\alpha_{k,i}^{0},\gamma_{k,t}^{0}\}\sim\mathcal{N}(0,0.01)$, $X_{it}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0}+\gamma_{1,t}^{0}+\gamma_{2,t}^{0},0.25)$, and $X_{i0}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0},0.25)$.
-The null hypothesis $H_{0}:(\beta_{0},\rho_{0})^{\prime}=(0.5+\delta,0.5+\delta)^{\prime}$ for the dynamic model, and $H_{0}:\beta_{0}=0.5+\delta$ for the static model, with a significance level $5\%,$ where $\delta\in\{\pm 0.2,\pm 0.15,\pm 0.1\}$.
-The number of replications is $1000$.
-Table B.9: Empirical Size of LM and Wald Tests Based on Corrected and Uncorrected Likelihoods for Probit Model, Design 2
-(a) LM Tests
-Static Model Dynamic Model
-$(N,T)$ Uncorrected Corrected Uncorrected Corrected
-$(30,30)$ $69\%$ $34\%$ $59\%$ $30\%$
-$(60,60)$ $74\%$ $18\%$ $62\%$ $17\%$
-$(90,90)$ $71\%$ $14\%$ $60\%$ $13\%$
-(b) Wald Tests
-Static Model Dynamic Model
-$(N,T)$ Uncorrected Corrected Uncorrected Corrected
-$(30,30)$ $37\%$ $29\%$ $25\%$ $26\%$
-$(60,60)$ $71\%$ $18\%$ $56\%$ $16\%$
-$(90,90)$ $72\%$ $14\%$ $60\%$ $13\%$
+Table F.10: Empirical Size and Power from LR Test for Probit and Logit Models with Heterogeneous Autoregressive Coefficient, Design 1
+(a) Empirical Size
+Size
+$(N,T)$ Uncorrected Analytical Bootstrap
+Static Logit Model
+$(30,30)$ $38$ $16$ 11
+$(60,60)$ $41$ $11$ 33
+$(90,90)$ $39$ 77 22
+Static Probit Model
+$(30,30)$ $78$ $41$ 99
+$(60,60)$ $81$ $26$ 33
+$(90,90)$ $79$ $15$ 33
+Dynamic Logit Model
+$(30,30)$ $62$ $37$ 88
+$(60,60)$ $61$ $16$ 22
+$(90,90)$ $55$ $10$ 22
+Dynamic Probit Model
+$(30,30)$ $93$ $78$ $52$
+$(60,60)$ $93$ $45$ 77
+$(90,90)$ $94$ $30$ 22
+(b) Empirical Power
+Power at $\delta$ (Analytical) Power at $\delta$ (Bootstrap)
+$(N,T)$ 0.2-0.2 0.1-0.1 $0.1$ $0.2$ 0.2-0.2 0.1-0.1 $0.1$ $0.2$
+Static Logit Model
+$(30,30)$ $83$ $44$ $13$ $37$ $36$ 88 11 11
+$(60,60)$ $100$ $77$ $49$ $98$ $100$ $62$ 11 2121
+$(90,90)$ $100$ $97$ $91$ $100$ $100$ $95$ 66 $91$
+Static Probit Model
+$(30,30)$ $99$ $85$ $12$ $27$ $61$ 1919 55 44
+$(60,60)$ $100$ $99$ $53$ $99$ $100$ $87$ 0 55
+$(90,90)$ $100$ $100$ $97$ $100$ $100$ $100$ 11 $88$
+Dynamic Logit Model
+$(30,30)$ $80$ $54$ $39$ $59$ $33$ $14$ 55 66
+$(60,60)$ $100$ $77$ $59$ $99$ $98$ $42$ 11 2020
+$(90,90)$ $100$ $97$ $93$ $100$ $100$ $87$ 66 $90$
+Dynamic Probit Model
+$(30,30)$ $99$ $91$ $67$ $71$ $81$ $62$ $44$ $40$
+$(60,60)$ $100$ $98$ $76$ $100$ $100$ $81$ 22 $11$
+$(90,90)$ $100$ $100$ $98$ $100$ $100$ $99$ 22 $84$
 […]
-Notes: Probit model with $\varepsilon_{it}\sim\mathcal{N}(0,1)$.
-Design 2: $\{\alpha_{k,i}^{0},\gamma_{k,t}^{0}\}\sim\mathcal{N}(0,1/4)$, $X_{it}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0}+\gamma_{1,t}^{0}+\gamma_{2,t}^{0},1)$, and $X_{i0}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0},1)$.
-The null hypothesis $H_{0}:(\beta_{0},\rho_{0})^{\prime}=(0.5,0.5)^{\prime}$ for the dynamic model, and $H_{0}:\beta_{0}=0.5$ for the static model, with a significance level $5\%$.
-The number of replications is $1000$.
-Table B.10: Empirical Power of LR, LM, and Wald Tests Based on Bias-corrected Likelihoods for Probit Model, Design 2
-(a) LR Tests
-$\delta$
-$(N,T)$ $-0.20$ $-0.15$ $-0.10$ $0.10$ $0.15$ $0.20$
-Static Model
-$(30,30)$ $97\%$ $90\%$ $77\%$ $14\%$ $13\%$ $15\%$
-$(60,60)$ $100\%$ $100\%$ $96\%$ $34\%$ $70\%$ $83\%$
-$(90,90)$ $100\%$ $100\%$ $99\%$ $85\%$ $98\%$ $96\%$
-Dynamic Model
-$(30,30)$ $95\%$ $86\%$ $74\%$ $25\%$ $28\%$ $38\%$
-$(60,60)$ $100\%$ $100\%$ $93\%$ $47\%$ $84\%$ $96\%$
-$(90,90)$ $100\%$ $100\%$ $100\%$ $90\%$ $99\%$ $100\%$
-(b) LM Tests
-$\delta$
-$(N,T)$ $-0.20$ $-0.15$ $-0.10$ $0.10$ $0.15$ $0.20$
-Static Model
-$(30,30)$ $96\%$ $87\%$ $73\%$ $12\%$ $13\%$ $17\%$
-$(60,60)$ $100\%$ $100\%$ $96\%$ $37\%$ $69\%$ $76\%$
-$(90,90)$ $100\%$ $100\%$ $99\%$ $86\%$ $91\%$ $88\%$
-Dynamic Model
-$(30,30)$ $94\%$ $83\%$ $70\%$ $21\%$ $27\%$ $38\%$
-$(60,60)$ $100\%$ $99\%$ $92\%$ $48\%$ $81\%$ $91\%$
-$(90,90)$ $100\%$ $99\%$ $99\%$ $90\%$ $94\%$ $94\%$
-(c) Wald Tests
-$\delta$
-$(N,T)$ $-0.20$ $-0.15$ $-0.10$ $0.10$ $0.15$ $0.20$
-Static Model
-$(30,30)$ $91\%$ $83\%$ $68\%$ $10\%$ $13\%$ $19\%$
-$(60,60)$ $99\%$ $99\%$ $95\%$ $39\%$ $76\%$ $92\%$
-$(90,90)$ $99\%$ $99\%$ $99\%$ $88\%$ $99\%$ $99\%$
-Dynamic Model
-$(30,30)$ $88\%$ $75\%$ $60\%$ $21\%$ $27\%$ $39\%$
-$(60,60)$ $100\%$ $99\%$ $91\%$ $49\%$ $85\%$ $98\%$
-$(90,90)$ $99\%$ $99\%$ $99\%$ $91\%$ $99\%$ $100\%$
+F.3 Probit and Logit under Design 2
+In this section, we provide the simulation results under the same settings as in Section 4, except that $\{\alpha_{k,i}^{0},\gamma_{k,t}^{0}\}\sim\mathcal{N}(0,0.25)$, $Z_{it}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0}+\gamma_{1,t}^{0}+\gamma_{2,t}^{0},1)$, and $Z_{i0}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0},1)$.
+Table F.11: Comparisons of Uncorrected and Corrected Logit and Probit Estimates, Design 2
+$(30,30)$ $(60,60)$ $(90,90)$ $(120,120)$ $(150,150)$
+Bias RMSE Bias RMSE Bias RMSE Bias RMSE Bias RMSE
+Static Logit Model
+$\widehat{\beta}$ $0.45$ $1.20$ $0.10$ $0.13$ $0.06$ $0.07$ $0.08$ $0.05$ $0.07$ $0.04$
+$\widehat{\beta}_{L}$ $0.10$ $0.20$ $0.02$ $0.06$ $0.01$ $0.04$ $0.01$ $0.03$ $0.01$ $0.02$
+Static Probit Model
+$\widehat{\beta}$ $0.94$ $2.28$ $0.15$ $0.25$ $0.07$ $0.08$ $0.10$ $0.05$ $0.08$ $0.04$
+$\widehat{\beta}_{L}$ $0.33$ $5.17$ $0.04$ $0.06$ $0.02$ $0.04$ $0.02$ $0.03$ $0.01$ $0.02$
+Dynamic Logit Model
+$\widehat{\beta}$ $0.53$ $1.58$ $0.11$ $0.21$ $0.06$ $0.07$ $0.08$ $0.05$ $0.06$ $0.04$
+$\widehat{\beta}_{L}^{(1)}$ $0.12$ $0.20$ $0.02$ $0.06$ $0.01$ $0.04$ $0.01$ $0.03$ $0.01$ $0.02$
+$\widehat{\beta}_{L}^{(2)}$ $0.13$ $0.28$ $0.03$ $0.06$ $0.01$ $0.04$ $0.01$ $0.03$ $0.01$ $0.02$
+$\widehat{\rho}$ 0.11-0.11 $0.24$ 0.06-0.06 $0.11$ 0.04-0.04 $0.07$ 0.05-0.05 $0.05$ 0.04-0.04 $0.04$
+$\widehat{\rho}_{L}^{(1)}$ 0.04-0.04 $0.19$ 0.02-0.02 $0.09$ 0.01-0.01 $0.06$ 0.01-0.01 $0.04$ 0.01-0.01 $0.03$
+$\widehat{\rho}_{L}^{(2)}$ 0.04-0.04 $0.20$ 0.01-0.01 $0.09$ 0.01-0.01 $0.06$ $0.00$ $0.04$ $0.00$ $0.03$
+Dynamic Probit Model
+$\widehat{\beta}$ $0.91$ $1.84$ $0.17$ $0.25$ $0.08$ $0.10$ $0.11$ $0.06$ $0.08$ $0.04$
+$\widehat{\beta}_{L}^{(1)}$ $0.21$ $0.40$ $0.05$ $0.08$ $0.02$ $0.04$ $0.02$ $0.03$ $0.01$ $0.03$
+$\widehat{\beta}_{L}^{(2)}$ $0.20$ $0.28$ $0.05$ $0.08$ $0.02$ $0.04$ $0.03$ $0.03$ $0.02$ $0.04$
+$\widehat{\rho}$ 0.03-0.03 $0.16$ 0.02-0.02 $0.07$ 0.01-0.01 $0.05$ 0.02-0.02 $0.03$ 0.02-0.02 $0.03$
+$\widehat{\rho}_{L}^{(1)}$ 0.02-0.02 $0.14$ 0.01-0.01 $0.07$ 0.00-0.00 $0.04$ 0.01-0.01 $0.04$ 0.01-0.01 $0.03$
+$\widehat{\rho}_{L}^{(2)}$ 0.01-0.01 $0.14$ 0.00-0.00 $0.07$ 0.00-0.00 $0.05$ $0.00$ $0.04$ 0.01-0.01 $0.04$
 […]
-Notes: Probit model with $\varepsilon_{it}\sim\mathcal{N}(0,1)$.
-Design 2: $\{\alpha_{k,i}^{0},\gamma_{k,t}^{0}\}\sim\mathcal{N}(0,1/4)$, $X_{it}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0}+\gamma_{1,t}^{0}+\gamma_{2,t}^{0},1)$, and $X_{i0}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0},1)$.
-The null hypothesis $H_{0}:(\beta_{0},\rho_{0})^{\prime}=(0.5+\delta,0.5+\delta)^{\prime}$ for the dynamic model, and $H_{0}:\beta_{0}=0.5+\delta$ for the static model, with a significance level $5\%,$ where $\delta\in\{\pm 0.2,\pm 0.15,\pm 0.1\}$.
-The number of replications is $1000$.
-C Additional Empirical Results
-C.1 Sample
-The data set is constructed from waves $20-30$ of PSID, covering the survey years 1987 to 1997.
+Table F.12: Empirical Size and Power from LR Test for Logit and Probit Model, Design 2
+Size Power at $\delta$
+$(N,T)$ Uncorrected Analytical 0.2-0.2 0.1-0.1 $0.1$ $0.2$
+Static Logit Model
+$(30,30)$ $47$ 2020 $79$ $45$ $12$ $26$
+$(60,60)$ $43$ $10$ $100$ $73$ $40$ $93$
+$(90,90)$ $42$ 99 $100$ $94$ $81$ $100$
+$(120,120)$ $40$ 77 $100$ $99$ $97$ $99$
+$(150,150)$ $39$ 55 $100$ $100$ $100$ $100$
+Static Probit Model
+$(30,30)$ $77$ $39$ $97$ $77$ $14$ $15$
+$(60,60)$ $76$ 2020 $100$ $96$ $34$ $83$
+$(90,90)$ $72$ $14$ $100$ $99$ $85$ $96$
+$(120,120)$ $72$ $10$ $99$ $99$ $98$ $98$
+$(150,150)$ $72$ 99 $99$ $99$ $99$ $99$
+Dynamic Logit Model
+$(30,30)$ $43$ $18$ $75$ $41$ 1919 $35$
+$(60,60)$ $38$ 99 $100$ $69$ $44$ $96$
+$(90,90)$ $34$ 77 $100$ $93$ $86$ $100$
+$(120,120)$ $33$ 66 $100$ $99$ $98$ $100$
+$(150,150)$ $31$ 55 $100$ $100$ $100$ $100$
+Dynamic Probit Model
+$(30,30)$ $71$ $37$ $95$ $74$ $25$ $38$
+$(60,60)$ $66$ 1919 $100$ $93$ $47$ $96$
+$(90,90)$ $63$ $14$ $100$ $100$ $90$ $100$
+$(120,120)$ $64$ $11$ $99$ $99$ $99$ $99$
+$(150,150)$ $62$ 88 $99$ $99$ $99$ $99$
+[…]
+Table F.13: Empirical Size and Power from LM Test for Logit and Probit Model, Design 2
+Size Power at $\delta$
+$(N,T)$ Uncorrected Analytical 0.2-0.2 0.1-0.1 $0.1$ $0.2$
+Static Logit Model
+$(30,30)$ $42$ $18$ $77$ $42$ $12$ $28$
+$(60,60)$ $42$ 99 $99$ $72$ $41$ $93$
+$(90,90)$ $41$ 99 $100$ $94$ $82$ $100$
+$(120,120)$ $40$ 77 $100$ $99$ $97$ $100$
+$(150,150)$ $39$ 55 $100$ $99$ $100$ $100$
+Static Probit Model
+$(30,30)$ $69$ $34$ $96$ $73$ $12$ $17$
+$(60,60)$ $74$ $18$ $100$ $96$ $37$ $76$
+$(90,90)$ $71$ $14$ $100$ $99$ $86$ $88$
+$(120,120)$ $72$ $10$ $99$ $99$ $96$ $91$
+$(150,150)$ $72$ 99 $99$ $100$ $98$ $94$
+Dynamic Logit Model
+$(30,30)$ $38$ $16$ $73$ $38$ 1919 $36$
+$(60,60)$ $37$ 99 $100$ $68$ $45$ $94$
+$(90,90)$ $34$ 77 $100$ $93$ $86$ $99$
+$(120,120)$ $33$ 66 $100$ $99$ $98$ $99$
+$(150,150)$ $31$ 55 $100$ $100$ $100$ $99$
+Dynamic Probit Model
+$(30,30)$ $59$ $30$ $94$ $70$ 2121 $38$
+$(60,60)$ $62$ $17$ $100$ $92$ $48$ $91$
+$(90,90)$ $60$ $13$ $100$ $99$ $90$ $94$
+$(120,120)$ $62$ $10$ $99$ $99$ $97$ $96$
+$(150,150)$ $62$ 77 $99$ $99$ $99$ $96$
+[…]
+Table F.14: Empirical Size and Power from Wald Test for Logit and Probit Model, Design 2
+Size Power at $\delta$
+$(N,T)$ Uncorrected Analytical 0.2-0.2 0.1-0.1 $0.1$ $0.2$
+Static Logit Model
+$(30,30)$ $32$ $15$ $75$ $39$ $10$ $29$
+$(60,60)$ $41$ 99 $100$ $72$ $41$ $95$
+$(90,90)$ $42$ 99 $100$ $94$ $81$ $100$
+$(120,120)$ $40$ 77 $100$ $99$ $97$ $100$
+$(150,150)$ $39$ 55 $100$ $100$ $100$ $100$
+Static Probit Model
+$(30,30)$ $37$ $29$ $91$ $68$ $10$ 1919
+$(60,60)$ $71$ $18$ $99$ $95$ $39$ $92$
+$(90,90)$ $72$ $14$ $99$ $99$ $88$ $99$
+$(120,120)$ $72$ $10$ $99$ $99$ $98$ $99$
+$(150,150)$ $72$ 99 $99$ $99$ $99$ $99$
+Dynamic Logit Model
+$(30,30)$ $27$ $15$ $69$ $34$ $18$ $36$
+$(60,60)$ $36$ 88 $100$ $68$ $45$ $97$
+$(90,90)$ $34$ 77 $100$ $92$ $86$ $100$
+$(120,120)$ $33$ 66 $100$ $99$ $98$ $100$
+$(150,150)$ $31$ 55 $100$ $100$ $100$ $100$
+Dynamic Probit Model
+$(30,30)$ $25$ $26$ $88$ $60$ 2121 $39$
+$(60,60)$ $56$ $16$ $100$ $91$ $49$ $98$
+$(90,90)$ $60$ $13$ $99$ $99$ $91$ $100$
+$(120,120)$ $64$ $10$ $99$ $99$ $98$ $98$
+$(150,150)$ $62$ 77 $99$ $99$ $99$ $99$
+[…]
+Table F.15: Comparisons of Uncorrected, Corrected, and Jackknife Estimates for Logit and Probit, Design 2
+$(N,T)$ $(30,30)$ $(60,60)$ $(90,90)$
+Bias RMSE Bias RMSE Bias RMSE
+Static Logit Model
+$\widehat{\beta}$ $0.400$ $1.435$ $0.101$ $0.125$ $0.060$ $0.072$
+$\widehat{\beta}_{L}$ $0.093$ $0.188$ $0.022$ $0.056$ $0.010$ $0.036$
+$\widehat{\beta}_{J}$ 2.314-2.314 $4.328$ 0.172-0.172 $0.406$ 0.031-0.031 $0.070$
+Static Probit Model
+$\widehat{\beta}$ $0.846$ $2.168$ $0.151$ $0.261$ $0.074$ $0.082$
+$\widehat{\beta}_{L}$ $0.182$ $0.339$ $0.041$ $0.068$ $0.019$ $0.033$
+$\widehat{\beta}_{J}$ 4.929-4.929 $14.383$ 0.504-0.504 $0.981$ 0.105-0.105 $0.214$
+Dynamic Logit Model
+$\widehat{\beta}$ $0.415$ $0.858$ $0.104$ $0.137$ $0.061$ $0.074$
+$\widehat{\beta}_{L}$ $0.104$ $0.201$ $0.023$ $0.060$ $0.012$ $0.043$
+$\widehat{\beta}_{J}$ 2.442-2.442 $4.223$ 0.231-0.231 $0.481$ 0.043-0.043 $0.105$
+$\widehat{\rho}$ 0.120-0.120 $0.257$ 0.055-0.055 $0.108$ 0.036-0.036 $0.071$
+$\widehat{\rho}_{L}$ 0.047-0.047 $0.205$ 0.016-0.016 $0.089$ 0.010-0.010 $0.059$
+$\widehat{\rho}_{J}$ $0.020$ $0.218$ $0.003$ $0.092$ $0.002$ $0.060$
+Dynamic Probit Model
+$\widehat{\beta}$ $1.040$ $2.343$ $0.184$ $0.287$ $0.080$ $0.090$
+$\widehat{\beta}_{L}$ $0.209$ $0.332$ $0.052$ $0.082$ $0.020$ $0.049$
+$\widehat{\beta}_{J}$ 3.352-3.352 $7.708$ 0.558-0.558 $1.125$ 0.118-0.118 $0.217$
+$\widehat{\rho}$ 0.046-0.046 $0.169$ 0.022-0.022 $0.072$ 0.014-0.014 $0.046$
+$\widehat{\rho}_{L}$ 0.028-0.028 $0.145$ 0.011-0.011 $0.067$ 0.009-0.009 $0.053$
+$\widehat{\rho}_{J}$ 0.029-0.029 $0.169$ 0.006-0.006 $0.068$ 0.002-0.002 $0.044$
+[…]
+Notes: Simulated separately from Table F.11.
+For dynamic models, We set $\tau=1$ in $L\left(\theta\right)$, producing corrected estimators $\widehat{\rho}_{L}$ and $\widehat{\beta}_{L}$ for $\rho_{0}$ and $\beta_{0}$ respectively.
+F.4 Two-way Fixed-Effect Probit
+In this section, we consider a dynamic probit model with two-way scalar fixed effects appearing only in the intercept.
+This is a special case of the two-way DN-HP model.
+We take Design 1 in Fernández-Val and Weidner (2016) and simulate our method, the jackknife, and the bootstrap here.
+We set $N=56$ and $T\in\left\{14,28,56\right\}$, and consider
+$Y_{it}=\mathds{1}\left\{\rho_{0}Y_{it-1}+\beta_{0}Z_{it}+\alpha_{i}^{0}+\gamma_{t}^{0}+\varepsilon_{it}>0\right\}$
+with $Y_{i0}=\mathds{1}\left\{\beta_{0}Z_{i0}+\alpha_{i}^{0}+\gamma_{0}^{0}+\varepsilon_{i0}>0\right\}$, where $\rho_{0}=0.5$, $\beta_{0}=1$, $\varepsilon_{it}$ and $\varepsilon_{i0}$ are standard normal, $\left\{\alpha_{i}^{0},\gamma_{t}^{0},\gamma_{0}^{0}\right\}\sim\mathcal{N}\left(0,0.0625\right)$, and $Z_{it}=Z_{it-1}/2+\alpha_{i0}+\gamma_{t0}+\nu_{it}$ with $\nu_{it}\sim\mathcal{N}\left(0,1/2\right)$ and $Z_{i0}\sim\mathcal{N}\left(0,1\right)$.
+We set $\tau\in\left\{1,2\right\}$.
+For the hypothesis tests, we only present the results from the LR tests, using the uncorrected likelihood $\widehat{l}\left(\theta\right)$, the corrected likelihood $L\left(\theta\right)$ (setting $\tau=1$), and the bootstrap.
+We set the number of repetitions to $499$ for the bootstrap and run $1000$ replications.
+The rest are the same as in Section 4.
+Table F.16 presents the simulation results.
+First, the biases and RMSEs of $\widehat{\beta}$, $\widehat{\beta}_{J}$, $\widehat{\rho}$, and $\widehat{\rho}_{J}$ are in line with the original simulations of Fernández-Val and Weidner (2016).
+For instance for $\left(N,T\right)=\left(56,14\right)$, the bias of $\widehat{\beta}$ is $0.182$, roughly $18\%$ relative to the true value of $1$, and its RMSE is $0.221$, roughly $22\%$.
+For $\widehat{\beta}_{J}$, the bias is $8\%$ and the RMSE is $11\%$.
+These are very close to the results of Fernández-Val and Weidner (2016).
+As for the corrected estimators, we find that the bootstrap estimators $\widehat{\beta}_{B}$ and $\widehat{\rho}_{B}$ are generally the best, outperforming the jackknife estimators $\widehat{\beta}_{J}$ and $\widehat{\rho}_{J}$ and our corrected estimators $\widehat{\beta}_{L}^{\left(\tau\right)}$ and $\widehat{\rho}_{L}^{\left(\tau\right)}$.
+The biases of $\widehat{\beta}_{B}$ and $\widehat{\rho}_{B}$ are also comparable to the parameter-based procedure of Fernández-Val and Weidner (2016).
+Our procedure outperforms the jackknife in estimating $\beta_{0}$ (but not $\rho_{0}$) when the sample size is small.
+However, the differences between all estimators (except the MLE) become small as the sample size increases.
+While all being analytical corrections, our corrected estimator performs slightly worse than the parameter-based corrected estimator of Fernández-Val and Weidner (2016).
+This can be anticipated, because our procedure produces the bias correction effect only indirectly via the likelihood, whereas the parameter-based approach applies directly to the estimator.
+However, we would like to point out that the procedure of Fernández-Val and Weidner (2016) does not apply to the DN-HP model without modifications.
+Table F.16: Comparisons of Uncorrected and Corrected Two-Way Fixed-Effect Probit Estimates
+$(N,T)$ $(56,14)$ $(56,28)$ $(56,56)$
+Bias RMSE Bias RMSE Bias RMSE
+$\widehat{\beta}$ $0.182$ $0.221$ $0.093$ $0.120$ $0.062$ $0.078$
+$\widehat{\beta}_{L}^{(1)}$ $0.063$ $0.156$ $0.020$ $0.073$ $0.012$ $0.047$
+$\widehat{\beta}_{L}^{(2)}$ $0.076$ $0.159$ $0.022$ $0.075$ $0.012$ $0.047$
+$\widehat{\beta}_{J}$ 0.083-0.083 $0.151$ 0.028-0.028 $0.081$ 0.007-0.007 $0.047$
+$\widehat{\beta}_{B}$ 0.045-0.045 $0.105$ 0.017-0.017 $0.069$ 0.002-0.002 $0.045$
+$\widehat{\rho}$ 0.226-0.226 $0.266$ 0.098-0.098 $0.134$ 0.047-0.047 $0.081$
+$\widehat{\rho}_{L}^{(1)}$ 0.100-0.100 $0.166$ 0.030-0.030 $0.090$ 0.016-0.016 $0.065$
+$\widehat{\rho}_{L}^{(2)}$ 0.102-0.102 $0.171$ 0.024-0.024 $0.091$ 0.011-0.011 $0.065$
+$\widehat{\rho}_{J}$ $0.046$ $0.161$ $0.015$ $0.095$ $0.000$ $0.067$
+$\widehat{\rho}_{B}$ $0.015$ $0.131$ $0.010$ $0.088$ 0.001-0.001 $0.064$
+[…]
+Table F.17 presents the empirical sizes and powers of the LR test using the uncorrected likelihood (size only), our corrected likelihood, and the bootstrap.
+Briefly speaking, our procedure delivers an empirical size close to the nominal level of $5\%$ and provide a reasonable empirical power, starting from $T=28$.
+On the opposite, the empirical size from the bootstrap is close to $5\%$ for all $T\in\left\{14,28,56\right\}$, but its power is relatively low, especially near the true hypothesis.
+Table F.17: Empirical Size and Power from LR Test for Two-Way Fixed-Effect Probit Estimates
+(a) Empirical Size
+Size
+$(N,T)$ Uncorrected Analytical Bootstrap
+$(56,14)$ $53$ $17$ 44
+$(56,28)$ $30$ 88 33
+$(56,56)$ $25$ 77 55
+(b) Empirical Power
+Power at $\delta$ (Analytical) Power at $\delta$ (Bootstrap)
+$(N,T)$ 0.2-0.2 0.1-0.1 $0.1$ $0.2$ 0.2-0.2 0.1-0.1 $0.1$ $0.2$
+$(56,14)$ $79$ $37$ $33$ $70$ $35$ $10$ 66 1919
+$(56,28)$ $97$ $46$ $45$ $95$ $87$ $25$ $10$ $56$
+$(56,56)$ $100$ $79$ $72$ $100$ $100$ $70$ $27$ $94$
+[…]
+G Additional Empirical Results
+G.1 Sample
+The data set is constructed from waves 20–30 of PSID, covering the survey years 1987 to 1997.
 We selected this time span because 1997 was the last year PSID conducted annual surveys, switching to a biennial schedule in 1999.
-Each survey wave includes questions about employment status in the prior calendar year, such as, ‘About how many weeks did (he/she) work on that job last year?’ and ‘During the weeks that (he/she) worked, about how many hours did (he/she) usually work per week?’ We use responses to these questions to construct individual labor force participation (LFP) status for the prior year.
+Each survey wave includes questions about employment status in the prior calendar year, such as, “About how many weeks did (he/she) work on that job last year?” and “During the weeks that (he/she) worked, about how many hours did (he/she) usually work per week?” We use responses to these questions to construct individual labor force participation (LFP) status for the prior year.
 Thus, wave 30 (1997) provides LFP data for 1996, wave 29 (1996) for 1995, and so on.
 In our final sample, labor force participation status is observed from 1987 to 1996, with lagged participation status observed from 1986 to 1995.
-We also restrict our sample to individuals whose LFP status changed at least once during the sample period.
-Individuals whose labor supply status never changed during the sample period are not informative for our analysis, because the dependent variable could be perfectly predicted by setting the individual fixed effects in the intercept of our models to plus or negative infinity.
-The final sample consists of a panel of $N=90$ individuals over $T=10$ years.
-Table C.1 presents the summary statistics.
-Figure C.1 shows the average percentage of individuals with given current and past LFP status.
-Figure C.2 plots the LFP transition rates by year.
-Table C.1: Descriptive Statistics
+We restrict the data to the informative sample of individuals whose current and lagged LFP status each varied at least once during the sample period.
+This ensures within-person variation in both variables, which is necessary for identification of the dynamic model.
+The final sample contains $N=86$ single mothers observed for $T=10$ consecutive years.
+Table G.1 reports the summary statistics.
+Figure G.1 displays the joint distribution of current and lagged LFP status, and Figure G.2 plots the year-specific transition rates between employment and non-employment.
+Table G.1: Descriptive Statistics
 Variable Mean Std. Dev. Min Max
-Labor Force Participation 0.482 0.500 0.000 1.000
-Lagged Participation 0.458 0.498 0.000 1.000
-Number of Children 2.24 1.098 1.000 6.000
-  • Notes: Summary statistics are computed based on a sample of 90 individuals observed from 1987 to 1996. All variables have 900 observations. Data source: PSID 1987 - 1997.
-Figure C.1: Labor Force Participation Contingency Table. Each cell contains the percentage of individuals with given current and past labor force participation status.
-Figure C.2: Labor Force Participation Status Transition Rates By Year
-C.2 Extra Empirical Results
-Table C.2: Estimated Heterogeneous Effects on Labor Force Participation
+Labor Force Participation 0.491 0.500 0.000 1.000
+Lagged Participation 0.467 0.499 0.000 1.000
+Number of Children 2.25 1.104 1.000 6.000
+  • Notes: Summary statistics are computed based on a sample of 86 individuals observed from 1987 to 1996. All variables have 860 observations. Data source: PSID 1987–1997.
+Figure G.1: Labor Force Participation Contingency Table
+[…]
+Notes: Each cell contains the percentage of individuals with given current and past labor force participation status.
+Figure G.2: Labor Force Participation Status Transition Rates By Year
+G.2 Extra Empirical Results
+Table G.2: Estimated Heterogeneous Effects on Labor Force Participation
 10th 25th 75th 90th
 Mean Median Std. Dev. Percentile Percentile Percentile Percentile
-$\alpha_{i}$ $0$0.975 6.203 -6.909 -2.443 4.292 6.774
-$\gamma_{t}$ 0.000 0.146 0.773 -1.151 -0.466 0.638 0.979
-$\beta_{it}$ 0.126 -0.504 6.250 -6.835 -2.628 4.050 6.982
+$\alpha_{i}$ 0.000 0.012 2.329 -2.477 -1.858 1.883 3.521
+$\gamma_{t}$ 0.000 -0.125 0.450 -0.435 -0.361 0.248 0.692
+$\beta_{it}$ 0.436 0.388 2.372 -2.186 -1.487 2.216 3.768
+Table G.3: Estimated Heterogeneous State Dependence in Labor Force Participation
+10th 25th 75th 90th
+Mean Median Std. Dev. Percentile Percentile Percentile Percentile
+$\zeta_{i}$ 0.000 -0.021 2.359 -2.831 -1.766 1.730 3.353
+$\eta_{t}$ 0.000 -0.149 0.812 -0.926 -0.708 0.245 1.358
+$\rho_{it}$ 0.253 0.083 2.495 -2.812 -1.635 2.090 3.725
+Figure G.3: Estimated Heterogeneous State Dependence in Labor Force Participation
+(a) Distribution of Estimated State-Dependence Parameters
+(b) Temporal Evolution of Estimated State Dependence
+[…]
+Notes: Panel (a) shows the distribution of the estimated individual-time state-dependence parameters $\rho_{it}$ from the dynamic heterogeneous-slope model.
+Panel (b) plots the median (solid), interquartile range (25th–75th, dashed), and decile range (10th–90th, dotted) of $\rho_{it}$ over time.
+H Average Partial Effects
+We discuss the average partial effects (APEs) in this section. Denote
+$\Delta\left(\theta,\phi\right):=\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}\Delta_{i,t}\left(\theta,\alpha_{i},\gamma_{t}\right)$
+where $\Delta_{i,t}\left(\theta,\alpha_{i},\gamma_{t}\right)$ is, for simplicity, a scalar partial effect function of interests.
+For the probit model in Section 4, $\Delta_{i,t}\left(\theta,\alpha_{i},\gamma_{t}\right)$ could be, e.g.,
+$\Delta_{i,t}\left(\theta,\alpha_{i},\gamma_{t}\right):=F\left(\rho+Z_{it}\left(\beta+\alpha_{1,i}+\gamma_{1,t}\right)+\alpha_{2,i}+\gamma_{2,t}\right)-F\left(Z_{it}\left(\beta+\alpha_{1,i}+\gamma_{1,t}\right)+\alpha_{2,i}+\gamma_{2,t}\right),$
+which is the partial effect of the binary variable $Y_{it-1}$, where $\theta=\left(\rho,\beta\right)$ and $F\left(\cdot\right)$ is the standard normal cumulative distribution function.
+Let $\overline{\mathbb{E}}$ denote the unconditional expectation w.r.t.
+the joint density of $\left(Y_{it},X_{it}^{\prime},\phi^{0\prime}\right)^{\prime}$, $\Delta:=\Delta(\theta_{0},\phi^{0})$, and $\overline{\Delta}:=\overline{\mathbb{E}}\Delta$.
+For nonlinear models with two-way fixed effects in the intercept only, Fernández-Val and Weidner (2016) show that the APE estimator $)\widehat{\Delta}:=\Delta(\widehat{\theta},\widehat{\phi}(\widehat{\theta}$)) is asymptotically unbiased.
+We briefly re-visit this result here and justify that it also applies to the DN-HP models.
+Consider the decomposition
+$\widehat{\Delta}-\overline{\Delta}=(\widehat{\Delta}-\Delta)+(\Delta-\overline{\Delta}).$
+Here $\widehat{\Delta}-\Delta$ is due to the estimation error of the parameters.
+As demonstrated by Fernández-Val and Weidner (2016), $\widehat{\Delta}$ is affected by the IPP in that it possesses a bias of order $1/\sqrt{NT}$ away from $\Delta$.
+On the other hand, $\Delta-\overline{\Delta}$ is due to the estimation of the population mean $\overline{\Delta}$ by the sample mean $\Delta$.
+The component $\Delta$ does not suffer from the IPP.
+However, its convergence rate depends on the data-generating process and could potentially be slower than $1/\sqrt{NT}$.
+To see this, consider $\left\{\alpha_{i}^{0},\gamma_{t}^{0}\right\}$ being i.i.d.
+across both $i$ and $t$.
+Calculating the variance of $(\Delta-\overline{\Delta})$, we have
+$\overline{\mathbb{E}}(\Delta-\overline{\Delta})^{2}=\frac{1}{N^{2}T^{2}}\sum_{i=1}^{N}\sum_{j=1}^{N}\sum_{t=1}^{T}\sum_{s=1}^{T}\overline{\mathbb{E}}\left[\widetilde{\Delta}_{i,t}(\theta_{0},\alpha_{i}^{0},\gamma_{t}^{0})\widetilde{\Delta}_{j,s}(\theta_{0},\alpha_{j}^{0},\gamma_{s}^{0})\right],$
+where $\widetilde{\Delta}_{i,t}\left(\theta,\alpha_{i},\gamma_{t}\right):=\Delta_{i,t}\left(\theta,\alpha_{i},\gamma_{t}\right)-\overline{\mathbb{E}}\Delta_{i,t}\left(\theta,\alpha_{i},\gamma_{t}\right)$.
+Here, $\widetilde{\Delta}_{i,t}(\theta_{0},\alpha_{i}^{0},\gamma_{t}^{0})$ and $\widetilde{\Delta}_{j,t}(\theta_{0},\alpha_{j}^{0},\gamma_{t}^{0})$, for $i\neq j$, depend on the same $\gamma_{t}^{0}$; while $\widetilde{\Delta}_{i,t}(\theta_{0},\alpha_{i}^{0},\gamma_{t}^{0})$ and $\widetilde{\Delta}_{i,s}(\theta_{0},\alpha_{i}^{0},\gamma_{s}^{0})$, for $t\neq s$, depend on the same $\alpha_{i}^{0}$.
+Since $\alpha_{i}^{0}$ and $\gamma_{t}^{0}$ are held as random variables in the APE context, $\widetilde{\Delta}_{i,t}(\theta_{0},\alpha_{i}^{0},\gamma_{t}^{0})$ exhibits dependencies across both $i$ or $t$.
+As a consequence, the rate of the variance $\overline{\mathbb{E}}(\Delta-\overline{\Delta})^{2}$ may be slower than $1/NT$.
+In other words, the bias $\widehat{\Delta}-\Delta$ is negligible relative to the variance of $\Delta-\overline{\Delta}$, implying $\widehat{\Delta}$ being asymptotically unbiased towards $\overline{\Delta}$.
+Of course, there still might be some finite-sample bias.
+For the two-way DN-HP model here, applying the same expansion as in Section 2 to $\Delta\left(\theta,\phi\right)$ would establish that the bias $\widehat{\Delta}-\Delta$ is also of order $1/\sqrt{NT}$.
+Therefore, the same arguments above still apply.
+Next, we conduct simulation studies to demonstrate that estimating the APEs using $\widehat{\theta}_{L}$ already leads to a considerable reduction of its finite-sample bias.
+Under the same settings in Section 4, Table H.1 compares the biases and RMSEs of $)\widehat{\Delta}^{Z}:=\Delta^{Z}(\widehat{\theta},\widehat{\phi}(\widehat{\theta}$)) v.s.
+$\widehat{\Delta}_{L}^{Z}:=\Delta^{Z}(\widehat{\theta}_{L},\widehat{\phi}(\widehat{\theta}_{L}))$ and $)\widehat{\Delta}^{Y}:=\Delta^{Y}(\widehat{\theta},\widehat{\phi}(\widehat{\theta}$)) v.s.
+$\widehat{\Delta}_{L}^{Y}:=\Delta^{Y}(\widehat{\theta}_{L},\widehat{\phi}(\widehat{\theta}_{L}))$, where
+$\Delta^{Z}(\theta,\phi)$ $:=\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}\Delta_{i,t}^{Z}\left(\theta,\alpha_{i},\gamma_{t}\right),\qquad\Delta^{Y}(\theta,\phi):=\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}\Delta_{i,t}^{Y}\left(\theta,\alpha_{i},\gamma_{t}\right),$
+$\Delta_{i,t}^{Z}\left(\theta,\alpha_{i},\gamma_{t}\right)$ $:=f\left(\rho Y_{it-1}+Z_{it}\left(\beta+\alpha_{1,i}+\gamma_{1,t}\right)+\alpha_{2,i}+\gamma_{2,t}\right)\left(\beta+\alpha_{1,i}+\gamma_{1,t}\right),$
+$\Delta_{i,t}^{Y}\left(\theta,\alpha_{i},\gamma_{t}\right)$ $:=F\left(\rho+Z_{it}\left(\beta+\alpha_{1,i}+\gamma_{1,t}\right)+\alpha_{2,i}+\gamma_{2,t}\right)-F\left(Z_{it}\left(\beta+\alpha_{1,i}+\gamma_{1,t}\right)+\alpha_{2,i}+\gamma_{2,t}\right),$
+$F\left(\cdot\right)$ is the cumulative distribution function of a standard normal (for the probit model) or a standard logistic (for the logit model) distribution, and $f\left(\cdot\right)$ is the probability density function corresponding to $F\left(\cdot\right)$.
+The biases and RMSEs are calculated relative to the Monte-Carlo means of the infeasible APEs $\Delta^{Z}:=\Delta^{Z}(\theta_{0},\phi^{0})$ and $\Delta^{Y}:=\Delta^{Y}(\theta_{0},\phi^{0})$, respectively, and the biases are reported in percentages relative to the Monte-Carlo means of $\Delta^{Z}$ and $\Delta^{Y}$ (for better comparison).
+Finally, we only report the APE estimators calculated using $\widehat{\beta}_{L}^{(1)}$ and $\widehat{\rho}_{L}^{(1)}$ (omitting those using $\widehat{\beta}_{L}^{(2)}$ and $\widehat{\rho}_{L}^{(2)}$ because they are similar).
+For the static logit model with $(N,T)=(30,30)$, we find that the bias of the plug-in estimator $\widehat{\Delta}_{L}^{Z}$ is only $2\%$, while the uncorrected APE estimator $\widehat{\Delta}^{Z}$ shows a $16\%$ bias.
+The RMSE of $\widehat{\Delta}_{L}^{Z}$ is also smaller than that of $\widehat{\Delta}^{Z}$.
+As the sample size increases, the RMSEs of both estimators decrease, with $\widehat{\Delta}_{L}^{Z}$ outperforming $\widehat{\Delta}^{Z}$ constantly.
+The results here suggest that estimating the APEs using $\widehat{\theta}_{L}$ already reduces the finite-sample bias without inflating the RMSE, compared to using $\widehat{\theta}$ directly, even that $\widehat{\Delta}_{L}^{Z}$ and $\widehat{\Delta}_{L}^{Y}$ are intrinsically not bias-corrected estimators.
+Table H.1: Comparisons of Average Partial Effects for Logit and Probit Models, Design 1
+$(N,T)$ $(30,30)$ $(60,60)$ $(90,90)$
+Bias RMSE Bias RMSE Bias RMSE
+Static Logit Model
+$\widehat{\Delta}^{Z}$ $16$ $0.028$ 77 $0.012$ 44 $0.008$
+$\widehat{\Delta}_{L}^{Z}$ 2-2 $0.018$ 3-3 $0.009$ 3-3 $0.006$
+Static Probit Model
+$\widehat{\Delta}^{Z}$ $18$ $0.057$ 77 $0.014$ 44 $0.009$
+$\widehat{\Delta}_{L}^{Z}$ 11 $0.019$ 2-2 $0.008$ 2-2 $0.006$
+Dynamic Logit Model
+$\widehat{\Delta}^{Z}$ $15$ $0.027$ 66 $0.011$ 55 $0.008$
+$\widehat{\Delta}_{L}^{Z}$ 2-2 $0.018$ 3-3 $0.009$ 2-2 $0.006$
+$\widehat{\Delta}^{Y}$ 32-32 $0.049$ 16-16 $0.024$ 10-10 $0.016$
+$\widehat{\Delta}_{L}^{Y}$ 19-19 $0.038$ 9-9 $0.019$ 5-5 $0.012$
+Dynamic Probit Model
+$\widehat{\Delta}^{Z}$ $26$ $0.121$ 88 $0.015$ 55 $0.010$
+$\widehat{\Delta}_{L}^{Z}$ 44 $0.027$ 0 $0.008$ 1-1 $0.005$
+$\widehat{\Delta}^{Y}$ 21-21 $0.048$ 10-10 $0.022$ 6-6 $0.015$
+$\widehat{\Delta}_{L}^{Y}$ 15-15 $0.039$ 6-6 $0.018$ 3-3 $0.012$
+[…]
+Notes: Bias is reported in percentage relative to the Monte-Carlo mean of the infeasible APEs $\Delta^{Z}$ or $\Delta^{Y}$.
+For the dynamic models, $\widehat{\Delta}_{L}^{Z}$ or $\widehat{\Delta}_{L}^{Y}$ are calculated using $\widehat{\beta}_{L}^{(1)}$ and $\widehat{\rho}_{L}^{(1)}$.
+I The Failure of the Jackknife
+In our simulation studies, we find that the jackknife does not perform well in reducing the bias of structural parameters associated with two-way heterogeneities for small $N,T$.
+This could be due to the incompatibility of the jackknife with the normalization for the identification.
+Note that the normalization is crucial for the identification under the DN-HP model.
+In this section, we discuss the possible cause of the failure of the jackknife.
+Consider even $N,T$ and a simple probit model
+$Y_{it}=\mathds{1}\left\{\left(\beta+\alpha_{i}+\gamma_{t}\right)X_{it}+\varepsilon_{it}>0\right\},$
+where $X_{it}$ is a scalar regressor and $\left\{\beta,\alpha_{i},\gamma_{t}\right\}$ are all scalar parameters.
+For simplicity, we assume that there is no intercept, which does not affect our discussion here.
+Recall that we set $N^{-1}\sum_{i=1}^{N}\alpha_{i}^{0}=T^{-1}\sum_{t=1}^{T}\gamma_{t}^{0}=0$ in the data-generating process and normalize $N^{-1}\sum_{i=1}^{N}\alpha_{i}=T^{-1}\sum_{t=1}^{T}\gamma_{t}=0$ during the estimation.
+The normalization causes no problem on the full data set.
+However, this is not the case for half panels.
+To see the problem, define
+$\mathcal{P}_{N,T/2}^{\left(1\right)}$ $:=\left\{\left(i,t\right):i=1,\ldots,N;t=1,\ldots,T/2\right\},$
+$\mathcal{P}_{N,T/2}^{\left(2\right)}$ $:=\left\{\left(i,t\right):i=1,\ldots,N;t=T/2+1,\ldots,T\right\},$
+$\mathcal{P}_{N/2,T}^{\left(1\right)}$ $:=\left\{\left(i,t\right):i=1,\ldots,N/2;t=1,\ldots,T\right\},$
+$\mathcal{P}_{N/2,T}^{\left(2\right)}$ $:=\left\{\left(i,t\right):i=N/2+1,\ldots,N;t=1,\ldots,T\right\},$
+where $\mathcal{P}_{N,T/2}^{\left(1\right)}$ and $\mathcal{P}_{N,T/2}^{\left(2\right)}$ are the two half panels obtained by splitting along the time dimension, and $\mathcal{P}_{N/2,T}^{\left(1\right)}$ and $\mathcal{P}_{N/2,T}^{\left(2\right)}$ are constructed by splitting along the individual dimension.
+Focus on the two half panels $\mathcal{P}_{N,T/2}^{\left(1\right)}$ and $\mathcal{P}_{N,T/2}^{\left(2\right)}$, and define $g_{T/2}^{\left(1\right)}:=\left(T/2\right)^{-1}\sum_{t=1}^{T/2}\gamma_{t}^{0}$ and $g_{T/2}^{\left(2\right)}:=\left(T/2\right)^{-1}\sum_{T/2+1}^{T}\gamma_{t}^{0}$.
+Here $g_{T/2}^{\left(1\right)}$ is the summation of $\gamma_{t}^{0}$ over the time periods in $\mathcal{P}_{N,T/2}^{\left(1\right)}$ and $g_{T/2}^{\left(2\right)}$ similarly over $\mathcal{P}_{N,T/2}^{\left(2\right)}$.
+Note that $g_{T/2}^{\left(1\right)}+g_{T/2}^{\left(2\right)}=0$ but they are generally nonzero themselves.
+When the model is estimated on $\mathcal{P}_{N,T/2}^{\left(1\right)}$, the normalization $\sum_{t=1}^{T/2}\gamma_{t}=0$ is imposed.
+However, this normalization does not agree with $g_{T/2}^{\left(1\right)}\neq 0$.
+Consequently, the effect from $g_{T/2}^{\left(1\right)}$ is absorbed into the common parameter.
+In other words, the MLE computed on $\mathcal{P}_{N,T/2}^{\left(1\right)}$, say $\widehat{\beta}_{N,T/2}^{\left(1\right)}$, is essentially estimating $\beta_{N,T/2}^{\left(1\right)}=\beta_{0}+g_{T/2}^{\left(1\right)}$.
+The same issue happens on $\mathcal{P}_{N,T/2}^{\left(2\right)}$, causing the MLE obtained on $\mathcal{P}_{N,T/2}^{\left(2\right)}$, say $\widehat{\beta}_{N,T/2}^{\left(2\right)}$, to be estimating $\beta_{N,T/2}^{\left(2\right)}=\beta_{0}+g_{T/2}^{\left(2\right)}$.
+Note that $\beta_{N,T/2}^{\left(1\right)}\neq\beta_{N,T/2}^{\left(2\right)}$ generally but $(\beta_{N,T/2}^{\left(1\right)}+\beta_{N,T/2}^{\left(2\right)})/2=\beta_{0}$.
+The same happens on $\mathcal{P}_{N/2,T}^{\left(1\right)}$ and $\mathcal{P}_{N/2,T}^{\left(2\right)}$, causing the corresponding MLEs, say $\widehat{\beta}_{N/2,T}^{\left(1\right)}$ and $\widehat{\beta}_{N/2,T}^{\left(2\right)}$, to be estimating some $\beta_{N/2,T}^{\left(1\right)}$ and $\beta_{N/2,T}^{\left(2\right)}$ generally different from $\beta_{0}$.
+Now, suppose that the MLE admits the expansion
+$\widehat{\beta}-\beta_{0}=\frac{b_{\alpha}\left(\beta_{0}\right)}{T}+\frac{b_{\gamma}\left(\beta_{0}\right)}{N}+o_{\mathbb{P}}\left(N^{-1}\vee T^{-1}\right)$ (I.1)
+for some bias terms $b_{\alpha}\left(\beta\right)$ and $b_{\gamma}\left(\beta\right)$, as in Fernández-Val and Weidner (2016).
+On half panels, because the estimated targets are generally different, we would have
+$\widehat{\beta}_{N,T/2}^{\left(1\right)}-\beta_{N,T/2}^{\left(1\right)}$ $=\frac{b_{\alpha}(\beta_{N,T/2}^{\left(1\right)})}{T/2}+\frac{b_{\gamma}(\beta_{N,T/2}^{\left(1\right)})}{N}+o_{\mathbb{P}}\left(N^{-1}\vee T^{-1}\right),$
+$\widehat{\beta}_{N,T/2}^{\left(2\right)}-\beta_{N,T/2}^{\left(2\right)}$ $=\frac{b_{\alpha}(\beta_{N,T/2}^{\left(2\right)})}{T/2}+\frac{b_{\gamma}(\beta_{N,T/2}^{\left(2\right)})}{N}+o_{\mathbb{P}}\left(N^{-1}\vee T^{-1}\right),$
+$\widehat{\beta}_{N/2,T}^{\left(1\right)}-\beta_{N/2,T}^{\left(1\right)}$ $=\frac{b_{\alpha}(\beta_{N/2,T}^{\left(1\right)})}{T}+\frac{b_{\gamma}(\beta_{N/2,T}^{\left(1\right)})}{N/2}+o_{\mathbb{P}}\left(N^{-1}\vee T^{-1}\right),$
+$\widehat{\beta}_{N/2,T}^{\left(2\right)}-\beta_{N/2,T}^{\left(2\right)}$ $=\frac{b_{\alpha}(\beta_{N/2,T}^{\left(2\right)})}{T}+\frac{b_{\gamma}(\beta_{N/2,T}^{\left(2\right)})}{N/2}+o_{\mathbb{P}}\left(N^{-1}\vee T^{-1}\right).$
+Since $\beta_{0}$, $\beta_{N,T/2}^{\left(1\right)}$, $\beta_{N,T/2}^{\left(2\right)}$, $\beta_{N/2,T}^{\left(1\right)}$, and $\beta_{N/2,T}^{\left(2\right)}$ generally differ from each other, the five $b_{\alpha}\left(\cdot\right)$ terms would also be different.
+The same happens to $b_{\gamma}\left(\cdot\right)$.
+Next, define
+$\widehat{\beta}_{N,T/2}$ $:=\frac{1}{2}\left(\widehat{\beta}_{N,T/2}^{\left(1\right)}+\widehat{\beta}_{N,T/2}^{\left(2\right)}\right)$
+$=\beta_{0}+\frac{b_{\alpha}(\beta_{N,T/2}^{\left(1\right)})+b_{\alpha}(\beta_{N,T/2}^{\left(2\right)})}{T}+\frac{b_{\gamma}(\beta_{N,T/2}^{\left(1\right)})+b_{\gamma}(\beta_{N,T/2}^{\left(2\right)})}{2N}+o_{\mathbb{P}}\left(N^{-1}\vee T^{-1}\right),$
+$\widehat{\beta}_{N/2,T}$ $:=\frac{1}{2}\left(\widehat{\beta}_{N/2,T}^{\left(1\right)}+\widehat{\beta}_{N/2,T}^{\left(2\right)}\right)$
+$=\beta_{0}+\frac{b_{\alpha}(\beta_{N/2,T}^{\left(1\right)})+b_{\alpha}(\beta_{N/2,T}^{\left(2\right)})}{2T}+\frac{b_{\gamma}(\beta_{N/2,T}^{\left(1\right)})+b_{\gamma}(\beta_{N/2,T}^{\left(2\right)})}{N}+o_{\mathbb{P}}\left(N^{-1}\vee T^{-1}\right).$
+Noting (I.1) we have
+$\widehat{\beta}_{N,T/2}-\widehat{\beta}$ $=\frac{b_{\alpha}(\beta_{N,T/2}^{\left(1\right)})+b_{\alpha}(\beta_{N,T/2}^{\left(2\right)})-b_{\alpha}\left(\beta_{0}\right)}{T}+\frac{b_{\gamma}(\beta_{N,T/2}^{\left(1\right)})+b_{\gamma}(\beta_{N,T/2}^{\left(2\right)})-2b_{\gamma}\left(\beta_{0}\right)}{2N}$
+$\quad\quad+o_{\mathbb{P}}\left(N^{-1}\vee T^{-1}\right),$
+$\widehat{\beta}_{N/2,T}-\widehat{\beta}$ $=\frac{b_{\alpha}(\beta_{N/2,T}^{\left(1\right)})+b_{\alpha}(\beta_{N/2,T}^{\left(2\right)})-2b_{\alpha}\left(\beta_{0}\right)}{2T}+\frac{b_{\gamma}(\beta_{N/2,T}^{\left(1\right)})+b_{\gamma}(\beta_{N/2,T}^{\left(2\right)})-b_{\gamma}\left(\beta_{0}\right)}{N}$
+$\quad\quad+o_{\mathbb{P}}\left(N^{-1}\vee T^{-1}\right).$
+Given the jackknife estimator $\widehat{\beta}_{J}:=3\widehat{\beta}-\widehat{\beta}_{N,T/2}-\widehat{\beta}_{N/2,T}$, we have
+[…]
+$=\frac{6b_{\alpha}(\beta_{0})-2b_{\alpha}(\beta_{N,T/2}^{\left(1\right)})-2b_{\alpha}(\beta_{N,T/2}^{\left(2\right)})-b_{\alpha}(\beta_{N/2,T}^{\left(1\right)})-b_{\alpha}(\beta_{N/2,T}^{\left(2\right)})}{2T}$
+$\quad\quad+\frac{6b_{\gamma}(\beta_{0})-b_{\gamma}(\beta_{N,T/2}^{\left(1\right)})-b_{\gamma}(\beta_{N,T/2}^{\left(2\right)})-2b_{\gamma}(\beta_{N/2,T}^{\left(1\right)})-2b_{\gamma}(\beta_{N/2,T}^{\left(2\right)})}{2N}$
+$\quad\quad+o_{\mathbb{P}}\left(N^{-1}\vee T^{-1}\right).$ (I.2)
+If $\beta_{0}$, $\beta_{N,T/2}^{\left(1\right)}$, $\beta_{N,T/2}^{\left(2\right)}$, $\beta_{N/2,T}^{\left(1\right)}$, and $\beta_{N/2,T}^{\left(2\right)}$ were the same, the terms in the numerator would cancel out; and $\widehat{\beta}_{J}-\beta^{0}$ would be $o_{\mathbb{P}}\left(N^{-1}\vee T^{-1}\right)$.
+For two-way DN-HP models, $\beta_{N,T/2}^{\left(1\right)}$, $\beta_{N,T/2}^{\left(2\right)}$, $\beta_{N/2,T}^{\left(1\right)}$, and $\beta_{N/2,T}^{\left(2\right)}$ only approach $\beta_{0}$, as $N,T$ increase, because $g_{T/2}^{\left(1\right)}\$and $g_{T/2}^{\left(2\right)}$ (and similar terms from subpanels $\mathcal{P}_{N/2,T}^{\left(1\right)}$ and $\mathcal{P}_{N/2,T}^{\left(2\right)}$) approach zero.
+When $N$ and $T$ are small, they remain different.
+This causes $\widehat{\beta}_{J}$ to perform less well under small $N,T$.